A Complete Bibliography of Publications in Contemporary Physics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

01 April 2025 Version 1.37

Title word cross-reference

$100 [4158]. $11.95 [4073]. $120.00 [4337, 4201]. $125.00 [4234, 4216]. $135 [4206, 4079]. $145 [4061]. $148.00 [4096]. $156.00 [4252]. $156.35 [4242]. $159 [4241, 4260]. $159.00 [4140, 4173, 4150, 4144, 4142]. $16.95/£ [4291]. $168.00 [3996]. $169 [4334]. $175 [4172]. $18.95/£ [4327]. $190 [4241, 4189, 4260, 4136]. $190.00 [4251, 4173, 4249, 4169, 4170, 4174]. $190.000 [4290]. $195.00 [4250]. 2 [2696, 3666]. $22.50 [555]. $24.99 [4105]. $26.95 [4292]. $27.95/£ [4327]. $28.00 [4132, 4024, 4200]. $29.95 [4316]. 3 [2779, 3328, 3781, 3145, 2223]. 3 + 1 [3185]. $32.00 [4143]. $32.00/£ [4356]. $32.22 [4317]. $32.99 [4328]. $33.00/£ [4355]. $34.99 [4322]. $38/£ [4114]. 4 [3432, 4413]. $40 [4303, 4233]. $40.00 [4244]. $45.00 [4095]. $48.00 [4070]. $50 [4233]. $50.00 [4244]. $56.00 [4354]. $58.00 [4324]. $64 [4283]. $68.00 [4010, 4086]. $69.95 [4186]. $69.99 [4146]. $7.00 [662]. $78.00 [4098, 4067]. $78/£ [4114]. $79.99 [4254, 4107]. $84.99 [4130, 4326]. $88.00 [4029]. $89.91 [4306]. $90.00 [4201]. $95 [4131, 4341]. $95.00 [4139]. $98 [4108]. $98.00 [4068]. ^{28} [2892]. ^{3} [730, 1107]. ^{4} [1737].

(R) [3438]. _{2} [1139, 4180]. _{3} [4180]. _{3}^{+} [1642]. _{4} [4180]. _{c} [1580]. a [3726]. α [2177, 2165]. β [141]. PT [2138]. E = mc^{2} [4067]. Λ [4076, 3222]. M [3420].

-D [2223]. -dimensional [3781]. -history [3420]. -particle [141]. -quartz [2177]. -symmetric [2138]. -theorem [3726].

//doi.org/10.1063/9780735423350 [4206]. //store.ioppublishing.org/page/detail/A [4201]. //store.ioppublishing.org/page/detail/A-Handbook-of-Mathematical-Methods-and-ProblemSolving-Tools-for-Introductory-Physics-Second-Edition/ [4201]. //www.cambridge.org/gb/academic/subjects/physics/particle [4211]. //www.cambridge.org/gb/academic/subjects/physics/particlephysics-and-nuclear-physics/ideas-particle-physics-4th-edition [4211]. //www.elsevier.com/books/space [4186]. //www.elsevier.com/books/space-science-and-public-engagement/kaminski/978-0-12-817390-9 [4186]. /978 [4156]. /978-0-19-885188-2 [4156].

0 [4225, 4284, 4038, 4005, 4285, 4148, 4130, 4222, 3994, 3957, 4283, 1562, 3956, 1844]. 0-19-506908-0 [1562]. 0-19-850551-5 [2041]. 0-19-851097-7 [2041]. 0-19-856479-1 [2057]. 0-226-79845-3 [2191]. 0-7503-3461-4 [4170]. 0-7503-3801-6 [4234]. 0-7923-3818-9 [1844]. 0/ISBN [4154].

1 [4137, 4271, 4221, 21, 4021, 4044, 4008, 4015, 4283, 4205, 4131, 4129, 4276, 2701, 11, 4039, 4033, 3187, 4306]. 1-032-2029-0 [4283]. 10.00 [834]. 100 [2312, 3171]. 100th [3461]. 105s [291]. 114 [112]. 11th [278]. 125-£ [4183]. 12th [277]. 13th [277]. 14.60 [834]. 15pp [4081]. 16th [276]. 1920s [3363]. 1925 [3048]. 1926 [2585]. 1931 [2778]. 1940 [3857]. 1944 [1136]. 1945 [1092]. 1959 [50]. 1960 [75, 51, 137, 89, 76]. 1961 [193, 156, 144]. 1962 [158, 180, 242, 214]. 1963 [221, 261, 278, 277, 276]. 1964 [366, 338, 308, 316]. 1975 [834]. 1981 [1131]. 19th [365]. 1st [2945, 3947, 4061, 4190, 4323, 4283].

2 [4211, 4273, 4272, 4030, 32, 4138, 3999, 4048, 4042, 4003, 4150, 4223, 34, 4023, 4136, 2765, 3200, 3172]. 2003 [3020]. 2008 [3063]. 2014 [3551, 3633, 3836]. 2018 [3877]. 2024 [3751]. 20th [365]. 21 [2926]. 21st [3945, 3989, 4347, 3753, 4042, 4186, 2449, 4234]. 25th [3133]. 2nd [4078, 3318, 2788, 2786, 2667, 2936, 3796, 2927, 4140, 2715, 3124, 4056, 4273, 2865, 3116, 2857, 3723, 3609, 4005, 4123, 2320, 2509, 2866, 2995, 3900, 3023, 4183, 2621, 2907, 2639, 2533, 3845, 2845, 3036, 2901, 2806, 2684, 3626, 3280, 3561, 3962, 4081, 4204, 4130, 4259, 4269, 4281, 180, 221, 3323, 3920, 3999, 3377, 3317, 3400, 3553, 3427, 3436, 3668, 3708, 3643, 3783, 3957, 3910, 4090, 4345, 4187, 4165, 3510, 2523, 2703, 3401, 4157, 4292, 2889, 3832, 2299, 3208, 3768, 3770, 3271, 3430, 3712, 2796, 2873, 2890, 2869, 2971, 4282, 4131, 3703, 3842, 4122, 3209, 2997, 3827, 3917]. 2nd [3949, 4080, 4234, 2436, 2728, 3109, 3347, 4144, 4351, 3731, 2976, 2511, 3645, 3996, 3995, 4077, 4167, 3138, 2557, 2390, 2701, 4253, 3672, 3285, 3528, 2579, 4216, 2422, 2982, 2370, 2665, 2666, 2453, 2388, 2475, 2470, 2646, 2663, 2802, 2827, 3134, 3443, 3621, 3667, 3818, 3850, 4004, 3916, 4233, 4175, 2919, 2586].

3 [4286, 4268, 4191, 2082, 4029, 4098, 4196, 4141, 4047, 4091, 3078, 4099, 4224, 4019, 4201, 4208, 4199, 4142]. 307pp [4156]. 331 [4028]. 3D [4283]. 3rd [3766, 2462, 2868, 4293, 2955, 2939, 3826, 4148, 3809, 2954, 2996, 3437, 3588, 3631, 4091, 4294, 2916, 2738, 2517, 3231, 2855, 3922, 4053, 2676, 4071, 4066, 2314, 3890, 2638, 2913, 2885, 3153, 3021, 3411, 2319].

4 [4275, 4000, 4210, 4035, 4025, 4001, 3459, 4187, 437, 4282, 3996, 4216, 4145, 2756]. 40 [3313]. 419-9 [4177]. 44th [38, 36, 37, 39]. 45th [97]. 461pp [4081]. 46th [158]. 48th [308]. 49 [2329]. 496pp [4175]. 4th [3986, 2463, 4211, 2828, 2792, 3016, 4040, 3429, 3973, 3701, 3979, 3215, 2882, 193].

5 [2329, 4198, 4043, 4037, 4200, 4144, 4219, 4094]. 50 [2755]. 50th [2238, 4346]. 5s.Od [554]. 5th [4133, 3395, 2829, 4275, 180, 221, 4189, 2952, 2736].

6 [4010, 2772, 4135, 4143, 3102, 4190, 4031, 4040, 4093, 4218, 4022, 4036, 4265, 4147, 4004]. 60th [4092]. 6th [2335, 2373].

7 [4018, 4182, 4184, 4027, 4274, 4002, 4041, 3993, 3998, 3423, 3995, 4188, 4092, 4215]. 7th [3721, 4092].

8 [4132, 2893, 4202, 4140, 4032, 3945, 4045, 4100, 4183, 4026, 4146, 4097, 4007, 4151, 4020, 4096, 4266, 4220, 4207, 4267]. 80 [3305, 2822]. 80th [648, 3681].

9 [4016, 4217, 4139, 4206, 3425, 4006, 4046, 4024, 4186, 2644, 4017, 4197, 4189, 4095, 4034, 3997, 1844, 4185]. 9.80. [869]. 90EUR [4232]. 90th [3412, 3976]. 978 [4132, 4010, 4016, 4211, 4137, 4225, 4286, 2893, 2772, 4217, 4284, 4275, 4202, 4268, 4135, 4140, 4139, 4038, 4000, 4032, 4210, 4273, 4271, 4206, 4035, 3425, 4005, 3945, 4045, 4018, 4028, 4182, 4184, 4191, 4221, 4027, 4100, 4143, 4183, 4272, 2329, 3102, 4030, 4285, 4138, 4026, 4190, 4148, 4198, 4006, 4021, 4130, 4222, 4274, 4025, 4002, 4001, 4046, 4043, 4041, 3999, 4029, 4044, 4008, 4024, 4098, 4196, 4141, 3459, 4031, 4037, 4047, 3993, 4040, 3994, 4048, 3957, 4042, 4003, 4015, 4091, 4146, 4187, 3998, 4097, 4150, 2644, 3423, 4007, 4017, 3078, 4099, 4093, 4200, 4151, 4205]. 978 [4224, 4019, 4282, 4131, 4197, 4201, 4223, 4218, 4020, 4022, 4036, 4189, 4144, 3996, 3995, 4129, 4208, 4188, 4219, 4276, 4199, 2701, 4095, 4096, 4094, 4265, 4266, 4220, 4092, 4142, 4147, 4216, 4034, 4004, 4039, 3956, 4033, 3997, 4023, 4136, 4207, 4215, 4145, 4185, 4267]. 978-0-12-804651-7 [4053]. 978-0-12-809279-8 [4054]. 978-0-12-809981-0 [4111]. 978-0-12-811991-4 [4009]. 978-0-12-817390-9 [4186, 4186]. 978-0-12-818642-8 [4183]. 978-0-12-821940-9 [4189]. 978-0-12-821987-4 [4216]. 978-0-19-193849-8 [4328, 4348]. 978-0-19-286417-8 [4311]. 978-0-19-289507-3 [4199]. 978-0-19-875986-7 [4162]. 978-0-19-878470-8 [4097]. 978-0-19-879651-0 [4064]. 978-0-19-879725-8 [4100]. 978-0-19-880911-1 [4137]. 978-0-19-881926-4 [4073]. 978-0-19-882193-9 [4017]. 978-0-19-883025-2 [4203]. 978-0-19-883026-9 [4203]. 978-0-19-883191-4 [4025]. 978-0-19-884222-4 [4109]. 978-0-19-885187-5 [4156]. 978-0-19-885188-2 [4156]. 978-0-19-885379-4 [4187]. 978-0-19-885836-2 [4158]. 978-0-19-886078-5 [4198]. 978-0-19-886188-1 [4276]. 978-0-19-886204-8 [4266]. 978-0-19-886472-1 [4178]. 978-0-19-886524-7 [4175]. 978-0-19-886910-8 [4320]. 978-0-19-890378-9 [4016]. 978-0-19-921561-4 [4000]. 978-0-19-958704-9 [2644]. 978-0-199-34457-4 [3459]. 978-0-224-06262-6 [3102]. 978-0-241-41113-1 [4117]. 978-0-323-85093-3 [4224]. 978-0-367-33891-6 [4269]. 978-0-367-40692-9 [4217]. 978-0-367-45736-5 [4257]. 978-0-367-68526-3 [4259]. 978-0-367-71039-2 [4305]. 978-0-367-76837-9 [4294]. 978-0-367-76998-7 [4319]. 978-0-367-77000-6 [4297]. 978-0-429-32952-4 [4269]. 978-0-521-19357-3 [4141]. 978-0-521-86273-8 [4026]. 978-0-521-87507-3 [3078]. 978-0-691-12588-6 [4340]. 978-0-691-15579-1 [4131]. 978-0-691-16292-8 [4007]. 978-0-691-16876-0 [4356, 4344]. 978-0-691-17487-7 [4027]. 978-0-691-17790-8 [3945]. 978-0-691-17973-5 [4085]. 978-0-691-18386-2 [4023]. 978-0-691-20040-8 [4274]. 978-0-691-20166-5 [4200]. 978-0-691-20177-1 [4310]. 978-0-691-20189-4 [4282]. 978-0-691-20612-7 [4184]. 978-0-691-20672-1 [4205]. 978-0-691-20710-0 [4285]. 978-0-691-20712-4 [4301]. 978-0-691-20734-6 [4190]. 978-0-691-20739-1 [4176]. 978-0-691-20983-8 [4151]. 978-0-691-21130-5 [4355]. 978-0-691-21523-5 [4316]. 978-0-691-21547-1 [4338]. 978-0-691-22365-0 [4292]. 978-0-691-23161-7 [4274]. 978-0-691-23504-2 [4339]. 978-0-691-23566-0 [4291]. 978-0-691-24007-7 [4327]. 978-0-691-24178-4 [4275]. 978-0-7354-2332-9 [4206]. 978-0-7503-1150-2 [4174]. 978-0-7503-1470-1 [4221]. 978-0-7503-1713-9 [4139]. 978-0-7503-2078-8 [4169]. 978-0-7503-2140-2 [4150]. 978-0-7503-2294-2 [4136]. 978-0-7503-2374-1 [4204]. 978-0-7503-2376-5 [4204]. 978-0-7503-2460-1 [4209]. 978-0-7503-2558-5 [4144]. 978-0-7503-2636-0 [4303]. 978-0-7503-2672-8 [4207]. 978-0-7503-2676-6 [4147]. 978-0-7503-2688-9 [4161]. 978-0-7503-3081-7 [4188]. 978-0-7503-3083-1 [4188]. 978-0-7503-3227-9 [4306]. 978-0-7503-3399-3 [4201]. 978-0-7503-3439-6 [4251]. 978-0-7503-3449-5 [4249]. 978-0-7503-3461-7 [4170]. 978-0-7503-3521-8 [4140]. 978-0-7503-3539-3 [4244]. 978-0-7503-3567-6 [4241]. 978-0-7503-3581-2 [4233]. 978-0-7503-3609-3 [4142]. 978-0-7503-3611-6 [4142]. 978-0-7503-3631-4 [4347]. 978-0-7503-3639-0 [4260]. 978-0-7503-3649-9 [4173]. 978-0-7503-3651-2 [4173]. 978-0-7503-3653-6 [4222]. 978-0-7503-3655-0 [4222]. 978-0-7503-3659-8 [4202]. 978-0-7503-3713-7 [4232]. 978-0-7503-3757-1 [4281]. 978-0-7503-3758-8 [4281]. 978-0-7503-3759-5 [4281]. 978-0-7503-3760-1 [4281]. 978-0-7503-3801-1 [4234]. 978-0-7503-3863-9 [4240]. 978-0-7503-3867-7 [4290]. 978-0-7503-3875-2 [4253]. 978-0-7503-4049-6 [4325]. 978-0-7503-4051-9 [4325]. 978-0-7503-4957-4 [4341]. 978-0-7503-5197-3 [4343]. 978-0-7503-5217-8 [4333, 4337]. 978-0-7503-6042-5 [4353]. 978-0-7503-6043-2 [4353]. 978-0-7503-6044-9 [4353]. 978-019-8726-50-0 [4038]. 978-019-8802-87-7 [4041]. 978-019-8803-80-5 [4043]. 978-019-8831-86-0 [4082]. 978-052-1117-11-1 [4021]. 978-1-003-18710-3 [4270]. 978-1-009-01369-7 [4182]. 978-1-009-16027-8 [4326]. 978-1-009-16028-5 [4322]. 978-1-009-22641-7 [4342]. 978-1-009-25315-4 [4346]. 978-1-009-27655-9 [4309]. 978-1-009-36406-5 [4345]. 978-1-032-00121-0 [4284]. 978-1-032-02116-4 [4270]. 978-1-032-02120-1 [4302]. 978-1-032-03393-8 [4270]. 978-1-032-10499-7 [4318]. 978-1-032-15169-4 [4349]. 978-1-032-19739-5 [4283]. 978-1-032-22146-5 [4308]. 978-1-032-25602-3 [4296]. 978-1-032-25607-8 [4267]. 978-1-032-26065-5 [4336]. 978-1-032-27589-5 [4269]. 978-1-032-33555-1 [4271]. 978-1-107-02761-9 [4107]. 978-1-107-03429-7 [3421]. 978-1-107-03629-1 [4015]. 978-1-107-07657-0 [4079]. 978-1-107-11074-8 [4032]. 978-1-107-12920-7 [4072]. 978-1-107-13476-8 [4146]. 978-1-107-13522-2 [4157]. 978-1-107-13739-4 [4001]. 978-1-107-15445-2 [4048]. 978-1-107-16014-9 [4006]. 978-1-107-16365-2 [4003]. 978-1-107-18229-5 [4037]. 978-1-107-18941-6 [4040]. 978-1-107-18961-4 [4035]. 978-1-107-18963-8 [4066]. 978-1-107-19488-5 [4165]. 978-1-107-68453-9 [3425]. 978-1-108-12049-4 [4210]. 978-1-108-41507-1 [4063]. 978-1-108-41531-6 [4052]. 978-1-108-41846-1 [4078]. 978-1-108-42209-3 [4019]. 978-1-108-42731-9 [4071]. 978-1-108-42756-2 [4060]. 978-1-108-42808-8 [4020]. 978-1-108-42985-6 [4061]. 978-1-108-42990-0 [4122]. 978-1-108-46650-9 [4034]. 978-1-108-47003-2 [4077]. 978-1-108-47122-0 [3994]. 978-1-108-47154-1 [4039]. 978-1-108-47164-0 [4130]. 978-1-108-47699-7 [4081]. 978-1-108-47700-0 [4005]. 978-1-108-47701-7 [4002]. 978-1-108-47743-7 [4092]. 978-1-108-48084-0 [4166]. 978-1-108-48368-1 [4168]. 978-1-108-48805-1 [4335]. 978-1-108-48998-0 [4309]. 978-1-108-49 [4177]. 978-1-108-49113-6 [4135]. 978-1-108-49145-7 [4248]. 978-1-108-49212-6 [4093]. 978-1-108-49493-9 [4197]. 978-1-108-49698-8 [4045]. 978-1-108-71751-9 [4046]. 978-1-108-72204-9 [4065]. 978-1-108-72740-2 [4211]. 978-1-108-81978-7 [4105]. 978-1-108-83497-1 [4129]. 978-1-108-84140-5 [4219]. 978-1-108-84142-9 [4185]. 978-1-108-84479-6 [4350]. 978-1-108-90616-6 [4143]. 978-1-108-94870-8 [4307, 4352]. 978-1-108-95315-3 [4286]. 978-1-1087-0130-3 [4084]. 978-1-118-17361-9 [4116]. 978-1-118-93906-2 [4334]. 978-1-118-93907-9 [4334]. 978-1-119-26770-6 [4235]. 978-1-119-27728-6 [4120]. 978-1-119-46852-3 [4354]. 978-1-119-50616-4 [4250]. 978-1-119-84086-2 [4273]. 978-1-119-88038-7 [4332]. 978-1-119-90923-1 [4304]. 978-1-138-49993-5 [4094]. 978-1-138-58387-0 [4317]. 978-1-138-59878-2 [4258]. 978-1-316-99574-7 [4215]. 978-1-4200-6471-1 [2701]. 978-1-78634-490-8 [4069]. 978-1-78634-511-0 [3956]. 978-1-78634-572-1 [4057]. 978-1-78634-704-6 [4070]. 978-1-78634-707-7 [4112]. 978-1-78634-712-1 [4114]. 978-1-78634-725-1 [4114]. 978-1-78634-788-6 [4265]. 978-1-78634-955-2 [4223]. 978-1-80061-342-3 [4243]. 978-1-84831-044-5 [2329]. 978-1-84832-615-6 [2772]. 978-110-7106-77-2 [4042]. 978-110-8426-71-8 [4058]. 978-110-8488-39-6 [4036]. 978-111-9425-39-7 [3995]. 978-178-6346315 [4083]. 978-178-6347-03-9 [4095]. 978-3-030-04078-9 [4115]. 978-3-030-05449-6 [4031]. 978-3-030-12877-7 [4018]. 978-3-030-13612-3 [4091]. 978-3-030-50083-2 [4149]. 978-3-030-53693-0 [4148]. 978-3-030-55936-6 [4118]. 978-3-030-61372-3 [4171]. 978-3-030-62518-4 [4160]. 978-3-030-63409-4 [4145]. 978-3-030-72530-3 [4208]. 978-3-030-75638-3 [4191]. 978-3-0304-6776-0 [4154]. 978-3-0304-6776-0/ISBN [4154]. 978-3-0304-6777-7 [4154]. 978-3-0305-1846-2 [4155]. 978-3-031-43955-1 [4321]. 978-3-319-70071-7 [4110]. 978-3-319-74783-5 [4059]. 978-3-319-96360-0 [3957]. 978-3-319-96835-3 [4047]. 978-3-527-34493-2 [4272]. 978-3-527-34759-9 [4242]. 978-3-527-40926-6 [4022]. 978-3-527-41405-5 [4295]. 978-3-527-69805-9 [4323]. 978-3-527-83366-5 [4252]. 978-3-527-83616-1 [4295]. 978-3-527-83617-8 [4295]. 978-3-658-32645-6 [4218]. 978-4-431-54170-7 [3993]. 978-9-81-461268-5 [3337]. 978-9-81-461270-08 [3337]. 978-981-12-0119-6 [4004]. 978-981-12-0186-8 [4172]. 978-981-12-0423-4 [4106]. 978-981-12-0507-1 [4044]. 978-981-12-0645-0 [4123]. 978-981-12-0769-3 [4113]. 978-981-12-1389-2 [4138]. 978-981-12-1634-3 [4196]. 978-981-12-2646-5 [4293]. 978-981-12-2654-0 [4225]. 978-981-12-3114-8 [4163]. 978-981-12-4790-3 [4268]. 978-981-12-4792-7 [4268]. 978-981-12-4928-0 [4268]. 978-981-12-4959-4 [4256]. 978-981-12-6177-0 [4255]. 978-981-12-6800-7 [4324]. 978-981-12-8741-1 [4351]. 978-981-1230-51-6 [4231]. 978-981-15-9049-8 [4220]. 978-981-15-9437-3 [4254]. 978-981-1590-61-0 [4159]. 978-981-322-794-1 [4033]. 978-981-322-923-5 [4164]. 978-981-323-518-2 [4086]. 978-981-323-568-7 [3998]. 978-981-323-577-9 [4121]. 978-981-323-642-4 [3996]. 978-981-323-685-1 [4008]. 978-981-323-691-2 [4030]. 978-981-323-959-3 [4099]. 978-981-3233-331 [4028]. 978-981-3233-39-3 [4029]. 978-981-3237-70-4 [4067]. 978-981-3239-87-6 [4080]. 978-981-327-002-2 [3999]. 978-981-327-075-6 [4068]. 978-981-327-133-3 [4098]. 978-981-327-206-4 [4062]. 978-981-327-252-1 [4119]. 978-981-327-399-3 [4108]. 978-981-327-417-4 [4055]. 978-981-327-738-0 [4056]. 978-981-327-796-0 [4167]. 978-981-327-831-8 [4132]. 978-981-327-850-9 [4024]. 978-981-3272-24-8 [4096]. 978-981-3276-82-6 [4010]. 978-981-3278-18-9 [3997]. 978-981-4350-51-8 [2893]. 978-981-4612-67-8 [3337]. 978-981-4616-96-6 [3423]. 978-981-4616-97-3 [3423]. 978-981-4616-99-7 [3423]. 981 [2082]. 981-238-004-3 [2082]. 98pp [4057]. 9th [193].

A-Handbook-of-Mathematical-Methods-and-ProblemSolving-Tools-for-Introductory-Physics-Second-Edition [4201]. Á.S [2979]. Aaron [2932]. Abbott [2419]. ABC [2989]. ABCs [4223]. Abdus [2870, 2239, 3976]. Abhay [3508]. Abingdon [4271, 4257, 4259, 4269, 4270]. above [2459, 2508, 2412, 2301, 2394, 2372, 2355, 2432, 2373, 2656, 2484, 2721, 2736, 3069, 3021, 4207, 4335]. Abragam [3026, 129]. Abraham [3081, 2748]. Abrahams [2755]. Absolute [1476, 430, 229]. Absorption [2680, 141, 573, 3966, 135, 911]. abstract [3525]. Absurdly [4411]. Academic [584, 4009, 3776, 4211, 2311]. academics [2396, 2884, 3432]. Accardi [2945]. Acceleration [2569, 2480, 1656, 1648, 430]. Accelerator [2757, 3371, 1660, 3208, 1276, 1029, 3921, 3484, 3052, 3986, 2885]. Accelerators [3442, 3618, 3988, 3173, 1959, 1054, 3350, 3349, 3371, 2489]. accessible [3244, 2343, 3078]. Accident [1154]. according [3469, 3949]. account [1098, 2392, 2534, 2906]. accounted [1171]. Accounts [2637]. Accreting [2441]. Accurate [125]. Acharya [2943]. Achim [4336]. acoustic [482, 191, 1203, 3360]. acoustics [4120, 1726, 969, 1000]. acquisition [3351]. Across [2350, 141, 3931, 3601]. actinide [636]. actinides [913]. Action [2301, 1839, 2005, 2884, 3762]. Active [3266, 3094, 2835, 1516, 3891, 2467, 2432]. Activity [2993, 3354, 2013, 2325, 470, 2440, 3897]. actual [3593]. Acuña [4170]. ad [2104]. Adachi [3083]. Adaptation [2695]. Adaptive [2589, 2638, 3388, 3701]. added [4117]. Addison [3768]. additive [626]. Adiabatic [227, 405, 723]. Adil [4290]. administrative [397]. administrators [2548, 3206]. Adrian [3017, 4198, 3188, 3673]. AdS [3592, 2893]. AdS/CFT [3592, 2893]. Advanced [3333, 2926, 3037, 2801, 4417, 3364, 3232, 2416, 3014, 2322, 3178, 2921, 4120, 4060, 2883, 3122, 3318, 3393, 3211, 3052, 3350, 2502, 2788, 3073, 4343, 3509, 4241, 4217, 4284, 4275, 2936, 3180, 4202, 3301, 2911, 3107, 2458, 3011, 2966, 3485, 3072, 4293, 2865, 3384, 2459, 3383, 4184, 3349, 3137, 4113, 4356, 2866, 3035, 2942, 3142, 3123, 3296, 3658, 3023, 3507, 2749, 2842, 2838, 2836, 2853, 3275, 3194, 3353, 3408, 3488, 4183, 4240, 4272, 3144, 2428, 3092, 2907, 3168, 2735, 2494, 2992, 3245, 4325, 2541, 2845, 2901, 4072, 4178, 4190, 4198, 4252, 3279, 3343, 3344, 3450, 3452, 4257, 4259]. advanced [4269, 4309, 4281, 4295, 4323, 3323, 3407, 2887, 3184, 4242, 4250, 3081, 3043, 2993, 3435, 3438, 3437, 3422, 4105, 4339, 2597, 3099, 4187, 4294, 3418, 2348, 2387, 3046, 3224, 3287, 3196, 3270, 2333, 4318, 4319, 3431, 3233, 3366, 3429, 2844, 4165, 3034, 3235, 3398, 3492, 2891, 3128, 2365, 2817, 3114, 3277, 3005, 3210, 3415, 3446, 4186, 2292, 2461, 3423, 3026, 2950, 2881, 2929, 2594, 2497, 2899, 2649, 2499, 3084, 3259, 3307, 4200, 4205, 3049, 3231, 3234, 3065, 3216, 2839, 3248, 3282, 3272, 2664, 4349, 2415, 2835, 2874, 2924, 2837, 2834, 2880, 3062, 3066, 3119, 2997, 3063, 4054, 4083]. advanced [4053, 4080, 4079, 4169, 4170, 4189, 4234, 2436, 2657, 2677, 2676, 2943, 2909, 2728, 2782, 3183, 4351, 3096, 4306, 3064, 3414, 3380, 3476, 4071, 4255, 4188, 4276, 4350, 4070, 4066, 4077, 4167, 4297, 4308, 2310, 2314, 2390, 2608, 2620, 2896, 3228, 3461, 4253, 4265, 4266, 4332, 2719, 2805, 3148, 3255, 3264, 3265, 3482, 3483, 3151, 4174, 2579, 4216, 2468, 2501, 3286, 2951, 2947, 2298, 2401, 2398, 2451, 2399, 2619, 3191, 3185, 3278, 4290, 2488, 2905, 3087, 2848, 2454, 2453, 2388, 2373, 2470, 2613, 2646, 2591, 2656, 2663, 2484, 2956, 2721, 2736, 3082, 3068, 3153, 3041, 3021]. advanced [3242, 3170, 3260, 3411, 3338, 3372, 3475, 3495, 4058, 4207, 4176, 4215, 4175, 4185, 2922, 2915, 2882, 2967, 2919, 2852, 2704, 4101, 2744, 3741, 3750, 2615, 3076, 3661, 3235, 2395, 2713, 2744, 2873, 2875, 4166, 2401, 3191, 3746]. advanced/research [2428]. Advances [3326, 1020, 4224, 2750, 3132, 771, 768, 963, 1203, 39, 2921, 3051, 2798, 2775, 4159, 3874, 3882, 3735, 3028, 44]. Advancing [90, 4125]. advent [4163]. Adventure [2933, 3603]. Adventures [2581, 2496, 3942, 4109, 2922]. Advice [4239, 2548]. Advisory [391]. Aegerter [4034]. Aeppli [3539]. Aerodynamics [2841, 3741]. aerosol [706]. aerosols [451]. Aert [4224]. Aerts [3320]. afraid [4145]. after [3468, 1628, 4048, 3465, 3949, 3267]. afterglow [736]. aftermath [3225]. AG [4208, 4171]. again [1543]. against [3603]. Agarwal [3019]. Age [3405, 4394, 4400, 4401, 18, 43, 3875, 1903, 1951, 1762, 1741, 2143, 1631, 3857, 3111, 3149, 3810]. Agency [834]. Agents [3095]. Ages [3044, 3802]. aggiornamento [86]. aggregates [1357]. Aggregation [1357, 1912]. Agio [3367]. Agop [3554]. Agranovich [3632]. Agrawal [2820]. Agresti [3562]. agricultural [4240]. Aharonov [1685]. Aharony [4096]. Ahmed [3401, 3432]. Ahufinger [3022]. AI [4349]. Aichelin [4159]. aids [307, 846]. Aihara [4173]. aims [1778]. AIP [4206]. air [517, 2085, 2216, 4049, 1050, 138, 1035]. Airless [4054]. Aitchison [3509, 3216, 2263]. Ajawad [3315]. Akhlesh [2546]. Akinwande [2838]. Akio [3319]. Akira [3501]. Al [4291, 2857]. Al-Khalili [4291, 2857]. Alain [2686, 3287, 2929, 2760]. Alamos [1092, 1092]. Alan [3013, 3562, 3066, 3506]. Alastair [2904, 4334]. Alavi [3504]. Albert [1165, 3473, 3461, 1165, 3473, 2912]. Alberto [2565, 3608, 2624]. alchemist [4027, 1972, 2026]. alchemy [4132, 1688]. Alcubierre [3185]. Alessandro [3317, 3666, 3997]. Alex [2774, 2933, 2496, 3215, 2496]. Alexander [3079, 2767, 3133, 3371, 3577, 3492, 3189, 3208, 2804, 3606, 4]. Alexei [3079]. Alexey [3530, 4234, 3788]. Alexia [3659]. Alfred [3558]. algebra [4317, 1730, 1704]. algebra-based [4317]. Algebraic [2414]. algebras [3409, 4008]. Algorithmic [2568]. algorithms [3927, 1729, 3782, 3023]. Ali [3154, 2787, 3534]. Aliaksandr [4257]. Alice [1092]. Alien [3378]. Alive [3056]. alkali [1271]. All-laser-driven [3387]. all-optical [1594]. Allan [3506, 2580]. Allday [2498]. Allemand [3487]. Allen [291, 4117, 3064, 2953, 3164, 3622, 274, 282]. Allori [4154]. Alloy [3889]. alloys [314, 777, 1007, 295, 638, 536, 3103]. almost [2231, 3867]. Alonso [4032]. Alpha [1465, 2024]. Alpha-particle [2024]. Also [4225]. Altarelli [3935]. Altena [3082]. alternating [5, 398]. Alternative [3089, 2789]. Alternatives [2412]. Altland [3994]. Alù [3367]. always [981]. Amabilino [2503]. Amador [2533]. Amaldi [3442]. Amanda [2984]. Amar [2760]. amateur [3478, 3773, 3732, 2438, 3345]. amateurs [4261]. Amati [2311]. Amazing [2645]. Ambjorn [4271]. Amen [2876]. American [3751, 151, 479, 675, 701]. Amir [3956]. Amiri [3504]. Amit [3539, 3431]. Ammon [3498]. Amoroso [3261]. amorphous [1049, 901, 962, 543, 1278, 1282, 809, 3764]. ampere [13, 238, 185, 202]. Amplifiers [2820]. amplitude [3614]. Amrein [2492]. Amy [4186]. Ana [2771, 1844]. Ana-Maria [1844]. analogies [4243]. Analogue [3232, 164, 2731, 3640]. analogues [2731, 1892]. Analysis [4360, 1131, 2901, 633, 4385, 2623, 2943, 3188, 2592, 4078, 328, 802, 3715, 3947, 1841, 3895, 3808, 4222, 3577, 1886, 4007, 768, 299, 2727, 2354, 2940, 3629, 3793, 3868, 3896, 3131]. analytic [2359]. Analytical [2781, 2757, 3895, 475, 3943, 2845]. Anand [4178]. Anantha [2394]. Ananthaswamy [2519]. Anatoliy [4061]. Anatoly [2942]. anatomy [3078]. Anca [3240]. Ancestry [4410, 3423]. ancient [3251]. ancillas [3545]. and/or [2864]. Anderson [2755, 4048, 3766, 2886]. Andrady [4235]. Andrea [3179, 3004, 4146, 2526, 3367, 4208, 2400]. Andreas [2745]. Andrei [3576, 2898, 3277, 3483, 3306]. Andrej [4293]. Andreotti [3213]. Andrés [3189]. Andrew [3037, 4045, 3349, 3150, 3577, 4318, 3018, 2539, 2983, 2839, 3166, 4320, 2620, 3281, 4335, 2787]. Andrews [3274, 3087]. Andrey [2913]. Andrzej [3350]. Andy [4144]. Anecdotal [2827]. Angelo [4154, 4208]. Anger [2870]. Angilella [3652]. Anglos [2803]. Angular [414, 2629, 3087, 609, 2374]. Angus [2828]. Anh [4302]. Anil [2788, 2519]. animal [1399]. animals [4219]. Anindya [4251, 4350]. Anirban [4350]. Anisotropy [464, 2126, 2546]. Anisovich [3218]. Ann [2537, 2933]. Anna [3022, 3472]. Anne [3433]. annealing [2170, 3782]. annealing-based [3782]. Annemarie [3408]. Annibale [3789]. Annick [4224]. Anniversary [2238, 3461, 4346, 4092]. annotations [3473]. Annual [391, 137, 97, 158, 180, 221, 289, 341, 308, 26, 98]. anomalous [781, 2275, 3866]. anomaly [2186]. Answer [2610, 3708]. answers [1545, 4114, 4080]. ant [1578]. Antarctica [2906, 3113]. antennas [1356, 3159, 2273, 3367, 3404]. anthology [3432]. Anthony [3509, 3097, 4235, 3216, 4253, 3528]. anthropic [1408]. anti [2120]. anti-neutrino [2120]. anticipating [4028]. Antiferromagnetic [943]. antiferromagnetics [912]. Antihydrogen [2212]. Antikythera [3249]. Antimatter [1781, 1647, 3981, 2392, 4025]. antineutrino [122]. antiparticle [4173]. Antiprotonic [1451]. antiquities [2405]. Antiquity [2887, 3044]. Antoine [2568]. Anton [2827]. Antonello [3179]. Anupam [2978]. anvil [1363]. any [2352]. anybody [2308]. anyone [2463, 2466]. Anyons [1521]. Aoki [2472]. Aoyama [2657]. Aperiodic [2389]. apertures [1737]. Apollo [2003]. App [3239]. Apparatus [268, 2463, 2463, 31]. apparent [457]. apple [1834]. Application [1913, 913, 425, 683, 2629, 2953, 3934, 832, 223, 587, 2961, 4095, 4265]. application-driven [3934, 4095, 4265]. Applications [2893, 554, 1371, 3371, 798, 2749, 2842, 2836, 3207, 3444, 885, 3165, 2541, 3930, 2394, 4380, 2457, 2566, 2771, 2716, 4227, 2584, 3058, 2517, 3310, 3276, 2924, 2782, 2592, 543, 593, 3766, 2094, 2170, 1340, 1177, 618, 818, 121, 4417, 422, 3638, 298, 4138, 567, 3858, 3557, 3617, 384, 3630, 134, 4353, 862, 3768, 2272, 2811, 759, 4255, 1099, 1480, 4266, 2147, 2454, 4215, 4168, 2883, 3051, 2386, 3038, 2942, 2948, 3408, 2975, 2342, 2791, 2625, 2965, 3280, 3204, 2726, 3003, 3447, 2803, 3084, 2998, 2784, 2943, 2825, 3183, 2607, 3014, 2717, 3059, 2579, 2799, 3082, 3068]. Applications [2598, 3767, 3498, 4284, 3619, 4254, 3646, 4240, 4026, 4323, 4242, 4250, 4158, 3531, 3486, 3665, 4282, 56, 4249, 3634, 3735, 4070, 4334, 3537, 3503, 3814, 3743, 3581, 3775, 3799, 3947, 4235, 2743, 2793, 3872, 3940]. Applied [205, 3493, 150, 175, 3953, 2486, 3747, 3653, 154, 3714, 3961, 3047, 3883, 2523, 3996, 3977, 2488]. Appraised [3071]. Approach [2686, 2966, 3089, 4365, 3243, 2422, 2398, 2669, 3300, 2025, 3480, 211, 152, 3914, 3920, 3563, 4321, 3466, 2811, 3572, 3880, 4306, 4129, 4219, 3221, 2117, 745, 2432, 2291, 4139, 4052, 4199, 3672, 4147, 4216, 3475, 3746, 3840, 3424, 3884]. approaches [2604, 2685, 2807, 2431]. appropriate [4173]. April [3063]. apse [59]. Aquatic [834]. Arabatzis [4347]. archaeological [2515]. Archaeology [2590, 150, 175, 577]. Archimedes [510]. Architectural [1000]. Architecture [1386, 3650, 3421]. Architectures [2795]. Arditty [1191]. area [4289]. Aretxabaleta [3656]. Argand [425]. Argon [3154]. Arieh [3645, 3654]. Ariel [261]. arising [764, 3418, 240]. aristocrat [765]. Aristotle [2353]. Arkadiusz [3497]. Arnab [2594]. Arnold [2585, 3235, 4336, 4224]. Arntzenius [3182]. Arregui [3179]. Arribas [2779]. Arrighi [2319]. arrow [1928, 1835]. Arrows [2843, 3335]. Art [2417, 3520, 1058, 3472, 871, 3945, 820]. Artés [2979]. Artful [1858]. Arthur [1136, 3246, 739]. article [4154, 4221, 3448, 4270, 3432, 4224, 4172]. articles [2506]. artificial [840, 262, 2326, 3084]. Artist [4407, 3869]. artistry [4091]. arts [393]. Arun [3746]. Arunachalam [3014]. Asa [2772]. Asher [4058]. Ashim [2556]. Ashok [2936, 2693, 3409]. Ashtekar [3508]. Asked [4269]. asking [1974]. Aslamazov [2913]. Aspect [2686, 645, 2250, 3287, 2249]. Aspects [2897, 755, 3176, 1074, 2681, 162, 578, 1104, 2483, 135, 962, 4226, 3836, 228, 2609, 2976]. aspiring [4084]. Aspuru [3506]. Aspuru-Guzik [3506]. Assael [4269]. Assembled [2461, 2014]. assessment [1903, 1068, 1703]. assisted [1452]. Association [151, 75, 221, 136, 169, 307, 289, 341, 26, 98, 3654, 79, 147, 180, 213]. Aste [2334]. Aster [4053]. asteroids [1640, 1892, 3705]. Asteroseismic [4007]. astra [2104]. astray [2023]. Astrobiology [2089, 3129, 3263]. astrochemistry [3993]. astrodynamics [3685]. astrometry [4264, 3069, 3082]. Astronautics [2889]. Astronomer [2299, 944, 3926]. astronomers [4285, 2438, 3773, 3732]. Astronomical [2788, 3435, 1608, 2442, 3039, 3267, 3705, 4112, 1516, 2788, 2504, 4112, 2835].

Astronomy [3472, 2727, 353, 929, 2682, 4114, 4200, 4205, 3243, 2778, 3183, 4125, 3215, 4133, 617, 844, 2439, 1587, 4143, 1551, 1783, 87, 3577, 2519, 3060, 954, 352, 1419, 2169, 1081, 3985, 1661, 1233, 3388, 2450, 2779, 3016, 2871, 2420, 2610, 3564, 3708, 3643, 4040, 3532, 4065, 3655]. Astroparticle [2881, 3593]. astrophotography [892]. Astrophotonics [4314]. Astrophysical [2750, 1351, 3978, 2404, 3858, 3286, 2424]. Astrophysics [2788, 3168, 2457, 2566, 4382, 2884, 2295, 3082, 710, 2252, 4325, 1279, 3854, 3824, 2251, 2023, 3926, 1716, 3228, 1734, 2738, 2972, 2415, 2310, 4339, 4063, 4174, 3113, 3435, 3438, 3865, 2682, 2594, 3533, 2442, 3621, 3985]. astrospheres [3891]. Asymmetric [788, 800]. asymmetry [4173, 1647]. asymptotic [3954]. asymptotics [3828, 2776, 4293]. Athanasios [3666]. Athanassoula [2467]. Atkins [2829, 1359]. Atlantic [789]. ATLAS [2647, 3020, 3732, 4031, 4015, 2505, 3066, 3648, 4345]. Atmosphere [2890, 1699, 3520, 2032, 3790, 1806, 178, 670, 1987, 923, 2223, 4104, 3211, 2593, 2288, 2295]. Atmospheres [2654, 1022, 3733, 2618]. Atmospheric [3520, 644, 639, 2807, 2905, 1406, 706, 978, 1480, 3610]. Atom [1634, 2128, 1832, 623, 637, 653, 3048, 2067, 1458, 4277, 3464, 1947, 3161, 4230, 1677, 1452, 2131, 4142, 1378, 1942, 1892, 1221, 3800, 3224]. Atom-atom [623, 637, 653]. atom-ion [623, 637, 653]. Atomic [2921, 2788, 4357, 4358, 3322, 1791, 2329, 38, 578, 2706, 2129, 2834, 3048, 834, 1913, 3946, 631, 1412, 3, 1641, 3938, 1270, 2763, 1349, 1503, 1776, 1602, 4112, 1664, 4224, 2445, 1277, 3939, 1055, 2277, 412, 311, 1817, 2788, 77, 241, 675, 2329, 3245]. atomic-resolution [4224]. Atomistic [3127, 4334]. Atoms [2220, 3393, 3022, 50, 4362, 1041, 3071, 1606, 4106, 3763, 1451, 1531, 2892, 1833, 565, 1720, 4118, 3659, 2175, 1119, 916, 1536, 1112, 800, 1162, 2287, 3631, 2107, 980, 1065, 1428, 623, 637, 1151, 1759, 2154, 4022, 2552, 3467, 1028, 1378, 1658, 2326, 2783, 2953, 4119, 2949, 3335, 3598, 3798]. Attacks [3517]. Attard [3038]. attenuation [203]. Attosecond [3011, 3729, 3550, 4172, 3696]. attributes [580]. audience [2427]. Audretsch [2371]. Auffèves [3659]. Auger [3792, 705]. August [3836]. Augustyn [4272]. Auletta [3282]. Aung [3319]. aurora [1648]. Auroral [3342, 604]. Australia [248]. authenticating [2405]. authentication [2228]. author [2074, 2151]. authority [810]. autobiography [3088, 3088]. Automation [323]. automobiles [1797]. available [4225]. avalanches [1792]. Avijit [2560]. Avraham [2991]. award [3476]. award-winning [3476]. awards [99]. awkward [735]. Axel [3531]. Axelsson [2441]. Axions [4261, 2225]. axis [4241]. Azároff [2639]. azoic [1622].

B [3111, 2989, 2911, 3750, 3723, 3692, 3800, 4123, 4069, 2412, 1040, 2428, 2342, 1978, 3616, 4006, 4043, 4196, 4141, 3689, 3754, 3797, 4040, 4048, 2609, 2765, 3135, 2243, 2817, 3200, 3277, 2435, 2367, 2247, 2525, 3187, 3026, 2269, 2940, 2971, 2874, 4053, 2846, 2976, 3996, 4071, 3014, 2390, 3148, 3742, 4086, 2718, 2952, 3236, 2398, 2503, 2590, 3284, 4136]. B-G [2846]. B.A. [232]. Baaquie [3527, 2521]. Babaev [3460]. Babiker [3087]. Babinsky [2974].

Babu [3183]. Babylonian [1766]. Bachelor [4265]. Bachmann [3663]. back [1844, 3311, 4316]. Back-of-the-envelope [3311]. Backer [4224]. background [3520, 2661, 1628, 298, 2886, 3120, 1811, 3907, 706, 2292, 4157]. backward [1674]. Bacon [3614]. bacterial [2304, 1622, 1782]. Bada [2387]. Badash [1092]. Badawi [3457]. Bagarello [4093]. Bagchi [3750]. baggage [2090]. Baggott [4137, 4328, 2329, 4348, 3763, 2694, 2986]. Bailer [3755]. Bailer-Jones [3755]. Bailin [3681]. Baillon [3480]. Bailyn [3383]. Baines [4042]. Bajnok [3591]. Bajorski [2888]. Bakar [4350]. Balachandran [4071]. Balakumar [4071]. balance [1924, 1987]. Balasis [3787]. Balázs [3031]. Ball [2289, 3253]. Ballet [3100]. Ballistics [2843]. Bandarenka [4257]. bandgap [1326]. bands [4247, 1575, 3451]. Banerjee [2817]. Bang [4123, 3442, 1904, 1708, 2435, 3106]. bangs [856]. Banhart [2395]. Banks [2893, 2346]. Baojiu [4108]. Bar [3458, 1356]. Bar-llan [3458]. Barber [2389]. Barbuy [2467]. Barenghi [3716]. Barford [3116]. Barger [3032]. Barica [4242]. barium [782]. Barkataki [4267]. Barman [4309]. Barnes [3695, 2709, 2631]. Barnett [2455]. Barnsley [3497]. Baron [2765]. Barr [3613]. Barrabès [3191]. Barrat [2352]. Barrett [4145]. barrier [1118]. barriers [640]. Barrow [2312, 2456]. Barry [3430, 2520]. Barthélemy [2352, 3707]. baryogenesis [1818]. Basa [2929]. Based [3265, 4317, 844, 2892, 604, 1276, 4081, 176, 3160, 2381, 4227, 2700, 3782, 3207, 3033]. bases [810]. Basic [3395, 2490, 3446, 2487, 2780, 364, 3953, 3520, 3480, 3895, 237, 245, 1529, 4122, 2454, 2897, 3072, 3240, 3675, 2769]. Basics [2927, 2717, 2420, 3494, 3581, 2510]. Basis [2427, 2409, 3365, 179, 360, 370, 3359, 3525]. Bass [1069, 2931]. Bassi [4154]. Basu [4007, 3141]. Batanouny [2432, 4166]. Bates [899]. Batterman [3091]. Battisti [2741]. Baturin [3276]. Baulieu [3736]. Baumann [3583]. Baumberg [3788]. Baumer [2627]. Baumgarte [2649]. Baumjohann [3072]. Bayes [2959, 2215]. Bayesian [3755, 2354, 2215, 3444, 2623]. Bayin [3995]. Be [2057, 1409, 4296, 2411, 3839, 4418, 2980, 3525, 530, 4145]. Beale [2676]. Beam [2465, 2146, 1177, 816, 1475, 1503, 924, 768, 1063, 1959, 3350]. Beams [3059, 1137, 1340, 761, 2, 1112, 1972, 524, 2121, 926, 2570, 1323, 1277, 3988]. beams-origins [1340]. bearings [614]. beat [1693]. Beating [2405]. Beatriz [2467]. Beautiful [2679]. Beauty [2279, 1975, 2103, 3462, 1753, 4395, 2894, 3254]. became [2127]. Bécherrawy [2900]. Beck [2992]. Beckers [2976]. Beethoven [2206]. beginners [2386, 2337, 3961, 1708, 1618, 1814, 4353, 4055]. beginning [3909, 2293, 2810, 4274, 2355, 2594, 2045, 2442, 2401]. Behaviour [2433, 1145, 887, 424, 3242]. Behind [3128, 4286, 4307, 4352, 3197]. Behnia [3580]. Bei-Lok [4141]. Bekenstein [2912]. Belair [3196]. Belal [3527, 2521]. Belavin [3079]. Believing [3968]. Belkic [2975, 2975]. Bell [3730, 2160]. Bellac [3287]. Bellet [2794]. Bellini [3207]. Bellos [2496]. Bells [4359, 3448]. Bely [2610]. Ben [4211, 2760, 3645, 3654, 2619]. Ben-Naim [3645, 3654, 2619]. Bénard [1234]. beneath [808]. benefits [3252, 2063]. Benign [4411]. Benini [2741]. Benjamin [3374, 2536]. Benkadda [3688]. Bennemann [3187]. Bennett [4275]. Benson [2421, 3100, 2946]. Beowulf [2062]. Berezinskii [3313]. Berger [3508]. Bergman [4120]. Berkeley [3492]. Berkowitz [584]. Berkshire [2952]. Berlin [4059]. Berman [2678]. Bernard [2459, 875, 3596, 3149, 3604]. Bernath [3631]. Bernevig [3277]. Bernhard [4197]. Bernstein [3448]. Berry [3903]. Bersuker [2428]. Bert [3036]. Berthier [3002, 2915]. Berthold [3398]. Berthold-Georg [3398]. Bertin [3377]. Bertolotti [3475]. Bertone [3197, 2630]. Bertrand [2884]. Bessel [2121]. Beta [1149, 2606, 1155, 2115]. beta-decay [1155]. Beth [2639, 2720]. Bethe [3346]. Bettini [3317]. Between [3213, 4401, 410, 667, 3875, 126, 407, 640, 3930, 2133, 2765, 2343, 204, 210, 623, 637, 3310, 735, 3654, 4104]. Bevan [3188]. Beverly [3508]. Beyond [3573, 3033, 3967, 4208, 3644, 4177, 3783, 4342, 2322, 2592, 2720, 2804, 2554, 3676, 2744, 3102, 2767, 2872, 2854, 2631, 2822, 3558, 4311, 4058, 3794, 2970]. Bhatt [2850]. Bialek [3080]. Bianisotropy [2546]. Biedenharn [2374]. Big [4123, 3442, 2435, 3283, 3106, 4221, 1708, 4347]. Bigi [2586]. Bigio [3565]. Bikas [3539]. Bilal [4254]. bilayers [1309]. Bilgili [3656]. Bill [2506]. billion [4399, 3876]. Binêtruy [4064]. Binaries [3067]. binary [1559, 3970, 59, 1512, 3075]. Bing [4107]. Binney [3260]. Binoculars [2439]. Binoth [3054]. Bio [2945, 2771, 647, 3584, 2945, 2771]. Bio-Applications [2771]. bio-engineering [647]. Bio-Informatics [2945, 2945]. Bio-Interfaces [2771]. Biochemical [2541]. biochemistry [3432]. biochemists [4097]. Bioelectromagnetism [4284]. bioengineering [3432]. Biographical [811]. biographies [2668]. Biography [2057, 4418, 2565, 2508, 2870, 3017, 4073, 3102, 4355, 2191, 2886, 2983, 4117, 3006, 3490, 4320, 3048, 3149, 2934, 2650]. Bioinspired [3496]. Biological [2396, 2837, 2760, 3629, 1841, 1240, 4090, 2734, 3331, 3678, 1871, 4093, 1220, 3929, 3940, 4034, 3951, 2853]. Biology [1582, 3058, 4217, 1796, 152, 2233, 3077, 3487, 3832, 3432, 746, 3640, 2173, 3937]. biomagnetic [3630]. Biomagnetics [3630]. Biomedical [4415, 2975, 3407, 2396, 3457, 2671, 4103, 3565, 3062, 3440]. Biomedicine [2880, 3795]. Biomimetics [3104]. biomolecular [3706]. Bionanomaterials [4240]. Biophotonics [3246, 3683, 2396, 3998]. Biophysical [2409]. Biophysics [2457, 2837, 3832, 3432, 2844, 3080, 3606]. biosciences [3392]. Biosensor [2673]. Biosensors [2673, 2991, 4290]. Biosystems [2844]. Biot [238, 133]. Biotechnological [3408, 4323]. Biphoton [2723]. Bipin [2490]. Birds [3551]. birdsong [3008]. Birkbeck [53]. Biroli [3002]. Birth [3004, 2562, 2982, 2934, 2175, 2201]. Birthday [3412, 648, 694, 3681, 3976]. birthplace [2404]. Bisi [3623]. bistability [1170]. Biswas [2379]. Bits [2421, 2946]. Bjelkhagen [3232]. Black [2078, 3383, 2597, 2562, 2716, 2916, 2914, 2651, 3191, 2934, 2013, 2479, 710, 2252, 3784, 3477, 3576, 2251, 1681, 3825, 1264, 4208, 3963, 1512, 4033, 4335, 2857, 3040, 3273, 3411]. blackbody [1850]. Blackett [3702]. bladder [4241]. Blagojević [3126]. Blahut [4060].

Blandford [4184, 4190, 4176]. Blank [2370]. Blanter [2402]. Blasio [4059]. Bleaney [3200, 3200, 3026]. Bletchley [2772, 4361]. blind [846]. block [3967]. Blockbuster [3841]. Blocks [2780, 3674]. Blodgett [1202]. Bloom [2688]. Blow [3403]. blue [1748]. Blundell [2511, 3345, 2511, 2413, 2990]. Boal [3005]. Board [2478, 2697, 3158, 3288, 3725, 2958, 3569]. Bob [3512]. Bobin [3578]. Boca [4317, 4217, 4284, 4296, 4259, 4269, 4270, 4294, 4318, 4319, 4302, 4336, 4094, 4258, 4267]. Bock [1069]. Bodies [1265, 2737, 666, 2603, 4054, 3705]. Body [3393, 3022, 2350, 3090, 4231, 3652, 3591, 1145, 3884, 3920, 3456, 3972, 2030, 3311, 165, 4166, 4398, 12, 2710, 3540]. Boetttcher [3094]. Bogolubov [2665]. Bohm [1685]. Bohmian [3047, 2979]. Bohr [2770, 4073, 4328, 4348, 3224, 3048, 4142, 2558, 1412, 1156]. Boi [2856]. Boisvert [2632]. Bojowald [4209, 2716]. Bok [2527, 2596]. Bolin [2367]. Boltzmann [3902, 3798, 368, 382]. Bomb [4357, 4358, 2329, 3226, 3524, 3102, 2018, 3938]. Bona [4293, 2776]. bonanza [776]. Bond [3375, 866]. bonding [901]. Bonds [3451, 2521, 1759]. Bonifacio [1140]. Bonilla [2711]. Bonnie [2839]. Book [3670, 3806, 3834, 3817, 3757, 3776, 3767, 3807, 3766, 3942, 4133, 4132, 3622, 3711, 3953, 4120, 4078, 4060, 4106, 3878, 4149, 3651, 3741, 3841, 4010, 4057, 3764, 3763, 45, 57, 67, 130, 149, 173, 181, 189, 199, 207, 215, 226, 233, 244, 257, 271, 280, 294, 301, 309, 318, 327, 336, 343, 351, 357, 367, 373, 383, 392, 400, 408, 418, 426, 433, 443, 449, 460, 467, 473, 480, 488, 489, 496, 497, 503, 512, 519, 533, 556, 562, 569, 575, 585, 591, 599, 612, 620, 627, 634, 641, 649, 656, 663]. Book [668, 676, 681, 688, 695, 702, 708, 715, 720, 726, 733, 741, 750, 766, 758, 773, 779, 786, 787, 794, 803, 804, 814, 821, 822, 828, 835, 841, 842, 849, 851, 858, 876, 882, 888, 897, 920, 930, 939, 948, 956, 966, 973, 995, 1003, 1016, 1024, 1031, 1043, 1051, 1060, 1071, 1077, 1086, 1094, 1102, 1109, 1115, 1125, 1134, 1141, 1147, 1158, 1167, 1173, 1184, 1195, 1200, 1208, 1216, 1222, 1230, 1238, 1245, 1252, 1260, 1266, 1274, 1280, 1286, 1287, 1293, 1299, 1306, 1312, 1319, 1329, 1337]. Book [1344, 1352, 1360, 1367, 1375, 1382, 1389, 1396, 1403, 1410, 1417, 1425, 1432, 1439, 1446, 1455, 1462, 1470, 1477, 1484, 1491, 1498, 1508, 1513, 1519, 1525, 1533, 1539, 1548, 1556, 1567, 1576, 1585, 1592, 1598, 1604, 1611, 1613, 1619, 1625, 1632, 1638, 1645, 1654, 1662, 1668, 1675, 1682, 1689, 1695, 1701, 1709, 1718, 1724, 1732, 1739, 1746, 1755, 1756, 1763, 1764, 1771, 1779, 1784, 1789, 1794, 1802, 1809, 1815, 1821, 1828, 1837, 1845, 1853, 1859, 1867, 1876, 1884, 1889, 1901, 1908, 1919, 1929, 1939, 1945, 1953, 1961, 1970]. Book [1976, 1985, 1990, 1996, 2004, 2011, 2020, 2028, 2036, 2042, 2051, 2052, 2058, 2065, 2066, 2072, 2084, 2092, 2100, 2106, 2112, 2119, 2125, 2130, 2135, 2140, 2145, 2150, 2158, 2168, 2172, 2178, 2182, 2187, 2192, 2197, 2203, 2207, 2210, 2214, 2218, 2222, 2226, 2232, 2237, 2274, 2280, 2284, 3618, 3986, 3988, 4317, 4324, 3790, 3868, 3805, 3774, 3946, 4016, 4062, 4110, 4211, 4137, 4225, 4286, 3885, 3896, 3560, 3574, 3721, 3583, 3719, 3541, 3559, 3830, 4082, 4343, 4241, 4217, 4284, 4275, 3796, 3974, 4202, 4232, 4268, 3785, 3927, 4135, 3698, 3699].

Book [3700, 3750, 3828, 3918, 3696, 3619, 3901, 3772, 3912, 4140, 4139, 4254, 4038, 4000, 3934, 3539, 3580, 3625, 3710, 3635, 3835, 3899, 3898, 3872, 3873, 3944, 3984, 4032, 4009, 4056, 4055, 4210, 4273, 4271, 4304, 4296, 4206, 3656, 3716, 3644, 3840, 3943, 4035, 4231, 4293, 3677, 3695, 3723, 3579, 3609, 3692, 3800, 3875, 3791, 3748, 3759, 3848, 3855, 3881, 4005, 3945, 4045, 3913, 3935, 3914, 4018, 3938, 3955, 4028, 4073, 4123, 4069, 4154, 4182, 4177, 4134, 4184, 4191, 4221, 3646, 3839, 3877, 4027, 4100, 4113, 4356, 4328, 3715, 3663, 3869, 4307, 4118, 3599, 3585]. Book [3900, 3738, 3947, 3919, 4061, 4333, 3554, 3544, 3657, 3658, 3659, 3744, 3747, 3538, 4143, 4209, 3628, 3629, 3706, 3889, 3961, 4183, 4240, 4272, 3530, 3652, 3591, 3648, 3844, 3758, 3931, 4030, 4084, 4155, 4248, 3647, 4251, 112, 326, 4325, 4327, 4355, 3620, 4291, 4303, 4285, 4138, 4173, 3650, 3637, 3858, 3845, 3859, 3784, 3856, 3987, 3989, 3958, 4347, 4344, 4348, 454, 3626, 3679, 3883, 3762, 3793, 3739, 3983, 3930, 4026, 4072, 4052, 4178, 4190, 4148, 4198, 4163, 4252, 3561, 3557, 3617, 3616, 3895, 3808, 3894, 3753, 3755, 3809, 3960, 4006, 4021, 3962, 4081]. Book [4156, 4204, 4130, 4203, 4162, 4235, 4222, 4257, 4259, 4269, 4309, 4281, 4270, 4295, 4274, 4305, 4323, 3771, 3884, 4025, 3948, 4002, 3915, 3925, 4001, 4046, 4043, 3975, 4041, 3920, 3954, 3999, 4029, 4044, 4008, 4024, 4085, 4108, 4068, 4119, 4098, 4196, 4141, 4242, 4250, 3552, 3553, 3564, 3705, 3630, 3669, 3668, 3588, 3709, 3576, 3671, 3708, 3577, 3627, 3631, 3865, 3816, 3754, 3886, 3861, 3854, 3891, 3860, 3804, 3801, 3797, 3752, 3769, 3773, 3751, 3787, 4031, 4037, 4047, 3993, 4040, 3994, 4048, 3957, 4042, 3936, 4003, 4090, 4091, 4105, 4146, 4345, 4338, 4342, 4346, 4339, 4340, 4187, 4283, 4294, 3661, 3551]. Book [4322, 4353, 4318, 4319, 3135, 3660, 3623, 3998, 4165, 3540, 3810, 3980, 3592, 4302, 547, 3563, 3794, 3921, 4097, 4107, 4157, 4336, 4292, 3825, 4150, 4186, 3578, 3587, 3608, 3678, 3597, 3832, 4007, 4017, 3982, 3525, 3717, 3614, 3632, 3862, 4112, 4114, 4115, 4117, 4159, 4161, 4158, 4160, 3720, 3768, 3770, 3843, 3952, 4321, 3562, 3531, 3653, 3691, 3589, 3874, 3882, 3789, 3740, 3950, 4099, 4093, 4200, 4151, 4205, 4224, 3819, 3978, 3533, 3532, 3697, 3849, 3926, 3527, 3712, 3756, 3973, 4164, 3316, 3665, 3749, 3846, 3905, 4019, 4282, 4326, 4121, 4131, 3718, 3666]. Book [3703, 3842, 4122, 4197, 4349, 4337, 4354, 4201, 4223, 4244, 4218, 4249, 4316, 3558, 3584, 3582, 3607, 3634, 3649, 3682, 3683, 3735, 3736, 3833, 3734, 3879, 3880, 3795, 3827, 3836, 3922, 3959, 3917, 4020, 4022, 4036, 3949, 4054, 4083, 4053, 4080, 4079, 4169, 4170, 4189, 4234, 3604, 3636, 3923, 4144, 4351, 4341, 4352, 3893, 3811, 3863, 3932, 3724, 3613, 3704, 3526, 4306, 3690, 3476, 3543, 3645, 3693, 3681, 3864, 3831, 3777, 3803, 3996, 3995, 3929, 3933, 4071, 4129, 4255, 4208, 4188, 4219, 4301, 4311, 4310, 4276, 4350, 4320, 4070, 4066, 4077, 4166, 4167, 3654]. Book [3760, 4199, 4260, 4297, 4308, 4334, 4095, 4096, 4094, 4253, 4265, 4266, 4332, 3537, 3566, 3536, 3565, 3567, 3534, 3575, 3595, 3610, 3612, 3674, 3673, 3596, 3611, 3672, 3903, 3814, 3887, 3904, 3742, 3745, 3888, 3902, 3813, 3890, 3743, 3812, 3939, 3941, 3940, 4065, 4064, 4063, 4220, 3555, 3528, 3529, 3702, 3701, 3871, 3838, 3870, 3837, 3979, 4086, 4092, 4142, 4147, 4171, 4174, 3786, 3788, 4216, 527, 3928, 540, 4290, 3977, 3937, 4034, 3976, 3713, 3733, 3815, 4059, 3985, 4109, 3605, 3542, 3568, 3535, 3556, 3598, 3621, 3606, 3675, 3624, 3667, 3615, 3722, 3581, 3684, 3707, 3586, 3662, 3798, 3818]. Book [3847, 3867, 3857, 3829, 3737, 3775, 3792, 3876, 3802, 3850, 3892, 3897, 3765, 4004, 4039, 3956, 4033, 3924, 3997, 4023, 3981, 3916, 4058, 4116, 4067, 4172, 4136, 4207, 4176, 4215, 4233, 4335, 3676, 4145, 4175, 4258, 3519, 3685, 3655, 3688, 3761, 4111, 4185, 4168, 3594, 3590, 3593, 3633, 3680, 3694, 3714, 3746, 3866, 3799, 3951, 4243, 4256, 4267, 2921, 3157, 3395, 3333, 2926, 2756, 3079, 3180, 2819, 3301, 3051, 3203, 2858, 2795, 2490, 2912, 3091, 3010, 3229, 3171, 3320, 3425, 3137, 2853, 3370, 2755, 3257, 2541, 3047, 3181, 3280, 2542, 3323, 3126, 3321, 3317, 3309, 2886, 2746, 3046, 3143, 3233, 3024, 2771]. book [2844, 3050, 3392, 3303, 3273, 3277, 3210, 3345, 2623, 3340, 3423, 3457, 3487, 2682, 2923, 2517, 2685, 2743, 2739, 2803, 2946, 3145, 3031, 3307, 3231, 3234, 3243, 2839, 2972, 3248, 3282, 3272, 2449, 3033, 3012, 2624, 3062, 2725, 3337, 3461, 3255, 2468, 3286, 3236, 3278, 2799, 3087, 2666, 3178, 3082, 3129, 3041, 3215, 3242, 3170, 3260, 2598, 2922, 4123]. Booklet [554]. Books [461, 474, 481, 504, 513, 520, 528, 534, 541, 548, 563, 576, 586, 592, 598, 606, 613, 621, 628, 635, 642, 650, 655, 661, 669, 674, 682, 689, 696, 703, 709, 714, 721, 727, 734, 742, 751, 757, 767, 774, 780, 795, 813, 829, 836, 852, 859, 870, 877, 883, 889, 898, 904, 905, 910, 921, 931, 940, 958, 967, 975, 985, 987, 997, 1005, 1033, 1044, 1052, 1062, 1072, 1078, 1088, 1095, 1103, 1110, 1116, 1126, 1135, 1142, 1148, 1159, 1168, 1174, 1185, 1196, 1201, 1209, 1223]. Books [1231, 1239, 1246, 1253, 1261, 1267, 1275, 1281, 1294, 1300, 1307, 1313, 1322, 1331, 1338, 1346, 1354, 1362, 1369, 1377, 1384, 1390, 1398, 1404, 1411, 1418, 1426, 1433, 1440, 1448, 1456, 1463, 1471, 1478, 1485, 1493, 1500, 1509, 1514, 1520, 1526, 1534, 1540, 1549, 1557, 1568, 1577, 1586, 1593, 1599, 1605, 1620, 1626, 1633, 1639, 1646, 1655, 1663, 1669, 1676, 1683, 1690, 1696, 1702, 1710, 1719, 1725, 1733, 1740, 1747, 1772, 1780, 1785, 1790, 1795, 1803, 1810, 1816, 1822, 1829, 1838, 1846, 1854, 1861, 1869, 1878, 1891, 1895]. Books [1900, 1910, 1920, 1931, 1941, 1946, 1954, 1963, 1971, 1977, 1991, 1997, 2029, 2074, 2075, 2037, 2043, 2073, 2151, 2152, 947, 957, 974, 986, 1018, 1025, 996, 1004, 1017, 1032, 1061, 1087, 4186]. booster [71]. Booth [3203]. Borchers [4053]. Boris [3114, 3460, 3610]. Born [4092, 4355, 965]. Bosanac [3744]. Bosco [2884]. Bose [2704, 2573, 1720, 2277, 4228, 2436, 2580]. Bose-Condensed [2580]. Boson [3403, 3731, 2810, 3390, 3465, 3225, 3128]. bosons [2183, 1301]. Bot [3715]. both [1388]. Botta [2387]. Böttcher [4185]. bottom [2733]. Bouchaud [3002]. Boucher [2372]. Boulevard [2349]. bound [722, 1831]. Boundary [2974]. bouquet [3918]. Bowen [3567]. Bowl [2637]. Bowler [3127]. Box [3713, 2641, 3713]. boy [2002]. Boyarkin [2744, 2744]. Boyd [2310]. Bozhevolnyi [2879]. Braat [4020]. Brading [2535]. Bradt [3435, 2442]. Bragg [605]. Brain [2470, 4241, 3009]. brain-bladder [4241]. Braithwaite [2920]. Brakke [2915]. branched [1430]. Brane [2786]. branes [1830]. brass [1887]. Brázdová [3127]. breakdown [2137, 3646]. breakers [1693]. Breakfasts [2639]. Breaking [4361, 2183, 3354]. breakthroughs [4356, 4344]. Brendon [2636]. Brennan [4343]. Breslin [2411, 2933]. Brézin [2587]. Brian [4307, 3257, 4156, 3438, 3405]. Bridges [3699, 2675, 3853, 3812]. Brief [2989, 2057, 4418, 2951, 1423, 1242, 4180, 3985, 4153]. briefest [3654]. Briggs [2772]. Bright [3135, 2289]. Brillouin [603]. Bringing [3600, 3416]. Brink [4098, 2330, 3337]. Bristol [4241, 4202, 4232, 4140, 4139, 4221, 4333, 4251, 4325, 4303, 4222, 4281, 4353, 4161, 4337, 4201, 4244, 4169, 4170, 4234, 4144, 4341, 4306, 4188, 4260, 4253, 4142, 4147, 4174, 4290, 4136, 4207, 4233]. Britain [808]. British [213, 41, 75, 147, 169]. Brittle [587]. Brizard [3644, 2418]. Broadening [2807]. Brochard [2527, 2596]. Brochard-Wyart [2527, 2596]. Brock [2751]. Broglia [2330]. Broida [1092]. Broken [1569, 2349]. bromide [432]. Bronnikov [3040]. Brosch [2438]. Brotherton [3232]. Brotherton-Ratcliffe [3232]. Brown [1798, 4058]. Brownian [3186, 1796, 2088, 2131, 3584]. Brownlee [4340]. Bruce [2847, 2796, 4142, 4171]. Brumer [3065]. Bruning [3618, 3594]. Bruno [3678, 3213, 3503, 3056, 2453]. Brunton [4019]. Bruß [2297]. Brussels [869]. Bruza [3589]. Bruzual [2626]. Bryan [4286, 4352, 2663, 2507]. Bryant [3275]. Bryson [2506]. Bubble [633]. bubbles [812, 845, 1717]. Buchwald [3748, 3312]. Budapest [3724]. Budhu [3786]. Budker [3492]. bugbears [3929]. Buhmann [2427]. Build [3239]. builders [2474]. Building [3853, 2780, 168, 349, 3776, 264, 3674, 2463, 3524]. Buildings [2675]. built [2384]. Buldyreva [2807]. Bulges [2467]. Bulk [1466, 1472, 531, 4193]. Bundle [3840]. Bunker [2680]. Burbidge [2343]. Burbine [3705]. Bureau [555, 2467]. Burguete [3664]. Burke [4040]. burning [4125]. Bursts [2688, 2070, 2813, 3039, 3057, 2929, 4107]. Burt [3035]. Burton [3113, 56]. Busch [3710]. Buscher [3697]. Buschow [2830]. Busemeyer [3589]. Business [2657, 4057, 56]. Butikov [3476]. Butt [4295]. Buttar [3054]. Buzug [2465]. Byrne [2508]. Byron [4148].

C [4060, 3025, 2433, 2463, 2756, 2374, 2486, 2593, 2949, 2966, 3716, 2356, 3056, 3412, 2347, 3738, 2057, 2255, 2301, 2561, 4026, 4072, 2191, 3808, 3755, 4269, 4270, 2567, 3054, 3323, 4024, 3057, 2993, 3437, 3708, 3627, 4048, 2300, 2430, 2623, 3100, 2379, 2461, 2525, 3263, 3823, 3525, 2692, 2881, 2773, 3770, 2803, 2839, 2651, 4354, 3735, 4053, 2436, 2380, 2846, 2440, 2419, 2315, 3536, 3674, 2532, 3555, 3528, 2501, 3286, 2947, 3236, 2298, 4109, 2454, 2453, 2507, 2663, 2721, 3153, 3684, 3775, 3688]. C. [1461]. C.H [648]. Cacialli [2795]. cackled [2772]. Cademartiri [2529]. cages [3467]. Cahill [4077, 3170]. Cahn [4332]. Cai [2791, 3118]. Cairo [3063]. Calculation [2566]. Calculations [2670, 3439, 4278, 4377]. calculators [807, 862]. Calculus [2754, 3256, 3347, 3485, 3922]. Caldeira [3303]. Calder [865]. Caldwell [3668]. Callaghan [3373]. Calle [2509, 2493]. Called [2403]. Calogero [3339]. calorimetry [227, 3818]. Calvin [4258]. Cambridge [4078, 4060, 4211, 4286, 662, 4135, 4032, 4009, 4210, 4035, 3425, 4005, 4045, 4182, 4177, 4307, 4061, 4143, 4084, 4248, 112, 4026, 4072, 4052, 4006, 4021, 4081, 4130, 4309, 4002, 4001, 4046, 4141, 4037, 4040, 3994, 4048, 4042, 4003, 4015, 4105, 4146, 4345, 4342, 4346, 4322, 4165, 3421, 4107, 4157, 2505, 3653, 4093, 1569, 4019, 4326, 2295, 4122, 4197, 3066, 4020, 4036, 4054, 4053, 4079, 4352, 4071, 4129, 4219, 4350, 4066, 4077, 4166, 4065, 4063, 4092, 4034, 3129, 4039, 4058, 4215, 4335, 4185, 4168, 255, 1045, 1053, 177, 1343, 2372]. Cameron [4142, 4171, 2839]. Camley [4009]. Campanelli [4161]. Campbell [3343, 4058]. Campbell-Brown [4058]. Can [1506, 4367, 4313, 1128, 2549, 3690, 530, 3839]. Canadell [2966]. Cancer [2991, 3432, 4300, 3603, 3990, 3596]. Canio [4147]. Canonical [2716]. Canseliet [3434]. Cantilevers [3331]. Cantor [4156]. capacity [1049]. Cape [3102]. capillarity [3573]. Cappelli [3004]. Capri [2827]. capture [977]. Car [2809]. Carballido [3678]. Carballido-Landeira [3678]. Carbon [3051, 2902, 2332, 3033, 3898, 3544, 1020, 4340, 2838, 3889]. Cardiff [75]. cardiovascular [1849]. Cardona [3189, 2792]. career [2921, 4120, 4078, 4060, 4106, 2686, 2883, 3157, 3395, 3052, 3350, 2989, 3061, 3098, 3201, 2667, 4241, 4275, 3022, 3180, 3506, 2819, 2648, 2911, 2840, 3107, 2593, 3298, 3365, 2830, 2715, 3124, 2544, 3376, 4231, 2795, 2955, 2865, 3116, 4154, 3349, 3137, 4113, 4356, 4328, 2866, 3038, 3095, 3035, 3247, 4118, 3296, 2939, 2836, 2853, 2785, 2988, 3194, 3179, 3353, 3408, 3499, 4240, 4272, 2617, 2775, 3207, 3370, 2621, 2907, 3168, 3245, 2791, 2841, 2999, 4072, 4052, 4178, 4190, 4198, 2965, 3279, 3501, 4235, 4257, 4309, 2693, 4323, 2660, 2800, 2954, 3054, 3407, 2887, 3126, 3205, 4085, 4108]. career [4068, 4119, 4196, 3184, 3007, 3081, 3074, 4339, 3099, 4187, 4294, 2538, 2820, 3046, 3431, 3143, 3233, 3366, 2670, 2771, 3050, 4165, 3303, 3398, 3492, 2716, 3128, 3274, 3382, 3080, 3005, 4157, 3230, 3210, 3446, 4186, 2539, 3340, 3423, 2526, 3026, 3003, 3424, 3189, 3511, 2680, 3015, 3162, 3447, 2929, 3367, 3232, 3307, 3449, 3065, 2622, 3430, 2839, 3272, 4121, 2784, 2712, 3127, 3033, 3012, 4122, 4349, 2815, 3094, 2780, 2624, 2835, 2874, 2924, 2837, 2834, 2760, 2750, 2880, 3062, 3119, 3409, 4054, 4083, 4189, 3347, 2546, 3414, 3476, 3489, 4255, 4219, 4301, 4310, 4276, 4350]. career [4166, 4167, 2759, 3138, 3014, 3346, 2725, 2557, 4199, 4297, 4334, 2896, 3228, 3461, 4253, 4265, 4266, 4332, 3148, 3146, 3141, 3404, 3391, 3410, 3440, 3482, 3441, 3481, 3483, 3439, 3504, 3503, 4063, 4220, 4174, 2758, 3191, 3185, 3278, 3508, 2799, 2807, 2826, 2654, 2665, 2666, 2832, 3306, 3495, 4058, 4215, 3397, 4185, 4168, 2920, 2882, 2967, 2919, 2586, 2769]. cares [3731]. Carl [2502, 2859, 1968]. Carlile [3236]. Carlo [3927, 4182, 3557, 4146, 2860]. Carlos [2509, 4119, 2493]. Carlson [3149]. Carol [3023]. Carr [2819, 2459]. Carreira [3300]. carried [1560]. Carroll [4036]. carry [4313]. carrying [126, 407]. Carsten [2965]. Cartesian [3399]. Cartier [2301]. Cartography [3044, 4091]. Cartoon [4258]. Cascade [3146]. Case [4113, 4369, 2333, 2512, 999, 1008, 3459]. Casimir [1040, 2164, 1579]. Casoli [2476]. Casquilho [3507]. Cassidy [2906]. Castañeda [3014]. Castellani [3004, 2535]. Castiglione [2351]. Castro [3557]. Cat [3056, 2261, 3727, 1235, 2262, 3600]. Cataclysmic [2180]. Catalin [3500]. catalysed [1479]. Catling [3263]. Causal [2153]. causality [3220, 2208, 3684, 2545]. Causation [4367, 3690]. Cause [3354, 3247]. caused [1703]. Causes [2721]. Cautionary [2548]. cave [2026]. Cavendish [380, 3612, 211, 1045, 1053, 765, 177]. Cavity [4212, 2381, 1606]. Cavity-based [2381]. Cavity-QED [4212]. Cayman [4150]. CDM [3222]. Cecchi [3648]. Cecotti [3511]. Celebrating [2599]. celebration [3131]. celebrations [64]. Celestial [3044, 3000, 2832, 3685]. cell [1363, 3005]. cells [891, 4252, 35, 514, 2259, 1199, 2260, 2891, 2834, 3028]. cellular [4179]. Cencini [4208]. census [3781]. centenary [3131, 946, 2934]. centennial [3508]. Centre [3102, 4388, 800, 275]. centred [2117]. centric [3160]. Century [3349, 3168, 3369, 3121, 2682, 3468, 1137, 3828, 3945, 23, 2548, 1902, 4186, 235, 4200, 1097, 2258, 1464, 2887, 2460, 2449, 3618, 2257, 4347, 3753, 4042, 4234, 3909, 3989]. Cepa [2779, 2779]. Cepa-Nogué [2779, 2779]. Cepheid [2732]. ceramic [505, 2405, 3319]. ceramics [4323]. Cerenkov [134]. CERN [680, 1495, 3389, 3731]. Certain [3414, 1705]. Cetto [1844]. Ceulemans [3758]. CFT [3592, 2893]. CH [4154, 4191]. Chabert [2920]. Chadwick [648, 749]. Chaichian [3240]. chain [2184]. chains [1879]. Chakrabarti [3539, 4350]. Chakraborti [4350]. Challenge [2917, 961, 1812, 4192, 3194]. Challenges [3267, 3989, 2294, 1538, 3618]. Cham [4110, 4154, 4191, 4118, 4155, 4031, 4047, 4115, 4160, 4321, 4145]. Chamber [633, 192]. chambers [230]. Champagne [3196]. Chance [2957]. Chandra [3423]. Chandrasekhar [2934, 2251, 1264]. Chandre [3167]. Chang [3412, 2866, 2787, 3011, 3412, 2567, 3046, 3720, 3030, 3693, 3536]. change [3333, 4103, 2204, 653, 165, 2367]. changes [1703, 290, 3372]. channel [4231]. channelling [478]. channels [1841, 3823]. Chao [3371, 3208, 3484]. Chaos [1618, 1664, 1546, 3471, 3638, 1391, 2016, 3235, 2416, 3996, 2351, 2972, 4004]. chaotic [2060, 3749]. Chaparro [4170]. Chaparro-Romo [4170]. Chapelle [3408]. Chaplin [4007]. Chaplot [2712]. Chapters [3143]. characterisation [3686, 3070, 3809]. characterising [4189]. Characteristics [2776, 4293]. Characterization [2614, 4203, 4180, 3015]. Charap [2848]. Charbonneau [3803]. Charge [1458, 1283, 2711, 1080, 637, 61, 402, 439]. Charge-density [1283]. Charged [552, 2388, 2569, 2480, 1542, 1948]. charging [303]. Charitos [4347]. Charles [4149, 4296, 3383, 1092, 2597, 2632, 785]. Charlier [4081, 3265]. Charlot [2626]. Charlotte [3434]. Chaudhuri [2750]. Chaumont [3138]. Chazelas [3434]. Checkerboards [3214]. Chekhova [2746]. Chelikowsky [4022]. Chemical [2433, 2853, 1112, 2806, 3407, 108, 242, 834, 3629, 3993, 4090, 3432, 4189, 2060]. chemist [1649]. Chemistry [3309, 3238, 2319, 249, 1580, 4118, 3889, 223, 1163, 4295, 1171, 3432, 685, 3937, 1659, 3310, 4116, 3930]. chemists [3974, 4097, 2487]. Chen [2836, 3629, 4250, 2355, 3629]. Cheng [3499, 3071]. Cheng-Chung [3499]. Chennupatti [2775]. cheque [1984]. Chern [3948]. Cherry [3457]. Chervon [4044]. Chiang [3299]. Chiang/translated [3299]. Chiao [3522]. Chiara [4242]. Chicago [2191]. Chichester [4272]. Chien [3299]. Chien-Shiung [3299]. Childs [2941]. China [4419, 187, 3441].

Chiou [3246]. chiral [4164, 4226]. Chirality [2503]. Chityala [3351]. Cho [4353, 2555]. Chorus [4359, 3448]. Chou [3371]. Choudhuri [2594, 3777]. Choudhury [2712]. Chris [4079, 3605]. Christensen [4015, 3384, 3001]. Christian [4241, 3320, 3659, 3708, 3660, 3340, 3490, 3482, 3217]. Christmas [253, 186]. Christof [4034]. Christophe [4081, 3265, 4004]. Christopher [2883, 2463, 3491, 3056, 3494, 3323, 3564, 3210, 2691, 2468, 1251]. Chromey [3643]. chromism [1370]. Chromodynamics [2609, 2822, 1066, 3825, 1394]. Chu [3066]. Chua [2932]. Chuan [3659]. Chuang [3649, 2656]. Chuck [2809]. Chumakov [2398]. Chun [3109]. Chung [3499]. Churchill [15, 3226]. Cimatti [4146]. Cinette [313]. Cipelletti [3002]. Circuits [3198, 2604, 3024, 2879]. Circumstellar [3309]. cities [3707]. civil [358, 397, 101]. Civilisation [2789]. Clarendon [1459]. Clark [3054, 2632]. class [211, 397, 2603]. Classic [3026]. Classical [2686, 2781, 1265, 4184, 4190, 3366, 3398, 1807, 2807, 511, 3470, 3750, 3971, 3944, 2320, 3585, 3386, 2127, 3895, 4187, 3419, 3331, 3906, 3417, 3736, 3864, 4129, 3887, 1262, 1892, 4176, 2048, 2648, 4035, 2999, 4002, 2726, 2692, 4161, 2391, 4351, 111, 2705, 2978, 2952]. Classification [2468]. Claude [2921, 2686, 3772, 2954, 3191]. Claudia [3611]. Claudio [3120, 4115]. Claus [2939, 2288]. clay [1766]. Clays [2614]. Clean [138]. cleaning [2090]. Clear [3730, 994]. clearing [16]. clearly [1806]. Clegg [3405]. Clerk [3281]. Cliff [4347]. Clifford [3840]. Clifton [3757]. Climate [2282, 2367, 3333, 2204, 1325, 1595, 1703, 1987, 2890, 3713, 4039]. Climates [2993]. Climatic [906]. Clinical [2842, 2876]. cloaking [3130]. clocks [38, 2327, 4074, 3966, 2060]. Close [2366, 4025, 2764, 2392, 2612, 3128, 2877, 2426]. Closures [3376]. Cloud [2663, 192, 1117, 1355]. Cloude [2584]. Clouds [2916, 3860, 3993, 1987, 3495]. clue [59]. Clustering [1879, 1465, 2024]. Clusters [1487, 4362, 2749, 500, 2780, 1641, 1070, 1791, 1036, 1649, 4022]. CMS [2647]. Co [4106, 4268, 1249, 4022, 740, 3794]. co-existence [3794]. Co-operative [1249, 740]. Coarse [2351]. Coastal [2484, 3656]. COBE [1628, 1652]. Cockcroft [1221]. Code [4361, 2562, 2228, 1693]. Code-Breaking [4361]. Codebreaking [2772]. Codes [209]. coding [3504]. Coevolution [2651]. coexist [1128]. Coey [4046, 2494]. cognition [3589]. Cohen [2921, 3718]. Cohen-Tannoudji [2921]. Coherence [2309]. coherences [3386]. Coherent [2198, 2797, 3601, 2869, 4138]. coherently [2179]. Cohomology [2444]. Coiffier [2956]. coil [126, 407]. Cold [3161, 2131, 4118, 1651, 2287, 4119, 4049, 2231, 1039, 4102]. Cole [3234, 2637]. Coleman [3911]. Coles [4015]. Coletta [4031]. Colin [3270, 4260]. collaboration [2050, 2064, 4, 719, 4161]. collapse [1415, 998]. collapsing [3477]. Collected [3079, 291, 380, 3432, 4413]. Collection [3114, 2350, 4154, 4221, 3448, 4270, 4187, 3432, 4224, 4172]. collective [1956, 3944]. College [54, 96, 53, 180, 221, 52, 106, 15]. colleges [79]. Collider [3403, 2542, 3731, 3594, 2012, 4138, 1064, 1495, 2183, 3465, 2471]. Colliders [2662, 2322]. colliding [1112]. Collier [3198]. collinear [4278].

Collins [2831]. collision [578, 711, 1677]. Collisional [2807]. collisions [1458, 1452, 623, 637, 653, 704, 1069, 623, 637, 653, 4168]. Colloidal [1144, 678]. Colloids [3002, 2443]. Colloquia [2467, 2737]. Colombo [3033]. Colomo [3004]. colonial [1782, 2002]. colonies [1622, 1782]. Colonna [4026]. Colosi [3765]. colossal [2209]. Colour [1370, 1227, 908, 3232, 139, 163, 1641, 885, 601, 37, 2289, 3575, 3624]. Colvin [3242]. Combinatorial [2685, 2743, 2739]. Combinatorics [2793]. Combined [834]. come [2305, 578]. comes [1762, 1631]. cometary [1510, 1535, 1644]. Comets [1113, 1640, 2996, 2353, 3355]. comings [1132]. Comment [1359]. Commentaries [3126]. commentary [3473]. Comments [2099, 2781]. commercialization [4252]. Commission [869, 1019]. commissioning [1495, 1054]. committee [30, 74, 40, 177]. Common [3247, 2257, 1097, 2258, 1883, 3645]. Commonly [4269]. Communicating [1915]. Communication [288, 3806, 2184, 3504, 3664, 254, 2985]. Communications [3362, 2486, 337, 1188, 4060]. Communist [187]. Communities [869]. Compact [2441]. compaction [744]. Companies [2657]. companion [3988, 4375, 3528]. Companions [3427, 4392, 4393, 3668, 4114]. company [2057, 1343]. Comparative [1157]. compendium [2422]. Compensated [2340]. Competitive [4086]. compilation [701]. Complementary [2770]. complete [1297, 3679, 3870]. completely [1677]. Complex [2113, 2352, 2766, 2657, 2745, 3380, 1782, 3747, 2730, 3787, 1192, 4017, 2208, 3665, 4249, 4350, 3202, 2652, 3814, 3229]. Complex-Valued [2766]. Complexity [2124, 3354, 1871, 2972, 4311, 4276, 1975, 3803, 2840, 3693, 4175, 3229, 4099]. Composite [950, 1481]. Composites [2817]. Composition [2826, 2188, 3990]. Compositional [3157]. compositions [4114]. Compound [3013, 2932]. compounds [1333, 3889, 1166, 1292]. Comprehensive [3375, 2905, 3851, 3975, 2372, 2348]. Compressible [2841]. compression [1114, 3919]. compressive [3673]. Compton [881, 739]. computable [2033]. Computation [2757, 3438, 2985, 2592, 2400, 2656, 1934, 1872, 3975, 4382, 2499, 3950, 2528, 2371, 2499, 3785]. Computational [2777, 2794, 2616, 2432, 2997, 3741, 2061, 2749, 3883, 3950, 2487, 3582, 4120, 2749, 3118, 3181, 3297, 3717, 3140, 3271, 2332, 2943, 4185]. Computed [2465]. Computer [3023, 1730, 2625, 711, 3127, 1684, 2317, 1091, 4217, 942, 2062, 2517, 2653, 4282, 1704, 2649, 3239, 2457, 3033, 3842]. Computer-Based [3033]. Computerized [1775]. computers [1750, 1365, 1371, 2306, 56]. Computing [2033, 3027, 2512, 2317, 2399, 1966, 2455, 3545, 4216, 4255, 4267]. concentrative [1249]. Concept [2856, 342, 710, 2252, 1143, 2251, 1621, 2661]. conception [732]. Concepts [3089, 2965, 3240, 2874, 3064, 2579, 4191, 3895, 3994, 1529, 4093, 3733, 2456, 139, 3109, 2696, 2529, 2348, 2739, 2511, 3847]. Conceptual [1661, 3791, 875]. Concise [2346, 3296, 2848, 3806, 4317, 3955, 3847]. Conclusion [1888]. condensates [2573, 2277, 4228]. Condensation [2436, 1720, 2516, 2704].

Condensed [472, 3124, 3279, 2850, 2616, 2472, 2432, 2389, 2349, 2401, 1137, 1402, 1487, 1862, 3491, 3873, 2961, 3557, 1145, 4001, 3060, 1960, 2487, 3718, 3879, 3880, 3836, 4166, 3505, 1414, 2580]. Condensed-Matter [2472]. condition [1745, 3851]. Conditional [3515]. Conditioning [3138]. conditions [1207, 530, 2340, 3242]. Condon [3655]. conductance [1728]. Conducting [1268, 1269, 4206, 1449]. Conduction [607, 506, 3501, 515, 439]. conductivity [677, 1049, 425, 542, 378, 537, 333, 1289]. conductor [558, 240, 3230]. conductors [918, 903]. Cone [2465]. Cone-Beam [2465]. Conference [143, 266, 3133, 3412, 156, 144, 3063, 137, 65, 79, 78, 554, 2727, 3261, 2984, 3075, 3054, 193, 2932, 2898, 157, 338, 3189, 4159, 2884, 415, 2651, 3012, 2780, 2760, 2750, 2441, 2822, 3039, 3067, 3267, 3348, 554, 51]. conferences [1070]. confined [1606]. Confinement [2646, 2928, 1825, 4164]. Conflicting [1836]. Conformal [2622]. Conformational [1449]. confusing [1989]. Congrès [366]. congress [464]. Conjecture [4299]. Conjugated [3116, 1370]. Conlon [3523]. connected [3456, 4398]. Connections [4104]. Connolly [3577]. Consales [3740]. Consciousness [1630, 2796]. consequences [23, 3689]. Conservation [3402, 2758, 414]. conservatives [1445]. considerations [3159]. Consoli [3997]. Constancy [1175]. Constant [4076, 2095, 632, 469, 712, 2254, 2253]. Constantine [2997]. Constants [3083, 632, 2038, 3292, 2087, 311]. Constitution [2335]. construct [2202]. construction [1495]. Consultants [555]. contact [114]. containing [991, 3889]. Conte [2489, 3052]. Contemporary [110, 2238, 303, 298, 306, 335, 1766, 95, 128, 55, 80, 269, 936, 409, 630, 2605, 185, 1014, 3327, 4153]. Context [4367, 1643, 3690]. Continental [4370, 3078]. Continua [2702]. Continuous [3431, 3233, 2971, 4376, 2176]. Continuum [3205, 2365, 3947, 4061, 4178, 2874, 3096]. contraction [1615]. contrast [2362]. Contreras [3189]. Contributions [2599]. Control [3374, 3065, 2782, 2235, 1503, 2016, 3434, 4405, 2530, 4019]. Controlled [2045, 1898, 193, 719, 853, 4116, 3578]. controls [2086]. Controversy [4370, 3363, 3078]. Contucci [3269, 3735]. convection [1234, 3413]. Convective [1422, 1427, 3211]. conversations [3834]. conversazione [214, 276, 350]. conversion [3822, 806]. converter [791]. convolution [745]. Cook [4069, 1083]. cookbook [4267]. cool [1378, 2441]. cooled [1720]. Coolen [3789]. cooling [723, 1536, 3852, 1864]. Cooper [1073]. Cooperative [1782]. Coopersmith [2661, 3762]. Copeland [3111]. Copenhagen [3513]. Copernican [4037]. Copernicus [3651]. Coplan [2463]. copy [3778]. Corbally [2468]. core [2188, 1290, 3722]. Cormack [4334]. Cornelius [3119, 4308]. Corni [4321]. Coronado^{T M} [2302]. Coronae [3007]. Coronal [3007]. correct [1989]. Correction [4011, 4012, 1775, 3991, 4216, 3098]. Correlated [3029]. correlation [4169]. correlations [1712, 1691]. Correspondence [2765, 2437, 126, 2481, 3592]. Corrigendum [3324, 3325, 3385]. Corso [86]. Cortez [3507]. Cosmic [617, 1799, 3168, 3958, 3268, 2597, 2143, 156, 2292, 3423, 1075, 2000, 4410, 1659, 22, 3965, 4149, 4082, 4135, 1628, 1731, 2141, 1656, 4355, 3650, 3477, 3861, 1811, 4157, 3907, 1768, 415, 716, 1995, 345, 2447, 3203, 2870, 4143, 3845, 3669, 2379, 2881, 4065, 3667, 2456, 2964, 4243]. Cosmic-ray [2143]. Cosmochemistry [2613]. Cosmogenesis [4384, 1562]. cosmogeny [1408]. cosmogonical [1466]. Cosmological [2095, 4076, 2815, 3859, 1550, 2343]. Cosmology [3040, 4375, 4420, 4421, 2716, 929, 2599, 919, 1555, 3784, 4196, 3797, 4339, 2519, 1617, 2435, 3390, 2126, 4121, 998, 2215, 4176, 4145, 3593, 2658, 3168, 3047, 3074, 3436, 2519, 3093, 2343, 2539, 4205, 2903, 2611, 2741, 2591, 2922, 3723, 4209, 3502, 3620, 4044, 3816, 1100, 3691, 4200, 3854]. Cosmos [4374, 3455, 3240, 3215, 4133, 4387, 4109, 1823, 3261, 3378, 3695, 3579, 4155]. cost [1516]. costs [2063]. Cotton [2663]. Cottrell [3805]. Coulomb [1657, 3294]. Council [391, 437, 3206]. Count [1937]. counter [1678]. counter-intuitive [1678]. counting [2892]. coupled [3820]. coupling [3659, 4119, 3202, 2209]. Coupon [2521]. Course [2790, 2365, 3108, 3304, 159, 206, 3974, 2490, 3692, 3488, 4072, 52, 4095, 4265, 4332, 3605, 74, 2875, 2559, 3679, 4294]. Courses [154, 342, 40, 248, 2320]. cousins [1235, 2262, 2261]. covalent [901]. Covariant [2848]. cover [4353]. covering [2320]. Cowie [2319]. CP [3061, 1978, 2586]. Cracking [2562]. Cracow [499]. Crafting [2585]. cranks [1445]. crashes [2047]. Craters [2672, 4031]. CRC [4317, 4217, 4284, 4271, 4259, 4269, 4270, 4283, 4318, 4319, 4302, 4094, 4258, 4267]. Crease [4303]. creates [4149]. Creation [4394, 1597, 3857, 3844]. Creativity [2679, 941]. Creator [4372, 3453]. Creep [428]. Cremona [2532]. Crescitelli [3740]. Cressman [4294]. Creutz [4164]. crime [823]. Cripps [3528]. crisis [18]. Cristian [3269]. criteria [1778]. Critical [2715, 1242, 2983, 2729, 521, 1903, 3516, 1703, 1912, 3929, 4147, 1737]. Criticality [3070]. Critique [2412]. cross [240, 3782]. cross-disciplinary [3782]. cross-section [240]. Crossed [816]. Crossing [4010]. Crossover [3417]. crowd [4109]. crud [1243]. Crusts [2826]. Cryogenic [1006]. cryogenics [411]. Cryptanalysis [2568]. cryptography [1692, 1693]. Crystal [3154, 4365, 2341, 800, 3930, 4379, 346, 3310, 2892, 1441, 994, 3900, 1163, 3626, 4198, 1503, 1894, 3424]. crystalline [381, 4247, 1254, 1831, 506, 3255, 1298, 4333, 4337]. crystallization [999, 1008]. crystallographer [4253]. Crystallographic [3930, 3310, 3131]. crystallography [3711, 630, 3494, 48, 583, 2769, 4083, 3599]. Crystals [1457, 3344, 2300, 2102, 790, 677, 1405, 1870, 1435, 546, 121, 1096, 2256, 908, 1332, 432, 1570, 1271, 1144, 1099, 550, 830, 478, 3294, 347, 2668, 2255, 3254]. CT [2465]. Cubero [3584]. Cugliandolo [2846, 3772, 3374, 2760]. Cui [3113]. Cullerne [2358]. Cultural [2803, 2211]. Culture [2884, 3945]. CUP [4352]. Curiosity [2849]. Curmudgeon [2886]. Current [2807, 5, 752, 126, 407, 3499, 2129, 1475, 1466, 447, 240, 3174, 139, 3184, 3827, 2628]. current-carrying [126, 407]. Current-Driven [2628]. currents [640, 398, 1073]. Curriculum [245, 479]. Curry [3495]. Curved [2648, 2524, 1278, 1282, 4141, 625]. Cusano [3179, 3740]. Cushman [2976].

Cushman-Roisin [2976]. Cutolo [3179]. cycle [3550, 1863, 4289, 3777]. cycles [3456, 4398]. cyclones [1091]. Cyclotron [47]. cylinders [4206]. Cyron [4241]. Czysz [2453].

D [3806, 3817, 4120, 4106, 2312, 3368, 1121, 4317, 4062, 2786, 2297, 2374, 440, 3785, 3699, 2385, 4140, 2334, 3656, 2311, 2471, 2912, 3171, 3383, 4069, 4184, 3744, 3322, 3628, 2696, 2775, 2494, 3620, 3793, 4072, 4190, 3753, 4021, 4046, 2737, 3769, 4003, 2597, 2330, 3661, 2318, 2765, 3254, 3810, 2524, 4097, 2379, 2525, 3457, 2692, 2899, 2427, 2803, 3145, 3307, 3691, 3589, 2689, 2985, 3316, 3500, 3127, 3094, 2633, 3607, 3683, 2676, 3183, 3704, 3526, 2419, 3471, 2793, 1562, 2259, 2683, 3087, 3032, 2470, 2559, 2934, 3021, 4039, 4033, 4023, 4058, 4176, 4215, 2456, 2041, 2779, 3328]. D [3432, 3666, 2223, 4413]. D-Brane [2786]. D.F [4066]. D.J [4066]. D.S.I.R. [29, 99]. Dae [3493]. Daglis [3787]. Daily [4052, 3490]. Dalarsson [2740]. Dalby [4097]. Dameris [3520]. Dams [2675]. Dan [4339, 2647, 3314, 3621]. Dance [2417]. Daniel [3583, 4232, 4268, 2897, 2944, 4105, 2771, 3186, 3285, 2632, 2768]. Daniele [2431]. Daniels [2354]. Danner [2932]. Danon [3076]. Dario [3659]. Dark [1637, 4076, 2714, 2664, 2576, 3197, 2630, 1392, 2450, 3670, 3464, 3547, 1992, 4160, 2046, 3956, 3268, 3136, 3542]. Darken [2843]. darkness [2570]. Darrigol [2887, 3798]. Darrin [2590]. Darwin [1343, 2534]. Dass [3526]. Data [2727, 2766, 3093, 2623, 2354, 3188, 701, 3577, 1896, 4007, 4207, 3259, 4078, 3577, 2901, 3808, 4019, 4350]. Data-driven [4019]. dating [1020]. Datta [2556]. Dattagupta [2403]. David [2921, 3115, 4202, 2911, 4273, 3376, 3133, 3195, 3579, 3055, 2775, 4285, 3343, 3451, 4006, 3438, 3479, 2816, 2548, 2668, 2746, 2690, 3169, 3469, 3121, 3005, 3263, 1057, 3232, 3532, 3697, 3584, 3471, 4332, 3505, 2503, 3260, 3372, 4058, 4207, 2598, 2922, 3355, 2919]. Davide [4242, 4208]. Davidson [3143]. Davies [3425, 2845, 3684]. Davis [2463, 3323, 3255]. Davy [331]. Dawid [3219]. Dawn [3322, 4213, 4355]. Dawood [3098]. Dawson [2322]. day [53, 4146, 278]. Days [2772, 4361, 1160, 89, 195, 242, 256, 262, 265, 277, 365, 334, 406, 662]. Dead [3056]. Dear [3574]. Death [3152, 2657, 3039]. Debate [2778, 2558, 4328, 4348, 747]. decade [1217, 1225, 707, 2626]. Decades [3020]. Decay [1149, 2606, 1155, 1958, 1650, 2115]. decays [1978]. December [1136]. decision [3589]. decisive [1156]. Decoding [2495]. Deconstructing [3816]. dedicated [2787]. Deen [3141]. Deep [4392, 4393, 3355, 1926, 4345, 1069, 3427, 3668]. Deep-Sky [4392, 4393, 3427, 3668]. Defect [2068]. Defects [1805, 464, 1449, 970, 1823, 4198]. defence [101]. defining [4102]. definition [2892, 13]. deformation [3544]. deformations [4231]. degree [2412]. Deji [2838]. Dekker [4224, 3044]. Délery [2841]. Delft [3994]. delight [2157, 1042]. delivered [50]. delocalised [4280]. Delsanto [4145]. delta [1588]. delta-doped [1588]. demon [3514, 1698, 2219]. demonstrating [3389]. Demonstration [151, 260, 861]. Demonstrations [3458, 236, 4383, 909, 285]. demystified [3645]. Demystifying [3336].

Denizli [4290]. Denny [2851, 2843, 2675]. dense [1291]. Density [2965, 3450, 2539, 141, 3853, 1283, 4373, 3824, 3770, 4273]. Density-Functional [2965]. Denstone [52]. Department [53]. departments [270, 79]. departure [554]. Dependability [2670]. Dependent [2965]. depolarization [653]. depths [1228, 1373]. Derek [3411]. Dereniak [2385, 2385]. Dereziński [3217]. Derived [3071]. Dermer [2597]. Dermott [4305]. Desai [2461, 2490]. Desbiolles [3487]. Desch [4150]. Description [2979]. Design [2757, 2836, 3159, 2660, 2797, 3362, 3300, 869, 54, 5, 4252, 4242, 1495, 1154, 444, 1713, 1054, 2493, 3605]. Designed [2753]. designer [420]. Designing [950, 2463]. Desorption [1244]. D’Espagnat [875]. destructive [2515]. detail [4201]. Detailed [631]. detect [3464]. detected [818]. Detecting [1523, 2194]. Detection [2076, 2476, 2839, 2032, 284, 2046, 4288, 823, 2224, 2954, 2387, 4290, 3792]. detector [3389]. Detectors [3204, 3013, 937, 273, 453, 1542, 2294, 1760, 1752, 4158, 4152, 2169, 3893]. determination [430, 469, 911, 141, 2526]. determining [1677]. Determinism [3235, 3220]. Deterministic [2016]. Detlef [4154]. Detonation [2393]. Development [4419, 2854, 1248, 3441, 802, 1194, 324, 1045, 1053, 2241, 707, 2109, 245, 1752, 1316, 322, 330, 2242, 1748, 792]. Developments [854, 1340, 1405, 1870, 630, 4314, 1881, 838, 537, 638, 12, 333, 1713, 4098]. Devenish [3613]. Device [2838, 3279]. Devices [2660, 2538, 2932, 652, 2320, 2567, 1635, 3940, 543, 2048, 3174, 3101, 2567, 3141]. Deville [2794]. Devoret [3374]. Dewar [394]. Dewhurst [3700]. DeWitt [2301]. DeWitt-Morette [2301]. Dexheimer [2386]. Deza [3132]. Dfl [662]. Dharma [3167]. Dharma-wardana [3167]. Dharmadasa [3028]. Dharmaraja [3014]. Diagnosing [456]. Diagnosis [850, 1048, 886]. diagnostic [1775]. diagnostics [589, 3990, 2991]. diagram [1964, 425]. Diagrammatic [3157]. dialogues [3834]. diameters [609]. Diamond [2818, 11, 34, 93, 1363]. diamond-anvil [1363]. Diamonds [799]. Dianne [2512]. Diaspro [3795, 3608, 2624]. diatomic [4110]. Dice [2933, 4364, 1844]. Dick [3676]. Dickerscheid [2525]. Dictionary [2871, 3721]. Did [2336, 662, 1101, 2748]. Didn’t [2312]. Diederik [3320]. Dielectric [2486, 1145]. dielectrics [832]. Diep [3209]. difference [407]. Differences [2345]. Different [2850, 2886]. Differentiability [3554]. Differentiable [2545]. Differential [2414, 2457, 2543, 2653, 3256, 3264, 542, 3480]. differentiation [3689]. difficulty [286]. diffracting [3370]. Diffraction [287, 3050, 2121, 471, 60, 246, 629, 938, 1295, 1759, 686, 212, 455, 1736, 3494]. Diffuse [3058]. Diffusion [4379, 1468, 1912, 1298, 817, 979, 3640, 3866, 3186, 3866]. Diffusion-limited [1912]. Diffusive [1622, 3413]. Digital [2022, 1316, 3232, 1188, 3559, 4144, 3101]. Dilute [777, 3103, 2436]. Dimakis [2842]. dimension [1503, 3433]. Dimensional [4385, 2622, 2316, 3868, 615, 3550, 3781, 3793, 3358, 1993, 1292, 4222, 2555].

Dimensions [3079, 4363, 3877, 1552, 2021, 1429, 3760, 60, 1199, 2260, 3040, 2902, 2622, 4210, 4121, 2259, 2895]. Dimitri [4029]. Dimitry [2815]. Dimo [2715]. Dimock [2687]. Dimoulas [3666]. Dina [2621]. Dine [2893, 3783]. Ding [4252]. Dinh [3440]. Diode [1528, 1270, 1748]. Diode-pumped [1528]. Diodes [2909]. diploma [154]. dipole [126, 664, 1948]. Diptiman [3539]. diquarks [2159]. Dirac [3840, 2350, 1569, 2953, 1933, 694, 2336, 3851, 2103]. Direct [1206, 2046]. Directions [2926, 2572]. Dirk [2615, 2823, 3144, 2937]. disappearance [2359]. disaster [763]. discharge [1046]. discharges [2137]. disciplinary [3782]. disciplines [2842, 2352]. discontinuous [515]. discord [2206]. discotic [1870]. discourse [50, 3157]. discovered [2735]. discoveries [4028, 2294, 1265, 2321]. Discovering [4101, 4322, 3676, 4256]. Discovery [3349, 2986, 1301, 3468, 4135, 2257, 1773, 2574, 3821, 1097, 2258, 3465, 1608, 4065, 2053, 247]. Discrete [4376, 4381, 3421, 2153, 3061, 3431, 2517]. discrimination [1935]. Discs [2441]. Discussing [2864]. discussion [261]. dislocation [1290]. dislocations [3826, 287, 1211, 3147]. Disorder [3139, 2007, 581]. disordered [3735, 2942]. Dispatches [2519]. Dispersion [3218]. Dispersive [2659, 1489]. display [1099]. disposal [1501]. Dissertori [4347]. dissipation [3639, 3303]. Dissipative [3162, 1140]. dissymmetry [470]. distance [1839]. Distances [1304, 840, 1731]. distant [2013, 1714, 262, 4399, 3876]. distributed [3745]. Distributing [4126]. distribution [368, 3517, 1776, 382]. distributions [1080, 2142]. disturbances [574]. dit [4313]. dited [4058]. Ditlevsen [2689]. Ditton [334]. Divakaran [3539]. diversions [3828]. diversity [4150]. Divide [3056]. division [264]. Djordjevic [4216]. Dmitry [3492]. DNA [2108, 2101, 2077, 2053, 1544, 3606]. do [3698, 4005, 4154, 3034]. Dobrosavljevic [3230]. Docktor [4225]. docs [3057, 2993, 3016]. documentary [2367]. documents [4113]. Dodd [4211]. Dodelson [3709]. dodging [1858]. D’Odorico [2833]. Does [3383, 1998, 2097, 2933, 1347, 3982]. doi.org [4206]. Dokshitzer [2822]. Dollis [89, 277]. Domain [748, 3207, 546, 791, 909]. domains [943]. Domine [4125]. Don [3403, 4340, 3258, 3378]. Donald [3647, 3588, 3100]. Dong [3542]. Donges [3531]. Donnachie [2366]. Donnelly [3754]. Donoghue [3430]. don’t [4325, 4145]. Door [2742]. doped [1588]. Doping [1119, 518, 532]. Doppler [574]. Dose [3444, 1444]. dosimetry [832]. dot [3482]. dots [2326, 2014]. Double [2606, 2505, 2115, 631, 752, 3778, 3413]. Double-diffusive [3413]. Doublet [2966]. doubt [4023]. Dowling [3239]. Down [4367, 1607, 3822, 3416, 3690]. down-conversion [3822]. Dr [3488]. Dr. [1034]. drag [2276]. drama [4328, 4348]. drawings [268]. dream [1972]. dressed [2175]. Dreyer [2340]. Drift [4370, 3078, 3640]. Driscoll [2749]. Drive [2849]. driven [3934, 2179, 3387, 4088, 2628, 4019, 4095, 4265]. Drives [3164]. driving [3563, 3360]. drop [1913]. droplet [1347, 2423]. drum [1627]. Drummond [3368]. dry [1571]. DSIR [127, 168, 213, 220, 264]. Dua [2943]. dualism [112]. duality [3498, 1830, 2961, 3510, 3863]. Duan [3489]. Dubus [3333]. Duck [3575]. Ducloy [2846]. Dudley [2725]. due [430, 165, 1737]. Dufrêne [2834]. dull [1409]. Dunajski [2544]. Duncan [4199, 3791]. Dunlop [3066]. Dunmur [2668]. Dunn [2860]. Dunningham [2655]. Duppe [3611]. duration [1903]. during [1442]. Dürr [4154]. Durrer [4157, 2292]. Dust [3309, 3348, 2469, 2345]. Dusty [2652]. Dutta [3539]. DVD [3276, 3314]. DVD-ROM [3276, 3314]. Dwarfs [4090, 1749, 1798]. Dwight [3574, 3485]. Dykman [3024]. dynamic [3919, 3672]. Dynamical [1391, 3340, 2869, 2695, 1474, 3905, 4019, 3967, 2352, 3002, 2315, 2859]. Dynamics [3350, 2786, 2840, 2727, 3373, 4366, 2333, 3366, 3202, 2839, 2821, 2500, 2976, 2416, 3471, 2181, 2184, 1887, 3554, 1422, 1427, 2604, 1994, 3515, 1805, 1416, 3552, 2734, 1350, 1503, 1893, 952, 3678, 3832, 2062, 2208, 3749, 3996, 3933, 651, 2219, 4088, 3707, 3867, 3761, 4168, 1817, 3025, 2561, 2806, 2950, 2997, 2316, 2376, 2540, 2663, 4190, 3604, 3489, 4172, 3377, 3479, 3366, 2461, 3430, 4301, 3482, 3921, 3932]. Dynamo [1768, 684, 403]. Dyson [3918, 3551, 3574, 3412]. Dzevad [2975].

E-Book [4286, 4118, 4143, 4209, 4183, 4155, 4291, 4285, 4347, 4196, 4187, 4302, 4186, 4354, 4189, 4144, 4352, 4297, 4308, 4265, 4220, 4243, 4307, 4318, 4142, 4185, 4254, 4272, 4303, 4149, 4275, 4268, 4139, 4154, 4356, 4325, 4327, 4344, 4148, 4235, 4222, 4257, 4269, 4270, 4295, 4274, 4336, 4150, 3423, 4200, 4205, 4224, 4349, 4337, 4201, 4244, 4260, 4334, 3337, 4171, 4145]. ear [421, 475]. earliest [1773]. Early [2815, 1214, 2854, 3014, 2921, 4120, 4078, 4060, 4106, 2686, 2883, 3157, 3395, 3052, 3350, 2989, 3061, 3098, 3201, 2667, 4241, 4275, 3022, 3180, 3506, 2819, 2648, 2911, 2840, 3107, 2593, 3298, 3365, 2830, 2715, 3124, 2544, 3376, 4231, 2795, 2955, 2865, 3116, 4154, 3349, 3137, 4113, 4356, 4328, 2866, 3038, 3095, 3035, 3247, 4118, 3296, 2939, 2836, 2853, 2785, 2988, 3194, 3179, 3353, 3408, 3499, 4240, 4272, 2617, 2775, 3207, 3370, 2621, 2907, 3168, 3245, 2791, 2841, 2999, 4072, 4052, 4178, 4190, 4198, 2965, 3279, 3494, 3501, 3450, 3452, 4081, 4204, 4203, 4162, 4235, 4257, 4259]. early [4269, 4309, 4295, 2693, 4323, 2660, 2800, 2954, 3054, 3407, 2887, 3126, 3205, 4085, 4108, 4068, 4119, 4196, 3184, 3007, 3081, 3074, 3689, 4339, 3099, 4187, 4294, 2538, 2820, 3046, 3431, 3143, 3233, 3258, 3366, 3429, 2670, 2771, 3050, 4165, 3303, 3398, 3492, 2716, 3128, 3274, 3382, 3080, 3005, 4157, 3230, 3210, 3446, 4186, 2539, 3340, 3423, 2526, 3026, 3003, 3424, 3189, 3511, 2680, 3015, 3162, 3447, 2929, 3367, 2512, 3232, 3307, 3449, 3065, 2622, 3430, 2839, 1160, 3272, 4121, 2784, 2712, 3127, 3033, 3012, 4122, 4349, 3094, 2068, 2780, 2624, 2835, 2874, 2924, 2837, 2834, 2760]. early [2750, 2880, 3062, 3119, 3409, 4054, 4083, 4189, 3347, 1647, 2546, 3414, 3476, 3489, 4255, 4219, 4301, 4310, 4276, 4350, 4166, 4167, 2759, 3138, 3346, 2725, 2557, 4199, 4297, 4334, 2896, 3228, 3461, 4253, 4265, 4266, 4332, 3148, 3146, 3141, 3404, 3391, 3410, 3440, 3482, 3441, 3481, 3483, 3439, 3504, 3503, 4063, 4220, 4174, 3963, 2758, 3191, 3185, 3278, 3508, 2799, 2807, 2826, 2654, 455, 2665, 2666, 2832, 3306, 3495, 4058, 4215, 3397, 4185, 4168, 2920, 2882, 2967, 2919, 2586, 2769]. early-career [3365, 4185]. Earth [2335, 3456, 2289, 2898, 156, 3415, 2909, 3096, 4398, 2982, 3085, 3722, 4058, 1101, 3889, 3597, 570, 1751, 4104, 2188, 252, 1522, 178, 1212, 1773, 1924, 1700, 3416, 4386, 1788, 3423, 2672, 2993, 2644, 3276, 3676, 4342, 3134]. Earthquake [1067]. earthquakes [456]. Earths [3876, 3889, 4399, 2190, 2603]. Easy [2639, 861, 3037]. eat [3009]. EBook [4341, 4323, 4319, 4225, 4241, 4293, 4173, 4204, 4309, 4281, 4353, 4233]. Echo [3373, 1076]. eclipse [3908]. eclipses [3751]. ecological [4093]. ecology [4219]. economics [410, 3899, 3503, 3248, 2782]. economists [3503]. econophysics [2960, 2634, 3248, 2657]. Ecosystems [3410]. ed [2509, 4090, 3196, 2971, 3209, 2436, 2511, 2422, 2453, 2475, 2470, 2802]. ed. [3900, 3979]. Eddington [1136, 1264]. Eddy [3488]. Edge [2519, 2647]. Edited [584, 2082, 555, 3157, 3333, 3618, 2414, 2893, 2926, 2756, 3079, 3574, 3098, 3201, 1191, 2296, 2297, 848, 2819, 2527, 2596, 2767, 3051, 2840, 3699, 2386, 2485, 3203, 3772, 2830, 2858, 3313, 4032, 4009, 2795, 2311, 2912, 2751, 2727, 3091, 3133, 3010, 3229, 3261, 3171, 3412, 3320, 3425, 3945, 4028, 4123, 4154, 4221, 3646, 3371, 3137, 4100, 3659, 3747, 1131, 2984, 2798, 2991, 3104, 2988, 3194, 3179, 3353, 3408, 3499, 4240, 4272, 2775, 2779, 3370, 2626, 3075, 2755, 2342, 2421, 4285, 2541, 3004, 3784, 1092, 3047, 3626, 2294, 4252, 4235, 4270, 2366, 2542, 3054, 1069, 3126, 4108]. edited [4119, 4196, 3184, 4242, 4250, 2850, 2737, 3113, 2993, 3436, 3630, 3689, 3787, 4031, 4048, 4015, 2674, 2918, 2387, 2506, 2765, 2709, 2886, 2746, 3312, 3143, 3374, 2396, 3024, 2771, 3050, 3392, 3135, 2431, 3273, 3108, 2437, 2714, 3230, 2371, 2435, 2623, 2379, 2616, 3187, 3340, 3457, 3487, 3678, 2682, 3003, 3189, 2923, 2983, 4159, 4161, 2288, 1057, 2671, 2973, 3246, 3367, 2395, 2647, 2685, 2743, 2739, 2803, 2884, 2946, 3145, 3031, 3307, 3653, 3740, 4099, 4224, 3305, 2839, 2652, 4131, 2467, 2651, 2472, 2712, 3033, 3012, 3666, 3209, 3094, 2780, 2624, 2834, 2760, 2876, 2750, 3062]. edited [3002, 3119, 3063, 3607, 3735, 3795, 2974, 2846, 2441, 3132, 2535, 2443, 3476, 3681, 3760, 2725, 1140, 4334, 2322, 3337, 3461, 2315, 2376, 2378, 2628, 2627, 2629, 2720, 2879, 3059, 3440, 3441, 3481, 3505, 3673, 3596, 3611, 4220, 3281, 4147, 2368, 2503, 2630, 2799, 3664, 2632, 2822, 3087, 2599, 2666, 2388, 2470, 2590, 2721, 3082, 3129, 3684, 3792, 4033, 4058, 4136, 2598, 2787, 3676, 3039, 3067, 3267, 3348, 3594, 3590, 3593, 2922, 1146, 44, 3633, 3714]. Edition [3318, 3201, 3349, 3257, 3494, 3323, 3377, 3317, 3400, 3553, 3427, 3437, 3436, 3588, 2706, 3163, 3429, 3274, 3200, 3231, 3234, 3065, 3271, 3216, 3347, 3461, 3255, 3285, 3286, 3443, 3817, 3766, 4078, 3395, 2463, 2462, 4211, 3721, 4343, 3509, 4275, 2560, 3458, 3796, 2828, 4140, 2335, 4056, 4273, 4293, 2857, 3723, 3609, 4005, 4123, 2320, 4061, 4183, 3647, 2845, 3626, 3826, 4190, 4148, 3280, 3561, 3809, 4081, 4204, 4130, 4259, 4269, 4281, 4295, 4323, 3999, 3668, 3708, 3631, 3643, 3783, 4040, 3957, 4091, 4345, 4346, 4187, 4283, 4294, 4353, 4165, 3510, 3401, 4157, 4292, 2299, 3162, 3208]. edition [3768, 3770, 3430, 3712, 2855, 4282, 4131, 3703, 4122, 4201, 4244, 4053, 4080, 4189, 4234, 4144, 4351, 3132, 3731, 3645, 3996, 3995, 4071, 4066, 4077, 4167, 2314, 2390, 4253, 3672, 3528, 3701, 4092, 4216, 2370, 2388, 2373, 3215, 3411, 3475, 3621, 3667, 3581, 3818, 4004, 4233, 4175, 2319, 4133, 3986, 3947, 3845, 3962, 3920, 3910, 3832, 3973, 3842, 3827, 3922, 3917, 3949, 3863, 3890, 3850, 3916]. Edition/ [4201]. Editorial [419, 922, 1558, 2478, 2697, 2958, 3158, 3288, 3725, 4312, 3289, 1232, 3569]. Edmondson [2809]. Edmund [3349]. edn [2788, 2829, 2786, 2667, 2936, 2927, 2715, 3124, 2868, 2955, 2865, 3116, 2866, 2945, 2995, 2939, 3023, 2621, 2907, 2639, 2792, 2533, 3036, 2901, 2806, 2684, 2954, 3016, 2996, 2523, 2889, 2916, 2738, 2517, 2796, 2873, 2890, 2869, 2997, 2676, 2728, 3109, 2976, 3138, 2557, 2701, 2638, 2913, 2579, 2952, 2982, 2665, 2666, 2646, 2663, 2885, 2736, 2827, 3134, 3153, 3021, 2882, 2919, 2586, 2703]. eds [3553]. Eduardo [3124]. educated [2353, 2438]. education [78, 43, 307, 1019, 289, 341, 2644, 4321, 389, 79]. educational [362]. Edward [3115, 2600, 3177, 1100, 4071, 2802]. Edwin [3411, 2175, 139]. effect [1372, 1933, 933, 1685, 2765, 2164, 909, 1450, 490, 759, 13, 61, 133, 6, 2428]. Effective [1178, 2349]. Effects [2425, 1012, 374, 2422, 834, 832, 3530, 4006, 2007, 1596, 4141, 425, 1987, 1602, 926, 2271, 2445, 542, 683, 507, 2248, 3546, 1635, 3482, 2721, 2247]. efficiency [1797]. Effrosyni [3186]. Efroimsky [2832]. Egdall [3302]. Egor [3460]. Egypt [3063]. Ehrlich [4294, 2371]. Ehsan [3504]. Ehud [3410]. Eicher [3579, 3355]. Eichhorn [3568]. Eigen [3380]. eightfold [3758]. Eiji [3184]. Einasto [3268]. Einstein [3840, 2565, 2912, 4328, 3987, 4348, 3009, 2562, 2864, 2649, 3006, 2436, 3473, 2718, 2558, 4067, 2534, 2704, 2573, 1720, 1950, 4076, 946, 1165, 3473, 3461, 3291, 2277, 4228, 2160, 1165, 3945, 3302, 3071, 3363, 3155]. Einsteinian [4356, 4344]. Eisberg [3379]. Eisberg-’t [3379]. EISCAT [833]. Either [1388, 3056]. Ekert [2846]. Ekpenyong [4318]. Ekspong [2082]. El-Batanouny [2432, 4166]. Elastic [2702, 491, 3900, 3544]. Elasticity [2835, 4198, 3359, 3864, 588, 4190]. Electric [2491, 1948, 2443, 311, 2304, 664, 398, 165, 4289, 61, 597, 2453]. Electrical [515, 3149, 560, 671, 374, 1401, 398, 1211, 1046, 1210, 3647, 3501]. Electricity [3200, 486, 4062, 374, 111, 3115, 3200, 2802]. Electrification [114]. Electro [3323, 2378, 3362]. Electro-optic [3362]. Electro-Optics [3323]. Electrochemical [1529, 4272, 2799]. electrocrystallization [448]. electrodynamic [202, 238]. Electrodynamics [2856, 2848, 4364, 3738, 2811, 1606, 1844, 3398, 2744, 3166]. electrojet [494]. electrokinetics [4153]. Electroluminescence [140]. Electrolysis [514]. Electromagnetic [105, 2757, 3118, 3407, 3198, 2900, 1722, 4106, 3176, 2347, 4234, 4289, 4116, 247, 4140, 4038, 3744, 3747, 2366, 3642, 3382, 2546]. Electromagnetics [3055, 2486, 2347, 3714]. Electromagnetism [176, 3233, 13, 61, 133, 2978, 164, 286, 363, 511, 4232]. Electromechanics [3148]. Electron [2375, 1648, 640, 1364, 3118, 693, 1712, 1670, 1495, 1063, 3059, 1814, 915, 3953, 631, 2221, 881, 476, 1045, 1053, 615, 1998, 845, 1143, 1308, 1956, 1064, 583, 1029, 1677, 2642, 1452, 524, 607, 437, 1759, 686, 953, 971, 2272, 3874, 3882, 4224, 502, 1120, 949, 705, 376, 651, 217, 2209, 455, 2275, 1225, 1131, 3026, 3937, 2344]. electron-atom [1677, 1452]. electron-hole [1308, 953, 949]. Electronic [2966, 3894, 2096, 3265, 1518, 1541, 1204, 4247, 4278, 2076, 4081, 1027, 807, 1807, 901, 558, 962, 3903, 543, 2209, 1055, 4215, 1244, 3799, 651, 3116, 2466, 56, 555, 2487, 4122, 3255]. Electronics [2795, 2998, 144, 771, 325, 1538, 65, 4229, 3101, 2746, 3446, 4260]. Electrons [3029, 2300, 1254, 860, 161, 174, 182, 190, 200, 788, 1285, 3779, 1151, 651, 3446, 1057, 4172]. electrophoresis [1544]. electrostatic [522, 164, 1841]. Electrostatics [3118, 205, 298, 285, 303, 4206]. electrothermal [3961]. Electrotransport [731]. Electroweak [1818, 2183, 3677]. Elegance [3100]. element [4340]. Elementary [2560, 2897, 1318, 4417, 3317, 354, 2244, 197, 285, 315, 46, 2240, 416, 296, 3913, 286, 1001, 3801, 3731, 4271, 4045, 2239, 2243, 2633]. Elements [2433, 938, 2769, 1204, 3889, 607, 1324, 2165, 119, 890, 3244, 2296, 3835, 3341, 2729]. Elena [3004, 2535, 2592]. Eli [2619]. Elias [4131]. Elihu [2755]. Elisa [3667]. Elizabeth [3203, 4150]. Elkins [3689]. Elkins-Tanton [3689]. Ellad [2874]. Ellen [3333]. Elliott [2874]. Ellis [2311, 4355, 4346, 2903, 3171, 2984, 3690, 3684]. Elsevier [662, 4183, 4186, 4224, 4054, 4053, 4189, 4216, 4111]. Elvira [2387]. embryonic [1487]. emergence [3131, 3794, 3597, 1960, 4276]. Emergent [4075, 3469]. Emidio [4160]. Emigration [222]. Emilio [1265]. Emin [3169]. Emission [3098, 2102, 314, 1922, 643, 123, 2030, 698, 699]. emissions [1703]. emitter [4051]. Emitting [2909]. Emlet [4258]. Emmanuel [3306]. Emmy [3817, 2683]. emphasizing [3926]. emulate [4243]. enabled [4236]. encounter [3680]. Encounters [2796, 4321]. Encrenaz [2476, 3415]. Encyclopaedia [3119]. encyclopedia [3119]. encyclopedic [2422, 4031]. end [2810]. endeavour [2278]. endor [631]. enduring [3612]. energies [680, 2000, 2866]. Energy [145, 3350, 2926, 3073, 2662, 3194, 2999, 2597, 2566, 3099, 4076, 2714, 1473, 2738, 2415, 83, 2722, 77, 241, 2789, 2928, 3593, 834, 2450, 3807, 3333, 171, 416, 1922, 4179, 4217, 1340, 3464, 3715, 4272, 3570, 2733, 3784, 1020, 1924, 764, 3824, 4294, 4049, 653, 686, 3352, 160, 1716, 1075, 137, 2831, 412, 3818, 1127, 2447, 2293, 3542, 2661, 4257, 3274, 3892]. engagement [4186]. engagement/kaminski/978 [4186]. Engel [3667]. engineer [4217, 3499, 3488, 2309, 2465, 2464, 2671, 2584, 4241, 4252, 4242, 4250]. Engineered [3374]. Engineering [3122, 2842, 450, 2396, 2797, 2590, 4078, 420, 428, 618, 2749, 4026, 4052, 1104, 647, 3768, 2653, 3712, 4019, 4265, 3434, 4216, 4405, 444, 2041, 2358, 3055, 2794, 3323, 2457, 2675, 3208, 2395, 3806, 3995, 3484, 2474, 2784]. Engineers [3014, 3178, 2967, 2883, 3122, 3395, 3333, 109, 4275, 2828, 3011, 3154, 2865, 3349, 3371, 3038, 3035, 2836, 3275, 3194, 3179, 3353, 3408, 4272, 2941, 2791, 2841, 4190, 4198, 2702, 2660, 2800, 2954, 2944, 3323, 3407, 3205, 2674, 3233, 3366, 2670, 3050, 2523, 3005, 4186, 2474, 3283, 3026, 3003, 3424, 3140, 3246, 2647, 3084, 3145, 3259, 3232, 3307, 3449, 3049, 3231, 3234, 2784, 2874, 2924, 2880, 3062, 3119, 2722, 4188, 2725, 2557, 4199, 4095, 4265, 4266, 3404, 3440, 3482, 3441, 3481, 3483, 3505, 3504, 3503, 2910, 2665, 2928, 2969, 3484, 3474, 3300, 4185, 2915, 2920, 2882, 2919, 2769, 4120]. engineers [4078, 4060, 4106, 4248, 4072, 4178, 4323, 4068, 4071, 4334, 4084, 4197, 4188, 3536, 3974]. engines [932, 3966, 2453, 3349]. Englert [2846, 3398]. English [4187]. Enhanced [3408]. enhancement [3159]. Enigma [2796, 4027, 1992, 3997]. Enlarged [3065, 4187]. enlightenment [2278]. enquiry [231]. Enric [4141]. Enrico [2674, 3810, 3214, 2634, 4400]. Entangled [1407, 2055, 2371]. entangled-photon [2055]. Entanglement [2071, 2854, 3175, 2862, 1840, 4192, 2935, 4328, 4348, 3510, 3863]. Enterprise [2449]. Entin [4096]. Entin-Wohlman [4096]. entropies [3221]. Entropy [1297, 2866, 1574, 581, 2281, 3654, 4175, 3258, 4217, 3645, 2994]. entry [3979]. envelope [3926, 3311]. Environment [834, 252, 896, 1510, 1535, 1522, 2224, 2404, 2798]. Environmental [1671, 4369, 2204, 2722, 2833, 3766, 4240, 4323, 3459]. environments [3891]. Enz [2057]. Eothen [313]. ePDF [4295]. epidemics [2762, 4219]. epitaxial [1341]. epitaxy [1475, 1503]. epitaxy-fundamentals [1475]. epithelium [4300]. EPR [1808]. ePub [4295]. Equation [2995, 2953, 2730, 625, 2154, 3640, 4105, 3902]. Equations [3840, 2457, 2543, 2649, 3256, 3573, 4156, 2605, 1259, 3866, 2414, 2653, 2874, 3402]. equatorial [691]. Equilibrium [2290, 2824, 2406, 3833, 297, 2369, 2219, 4168, 4095, 3038]. equinox [3394]. Equipment [2670, 4244]. equivalence [1614, 238]. Equivalent [3198]. era [3518, 2045, 3101, 3613, 3825]. Erasmo [4113, 3370]. erasure [1943]. Erdmenger [3498]. Eren [3407]. Erenso [4232]. Eric [3244, 3597, 2983, 3183, 2582, 3042]. Erik [3119, 2804]. Ernest [3171]. Ernesto [3642, 3665]. erosion [2213]. Erratum [198, 427, 1224, 1321]. error [3991, 4216]. Errors [396]. eruption [2001]. Eruptive [2223]. Erwin [3309]. Escribano [3678]. especially [2811]. Espenak [3751]. Esposito [2674]. Essay [2099, 3520, 3909, 155, 1121, 1335, 2893, 2772, 3523, 1265, 1778, 1191, 1808, 1800, 440, 675, 848, 1820, 2034, 3522, 1836, 3911, 205, 3425, 1040, 1083, 1777, 1852, 2057, 1131, 223, 2329, 3102, 3524, 1783, 1904, 1857, 146, 1938, 1801, 965, 1092, 2191, 875, 1770, 1069, 4015, 3910, 1843, 2035, 2765, 316, 850, 3421, 2644, 3423, 1100, 1851, 3078, 1057, 1819, 56, 1813, 1165, 1653, 129, 3521, 1858, 2552, 1937, 1140, 2701, 3337, 1039, 1562, 1814, 1844, 1827, 188, 2041, 1146, 44, 834, 1835, 1936, 584, 2426, 2413, 2710]. Essays [3381, 2311, 4196, 3681, 2983]. Essence [2486, 1504, 4083]. Essential [2312, 3212, 3067, 1442, 3541, 3212, 3315, 3552, 3887, 3655]. Essentials [4415, 3139, 4182, 4304, 2320, 3055, 3457, 2950, 4174, 3068]. Essig [4028]. Establishment [77, 241]. Establishments [213, 127, 168, 220, 264]. estimation [4331, 4053]. Estrada [3665]. Ether [1132, 4100]. Etienne [3500]. Ettore [2359]. Eugene [3430, 2924, 2580, 3476, 3505]. Eugenio [3204]. Euler [3235]. EUR [4148, 4208]. EUR139.13 [4224]. EUR150.00 [4284]. EUR35.00 [4307]. EUR39.99 [4321]. EUR46.30 [4325]. EUR87.52 [4307]. Euro [4253]. Europe [3956, 740, 3001]. European [869, 667]. evaluation [1503]. Evans [2542]. evaporatively [1720]. evening [50]. events [3800, 2223]. ever [1916]. Everett [2508, 3469, 3164, 3469]. Everettian [4014]. everybody [2080, 3752]. Everyday [3302, 2971, 4322]. Everyone [4396, 4285, 3871, 4145]. Everything [3283, 4400, 2124, 3295, 3856, 3810, 1601, 3867, 3931]. Everywhen [2581]. Everywhere [2581]. Evidence [3093]. Evolution [2843, 4378, 3266, 3364, 3422, 2709, 2839, 2467, 2620, 2501, 2994, 3134, 755, 4123, 374, 3479, 3861, 3993, 3970, 4146, 81, 1851, 1911, 4114, 3533, 1220, 2960, 890, 2621, 2826, 3928]. Evolving [2993]. exact [4210, 2718]. Exactly [3652]. exam [3492]. examination [2359]. Example [3052]. Examples [1678]. EXCEL [4353]. excellence [1927]. Excel(R) [2901]. Exchange [1515, 1458, 653]. Excitability [2409]. excitable [2325]. excitation [1518, 1541, 831, 623, 637, 653, 1517]. excitations [1027, 2897, 3214, 3632]. excited [980, 1776]. Exciting [1531, 3516]. Exciton [2516]. Excursions [3262]. exegesis [3798]. exercises [4324, 3138]. Exhibition [38, 144, 65, 36, 37, 97, 158, 441, 485, 308, 526, 568, 619, 442, 39]. exhibits [214, 276, 350]. existence [2161, 3794]. Exner [2370]. exobiology [4123]. Exofrontiers [4221]. exoplanet [3698, 3910, 2654]. exoplanetary [4221, 3733]. Exoplanets [3067, 4399, 2709, 3415, 4338, 3876]. Exotic [722, 565, 2971, 1879, 2514]. Expanded [3349]. Expanding [2778, 3620]. expansion [837, 1615]. expeditions [3908]. expenditure [675]. experienced [3967]. Experiment [2164, 2712, 3853, 2383, 414, 2120, 789, 545, 3916, 2992]. Experimental [3395, 1255, 2837, 2750, 1149, 982, 3480, 785, 2294, 3824, 2519, 3510, 1529, 1812, 3863, 4078, 4309, 2616, 3555, 3236]. experimentalists [4202]. Experiments [2927, 2662, 4383, 3006, 2924, 151, 646, 5, 3696, 2339, 861, 816, 3161, 980, 1811, 4087, 4089, 3997, 3590, 2647, 3696, 4318, 3458]. experts [2659, 3206, 2453, 2388, 2590, 2663]. experts/advanced [2388]. Explained [3071, 2453]. explaining [3731]. explanation [3649]. Exploding [2929]. Exploration [3020, 1974, 482, 1021, 281, 305, 312, 329, 337, 344]. Explorations [412, 435]. explorer [1628]. explores [4172]. Exploring [2195, 1590, 602, 3206, 3433, 3151, 3270, 3608, 19, 4187, 3114, 2899, 2860, 2608, 4039]. explosion [1657]. explosions [820, 456, 3586, 3075, 3761]. explosive [459]. exposition [440]. Exposure [3575, 1596]. extending [3731]. Extensions [3311]. external [3482]. extinction [1629]. Extra [3040, 2839]. Extra-Solar [2839]. Extragalactic [752, 3907]. Extraordinary [3403, 3128, 4397, 4047]. Extrasolar [2476, 3334]. Extraterrestrial [3195, 3378]. Extreme [3356, 3242, 3242]. extremely [142, 4263]. extremes [2110]. Eye [2943]. eyes [1399, 3615]. Ezawa [2422]. Ezekiel [1191].

F [899, 4324, 2357, 4268, 2486, 2485, 3772, 2375, 3539, 2425, 3716, 2795, 3695, 2984, 3322, 2988, 2764, 3308, 3212, 3193, 2779, 2931, 2255, 2421, 2392, 4081, 2383, 2366, 3113, 3016, 3341, 3553, 3631, 3643, 3773, 3751, 4040, 4042, 4346, 2387, 2333, 3374, 2562, 2476, 2526, 3189, 2946, 4093, 3697, 3430, 3712, 2903, 2432, 4201, 2846, 3255, 3265, 3566, 3555, 2422, 4059, 2632, 2426, 3082, 3535, 3484, 3684, 4172, 3397, 2768, 4231]. F-theory [2383]. F.R.S [1136]. F.R.S. [648, 796, 1034]. Fabbrichesi [4160]. Fabian [3520, 4323]. Fabre [3772, 2686]. Fabrice [3788]. Fabry [3862, 657]. Faces [3840, 894, 2441]. facilities [2355]. facing [270]. Facsimile [3473]. Fact [2548]. Factors [387]. Facts [3913, 2343]. Fadeev [3885]. Fadin [2609]. Fähnle [2522]. Fai [4308]. Failed [2658, 3502]. Failure [3384]. Faist [3146]. Faith [3426, 3636]. Falcioni [2351]. Falkovich [2790]. fall [1265]. Fallacy [2753]. Fallani [3245]. Fallible [1644]. Falling [1265]. false [2133, 1171]. Familiarity [3380]. family [1139]. fantasy [3636]. Fantini [3565]. Far [495, 2272, 582, 602, 3833]. Far-infrared [2272]. Faraday [2650, 165, 84, 247]. Faraoni [4243]. Faria [2681, 2316]. Farr [3264]. Fascinating [4322, 4326]. Fascination [2371]. Fashion [3636]. Fasolino [3033]. Fast [903, 283, 1594, 887, 2809]. fastest [4172]. Fatermans [4224]. Father [4400, 3810]. Fathi [2315]. fatigue [428]. Fauchet [2396]. favourite [2034]. Fayyazuddin [3401, 2855]. Fe [4180, 4180]. Feats [2675]. feature [1629]. features [4149]. Featuring [2901]. Federico [4321]. Fedor [2719]. feedback [2235, 2039, 3482]. feet [3492]. Feigelson [3183]. Felder [2368]. Feldman [2295]. Feldmeier [4336]. Felice [2648]. Feliciano [3450]. Fellows [811]. Femto [3516]. Femto-world [3516]. Femtosecond [2233, 2924]. Fereira [2409]. Fermi [4003, 3810, 1057, 4368, 1143, 1855, 4400, 161, 174, 182, 190, 200, 2674]. Fermilab [2356]. Fermion [1817, 1333]. Fernández [2737, 3706]. Fernando [2648]. ferrite [782]. Ferroelasticity [1897]. Ferromagnetic [1459, 2522, 484]. ferromagnetism [1578]. Ferruccio [3584]. Ferry [3451]. Festschrift [1057, 2787]. Few [1863, 2034]. Feynman [2350, 2864, 1833, 1872, 4397, 2583, 4047]. fi [1243]. Fiber [1191, 2719, 3740]. Fiber-Optic [1191]. Fibers [2725]. fibre [1820, 4126, 1438, 3745, 3862]. Fibre-optic [3862]. Fibrous [544]. fiction [2452, 3841]. Field [2346, 3079, 2290, 2715, 2490, 2801, 2999, 2687, 2524, 3345, 3189, 2587, 2622, 3125, 2681, 2875, 3316, 2546, 2349, 2422, 557, 2369, 2666, 1844, 1263, 2264, 2221, 1424, 113, 3539, 3491, 3899, 3911, 4035, 3791, 3914, 4177, 2382, 3585, 3159, 684, 1212, 403, 2394, 1697, 2366, 3884, 4002, 4044, 4141, 4050, 1700, 3511, 2263, 3819, 1566, 165, 3879, 3880, 1467, 4306, 240, 3225, 3672, 597, 2453, 1483, 2469, 1213, 3124, 2757, 2824, 2291]. Field-ion [557]. field-theoretical [3884]. Fields [554, 2524, 2744, 826, 3176, 4018, 4061, 1657, 1791, 1582, 445, 1686, 4041, 4029, 2519, 4315, 1768, 272, 429, 4075, 61, 651, 4116, 1691, 2949, 2525, 2443, 3217, 3763, 3999, 3669, 3736, 3944]. Fifth [3257, 2206, 1504]. Fifty [2850, 3885]. Fight [2980]. Figueroa [3370]. Figures [3083, 4140, 4139, 3083]. Fijiwara [2657]. Filippenko [3215]. Filippo [3004, 4146]. Filling [204, 210]. Film [633, 3028, 313, 1886, 241, 2828]. Films [279, 1441, 1457, 515, 1666, 1503, 484, 1202, 1511, 3105]. Filomena [2566]. filter [1788]. filters [62, 2828]. final [2765, 1603, 2396, 2300]. final-year [2765]. finance [1965, 2521, 3256, 3768, 3899]. financial [3252]. find [3698, 2235]. Finding [4037, 2435]. Fine [2680, 2753, 393]. Fine-Tuning [2753]. finely [3695, 4387]. finely-tuned [4387]. Fingerprinting [1811]. fingerprints [2171]. Finitary [2634]. Finite [4008, 2330, 1467, 2580, 1317]. Finite-temperature [1467]. Finland [834]. Finn [3901]. Finzi [4115, 2380]. fire’ [4027]. fireworks [1799, 3956]. firmament [1084]. First [4357, 4358, 2870, 3875, 4213, 3081, 2980, 3369, 2365, 2748, 3299, 2126, 3239, 2559, 3020, 3909, 1137, 3974, 1217, 1225, 3692, 1026, 2329, 3193, 3797, 4294, 3335, 2300, 3428, 2571, 2460, 1642, 4033, 1464, 3680, 4401, 3886, 3135, 2825]. first-year [3335]. Fischer [3015, 3528]. Fischer-Cripps [3528]. fisherman [2735]. Fisk [2858]. Fitzpatrick [3199, 3000, 3734, 3556]. Five [3020, 1131, 1254, 4121]. fix [2549, 1694]. fixes [1624]. FL [4259, 4269, 4270, 4294, 4318, 4319, 4302, 4267]. FL/Abingdon [4259]. flagellar [2304]. flames [817]. Flamini [4031]. Flaser [4042]. Flash [273]. Flat [2668]. Flat-Screen [2668]. Fleisch [4105, 2944]. Flewitt [3809]. flexible [4251, 4250]. Flight [3641, 517, 1164, 283, 2889]. flights [2763]. Flood [3281]. floor [482, 743, 493]. floppy [2079]. Flow [2079, 2500, 3402, 1852, 2406, 1737, 2941, 3429, 2513, 3147]. Flower [3201]. flowing [240, 2213, 736]. Flows [2340, 2441, 2934, 1466, 1472, 1328]. Fluctuating [3024]. Fluctuation [3301, 3639, 2466]. Fluctuations [2715, 2466, 3250, 913, 1697, 1387, 4000]. Fluid [3025, 3233, 3366, 3213, 2997, 3172, 2500, 2513, 2976, 3741, 595, 4231, 1457, 4190, 3980, 952, 1328, 3867, 4168, 2790, 2333, 3366, 3604, 2934]. Fluidics [551]. Fluids [2883, 516, 861, 3864, 3716]. Fluorescence [2624, 2032, 4051, 3608, 4169, 1028, 2173]. fluoroscopy [1316]. flux [3957]. Flying [3328, 1998]. Foa [4081, 3265]. Foams [4409, 2041]. Fock [3558]. Focused [3059, 4199]. Foden [2452]. Foe [3426]. foils [217]. folding [2189]. following [114]. Fomin [4044]. Foner [848]. Fong [3299]. food [1842]. Force [4372, 3453, 2834, 113, 3563, 4280, 165]. forces [993, 718, 1579, 1378, 3360, 579, 3097, 4018, 3962, 3703]. Ford [3524]. forecasting [572, 952]. forecasts [1082]. forefront [2303]. Foreign [1133, 1529]. Foreword [1, 409, 3473]. Forgan [4003]. forger [2405]. forgetting [1943]. forgotten [1937]. Forinash [2722]. Form [2748, 1487, 775]. Formaggio [3754]. format=PB [4211]. Formation [1435, 2476, 2839, 2501, 3443, 3348, 3602, 1622, 1782, 2122, 4146, 1529, 4114, 610, 2068, 3360, 1073, 4412, 3993, 2709, 2467, 4033]. formative [3987]. formed [530]. forming [999, 1008, 3860, 3202]. Formulation [2241, 3402, 322, 4218]. Forneris [4347]. Forterre [3213]. Fortov [2652, 3276]. Fortson [4256]. fortunate [3695, 4387]. Fortunato [3282]. forward [249]. Foster [3806]. Fotakis [2803]. Foundation [3473, 2358, 3824, 3160, 42, 3715, 3193]. Foundational [2816]. Foundations [17, 28, 1437, 2911, 3038, 2617, 875, 3205, 2746, 3210, 3125, 3498, 3559, 4284, 3587, 4007, 2455, 2615, 3144, 2937, 3733, 3365, 2984, 4209, 3190, 3754, 3562, 2415, 2722, 2831, 3934]. Four [2622, 3550, 4274]. four-dimensional [3550]. Fourier [2543, 1489, 745, 4353, 3582]. Fourier-transform [745]. Fourth [3494, 3162, 3216, 3509, 4295, 3597, 144]. Fox [3459, 2667, 3748, 3312]. fractal [1357]. fractality [3554, 3497]. Fractals [4360, 2039, 2416, 3497, 3749, 3896]. fractional [2195, 2754, 3922, 3347, 3866]. Fracture [444, 587, 2925]. Fradkin [3124]. Frames [2727]. framework [3791]. Franca [3659]. Francesco [4210, 2833, 3339]. Francis [4296, 4294, 4297, 3615, 4257, 4305, 4283, 4336, 4349, 4308, 4267]. Francisco [3179]. Franck [4005, 3034]. Franco [3125, 3119]. Francois [3487]. Francoise [2527, 2596]. Frank [4025, 2612, 3128, 3182, 3208, 2877, 3483, 2952, 2632, 2650, 2255]. Frankel [3078, 2868]. Franklin [4035, 3607, 3181, 2713]. Frans [2534]. Fraser [2870, 2460]. Fraternali [4146]. Fraud [2548]. Fred [2796]. Frederico [3122]. Free [2866, 1967, 1650, 1265, 4217, 566, 2642, 971, 2272, 3746, 2340]. free-electron [971, 2272]. Freegarde [3107]. Freeman [4038, 3412, 3551]. Freericks [3672]. Freezing [2872, 2287]. Freidberg [2293, 3519]. French [2474, 442]. frequencies [353, 1956]. Frequency [2038, 3207, 2920, 191, 395, 2087, 2327]. Frequentism [2959]. Freudenberg [2945]. Friction [624, 2114, 3780, 92, 320, 6, 3939]. Friday [50]. Friedman [2829]. Friedrich [3126]. Friedwardt [2928]. Friend [3426]. friendship [4117]. frills [4317]. Frisch [2245]. Frishman [2622]. Fritzsch [2864]. frogs [3551]. Fröhlich [2368]. Frolov [3576, 2914]. Front [2548]. Frontier [2356, 2471, 3570]. Frontiers [2519, 2283, 2704, 3784, 2617]. fronts [90]. Frostig [2470]. froth [2041]. frustrated [2283, 3209]. FSO [3746]. Fu [3629]. Fuchs [4321]. Fuel [1797]. Fuels [2849]. Fujii [4150]. Fujimoto [4281, 3404]. Fujita [3501]. Fukata [4220]. Fulde [3029]. full [1882]. Fullerene [1738]. Funaro [4106]. Function [2470, 3832, 3819, 2030, 1793, 2266, 3811, 2265]. Functional [2965, 3450, 4273, 3853, 4373, 3954, 4242, 4236, 2795, 2301, 3264]. functionalized [4323]. functions [4154, 831, 458, 2427, 2632]. Fundamental [3470, 2727, 2141, 1767, 3381, 3431, 3305, 2874, 2608, 2669, 3719, 2339, 4277, 4191, 632, 2038, 3292, 2087, 3739, 4376, 3801, 3220, 2554, 4229, 3470, 3238, 4220, 3714]. Fundamentals [459, 2791, 3626, 3323, 2979, 2333, 3141, 2938, 3706, 4026, 1475, 4158, 3531, 4334, 2098, 3657, 4319, 3786, 3025, 3796, 3011, 3580, 2341, 2792, 3617, 2567, 2816, 3608, 2464, 3307, 3718, 3893, 4266, 3362, 2956, 3069, 2704]. funded [3776]. Furlanetto [3081]. furnaces [224]. Furst [3813, 4258]. Further [2506]. Fusion [2293, 1069, 2050, 2064, 961, 2009, 498, 1114, 1651, 3642, 3824, 2129, 193, 3416, 867, 2027, 719, 853, 1479, 2045, 2231, 1825, 3850, 2646, 3578, 3099, 2646]. Future [3241, 567, 1811, 2599, 3333, 3468, 1974, 1674, 4254, 2012, 1903, 639, 891, 4314, 2384, 1302, 1183, 4262, 654, 4399, 4095, 4265, 3904, 3702, 2031, 3876, 2453]. Fuxi [3441]. Fuxiang [3108, 2987].

G [3766, 3711, 4060, 2883, 3741, 4010, 2297, 3509, 2527, 2596, 4038, 2375, 4296, 3716, 2955, 3695, 1040, 3023, 3652, 2626, 3648, 3040, 2294, 2366, 3407, 3771, 3113, 2993, 3074, 3627, 3752, 3787, 4187, 2783, 3381, 2898, 3235, 2343, 2808, 3210, 2299, 2692, 2881, 3486, 3231, 2729, 2903, 3090, 3019, 3749, 3718, 2835, 2924, 3682, 4170, 2854, 2846, 3183, 3613, 2546, 3690, 2419, 4070, 4199, 2620, 3148, 3567, 3674, 3814, 3555, 2530, 2409, 2368, 2423, 2417, 2488, 2265, 2253, 2599, 2373, 2590, 2663, 2721, 3001, 2969, 3684, 3792, 4207, 2319, 2344]. G. [1497]. Gábor [3247]. Gabriel [3539, 2765]. Gabrielli [4160]. Gagnon [3731]. Gaia [4160, 3781]. Gail [3309, 3713]. Gain [2820]. Gaisser [3667]. Galactic [1416, 3266, 3094, 2780, 3479, 2732]. Galaxies [3081, 3436, 4420, 4421, 2748, 3348, 1836, 1847, 2305, 1714, 4355, 3860, 2552, 3377, 2651, 2552, 2938, 2450, 4146, 3861, 2434]. Galaxy [2467, 1059, 4146, 3781]. Galileo [1265, 2134, 3931, 2968]. Galina [4058]. Galison [3945]. Galitski [3114]. Gallager [3486]. galleries [770]. Galtier [3671]. Gamaly [2924]. gambling [2088]. game [1744, 3463]. games [2088]. Gamma [1551, 2597, 2688, 2929, 3039, 2070, 2813, 4107, 3057]. Gamma-Ray [2688, 3039, 1551, 2070, 4107, 3057]. Gamow [2908]. Gan [3441]. Ganeev [3298]. Gangopadhyaya [2947]. Gao [3811, 3633]. gap [204, 210, 679]. Gaponenko [2563]. Garcia [3194, 2314]. Garcia-Martinez [3194]. Garfinkle [2318, 2318]. Garg [2978]. Garibaldi [2634]. garnet [791]. garnets [438]. Garnett [3275]. Garnier [2517]. GARP [789]. Garrison [3522]. Garry [3099]. Garston [168, 264]. Gary [2613]. Gas [3348, 1204, 1217, 3573, 880, 838, 4146, 550, 2219, 3761]. gas-phase [1204]. Gaseous [3204, 1981]. Gases [2580, 2846, 2807, 1379, 472, 659, 578, 657, 1703, 593, 510, 2436, 4068]. Gaston [2528]. Gastone [3375]. Gaudin [3346]. gauge [2193, 1241, 1247, 3518, 4138, 2961, 4085, 1488, 1138, 4075, 1282, 3498, 3509, 3126, 3216]. gauge-gravity [2961]. Gauge/gravity [3498]. Gauntlett [3305]. Gauss [239]. Gaussian [3515]. Gautam [4159]. Gavrilenko [2800]. gb [4211]. GBP [4335]. GBP26.99 [4348]. Gbur [3682]. Ge [4024, 3030, 3681]. geese [2772]. gelatin [1386]. gelation [1430]. Geller [4294]. gels [1386]. General [2911, 3150, 4303, 3043, 2713, 1631, 3473, 3461, 2938, 2718, 3191, 4120, 4106, 2312, 3318, 3393, 3111, 4057, 4317, 4324, 2462, 2661, 4211, 4286, 2772, 2809, 2989, 3295, 4082, 4343, 4241, 4275, 2491, 2527, 2851, 2843, 3107, 2439, 2403, 2334, 2411, 2585, 2311, 2356, 2471, 2679, 2495, 2658, 2600, 2565, 2564, 2498, 2857, 2870, 2742, 2707, 2904, 2770, 3056, 3229, 3167, 3195, 3403, 3406, 3462, 3425, 3692, 3955, 4073, 2509, 3349, 3137, 4113, 4356, 2977, 3035, 1026, 3442, 2672, 3322, 2673, 2853, 3104, 3488, 4240, 2764, 3102, 3226, 3308, 3444, 2504, 3152, 3139, 3192, 3302, 3502, 4155, 2906]. general [4327, 4355, 4291, 4285, 2392, 3004, 4344, 3009, 2684, 1686, 3177, 4156, 4274, 2302, 2542, 2726, 3268, 3219, 3400, 3394, 3478, 3427, 3436, 3456, 3477, 3455, 3459, 4345, 4338, 4342, 4340, 2289, 2318, 2402, 2452, 2519, 2548, 2506, 2668, 2675, 2849, 2886, 3131, 3088, 2980, 3120, 3254, 3224, 3262, 3253, 3244, 3196, 3375, 3335, 3405, 3381, 3270, 3312, 3453, 2397, 2612, 4322, 3135, 2562, 3121, 3363, 3128, 3018, 3274, 3202, 2714, 3354, 2473, 2430, 2691, 3345, 3415, 3445, 4186, 2371, 2435, 3100, 2367, 2379, 2644, 2493, 2863, 3263, 3283, 3423, 3487, 2299, 2505, 2916, 2881, 2929, 2933, 2864, 3078]. general [4115, 4117, 2466, 2899, 3428, 4321, 2753, 2427, 3084, 3031, 3351, 3234, 3299, 3926, 3271, 2811, 3756, 2694, 2796, 3006, 2908, 3276, 2877, 2434, 2449, 2778, 3106, 4223, 4218, 4316, 3066, 3002, 4036, 2380, 2854, 2520, 2513, 3239, 2695, 2833, 4208, 4311, 4320, 2717, 2149, 2496, 2608, 2814, 2895, 3197, 3225, 2331, 3441, 4065, 4064, 2532, 2582, 2589, 2588, 2913, 3048, 3151, 3149, 3281, 4086, 4171, 3286, 2982, 2417, 2645, 2558, 3136, 2789, 2321, 4109, 2986, 2599, 2934, 2871, 2957, 2827, 2970, 3134, 3001, 3020, 3085, 3129, 3042, 3215, 3284, 3379, 3443, 4233, 2534, 2456, 4258, 3300, 2917]. general [2964, 3378, 3355, 2420, 2650, 2610, 2768, 3164, 4256, 2990, 2988, 4059, 4196, 4182, 3508]. Generalised [2281, 3165, 2362, 3165]. generalized [3562]. generally [2343]. generalreaders [2968, 2986]. generated [423]. Generating [3789]. Generation [2094, 3298, 2725, 107, 1768, 3174, 3590]. generator [73]. generators [522]. genesis [3224]. Genetic [1729, 2604, 3640]. genetics [1988]. genie [1723]. genius [1937, 2896]. Gennaro [3282]. Genne [2527, 2596]. Gennes [2977]. geochemistry [3766]. geodesy [337]. Geodynamics [3588]. geodynamo [1787]. Geoff [3805]. Geoffrey [2529, 4269, 2752]. geography [4015]. Geological [1547, 3705, 4090, 2500]. geology [672, 4015]. geomagnetic [684]. Geomagnetism [3085]. Geometric [1278, 3900, 3894, 3189]. geometrical [3953, 3470, 3661, 2385]. geometrisation [2481]. Geometry [2868, 2856, 2540, 2550, 2894, 4271, 3264, 3540, 4036]. Geophysical [3025, 482, 2976, 4, 102, 879, 3867]. geophysicists [2472]. geophysics [1766, 3595]. Georg [2840, 3398]. George [796, 291, 2984, 3637, 4346, 2765, 3421, 3049, 3234, 3243, 2908, 4354, 2832, 2637, 4030]. Georgiou [2803]. Geosciences [2472]. geospace [3787]. geosphere [3597]. Gerace [3659]. Gerald [3945, 3264]. Gerard [3196, 2397, 3379, 3217]. Gerini [3490]. German [9, 3473, 2018]. germanium [1007]. Germany [4295]. Geroch [2499]. Gerrard [3488]. Gerry [3056]. Gertsman [3596]. Get [3776]. Getting [2420]. GeV [437]. Ghione [2538]. Ghirardi [4154]. ghost [2699, 1691]. GianCarlo [4154]. Gianfranco [3197, 2630]. Giani [2881]. giant [1515, 4316]. Giardina [3735, 3269]. Gibbons [2599]. Gibson [2347]. Gifted [3345]. Gilbert [2686]. Giles [2452]. Gilles [2977]. Gillespie [3186]. Gillet [4095, 4265]. Gilli [3375, 3375]. Gingerich [3651, 2682]. Ginoux [3490]. Ginzburg [2730, 2324]. Giordano [3179]. Giorgio [4210, 2954, 3282]. Giosuè [2884]. Giovanni [2679, 2538, 2526, 3634, 2741]. girl [231]. Girvin [4001]. Gitterman [3280, 2331, 2951]. Giuggioli [4219]. Giuliani [2375]. Giulio [3002]. Giuseppe [3377]. Giustino [3450]. Glasgow [766, 76]. Glass [3254, 3202, 558, 999, 1008, 24, 346, 860, 1028, 530]. Glass-forming [3202, 999, 1008]. Glasses [3002, 991, 4210, 990, 1364, 1106, 558, 559, 4334, 3269]. Glassy [1893, 564]. Glazer [4253, 3599]. Gliding [2851]. global [1699, 1595]. Glover [3054]. glow [3662]. Gluck [3529]. Glueballs [1765]. Gluon [2756, 2313, 1189]. gluons [3778]. GmbH [4022]. go [2146, 2549, 1506]. goals [3618, 1600]. God [2564, 2933, 2986, 3521]. Gödel [3220]. goes [3774, 2893, 2023]. Goethe [3575]. Goff [2582]. going [1960]. goings [1132]. Goins [2720]. Gold [2851, 2582]. Goldberg [3769]. Goldsmith [4016]. golf [1667]. Gołkowski [2967]. Golowich [3430]. Golub [3400]. Golubovic [2740]. Gómez [4032, 2387]. Gómez-Elvira [2387]. Gómez-Pedrero [4032]. Goncalves [3628]. Gonnet [2528]. Gonzalez [2365, 2769, 4170]. González-Acuña [4170]. Gonzalo [3300, 2964]. good [662, 1412, 2520]. Goodchild [1092]. Goodson [3749]. Goodstein [2548, 2922]. Goody [3468]. Gorbunov [2815]. Gordon [2926, 2870, 2460, 3731, 4233, 2154]. Goriely [2760]. Goringe [1131]. Gösta [2082]. Got [3262, 2019]. Gott [3650]. Gottvald [4030]. Götze [3202]. Gould [2625, 4282]. Governing [2874]. government [324, 792]. Govind [2597, 2820]. Govindjee [4178]. Gradient [3373]. gradiometry [4234]. Graduate [3142, 2659, 2485, 2375, 3491, 4273, 2425, 2868, 2411, 2311, 2350, 2490, 3738, 4209, 2749, 2842, 2352, 2341, 2342, 2421, 2563, 2301, 2294, 4259, 4269, 4309, 4281, 4295, 2366, 2567, 3438, 2597, 2489, 2566, 2609, 2528, 2816, 2706, 2918, 3013, 2396, 3492, 2365, 2817, 2438, 3454, 2497, 2309, 2465, 2464, 2671, 2584, 2455, 2499, 2512, 2681, 2713, 2744, 2875, 2729, 2824, 2821, 2818, 2972, 3045, 3019, 2467, 2472, 2487, 4079, 4169, 4170, 4234, 2580, 2825, 2909, 3183, 2441, 2440, 2500, 2418, 2416, 2419, 2510, 2511, 2833, 2711, 2745, 3064, 2895, 3228, 2377, 2331, 2568, 2859, 2982, 2298, 2401]. graduate [2370, 2369, 2503, 2488, 2444, 2683, 2831, 2291, 2470, 2885, 2345, 2704, 4158]. graduates [2993, 2355, 2647, 3216, 2903]. Grady [3648]. Grae [2513]. Graf [4295, 2368]. Graham [2684, 4040, 3016, 596, 3615]. Graham-Smith [4040, 3016, 3615]. Graining [2351]. grains [1571]. grammar [117, 231]. Granada [169]. Grand [2658, 1305, 3021, 3502, 4155, 1957, 3989]. Grandy [2994]. Grant [2680]. grants [29, 99]. Granular [3002, 2219, 2059, 3607, 3213]. Graphene [2838, 3894, 3265, 2573, 3628, 3501, 4081, 3358, 3819, 2271, 3101, 4130, 3279, 2902, 3941]. Graphene-Based [3265, 4081]. Graphene-Related [3894]. graphic [4258]. Graphs [2808, 3789]. gratings [471]. gravitas [2104]. Gravitation [2911, 3126, 1013, 4389, 4390, 4391, 2611, 1906, 469, 1334, 3508, 1219, 2693, 2617, 3437, 3846, 3153]. Gravitational [646, 977, 4329, 645, 2250, 1760, 1550, 1415, 2023, 3546, 998, 3291, 1356, 3858, 4101, 1812, 3893, 2069, 4064, 3792, 3916, 3709, 3854, 2249]. Gravity [2481, 4372, 3093, 3453, 2340, 2716, 3155, 4299, 3191, 243, 3774, 2193, 1250, 430, 819, 1777, 2731, 1707, 2961, 3954, 4085, 4108, 4141, 4230, 843, 978, 4280, 1262, 597, 2454, 3802, 3778, 2425, 2939, 2984, 3190, 2431, 2524, 2716, 2714, 2298, 4043, 3477, 3757, 4234, 4064, 3498, 3948]. gravity-induced [597]. Gray [2883, 2468, 3577]. Greanya [3496]. Great [2778, 2558, 614, 662, 4222, 3751, 4325]. greatest [1916]. Greco [4138]. Greek [3763, 2887]. Green [3553, 3819, 2647, 3691, 3314, 458]. Greenberg [3656]. greenhouse [1450, 1703]. Greenstein [3243]. Greenwich [793]. Greer [2463]. Gregersen [3425, 3483]. Gregor [2652, 3276]. Gregory [2790, 2623]. Grémaud [2846]. grenades [2128]. Gribbin [2434]. Gribov [2822]. Gribov-80 [2822]. Grier [4054]. Griffin [2580]. Griffiths [4066, 2718, 1459, 3121]. Grigory [3233]. Grijs [2780]. Grimmett [2808]. Gripaios [4211]. Grissom [2867]. Gronchi [2526]. Gross [3133]. Grossmann [3924]. Ground [844, 604, 3027]. Ground-based [604]. Group [4296, 4365, 3930, 4257, 4305, 4283, 4294, 4336, 3424, 3310, 4349, 4297, 4308, 2758, 2398, 4267, 3758, 3954, 37, 3073, 3625, 4056, 3960, 3959, 4167, 2631]. Group-Theoretical [2398]. Groups [3382, 446, 1575, 4008, 3409]. Groves [3699]. grow [4276]. Growth [4379, 4384, 1562, 1430, 141, 584, 848, 4241, 1622, 1782, 3626, 1503, 1341, 346, 2982]. growth-point [584]. Grynberg [2686]. Grzegorczyk [2394]. Gubbels [2525]. Gubbins [2883]. Guéry [2921]. Guéry-Odelin [2921]. Guest [1232, 2444]. Guichun [3649]. Guide [2509, 3150, 3105, 2944, 3258, 2635, 4385, 2680, 2807, 2957, 2610, 3164, 3806, 3776, 4120, 4078, 4060, 4317, 3868, 3946, 3830, 3458, 2851, 2843, 3955, 4123, 1059, 1510, 1535, 2337, 3961, 2779, 4325, 4327, 1229, 3793, 3343, 3496, 2302, 3478, 3708, 3577, 3936, 4105, 4187, 3018, 4292, 3415, 2299, 2899, 3756, 4218, 2624, 2834, 2880, 3991, 1015, 4070, 3138, 3870, 2987, 3397, 4258, 1588, 2420, 161, 174, 182, 190, 200, 2989, 3076, 3192, 2299, 3127, 2546, 2848, 4333, 3804, 3979, 4142, 3959]. guidebook [3478, 3724]. Guided [2567, 3138, 3929]. Guided-Wave [2567]. Guiding [2341]. Guido [3935]. Guillaume [3788]. Guitar [2474]. Gulbenkian [42]. Gun [3262]. Gunapala [2775]. Gunn [4098]. Gunnar [3070]. Gunther [4347]. Guo [3649]. Guozhen [4250]. Gupta [4302, 3540]. Gurfil [3685]. Gutfreund [4356, 4344, 3473]. Guyon [3500]. Guzik [3506]. Gyaneshwar [4259]. gyromagnetic [115]. Gyroscopes [340].

H [3806, 2463, 4286, 2414, 1459, 1191, 2296, 440, 2828, 3203, 2830, 2858, 2727, 3261, 3412, 4307, 1040, 2257, 1083, 2708, 2838, 2988, 3524, 3652, 4072, 4222, 3407, 2601, 4024, 3113, 3553, 3436, 3459, 3705, 3816, 3804, 4042, 4003, 2528, 2372, 2898, 3592, 2539, 2525, 3187, 3717, 2355, 2950, 3246, 2940, 3432, 3049, 3234, 2729, 2818, 2903, 2664, 2295, 2332, 3209, 3094, 3634, 4053, 2393, 2436, 2391, 2782, 4352, 3745, 2442, 2530, 2952, 2409, 2451, 3178, 2373, 2663, 2484, 3484, 3675, 3722, 2637, 3761, 2344, 44, 1642]. H.-S [2838]. H.M.S.O [834]. Habitability [2476, 3897, 3238]. Habitable [4407, 3134, 3869]. Habitat [3958]. Hackensack [4132, 4167]. Hadron [2471, 3403, 3442, 2542, 3173, 2183, 3465, 3731, 3594, 1488]. Hadronic [2366]. hadrons [722, 2229]. Hagai [4217, 3266]. Hagar [3431]. Haija [3315]. Haken [3095]. Hale [3435]. Half [4289, 1137, 3828, 3797, 2278]. half-century [1137, 3828]. Half-cycle [4289]. Halide [4246, 1271]. halides [991]. Hall [2437, 2422, 1728, 759]. Hamiltonian [2950, 3932]. Hamiltonians [3527]. Hamm [2696]. Hammond [3494]. Han [4068, 3318, 3609, 3108, 2987]. Hancock [3653]. Handbook [3091, 3083, 3407, 2372, 3312, 2632, 4078, 4060, 3395, 4317, 4202, 2527, 2830, 3748, 3035, 3353, 2790, 2494, 2791, 2841, 4187, 2724, 3050, 4165, 2623, 2505, 3208, 3768, 2671, 2973, 3246, 3367, 2518, 3232, 3486, 3653, 3231, 4201, 2837, 2834, 2876, 4169, 4170, 2511, 2725, 4266, 2568, 3440, 3481, 3505, 2492, 2388, 2590, 3484, 2769, 4267, 3035, 3207, 3440, 3771, 3910, 3614, 3803, 2830, 2673, 3353, 3889, 2841, 3626, 3823, 3208, 2671, 2973, 3246, 3145, 31, 3607, 2627, 3481, 2388, 2590, 3484, 2919, 3849]. handbook/reference [2388]. handbook/textbook [2568]. Handed [2300]. Handfield [2957]. handle [1378]. Hands [2899]. Hands-On [2899]. Hanel [4017]. Hanoch [4356, 4344, 3473]. Hans [3241, 2865, 4295, 3309, 3437, 4321, 3232, 3272, 3611, 3153, 4185]. Hans-Jürgen [4295]. Hans-Peter [3309]. Hao [3681, 3996]. Haphazard [1180]. Happer [2783]. Harald [2995, 2864, 4351, 2510]. hard [2962, 4353]. Hardback [4138, 4052, 4068, 3459, 4322, 4318, 4117, 4158, 4121, 4311, 4253, 4174, 4078, 4060, 4106, 4010, 4057, 4062, 4110, 4137, 4225, 4082, 4343, 4241, 4217, 4284, 4232, 4268, 4135, 4140, 4038, 4000, 4032, 4009, 4056, 4055, 4035, 4231, 4005, 4045, 4018, 4028, 4123, 4069, 4177, 4221, 4100, 4307, 4061, 4333, 2057, 4240, 3102, 4030, 4248, 4251, 4355, 4173, 4348, 4026, 4072, 4178, 4198, 2191, 4163, 4252, 4006, 4021, 4081, 4156, 4130, 4162, 4222, 4259, 4269, 4281, 4270, 4274, 4305, 4002, 4001, 4043, 4041, 3999, 4044, 4008, 4108, 4119, 4098, 4141, 4242, 4250, 4031, 4037, 4047, 3993, 4040, 3994]. hardback [4048, 3957, 4042, 4003, 4015, 4091, 4146, 4345, 4338, 4342, 4346, 4339, 4340, 4319, 3998, 4165, 3421, 4107, 4157, 4007, 4017, 4114, 4159, 4161, 4099, 4093, 4200, 4164, 4019, 4282, 4131, 4197, 4201, 4223, 4244, 4249, 4316, 4020, 4036, 4083, 4080, 3996, 3995, 4071, 4129, 4255, 4208, 4188, 4219, 4310, 4276, 4350, 4320, 4066, 4077, 4166, 4199, 4260, 4334, 3337, 4096, 4094, 4332, 4064, 4063, 4092, 4147, 4109, 4004, 4039, 3956, 4033, 3997, 4058, 4116, 4067, 4172, 4136, 4215, 4233, 2041, 4256, 4267]. hardcover [4120, 4210, 4273, 4271, 4154, 4027, 4328, 4327, 4029, 3423, 4234, 4306, 4290, 4023, 4207, 4335, 4168]. hardening [873]. hardness [873]. Hardy [2373]. Hargittai [3724, 2600, 2849, 3724]. Hari [3526]. Harkness [4343]. Harkness-Brennan [4343]. Harmonic [3298]. harmony [2206]. Harold [2335, 4294, 3597, 3598]. Harper [3093]. Harrie [1205]. Harris [3094, 2449]. Harrison [4132, 1100]. Harry [4347, 2613]. Hartle [4196]. Hartnett [4150]. Hartog [3745]. harvesting [2733]. Harvey [2625, 4282, 2974]. Harwell [159, 206, 1029, 241]. Harwit [4143, 4065]. Hassani [2552]. Hatsuda [2313]. Haupt [3333]. Hauptman [2662]. Haur [2773]. Haus [3796, 2717]. Haven [3306]. Havlicek [2370]. Hawkes [4224]. Hawking [4346, 4117, 2599, 2731, 2249]. Hayward [3273]. Haywood [2758]. hazard [1067, 850]. hazards [1068]. hbk [4293, 4156, 4203, 4309, 4281]. Heacox [3620]. Healey [3759]. health [1596]. Hear [3042, 3042]. hearing [421, 1627]. Hearnshaw [2504]. Heart [3206, 2083, 2086]. Heat [677, 3068, 1342, 3250, 1049, 3760, 4111]. Heather [4201]. heats [935]. Heaven [2742]. heavens [3478]. Heavy [1333, 1069, 1130, 1069]. Heavy-fermion [1333]. Hecht [3036]. Héctor [4170]. Heever [2663]. Hehl [3126]. Heidarzadeh [2353]. heights [604]. Heikkilä [3301]. Heilbron [4073, 4328, 4348, 2968]. Heinemann [3246]. Heinosaari [2935]. Heinz [2840]. Heisenberg [3497, 2864, 2018, 2677]. held [834]. Helge [2658, 3502, 3048]. Heliophysics [2993, 3891, 2721]. Helios [1029]. Helium [462, 595, 1332, 2423, 1680, 2128, 2573, 723, 1172, 724, 1532, 2277]. helium-3 [723]. Heller [3042]. Helliwell [4129]. Helmut [2522, 3510, 4311]. Helrich [2502]. Hemley [2472]. Henchman [3472]. Hendry [2839]. Heng [3733]. Henley [3171, 2314]. Henneaux [3133]. Henneberger [2421, 2946]. Henri [2861, 3490]. Henrik [3425]. Henry [3078]. Hepp [2368]. her [2261, 1235, 2262]. Herbert [1092]. Here [2430, 1506, 2019]. Heritage [2803, 2211, 4065, 2590]. Hermann [3095, 4158]. Hermitian [2759]. Hernandez [3370]. Hernandez-Figueroa [3370]. Heroic [4394, 3857]. Herold [2991]. Herrmann [4185, 2754, 3347]. Herschel [3192]. Heterogeneities [3002]. heterogeneous [999, 1008]. Heusler [536]. hexagonal [782]. Hexatic [1441]. Hey [3509, 3216]. Hickey [4342]. Hidden [3254, 3226, 3386]. Hide [3163]. Hideaki [2657]. Hideo [2472]. Higgs [3318, 3403, 3609, 3128, 3731, 3197, 3225, 3129, 2810, 3390, 2183, 3465, 2986]. High [145, 171, 3350, 2926, 3073, 652, 416, 554, 2662, 1580, 1176, 1769, 2999, 3321, 3824, 2538, 434, 1560, 1494, 352, 429, 1120, 2415, 4223, 3594, 3593, 1146, 1213, 2220, 375, 1007, 523, 1501, 680, 191, 476, 1012, 517, 2217, 3570, 3784, 477, 1543, 208, 1020, 707, 463, 764, 3861, 4283, 657, 960, 4117, 1076, 1203, 160, 4326, 854, 1716, 833, 1075, 137, 142, 1363, 2831, 2447, 3298, 2597, 4283, 2738, 3012]. High-Energy [2415, 3570, 1716]. High-energy-density [3824]. high-level [1501]. high-N00N [2217]. High-order [3298]. High-Pressure [4223, 3012]. High-Speed [2538, 652, 1176, 1012]. High-temperature [1769, 1007, 1543, 208]. high-voltage [476]. higher [4029, 2658, 3502]. highest [2000]. highlights [4345]. highly [980, 1428]. Hilairy [4150]. Hilbert [2357, 2370, 2492]. Hill [89, 277, 3660, 2691]. Hillery [3368]. Hinode [3886]. Hirakata [2925]. Hirishi [2657]. Hiroaki [4173]. Hirose [3439]. Hiroyuki [3544, 2925]. Hirschfelder [1092]. Hirshfeld [2953]. Hisashi [3404]. historian [4113]. historians [2504, 2353]. Historical [1149, 3520, 2356, 4356, 1194, 1242, 4344, 3381, 3475]. histories [3654]. History [4357, 4358, 2335, 1880, 4363, 2329, 3168, 4419, 2887, 2668, 2353, 2430, 2367, 2694, 2827, 2456, 3300, 3468, 1778, 1865, 3602, 468, 4284, 1169, 328, 1875, 3748, 3877, 1952, 1423, 3226, 3308, 3420, 1799, 21, 32, 3856, 3987, 1773, 1143, 2577, 4091, 2548, 2506, 2849, 3335, 3312, 1834, 2437, 2167, 2435, 3907, 4200, 3654, 3473, 3612, 3664, 1616, 3985, 455, 3802, 3997, 2117, 2160, 2504, 3524, 3441]. Hitler [3253]. hobbyists [2387]. Hoboken [4120, 4273, 4304, 4235, 4250, 3995, 4334, 4332]. Hobson [3212, 3193, 3739]. Hoddeson [2356]. Hodge [2672]. Hofer [3247]. Hofer-Szabó [3247]. Hoffman [3653]. Hofmann [2948]. Hofmeister [4111]. Hogan [2354, 3191]. Hole [3383, 1309, 2562, 2914, 710, 2252, 3576, 2251, 1308, 953, 949]. Hole-mediated [1309]. Holes [2857, 3040, 2597, 2716, 3273, 2651, 2934, 3411, 2013, 2078, 2479, 3784, 3477, 1681, 432, 1264, 4208, 3963, 1512, 2916, 3191, 4033, 4335]. Holgate [2537]. Holger [2974]. Holly [739]. Holmes [2869]. Hologram [928]. holographic [3329, 3915]. Holography [422, 525, 3418, 2429, 3232, 2852]. Holstein [3754, 3430]. Holthuijsen [2484]. Holton [3945]. home [1596]. Homogeneous [3376]. Honegger [3558]. Honig [4189]. Honor [3308]. Honour [3412, 1057]. Hooft [3379, 2397, 3379]. Hoover [3023, 3023]. hope [983]. hopes [1560]. Hopper [4339]. Horace [1343]. Horacio [3123, 3132]. Horatiu [4002]. Horing [3819]. Horizons [3273]. hormone [600]. Horton [3688]. Horwitz [3538, 2454]. Hosepipes [1832]. Hot [1749, 2190, 1285, 2441, 4102, 4049]. Hotakainen [2321]. Houches [3772, 2760, 3836, 2846]. Houle [2333]. House [4079, 16]. Houssa [3666]. Howell [2488, 4136]. Hsiung [3720]. Hsu [2777]. https [4211, 4206, 4186, 4201]. Hu [4141]. Huang [3635, 2475]. Huard [3374]. Hubble [4041, 1874, 1714, 1482]. Hubeny [3067]. Hubert [3208]. Huggett [2581]. Hugh [2508]. Hughes [3277]. Hugo [3370, 3240]. Hulin [3500]. Human [2350, 4367, 3084, 2943, 2722, 2710, 3520, 1745, 3456, 3690, 4398, 1849, 3802]. humanity [3904]. humankind [4316]. Humboldt [4]. Humphreys [3203]. Humphry [331]. Hundred [3314]. Hungarian [2363]. Hunt [2564, 2707]. Huss [2613]. Hutzler [2041, 3248]. Hwa [2756]. Hwang [2927, 3999]. Hybrid [3545, 3161, 3742]. hybrids [1765]. hydras [2163]. Hydraulics [278]. Hydroclimatology [2607]. hydrodynamics [3586, 3278, 3500]. Hydrogen [3375, 123, 1105, 1349, 866, 4245, 3767]. Hydrogenic [2087, 2953]. Hydrostatic [614]. Hyperbolic [2653]. hyperfine [826]. Hypernuclei [1288]. Hypersensitizing [892]. Hyperspace [3923]. hysteresis [6]. Hÿtch [4224].

IA [3626, 3075]. IAC [2453]. Ian [2462, 3509, 2564, 4028, 3205, 3478, 2566, 3216, 2871, 3021, 1507]. IAU [3203, 2727, 2737, 2898, 2467, 2780, 3039, 3067, 3267, 3348]. Iayers [1257]. IB [3626]. IBIS [283]. Ice [2906, 1090]. Icons [4124]. Icy [2737]. ideal [3573, 3519]. ideas [4211, 363, 3856]. ideas-particle-physics-4th-edition [4211]. idiosyncracies [2082]. Idrus [3504]. Igel [3717]. Ignition [2928]. ignorance [387, 3514, 1393]. II [2926, 4232, 3658, 2744, 61, 4297, 175, 1782, 1225, 366, 1247, 1053, 1427, 355, 370, 303, 153, 1008, 245, 369, 532, 401, 210, 637, 559, 1748, 699, 1282, 2596, 3322]. III [380, 653, 133, 2508]. Ikeda [2657]. Ikesue [3319]. Iliopoulos [3736, 3801]. Illuminate [2742]. Illuminating [1799, 3594]. Illumination [3425]. illus [4283]. Illusion [2904, 2017, 1743]. illustrate [414]. Illustrated [2372, 2635, 2705, 2473, 3768, 3849]. Illustrating [3044]. Illy [3006]. Image [815, 2884, 791, 4165, 673, 3351]. Imagery [2456]. Images [2456, 615, 3637]. Imagination [2679]. imaginative [4321]. Imaging [582, 4415, 2842, 2684, 3232, 4224, 2888, 2938, 2470, 1226, 3559, 3550, 208, 1721, 3478, 1316, 839, 3601, 3432, 3874, 3882, 3697, 2006, 3795, 1973, 3201, 3457, 2464, 2973, 3145, 3630, 3582, 3673, 3204, 4020, 3362, 3870]. Imke [3238]. Impact [2527, 2596, 2672, 2898, 1704, 4031, 3970, 577, 3676]. impacts [3891, 3274, 2722, 3520, 834]. imperfect [908]. Imperfections [4407, 3869]. Imperial [54, 96, 180]. implantation [518, 532, 1296]. implementation [4120]. Implications [2126, 2181, 1472, 102]. Imponente [2741]. importance [3333, 778]. important [537]. impressions [36]. imprint [2896]. Improper [2766]. improve [2174]. improved [1516]. Improving [2670, 2552]. impurities [464]. in-between [735]. Inan [2967]. inanimate [3221]. inclusions [2224]. Incoherent [729]. Incompressible [3429]. Index [2394, 2074, 2075, 2151, 2152]. indirect [3547]. Indistinguishable [2148]. Induced [999, 1008, 1518, 1541, 3728, 725, 934, 597, 1244, 2833, 3646]. induction [105, 165, 247]. Industries [136]. Industry [2804, 1119, 2889]. Inelastic [603, 1069, 704]. Inert [550, 593]. inertia [1265]. Inertial [2928, 3824, 340]. Infared [1587]. Inference [3259, 3755, 2208, 2215]. infinite [4210]. Infinity [2764, 3128]. Inflation [1628, 2539, 2815, 2126, 3583]. inflationary [2156]. Informatics [2945, 2945]. Information [3111, 3180, 3425, 2945, 2636, 4416, 1850, 3160, 3210, 2371, 1271, 2985, 2846, 3380, 2536, 2281, 2078, 3506, 4298, 1083, 3518, 4313, 3975, 2107, 3548, 2455, 1943, 1969, 2698, 2055, 4216, 2147, 3684, 4267, 2296, 2297, 2495, 2692, 2455, 2656, 3425, 2400]. Information-based [3160]. informed [2450]. infra [495, 204, 210, 135, 258]. infra-red [495, 204, 210, 135, 258]. Infrared [2696, 2775, 806, 3151, 2220, 937, 582, 1587, 791, 989, 4085, 602, 2272, 3697, 447, 2006]. Inguscio [3245]. inhibition [2102]. Initial [1207]. injection [3482]. inner [4054, 274, 3722]. Innocence [4401, 3875]. Innovations [2804]. Inquiries [4359, 3448]. Inquiry [3243]. insider [3776, 3804]. insight [4149]. insights [4110, 2216, 1653, 2223]. insoluble [850]. inspirational [3719]. inspired [2193, 3496]. instabilities [1391]. instability [1009]. Instantaneous [817]. Instantons [2544, 2999]. Institute [145, 554, 4417, 4240, 605, 4201, 2322, 4147, 27, 97, 143, 158, 40, 308, 4124, 499, 76]. Institution [186, 253, 397, 53, 394]. Instrument [1343]. Instrumentation [3407, 171]. instruments [1887, 31, 4244]. insulating [790]. insulation [303]. Insulator [3230, 697, 2209]. Insulators [3277, 1289]. insulin [600]. Integrable [3311, 2444, 3591]. Integral [2291, 2995]. Integrals [3123, 1402, 3527]. Integrated [1193]. Integration [2301, 2998, 2932]. integrity [950]. Intellect [4404, 2191, 2191]. intelligence [4227, 3195]. Intelligent [2988]. Intelligible [3300]. Intense [2949, 1657, 1198]. intensifiers [815]. intensity [2220, 1791, 2482]. Interacting [3067, 3598, 3833]. Interaction [3122, 2954, 2048, 790, 895, 3712, 3225]. Interactions [2366, 2703, 2924, 680, 1001, 1515, 626, 1328, 3937, 2974, 3744, 2355]. Interactive [3343, 4333]. interdisciplinary [3454, 4219]. interest [4120, 4106, 3318, 3111, 4057, 4324, 2809, 2989, 3295, 2851, 2843, 2585, 2471, 2679, 2658, 2600, 2564, 2498, 2857, 2742, 2707, 2770, 3056, 3167, 3195, 3403, 3406, 3462, 4073, 3349, 4113, 2988, 3139, 3192, 3302, 4155, 2906, 4291, 4285, 3009, 2684, 3177, 4156, 4274, 2726, 3219, 3400, 3456, 3477, 3455, 3459, 4345, 4338, 4342, 4340, 3131, 3120, 3254, 3253, 3196, 3375, 3405, 3270, 4322, 3135, 3121, 3363, 3128, 3018, 2691, 3415, 3445, 2435, 3100, 3423, 2916, 2881, 2929, 2933, 2864, 3078, 4115, 2899, 3428, 4321, 2753, 3234, 2694, 2796, 3006, 2908, 2877, 447, 2778, 4316, 3066, 3002, 2513, 3239]. interest [4208, 4311, 2717, 2496, 2608, 2814, 3197, 3225, 3461, 2532, 3048, 3151, 4086, 4171, 3286, 2982, 2645, 2558, 2789, 4109, 2986, 2970, 3001, 3020, 3042, 3215, 3284, 3379, 4258, 3300, 2917, 2964, 3378, 3355, 2650, 2610, 2768, 3164, 4256, 2990, 2521]. interested [2294, 3206, 2466, 2426, 2413, 2450]. Interesting [1783]. interface [3706, 1898]. interfaces [2216, 2771, 4295]. interfacial [3962]. Interference [103, 62]. interferometer [260, 657]. interferometric [1303, 3893, 1262]. interferometry [4277, 3464, 1947, 3510, 4230, 3697, 3863, 1999]. interior [59]. Interiors [1083, 1074, 3422]. Intermediate [2561]. intermetallic [1333]. Intermetallics [1980]. Intermolecular [3097, 4110, 3703]. Internal [826, 666, 653]. International [78, 266, 4254, 3116, 4191, 3289, 719, 3276, 4218, 3063, 137, 65, 4220, 3020, 3039, 3267, 3593, 834, 3711, 70, 464, 366, 163, 1019, 2050, 2064, 4, 157, 338, 415, 560, 156, 144, 3633]. internet [4248, 1951, 2732]. Interplay [2462, 2343, 3787]. Interpretation [1093, 3086, 1145, 3469, 559, 322, 3469, 4069, 2241]. interpreter [3614]. Interpreting [2983]. intersecting [680]. Interstellar [4238, 4403, 2469, 4411, 1406, 123, 3860, 3993, 1629, 3445]. intra [4110]. intra- [4110]. Intriguing [2491, 2481]. Introducing [368, 409, 382]. Introduction [3393, 2788, 2786, 2138, 2573, 4271, 2754, 2189, 2238, 2621, 3297, 3436, 2489, 4420, 2402, 4421, 2844, 2514, 3303, 2817, 3358, 2642, 3415, 3140, 3449, 3125, 2855, 3000, 2784, 2595, 2761, 3347, 3132, 4299, 2556, 2418, 2960, 2310, 2531, 3285, 2723, 2758, 2618, 4406, 4088, 3178, 2373, 2646, 3397, 2469, 3101, 4364, 4078, 2102, 2569, 2304, 4343, 561, 3974, 4139, 4000, 2403, 1241, 1247, 2480, 4118, 4298, 3687, 2812, 3628, 3329, 326, 2730, 3856, 2961, 3762, 1413, 2394, 3686, 3925, 3553, 3643, 4040, 2176, 2605, 2333, 4112, 3819, 3926, 2482, 2148, 878, 3905, 2332, 4197, 4218, 3922]. introduction [4036, 4169, 1517, 3489, 3693, 4188, 4245, 3330, 4260, 3745, 3888, 3940, 2147, 1296, 3221, 3782, 3847, 3775, 3850, 2345, 4153, 3199, 2346, 2614, 2868, 2865, 3123, 2684, 3120, 3018, 3263, 3216, 3090, 2434, 2390, 2947, 2905, 2426, 2650, 2990, 3757, 3622, 3651, 3790, 3805, 4016, 4286, 3509, 4273, 4045, 4073, 3599, 4303, 3627, 4099, 4352, 3543, 4310, 2413, 3556, 2686, 4062, 3180, 3785, 3107, 2715, 2425, 2955, 3723, 4177, 2995, 2636, 2823, 3507, 2842, 3275, 3591, 2937, 2563, 2533, 2655, 4081, 4281, 3317, 3576, 3671, 3993, 4146, 3592, 3382, 3921, 4336, 4017, 2680, 2309, 2587]. Introduction [2914, 2815, 2837, 2997, 3109, 2976, 3471, 4066, 3014, 4199, 4096, 2377, 3265, 3185, 2665, 2475, 2507, 3829, 3737, 2752, 2852, 3955, 3862, 3879]. introduction/reference [2394, 2332]. Introductory [2801, 4362, 2382, 3328, 2773, 3110, 2637, 2699, 4232, 2184, 2320, 2763, 3738, 2302, 3994, 2318, 2054, 2268, 3768, 2267, 4201, 3991, 2391, 2361, 531, 2420, 2740, 3493, 3917, 4022, 4034]. intuitive [4045, 1678]. invaders [1133]. invariance [405]. invariant [3560, 1728]. invariants [405]. invented [2735]. Invention [2289, 2986]. Inventions [3006]. Inventor [3149]. inventory [3907]. inverse [1334, 4053, 1190]. inverse-square [1190]. inversion [1621, 878, 1257]. Inverted [2637]. investigation [372]. Investigations [2137, 499, 2515, 1277]. investments [3252]. invincible [1393]. Invisible [2679, 3623]. Invitation [4268]. involving [980, 653]. Ioan [3554]. Iodine [917, 4110]. Ioffe [2609, 3800]. Ion [1240, 1069, 518, 532, 3823, 768, 3059, 3294, 1296, 1458, 3161, 1385, 1130, 2640, 2177, 623, 637, 653, 926, 2641, 557]. ion-atom [1458]. ion-irradiated [2177]. Ionic [1406, 1841, 1027, 903]. ionised [1428]. ionising [1444]. Ionization [303, 737]. ionizing [209]. ionosphere [549, 644, 691, 729, 685, 833, 266]. Ionospheric [783, 574, 403]. ions [4118, 1069, 1767, 4212, 637, 1107, 4245, 3467, 1256, 1069, 1069, 3026]. IOP [4343, 4241, 4202, 4232, 4140, 4221, 4333, 4209, 4251, 4325, 4303, 4173, 4347, 4204, 4222, 4281, 4353, 4150, 4161, 4337, 4201, 4244, 4249, 4169, 4170, 4234, 4144, 4341, 4306, 4188, 4260, 4253, 4142, 4174, 4290, 4136, 4207, 4233, 4139]. Ira [3302]. Iraj [3504]. iron [791, 295, 4180]. irradiated [2177]. irradiation [1372]. irregularities [783]. Irudayaraj [3062]. Irving [3565]. Isaac [3878, 2708, 3270, 2437, 1272, 3093]. Isaev [4008]. ISBN [4132, 4120, 4078, 4060, 4106, 4149, 4010, 4057, 4317, 4324, 4016, 4062, 4110, 4211, 4137, 4225, 4286, 2893, 2772, 4082, 4343, 4241, 4217, 4284, 4275, 4202, 4232, 4268, 4135, 4140, 4139, 4254, 4038, 4000, 4032, 4009, 4056, 4055, 4210, 4273, 4271, 4304, 4296, 4206, 4035, 4231, 4293, 3425, 4005, 3945, 4045, 4018, 4028, 4073, 4123, 4069, 4154, 4182, 4177, 4184, 4191, 4221, 2082, 4027, 4100, 4113, 4356, 4328, 4307, 4118, 4061, 4333, 2057, 4143, 4209, 4183, 4240, 4272, 2329, 3102, 4030, 4084, 4155, 4248, 4251, 4325, 4327, 4355, 4291, 4303, 4285, 4138, 4173, 4347, 4344, 4348]. ISBN [4026, 4072, 4052, 4178, 4190, 4148, 4198, 2191, 4163, 4252, 4006, 4021, 4081, 4156, 4204, 4130, 4203, 4162, 4235, 4222, 4257, 4259, 4269, 4309, 4281, 4270, 4295, 4274, 4305, 4323, 4025, 4002, 4001, 4046, 4043, 4041, 3999, 4029, 4044, 4008, 4024, 4085, 4108, 4068, 4119, 4098, 4196, 4141, 4242, 4250, 3459, 4031, 4037, 4047, 3993, 4040, 3994, 4048, 3957, 4042, 4003, 4015, 4091, 4105, 4146, 4345, 4338, 4342, 4346, 4339, 4340, 4187, 4283, 4294, 4322, 4353, 4318, 4319, 3998, 4165, 3421, 4302, 4097, 4107, 4157, 4336, 4292, 4150, 4186, 2644, 3423, 4007, 4017, 3078, 4112, 4114, 4115, 4117]. ISBN [4159, 4161, 4158, 4160, 4321, 4099, 4093, 4200, 4151, 4205, 4224, 4164, 4019, 4282, 4326, 4121, 4131, 4122, 4197, 4349, 4337, 4354, 4201, 4223, 4244, 4218, 4249, 4316, 4020, 4022, 4036, 4054, 4083, 4053, 4080, 4079, 4169, 4170, 4189, 4234, 4144, 4351, 4341, 4352, 4306, 3996, 3995, 4071, 4129, 4255, 4208, 4188, 4219, 4301, 4311, 4310, 4276, 4350, 4320, 4070, 4066, 4077, 4166, 4167, 4199, 4260, 4297, 4308, 4334, 3337, 4095, 4096, 4094, 4253, 4265, 4266, 4332, 4065, 4064, 4063, 4220, 4086, 4092, 4142, 4147, 4171, 4174, 4216, 1562, 4290, 4034, 4059, 4109, 4004, 4039, 3956, 4033]. ISBN [3997, 4023, 4058, 4116, 4067, 4172, 4136, 4207, 4176, 4215, 4233, 4335, 4145, 4175, 1844, 4258, 4111, 4185, 4168, 2041, 4243, 4256, 4267]. ISBN-10 [4170, 4234]. ISBN-13 [4170, 4234, 4168]. Isenberg [3508]. Ishtiaq [3401]. Islamic [3326]. ISN [2701]. Isochronous [3339]. Isotope [172, 462, 595, 1496]. Isotropic [1121, 3376]. Issue [2238]. issues [2129]. Istvan [2600, 2849, 3724]. ITER [3688]. Itoh [2818, 2818]. Iung [2966]. IV [2945]. Ivan [3506, 2878, 2695, 4216, 3189]. Ivancevic [2350, 2350]. Ivezić [3577]. Ivo [3507, 3059]. ix [2082]. Iyetomi [2657].

J [3776, 2312, 3211, 3763, 4062, 2781, 2786, 3721, 1191, 3509, 2374, 4241, 2491, 3115, 3506, 3796, 2290, 3698, 3699, 2949, 3696, 4038, 2830, 2439, 3710, 3656, 3644, 2311, 3579, 3748, 4045, 4073, 2293, 4100, 2995, 1131, 2708, 3179, 2329, 3102, 2779, 2261, 2504, 2766, 2341, 2342, 2494, 2999, 3650, 3047, 3762, 4052, 2777, 3616, 3809, 4021, 4269, 4305, 3054, 3407, 4046, 2737, 3057, 2993, 3074, 3552, 3436, 3668, 3708, 3577, 3627, 3754, 3801, 3797, 3994, 4048, 2566, 2816, 2918, 2387, 3551, 3050, 2898, 3121, 2343, 3345, 4150, 2371, 2435, 2354, 3187, 3597, 4007, 2505, 2950, 4112, 4159, 2263]. J [2269, 2753, 2395, 3653, 4200, 4151, 4205, 3819, 3712, 3218, 2713, 2985, 3238, 4019, 56, 2267, 4121, 2295, 2434, 2472, 3521, 3682, 3736, 3795, 4020, 4022, 4054, 2393, 2436, 2391, 2825, 4351, 2611, 3704, 2860, 2976, 2418, 2510, 2511, 2725, 2717, 2620, 4095, 2315, 2376, 3565, 3567, 3812, 4086, 3788, 2468, 2501, 2349, 2631, 2530, 2947, 3236, 2370, 2369, 2488, 2807, 2968, 2632, 2822, 2986, 2599, 2388, 2507, 2559, 2934, 2721, 3134, 3042, 3284, 3495, 3662, 3765, 4058, 2598, 3397, 3676, 3655, 3761, 4185, 2319, 3590, 3355, 1146, 2919, 2358, 2650, 3694, 3799, 2450]. J-P [2898]. J-R [3708]. J. [1459, 1092, 4388]. J.-C [2315]. J.-M [2976]. J.-P [3710, 2501]. Jaan [3268]. Jack [3111, 3238, 3264]. Jackson [4332]. Jacob [3506, 2912, 2655, 3577, 2622]. Jacobs [3447]. Jacobsen [4079]. Jacques [2527, 2596, 2780, 2917]. Jadczyk [3497]. Jagadish [2775]. Jagit [4272]. Jahn [2428]. Jai [3488]. Jaksch [2985]. Jalil [3534]. Jamal [3141]. James [3025, 648, 4211, 4202, 3945, 3192, 2777, 4196, 3190, 3135, 3230, 3672, 2793, 3508, 3260, 4175, 3655, 3821, 3525, 3397, 2650, 3281]. Jamieson [2767]. Jan [4271, 2878, 2625, 4282, 3217]. Janiak [2983]. Janice [4273]. January [158, 180, 221, 308]. Japan [3993, 156]. Jarmo [4224]. Jaroszkiewicz [3421, 3637]. Jascheck [2433]. Jasheck [2433]. Jason [4347]. Jastrow [2473]. Jau [2783]. Javaux [2387]. Javier [3194]. Jay [3400, 2837, 3264, 3215]. Jayant [2916, 2531]. Jean [2841, 4081, 3446, 3487, 3578, 3500, 3002, 3490, 4265, 3265, 3434, 2956, 3069, 2915]. Jean-Christophe [4081, 3265]. Jean-Marc [3490]. Jean-Michel [4265]. Jean-Philippe [3002]. Jean-Pierre [3500]. Jed [3312]. Jeff [2984, 3242]. Jeffrey [4275, 4222, 2616, 3519]. Jeffreys [2335]. Jelley [3274]. Jenkins [4207]. Jennifer [2661, 4225]. Jenny [2891]. Jérôme [3146]. Jeremy [3448, 3788, 3110]. Jeroen [3076]. Jerome [3589, 3305]. Jerry [2718]. Jersey [4085]. Jesper [3483]. Jesse [4292]. Jesus [3541, 4118]. JET [869]. Jets [3094]. Jeu [3675]. Jianbai [3030]. Jiang [3098]. Jie [3823]. Jim [4137, 2857, 4328, 4291, 4348, 2694]. Jin [3165, 2932]. Jin-Jin [3165].

Jincheng [4334]. Jinghua [2932]. Jingsong [3537]. Jining [2836]. Jinqiao [3489]. Jinyi [3457]. Jirí [2718]. Joachain [2949]. Joannopoulos [4021]. João [3507, 2608]. Joel [3506, 4035, 3181]. Joeri [3267]. Jogesh [3183]. Johanna [3498]. Johansson [2424]. John [3122, 766, 2463, 2593, 3522, 4206, 2615, 2662, 4328, 2823, 3144, 2504, 2937, 4348, 4098, 3315, 2816, 4294, 3274, 2517, 2653, 3430, 3012, 2974, 4260, 4334, 3473, 4332, 2582, 2588, 3529, 2831, 2848, 2669, 4207, 2456, 1653, 4332]. Johns [2845]. Johnson [3506, 4327, 3698, 2726]. Johnston [2462]. Joint [869]. Jon [3242]. Jonathan [2788, 2498, 2687, 3239]. Jonathon [3102, 3129]. Jones [3553, 3436, 2985, 3634, 3148, 3755, 3797]. Joon [2866]. Jordi [3586]. Jörg [4047]. Jorge [3313, 4098, 3678, 3058]. Jorgensen [3662]. Jos [3271]. Jose [4098, 4225, 4032, 3586, 3313]. Joseph [3523, 2749, 4183, 1092, 4316, 3062, 3256, 3512]. Josephson [2247, 507, 2248, 2048]. Joshi [3477]. Joshua [4338, 2688, 4201]. Joslin [2883]. Journey [3017, 1461, 2368, 4253]. Joux [2568]. joy [4291]. József [3006]. Jr [3656, 2435, 2613, 3688]. Jr. [3114, 3230, 2994, 2665]. Juan [4032, 3132, 2629]. Judge [4057]. Judging [2600]. Julien [2527, 2596]. Juliette [4347]. Julio [3300, 2964]. July [366, 834]. jump [4154]. Jumps [1786]. Jun [3585, 3103, 2982]. June [1569, 834]. Jupiter [1022, 1642]. Jürg [2368]. Jurgen [4189, 4356, 4344, 4295, 3246, 3248, 3473, 3348]. just [4076]. Justin [2598]. Jutamulia [2485]. Juyoung [2798].

K=9780750333993 [4201]. Ka-Di [3165]. Kachelriess [4041]. Kai [3470, 3030, 2804, 2540]. Kai-Erik [2804]. Kaler [3135]. Kallosh [3784]. Kálmán [2749]. Kamenev [2824]. Kameshwar [2532]. Kaminski [4186]. KamLAND [2120]. Kamran [3580]. Kanas [4091]. Kane [2294, 4233, 2926, 2533, 3731]. Kaneda [3143]. Kannan [3617, 4203, 4266]. Kapitza [1933, 1040]. Kaplan [2832]. Kappl [4295]. Kapral [2461]. Kardar [2497]. Karen [2975]. Karl [4204, 3208, 2910, 1153, 1181]. Karlheinz [4295]. Karnakov [3114]. Karsten [3619]. Kartashova [2592]. Karthaus [3104]. Kasap [2372]. Kassal [3506]. Kastner [3086]. Katepalli [3143]. Katherine [2535, 2511]. Katsnelson [4130, 2902]. Kauffman [3261]. Kaufman [3644, 3139]. Kaviany [3068]. Kavokin [3788]. Kawai [3039]. Kaxiras [4021]. Kazuya [4108]. Kedar [3582]. Keith [2883, 2991]. Keller [3338]. Kelly [4183, 3341, 3535]. Kelvin [2765, 625]. Kempe [2267]. Ken [2631]. Kenji [3439]. Kenkre [4219]. Kennefick [4023]. Kenneth [3322, 3524, 3152, 2860, 3069, 2915]. Kenya [879]. Kepler [4136, 3532]. kept [2070, 3563]. Kerr [909]. Kerson [2475]. Ketterson [3616, 3187]. Kevin [4077, 3170]. Key [1279, 2456, 865, 1143, 3517, 1516]. KGaA [4022]. Khalili [4291, 2857]. Khan [4235]. Khanna [2853]. Khare [3582]. Khomskii [2897]. Khoo [4098]. Khrennikov [3306]. Khurgin [2342]. Khvorostyanov [3495]. Kibble [3305, 2952, 2243]. Kiefer [2939]. Killer [3239]. kilogram [2892]. kilometre [670]. Kim [4067, 2798, 3493]. Kimball [1092, 2588]. Kimberley [3056]. Kimura [4140, 3184]. Kinder [4292]. Kinematics [2565]. Kinetic [4379, 2619, 3833, 1912, 501]. Kinetics [2818, 3626, 3495, 2673]. King [4038, 4354, 4335, 2357, 3018, 3529]. Kingdom [222, 358]. Kinzer [2420]. Kip [4184, 4190, 4176]. Kira [2805]. Kirill [3040]. Kirin [4258]. Kiritsis [4131]. Kirkup [4078, 2901]. Kirkwood [2195]. Kirsten [4351, 2995, 2391, 2510]. Kiselev [3593]. Kisslinger [3533]. Kitamura [2925]. Kitchin [3286]. Kitsinelis [3581]. Kittell [2547]. Kjellander [3563]. Klafter [2825]. Klages [2416]. Klaus [3098, 3307]. Klebanoff [3767]. Klein [2969, 1880, 2154]. Kleppner [3285]. Klimek [4017]. Klimov [2398]. Klioner [2727]. Kloeden [3340]. Klotz [3321]. Klyatskin [2821]. Klyshko [2746]. Knecht [4349]. Kneller [584]. Knew [4400, 3810]. Knocking [2742]. Knots [1707]. Knotted [2570]. Know [2312, 4396, 2312, 4325, 3752, 3871]. knowledge [2312, 1297, 874, 1951, 2867]. Kobayashi [2988, 2940, 3439]. Koberlein [3438]. Koch [2805]. Kochelap [2377]. Koen [2614]. Kogan [3114, 2466]. Koichi [3764]. Koikegami [3103]. Kok [2636, 4098]. Kolanoski [4158]. Kolb [2356]. Kolenkow [3285]. Kolmogorov [4336]. Kondo [3103, 777]. Kool [2774]. Kopeikin [2832]. Korenaga [2982]. Kosovichev [2898]. Kosso [3774]. Kosterlitz [3313, 4098]. Kosuke [3103]. Kotkin [4187]. Kotlarchyk [2355]. Kounai [2787]. Kouveliotou [3057]. Kovalevsky [3069]. Kow [3046]. Koyama [4108]. Kozhevnikov [4094]. Kozyreff [2488]. Kragh [2658, 3502, 3048]. Krapivsky [2619]. Krausz [4172]. Krawczynski [3094]. Krems [4116]. Kreuzer [2708]. Krich [3506]. Krishna [2996, 2345]. Krishnan [4203, 3998, 4266, 3617]. Kronberg [3669]. Kronmüller [2522]. Krugel [2469]. Krupinski [4165]. Krupp [4042]. Kshitij [4240]. Kubin [3147]. Kugler [3764]. Kulik [2746]. Kumar [2853, 3540, 4159, 2558]. Kunihiko [3103]. Kurt [3447, 1034]. Kuttner [2796]. Kutz [4019]. Kuzelev [2659]. Kuzmany [2865]. Kwek [3412, 3659, 2846]. Kyle [2722]. Kylstra [2949]. Kyohei [3404]. Kyoto [156, 157].

L [3767, 4078, 899, 4225, 2781, 2926, 2374, 2385, 2386, 2458, 3203, 2966, 4009, 3656, 2955, 2356, 3261, 3412, 3695, 3945, 4073, 4328, 3747, 3629, 2766, 4348, 3279, 3177, 3755, 4235, 2429, 1507, 4024, 2993, 3588, 3689, 4048, 4015, 2609, 2816, 4187, 2387, 3093, 3429, 2396, 2771, 2524, 3277, 2437, 2435, 2649, 2940, 3653, 2890, 4019, 3272, 2712, 3033, 3718, 3736, 2846, 3256, 2607, 2860, 2711, 3064, 4276, 3138, 3014, 2401, 2619, 2968, 2822, 3087, 2454, 2590, 2656, 2721, 3001, 4207, 3594, 2650]. L-C [2846]. lab [1823, 3529, 3568, 3740]. Lab-on-fiber [3740]. laboratories [264]. Laboratory [145, 1459, 468, 1045, 1053, 177, 2652, 108, 142, 364, 365, 3397, 2344, 1406, 323, 959, 3488, 2261, 1774, 1235, 2262, 3978, 195, 219, 256, 262, 265, 406, 3850, 380, 211, 765, 242, 3612, 3555, 3458]. labs [4232]. Ladd [4083, 3344]. Ladislas [3147]. Ladislav [3591]. Lady [3299]. Laegsgaard [3483]. Lafyatis [4038]. Lagrangian [2418, 3932]. Lahiri [2560]. Lahti [3710]. Laio [2833]. Laird [4135]. Lakhtakia [2546]. Lallit [4178]. Laloë [4005, 3034]. Lam [3664, 3470, 2540]. Lambert [4260]. Lambourne [3436, 2611]. Lamy [3408]. Lancaster [3345]. Land [139]. Landau [2730]. Landauer [1943]. Landé [112]. Landeira [3678]. Landi [2295]. landscape [2163]. Lane [4117]. Laneuville [4150]. Lanfranchi [4160]. Lang [3152]. Lange [2781]. Langmuir [1202]. Language [2935, 2358, 4342, 3302]. Large [4277, 2471, 3403, 2542, 4263, 3731, 3964, 321, 3594, 1358, 1758, 942, 4346, 1064, 1171, 1495, 4151, 1957, 1054, 1516, 1355, 2183, 3465]. Large-scale [4277, 321, 1758, 4151]. Largs [2765]. Larkoski [4045]. Lars [4098, 3337]. Larsen [3242]. Laser [1518, 1541, 3395, 2949, 3298, 1047, 2228, 1762, 2731, 1349, 1536, 2847, 1496, 2678, 1698, 1452, 2924, 1197, 934, 3852, 1023, 2146, 1177, 1270, 1657, 2007, 4050, 1198, 284, 1677, 971, 1378, 3387, 3395, 3646, 3207, 3531, 3568]. Laser-assisted [1452]. Laser-Based [3207]. Laser-induced [1518, 1541, 934, 3646]. Laser-produced [1047]. Lasers [1114, 184, 622, 1035, 4202, 771, 1742, 2233, 2211, 1413, 1391, 838, 1385, 1443, 1528, 2642, 2272, 1748, 2060, 1480, 3482, 3979, 3475, 3146, 3319, 2424, 2736, 4140, 3323, 3428, 2803, 3888]. Lasing [1621]. Last [4400, 707, 3810, 1254]. Laszlo [3647]. Latif [4033]. latitude [833]. Latora [3814]. Lattice [3518, 4365, 1089, 3424, 3902, 3927, 1241, 1247, 1488, 1009, 2209, 1856]. Lattices [3022, 1241, 1247, 2107, 1831, 2808]. lattices-crystalline [1831]. Launay [3446]. Laura [4343, 4323]. Laureate [3432, 965]. Laurence [2977]. Laurent [3481]. Laussy [3788]. Lautrup [2971]. lava [2213]. Lavega [2618]. Lavrentiveva [2807]. Lavrinenko [3483]. law [1265, 1334, 2142, 165, 133, 3645, 1864, 3721, 238, 3587]. Lawless [2707]. Lawrence [3538, 1092]. Lawrie [3021]. Laws [1175, 2758, 2417, 3839, 164, 596, 2142, 3949]. lay [2353]. Layer [2974, 782]. layers [878]. layman [1687, 1015]. laymen [1241, 1247]. Lazarro [2737]. Lazer [1562]. Lázló [3247]. lazy [3762]. Leading [2647]. Leap [2350, 4316]. learn [3964, 4399, 3876]. Learning [3259, 4225, 3577, 4214, 4019, 4197, 2700, 4341, 3330, 3660]. least [2549, 3762, 4262]. Leather [213]. leave [2026]. Lebedev [4124]. Lecture [2875, 2760, 2846, 3772, 131, 1569, 3836, 3491]. lecturer [4239, 2398]. Lecturers [3114, 3043, 4187, 4283, 2543, 3064, 2723, 4326]. Lectures [4414, 2082, 3337, 2297, 253, 3911, 397, 169, 186, 2912, 2322, 2297, 2936, 3699, 3561, 2693, 4085, 3398, 3770, 2821, 3045, 4063, 3987]. Lederman [2691]. Lee [4333, 4337, 2866, 3499, 2930, 2923, 2393, 2885, 3761]. Leek [4333, 4337]. Leeuwen [3090]. Leeuwen. [3267]. Left [2300]. Left-Handed [2300]. Lega [3189]. legacy [3945, 2003, 2103, 3612, 3935]. Leggett [2261]. legs [2123]. Lekner [4206]. Lemaitre [2835]. Lemons [3793, 3258]. Length [3431, 4376, 2161, 3601]. Lenihan [766]. lens [3605]. Lensing [3291, 1550, 2023, 3709, 3858]. Lenton [2644]. Leo [2484]. Leon [2907, 3400, 2691]. Leonard [4117, 3533]. Leonardo [2740, 3245]. Leong [3659]. Leonid [3559, 2639]. LEP [1064]. Lépine [2780]. Lepri [3760]. leptogenesis [2812]. leptons [1318]. Leroy [2954, 2881]. Lesgourgues [3074]. Lesne [2351]. less [1552, 3473]. lesson [3702]. lessons [3510, 3863].

Leszek [2767]. Let [2411]. Letellier [4004]. Leticia [3772, 3374, 2760]. Letokhov [2424]. Letters [1092, 2765, 4113]. Leuchs [2297]. Lev [2989, 3984, 2913]. Lévai [2822]. Level [4120, 4078, 4060, 4106, 4057, 4317, 4324, 4062, 4110, 4286, 3480, 3498, 4082, 4343, 3509, 4241, 4284, 4275, 3506, 4232, 4268, 4254, 3491, 3485, 4056, 4055, 4273, 4271, 4304, 4296, 4231, 4293, 4073, 4123, 4069, 4154, 4177, 4113, 4356, 4328, 4307, 4118, 4061, 4333, 2057, 3507, 3499, 3493, 3488, 4240, 4272, 3502, 4084, 4155, 4248, 4251, 4325, 4327, 4355, 4291, 4303, 4285, 4347, 4344, 4348, 4072, 4052, 4178, 2191, 4163, 4252, 3494, 3496, 3501, 4081, 4156, 4162, 4235, 4257, 4259, 4269, 4309, 4281, 4270, 4295, 4274, 4305, 4323, 4085, 4108, 4068, 4119, 4242, 4250, 3478, 3479, 3477, 4105]. Level [4345, 4338, 4342, 4346, 4339, 4340, 4283, 4294, 4322, 4353, 4318, 4319, 4165, 3492, 3510, 3421, 3497, 4302, 3200, 4107, 4157, 4336, 4292, 3487, 3003, 3511, 4112, 4114, 4115, 4117, 4159, 4161, 4158, 4160, 4321, 3486, 4164, 3238, 3500, 4282, 4326, 4121, 4122, 4349, 4337, 4354, 4244, 4249, 4316, 4054, 4083, 4053, 4080, 4079, 4169, 4170, 4234, 3490, 4351, 4341, 4352, 4306, 3476, 3489, 4071, 4255, 4301, 4311, 4310, 4276, 4350, 4320, 4070, 4066, 4077, 4166, 4167, 4260, 4297, 4308, 4334, 2701, 4253, 4265, 4266, 4332, 3482, 3481, 3483, 3505, 3504, 3503, 4065, 4064, 4063, 4086, 4171, 4174]. Level [2451, 3508, 4290, 4059, 4109, 2894, 3443, 3484, 3475, 3474, 3495, 4058, 4116, 4067, 4172, 4176, 4233, 4335, 4175, 4258, 4111, 4168, 869, 4243, 4256, 4267, 2414, 1501, 2301, 2567, 237, 245, 3190, 2528, 2519, 2472, 2487, 3979, 2424, 2423, 3372, 2921, 2312, 2686, 2883, 3122, 3157, 3395, 3318, 3393, 3368, 3025, 3111, 3211, 3333, 3052, 3112, 3199, 3350, 2433, 2463, 2462, 2661, 2502, 2788, 2829, 4211, 4225, 2414, 2781, 2346, 2772, 2809, 2786, 2926, 2756, 2989, 3061, 3073, 3079, 3295, 3098, 3201, 2667, 2296, 2297, 2357, 2374, 4217, 2491, 2560, 2614, 2555, 2936, 3115, 3022, 3180, 3458, 4202]. Level [2290, 2897, 2819, 3301, 2648, 2583, 2911, 3241, 2527, 2596, 2767, 2927, 3051, 3097, 2851, 2840, 2843, 3107, 2385, 2386, 2458, 2486, 2659, 2485, 2593, 2949, 2828, 3011, 3203, 3298, 3365, 2830, 2335, 2439, 2403, 2334, 2375, 2522, 2715, 2774, 2858, 3029, 3124, 2966, 3313, 4210, 2425, 2868, 4206, 2411, 3072, 2544, 3376, 2585, 3154, 2795, 2955, 2865, 3116, 3384, 2311, 2356, 2471, 2350, 2459, 2615, 2679, 2545, 2495, 2658, 2600, 2565, 2564, 2508, 2662, 2490, 2498, 2912, 2857, 2870, 2742, 2707, 2751, 2904, 2770, 2856, 2727, 3091, 2981, 3089, 3056, 3017, 3037, 3133]. Level [3010, 3229, 3237, 3261, 3167, 3195, 3171, 3403, 3406, 3412, 3320, 3383, 3462, 3425, 4182, 4184, 4191, 4221, 2293, 2320, 2412, 2509, 2754, 2347, 2757, 3349, 3371, 3137, 2866, 3038, 2977, 3095, 3035, 3247, 2942, 2740, 2867, 2945, 2801, 2948, 3055, 2995, 3142, 3150, 3442, 3123, 3296, 2823, 2939, 2984, 3023, 2672, 2708, 3322, 3373, 4209, 2749, 2842, 3396, 2673, 2798, 2838, 2836, 2853, 2785, 3104, 2988, 3275, 3194, 3179, 3353, 3408, 4183, 2764, 3102, 3226, 3308, 2975, 3076, 3144, 3212, 3193, 2617, 2747, 2428, 2696, 2775, 2779, 3207, 2931, 3092, 3370, 3444, 2941, 2504, 2621, 2907]. Level [3075, 3152, 3168, 2735, 2790, 2937, 2872, 2755, 3139, 3192, 3302, 3448, 2352, 2906, 2766, 2341, 2342, 2421, 2563, 2494, 2792, 2878, 2847, 3083, 3105, 2992, 3245, 3165, 3257, 2791, 2841, 2541, 2533, 2655, 2301, 2392, 3040, 3004, 2999, 2561, 2687, 2845, 3036, 2901, 2806, 2930, 3047, 3009, 3117, 2684, 4190, 4198, 2294, 2394, 2529, 2625, 2537, 2777, 2794, 2776, 3118, 2965, 3181, 3279, 3280, 3177, 3343, 3344, 3450, 3452, 3451, 4204, 4203, 4222, 2702, 2693, 2302, 2366, 2429, 2567, 2542, 2660, 2800, 2954, 2944, 3054, 3204, 3323, 3407, 2601, 2887, 3126, 2979, 3205, 4196, 3184, 3321, 2850, 3297, 2726]. Level [2737, 3113, 3007, 3081, 3057, 3043, 3044, 3074, 3016, 2996, 3268, 3266, 3219, 3377, 3341, 3315, 3317, 3400, 3309, 3394, 3364, 3435, 3438, 3427, 3437, 3436, 3456, 3422, 3455, 3459, 2289, 2313, 2457, 2597, 2489, 2674, 2566, 2609, 2528, 2816, 2706, 2918, 3013, 3099, 4187, 2330, 2318, 2372, 2348, 2402, 2387, 2452, 2519, 2548, 2678, 2538, 2506, 2668, 2675, 2932, 2849, 2765, 2709, 2820, 2886, 2783, 2746, 3131, 3088, 3046, 2980, 3120, 3254, 3224, 3262, 3163, 3287, 3253, 3244, 3196, 3375, 3369, 3335, 3405, 3381, 2690, 3093, 3270, 3312, 3453, 2340, 2333, 2397, 2612, 3431, 3143, 3233, 3258, 3366, 3429, 2670, 3374, 2396, 3024, 2771]. Level [2844, 2724, 3050, 3392, 3135, 2898, 2562, 3034, 3071, 3235, 3169, 3303, 3398, 2431, 2524, 2716, 3273, 2891, 3121, 3363, 3399, 3128, 3018, 3274, 3227, 2300, 2365, 2353, 2523, 2817, 3382, 2703, 3114, 3108, 3202, 3269, 3277, 3401, 3080, 3005, 2437, 2343, 2714, 2808, 3354, 3230, 3210, 2473, 2430, 2691, 3345, 3415, 3446, 3445, 4186, 2292, 2371, 2435, 2623, 2539, 3100, 2367, 2351, 2379, 2461, 2474, 2354, 2476, 2438, 2616, 2644, 2493, 2525, 2889, 2902, 2863, 3263, 3283, 3187, 3186, 3182, 3340, 3423, 3457, 3454, 3460, 2299, 2682, 2505, 2526, 3026, 3424]. Level [3189, 2355, 2680, 3015, 2692, 2950, 2797, 3162, 3447, 2688, 2916, 2881, 2929, 2748, 2738, 2923, 2933, 2864, 3078, 2983, 3208, 3206, 2288, 2466, 2594, 2497, 2899, 2773, 3428, 2309, 2465, 2464, 2671, 2584, 2753, 3140, 2973, 3058, 3246, 3367, 3027, 3030, 2395, 2455, 2427, 2517, 2647, 2543, 2518, 2649, 2653, 2685, 2499, 2512, 2743, 2739, 2803, 2940, 2884, 2946, 3084, 3145, 3031, 2998, 3259, 3232, 3307, 3351, 3449, 3432, 4200, 4205, 4224, 3049, 3231, 3234, 3243, 3304, 3299, 3065, 3271, 2587, 2622, 3430, 3125, 3216, 3218, 3305, 3310, 2694, 2796, 3006, 3053, 2839, 2908, 2681, 2689, 2713]. Level [2744, 2873, 2875, 2729, 2824, 2914, 2821, 2855, 2935, 2818, 2890, 2903, 2869, 2972, 3155, 3000, 3045, 3090, 2971, 2985, 3019, 3070, 3248, 3166, 3282, 3316, 2652, 3276, 3272, 3103, 2460, 2664, 2877, 3240, 3311, 555, 2295, 2467, 2434, 2449, 2651, 2778, 3106, 2332, 2472, 2432, 2487, 2784, 2712, 3127, 3033, 3012, 3086, 4197, 3209, 3213, 3214, 2415, 2595, 2815, 3094, 3332, 4201, 4223, 4218, 2780, 2624, 2835, 2874, 2924, 2837, 2834, 2760, 2876, 2750, 2880, 3062, 3066, 3002, 3119, 2997, 3063, 3172, 3409, 4189, 3156, 2393, 2974, 3413, 2436, 2391, 2380, 2580, 2657, 2581]. Level [2677, 2521, 2676, 2943, 2825, 2909, 2854, 2728, 2782, 2846, 3028, 3109, 3183, 3256, 3188, 3347, 3314, 2441, 2440, 3132, 2611, 2500, 2520, 2607, 2556, 2513, 2705, 2860, 2722, 3096, 2976, 3402, 3239, 2389, 2410, 2546, 2535, 3198, 2443, 2418, 2416, 2419, 2695, 2510, 2511, 2833, 2711, 2745, 3064, 3414, 3380, 4208, 4188, 4219, 2759, 3138, 3014, 3346, 2725, 2717, 2557, 4199, 2322, 2310, 2314, 2390, 2496, 2608, 2620, 2896, 2814, 2741, 2895, 3197, 3225, 3228, 3461, 2315, 2316, 2317, 2377, 2376, 2331, 2378, 2568, 2592, 2925, 2719, 2720, 2804, 2879, 2888, 2805]. Level [3148, 3059, 3146, 3141, 3255, 3264, 3265, 3319, 3362, 3339, 3404, 3391, 3410, 3440, 3441, 3434, 3439, 3433, 4220, 2532, 2582, 2531, 2589, 2588, 2938, 2910, 2913, 3048, 3151, 3149, 3285, 3281, 2718, 2579, 2723, 2978, 4216, 2442, 2468, 2501, 3286, 2987, 2349, 2422, 2540, 2530, 2793, 2951, 2952, 2859, 2758, 2947, 3236, 2982, 2409, 2298, 2401, 2424, 2398, 2370, 2400, 2368, 2399, 2369, 2423, 2619, 3217, 3191, 3185, 3278, 2503, 2417, 2645, 2558, 3136, 2799, 2789, 2488, 3147, 2321, 2618, 2807, 2826, 2905, 2994, 2968, 2654, 2444, 2683, 2492, 2831, 2822, 2953, 3087, 2986, 3032, 2848, 2454]. Level [2669, 2536, 2599, 2291, 2665, 2666, 3178, 2453, 2388, 2475, 2373, 2426, 2413, 2470, 2613, 2646, 2590, 2507, 2591, 2559, 2656, 2663, 2484, 2885, 2802, 2934, 2956, 2721, 2900, 2871, 2957, 2710, 2736, 2827, 2832, 2928, 3082, 2970, 3134, 3068, 3001, 3069, 3153, 3020, 3110, 3085, 3129, 2969, 3041, 3021, 3042, 3215, 3242, 3170, 3260, 3284, 3411, 3338, 3306, 3372, 3379, 4207, 4215, 2534, 2598, 2787, 3397, 2456, 3039, 3067, 3267, 3300, 3348, 4185, 2319, 2344, 2752, 2345, 2469, 2922, 2917, 2964, 3378, 3355, 2547, 2915, 2920, 2882, 2967, 2919, 2852, 2586, 2358, 2420, 2650, 2610, 2768, 3164, 2769, 2704, 2990]. Level [2450, 4173, 3939, 3372]. levels [535, 1752, 3078]. lever [509]. Levi [3095, 2518]. Levinson [3956]. levitated [3293]. Lévy [2763]. Lewenstein [3022]. Lewis [3695, 3119, 3117]. LHC [2294, 3825, 2647, 3613, 2294, 2044, 2360, 3677, 3054]. Li [3165, 4347, 4108, 4119, 3542, 2660]. Li-Ying [4347]. Liam [3583]. Library [555]. Licata [3794]. Lichtenberg [3619]. Lidar [3362, 2032]. Lidar. [2288]. Liddle [2539]. Lie [4008, 3409, 3409]. Lieb [2601]. Lieber [3674]. Life [2977, 3102, 3152, 2387, 2473, 3340, 4388, 1165, 4397, 2657, 3380, 4400, 2804, 3281, 2970, 3378, 3984, 3695, 2945, 4237, 1623, 3931, 965, 3456, 4047, 2364, 3060, 3810, 4387, 2049, 2483, 2863, 3423, 3608, 3597, 1744, 4398, 1327, 3684, 3676, 2173, 1237, 3600, 1461, 3129, 2637, 4090, 3487, 3525, 4275, 4342, 2834]. lifetime [3590]. Liger [3196]. Liger-Belair [3196]. Light [3393, 2227, 2342, 1697, 1175, 3289, 1923, 3058, 1921, 2551, 2909, 2629, 1378, 1042, 1373, 1856, 3327, 7, 2450, 2032, 844, 3290, 4287, 4213, 3659, 1791, 3396, 3386, 3570, 3550, 603, 806, 2287, 658, 756, 260, 1198, 3906, 3623, 1752, 1885, 1387, 4302, 725, 3907, 2121, 2023, 1831, 3770, 3356, 2570, 4194, 1461, 3575, 2194, 622, 2155, 139, 4256, 2686, 3318, 2411, 2309, 3087, 3609, 3608, 3598, 3122, 2820, 3712, 3338, 3581]. Light-Emitting [2909]. Light-field [1697]. light-matter [3659, 3712]. Light/Dark [2450]. Lighter [2589]. Lighthill [2376]. lighting [3581]. lightning [2027]. Lights [2005]. Lightwave [4060]. LIGO [4101]. Like [3383]. Lima [2918]. Liming [4252]. Limit [3276]. Limitations [2862]. limited [1912, 1658]. Limits [2867, 1538, 4143, 1444]. Lin [3560, 3319, 2927, 3661]. Lincoln [2819, 2608, 2471, 3403, 3378]. Linden [3444]. Lindesay [3190]. Line [2807, 4010]. linear [2012, 2734, 88, 1029, 3562, 310, 2543]. Lines [2548, 3198, 553, 753]. Linfeng [2836]. Linguistics [3157]. Link [4351, 667, 3504]. linking [4194]. Lintott [4109]. Lions [3172]. Lipatov [2609]. liquefaction [880]. Liquid [2668, 1913, 1441, 1405, 1870, 3900, 452, 1308, 500, 730, 4049, 1673, 953, 949, 1099, 1894, 724, 501, 2375]. liquid-simulator [501]. Liquids [2872, 3202, 735, 2443, 2573, 728, 339, 1187, 603, 4279, 847, 857, 2062, 297, 1855, 304, 424, 491, 538, 805, 531, 355, 3842]. Lisa [2742]. Lissauer [3238]. List [1018, 1025]. literacy [2552]. Literature [3472]. lithography [1063]. little [4338]. Littmann [3751]. Liu [3649]. Live [2834]. lived [981]. Liverpool [51]. lives [2057, 3405]. Living [2449, 4398, 872, 4274, 3221, 3456]. Livio [2714]. LL.D [648]. llan [3458]. Llewellyn [3092]. Lloyd [4325]. Lloyd-Ronning [4325]. Lluis [2629]. Local [901, 2161, 2224, 1691, 3354]. locality [2160]. Localization [1257, 2755]. Locating [2307]. locations [4114]. locomotion [2337]. lodestone [1169]. Loeb [3081, 2748]. logic [2133]. Logical [3320, 2623, 363]. Lok [4141]. London [4057, 291, 3102, 2191, 4349, 4223, 4249, 4189, 89, 4070, 4308, 3337, 4095, 4265, 3956, 834, 4243, 53]. Long [3088, 893, 782, 2282, 240]. Long-range [893]. Longair [3089, 3168, 2738, 3612]. Longer [1182]. Longitudinal [2354]. Look [3383, 1674]. Looking [249, 131, 1973, 3206]. loops [313]. López [3132]. Lord [380, 291, 1165, 3702]. Lorenz [4039, 3220]. Lorenzo [3481, 3058]. Lorrain [2333]. losses [122, 6]. Lost [2565]. Lotfia [3063]. Lou [4250]. Louck [2374, 2793]. Louie [3718]. Louis [3261, 2766, 3578, 3172]. love [3924, 3532]. Loveday [3012]. Lovelock [2364]. Loves [3163]. Lovett [2636]. Low [1340, 3301, 4028, 2566, 686, 94, 2547, 844, 1049, 2337, 1422, 1427, 1752, 3760]. Low-Energy [2566, 1340, 686]. lowdown [2217]. Lowenstein [2950]. lower [1516]. Lozier [2632]. Ltd [4106, 4268, 4240, 4204, 4222, 4169, 4170, 4234]. lubrication [162]. Luc [3500]. Luca [3393, 3002, 2833, 4219, 4310]. Lucas [4057, 4283, 4185]. Luciano [2856, 3278]. Ludovic [3002]. Ludovico [2529]. Ludwig [3885, 3798]. Luhmann [3627]. Lui [3664]. Luigi [2945]. Luis [4081, 2651, 2711, 3265, 1844]. Luiz [2909]. Lukas [3036]. Lukashov [4110]. Lukong [4308]. Luminescence [2177, 2224, 2878]. Luminosity [3594]. lunar [717, 3870, 337]. Lung [3046]. Lunine [3134]. l’univers [81]. Lupei [3319]. Lynch [3656]. Lyndon [2542]. Lyne [3016]. Lyth [2539].

M [3520, 2433, 2414, 2893, 3201, 3115, 2767, 3699, 2659, 2949, 3203, 2966, 4055, 3384, 2904, 2727, 3056, 3037, 3171, 3677, 4123, 3150, 694, 3599, 1131, 3628, 3212, 3193, 2779, 3648, 4030, 2352, 2341, 2494, 4325, 4138, 1830, 3679, 2394, 2794, 2776, 3343, 3617, 4203, 4001, 4046, 4041, 4024, 3007, 3057, 3400, 3752, 3773, 3751, 4031, 3957, 4042, 2566, 2330, 2402, 2340, 2396, 2844, 2716, 2300, 4292, 3230, 2473, 2691, 2351, 2354, 2476, 2525, 3632, 2355, 2692, 2797, 4114, 1057, 2465, 2455, 2427, 3653, 3740, 4099, 3819, 3234, 3065, 3218, 4164, 3053, 2744, 2875, 2935, 2890, 2903, 4121]. M [3106, 2432, 2487, 3718, 4122, 2415, 3094, 4201, 4223, 2750, 3063, 3634, 3795, 2825, 2846, 3028, 2441, 3132, 2500, 2976, 2389, 2410, 2419, 2511, 3681, 4129, 4219, 4070, 3138, 2725, 2314, 4095, 4266, 2317, 2377, 2331, 2378, 3141, 3566, 3505, 3610, 3813, 4065, 3702, 4092, 2349, 2530, 3236, 2409, 2401, 2398, 2370, 2368, 2417, 4034, 2321, 2826, 2265, 2444, 3087, 2848, 2373, 2646, 2656, 3215, 3484, 3707, 3737, 4033, 4067, 3067, 3300, 3655, 4111, 2319, 4267, 2450].

M-theory [1830]. M.-L [2966]. M.K.S.A. [387, 176]. MA [4009, 4056, 4167]. Maartens [2903]. MacCallum [2414, 2903, 3508]. Maccato [4242]. MacDonald [3775]. MacGregor [4275]. Mach [1342]. Machacek [2358]. Machán [4169]. Machine [3145, 3577, 4019, 3330, 3594, 4197, 4341]. Machines [3374, 1949, 4270, 2286, 4214, 3966, 2698, 2857]. Maciej [3022, 2544]. MacKay [3052, 2546, 2489]. Mackillo [2805]. Macleod [2828]. Macmillan [313]. macro [3942, 3252]. macromolecular [3663]. macromolecules [1220]. Macroscopic [3303, 2971, 2994, 1635, 1055]. Macroscropic [3738]. Madame [3299]. Maddaloni [3207]. Made [3037, 4386, 1570, 1703, 2644]. Madhusudhan [4221]. Maeder [2501]. Maekawa [3184]. Maffucci [3714]. Magdolna [3724]. Magellanic [1355]. magic [3926, 2399]. magma [2001]. Magnesium [2789]. magnet [407]. Magnetic [571, 554, 790, 2830, 3373, 446, 2836, 2975, 791, 812, 2494, 845, 2408, 535, 909, 3103, 3214, 2440, 402, 2646, 3085, 826, 1225, 126, 546, 1934, 818, 872, 1212, 1582, 873, 403, 445, 3617, 4309, 4046, 1129, 359, 369, 3669, 3957, 1700, 4315, 3992, 839, 1515, 607, 1768, 272, 863, 429, 781, 683, 1421, 2761, 4234, 61, 240, 3964, 4266, 651, 560, 1825, 2275]. magnetically [935]. magnetically-ordered [935]. Magnetism [251, 1666, 3200, 388, 464, 561, 584, 667, 848, 3898, 4309, 1128, 157, 338, 2204, 2283, 982, 3802, 1213, 3115, 2403, 3142, 2907, 129, 2802, 4062, 440, 584, 848, 2494, 4046, 2990, 111]. magnetization [165]. magnetizing [4094]. magneto [425, 909, 438, 683, 2333]. magneto-conductivity [425]. Magneto-Fluid [2333]. magneto-optic [909, 683]. magneto-optical [438]. magnetoencephalography [4236]. Magnetohydrodynamic [49]. magnetohydrodynamics [3671, 3978, 3332]. magnetometers [4236]. Magnetometry [724, 3156]. magnetoresistance [1515, 2209]. magnetosphere [855, 4104]. magnetospheres [1530]. Magnets [2757, 1843, 2205]. Magnificent [4128, 3270]. Magnitude [3135, 3926]. Magrab [4071]. Mahdaliza [3504]. Maillé [3503]. maintenance [687]. Maiti [4169]. Maitland [4269]. major [1068]. Majorana [4113, 2359]. Majumdar [3746]. Make [4407, 3869, 2962, 3949]. Making [3731, 1918, 4411, 988]. Maksimenko [3714]. Malament [2911]. Malcolm [3089, 3168, 2738, 3612, 3508]. Malin [2820, 3163]. Malinovsky [2678]. Malkan [4123]. Mallik [2379]. Mallow [2947]. Malyarenko [4061]. Man [654, 3106, 4397, 4400, 1672, 662, 1510, 1535, 3810, 1570, 1703, 3532, 1883, 1957, 1588, 4047]. man-made [1703]. management [925, 3514]. Managing [4192]. Manasa [3177]. Manchester [291, 221, 148, 232, 2880]. Manfred [2522, 3380, 2905]. Mangano [3074]. Manhattan [4171]. Manifestations [2271]. manifolds [3480]. Manipulating [2847]. manipulation [4103, 2362]. Manjit [2558]. Mann [3493, 3787, 2595]. Manning [3656]. Mansfield [3088]. Manton [3719]. Manual [3212, 3401, 3458, 2615, 3692, 2509, 3095, 3035, 2708, 3488, 3193, 2563, 2792, 2706, 2837, 2719, 3555, 3397, 2380, 2706]. Manucci [2317]. Manuel [2792, 3300]. Manufacturer [213]. Manuscript [3473]. Many [3393, 2221, 3022, 3840, 2508, 2962, 3090, 3311, 1469, 3010, 3652, 3591, 1145, 3884, 3920, 2371, 3972, 4276, 4166, 12, 3540]. Many-Body [3393, 3022, 3090, 3311, 3652, 3591, 1145, 3884, 3920, 3972, 4166, 12, 3540]. Many-electron [2221]. many-particle [1469]. Maoz [3621]. mapping [631, 4015]. maps [4091]. Maragò [3634]. Marc [3133, 3408, 4269, 3490, 2400, 3568]. Marcel [3924]. Marcelo [3659]. March [3652, 51]. Marco [3207, 4160, 3031, 2741]. Marcolli [2583]. Marcovich [3433]. Marcus [2543]. Marder [2724]. Marek [3297, 3140, 2967]. Marfatia [3032]. Margaret [4058]. Maria [2746, 1844, 3664]. Mariana [2740]. Marianne [2880]. Maricel [3554]. Marina [3075]. Marinescu [2692, 2692]. Mario [3052, 2489, 2714, 4161, 3367, 3066, 3475]. Mark [3368, 2667, 662, 2851, 2843, 3237, 3302, 3344, 3553, 3436, 2675, 3024, 2474, 2518, 2940, 3607, 4083, 3414, 2938, 3281, 3624]. markets [3252]. Markos [2300]. Marlyn [2607]. Marotta [3343]. Mars [4015, 3020, 4059, 2321]. Marsh [4010]. Marshall [2409]. Martin [3498, 4061, 4143, 4209, 2696, 4224, 2839, 2467, 4122, 4218, 2760, 440, 2487, 2390]. Martine [2760]. Martinez [3194, 2565]. Martino [4031]. Martins [4304]. Marvel [2542]. Masaki [3630]. Masanori [2945]. Masers [3203, 68, 236, 258, 3475]. Mashoon [4043]. mask [2097]. Maskow [4269]. mass [1020, 1178, 3801, 1944, 2199, 3174, 1734, 3763]. Masses [170, 1152, 2071]. Massey [1205]. Massimo [3245, 4208]. Massive [3039, 2564]. Massively [1591]. Massoud [3068]. Master [180, 221, 2455, 206, 26, 98, 4265]. masterpiece [3270]. masters [159, 2412]. Masud [3240]. Masujima [2523]. matching [3853]. Material [3015, 2749, 4173, 2384]. Materials [3384, 2792, 4371, 4373, 2794, 4379, 2932, 2771, 2395, 2880, 3147, 3068, 3767, 410, 754, 3974, 1007, 1049, 420, 428, 618, 1276, 409, 935, 505, 450, 4198, 4252, 3809, 4204, 4203, 2007, 1898, 4250, 2059, 434, 961, 2096, 3992, 3358, 1254, 887, 950, 506, 683, 2487, 3666, 4094, 1464, 3799, 2830, 2858, 2494, 2394, 2300, 2410, 3255, 2319, 4183, 3647, 3617, 4021, 4156, 4257, 4309, 4046, 3607, 4266, 3535, 4215, 3450, 3753, 3452]. Mathematica [3438, 4382, 3960]. Mathematical [2491, 4061, 3212, 2687, 3421, 2523, 2973, 2518, 3304, 2935, 2782, 3109, 2632, 3885, 3830, 4281, 3952, 4201, 4218, 4070, 1988, 3046, 4318, 2681, 3172, 3995, 2316, 3595]. Mathematica^{T M} [3343]. Mathematician [3100]. mathematicians [2518]. Mathematics [109, 3193, 2794, 2517, 2451, 27, 167, 1753, 2427, 2543, 2117, 3170, 4219, 4077, 3295, 3994, 3849, 3217, 3924]. Mathur [2348, 3683]. Matilde [2583]. MATLAB [3541, 4022, 4362, 4375, 3988, 3297, 3140, 3691, 3559, 3314]. Matolyak [3315]. matrix [3770]. Matter [3122, 3393, 3124, 3261, 3167, 3279, 2954, 2601, 2850, 3268, 2714, 2616, 2971, 2664, 2472, 2432, 2924, 1647, 3380, 2401, 3136, 4402, 2450, 3763, 1137, 3785, 1402, 1487, 1862, 3491, 3873, 1011, 3919, 1765, 3659, 1637, 1879, 3547, 1774, 2961, 1189, 1145, 1080, 4001, 1291, 3060, 1992, 690, 2083, 3982, 2046, 3432, 1960, 4121, 4075, 2487, 3718, 3879, 3880, 3836, 4166, 321, 486, 1392, 3221, 3242, 3684, 3981, 84, 1414, 4413, 3029, 3245, 3460, 2355, 3206, 2389, 3197, 2349, 2630, 4106, 3557, 4354, 4320, 3505, 3675, 19, 3396, 3915, 3712].

Matter-antimatter [1647]. Matters [2438]. Matthew [4317, 3275, 3823, 3316]. Matthieu [4150]. Matveev [3218]. Maudlin [2981]. Maupuech [3788]. Maurizio [3997]. Mauro [3282]. Maury [3208]. Max [965]. Maxim [3311]. Maxime [3659]. Maxwell [3840, 3075, 3281, 3514, 1698, 1259, 2219, 3612]. May [195, 256, 262, 276, 365, 261, 214]. maybe [3295]. Mazure [2929]. Mazzucato [3642]. McAllister [3583]. McCall [2710]. McCartney [3281]. McCauley [3256]. McCluskey [4317]. McComb [3376]. McCracken [3099]. McDonnell [2419]. McGill [196]. McGillicuddy [3656]. McGloin [2598]. McLennan [2826]. McRae [3322]. McSween [2613]. McWilliams [3025]. Mean [2715, 3372, 3672]. mean-field [3672]. Meaning [3320, 2534, 1671, 3987, 3811, 3473]. means [141]. Measure [3138, 3643, 4230, 3590]. Measured [2462]. Measurement [2545, 840, 4380, 3374, 3447, 2557, 2804, 2530, 468, 3881, 778, 874, 2174, 395, 3328, 2176, 960, 3531, 2482, 378, 1743, 2055, 1480, 1442, 3407, 3710, 3112, 3407]. Measurements [2648, 1671, 3207, 3245, 2670, 3961, 817, 2338, 438, 3818, 4111]. measures [2281]. Measuring [609, 445, 4051, 1793, 2266, 1925, 1658, 4102, 3049, 2265]. Mechanic [2518, 4296]. Mechanical [505, 398, 2213, 1290, 3249, 873, 4280, 1793, 2266, 2265, 3798, 1464, 2900]. Mechanics [2781, 2560, 2490, 3320, 3038, 2801, 2995, 4362, 3296, 3493, 3076, 2790, 2992, 2687, 2845, 3047, 2726, 4381, 3233, 3024, 3034, 3235, 3114, 2713, 3311, 3086, 3214, 4394, 2874, 2760, 3172, 3949, 1553, 2391, 2854, 3132, 2705, 3096, 2410, 2416, 2579, 2947, 3178, 2832, 3470, 1409, 1437, 4217, 4014, 4005, 4069, 3663, 3657, 3658, 360, 370, 2090, 3820, 4178, 1839, 3557, 875, 3513, 3999, 4196, 4187, 4319, 3906, 3510, 1171, 1460, 4161, 539, 2148, 1693, 3546, 1303, 4218, 4022, 1581, 4351, 3863, 1955, 4129, 4188, 529, 4297, 4308, 4086, 3786, 907, 3798, 3857, 4145, 4175, 1464]. mechanics [3685, 3600, 4129, 2502, 2829, 2936, 3089, 2823, 2845, 875, 2601, 3205, 3341, 2348, 3046, 3421, 2365, 3401, 3210, 2351, 2864, 3000, 3045, 3282, 2677, 2676, 2418, 2759, 3285, 2951, 2952, 2488, 2492, 2454, 2665, 3260, 4268, 3750, 3585, 3947, 3538, 3561, 3734, 3811, 4066, 4332, 4231, 3005, 2645, 3943, 4134, 3980, 3843, 3887]. mechanism [3249, 3225]. Mechanisms [3154, 1341, 744, 665, 2208, 4288, 4111]. Mechoulam [2912]. Medal [605, 2582]. Meder [4323]. Media [2820, 3233, 3213, 3402, 3988, 4375, 1923, 2659, 3002]. mediated [1309, 2285]. Mediavilla [2779]. Medical [3201, 2842, 2396, 2464, 144, 3119, 65, 1048, 1721, 886, 4165, 1316, 797, 1973, 3951]. Medicine [2918, 3768, 2533, 3058, 4343, 3832]. medieval [3326]. Medikonda [3177]. meditations [4270]. Mee [4082, 3719]. meeting [75, 147, 180, 221, 289, 341, 48, 26, 98, 76]. meetings [163]. Meets [2472, 2534]. Megascience [2356]. Meglicki [2399]. Mehlig [4197]. Mehran [2497]. Meirovitch [4217]. Meisel [3438]. Mekkaden [2750]. mekometer [166]. Melia [2562]. Melissinos [3017]. Melo [4119, 2681, 2316]. Melting [2021, 1420, 608]. Melville [4206]. members [397]. membranes [3900, 1841, 1240]. membranes- [1841]. memoir [4117, 4132]. memoirs [811]. Memorial [2912, 2822, 1569, 3976]. memory [2059]. memristors [4227]. MEMS [3148]. Mendel [2747]. Mendelssohn [1034]. Meng [4333, 4337]. Menon [2597]. Merches [3554]. Mercury [4048]. Meredith [4275]. merely [1885]. Merging [2479]. Meron [3410]. Merrill [3085]. Merritt [3479]. Mes [3208]. meson [4173, 1978]. mesons [1995]. Mesoscale [2061, 3147, 2114, 3211]. Mesoscale-Convective [3211]. Mesoscopic [2806, 2605, 1712, 2048]. Mess [3484]. MESSENGER [4048]. messengers [4143]. Messier [3427, 4345, 4392]. Mestel [2907]. Mestre [4225]. metabolism [4179]. Metaheuristic [2784]. Metal [2032, 3275, 697, 1186, 2209, 777, 1641, 4272, 3530, 1385, 515, 731, 217, 1511]. Metal-insulator [697]. Metal-nanoparticle [3275]. metal-semiconductor [3530]. Metallic [990, 651, 677, 58, 1364]. metallurgy [584]. Metals [3083, 314, 824, 640, 913, 452, 1143, 788, 500, 131, 926, 320, 378, 638, 1431, 570, 483, 1210, 161, 174, 182, 190, 200, 2909]. Metamaterials [2791, 2817, 2458, 2412]. metaphysical [2090]. Metaphysics [4416, 3425]. Metastable [2128, 1105, 1898]. metasurface [4193]. Metcalf [3598]. Meteorite [2672]. meteorites [3648, 2906]. Meteoroids [4058]. meteorological [413]. meteorology [864, 1314, 2452, 337, 90]. meteors [4058]. Method [3909, 2347, 3093, 2549, 518, 532, 3819, 385, 736, 3219, 3774]. methodologies [4219]. methodology [3649]. Methods [2757, 3212, 2696, 3118, 3181, 2724, 2523, 3189, 2973, 2395, 3304, 2837, 2634, 3109, 3183, 2711, 2557, 2492, 3082, 2598, 4078, 3830, 1778, 3927, 3873, 2193, 3644, 482, 3900, 1276, 3809, 1886, 300, 4201, 3995, 4070, 3483, 3595, 3814, 3050, 2487, 2835, 2860, 3994, 3527, 4122, 2659, 3832, 3961, 3895]. metre [183]. metric [390]. Metrology [3105, 3145, 2038, 2217, 2446, 4193, 1610, 4203]. meV [2893, 4417]. Mézard [2400]. MHD [3233, 3519]. Miake [2313]. Miao [3542]. Michael [2463, 4293, 3663, 3105, 4295, 2702, 4098, 3113, 3459, 3783, 2724, 4114, 3234, 3106, 2760, 4080, 4306, 4166, 4253, 3575, 2938, 3151, 3713, 2536, 2832, 3443, 3624, 2650]. Michel [3370, 3287, 3374, 3446, 3666, 3346, 4265]. Michele [3179]. Michelson [260, 3997]. Michio [2523]. Michler [2344]. Micro [3252, 3942, 2733, 1249, 3611]. micro-optics [3611]. micro-organisms [1249]. microanalysis [1120]. microbe [3600]. Microcavities [3788]. microcavity [1922]. Microdroplets [2915]. Microfabricated [2640]. micrographs [615]. Micromagnetism [2522]. Microparticles [2787]. microphysics [3495, 3610]. Microrheology [3813]. microring [3743]. microscale [4087]. microscope [476, 845]. Microscopic [2856, 2123, 492, 2416]. Microscopy [1131, 2834, 4079, 2344, 4287, 1045, 1053, 2076, 2572, 3418, 3608, 1203, 91, 502, 2362, 1120, 1502, 217, 557, 1563, 1814, 3327, 2173, 2624, 4224, 3622]. Microseisms [431]. Microsoft [4353]. Microstructural [2410]. Microstructure [2522]. microtron [104]. Microwave [1217, 1225, 2292, 2126, 508, 1811, 4157, 166, 3434, 4405, 3951]. microwaves [204, 210, 2048]. mid [2220, 3454]. mid-infrared [2220]. Mid-Ocean [3454]. Middle [3044, 2032]. Miele [3074]. Miftahof [4241]. Miguel [4324, 3185]. Mike [4016, 4098, 3614]. Mikhail [4130, 2902]. Mikhailov [2414]. Miklós [3247]. Milani [2526]. Milano [3119]. Milburn [3567, 2530]. Miles [2598]. Milky [3804, 3781]. millennia [2278]. Millennium [3215, 4133]. Miller [2874]. Millimetre [1233]. million [3797]. millisecond [1559]. Mills [2948, 56]. Milner [3754]. Milonni [4000]. Milutin [3126]. Mimkes [3248]. Mind [3472, 4367, 4149, 2408, 3690, 4320, 2350, 3167, 3403, 2917, 2768, 3637, 4149, 3406]. Minds [3378]. mineralogists [2472]. Mineralogy [2472]. minerals [493]. mines [2408]. Minhaeng [2555]. Miniatura [2846, 3659]. Minimal [2161]. mining [3577]. miniscule [1358]. Minoru [4281, 2982]. Minos [2200]. Miret [2979]. Miret-Artés [2979]. misconceptions [4270]. misconstrued [3654]. Mishra [2988]. missed [1905]. Missing [2354]. mission [3687, 4136]. Mitescu [3500]. mitigation [1067]. Mitin [2377]. Mitschke [2719]. Mittal [2712]. Mittemeijer [3050]. mixtures [595, 1942]. MKS [374]. Mlodinow [4117]. mobi [4281, 4353]. mobility [1107]. Mode [3202]. Mode-coupling [3202]. Model [988, 2744, 2322, 3048, 2631, 1913, 1412, 1049, 4177, 3769, 1192, 2215, 1988, 202, 4366, 1804, 3731, 3316, 2595, 3783, 3430, 3847]. Model-making [988]. Modeling [2794, 2660, 2354, 2782, 2556, 2712, 4292, 3936]. Modelling [2123, 3450, 2762, 3033, 3147, 2807, 3067, 2347, 2061, 4373, 4198, 3895, 3686, 711, 2223, 3803, 1684]. Models [3142, 2816, 3070, 3082, 1091, 3652, 2408, 2343, 3589, 1278, 4111, 2554, 2816, 2689, 3539, 3562, 3927, 3867]. Modern [3393, 522, 2615, 2742, 423, 1555, 3144, 3092, 2937, 4419, 2533, 4190, 1542, 3280, 1617, 3108, 2299, 2465, 2855, 3090, 378, 3063, 2705, 3471, 3441, 2951, 1431, 1610, 2669, 2319, 1263, 2264, 254, 4232, 4032, 3945, 578, 4072, 4376, 3671, 2201, 2519, 917, 3525, 2263, 4200, 3905, 3831, 4129, 288, 120, 4176, 4184, 3731, 4129, 2502, 2346, 3237, 4134, 4001, 2706, 3050, 4097, 2818, 3166, 3063, 3183, 3864, 3404, 3890, 4147, 3110]. modernity [4100]. modes [3175]. modified [2181, 4108]. modifying [2463]. modular [3358]. modulation [3812]. Moe [3075]. MoEDAL [3570]. Moggi [3648]. Mohammad [3073, 3098]. Mohapatra [4155, 2322]. Mohsen [2677]. Moiseyev [2759]. Molecular [2883, 2788, 675, 2927, 2205, 2991, 2, 1475, 1503, 3446, 2062, 3065, 1323, 1635, 1913, 1379, 152, 4298, 3223, 3706, 2963, 761, 3359, 2129, 2408, 3563, 470, 3832, 4112, 4245, 4260, 1817, 202, 2829, 3940, 3348]. molecule [4110, 2108, 917]. Molecules [4362, 2749, 660, 3446, 2745, 1659, 50, 4118, 1657, 1879, 1994, 1112, 989, 2096, 3563, 1065, 1151, 3487, 4022, 4194, 3964, 4260, 893, 3852, 3598, 3344, 4110, 3631, 555, 4116]. Molle [3666]. Molloy [2598]. moment [664, 781, 2275]. Moments [2347, 790, 1948, 2694]. Momentum [2374, 3087, 881, 414, 1670, 2629]. Mompart [3047]. MOND [2181]. Mondal [3608]. money [1918]. monitoring [3331]. Monk [3102]. monodromy [3223]. Monograph [4079, 2921, 4078, 4060, 4106, 2686, 2883, 3157, 3395, 3368, 2462, 2661, 2786, 2756, 2667, 2374, 4241, 2560, 2555, 3022, 3506, 2897, 2648, 2583, 2911, 3241, 2851, 2659, 2593, 2949, 3011, 4254, 2335, 2334, 2522, 2715, 2774, 3029, 3124, 3313, 4056, 4206, 2411, 3376, 4231, 3384, 2350, 2545, 2495, 2856, 3010, 4069, 4191, 2293, 2412, 2757, 3038, 3095, 3247, 4307, 4061, 3442, 2823, 2939, 3373, 4209, 2673, 2838, 2836, 3194, 3179, 3353, 3408, 3499, 4272, 2428, 2696, 2775, 3207, 2504, 2621, 3152, 3168, 2735, 2872, 3502, 4155, 4248, 2906, 2342, 2421, 2847, 3105, 3245, 3165, 4327, 2999, 4348]. monograph [2561, 2845, 2806, 2930, 4198, 4163, 2529, 2625, 2537, 2777, 2794, 2776, 3118, 2965, 3279, 3450, 4257, 4259, 4309, 4295, 2702, 4323, 2429, 2660, 2800, 2954, 3204, 2601, 2887, 2979, 3205, 3184, 3007, 3057, 3044, 3074, 3016, 2996, 3268, 3266, 3394, 3364, 3422, 4346, 2289, 2313, 2457, 2597, 2489, 2674, 2566, 2609, 2528, 2816, 2918, 3013, 3099, 4187, 4283, 2538, 2709, 2783, 3131, 3244, 3196, 3369, 3270, 3453, 3431, 3374, 2562, 3034, 3071, 3169, 2431, 2524, 2716, 2891, 3363, 3399, 3510, 3421, 3497, 2353, 3382, 2703, 2808, 3354, 4107, 4336, 2539, 2351, 2461, 2474, 2354, 2476, 2438, 2616, 2644]. monograph [2493, 2525, 2902, 3263, 3283, 3187, 3182, 3340, 3423, 3454, 3460, 3026, 3003, 2680, 2692, 2797, 3162, 2738, 3078, 4158, 2594, 2899, 2773, 3027, 2395, 2427, 2543, 3084, 2998, 4200, 3430, 3125, 3218, 3310, 4164, 2681, 2744, 2869, 3070, 2652, 3276, 3103, 2295, 2434, 2778, 4197, 3209, 3213, 2815, 3094, 3332, 2924, 2880, 3172, 3409, 4054, 4053, 4080, 4234, 2393, 2974, 3413, 3490, 2521, 2722, 3402, 2389, 2410, 2546, 2416, 2419, 2695, 2510, 2833, 2711, 3414, 3380, 4255, 4219, 4301, 4276, 4350, 2759, 3346, 2717, 2557, 2608, 2315, 2316, 2376, 2331, 2378, 2592, 2925, 2720, 2804]. monograph [2879, 2888, 2805, 3059, 3146, 3319, 3362, 3339, 3404, 3391, 3410, 3482, 3441, 3481, 3434, 3483, 3439, 3504, 3433, 3503, 4065, 4064, 4220, 2723, 2468, 2501, 2540, 2530, 2793, 2758, 2424, 2398, 2400, 2368, 2451, 2369, 2423, 2619, 3217, 3278, 3508, 2503, 2417, 3136, 2321, 2618, 2826, 2968, 4059, 2683, 2831, 2894, 3032, 2599, 2665, 2470, 2613, 2591, 2559, 2663, 2802, 2956, 2721, 2957, 2710, 2736, 2827, 2832, 2928, 3134, 3069, 3110, 3085, 3042, 3338, 3306, 3372, 3443, 3475, 3474, 3495, 4116, 4067, 4215, 4233, 4335, 3397, 3300, 4111, 4168, 2547, 2920, 2882, 2852, 2586, 4243, 2704, 2450]. monograph/essay [2710]. monograph/general [2906]. Monographs [3116, 4061, 3421]. monolayers [1186]. monopoles [2761]. Monsoon [2376]. Montanari [2400]. Montani [2741]. Monte [3927, 2860]. Montwill [2411, 2933]. Moo [3318, 3609]. Moo-Young [3318, 3609]. Moon [1974, 672, 2299, 3066, 3870, 4316, 840, 435]. Moore [2463, 3043]. Morette [2301]. Morf [3015]. Morfill [3276, 2652]. Morimoto [2782]. Morin [3115]. Morishita [3404]. Morley [3997]. Morowitz [3597]. Morrison [3478]. Morton [2980]. Moser [4336]. Moshe [3280, 2951]. Moshkalev [3059]. Mössbauer [760, 1881, 490]. most [3739]. Motion [2867, 1581, 339, 355, 3328, 59, 1476, 1151, 385, 2088, 4172, 1737, 3100]. motion-robust [3328]. Motives [2583]. Motohiro [3544]. Motoichi [3353]. motor [2304, 2123]. motors [2131, 3584]. Mott [3255, 1757]. Mou [3720]. Mou-Hsiung [3720]. movement [4219, 611, 2417]. moves [1914]. Moving [2123, 1923, 457, 61, 1658]. Mozer [3677]. Mr [2864, 2908]. MRI [3088]. Mrs. [2491]. MSc [2743, 2739]. Mu [3560]. Mu-Lin [3560]. much [919, 981]. Mudry [3491]. Muhammad [3073, 4254]. Mukhanov [2425]. Mukhopadhyay [4251]. Mullan [4305, 2559]. Mullaney [2505]. Mullard [255]. Muller [4351, 2774, 2760, 2995, 3310, 2391, 2510]. Muller-Kirsten [4351, 2995, 2391, 2510]. Mullin [3694]. Mullylä [2804]. multi [4242]. multi-purpose [4242]. Multidimensional [2624]. multidisciplinary [3572, 4306]. Multilayered [1537, 3672]. multilayers [1666, 1515]. Multiphase [3402, 4331]. multiphonon [1089]. Multiphoton [1379, 934]. Multiple [989]. Multistep [1400]. Multiverse [2459, 4127, 3469, 3521]. MULTIverses [2814]. Multiwavelength [3049]. Muni [3786]. Munk [2412]. Muon [1479, 781]. Muons [3015, 2094, 1862]. Murdoch [3205]. Murugan [2984]. Musazzi [3646]. music [4286, 1590, 4244, 4352, 3357, 2462, 2520, 4307]. Musical [969, 1887, 399]. musicologists [2474]. musicum [2553]. Musings [3100]. Muslim [2870]. MXene [4227]. MXene-based [4227]. my [2034, 3935, 3702, 965]. Myers [3618, 2980]. myPrint [4281]. Myridis [3432]. Myron [3139]. mysteries [3913, 4325, 2576, 3594, 4059]. mysterious [3650]. Mystery [2321, 4082, 239]. mystique [1995]. Myth [2321, 1672].

N [2788, 3719, 4082, 4241, 2949, 3716, 3644, 2545, 3412, 4221, 2673, 3628, 3652, 4155, 2806, 3054, 4024, 2850, 4040, 4042, 4091, 4338, 2609, 3810, 2438, 3819, 4019, 4249, 3526, 2419, 4334, 2322, 2310, 4332, 2317, 3255, 2259, 2807, 2665, 3039, 2547, 1146]. N.W.2 [89]. N.Y. [137]. N00N [2217]. Nadeau [2837]. Nadi [3063]. Nag [4251]. Nahar [2788]. Nahin [2491]. Naim [3645, 3654, 2619]. Nakariakov [3007]. naked [3477]. Nanda [4272]. Nandini [3230]. Nano [3353, 3024, 3293, 2378, 2128, 2733, 3537, 3036, 3584, 3714]. Nano-Electro-Optics [2378]. nano-electromagnetics [3714]. nano-grenades [2128]. Nano-Mechanics [3024]. nano-motors [3584]. Nano-optics [3353, 3537, 3036]. Nano-optomechanics [3293]. Nanoantenna [3408]. nanoapplications [3819]. Nanochemistry [2529]. Nanocomposite [2771]. Nanocomposites [2767]. Nanocomputing [2777]. Nanocrystals [4362, 4022]. nanodevices [4405, 3434]. Nanoelectromechanical [2136]. Nanoelectronics [3301, 2785, 2377, 2628, 3666]. Nanofabrication [2799, 3059]. nanofibers [4235]. Nanoguides [2879]. Nanoimaging [2720]. Nanomagnets [1896, 4298]. Nanomaterials [3051, 2836, 3265, 4081, 4242, 2988, 2800]. Nanomechanics [2925]. Nanomedicine [2836]. Nanometrology [2230, 2919]. nanoparticle [3275, 2224]. Nanoparticles [2880, 4180, 3293]. nanophotonics [3772, 3353, 2563, 3977]. nanoprobes [3629]. Nanosatellites [3571]. Nanoscale [3047, 2773, 2834, 4405, 2304, 3961, 2233, 3352, 3760, 2015, 3122, 3433, 2503, 4162, 3712, 3434]. Nanoscience [2749, 2777, 3488, 2234, 3674, 2402]. Nanoscopy [2624]. Nanosensors [2836, 3179, 3062, 2853]. Nanostructures [3003, 3696, 4260, 3672, 2787, 3530, 3742, 3898]. nanosystems [4377, 2836, 3439]. Nanotechnology [1905, 2798, 3177, 3982, 4199, 2795, 3493, 2777, 3488, 3674, 4254, 3194, 4302]. Nanotube [2838, 1841]. nanotubes [1956, 2332, 3033, 3544, 3501]. nanowires [4220, 3674]. Naoki [4220]. Nappi [3204]. Narayanan [3543, 2750]. Narayani [2712]. Narlikar [2343, 2916, 2531, 3474]. narrow [679]. NASA [4136, 1098]. Nastase [4002, 3592]. Natale [3207]. National [145, 468, 3206, 108, 142, 364, 365, 195, 219, 256, 262, 265, 406, 242]. Natural [2333, 2451, 3379, 3766, 3520, 932, 1841, 880, 1570, 577, 4243, 3803]. Nature [3290, 2056, 3425, 4154, 2228, 4416, 4148, 3163, 3375, 4159, 4160, 3243, 3414, 4208, 4171, 2558, 4145, 3834, 4106, 2099, 3878, 2081, 487, 993, 4027, 732, 1926, 718, 3597, 1123, 1111, 3543, 124, 1073, 4256, 3496, 4321, 4276, 4004, 3163, 3777]. Navarro [4100]. Navier [3573]. navigation [340, 337]. navigational [3251]. Nazarov [3076, 2402]. near [1430, 3159, 1326, 1028]. near-field [3159]. Nearest [3400]. Necessity [3399]. Need [4369, 1014, 1409, 3459, 2886]. Needs [4396, 3752, 83, 3871, 3333]. Negative [2394, 542, 2093, 2270, 229, 1107, 2269]. Neiner [2501]. Neiner/Physics [2501]. Nellis [2969]. Nelson [380, 291, 4292, 2891]. Nematic [861]. Nenad [4322, 4326, 3148]. Nenadic [3148]. Neoncinegraphy [539]. Neptunes [2190]. net [2735]. network [3503]. Networks [3471, 2113, 2762, 3789, 3665, 4197, 1248, 3814, 2352, 2657, 3031]. Netzer [3266]. Neuenschwander [3574, 3485, 2683]. Neunenschwander [3817]. neural [2325, 4197, 1248, 2285]. neuroimaging [4236]. neuromorphic [4227]. neutral [123, 623]. Neutrino [1152, 2199, 122, 1734, 2812, 4155, 2120, 1160, 3074, 2612, 2877, 2701]. Neutrinos [762, 2322, 3136, 1523, 2597, 3032]. Neutron [1137, 2955, 1187, 3321, 386, 401, 1530, 1303, 2633, 1149, 3236, 1512, 3835, 749, 283, 664, 924, 2355, 1076, 4288, 71, 3590, 1414, 1650, 1722, 2674, 3510, 3863, 3267]. Neutrons [2355, 3015, 566, 21, 32, 1039, 830]. never [2772]. Nevill [1757, 2769]. newcomer [2394]. Newman [4027, 3096, 3595]. Newton [3093, 3270, 2864, 2983, 3575, 3381, 1643, 1688, 1834, 1273, 1150, 1272, 1190, 1864, 2437, 3878, 4027]. Newtonian [2911, 2181, 360, 370]. Next [2626, 4028, 4316, 3174, 3590]. Ng [2943]. Nguyen [4302]. Ni [3585]. Nian [2756]. Nicholas [4275, 3113, 2920]. Nicholls [3699]. Nick [3274, 2581]. Nicolas [3772]. Nicole [4325, 2880]. Nicosia [3814]. Niels [3425, 1412, 4073, 3048]. Nielsen [2656]. Nigel [865]. Night [2439, 3151]. Niklas [2910]. Nikolay [3460]. Nikolova [2429]. Nikonov [3218]. Nikos [3051]. Nikuni [2580]. Niles [3483]. Nils [2740]. Nimrod [2759]. NINA [437]. Nineteenth [2887]. Nino [4154]. ninth [3647, 415]. Nipoti [4146]. Nir [3110]. Nishimori [2729]. Nishinaga [3626]. NIST [2632]. Nitant [4219]. Nittler [4048]. NJ [4132, 4120, 4273, 3945, 4355, 4235, 4274, 4292, 4200, 4205, 3995, 4301, 4310, 4167, 3337, 4332]. NMR [1226, 3373]. No [159, 554, 4418, 3419, 3129, 623, 637, 4317, 2057, 4023]. No-frills [4317]. Noakes [1497]. Nobel [2870, 965, 3432, 4200, 4151, 4205, 4414, 2082, 945, 3337, 2082, 3337]. Noble [545]. Nobuhiko [3439]. Noce [4147]. Noether [3817, 2683]. Nogué [2779, 2779]. Noise [1474, 2466, 2733, 2016, 1807, 2116, 2833]. Noise-Induced [2833]. Noisy [2331]. Noll [3531]. Nolte [3471]. Non [3333, 2290, 4278, 2734, 88, 3125, 2824, 626, 1855, 3255, 2369, 1691, 310, 3318, 4211, 4286, 2989, 2296, 1754, 2851, 2843, 677, 2912, 3056, 1241, 1247, 3349, 3137, 3095, 3035, 4307, 3150, 916, 2838, 3104, 2988, 3444, 3152, 4155, 2406, 1229, 2684, 3496, 2542, 3407, 3262, 3163, 2033, 3121, 3128, 3018, 3210, 2473, 3415, 3445, 4186, 1254, 3283, 3423, 2688, 2916, 2748, 3206, 506, 4200, 4205, 3234, 2908, 2515, 2778, 3066, 2443, 4311, 3014, 2717, 4308, 3225, 3461, 2582, 2913, 2789, 2826, 2986, 2219, 3178, 3134, 2048, 2964, 2160, 2650, 2610, 2769, 4254, 2057, 4355, 4303, 4338, 4342]. non [4340, 4320, 3139, 3038, 3370, 3843, 2622, 2759, 2540]. Non-additive [626]. non-classical [2048]. Non-collinear [4278]. non-computable [2033]. Non-crystalline [3255, 1254, 506]. non-destructive [2515]. Non-diffracting [3370]. Non-Equilibrium [2290, 2824, 2369, 2406, 2219, 3038]. non-experts [3206]. Non-Fermi [1855]. Non-Hermitian [2759]. Non-linear [2734, 88, 310]. Non-local [1691]. non-locality [2160]. non-metallic [677]. Non-Perturbative [3125, 2622]. non-relativistic [4308, 3843, 2540]. non-Rydberg [916]. non-science [2473]. Non-Scientist [3139]. non-scientists [1754]. non-specialist [2296, 1241, 1247, 2838, 1229, 2443, 2057]. Non-specialists [3333, 3318, 4211, 4286, 2989, 2851, 2843, 2912, 3056, 3349, 3137, 3095, 3035, 4307, 3150, 3104, 2988, 3444, 3152, 4155, 2684, 3496, 2542, 3407, 3262, 3163, 3121, 3128, 3018, 3210, 3415, 4186, 3283, 3423, 2688, 2916, 2748, 4200, 4205, 3234, 2908, 2778, 3066, 4311, 3014, 2717, 3225, 3461, 2582, 2913, 2789, 2826, 2986, 3178, 3134, 2964, 2650, 2610, 2769, 4254, 4355, 4303, 4338, 4342, 4340, 4320]. nonadditive [3221]. Nonautonomous [3340]. Noncircular [2766]. Nonequilibrium [3132, 2416, 3820, 3090, 3831]. Nonlinear [2067, 3368, 2614, 2840, 1887, 2541, 3024, 2571, 2416, 2752, 2015, 3796, 2573, 2733, 4281, 1474, 3770, 3466, 1635, 3812, 4088, 3619, 3343, 3678, 2711, 2592, 3391, 3410, 3537, 3566]. nonlinearity [1326, 2208]. Nonlocal [4043]. Nonperturbative [2609]. nonrelativistic [3704]. nonsense [3848]. nonspecialist [2191]. nonspecialists [3329]. Norbert [4158]. Nordlund [2844]. North [2299]. Norton [2620]. Norval [4256]. Norwich [147]. note [545]. Notebook [3343]. Notes [2886, 583, 2760, 2846, 3772, 3491, 295, 3836]. Nothing [1263, 2264, 4194, 2263, 2426]. Notice [2111, 1962, 2139, 2324]. Notices [1868, 947, 957, 974, 986, 996, 1004, 1017, 1032, 1061, 1087, 1320, 1330, 1345, 1353, 1361, 1368, 1376, 1383, 1397, 1447, 1492, 1499, 1612, 1860, 1877, 1890, 1909, 1930, 1940, 1986, 2091, 2105, 2118, 2144, 2323, 2477]. Notions [1002]. Nottingham} [554, 338]. Nouredine [2579]. Novel [3033, 2096, 4258, 3629, 3187, 3402]. November [1136, 50]. Novotny [3036]. Noz [4067]. NPL [166]. NTT [1516]. nucléaire [366]. Nuclear [1578, 145, 1735, 1155, 3875, 1501, 1276, 1934, 925, 4369, 1879, 872, 1624, 1866, 1080, 1381, 1573, 1665, 359, 369, 1175, 2674, 2816, 2706, 2035, 1104, 3416, 4315, 1885, 972, 499, 2445, 129, 738, 4400, 2310, 4401, 2507, 1527, 827, 1336, 834, 498, 1913, 4010, 143, 4211, 631, 4343, 1206, 1723, 4030, 1068, 319, 51, 1400, 1486, 2024, 2384, 3459, 3754, 193, 4318, 3810, 122, 1886, 839, 2027, 3206, 4159, 820, 853, 1421, 2236, 3228, 4174, 3829, 4168, 76, 3322, 4204, 1069, 2566, 2918, 2330, 3206, 2390, 4207]. nucleation [999, 1008]. Nuclei [927, 319, 2923, 1641, 21, 32, 1879, 1178, 1348, 1465, 2514, 1436, 170, 1958, 3266, 3094, 3479]. nucleon [762]. nucleosynthesis [4174, 2620, 3586]. nucleus [3, 959, 129]. number [2337, 1422, 1427, 4029]. Numberland [2496]. numbers [3704, 2373]. Numerical [2781, 4330, 2757, 1310, 2976, 3185, 2956, 4108, 2223, 2794, 2649, 2653, 3483]. Nunes [3122, 2566, 3712]. nutshell [3625, 1100, 3972, 3155, 2978, 2666, 3769, 4131, 3621]. NY [4206, 3459, 3421]. Nyiri [3218, 2822].

O [2781, 2421, 1092, 3184, 2674, 2387, 4180, 3303, 3492, 2365, 2682, 2245, 2946, 2744, 2972, 2500, 4096, 2468, 2492, 2736, 3798, 4058, 3594]. O.E.E.C. [69]. O.M [1136]. Obituary [796, 899, 976, 2324]. object [4267]. object-oriented [4267]. Objectives [237]. Objects [4392, 4345, 457, 4238, 3427, 2441, 3668]. obscure [3525]. observation [845, 2343, 396]. Observational [954, 1749, 752, 4108, 3643, 4150, 3722, 2591]. observations [574, 604, 1714, 1472]. Observatory [255, 793, 3886]. Observer [3035, 3160]. observer-centric [3160]. Observing [1731, 1874, 3192, 1151, 3478, 572, 2439, 2302, 2299]. obtained [3938]. occasion [3681]. occurrence [1596]. occurs [623, 637]. occurs-symmetrical [637]. Ocean [1236, 3284, 1699, 423, 482, 743, 3765, 3656, 3454]. Oceanic [2484]. oceans [1595]. Ochkin [2547]. Ockendon [2488]. Octave [3541]. Odagiri [3103]. Odd [3262]. Odelin [2921]. O’Donnell [3552]. Odyssey [3516, 4030]. Oertzen [1069]. off [800]. off-centre [800]. Offbeat [4270]. Office [89, 277, 413]. Offworld [2449]. Ogilvie [4306]. OH [3203]. Ohanian [3437, 3153]. Ohlsson [2801]. Ohtsu [3353, 2378]. Ohya [2945]. Oil [808, 914, 2789]. Okubo [3409]. Okun [2989]. Olaf [3104]. Olafsen [2616]. Old [2690, 3443, 2216, 1917, 4412]. Ole [3338]. O’Leary [2512, 2590]. Oleg [4231]. Olfaction [2602]. Olindo [3278]. Oliphant [662]. Oliveira [2561]. Oliver [3618, 2845]. Olivier [2887, 3213]. Ollivier [2476]. Olmes [3623]. Olshanii [3311]. Olver [2632]. O’Meara [2439, 3192, 3668]. on-orbit [3886]. Once [2008, 1543, 1183]. Oncology [2876]. One [1672, 4399, 2316, 1292, 3953, 2021, 4155, 237, 245, 3532, 4166, 3314, 3876]. one-body [4166]. One-Dimensional [2316, 1292]. oneness [3227]. Ong [2850]. ongoing [3327]. Online [4142]. Onofrio [3634]. ontology [3811]. open [4139, 2085, 89, 195, 242, 256, 262, 265, 278, 277, 365, 334, 406, 713]. Opening [77]. operation [201]. Operational [801, 792, 1091]. operative [740, 1249]. operator [2446]. Operators [2370]. ophthalmic [4032]. Oppenheimer [1092, 3102, 1092, 4388, 4404, 3102, 2191]. Opponents [3363].

Opportunities [1905, 3267, 3418, 3331, 3465, 4186]. optic [909, 683, 3362, 1191, 3862]. Optical [54, 937, 1191, 2614, 2555, 3022, 1820, 2828, 3116, 4103, 2636, 4365, 2085, 3083, 3105, 3165, 2541, 1385, 3407, 1326, 2107, 772, 2820, 1170, 2327, 470, 3424, 1831, 2288, 2005, 3145, 2273, 1438, 2624, 1517, 258, 2725, 2378, 2804, 3434, 2470, 1922, 2486, 3629, 2217, 3328, 3159, 2446, 3496, 831, 236, 1486, 1594, 438, 3697, 1606, 2055, 1079, 1635, 3482, 3504, 3745, 1272, 4405, 2326, 1817, 1516, 2667, 2485, 4183, 2791, 4006, 2703, 3367, 2998, 3634, 3156, 2717, 2557, 2804, 3673, 3743, 2598, 2787]. Optically [818, 3293, 4236, 2783]. Optically-detected [818]. Optics [2686, 96, 4419, 3323, 2887, 3145, 2835, 2378, 2719, 2888, 3441, 2723, 310, 1634, 3326, 3953, 2067, 2128, 4202, 4232, 3290, 3619, 3772, 4139, 3365, 4000, 3353, 3499, 2233, 4193, 1979, 1399, 3130, 963, 3770, 2571, 1813, 1959, 3582, 3864, 1743, 2069, 3391, 3537, 3536, 3555, 1193, 2454, 3775, 745, 1516, 3388, 2015, 3746, 3368, 2385, 3011, 3036, 2678, 3262, 3019, 1140, 2627, 2805, 2589, 2638, 2398, 2388, 3397, 2752, 3796, 3522, 4032, 3661, 3682, 4020, 4170, 3565, 3534, 3611, 3701, 4092, 2800, 3953]. optimal [2235]. Optimisation [2784]. Optimization [2170, 2685, 2757, 2743, 2739]. option [4010]. opto [771]. opto-electronics [771]. optoacoustic [933]. Optochemical [3179]. Optoelectronic [2567, 2660]. Optoelectronics [1727, 4419, 2372, 3145, 3441, 2538]. Optogalvanic [1056]. Optomechanical [3165]. Optomechanics [3165, 3293, 3567]. Opus [2553]. Orazi [3784]. Orbit [2526, 4119, 3886, 59, 1560]. Orbital [2629, 2966]. orbits [102, 1190]. Order [2897, 4384, 2493, 1562, 782, 4149, 1129, 13, 61, 133, 4175, 2614, 3139, 3298]. ordered [935]. ordering [1441, 665]. orders [3926]. ordinary [2549]. Organic [824, 2795, 1139, 3632, 3888, 1993, 4252, 2909]. Organised [3070]. organisms [1249]. organization [358, 4]. organizations [2025]. Organized [2461, 2998]. oriented [4267]. Origin [1501, 2539, 1460, 2575, 2826, 710, 2252, 1623, 3801, 134, 2251, 3597, 218, 4280, 345, 2165, 2335, 4123]. Origins [2565, 2970, 3129, 1340, 3423, 792, 3844, 2473]. Orio [3075]. Oriols [3047]. Oriti [2431]. Orloff [2388]. Ortiz [2729]. oscillating [817]. oscillation [2120]. Oscillations [3007, 3272, 1152, 1956, 398, 3543, 3476, 3199]. Oscillators [3024, 3820, 3466, 2060]. Oscillatory [2325, 3476]. Ostoja [4061, 2410]. Ostoja-Starzewski [4061, 2410]. Otaniemi [834]. Other [2491, 4359, 2774, 3403, 2509, 3448, 2675, 3828, 3918, 1879, 2519, 2096, 3358]. Otmar [2973]. Ott [2858, 3348]. Öttinger [4072]. Otto [3482]. OUP [4348]. Our [3308, 3302, 3405, 3136, 4407, 2970, 3378, 4410, 3333, 810, 3290, 3869, 1799, 2211, 4274, 3400, 4037, 4340, 3423, 4200, 4399, 3702, 3876, 3958, 2814, 3085]. Ours [3415]. Outline [3375, 2560]. outlook [4108]. Overdamping [2179]. Overduin [4121, 2450]. overview [2296, 2486, 2184, 2054, 2268, 3206, 2231, 2326, 3792, 2921, 2267]. Owen [3651]. Owens [3013]. own [2228, 4285]. Oxford [4016, 4137, 4082, 4038, 4000, 3091, 3748, 4073, 4027, 4100, 4356, 4328, 3142, 2057, 4291, 4285, 4348, 4178, 4198, 4156, 4203, 4162, 4025, 4043, 4041, 3459, 4187, 3312, 4097, 4017, 3026, 4158, 2455, 4282, 4311, 4276, 4320, 4199, 4266, 3505, 4064, 3147, 4109, 4175, 1459, 1217, 1225, 819, 2871]. Oxfordshire [4274]. oxide [4242, 4180]. oxides [4272, 2774]. Ozima [2982]. Ozin [2529]. Ozone [178, 3520, 1788, 3520]. ozone- [1788].

P [3807, 3520, 3741, 3790, 3774, 4225, 3559, 2296, 3796, 3699, 3700, 4038, 4000, 3710, 3656, 2795, 2727, 3412, 2057, 3538, 2988, 3628, 3212, 3193, 2872, 3758, 4248, 4303, 2301, 1092, 4026, 4148, 4198, 3809, 4259, 4270, 2302, 2429, 3054, 3999, 4008, 3341, 3669, 3576, 3689, 4037, 4040, 4015, 4340, 2289, 3661, 2820, 2333, 2396, 2724, 2898, 2300, 4097, 2435, 2623, 2351, 2881, 3768, 2512, 3653, 4200, 4151, 4205, 3065, 2689, 2873, 2914, 2818, 2869, 4121, 3683, 3735, 4020, 2676, 3239, 3414, 3803, 2608, 2629, 3141, 3567, 3788, 2501, 2370, 2488, 3713, 2822, 2454, 2599, 2453, 2710]. P [2970, 3069, 3535, 3684, 4175, 3039, 3519, 3685, 2358, 2420, 2450]. P. [694, 1359, 2527, 2596]. pack [3988, 4375]. packet [1994]. packets [1518, 1541]. Packing [2334]. Padgett [2598]. Padmanabhan [2617]. Paech [3066]. Page [2435, 4201]. pages [4210, 4052, 4146, 4169, 4170]. Paget [796]. pair [4050, 376]. Pairing [2330]. pairs [2055, 1073]. Pais [1165]. Palmer [3790]. Panagiotis [4347]. Pankaj [3477]. Panofsky [2257]. Panos [4244]. Panton [3429]. Paola [3375]. Paolo [3711, 3207, 3004, 2833, 4145]. Papen [4060]. paper [3463, 133]. Paperback [4085, 4105, 4283, 4318, 4054, 4266, 4132, 4317, 4324, 4016, 4211, 4268, 4140, 4304, 4296, 4206, 3425, 4073, 4182, 4184, 4191, 4113, 4307, 4183, 4084, 4190, 4156, 4269, 4270, 4295, 4025, 4046, 4024, 4294, 4097, 4292, 3423, 4112, 4114, 4115, 4160, 4321, 4151, 4218, 4022, 4053, 4301, 4070, 4167, 3337, 4095, 4065, 4086, 4216, 4034, 4059, 4175, 4258, 4111, 2041, 2908]. paperback/hardback [4140]. Papers [291, 380, 2387, 3551, 739, 1272, 3490]. Paradigm [2403, 3279, 3359, 3078]. paradigm-shift [3078]. paradigms [1907, 2685]. Paradox [1808, 4368, 3780, 4195, 1880, 4003]. Paradoxes [4408, 1907, 1967, 3815, 2964, 3845]. paradoxical [2088]. parallel [1365, 1371, 942, 1591]. Paramagnetic [3026, 347]. Parameter [4053]. parameters [1677, 1480, 4175]. parametric [3822]. Parasuraman [3653]. paraxial [745]. Pareto [2142]. Paris [366]. Parisi [4210, 3282]. parity [2446]. Park [334, 4361, 2772]. Parker [2524, 3716, 2520]. Parliament [293, 292]. Part [2241, 1466, 1472, 3366, 2682, 175, 1365, 1371, 1241, 1247, 1045, 1053, 999, 1008, 245, 359, 369, 401, 1171, 847, 857, 558, 559, 699, 150, 1622, 1217, 1225, 1422, 1427, 21, 32, 360, 370, 132, 153, 237, 386, 204, 210, 322, 651, 11, 34, 698, 1278, 1282]. Partha [3608]. Parthasarathy [4274]. Partial [2543, 2653]. Particle [3350, 604, 3017, 3237, 3349, 4417, 3317, 2489, 1902, 3382, 1108, 2881, 3216, 1097, 2258, 2744, 2855, 1959, 2595, 3465, 299, 2390, 2631, 2758, 2388, 141, 4211, 3509, 3250, 4330, 1469, 1875, 4045, 3913, 4028, 4155, 1229, 3547, 1542, 1770, 2366, 2024, 1001, 3754, 2519, 1255, 354, 2244, 1064, 2243, 3510, 2959, 552, 2647, 1214, 3863, 3731, 3667, 3818, 3847, 1481, 3593, 1054, 1734, 2986, 2564, 2662, 2257, 3442, 4173, 4339, 4158, 3218, 3613, 2630]. particle-physics-and-nuclear-physics [4211]. Particles [113, 1318, 3381, 1488, 2771, 3382, 3015, 2744, 2938, 3191, 2569, 416, 927, 296, 2480, 951, 3801, 3787, 197, 4087, 2148, 4121, 4075, 267, 46, 2240, 1948, 1380, 478, 22, 2497, 3053, 2608, 2239, 3656, 4018]. particular [1068]. Partridge [2435]. parts [2886, 2371, 4311]. party [4192]. Pasachoff [3215, 3400]. Pascal [2920]. Paschos [2685, 2743, 2739]. Pask [3270]. Pasquale [3207]. Passion [2849]. Past [3399, 667, 2211, 1302, 4399, 2031, 3876]. Pastor [3074]. Patents [3006]. Pater [3238]. Path [1402, 2995, 2291, 3931, 4262, 3123, 3527]. Pathria [2676]. Paths [3075]. Patrick [3503, 3702]. Pattern [2485, 3360, 1422, 1427, 212]. Pattern-formation [3360]. patterning [1622]. Patterns [2116, 1782, 3462, 2016, 4395, 2259, 1199, 2260]. Paul [3480, 2491, 3425, 3095, 2672, 3373, 3092, 2678, 3283, 2464, 3529, 3129, 4168, 2103]. Pauli [2057, 4418, 356, 2057]. Pauline [3731]. Paulo [3507]. Pavel [2619]. Pavesi [2396, 3481]. pbk [4156, 4203]. PDEs [3311]. Peaceful [820, 3794]. Pearce [2419, 2376]. Pécseli [3272]. Pedrero [4032]. Peebles [2435, 4200, 4151, 4205]. Pei [3071]. Peierls [2241, 1711]. Peiponen [2804]. Pelant [2878]. Peletier [4099]. Peliti [4310]. Pellonpää [3710]. Pelton [3275]. Peña [1844]. Pendry [2269]. pendulum [2060, 2331]. Penning [2339, 737]. Penrose [3220, 3636]. pentaquarks [2159]. people [765, 1973, 3800]. perception [4165, 3653, 775, 139]. percolating [2025]. Percolation [1430, 1036, 1037, 1511]. Percolation-type [1430]. Peregrinations [3681]. Pereira [2909]. Peres [3628]. Perez [3240, 4118]. Perfect [2334, 3700, 3949]. performance [2031]. Perini [3646]. periodic [3758, 725]. Perkins [2491]. permanent [407]. permeability [1309]. permissiveness [1123]. Perot [3862, 657]. perovskite [4246]. Perovskites [2774]. Perry [4139, 3119, 3027]. person [2353]. Personal [2906, 1117, 3524]. Personalities [3128]. Perspective [3163, 3470, 2959, 801, 4013, 3470, 2979, 3049, 4253, 3508, 3831]. Perspectives [4154, 2607, 3281, 3784, 4186, 3950, 3220, 4395, 4067, 3320, 2379, 2750, 4166, 2926, 2294, 3269, 2499, 3462]. Perturbation [2539, 4110]. Perturbations [2815]. Perturbative [2609, 3125, 2831, 2622]. Pervasive [3686]. Peskov [3204]. Peter [3670, 3368, 2829, 3619, 3029, 2696, 2941, 4248, 2766, 2792, 1092, 3309, 3552, 3631, 3099, 3088, 3143, 3084, 3589, 4224, 3248, 3604, 2888, 3191, 3598, 1040]. Pethick [2436]. Petit [3500]. Petkov [4191]. Petley [2247]. Petrini [4070]. petroleum [914]. Petrov [4110, 3593]. Petry [3575]. Pfeffer [3110]. Ph.D. [2351, 2461, 2354, 2743, 2739, 2894]. Ph.D.s [3057, 3016]. Phaenomena [3044]. Phase [2819, 3296, 4371, 1824, 4379, 1366, 3230, 3466, 1429, 2729, 2728, 1737, 2195, 1204, 3539, 3644, 4329, 3820, 1805, 1189, 1964, 3920, 4098, 1529, 2482, 2362, 3812, 2209, 907, 2224, 4055, 3280, 3452]. phases [1898, 730, 3903, 1894, 3900]. Phenomena [3301, 2715, 3118, 3303, 2971, 2833, 2628, 3379, 2137, 2573, 2079, 1242, 1179, 2085, 578, 466, 1357, 4322, 1249, 2466, 302, 521, 3435, 2333, 2729, 2442, 4072, 3623, 3610]. Phenomenology [3376, 4177]. phenomenon [1912, 494, 2393]. Phil [4215]. Philip [2838, 3669, 2886, 3253, 4292, 2449, 3634, 3372]. Philippe [3002]. Phillip [3059]. Phillips [2295, 2873, 2500, 2720]. philosopher [785, 2508]. philosophers [2459]. Philosophical [2856, 3320, 2535, 2957, 179, 3914]. Philosophy [3091, 3969, 3300, 4127, 932, 1899, 2506, 2435, 3664, 2498, 2581, 3759, 2981, 4303, 2451]. Phonon [1161]. phonons [3742, 4259]. phony [2206]. phosphine [4237]. Photinos [4244]. Photo [3728]. Photo-induced [3728]. Photocathodes [2031]. Photochromism [991]. Photodetectors [2775]. Photoelectric [1372]. photoelectricity [377]. Photoelectron [692, 1295, 1204]. photoexcited [432]. Photographic [3066, 385]. photographs [2417]. photography [1176, 4144]. Photoionisation [3467]. photometry [3737]. photomultiplier [3771]. Photon [2465, 2723, 361, 2246, 989, 2381, 4160, 2245, 4152, 2132, 2055, 3338]. Photonic [2102, 2341, 3548, 2300, 3440, 2320, 1635]. Photonics [3297, 2932, 2671, 3140, 3141, 2486, 3496, 3960, 2320, 3104, 2372, 2888, 3499, 3481, 3483, 3369, 3912, 3391, 3941]. Photons [3318, 3393, 2211, 2355, 3015, 2629, 2326, 4106, 3609, 580, 3198, 3558]. Photostimulated [900]. photovoltaics [4246]. Phua [3412, 4098, 3681, 2846]. Phuan [2850]. Phyl [3038]. phylogeny [3681]. physic [3739]. Physical [97, 143, 158, 468, 554, 2335, 3320, 2853, 3212, 3193, 3444, 38, 298, 2604, 2901, 3422, 308, 14, 2353, 4124, 2623, 3026, 235, 2518, 605, 2818, 3248, 442, 2837, 3188, 2976, 142, 364, 365, 36, 2654, 2554, 3174, 106, 76, 37, 39, 3719, 687, 3828, 3554, 874, 162, 4084, 360, 370, 4148, 3808, 3755, 4274, 201, 3936, 3120, 30, 475, 4292, 3587, 4321, 3432, 4205, 2208, 4288, 4189, 195, 219, 256, 262, 265, 406, 2960, 776, 231, 79, 74, 40, 51, 177, 3809, 3205, 2746, 3950, 3231, 3500, 4077, 3170]. physically [1989]. Physicist [3167, 2867, 2532, 2311, 2508, 732, 4030, 1801, 2959, 72, 185, 3942]. Physicists [1265, 27, 41, 109, 3830, 3974, 2459, 2189, 3994, 2674, 2134, 2402, 2363, 2519, 4322, 2351, 2461, 2474, 2476, 2616, 2525, 3487, 2647, 2518, 389, 2472, 2487, 3603, 4188, 4070, 3505, 3633, 3073, 2754, 2790, 2594, 3347, 2758, 3480, 3625, 4056, 4326, 3409, 3922, 4341, 4167, 3304]. physicists/computational [2487]. Physics [2921, 150, 175, 3122, 3157, 3393, 3468, 97, 143, 151, 158, 3052, 2025, 2462, 2788, 4357, 4358, 2926, 2989, 3201, 4179, 1754, 75, 554, 3458, 3301, 2767, 2851, 3107, 2868, 307, 1365, 1371, 4414, 372, 3116, 53, 2615, 2658, 2578, 2662, 2742, 2904, 2981, 3089, 3261, 3425, 2082, 2293, 248, 3839, 2866, 2942, 2801, 1552, 4061, 3442, 4416, 4240, 3144, 4367, 2735, 2937, 2872, 3342, 2792, 3245, 968, 2533, 232, 2655, 2806, 4190, 2171, 2625, 2777, 1924, 4378, 1175, 1291, 3266, 3364, 3435, 3422, 1965, 2457, 2489, 2674, 2706, 2519, 3163, 3253, 3335, 3405, 1600, 2306, 1104, 2303, 3071]. Physics [2891, 308, 3421, 1565, 4124, 2691, 2889, 747, 3208, 3206, 4114, 2497, 2899, 2455, 4200, 4151, 4205, 4224, 117, 3299, 499, 605, 3216, 3218, 3305, 4300, 2796, 2744, 2818, 3238, 3103, 2460, 3240, 1960, 2332, 2472, 2432, 2595, 4396, 4201, 4397, 2760, 3063, 685, 2581, 3109, 3256, 3314, 2520, 2705, 2722, 2535, 3064, 3603, 823, 2322, 2390, 3337, 2719, 3410, 2582, 713, 1894, 493, 2442, 2987, 2631, 2758, 1849, 2400, 2368, 4401, 3032, 2669, 2599, 2646, 3110, 3260, 3338, 2319, 2285, 4409, 2469, 2041, 2915, 2967, 1395, 2358, 3807, 1913, 78, 270, 379, 417, 4317, 4324]. physics [4211, 3885, 479, 2304, 4343, 1750, 416, 3509, 675, 701, 2281, 2221, 2181, 4330, 3974, 4232, 3971, 1796, 1402, 1169, 3934, 2339, 1862, 1236, 2411, 2602, 179, 487, 4277, 1875, 249, 1011, 296, 1453, 2196, 2279, 3462, 3677, 3875, 4045, 3969, 4127, 1980, 1880, 1270, 1405, 2212, 3585, 4333, 1074, 1276, 778, 409, 992, 1589, 590, 452, 4209, 2749, 3889, 223, 2329, 3591, 16, 3502, 2406, 784, 1019, 1229, 4303, 717, 225, 3856, 1258, 51, 177, 4376, 1218, 707, 3616, 3960, 4130, 4162, 4259, 4281, 4270, 2077, 317, 2109, 1770, 2366, 1381, 1573, 237, 245, 1767, 3686, 4196]. physics [384, 764, 1279, 1717, 3627, 3752, 3824, 4047, 4283, 2129, 193, 2143, 2201, 2548, 2886, 1583, 1255, 1902, 4322, 4318, 3331, 316, 3160, 4383, 3540, 1064, 1885, 441, 485, 1108, 4107, 399, 1171, 192, 2167, 797, 2103, 526, 568, 619, 4214, 1660, 4112, 4159, 4160, 3008, 93, 3972, 1664, 843, 3952, 2647, 2543, 3356, 3849, 290, 3712, 3220, 863, 988, 1160, 154, 964, 4282, 854, 1943, 1969, 1959, 716, 2183, 3465, 2700, 4395, 393, 4244, 3572, 3584, 3836, 3959, 3917, 3949, 1812, 1214, 3636, 3967, 1797, 601, 746, 3613, 3731, 137, 3690, 3476, 3681, 3864, 3831, 3929, 4255]. physics [4208, 4219, 4311, 264, 364, 4166, 3760, 529, 1099, 1480, 2157, 2236, 4095, 4265, 94, 1678, 3596, 3528, 3871, 4086, 4147, 4171, 1363, 120, 1918, 4034, 955, 228, 66, 1973, 2683, 2554, 2831, 3336, 3357, 1014, 1544, 3242, 3568, 3556, 3484, 3847, 3775, 3850, 3997, 4116, 4172, 4145, 4258, 1704, 1667, 3389, 3593, 1817, 76, 4229, 1146, 1734, 1284, 2173, 39, 1936, 3680, 2238, 27, 2751, 4184, 4417, 2765, 4124, 2501, 2809, 2374, 2491, 3124, 3072, 2356, 2679, 3091, 3017, 3133, 3237, 3171, 4184, 2509, 2842, 2838, 3092, 2639, 2999, 3117, 2294, 2537, 3181, 3279, 3054, 2850, 3315]. Physics [3317, 2918, 3381, 3431, 2396, 2562, 3121, 2523, 2616, 2881, 3271, 2873, 2914, 2855, 3012, 3119, 3132, 2510, 2511, 2314, 2701, 2910, 2913, 2401, 2370, 2619, 2669, 2599, 2475, 2885, 2827, 3021, 3593, 4149, 3986, 4317, 4211, 4225, 3719, 3491, 3485, 4009, 3748, 4028, 4191, 4061, 3507, 112, 4138, 3784, 4001, 4024, 3576, 3754, 3312, 4117, 3973, 3718, 4349, 4337, 4201, 3704, 4096, 3595, 3667, 3818, 3916, 4176, 3688, 4185, 4286, 675, 4307, 3706, 3308, 2747, 2639, 4084, 4198, 2394, 4295, 4305, 4068, 2996, 3309, 3957, 3233, 3492, 3399, 3487, 2890, 2971, 2472, 4354, 2876, 4144, 4352]. Physics [2588, 3529, 2723, 2417, 3713, 2559, 2710, 3892, 2920, 1146, 3799, 8, 3946, 3901, 3873, 3913, 3883, 3874, 3882, 3950, 160, 3880, 3864, 3837, 2894, 3829]. physics/ideas [4211]. physique [366]. Pian. [3039]. Pickering [4202]. Picoscience [3307]. picturalism [2328]. picture [2519]. Pieper [4218]. Pier [2954, 4145]. Pier-Giorgio [2954]. Pierce [2294]. Pierfrancesco [4210]. Pierluigi [3269]. Pierre [3446, 3487, 3500, 3172, 4064, 2610, 2977].

Pierre-Gilles [2977]. Pierre-Louis [3172]. Pierre-Yves [2610]. Pierrus [2781]. Pieter [2636]. Pigolotti [4310]. Ping [2905]. pinning [748]. Pinter [4149]. pion [1481]. Pioneer [3111, 2186]. pions [4164]. Pisanti [2674]. Pivato [2543]. pixellated [2039]. place [4037, 48, 393]. Placzek [4030]. Plagnol [3792]. plain [1510, 1535, 2430, 3645, 1588]. Planck [2893, 4417, 4041, 3687, 2195]. plane [4231]. planet [3993, 4150, 4399, 3876]. Planetary [1021, 3100, 2839, 3238, 2618, 666, 1074, 717, 2603, 3891, 337, 2166, 4301, 3737, 4039, 3394, 4150, 2476, 3234, 2826, 3129, 3897]. Planetesimals [3689]. Planetodiversity [2166]. planetology [1157]. Planets [3456, 3415, 3100, 3234, 2839, 3334, 2166, 4398, 2575, 1083, 2397, 2898, 2476, 3689]. Planning [2626, 74]. Plant [2910]. Plasma [2756, 3298, 3072, 20, 384, 466, 589, 2818, 2967, 3644, 1522, 1589, 707, 193, 234, 4125, 3693, 124, 2313, 2547, 2293, 3276, 3556, 3850]. Plasmas [1198, 2652, 1406, 1047, 3864, 2818, 3272, 2920, 3850, 3978]. Plasmon [3408]. Plasmon-Enhanced [3408]. plasmonic [3696, 3530, 2879]. plasmonics [2273, 3275, 3628]. plasmons [1517]. Plastic [2627, 3147, 3544]. Plasticity [3154]. plate [525]. Play [2933, 3477, 1843]. players [2474]. playground [765]. playing [1684, 2397]. plenty [1263, 2264, 2733, 2263]. Plenum [555]. plethora [2814]. Plévert [2977]. Pluchino [3997]. Plumes [3298]. plutonium [850]. pocket [807]. Podolský [2718]. poetry [2578]. Poets [2691]. Poincaré [3490, 2861]. Poinern [3488]. point [584, 420, 3516, 972, 521]. points [8, 848, 262]. Pokrovskii [4249, 2806]. Polarisation [2584]. polarised [3251]. polariton [2516]. Polarization [2309, 2429]. Polarized [524, 2094, 1862, 260]. polaron [1494]. Polarons [3169]. pole [944]. polemic [3575]. Policy [3031, 740, 391]. Politically [1989]. politicians [3206]. Politics [3128, 2367, 3226]. pollution [639, 4302, 868, 936]. Polyakov [4029]. polyatomic [989]. Polycrystalline [1311]. Polymer [2806, 1350, 567, 4235, 1449, 2767, 2927]. Polymeric [754, 450]. polymers [1430, 1268, 1370, 564, 1269, 979, 1407, 544, 3116, 2344, 2443, 2319]. polywater [594]. Pomeranchuk [723]. Poon [2787]. Popov [2942]. Popp [3246]. Popper [1153, 1181]. Popular [1770, 2439, 2392, 2397, 2371, 2380, 2581, 3197, 2534]. popularizer [1968]. population [1988]. Populations [2626]. Porous [3402]. Portable [2430]. Porter [2726]. Portolan [3659]. position [752, 1658]. positions [631]. Positron [1495, 2030, 1340, 2336, 1064]. positronics [1981]. positrons [659]. possibility [4237, 3086]. possible [919, 2294, 1596]. post [2659, 3057, 2993, 3016, 3394, 2859, 89, 277]. post-docs [3057, 2993, 3016]. post-equinox [3394]. post-graduate [2659, 2859]. postdoctoral [2330, 2395]. Postgraduate [3035, 2887, 3178, 4120, 4078, 4060, 4106, 2686, 2883, 3157, 3395, 3318, 3368, 3211, 3333, 41, 3052, 3350, 2433, 2502, 2788, 2414, 2346, 2786, 3061, 3073, 3480, 3498, 3098, 3201, 2667, 3509, 2374, 4241, 4217, 4284, 4275, 2614, 2555, 3022, 3180, 3506, 4202, 2897, 2819, 3301, 3097, 2840, 3107, 2593, 2949, 3011, 3365, 2830, 2403, 2334, 3029, 3124, 2966, 3485, 2544, 3376, 4231, 4293, 2795, 3116, 3384, 2615, 2564, 2662, 3091, 3017, 3037, 3133, 3237, 3412, 3383, 4182, 4184, 2320, 2347, 2757, 3349, 3137, 4113, 4356, 4328, 3038, 3095, 3247, 4118, 2995, 4061, 3442, 3123, 3296, 2823]. postgraduate [3023, 3396, 2838, 2836, 2853, 2785, 3104, 3275, 3194, 3179, 3353, 3408, 3499, 3493, 3488, 4183, 4240, 4272, 3076, 3144, 2696, 3207, 2621, 3168, 2735, 2790, 2872, 2755, 3448, 4248, 2766, 2792, 2878, 2847, 3105, 3245, 3257, 2791, 2841, 2541, 3040, 2999, 2561, 2687, 2845, 3036, 4072, 4052, 4178, 4190, 4198, 4252, 2529, 2777, 2794, 3118, 2965, 3279, 3280, 3343, 3344, 3494, 3501, 3450, 3452, 3451, 4204, 4203, 4257, 2702, 2693, 4323, 2660, 2800, 3054, 3204, 3407, 2601, 3126, 2979, 3205, 4085, 4108, 4068, 3184, 4242, 3321, 4250, 2726, 3081, 3377, 3435, 3437, 4339, 2313, 3099, 4294, 2330, 2348, 2402]. postgraduate [2678, 2538, 2820, 2783, 2746, 3046, 3224, 3196, 3369, 3270, 2612, 3431, 3143, 3233, 3258, 3366, 3429, 2670, 3374, 3024, 2771, 3050, 2898, 4165, 3169, 3303, 3398, 3492, 2524, 2716, 3128, 3497, 3382, 3108, 3269, 3080, 3005, 2808, 4157, 3230, 3210, 3345, 3446, 4186, 2623, 2539, 2476, 2616, 2525, 2902, 3187, 3182, 3340, 3423, 2682, 2526, 3026, 3003, 3424, 3189, 3511, 2680, 3015, 2692, 2950, 2797, 3162, 3447, 3208, 3206, 2466, 2594, 3140, 3058, 3367, 2395, 2455, 2685, 2803, 2946, 3084, 3145, 2998, 3232, 3307, 3449, 4200, 4205, 3049, 3231, 3234, 3065, 2622, 3430, 3310, 2839]. postgraduate [2873, 2914, 2855, 3090, 3166, 3316, 2652, 3272, 3103, 3311, 555, 2651, 2332, 2784, 2712, 3127, 3033, 3012, 4122, 4197, 4349, 3213, 3214, 2415, 2595, 3094, 3332, 2780, 2624, 2835, 2874, 2837, 2834, 2760, 2750, 2880, 3062, 3119, 2997, 3063, 3172, 3409, 4054, 4083, 4189, 3156, 2393, 2974, 3413, 3490, 2436, 2657, 2676, 2909, 2728, 2782, 2846, 3028, 3256, 3347, 2441, 2440, 3402, 2389, 2410, 2546, 2535, 2695, 3414, 3476, 3489, 4071, 4255, 4219, 4301, 4310, 4276, 4350, 4077, 4166, 4167, 2759, 3138, 3346, 2725, 4199, 4334, 2322, 2310, 2314, 2390, 2701, 2896, 3461]. postgraduate [4253, 4266, 4332, 2592, 2719, 2720, 2804, 2888, 2805, 3148, 3146, 3141, 3265, 3391, 3482, 4063, 4220, 2531, 4174, 2718, 2978, 4216, 2442, 3286, 2987, 2349, 2540, 2530, 2951, 2758, 2947, 2409, 2619, 3191, 3185, 3278, 3508, 4290, 3136, 3147, 2807, 2826, 2654, 2536, 2665, 2666, 2470, 2832, 3411, 3338, 3306, 3475, 3495, 4058, 4116, 4215, 4175, 3039, 3067, 4111, 4185, 4168, 2344, 2752, 2882, 2852, 2769, 4267, 2704, 4343, 4055, 4177, 4251, 4325, 4081, 4162, 4305, 4318, 4164, 4053, 4080, 4070]. postgraduate/research [2414]. postgraduate/specialist [2433, 3039]. postgraduates [2921, 2936, 2911, 2865, 2293, 2866, 2939, 2931, 2941, 2907, 2494, 3004, 3496, 2954, 2737, 3113, 2387, 2340, 2333, 2300, 2523, 2889, 3186, 3457, 3487, 3460, 3432, 2935, 2815, 2924, 2943, 2833, 2468, 2501, 3474, 2915, 2920, 2967, 2919]. postgraduates/specialists/advanced [2387]. postgraduation [2432]. postscript [1153, 1181]. potential [1486, 4090, 853, 2031]. Potter [4183, 2430]. Potvliege [2949]. Pötzsche [3340]. Pouliquen [3213]. powder [744]. powders [1571]. Power [1082, 4369, 2142, 107, 925, 201, 3459, 972, 638, 738, 1336, 2175]. Powers [3796, 3379]. Pozrikidis [2997]. Pp [554, 662, 291, 2082, 2057, 112, 555, 869, 4132, 4120, 4078, 4060, 4106, 4010, 4317, 4324, 4016, 4062, 4110, 4225, 4082, 4343, 4241, 4217, 4284, 4275, 4202, 4232, 4268, 4135, 4140, 4139, 4254, 4038, 4000, 4032, 4009, 4056, 4055, 4273, 4271, 4304, 4296, 4206, 4035, 4231, 4293, 3425, 4005, 4045, 4018, 4028, 4073, 4123, 4069, 4154, 4182, 4177, 4184, 4191, 4221, 4027, 4100, 4113, 4356, 4328, 4307, 4118, 4061, 4333, 4143, 4183, 4272, 3102, 4030, 4084, 4155, 4248, 4251, 4325, 4327, 4355, 4291, 4303, 4138, 4344, 4348, 4026, 4072, 4178, 4190, 4148, 4198, 2191, 4163, 4252]. pp [4006, 4021, 4204, 4130, 4203, 4162, 4235, 4222, 4257, 4259, 4269, 4309, 4281, 4270, 4295, 4274, 4323, 4025, 4002, 4001, 4046, 4043, 4041, 3999, 4029, 4044, 4008, 4024, 4085, 4108, 4119, 4098, 4196, 4141, 4242, 4250, 3459, 4031, 4037, 4047, 3993, 4040, 3994, 4048, 3957, 4042, 4003, 4015, 4091, 4345, 4338, 4342, 4346, 4339, 4340, 4187, 4283, 4294, 4322, 4353, 4318, 4319, 3998, 4165, 3421, 4302, 4097, 4107, 4157, 4292, 4150, 4186, 3423, 4007, 4017, 4112, 4115, 4117, 4159, 4161, 4158, 4160, 4321, 4099, 4093, 4200, 4151, 4205, 4224, 4164, 4019, 4282, 4326, 4121, 4131, 4197, 4337, 4354, 4244]. pp [4218, 4249, 4316, 4020, 4022, 4036, 4054, 4083, 4053, 4080, 4079, 4234, 4144, 4351, 4341, 4352, 4306, 3996, 3995, 4071, 4129, 4255, 4208, 4188, 4219, 4301, 4311, 4310, 4276, 4350, 4320, 4070, 4066, 4077, 4167, 4199, 4260, 4308, 4334, 3337, 4095, 4096, 4094, 4253, 4266, 4332, 4065, 4064, 4063, 4220, 4086, 4092, 4142, 4147, 4171, 4174, 4216, 4290, 4034, 4059, 4109, 4004, 4039, 3956, 4033, 3997, 4023, 4058, 4116, 4067, 4172, 4136, 4207, 4176, 4215, 4233, 4335, 4145, 4258, 4111, 4185, 4168, 834, 4243, 4256, 4267]. PPSE [4321]. Practical [211, 3076, 3105, 2680, 3216, 3006, 2487, 3127, 2807, 3397, 3509, 4273, 3577, 3517, 4122, 3568, 4132, 3755, 3697]. practicals [4283]. Practice [2585, 209, 259, 2474, 3064, 3685]. practicing [2464]. practising [2309, 2465, 2671, 2584, 3140]. Practitioner [1968, 3959]. practitioners [2743, 2739, 2803, 2787]. Pradhan [2788]. Pradisi [4070]. pragmatists [1445]. Pranay [4267]. Prandtl [1422, 1427]. Pratap [4240]. Pratesi [3648]. Pratim [3608]. Pravilov [4110]. Praxis [4059]. pre [2137, 2809, 4343, 4232, 3349, 3137, 3035, 3152, 4084, 4156, 2944, 3315, 3405, 4322, 3135, 3121, 3128, 3018, 3415, 4186, 3100, 3423, 2682, 2881, 2899, 3234, 3243, 2908, 3066, 3461, 2938, 2913, 3285, 4086, 2826, 3134, 3215, 3164, 4317, 4163]. pre-breakdown [2137]. pre-undergraduate [3018]. pre-University [3035, 2809, 4343, 4232, 3349, 3137, 3152, 4084, 4156, 2944, 3315, 3405, 4322, 3135, 3121, 3128, 3415, 4186, 3100, 3423, 2682, 2881, 2899, 3234, 3243, 2908, 3066, 3461, 2938, 2913, 3285, 4086, 2826, 3134, 3215, 2358, 3164, 4317, 4163]. Prebiological [700]. precedence [1827]. Precipitation [873, 604]. Precise [3150, 4363, 3245, 2339, 778, 3756, 2055, 3877]. Precision [4193, 4245, 2212, 2174, 1658, 3797]. predict [2336]. Predicting [1903]. prediction [1091, 1310, 2956]. Predictions [1703, 4373, 3450]. Preecha [3534]. preferably [3120]. Prehistoric [1175]. Premaratne [2820]. Premelting [1315]. preparatory [52]. present [3333, 667, 1903, 639, 891, 981, 4146, 1302, 3399]. present-day [4146]. presentation [4078]. Preservation [2803, 2211]. Press [4078, 4060, 4317, 4016, 4211, 4137, 4286, 4082, 584, 4217, 4284, 4275, 4135, 4038, 4000, 4032, 4009, 4210, 4271, 4035, 3425, 4005, 3945, 4045, 4073, 4182, 4177, 4184, 4027, 4100, 4356, 4328, 4307, 4061, 2057, 4143, 4084, 4248, 112, 4327, 4355, 4291, 4285, 4344, 4026, 4072, 4052, 4178, 4190, 4198, 2191, 4006, 4021, 4081, 4156, 4130, 4203, 4162, 4259, 4269, 4309, 4270, 4274, 4025, 4002, 4001, 4046, 4043, 4041, 4085, 4141, 3459, 4037, 4040, 3994, 4048, 4042, 4003, 4015, 4105, 4146, 4345, 4338, 4342, 4346, 4339, 4340, 4187, 4283, 4322, 4318, 4319, 4165, 3421, 4302, 4097, 4107]. Press [4157, 4292, 4007, 4017, 4158, 4093, 4200, 4151, 4205, 4019, 4282, 4326, 555, 4131, 4122, 4197, 4316, 4020, 4036, 4079, 4071, 4129, 4219, 4301, 4311, 4310, 4276, 4350, 4320, 4066, 4077, 4166, 4199, 4094, 4266, 4065, 4064, 4063, 4092, 4034, 4109, 4039, 4023, 4058, 4176, 4215, 4335, 4175, 4258, 4185, 4168, 4267]. Pressing [2110]. Pressure [3321, 4223, 375, 625, 854, 1363, 404, 1146, 1146, 3012]. pressures [523, 142]. prevent [3517]. prevention [823]. Prialnik [2737, 2621]. Pride [1827]. Priest [3878]. Priezzhev [2804]. Primary [35, 22, 4321]. primer [1351, 3716, 3620, 3755, 3825, 2687, 4160, 4207]. primeval [1847]. Primordial [2539, 4146, 2741]. Princeton [4275, 3945, 4184, 4027, 4356, 4327, 4355, 4291, 4285, 3987, 4344, 4190, 4274, 4085, 4338, 4339, 4340, 4292, 4007, 4200, 4151, 4205, 4282, 4131, 4316, 4301, 4310, 4023, 4176, 4205, 2666]. Principia [3270, 4128, 1150, 1190]. Principle [3247, 3354, 3328, 3762, 4222, 1614, 1705, 2861, 1943, 1658, 510]. Principles [3395, 2386, 3373, 4381, 3274, 3446, 2803, 3231, 3053, 129, 3059, 2799, 3068, 3520, 4324, 3470, 1342, 3696, 3619, 3462, 4162, 4274, 201, 3630, 3926, 3527, 4395, 4249, 3634, 4189, 738, 3526, 3814, 1642, 3568, 3174, 3799, 2566, 3080, 3953, 3585, 3036, 2930, 3983, 4203, 2954, 3975, 2678, 3421, 3058, 4205, 4121, 3536, 2638, 3701, 4092, 2736, 2967]. Print [4142, 4325, 4204, 4353]. Printed [4251]. Printing [4079]. Prize [605, 4200, 4151, 4205, 3837, 2980]. prizes [945]. Probabilistic [2634]. Probability [4360, 3444, 3138, 3014, 2888, 3896, 3798, 3112, 3041, 4148, 2808, 2940]. Probe} [2126]. probes [337, 1035]. Probing [1776, 847, 857, 3320]. Problem [2700, 777, 2095, 2406, 684, 509, 477, 2216, 3895, 3960, 3727, 952, 4201, 4199, 12, 2117, 1118, 2664, 3881]. Problem-based [2700]. problem-centred [2117]. problem-solving [4201]. Problems [270, 3052, 2781, 2786, 2706, 3233, 3114, 2685, 2705, 1913, 3830, 2347, 639, 4084, 1279, 4187, 4322, 4114, 3952, 4326, 1627, 3593, 3633, 2518, 3031, 3747, 4053, 4258, 2987]. ProblemSolving [4201]. proceeding [3039, 3348]. Proceedings [3133, 4417, 3308, 2737, 415, 2467, 2322, 3039, 3267, 3593, 834, 2414, 2727, 3261, 3412, 2984, 3075, 3054, 2932, 2898, 3189, 4159, 2884, 2651, 3012, 2780, 2760, 2750, 3063, 2441, 137, 2822, 3067, 266, 3412]. process [1743, 2165]. Processes [3211, 3123, 2352, 3257, 2979, 3435, 3422, 4379, 2808, 2940, 3132, 3138, 3014, 3255, 2442, 4406, 2787, 1379, 892, 989, 1385, 1400, 3860, 711, 448, 122, 4150, 3486, 3735, 304, 347, 2931, 3065, 2556, 2654, 2536, 4189, 3449]. Processing [2975, 2766, 1177, 4298, 2107, 3548, 3351, 2055, 4216, 2147, 2636]. processive [2123]. produced [1047]. producing [1792]. production [4050, 3008, 212]. professional [2350, 2347, 2351, 2461, 2354, 389, 3228, 2888]. professionals [2889, 2877, 4114]. Professor [3574, 1057, 4161, 766, 899, 1757, 3512, 1205]. profile [3856]. program [3776]. programmable [862]. programme [1699, 1029, 1325]. programmes [3206]. programming [3660]. Progress [498, 1825, 379, 417, 2296, 225, 317, 316, 636, 583, 716, 66, 2378]. progresses [4108]. project [3938, 843, 833, 3529, 869, 4171]. projected [1098]. projects [479, 3800, 2899]. Prokof’ev [3460]. prologue [801]. proof [4336, 2517]. Propagation [2820, 2300, 3058, 2482, 3765]. Properties [2667, 420, 3116, 3165, 3450, 3894, 432, 559, 2712, 3033, 3872, 505, 2812, 1290, 3647, 4373, 260, 236, 134, 475, 438, 1211, 901, 962, 1248, 4220, 321, 1055, 3242, 4215, 124, 1464, 1722, 2792, 2774, 3535]. proponent [1968]. proposal [869]. Propulsion [2453, 671]. prospect [2009]. Prospects [1947, 3904, 1081]. Prost [2527, 2596]. protein [600, 630, 2189]. proteins [2123]. protocols [2235]. Proton [839, 1706, 115]. Pruessner [3070]. Pruneau [3808]. Prutchi [2899, 2899]. psychological [362]. psychology [941]. PT [3971]. Pte [4106, 4268]. Public [3363, 4186, 2438, 2380]. publications [30]. Publishing [4106, 4241, 4268, 4139, 4254, 4206, 4191, 4221, 4240, 4251, 4325, 4303, 4204, 4222, 4281, 4150, 4201, 4223, 4218, 4169, 4170, 4234, 4188, 4260, 4220, 4147, 4290, 4136, 4207, 4233]. Pucci [3408]. Pudipeddi [3351]. Pudritz [3129]. Pugai [3079]. Pugh [3372, 2302]. Puglisi [4208]. pull [3802]. Pulsar [3016]. Pulsars [769, 3267, 1607, 1559, 1111]. pulse [2731, 4289]. Pulsed [3373, 867, 1414, 4315]. pulses [2220, 1863, 4289, 2847]. Pumped [2783, 1528, 4236]. pumping [1240, 1079]. Purcell [3115, 2802]. Puri [4159]. purpose [4242]. purposeless [1597]. Pursuit [2334, 2894]. Pursuits [3399]. put [1378]. Puzzle [3128, 2764]. Puzzles [1851, 3780]. Pyroelectric [673]. Python [3577, 3936, 4267, 4292, 3660, 3351].

QCD [2622, 3329, 3516, 1964, 2831]. QED [4212, 1243]. Qi [4167, 3457]. Qing [2982]. Qing-Zhu [2982]. quality [1516]. Quanta [3071, 2827, 3077]. Quantics [1524]. quantisation [3851]. Quantitative [1759, 285, 2844, 4224, 3565]. Quantization [3217, 1257, 2419]. Quantized [2686, 2524]. quants [2047]. Quantum [2686, 3368, 2829, 2346, 4331, 1750, 2296, 2297, 2374, 2936, 3022, 3180, 2897, 2819, 516, 1317, 1469, 3029, 4271, 2425, 2184, 4287, 2585, 2471, 2448, 2495, 2856, 3089, 3056, 3133, 3320, 2945, 2801, 2948, 2995, 1250, 1935, 4362, 2636, 4298, 3223, 2328, 3296, 819, 2984, 3493, 2217, 3076, 2421, 2847, 3165, 1888, 2090, 2655, 2999, 2561, 2687, 2845, 1966, 1401, 325, 2286, 4377, 1899, 2601, 2979, 1692, 3190, 2235, 4380, 2348, 2678, 2746, 3046, 3163, 3287, 3405, 2054, 2268, 3374, 3906, 3024, 3034, 3303, 4279, 2431, 2716, 3399, 1066, 3114, 3108, 3401, 1171, 3230, 3210, 2371, 2483, 2525]. Quantum [3189, 2234, 2692, 2797, 4087, 2933, 2864, 2899, 3966, 3972, 2946, 1030, 3125, 2694, 2681, 2744, 2875, 2935, 3045, 3090, 1693, 3311, 1813, 3086, 4394, 4396, 1606, 3949, 3991, 2677, 2521, 2846, 3239, 4236, 2759, 2055, 1140, 2069, 2805, 3265, 3048, 2723, 1394, 2540, 2951, 2947, 1630, 2398, 2370, 3217, 2645, 2444, 2492, 2822, 2277, 3549, 2665, 2656, 3260, 1464, 1844, 4229, 4364, 1263, 2264, 3763, 2170, 3250, 3506, 3974, 4268, 2138, 3785, 3971, 4139, 595, 3365, 4000, 3934, 3944, 4296, 3716, 4014, 3791, 3759, 3881, 4005, 3914, 3969, 4177, 4134, 2017, 1194, 2382, 3585, 1934]. quantum [3518, 3538, 4209, 2174, 4248, 1727, 3515, 2446, 1707, 2241, 3739, 2127, 1839, 3561, 4081, 875, 2577, 3513, 3915, 3925, 3975, 3920, 3999, 4068, 4196, 4141, 3477, 2133, 2176, 3517, 2107, 2338, 3331, 1979, 3548, 3469, 3510, 1387, 4051, 3227, 3825, 1807, 2263, 3417, 302, 3770, 3843, 2455, 4089, 3571, 2148, 4280, 322, 330, 2242, 616, 3546, 1840, 2040, 4075, 3545, 2698, 1303, 1793, 2266, 3736, 3833, 3917, 4022, 3811, 3863, 3704, 4306, 4066, 4166, 1743, 3330, 4199, 3225, 4095, 4265, 4332, 2179, 3482, 3838, 2194, 4216, 2361, 2147, 1262, 1786, 2265, 3336, 4013, 2326, 2454, 3782]. quantum [3857, 4145, 2014, 1988, 2160, 1691, 3600, 1835, 1737, 4267, 3794, 3157, 3393, 4137, 3079, 2290, 3927, 3522, 3772, 3912, 2375, 3539, 3710, 3899, 3911, 2311, 2350, 2490, 2498, 2904, 3089, 3848, 3855, 4328, 2945, 3657, 3658, 2939, 2992, 3257, 4348, 2687, 4021, 4041, 3920, 3954, 3752, 2609, 2402, 4319, 3374, 2524, 3497, 3210, 2691, 3345, 2797, 3162, 3447, 3720, 3027, 2455, 3589, 4093, 3819, 3065, 2796, 3053, 2985, 3019, 3282, 3316, 2267, 4218, 3734, 3879, 3880, 2854, 2728, 4260, 4297, 4308, 2317, 3146, 3567, 3439, 3871, 2579, 4216, 2422, 2530, 2298, 2399, 2369, 2558, 2454, 2669]. Quantum [2536, 2666, 2656, 3306, 3680, 3694]. quantum-field [3225]. quark [1679, 1189, 2756, 2313]. quark-gluon [1189]. Quarks [2159, 1318, 690, 4164, 3240]. quartz [2177]. Quasi [436, 1993, 3895, 1602]. quasi-analytical [3895]. quasi-atomic [1602]. Quasi-stellar [436]. Quasi-two-dimensional [1993]. Quasicrystals [1339]. qubit [2641]. qudit [4313]. Quest [2770, 3431, 2532, 3763, 4027, 4355, 4376, 4064]. Question [2610, 3708]. Questioning [2768]. Questions [4269, 1545, 4221, 3929, 3492, 4080]. quicksands [2448]. Quiet [250]. Quilt [2491]. Quinten [3105]. Quips [2827]. Quiroga [4032]. Quotes [2827].

R [3776, 4120, 4060, 1121, 4324, 3509, 2949, 4038, 2858, 4009, 3656, 3644, 2545, 2490, 3091, 3748, 3759, 4027, 2984, 3628, 2775, 3152, 3647, 2342, 3784, 2561, 2241, 3117, 2794, 3809, 1069, 2601, 4119, 2737, 3081, 3057, 3708, 3627, 3643, 3816, 3754, 3773, 3787, 3732, 4048, 4015, 4346, 4294, 2330, 2318, 2387, 2678, 3262, 4319, 3143, 2670, 2473, 2292, 2371, 2435, 2539, 3457, 1100, 3078, 4159, 2263, 2245, 2899, 3653, 3589, 4099, 3430, 2903, 3000, 3090, 3749, 56, 2449, 3127, 2835, 2750, 3736, 4022, 4053, 2676, 3183, 3132, 3613, 2416, 3777, 2725, 1140, 2557, 4297, 2322, 2310, 2390, 2376]. R [3286, 1497, 2613, 2663, 3684, 3818, 4039, 4116, 3593, 1146]. R.G [3754]. R.K [4159]. Rabindra [4155]. Raby [4177]. Race [2980, 3239, 3322]. Rachael [3680]. Rached [3370]. Radaelli [3711]. Radar [87, 1314, 2307]. Radek [4169]. Radial [2427]. Radiation [2763, 2954, 3407, 2597, 3013, 2811, 404, 2721, 884, 646, 1356, 832, 3176, 2347, 453, 2731, 1850, 1444, 2198, 645, 2250, 602, 134, 772, 2035, 2249, 1677, 926, 274, 282, 3467, 1140, 1995, 3713, 4215, 22, 4038, 3662]. radiations [209]. Radiative [2905, 1924, 1987]. Radio [1265, 123, 4374, 2515, 334, 2920, 70, 255, 1607, 752, 783, 353, 4040, 436, 352, 3655, 3455]. Radio-Frequency [2920]. Radioactive [451, 834, 1119, 925, 1972]. Radioactivity [1773, 1671, 566, 2035, 3120]. radioisotopes [1048, 7]. radiology [1775]. radiolysis [1027]. Radionuclides [850]. radiotherapy [3173]. Radko [3413]. Radon [1596]. Rae [2904]. Raeisi [4341]. Rafael [4170]. Rafii [2332]. Rafii-Tabar [2332]. Rafique [4254, 3073]. Raghuveer [4274]. Rai [2594]. Raimondi [3557]. Rain [3411]. Raine [1121]. Rajasekar [4324, 3657, 3658, 4319, 4297, 3566]. Rajeev [3235]. Rajendra [2943]. Rakesh [2556]. Ralf [2948, 2528]. Ralph [3142, 3092, 3129, 3667]. Ram [4302]. Ramakrishna [2394]. Raman [657, 1461]. Ramanujam [2429]. Ramiah [3657, 3658]. ramifications [1037]. Ramkarthik [4267]. Rammer [2290, 2369, 3799]. Rampino [2473]. Ramsey [3457]. Rancoita [2954, 2881]. Rand [3770]. Randall [2742]. Random [2808, 2940, 2825, 3414, 3138, 3252, 4061, 2054, 2268, 3789, 2267, 3735, 4208, 2259, 3640, 1199, 2260]. Randomness [2410, 2047, 3686]. Range [2288, 782, 2506, 893]. Range-Resolved [2288]. ranging [2032, 2737, 3113, 2467]. Ranjan [2712]. Rankin [2599]. Ransom [3655]. Raphael [2912, 2785]. rapidly [457]. Rappaz [2794]. Rare [2909, 472, 659, 3889, 570]. Rare-Earth [2909, 570]. Rashid [3298]. Rashmi [2461]. Rasinariu [2947]. Rasooly [2991]. Rastelli [3214]. ratchets [1796, 2088, 3584]. Ratcliffe [3232]. Rates [2521]. ratio [115]. Rationalist [3399]. rationalized [387]. Raton [4317, 4217, 4284, 4296, 4259, 4269, 4270, 4294, 4318, 4319, 4302, 4336, 4094, 4258, 4267]. Rauch [3510]. Rauscher [4174]. Raveendran [2750]. Ravin [2850]. Ravindra [4240]. Ravishankar [3351]. Ray [3102, 3262, 2688, 2929, 3039, 617, 314, 1204, 328, 802, 1551, 553, 2070, 2572, 3550, 629, 2813, 2143, 3131, 48, 583, 938, 4107, 2642, 2680, 3601, 91, 1419, 2295, 716, 2633, 4079, 911, 3387, 698, 699, 3675, 3775, 1736, 3644, 3057]. Rayleigh [1234]. Raymer [2265, 3752]. Raymond [3522, 4342, 2461, 3281]. Rays [2776, 2702, 2597, 156, 2881, 2013, 4293, 1656, 2550, 573, 415, 4080, 1075, 345, 3667, 2000, 2447, 3015]. Razavy [2677]. RB [4240]. Reach [3472]. reaction [3176, 2811]. Reactions [549, 1069, 2566, 1406, 1206, 319, 1112, 1130, 1400, 2024, 900, 499, 290, 2500, 4159]. Reactive [761]. Reactor [159, 206, 275, 498, 1501, 1104, 3829]. Reactors [1175, 2674, 1068, 961, 867, 887]. Reader [3126, 2312, 3111, 2462, 2661, 2334, 2411, 2356, 2495, 2658, 2742, 2707, 3229, 3403, 3406, 3462, 2509, 3442, 2764, 3226, 3308, 2906, 4274, 2542, 2289, 2318, 2452, 2548, 2506, 2668, 2675, 2849, 3131, 3088, 3120, 3335, 3453, 2397, 3354, 2473, 2430, 3445, 2371, 2435, 2367, 2379, 2644, 2493, 2863, 2505, 3078, 2466, 2877, 2434, 2449, 2520, 2695, 2331, 2417, 2321, 2426, 2413, 2827, 2534, 2456, 2420, 2450, 3150, 3756]. readers [3004, 2519, 2980, 3018, 2864, 4117, 3031, 3006, 2833, 2970, 3020]. readership [3393, 4317, 4211, 4286, 2772, 3295, 4343, 4241, 4275, 2527, 2851, 2843, 3107, 2585, 2471, 2679, 2600, 2565, 2564, 2498, 2857, 2870, 2904, 3195, 3425, 3349, 3137, 4356, 2977, 3035, 2672, 3322, 2673, 2853, 3104, 3488, 4240, 3102, 3444, 3152, 3139, 3192, 3302, 3502, 4327, 4355, 4303, 4285, 2392, 3009, 2684, 3177, 3268, 3400, 3394, 3478, 3427, 3436, 3456, 3455, 3459, 4338, 4342, 4340, 2886, 3254, 3224, 3262, 3253, 3244, 3405, 3381, 3270, 3312, 2612, 4322, 3135, 2562, 3121, 3363, 3128, 3274, 3202, 2714, 2691, 3345, 3415, 4186, 3263, 3283, 3423, 2916, 2929, 2933, 3428, 4321, 2753, 3084, 3351]. readership [3234, 3299, 3271, 2694, 2796, 2908, 3276, 2778, 3106, 4223, 4218, 3066, 3002, 3239, 4208, 4311, 4320, 2717, 2496, 2814, 2895, 3197, 3225, 3461, 3441, 2532, 2582, 2589, 2588, 2938, 2913, 3048, 3151, 3149, 3281, 4086, 4171, 3286, 2982, 2645, 2558, 3136, 2789, 2599, 2934, 2871, 2957, 3134, 3001, 3085, 3129, 3042, 3284, 3379, 3443, 4258, 3300, 3378, 3355, 2650, 2610, 2990, 4057, 4324, 4082, 4291, 4344, 4345, 4115, 4316, 4065, 4064, 4109, 4233, 4256]. Reading [9, 110, 2228, 306, 335, 95, 128, 55, 80, 269]. Real [3841, 4232, 580, 4233]. Realistic [3300, 3232]. Reality [2495, 2498, 2904, 2770, 3261, 3406, 3425, 4416, 1984, 3086, 2017, 3637, 3419, 1636, 1180, 1743, 2554, 3163, 2558, 2707, 3855]. realizations [2131, 2147]. Really [3034, 4005, 2305, 3982]. realm [2123]. reappraised [755]. Reasoning [2518]. Rebases [834]. Recami [4113, 3370]. Received [931, 2052, 2066, 461, 474, 481, 489, 497, 504, 513, 520, 528, 534, 541, 548, 556, 563, 576, 586, 592, 598, 606, 613, 621, 628, 635, 642, 650, 655, 661, 669, 674, 682, 689, 696, 703, 709, 714, 721, 727, 734, 742, 751, 757, 767, 774, 780, 787, 795, 804, 813, 822, 829, 836, 842, 852, 859, 870, 877, 883, 889, 898, 904, 905, 921, 940, 958, 967, 975, 987, 997, 1005, 1033, 1044, 1018, 1025, 1052, 1062, 1072, 1078, 1088, 1095, 1103, 1110, 1116, 1126, 1135, 1142]. received [1148, 1159, 1168, 1174, 1185, 1196, 1201, 1209, 1223, 1231, 1239, 1246, 1253, 1261, 1267, 1275, 1281, 1287, 1294, 1300, 1307, 1313, 1322, 1331, 1338, 1346, 1354, 1362, 1369, 1377, 1384, 1390, 1398, 1404, 1411, 1418, 1426, 1433, 1440, 1448, 1456, 1463, 1471, 1478, 1485, 1493, 1500, 1509, 1514, 1520, 1526, 1534, 1540, 1549, 1557, 1568, 1577, 1586, 1593, 1599, 1605, 1613, 1620, 1626, 1633, 1639, 1646, 1655, 1663, 1669, 1676, 1683, 1690, 1696, 1702, 1710, 1719, 1725, 1733, 1740, 1747, 1756, 1772, 1780, 1785, 1790, 1795, 1803]. received [1810, 1816, 1822, 1829, 1838, 1846, 1854, 1861, 1869, 1878, 1891, 1895, 1900, 1910, 1920, 1931, 1941, 1946, 1954, 1963, 1971, 1977, 1991, 1997, 2029, 2037, 2043, 2073]. reciprocal [4253]. Recognition [2485]. Recollections [662, 965, 1136, 1092]. reconstruction [615]. record [118]. recording [928]. recovery [914]. Recreational [2420]. rectangular [240]. rectifier [201]. Recycling [1140]. red [4010, 495, 1182, 3865, 204, 210, 135, 258, 4090]. Rédei [3247]. Redner [2619]. redshift [3861]. reduced [3645]. reduction [3466, 4302]. Reed [4142, 4171, 1887]. Rees [3231]. Reference [2727, 4120, 4106, 3395, 3393, 3368, 2433, 2463, 2357, 4293, 125, 2565, 2751, 2757, 3038, 3095, 2672, 2838, 2836, 2853, 2785, 3194, 3179, 3353, 3408, 4240, 4272, 2764, 3207, 3092, 3168, 4084, 4251, 1068, 3083, 4173, 4347, 2687, 4072, 4190, 2394, 4323, 2366, 2567, 2954, 3407, 3126, 3427, 3455, 4339, 2372, 3312, 4353, 4318, 4319, 3143, 3233, 2523, 3269, 2437, 3230, 3446, 2292, 2299, 2526, 2309, 2465, 2464, 2584, 3246, 3367, 3030, 2647, 3449, 4205, 2587, 3430, 4326, 2332, 2784, 3127, 3033, 3012, 3086, 4122, 4349, 3213, 2415, 4223, 2835, 2837, 3062, 3119, 3156, 2581, 3183, 3347, 2607]. reference [2556, 2705, 2722, 2695, 4071, 4166, 2759, 2725, 2557, 4199, 4266, 2805, 3481, 2718, 2468, 2501, 2349, 2807, 2654, 3087, 2669, 2388, 2871, 3170, 3484, 2752, 2769, 4174]. reference/textbook [2357]. reflection [212]. Reflections [1778, 1808, 2984, 1819, 3918, 1117, 2311, 2535, 3909]. refraction [2093, 2270, 2269]. Refractive [2394]. Refrigeration [837, 462]. refrigerators [3966]. refugee [2363]. Regev [2972]. regimen [2149]. regional [2109]. Reich [3253]. Reiner [1069]. Reinhard [3531, 3558]. Reinventing [2963]. reionisation [3965]. Reissue [3308]. Related [1191, 3894, 2422, 3780, 1045, 1053, 2842, 1627]. relating [262]. Relation [3218, 1297, 1068, 1988]. Relationships [3930, 3310]. relative [1476]. Relatively [3037, 3302]. Relativistic [1906, 1979, 2698, 3620, 4308, 3473, 4168, 2801, 3538, 3884, 3999, 2903, 3094, 3278, 2832, 4168, 3843, 2540]. Relativity [243, 2911, 3150, 509, 2845, 3043, 2562, 3071, 3363, 2649, 3191, 3955, 4182, 1026, 132, 153, 3987, 1686, 457, 2861, 3756, 4036, 1631, 3864, 2149, 3473, 3508, 511, 2155, 1219, 4145, 3473, 2565, 2655, 2635, 2713, 2391, 2531, 2718, 3185, 3560, 3692, 326, 4163, 3552, 3917, 4351, 2727, 3037, 2611, 3461, 4176, 2768, 82]. relativity/General [3955]. Relaxation [347, 1089]. Release [2928]. Releases [834]. Releasing [1723]. relevance [2209]. Reliability [2670]. Religion [3195, 1983, 1950, 1761]. religious [3878]. Remarkable [2134]. reminiscences [1082, 1092, 3017]. Remo [3437, 3153]. remodelling [4241]. Remote [896, 2288, 2584, 3231, 1480, 1572, 1751]. renaissance [1138]. Renee [3525]. Renewable [4294]. Renn [4356, 4344, 3473]. Renormalization [3376, 3954, 2715]. Renzoni [3584]. repertoire [3935]. Report [144, 869, 293, 27, 2296, 464, 292, 3206, 391, 334, 2298]. Report/Survey [869]. Reports [379, 417, 225, 317, 66, 263, 554, 42, 316]. requirement [1442]. Resag [4047]. Resconi [3667]. Research [145, 53, 2670, 275, 437, 3206, 2395, 3276, 3063, 89, 136, 213, 241, 277, 259, 3776, 54, 194, 263, 2414, 675, 740, 2715, 2350, 2412, 3137, 4113, 20, 324, 1119, 223, 2428, 2342, 2421, 208, 942, 2129, 193, 2050, 2064, 4, 636, 924, 1325, 10, 2354, 3457, 3460, 3525, 3432, 719, 638, 2700, 801, 127, 168, 220, 264, 278, 349, 334, 2231, 2592, 2720, 2409, 2424, 2368, 2369, 2423, 2503, 792, 228, 1825, 119, 1414, 3243, 3299, 4057, 3642, 3653, 2724, 3596]. researcher [4120, 4078, 4060, 4106, 2686, 2883, 3157, 3395, 3368, 3211, 3333, 3052, 3350, 4110, 2926, 3061, 3079, 3098, 3201, 2667, 4241, 4217, 4284, 4275, 2614, 2555, 3022, 3180, 3506, 2819, 2648, 2583, 2911, 3241, 2767, 3097, 2840, 3107, 2593, 2949, 3298, 3365, 2830, 2522, 2715, 2774, 2858, 3124, 4210, 4273, 4271, 2544, 3376, 4231, 2795, 3116, 2751, 2856, 2727, 3091, 3017, 3406, 3412, 3320, 3425, 4154, 4221, 2757, 3349, 4356, 4328, 3038, 3095, 3035, 3247, 4307, 4118, 3296, 2984, 3023, 3373, 2798, 2838, 2836, 2853, 2785, 3275, 3194, 3179, 3353, 3408, 3499, 4240, 4272, 2617, 2775, 2779]. researcher [3207, 3370, 2621, 3075, 3168, 2735, 2872, 2755, 3502, 4248, 2766, 3083, 3245, 3165, 2791, 2841, 2541, 2999, 3036, 3047, 4072, 4178, 4190, 4198, 4252, 2776, 3118, 2965, 3279, 3494, 3501, 3450, 3452, 4081, 4235, 4269, 2693, 4323, 2660, 2800, 3054, 3204, 3407, 2887, 3126, 3205, 4085, 4108, 4068, 4119, 4196, 3184, 4242, 4250, 2850, 3081, 3074, 4339, 3099, 4187, 2538, 2820, 3046, 3196, 3431, 3143, 3233, 3366, 2670, 3024, 2771, 3050, 2898, 4165, 3071, 3169, 3303, 3398, 3492, 2524, 2716, 3510, 3128, 3274, 3497, 3382, 2703, 3080, 3005, 2808, 4157, 3230, 3210, 3446, 4186, 2539, 3340, 3423]. researcher [2682, 2526, 3026, 3003, 3424, 3189, 3511, 2680, 3015, 3162, 3447, 3208, 3206, 4321, 3367, 3030, 2685, 3084, 3145, 3031, 3259, 3232, 3307, 3449, 3486, 4200, 4205, 4224, 3065, 2622, 3430, 2839, 3070, 2652, 3272, 2651, 2432, 2784, 2712, 3127, 3033, 3012, 4122, 4197, 4349, 3209, 3213, 3094, 3332, 2780, 2624, 2835, 2874, 2837, 2834, 2760, 2750, 2880, 3062, 3119, 3063, 3172, 3409, 4054, 4083, 4189, 2580, 3256, 3347, 3402, 2546, 3414, 3476, 3489, 4255, 4219, 4301, 4311, 4310, 4276, 4350, 4166, 4167, 2759, 3138, 3346, 2725, 2557, 4199, 4334, 2701, 2896, 2741, 3228, 3461]. researcher [4253, 4265, 4266, 4332, 2879, 3148, 3059, 3146, 3141, 3319, 3362, 3339, 3391, 3439, 4063, 4220, 4174, 2422, 2530, 2758, 2398, 2400, 2619, 3191, 3185, 3278, 3508, 4290, 3136, 2799, 3147, 2807, 2826, 2654, 3087, 3032, 2665, 2666, 2832, 3170, 3306, 3484, 3475, 3474, 3495, 4058, 4116, 4067, 4172, 4215, 3397, 3067, 3267, 3300, 3348, 4111, 4185, 4168, 2547, 2882, 2852, 2586, 2769, 4056, 4231, 4356, 4118, 4052, 4235, 4257, 4259, 4309, 4295, 4108, 4119, 4339, 4294, 4157, 4159, 4160, 4121, 4349, 4169, 4170, 4234, 3476, 4077, 4166, 4167, 4260, 4297, 4253, 4063]. researcher/scientist [3348]. researchers [2921, 2989, 2596, 2927, 3051, 2458, 2486, 2659, 2485, 2828, 3011, 2425, 2868, 2955, 2865, 2293, 2866, 2945, 2948, 2939, 2988, 2931, 2907, 2341, 2342, 2421, 2563, 2494, 3004, 2930, 2366, 2954, 3007, 2597, 2489, 2674, 2566, 2609, 2816, 2706, 2918, 3013, 4283, 2932, 2340, 2351, 2461, 2476, 2616, 2525, 2902, 3187, 3186, 3182, 3454, 2929, 2649, 2653, 2743, 2739, 2803, 2940, 2884, 2946, 2998, 3432, 2869, 2295, 2467, 2332, 2815, 2924, 2943, 2909, 2416, 2419, 3014, 2925, 3404, 3410, 3440, 3482, 3441, 3481, 3483, 3504, 3503, 2723, 2451, 2399, 2444, 2831, 2894, 2454, 2470, 2915, 2920]. researchers [2967, 2919, 2704, 4107, 3114, 4187]. reservoirs [914]. resistance [607, 4262]. resistivity [1210]. resolution [2382, 1967, 3727, 657, 4224, 1076, 1203, 352, 1120, 3537, 3795]. Resolved [2288, 1736]. Resolving [4195, 4013]. Resonance [3373, 2975, 3026, 2419, 2592, 1459, 631, 2699, 1225, 1934, 818, 872, 359, 369, 47, 4315, 839, 1421, 2285]. Resonances [831, 3566]. Resonant [1356, 1677]. Resonant-bar [1356]. resonators [3743]. resource [776]. respectable [2089]. results [1529, 2126, 3593]. Retelling [1917]. Retirement [2787]. return [3419]. Revealed [3302, 3722]. reversal [1726, 1647]. Reversibility [3023]. Review [3670, 3806, 3834, 3817, 3757, 3776, 3767, 3807, 3766, 3942, 4133, 4132, 3622, 3711, 3953, 4120, 4078, 4060, 4106, 3878, 4149, 3520, 3651, 3741, 3841, 4010, 3909, 4057, 3764, 3763, 1121, 3618, 3986, 3988, 4317, 4324, 3790, 3868, 3805, 3774, 3946, 4016, 4062, 4110, 4211, 4137, 4225, 4286, 3885, 3896, 2893, 2772, 3560, 3523, 3574, 3721, 3583, 3719, 3541, 3559, 3830, 4082, 4343, 1265, 1191, 440, 662, 675, 848, 4241, 4217, 4284, 4275, 3796, 3974, 4202, 4232, 4268, 3785, 3927, 4135, 3698, 3699, 3700, 3750, 3828, 3918, 3522, 3696, 3619, 3901, 3772, 3912, 4140, 4139, 4254, 4038]. Review [4000, 3934, 3539, 3580, 3625, 3710, 3635, 3835, 3899, 3898, 3872, 3873, 3944, 3984, 4032, 4009, 3911, 4056, 4055, 4210, 4273, 4271, 4304, 4296, 4206, 3656, 3716, 3644, 3840, 3943, 4035, 4231, 4293, 205, 3425, 3677, 3695, 3723, 3579, 3609, 3692, 3800, 3875, 3791, 3748, 3759, 3848, 3855, 3881, 4005, 3945, 4045, 3913, 3935, 3914, 4018, 3938, 3955, 4028, 4073, 4123, 4069, 4154, 4182, 4177, 4134, 4184, 4191, 4221, 3646, 3839, 3877, 4027, 4100, 4113, 4356, 4328, 3715, 3663, 3869, 4307, 1040, 4118, 3599, 3585, 3900, 3738, 3947, 3919, 4061, 4333, 1083, 3554, 3544, 3657, 3658]. Review [3659, 3744, 3747, 2057, 1131, 3538, 4143, 4209, 3628, 3629, 3706, 3889, 3961, 4183, 4240, 4272, 223, 2329, 3102, 3524, 3530, 3652, 3591, 3648, 3844, 3758, 3931, 4030, 4084, 4155, 4248, 3647, 4251, 112, 326, 4325, 4327, 4355, 3620, 4291, 4303, 4285, 4138, 4173, 3650, 3637, 3858, 3845, 3859, 3784, 3856, 3987, 3989, 3958, 4347, 4344, 4348, 965, 1092, 3626, 3679, 3883, 3762, 3793, 3739, 3983, 3930, 4026, 4072, 4052, 4178, 4190, 4148, 4198, 2191, 4163, 4252, 3561, 3557, 3617, 3616, 3895, 3808, 3894, 3753, 3755, 3809, 3960, 4006, 4021, 3962, 4081, 4156, 4204, 4130]. Review [4203, 4162, 4235, 4222, 4257, 4259, 4269, 4309, 4281, 4270, 4295, 4274, 4305, 4323, 875, 3771, 3727, 1069, 3884, 4025, 3948, 4002, 3915, 3925, 4001, 4046, 4043, 3975, 4041, 3920, 3954, 3999, 4029, 4044, 4008, 4024, 4085, 4108, 4068, 4119, 4098, 4196, 4141, 4242, 4250, 3552, 3553, 3564, 3705, 3630, 3669, 3668, 3588, 3709, 3576, 3671, 3708, 3577, 3627, 3631, 3865, 3816, 3754, 3886, 3861, 3854, 3891, 3860, 3804, 3801, 3797, 3752, 3769, 3773, 3751, 3787, 4031, 4037, 4047, 3993, 4040, 3994, 4048, 3957, 4042, 3936, 4003, 4015, 3910, 4090, 4091, 4105, 4146, 4345, 4338, 4342, 4346, 4339, 4340, 4187, 4283, 4294]. Review [3661, 2765, 3551, 4322, 4353, 4318, 4319, 316, 3660, 3623, 3998, 4165, 850, 3540, 3810, 3980, 3592, 3421, 4302, 3563, 3794, 3921, 4097, 4107, 4157, 4336, 4292, 3825, 4150, 4186, 2644, 3423, 3578, 3587, 3608, 3678, 3597, 3832, 4007, 4017, 3982, 1100, 3525, 3717, 3614, 3632, 3862, 3078, 4112, 4114, 4115, 4117, 4159, 4161, 4158, 4160, 3720, 1057, 3768, 3770, 3843, 3952, 4321, 3562, 3531, 3653, 3691, 3589, 3874, 3882, 3789, 3740, 3950, 4099, 4093, 4200, 4151, 4205, 4224, 3819, 3978, 3533, 3532, 3697, 3849, 3926, 3527, 3712, 3756, 3973, 4164, 3665, 3749, 3846]. Review [3905, 4019, 4282, 4326, 56, 1165, 4121, 4131, 129, 3718, 3666, 3703, 3842, 4122, 4197, 4349, 3521, 4337, 4354, 4201, 4223, 4244, 4218, 4249, 4316, 3558, 3584, 3582, 3607, 3634, 3649, 3682, 3683, 3735, 3736, 3833, 3734, 3879, 3880, 3795, 3827, 3836, 3922, 3959, 3917, 4020, 4022, 4036, 3949, 4054, 4083, 4053, 4080, 4079, 4169, 4170, 4189, 4234, 3604, 1149, 2552, 3314, 3636, 3923, 4144, 4351, 4341, 4352, 3893, 3811, 3863, 3932, 3724, 3613, 3704, 3526, 4306, 3690, 3476, 3543, 3645, 3693, 3681, 3864, 3831, 3777, 3803, 3996, 3995, 3929, 3933, 4071, 4129]. Review [4255, 4208, 4188, 4219, 4301, 4311, 4310, 4276, 4350, 4320, 4070, 4066, 4077, 4166, 4167, 3654, 3760, 1140, 4199, 4260, 4297, 4308, 4334, 2701, 3337, 4095, 4096, 4094, 4253, 4265, 4266, 4332, 3537, 3566, 3536, 3565, 3567, 3534, 3575, 3595, 3610, 3612, 3674, 3673, 3596, 3611, 3672, 3903, 3814, 3887, 3904, 3742, 3745, 3888, 3902, 3813, 3890, 3743, 3812, 3939, 3941, 3940, 4065, 4064, 4063, 4220, 3281, 3555, 3528, 3529, 3702, 3701, 3871, 3838, 3870, 3837, 3979, 4086, 4092, 4142, 4147, 4171, 4174, 3786, 3788, 4216, 1039, 3928, 1562, 4290, 3977, 3937, 4034, 3976, 3713, 3733, 3815, 4059, 3985]. Review [4109, 3605, 3542, 3568, 3535, 3556, 3598, 3621, 3606, 3675, 3624, 3667, 3615, 3722, 3581, 3684, 3707, 3586, 3662, 3798, 3818, 3847, 3867, 3857, 3829, 3737, 3775, 3792, 3876, 3802, 3850, 3892, 3897, 3765, 4004, 4039, 3956, 4033, 3924, 3997, 4023, 3981, 3916, 4058, 4116, 4067, 4172, 4136, 4207, 4176, 4215, 4233, 4335, 3676, 4145, 4175, 1844, 4258, 3519, 3685, 3655, 3688, 3761, 4111, 4185, 4168, 3594, 3590, 3593, 2041, 1146, 44, 834, 3633, 3680, 3694, 3714, 3746, 3866, 3799, 3951, 4243, 4256, 4267, 2099, 3468, 19, 18, 8, 82, 149, 155, 266, 656, 758, 835, 1141, 1147, 1625, 1996, 3052]. review [2781, 3470, 2699, 3098, 3201, 1778, 1808, 1800, 584, 2614, 2596, 2767, 1820, 2034, 2840, 2485, 3298, 1836, 3472, 2868, 1177, 2544, 3154, 3116, 2459, 2662, 2904, 3091, 2082, 2754, 3371, 2866, 2977, 2942, 2740, 2867, 2945, 2801, 2948, 3055, 2995, 3142, 3150, 2382, 2763, 3442, 100, 3296, 1777, 1852, 3322, 2749, 2798, 2853, 3104, 3275, 3194, 2975, 2617, 2747, 1783, 3207, 3444, 553, 2941, 3168, 9, 3328, 2791, 1857, 3547, 3040, 3004, 146, 1938, 1801, 2693, 3851, 1770, 3126, 3184, 2850, 2737, 3113, 3642, 3643, 3689, 3732, 3783, 1359, 1843, 2035, 2678, 2538, 2675, 2886, 2980]. review [3224, 3262, 3163, 3287, 3381, 2690, 3093, 2340, 2612, 3143, 2670, 4180, 3469, 4051, 3274, 3227, 81, 3108, 3202, 1560, 2714, 3230, 3210, 3345, 2379, 2863, 3823, 3003, 3189, 3511, 1851, 2688, 2748, 3206, 2288, 3428, 2653, 3049, 3304, 3299, 3065, 3271, 31, 1093, 2622, 3305, 1819, 160, 2664, 3240, 2001, 1813, 2467, 3106, 537, 1653, 2712, 1716, 2595, 2624, 2880, 3062, 3172, 1858, 2580, 2657, 2943, 2825, 2909, 2782, 3109, 3256, 111, 3132, 1937, 2611, 3731, 2860, 2976, 3198, 2745, 3471, 2725, 2620, 2741, 2895, 3473, 2582, 2531, 2589, 2588, 2913, 3285, 1993]. review [2361, 3191, 3185, 1814, 3032, 2848, 2536, 2787, 1827, 3355, 869, 188, 2915, 2919, 1835, 1936, 3102, 3934, 3092]. Reviewed [2074, 2075, 2151, 2152]. Reviews [2051, 2058, 2065, 2072, 2100, 2106, 2112, 2125, 2130, 2135, 2140, 2214, 2218, 2222, 2232, 2280, 2284, 45, 57, 67, 130, 173, 181, 189, 199, 207, 215, 226, 233, 244, 257, 271, 280, 294, 301, 309, 318, 327, 336, 343, 351, 357, 367, 373, 383, 392, 400, 408, 418, 426, 433, 443, 449, 460, 467, 473, 480, 488, 496, 503, 512, 519, 533, 562, 569, 575, 585, 591, 599, 612, 620, 627, 634, 641, 649, 663, 668, 676, 681, 688, 695, 702, 708, 715, 720, 726, 733, 741]. reviews [750, 766, 773, 779, 786, 794, 803, 814, 821, 828, 841, 849, 851, 858, 876, 882, 888, 897, 910, 920, 930, 939, 948, 956, 966, 973, 985, 995, 1003, 1016, 1024, 1031, 1043, 1051, 1060, 1071, 1077, 1086, 1094, 1102, 1109, 1115, 1125, 1134, 1158, 1167, 1173, 1184, 1195, 1200, 1208, 1216, 1222, 1230, 1238, 1245, 1252, 1260, 1266, 1274, 1280, 1286, 1293, 1299, 1306, 1312, 1319, 1329, 1337, 1344, 1352, 1360, 1367, 1375, 1382, 1389, 1396, 1403, 1410, 1417, 1425, 1432, 1439, 1446, 1455, 1462, 1470, 1477]. reviews [1484, 1491, 1498, 1508, 1513, 1519, 1525, 1533, 1539, 1548, 1556, 1567, 1576, 1585, 1592, 1598, 1604, 1611, 1619, 1632, 1638, 1645, 1654, 1662, 1668, 1675, 1682, 1689, 1695, 1701, 1709, 1718, 1724, 1732, 1739, 1746, 1755, 1763, 1764, 1771, 1779, 1784, 1789, 1794, 1802, 1809, 1815, 1821, 1828, 1837, 1845, 1853, 1859, 1867, 1876, 1884, 1889, 1901, 1908, 1919, 1929, 1939, 1945, 1953, 1961, 1970, 1976, 1985, 1990, 2004, 2011, 2020, 2028, 2036, 2042, 2084, 2092, 2119, 2145, 2150, 2158, 2168, 2172, 2178, 2182, 2187, 2192, 2197]. reviews [2203, 2207, 2210, 2226, 2237, 2274, 1335, 1904, 454, 3335, 547, 527, 540, 2840, 3371]. Revised [3349, 2706, 3163, 3196, 3065, 3817, 2560, 4187, 3132]. revisited [1835]. revivals [2195]. revolution [3759, 4356, 3908, 23, 4344, 1721, 2554, 3327, 3177, 4248, 4037, 3982]. revolutionaries [1445]. revolutionized [3532]. Revolutions [2658, 4386, 1907, 3502, 3121, 2644]. Reyes [3189]. Reyes-Lega [3189]. Reynolds [4026, 2337]. Rez [3807]. Rezzolla [3278]. Rhiger [2775]. Riazuddin [3401, 2855]. Ricardo [4086]. Riccardo [2741, 4220]. Ricciardi [3740]. Rice [4139]. Richard [4057, 3199, 2786, 3384, 3261, 2754, 3659, 2931, 3647, 4355, 3650, 3219, 2474, 2487, 4122, 2780, 3734, 3347, 3198, 2468, 2710, 3556, 2787, 1872]. Richards [3067]. Richardson [3644]. Richmond [3248]. riddle [4328, 4348, 3336]. Ridges [3454]. Ridley [3257, 3742]. Ridolfi [2833]. Ridpath [2871]. Rieckers [3558]. Rieke [3049]. Rift [879]. rigid [4231]. Riley [3212, 3193, 4084, 3027]. Ring [284, 1262]. Ring-laser [284]. rings [141, 680, 3394]. Ringström [3241]. Ríos [4118]. Ripoll [3058]. ripple [464]. Ripples [3284]. Risk [2047, 1067, 1068]. Risto [2804]. rivalries [410]. Rivalry [2980]. river [1873]. Rivier [2259]. Rivkin [4054]. Road [3088, 2854, 348, 3954, 3473]. Rob [3614]. Robbins [293, 292]. Robert [3211, 3692, 3102, 4303, 1092, 3436, 4283, 4294, 3312, 3374, 4388, 2499, 3486, 2664, 2778, 2595, 2611, 4332, 2589, 2638, 3285, 3701, 3178, 1092]. Roberto [3557, 3132]. Roberts [3789]. Robin [3112]. Robinson [3151]. robust [3328]. Roche [4081, 3265]. Rochester [137]. rock [3463]. Rocket [216]. rockets [2453]. rocky [4150]. Rodriguez [2614]. Roe [4148]. Rogel [3541]. Rogel-Salazar [3541]. Roger [4184, 4190, 3381, 3636, 4176]. Rogers [2320]. Rogue [3572, 2643]. Rohde [4248]. Roisin [2976]. Rojas [3240]. Roland [3563]. Role [1595, 710, 2252, 324, 4155, 525, 3358, 1677, 572, 1987, 2236, 3357, 2251]. ROM [3276, 3314]. Roman [3164]. Romatschke [4168]. Roming [3039]. Romo [4170]. Ron [2470]. Ronald [2829, 4345, 3429, 2874, 2632, 3085]. Ronde [3320]. Ronning [4325]. roots [4356, 4344]. Roques [2476]. Rory [2709]. Rosanne [3075]. Rosenbaum [3539]. Rosenblum [2796]. Ross [2826, 4145]. Rosser [326]. Rossi [3594]. rotary [2304]. Rotating [2501, 652, 984, 229, 3867, 2941]. Rotation [284, 2501, 1191]. rotations [1735, 4231]. Roth [2779]. rotons [1955]. Rottwitt [3619]. rough [1215]. roughness [1023]. Roustem [4241]. Routledge [4296, 4257, 4305, 4294, 4336, 4349, 4297, 4308]. Rouven [4028]. Rovelli [4182, 2298]. Rowan [2517, 3151]. Rowan-Robinson [3151]. Rowe [2816]. Rowlands [4144, 3261, 3587]. Roy [3203, 4309, 3708]. Royal [253, 811, 2506, 194, 64, 397, 53, 394, 261, 186, 214, 276, 350]. Rozelot [2501, 2898]. Rubakov [4008, 2815]. Rubber [588, 3359, 6]. Rubin [3040]. Ruda [2372]. rudiments [1524]. Rudolf [3072, 4017, 3015, 1711]. Rudolph [2756]. Ruffini [3437, 3153]. Rukhadze [2659]. rules [2219]. Rumford [1937]. Rumsfeld [2229]. Rupert [2437]. Rurali [4220]. Russell [3627, 2472, 3059]. Russenschuck [2757]. Russian [100, 4124]. Russo [3814]. Ruth [4157, 3086]. Rutherford [380, 291, 145, 291, 196, 148, 662]. Ryabova [4058]. Ryan [2874, 2620, 3815]. Rydberg [2253, 1518, 1541, 916, 712, 2254]. Rydbergs [4118]. Ryden [3723]. Ryutin [3593]. Ryutova [3957].

S [4324, 2502, 4110, 3470, 2893, 1191, 3506, 2386, 2485, 3203, 2439, 2403, 4055, 4273, 2425, 3644, 2770, 2727, 3945, 4184, 3123, 3744, 2838, 2779, 2626, 3168, 2872, 2342, 4355, 3793, 4190, 2394, 4162, 4257, 2366, 2429, 2567, 4001, 4044, 3297, 3057, 3074, 2996, 3477, 3668, 3709, 3993, 4048, 4015, 2674, 2609, 2372, 2348, 2678, 2251, 2249, 2333, 4319, 3258, 3235, 2379, 2355, 2688, 2881, 2738, 4159, 2899, 2773, 3768, 3770, 3140, 2649, 2803, 3533, 2818, 3045, 3019, 4019, 4121, 3718, 2815, 2633, 2874, 4036, 4054, 2393, 3604, 2728, 3811, 3132, 3995, 4276, 4350, 3014, 3760, 2557, 4297]. S [2322, 2317, 3566, 3536, 3596, 4063, 3702, 2540, 2424, 2398, 2669, 2599, 2413, 2591, 2663, 2885, 2934, 2969, 3041, 3684, 4067, 4176, 3676, 3761, 2345, 2041, 2547, 2967, 1146, 3633, 3714, 4267, 2450]. S.-H [2355]. S.-I [2818]. S.I. [390]. S245 [2467]. S263 [2737]. S264 [2898]. S266 [2780]. S279 [3039, 3267]. S282 [3067]. S287 [3203]. S291 [3267]. S292 [3348]. Sá [4119]. Saad [4235]. Saarloos [3002]. Sabine [2923]. Sachdev [2893, 2728]. Sachs [2747]. Sadamichi [3184]. Sadana [2673, 2673]. Sadao [3083]. Sadegh [4341, 3504]. Sadenius [2751]. Sadri [2552]. Sadruddin [3688]. safe [1501]. safety [1866, 2035, 1104, 348, 3954]. saga [2407]. Sagan [1968]. Sagir [4254]. Saha [3543]. Sahakian [4129]. Saikia [2934]. Sailing [2407, 2588]. Saitoh [3184]. Salam [2870, 3976, 2239]. Salasnich [3393]. Salazar [3541]. Salby [2890]. Saleem [3073]. Salzburg [193]. Samaj [3591]. Samei [4165]. Samorì [2795]. Sample [2564]. Samrath [2712]. Samuel [3510]. Samya [4188]. Sanchez [2618]. Sanchez-Lavega [2618]. sand [1792]. Sanda [2586]. Sander [2401]. Sanders [3816, 2664]. Sandoghdar [3772]. Sándor [3764]. Sandra [2463, 4224]. Sanjay [4178]. Sanjuan [4324, 3566]. Sankar [3540]. Sanpera [3022]. Santen [4099]. Santos [3659]. Santosh [3041]. Sanz [2979]. Sara [2654]. Saraceno [2970]. Sarantsev [3218]. Sarath [2775]. Sarbani [4007]. Saris [2534]. Saskia [3379]. Satellite [102, 553, 261]. satellites [840, 572, 262]. Sato [3544]. Sattler [3307]. Saturn [4042]. Satya [2750, 3543]. Satz [4311]. Savart [238, 133]. Saxby [2684]. Sayed [3504]. Saylan [4290]. Scalar [4044]. Scalas [2634]. scale [372, 4277, 1758, 4417, 2733, 4346, 4151, 3534, 321, 2893, 4041]. scales [2161, 772, 3601]. Scaling [2410, 3920]. scalpel [116]. Scanning [502, 1502, 1563, 2076]. Scarani [3680]. scatter [729, 1775]. Scattering [3118, 756, 3382, 725, 2633, 2227, 2347, 881, 1187, 603, 761, 693, 788, 607, 1076, 622, 830, 3675, 2955, 3321, 3236, 3835]. Scenes [3197]. sceptical [72]. Scerbo [3653]. Scerri [3244]. Schadschneider [2745]. Schäfers [3098]. Schakel [2349]. Scharf [2766]. Scherzer [2973]. Schiavone [2417]. Schieve [2454]. Schilling [3001, 1146]. Schleich [2296]. Schleicher [4033]. Schlieren [300]. Schliesser [2983]. Schmidt [3483, 3175]. Schmitz [4018]. Schneider [4275]. Schoelkopf [3374]. Scholarly [2041]. Scholl [2528]. School [4124, 2846, 372, 4283, 4117, 290, 4326, 3836, 2322, 159, 206, 3389, 3772, 2760]. schools [194, 176, 52, 131, 117, 120, 231]. Schörner [2751]. Schörner-Sadenius [2751]. Schottky [1118]. Schötz [3696]. Schreier [2766]. Schrijver [2993, 2440, 2721]. Schrödinger [3056, 2995, 3727, 4105, 1235, 2262, 3640, 3600, 4069, 2261, 3239]. Schroeter [4066]. Schulman [4276]. Schulz [2737]. Schumacher [2536]. Schuster [2840]. Schwartz [3810, 3316]. Schweber [3945]. Schwinger [3819]. sci [1243]. sci-fi [1243]. Science [3841, 43, 159, 206, 1865, 4202, 4232, 2527, 2596, 3051, 2843, 4140, 740, 40, 2707, 3371, 2977, 1952, 4209, 3139, 2342, 1972, 4173, 4347, 2930, 1687, 2684, 177, 2794, 3343, 221, 1983, 3400, 2506, 2668, 2675, 3196, 3128, 3274, 3445, 2367, 101, 4161, 2884, 3084, 3234, 3238, 1761, 1165, 1883, 4201, 4244, 4249, 4397, 2880, 4144, 4306, 3426, 4253, 2582, 4142, 4174, 85, 4403, 3085, 3284, 3372, 4411, 868, 3806, 4078, 2010, 4057, 69, 167, 342, 358, 4225, 1778, 3700, 3918, 3472, 3984, 307, 397, 3945, 4103, 4027, 4113, 2945, 2061, 3908, 2749, 3226, 3308, 1982, 885]. science [2089, 2504, 3637, 3989, 1915, 289, 341, 52, 4047, 2407, 2765, 3120, 3335, 2397, 1636, 3998, 14, 30, 2353, 26, 98, 2437, 2642, 1295, 2473, 4186, 2371, 2435, 2644, 2206, 235, 4117, 945, 3768, 4321, 2517, 4099, 279, 3571, 315, 4019, 1393, 2380, 2581, 2552, 2854, 4125, 3995, 4350, 25, 1461, 3702, 231, 1968, 3664, 1616, 1327, 3729, 1827, 106, 74, 362, 4256, 145, 180, 221, 437, 3180, 3444, 2548, 2980, 2724, 2671, 4223, 2451, 2507, 3306, 2534, 2456, 3300, 3766, 3721, 4221, 4291, 4204, 3649, 3890, 3997, 3195, 3525, 2773, 2884]. Sciences [3212, 3026, 187, 2288, 2378, 2804, 237, 245, 3060, 48, 2960, 577, 79, 2173, 3212, 3193, 2623, 3340, 3188, 3096, 2833, 2804, 4148, 4093]. Scientific [4132, 4106, 4010, 3909, 4324, 2463, 4062, 4225, 1907, 4268, 4359, 4056, 4055, 4231, 4293, 2742, 2856, 4028, 4123, 4069, 2082, 4113, 2057, 846, 4418, 3448, 4030, 4138, 4163, 3999, 4029, 4044, 4008, 4024, 4108, 4068, 4119, 4098, 4196, 3219, 3093, 3821, 2670, 3998, 1343, 739, 4124, 3423, 4112, 4114, 4099, 3243, 391, 4164, 4121, 4223, 4083, 4080, 3996, 4255, 4070, 4167, 3337, 4095, 4096, 4265, 4086, 4004, 3956, 4033, 3997, 4172, 3164, 4243, 4057, 18, 70, 43, 2661, 3774, 2509, 324, 2672, 23, 2504, 3844, 2352, 331, 4037, 2050, 2064, 807, 3660, 1029, 3525, 941, 31]. scientific [3724, 2278, 3612, 2426, 2413, 4067, 2528, 2512, 3259, 3649]. Scientist [3178, 4120, 4078, 4060, 4106, 2686, 2883, 3122, 3395, 3318, 3368, 3333, 3201, 4241, 4217, 4284, 4275, 2851, 2843, 2949, 4206, 3154, 3116, 2770, 3091, 3167, 4191, 3349, 3038, 3095, 3035, 3247, 2798, 2785, 3104, 3275, 3194, 3179, 3353, 3408, 3499, 3488, 4272, 2747, 3448, 3502, 2791, 4072, 4190, 4198, 4252, 2702, 4323, 2660, 2800, 3407, 3205, 4068, 3184, 4242, 4250, 2709, 3088, 3196, 3375, 2612, 3143, 3366, 2670, 2771, 3050, 2891, 3128, 3227, 3080, 3005, 3230, 3210, 4186, 3026, 3003, 3424, 2680, 2881, 2309, 2465, 2464, 2671, 2584, 3246, 3367, 2685, 3145, 3031, 3259, 3232, 3307, 3449, 4200]. scientist [4205, 4224, 3049, 3234, 3065, 3430, 3125, 3305, 2784, 2780, 2624, 2835, 2874, 2837, 2834, 2750, 2880, 3062, 3119, 3063, 3172, 3409, 4189, 2782, 3347, 3132, 2722, 3414, 4219, 4276, 3346, 2725, 4199, 4334, 2608, 2620, 2896, 2814, 2741, 3197, 3461, 4266, 2503, 2799, 2994, 2665, 2666, 2832, 3484, 3474, 4215, 3067, 3267, 3300, 3348, 4185, 2882, 3164, 2769, 4110, 4073, 4154, 4240, 4155, 4085, 3421, 4302, 4054, 4255, 4301, 4311, 4350, 4332, 4290, 4058, 4116, 4067, 4111, 4168, 2870, 3139]. scientist/engineer [2309, 2465, 2464, 2671, 2584]. Scientists [2901, 2921, 222, 4211, 1754, 3974, 2955, 2865, 3371, 100, 2941, 2907, 9, 4084, 2294, 2954, 2944, 3219, 2548, 2849, 2980, 961, 3283, 3457, 3460, 2355, 2929, 3140, 2518, 2649, 2653, 2499, 2998, 2487, 4197, 2924, 2825, 2895, 4095, 4265, 2928, 1644, 2915, 2920, 2967, 2837, 2317, 2967, 3808, 3755, 4034, 3918]. scientists/condensed [2487]. scintillations [783]. scissors [3463]. Scope [4120, 4078, 4060, 4106, 4057, 4317, 4324, 4062, 4110, 4082, 4343, 4241, 4232, 4254, 4056, 4055, 4304, 4296, 4231, 4293, 4073, 4123, 4069, 4154, 4177, 4184, 4113, 4356, 4328, 4307, 4118, 4061, 4333, 2057, 4240, 4084, 4155, 4248, 4251, 4325, 4327, 4355, 4291, 4303, 4173, 4347, 4344, 4348, 4072, 4052, 4178, 2191, 4163, 4252, 4081, 4156, 4162, 4235, 4257, 4259, 4309, 4295, 4305, 4323, 4085, 4108, 4068, 4119, 4242, 4250, 4105, 4345, 4338, 4342, 4346, 4339, 4340, 4294, 4322, 4353, 4318, 4319, 4165, 3421, 4302, 4107, 4157, 4336, 4292, 4112, 4114, 4115, 4117, 4159, 4161, 4158, 4160, 4321]. Scope [4164, 4326, 4121, 4122, 4349, 4337, 4354, 4244, 4249, 4316, 4054, 4083, 4053, 4080, 4079, 4169, 4170, 4234, 4351, 4341, 4352, 4306, 3476, 4071, 4255, 4301, 4311, 4310, 4350, 4320, 4070, 4066, 4077, 4166, 4167, 4260, 4297, 4308, 4334, 4253, 4332, 4064, 4063, 4086, 4171, 4174, 2501, 4290, 4059, 4109, 4058, 4116, 4067, 4172, 4176, 4233, 4335, 4175, 4258, 4111, 4168, 869, 4243, 4256, 4065, 2921, 2312, 2686, 2883, 3122, 3157, 3395, 3318, 3393, 3368, 3025, 3111, 3211, 3333, 3052, 3112, 3199, 3350, 2433, 2463, 2462, 2661, 2502, 2788, 2829, 4211, 4225, 4286, 2414, 2781, 2346, 2772, 2809]. Scope [2786, 2926, 2756, 2989, 3061, 3073, 3079, 3295, 3480, 3498, 3098, 3201, 2667, 2296, 2297, 2357, 3509, 2374, 4217, 4284, 4275, 2491, 2560, 2614, 2555, 2936, 3115, 3022, 3180, 3458, 3506, 4202, 4268, 2290, 2897, 2819, 3301, 2648, 2583, 2911, 3241, 2527, 2596, 2767, 2927, 3051, 3097, 2851, 2840, 2843, 3107, 2385, 2386, 2458, 2486, 2659, 2485, 2593, 2949, 2828, 3011, 3203, 3298, 3365, 2830, 2335, 2439, 2403, 2334, 2375, 2522, 2715, 2774, 2858, 3029, 3124, 2966, 3313, 3491, 3485, 4210, 4273, 4271, 2425, 2868, 4206, 2411, 3072, 2544, 3376, 2585, 3154, 2795]. Scope [2955, 2865, 3116, 3384, 2311, 2356, 2471, 2350, 2459, 2615, 2679, 2545, 2495, 2658, 2600, 2565, 2564, 2508, 2662, 2490, 2498, 2912, 2857, 2870, 2742, 2707, 2751, 2904, 2770, 2856, 2727, 3091, 2981, 3089, 3056, 3017, 3037, 3133, 3010, 3229, 3237, 3261, 3167, 3195, 3171, 3403, 3406, 3412, 3320, 3383, 3462, 3425, 4182, 4191, 4221, 2293, 2320, 2412, 2509, 2754, 2347, 2757, 3349, 3371, 3137, 2866, 3038, 2977, 3095, 3035, 3247, 2942, 2740, 2867, 2945, 2801, 2948, 3055, 2995, 3142, 3150, 3442, 3123, 3296, 2823, 2939, 2984, 3023, 2672, 2708, 3322, 3373, 3507, 4209, 2749, 2842, 3396]. Scope [2673, 2798, 2838, 2836, 2853, 2785, 3104, 3275, 3194, 3179, 3353, 3408, 3499, 3493, 3488, 4183, 4272, 2764, 3102, 3226, 3308, 2975, 3076, 3144, 3212, 3193, 2617, 2747, 2428, 2696, 2775, 2779, 3207, 2931, 3092, 3370, 3444, 2941, 2504, 2621, 2907, 3075, 3152, 3168, 2735, 2790, 2937, 2872, 2755, 3139, 3192, 3302, 3448, 3502, 2352, 2906, 2766, 2341, 2342, 2421, 2563, 2494, 2792, 2878, 2847, 3083, 3105, 2992, 3245, 3165, 3257, 2791, 2841, 4285, 2541, 2533, 2655, 2301, 2392, 3040, 3004, 2999, 2561, 2687, 2845, 3036, 2901, 2806, 2930, 3047, 3009, 3117, 2684, 4190, 4198, 2294, 2394, 2529, 2625, 2537]. Scope [2777, 2794, 2776, 3118, 2965, 3181, 3279, 3280, 3177, 3343, 3344, 3494, 3496, 3501, 3450, 3452, 3451, 4204, 4203, 4222, 4269, 4281, 4270, 4274, 2702, 2693, 2302, 2366, 2429, 2567, 2542, 2660, 2800, 2954, 2944, 3054, 3204, 3323, 3407, 2601, 2887, 3126, 2979, 3205, 4196, 3184, 3321, 2850, 3297, 2726, 2737, 3113, 3007, 3081, 3057, 3043, 2993, 3044, 3074, 3016, 2996, 3268, 3190, 3266, 3219, 3377, 3341, 3315, 3317, 3400, 3309, 3394, 3364, 3478, 3435, 3438, 3479, 3427, 3437, 3436, 3456, 3477, 3422, 3455, 3459, 2289, 2313, 2457, 2597, 2489, 2674, 2566, 2609, 2528, 2816, 2706, 2918, 3013, 3099, 4187, 4283, 2330, 2318]. Scope [2372, 2348, 2402, 2387, 2452, 2519, 2548, 2678, 2538, 2506, 2668, 2675, 2932, 2849, 2765, 2709, 2820, 2886, 2783, 2746, 3131, 3088, 3046, 2980, 3120, 3254, 3224, 3262, 3163, 3287, 3253, 3244, 3196, 3375, 3369, 3335, 3405, 3381, 2690, 3093, 3270, 3312, 3453, 2340, 2333, 2397, 2612, 3431, 3143, 3233, 3258, 3366, 3429, 2670, 3374, 2396, 3024, 2771, 2844, 2724, 3050, 3392, 3135, 2898, 2562, 3034, 3071, 3235, 3169, 3303, 3398, 3492, 2431, 2524, 2716, 3273, 2891, 3121, 3363, 3399, 3510, 3128, 3018, 3274, 3227, 3497, 2300, 2365, 2353, 2523, 2817, 3382, 2703, 3114, 3108, 3202, 3200, 3269, 3277, 3401]. Scope [3080, 3005, 2437, 2343, 2714, 2808, 3354, 3230, 3210, 2473, 2430, 2691, 3345, 3415, 3446, 3445, 4186, 2292, 2371, 2435, 2623, 2539, 3100, 2367, 2351, 2379, 2461, 2474, 2354, 2476, 2438, 2616, 2644, 2493, 2525, 2889, 2902, 2863, 3263, 3283, 3187, 3186, 3182, 3340, 3423, 3457, 3454, 3487, 3460, 2299, 2682, 2505, 2526, 3026, 3003, 3424, 3189, 3511, 2355, 2680, 3015, 2692, 2950, 2797, 3162, 3447, 2688, 2916, 2881, 2929, 2748, 2738, 2923, 2933, 2864, 3078, 2983, 3208, 3206, 2288, 2466, 2594, 2497, 2899, 2773, 3428, 2309, 2465, 2464, 2671, 2584, 2753, 3140, 2973]. Scope [3058, 3246, 3367, 3027, 3030, 2395, 2455, 2427, 2517, 2647, 2543, 2518, 2649, 2653, 2685, 2499, 2512, 2743, 2739, 2803, 2940, 2884, 2946, 3084, 3145, 3031, 2998, 3259, 3232, 3307, 3351, 3449, 3486, 3432, 4200, 4205, 4224, 3049, 3231, 3234, 3243, 3304, 3299, 3065, 3271, 2587, 2622, 3430, 3125, 3216, 3218, 3305, 3310, 2694, 2796, 3006, 3053, 2839, 2908, 2681, 2689, 2713, 2744, 2873, 2875, 2729, 2824, 2914, 2821, 2855, 2935, 2818, 2890, 2903, 2869, 2972, 3155, 3000, 3045, 3090, 2971, 2985, 3019, 3070, 3248, 3238, 3166, 3282, 3316, 3500, 4282, 2652, 3276, 3272, 3103, 2460, 2664, 2877]. Scope [3240, 3311, 555, 2295, 2467, 2434, 2449, 2651, 2778, 3106, 2332, 2472, 2432, 2487, 2784, 2712, 3127, 3033, 3012, 3086, 4197, 3209, 3213, 3214, 2415, 2595, 2815, 3094, 3332, 4201, 4223, 4218, 2780, 2624, 2835, 2874, 2924, 2837, 2834, 2760, 2876, 2750, 2880, 3062, 3066, 3002, 3119, 2997, 3063, 3172, 3409, 4189, 3156, 2393, 2974, 3413, 3490, 2436, 2391, 2380, 2580, 2657, 2581, 2677, 2521, 2676, 2943, 2825, 2909, 2854, 2728, 2782, 2846, 3028, 3109, 3183, 3256, 3188, 3347, 3314, 2441, 2440, 3132, 2611, 2500, 2520, 2607, 2556, 2513, 2705]. Scope [2860, 2722, 3096, 2976, 3402, 3239, 2389, 2410, 2546, 2535, 3198, 2443, 2418, 2416, 2419, 2695, 2510, 2511, 2833, 2711, 2745, 3064, 3414, 3380, 3489, 4208, 4188, 4219, 4276, 2759, 3138, 3014, 3346, 2725, 2717, 2557, 4199, 2322, 2310, 2314, 2390, 2496, 2608, 2620, 2701, 2896, 2814, 2741, 2895, 3197, 3225, 3228, 3461, 4265, 4266, 2315, 2316, 2317, 2377, 2376, 2331, 2378, 2568, 2592, 2925, 2719, 2720, 2804, 2879, 2888, 2805, 3148, 3059, 3146, 3141, 3255, 3264, 3265, 3319, 3362, 3339, 3404, 3391, 3410, 3440, 3482, 3441, 3481, 3434, 3483, 3505, 3439, 3504, 3433, 3503, 4220, 2532, 2582, 2531, 2589]. Scope [2588, 2938, 2910, 2913, 3048, 3151, 3149, 3285, 3281, 2718, 2579, 2723, 2978, 4216, 2442, 2468, 2501, 3286, 2987, 2349, 2422, 2540, 2530, 2793, 2951, 2952, 2859, 2758, 2947, 3236, 2982, 2409, 2298, 2401, 2424, 2398, 2370, 2400, 2368, 2399, 2369, 2423, 2619, 3217, 3191, 3185, 3278, 3508, 2503, 2417, 2645, 2558, 3136, 2799, 2789, 2488, 3147, 2321, 2618, 2807, 2826, 2905, 2994, 2968, 2654, 2444, 2683, 2492, 2831, 2822, 2953, 2894, 3087, 2986, 3032, 2848, 2454, 2669, 2536, 2599, 2291, 2665, 2666, 3178, 2453, 2388, 2475, 2373, 2426, 2413, 2470, 2613, 2646, 2590, 2507, 2591, 2559, 2656]. Scope [2663, 2484, 2885, 2802, 2934, 2956, 2721, 2900, 2871, 2957, 2710, 2736, 2827, 2832, 2928, 3082, 2970, 3134, 3068, 3001, 3069, 3153, 3020, 3110, 3085, 3129, 2969, 3041, 3021, 3042, 3215, 3242, 3170, 3260, 3284, 3411, 3338, 3306, 3372, 3379, 3443, 3484, 3475, 3474, 3495, 4207, 4215, 2534, 2598, 2787, 3397, 2456, 3039, 3067, 3267, 3300, 3348, 4185, 2319, 2344, 2752, 2345, 2469, 2922, 2917, 2964, 3378, 3355, 2547, 2915, 2920, 2882, 2967, 2919, 2852, 2586, 2358, 2420, 2650, 2610, 2768, 3164, 2769, 4267, 2704, 2990, 2450]. Scott [3607, 4301, 2826, 4258, 3655, 3692]. Scotti [3031]. Scottish [275]. scratch [2162, 1488]. Screen [2668]. Sea [3372, 493, 776, 1035, 3372]. sea-floor [493]. Sea-level [3372, 3372]. Seager [2654]. Sean [3273, 2620]. Search [3195, 2044, 3423, 2186, 2970, 4410, 2225, 1189, 690, 3227, 1812, 3811, 4065]. Searches [1847, 3547, 1380]. Searching [3080]. Searle [3555]. seas [808, 3284]. season [1085]. Seattle [4256]. seaweed [1561]. Sebastian [3343]. Sébastien [3671]. Second [3201, 3875, 3274, 3234, 3065, 3863, 3255, 4343, 3458, 4248, 1721, 4353, 4201, 4244, 13, 61, 133, 3645, 1916, 3475, 3581, 4401, 2614]. second-order [13, 61, 2614]. Secret [4357, 4358, 4361, 2329, 4027, 2772]. Secrets [2532, 2070, 2813, 1826]. section [240, 79]. sector [4233]. secured [3504]. security [2228]. Sedighe [4341]. Sedlmayr [3309]. see [3624]. Seed [4050]. seeding [1117]. Seeing [1806, 3472, 3968, 2506]. seeking [1936]. seemingly [3525]. seen [1801, 362]. Seestrom [3590]. Segal [3753]. Segrè [1265].

Seidelmann [2727, 3069, 3685]. Seideman [2628]. Seiringer [2601]. Seismic [3614]. Seismology [3007, 4080, 3717, 3722]. Seitaridou [3186]. Sekino [3630]. Selcuk [3995]. Selecta [2368]. Selected [848, 3885, 3747, 3551]. selection [5, 2215]. selenium [962]. Self [1027, 2461, 2998, 2014, 979, 3070]. Self-Assembled [2461, 2014]. self-diffusion [979]. Self-Organised [3070]. Self-Organized [2461, 2998]. Self-trapping [1027]. Selling [4057]. Selsis [2387, 2476]. Semi [2686, 2807]. Semi-Classical [2686, 2807]. Semiclassical [1955, 3933, 4141]. Semiconducting [558, 559]. Semiconductor [3013, 2932, 964, 3530, 201, 1503, 4220, 2014, 2421, 4162, 2538, 3030, 2946, 2805]. Semiconductors [2942, 2792, 3083, 1363, 571, 2573, 1326, 1308, 1285, 2096, 518, 1030, 1211, 901, 542, 135, 1748, 1311, 679, 543, 1289, 1588, 532, 2878, 3257, 3764, 3451]. Semimetals [679]. seminar [69]. Sen [3540, 3539, 2545]. Sénéor [3736]. senior [2293, 2525, 2889, 3186, 2295, 2442, 2683]. sense [3848, 3949, 3731, 3645]. Sensing [2288, 896, 1572, 1751, 2584, 3231]. Sensitivity [2354, 4280]. sensor [4251]. Sensors [3407, 3084, 2717, 1191, 1820, 1896, 3862, 3951, 3013, 3745]. sensorship [1058]. separation [172, 1496, 863, 2209]. September [193, 156, 89, 278, 277, 76, 242]. sequence [363]. Serbo [4187]. Sergei [4223, 2879, 2832]. Sergey [2563, 3040, 2746]. Sergio [3646, 3184, 3511, 3565]. Series [2926, 3116, 2806, 4190, 2794, 3013, 2288, 4321, 2455, 2884, 3145, 4205, 2295, 3132, 2416, 2378, 2804, 2888, 3147, 409, 4198, 2344, 2253]. Serjeant [3436, 2591]. Serpengüzel [2787]. service [397, 2211, 792]. Serving [3253]. session [177]. Sessler [3349]. Set [3353, 3246, 3119, 2086, 3509, 2795, 3626, 4255]. Seth [4275, 2585]. Sethna [4175]. sets [2153]. Seven [3244, 1833, 4191]. seventeenth [23]. seventieth [694]. Sevia [3504]. Sevrin [3133]. Sh [2466]. Sha [3402]. shadow [1273, 4023]. shake [1101]. Shaking [1571]. Shalaev [2791]. shallow [3867]. Shamonina [2458]. Shane [2584]. Shanmuganathan [3657, 3658]. Shape [2807, 4274, 457, 1627]. shapes [4340]. shaping [3418, 3357, 483]. Shapira [4255]. Shapiro [3065, 2649]. sharing [2862]. Sharon [2537]. Shattered [3758]. Shatterer [1092]. Shattuck [3607]. shear [2406]. Shedding [3396]. Shell [2689, 2514]. Shellard [2599]. Shells [3309]. Shelton [2607, 1146]. Shen [4250]. Shepard [3737]. Sherwood [4097]. Shifman [4030, 2875]. shift [3078]. Shifting [2807]. Shigeji [3501]. Shigemitsu [4284]. Shigeru [3103]. Shik [2930]. Shima [3544]. Shimabukuro [2486]. Shimada [2925]. Shimakawa [3764]. Shimon [3163]. Shimony [3386]. shine [3228]. Shinil [4353]. shining [2551]. Shinn [3433]. Shiung [3299]. Shlomo [3110]. Shock [234, 223, 2974]. Shoenberg [1057]. Sholl [4273]. Shong [2773]. Shoogo [4284, 3630]. Shoot [3870]. Shore [2847, 2366]. Short [947, 957, 974, 986, 996, 1004, 1017, 1032, 1061, 1087, 1320, 1330, 1345, 1353, 1361, 1368, 1376, 1383, 1397, 1447, 1492, 1499, 1612, 1860, 1868, 1877, 1890, 1909, 1930, 1940, 1962, 1986, 2091, 2105, 2111, 2118, 2139, 2144, 2790, 2323, 2477, 3120, 3018, 3263, 2434, 2426, 2650, 2990, 3757, 3622, 3651, 3790, 3805, 4016, 468, 464, 4073, 3599, 3856, 4095, 4265, 2413, 161, 174, 182, 190, 200]. Shortcuts [3164]. Shostak [4275]. Should [3009]. Shouyun [3441]. Shuang [3542]. Shubina [3530]. Shubnikov [3984]. Shultis [2860]. Shuvo [4350]. Si [2892, 561, 797]. Sibtain [2876]. Siddhartha [3540]. Side [2589, 3670]. Sidney [3911, 2619]. Signal [2766, 2419, 2975]. signals [4304, 2766]. Signatures [4407, 4150, 3869]. significance [951, 1652]. Signs [2049]. Silicon [1007, 1177, 287, 878, 1257, 3481, 3141]. Silk [4316]. Silva [2771, 2608]. Silvan [3945]. silver [991, 432]. Silverman [3414]. Similarities [2345]. similitude [4222]. Simm [3614]. Simmons [4183]. Simon [848, 3322, 3553, 3457, 3259]. Simone [4310]. Simons [3948]. Simple [3302, 3807, 410, 4210, 1657, 339, 355, 4274, 3563, 385, 514, 1210, 3142, 3186]. simpler [3222]. Simplicity [3380]. Simulating [3022, 1787]. Simulation [3023, 1091, 2061, 3842, 2660, 4217]. Simulations [3127, 3147, 3518, 4334, 2927, 4283, 3476]. simulator [501]. Sindhunil [4309]. Sine [2206]. Sine’-phony [2206]. Singapore [4106, 4010, 4324, 4062, 4225, 4268, 4254, 4056, 4055, 4231, 4293, 4028, 4123, 4069, 4113, 4030, 4163, 3999, 4029, 4044, 4024, 4108, 4068, 4119, 4098, 4196, 3998, 3423, 4112, 4114, 4099, 4121, 4083, 4080, 4351, 3996, 4255, 3337, 4096, 4220, 4086, 4004, 4033, 3997, 4067, 4172, 2846]. Singh [4240, 2348]. Single [3250, 3364, 2132, 2108, 121, 2076, 4212, 2381, 212, 2814]. Single-photon [2132, 2381]. Singular [3682]. singularities [3477]. singularity [3859]. Sir [2335, 1757, 2765, 1205, 648, 796, 331, 1136]. Sirenko [3747]. Sirius [2438]. Sirohi [2557]. Siscoe [2993, 2721]. Sitter [3560]. Sivia [2633]. Sixth [3286, 4340]. size [299, 1237, 3982]. Sizes [3283]. skeletons [2041]. Skidding [6]. skills [4057]. Skinner [3260]. skirmish [2167]. Skobel’tsyn [555]. Skomski [3142]. Skorobogatiy [2341]. Sky [2439, 3472, 2022, 4392, 4393, 4285, 3643, 2215, 3615, 3135, 4345, 3427, 3668]. skylight [3251]. Skyrmions [3992, 3872]. Slavik [3119]. Slawinski [4293, 4080, 2776, 2702]. slip [1737]. slit [212]. Sloan [2022]. Slough [334]. Slow [659, 2094, 3835, 2155, 2342]. slow-neutron [3835]. Sluckin [2668]. Small [3405, 3283, 3404, 1358, 1833, 372, 546, 3534]. small-scale [372]. Smart [2006]. smell [2602]. Smilga [4163]. Smirnov [3610]. Smith [1092, 4040, 3597, 3016, 3283, 2778, 2436, 3615]. smooth [1215]. Snow [1096, 2256, 4340, 1570, 2255]. snowflake [1622, 1782]. Snyder [2794]. Soap [1199, 2260, 2668, 2259]. Social [3031, 3306, 1952, 4093]. Society [158, 554, 38, 51, 36, 76, 37, 39, 97, 143, 194, 64, 811, 308, 605, 442, 261, 214, 276, 350, 2506]. sociological [2191]. sociology [3664]. Soffel [2727]. Soft [2616, 314, 1960, 3505, 3675]. softback [4225, 4123]. Softcover [4351, 4176]. software [3476, 2670]. soil [529, 3786]. Sokolov [2825]. Solar [224, 2737, 2993, 1583, 2891, 2295, 3028, 3136, 2832, 3443, 891, 4252, 3886, 4037, 218, 1911, 4054, 2223, 1572, 2404, 4412, 2898, 2839, 2440]. Solid [453, 2537, 2873, 2987, 1055, 2488, 4333, 3659, 1742, 3960, 466, 731, 1528, 900, 1602, 1670, 1243, 2040, 4337, 4096, 39, 3213, 4009, 2865]. Solid-state [453, 3659, 1742, 1528, 39, 2865]. solidification [723]. Solids [375, 2897, 4365, 3424, 2712, 142, 1615, 1146, 826, 4247, 92, 114, 756, 1027, 602, 1129, 711, 1010, 47, 386, 401, 847, 857, 552, 970, 1298, 809, 827, 2667, 440, 2522, 2966, 2749, 2703, 3108, 2466, 4178, 3616, 4006, 3632]. Soliton [3504]. Solitons [2999, 2544]. Solomon [4048]. Solution [3212, 2653, 3193, 2186, 3401]. Solutions [2781, 2999, 2457, 2706, 3233, 4210, 1435, 3895, 2235, 3633, 3052, 2786, 2987, 2705, 3830, 3952]. solvable [3652]. Solvation [3118]. Solvay [3133]. Solve [2518]. Solved [3114, 4187, 2781]. Solving [4368, 2649, 3960, 4322, 4201, 4258, 4003]. Solymar [2458, 3647]. Some [771, 861, 162, 630, 838, 236, 295, 862, 580, 1602, 963, 135, 1421, 447, 214, 276, 350, 12, 228, 2661, 2170, 438, 1529, 537, 3735, 738, 1082, 3929]. Something [1050, 4194]. somewhere [2133]. Sommerfeld [2585]. sonar [2307]. Sondergaard [3483]. Song [3145]. Sonic [856]. Sonnenschein [2622]. Sonoluminescence [1848]. Sons [4334]. Soo [2932]. Soo-Jin [2932]. Soukoulis [2300]. Soul [3253]. Souma [2657]. Sound [2216, 1792, 2550, 4244, 1527, 3765]. Sound-producing [1792]. sounding [729]. sounds [399]. Source [2451, 2789, 658, 1885, 2449]. sources [3250, 752, 2381, 436, 924, 978, 2132, 3387, 3581, 3792, 4058, 4215, 1414, 7]. South [315]. Soviet [3938, 2008]. Space [1098, 1164, 2648, 3072, 2545, 2981, 3261, 3406, 3137, 1745, 2984, 1589, 1874, 1714, 2993, 3821, 10, 281, 305, 312, 329, 337, 344, 2889, 4264, 3276, 3106, 3471, 2718, 1482, 2370, 2492, 439, 2590, 2721, 1974, 2195, 3644, 671, 2382, 1002, 2063, 445, 1697, 1760, 660, 2143, 3794, 1228, 654, 4121, 1606, 3558, 3693, 3467, 4253, 3595, 1304, 85, 1751, 228, 1278, 1282, 2453, 3850, 1659, 3746, 2652, 2780, 3355, 3164, 4137, 4346, 3296, 3627, 4186, 3182, 2449]. Space-charge [439]. space-science-and-public-engagement [4186]. space-time [4346]. space-time-matter [4121]. Spacecraft [2453, 435, 1021, 1006]. Spacelab [843]. Spacetime [2524, 4389, 4390, 4391, 4191, 2153, 4036, 3437, 3153, 4141]. spacious [1084]. Spain [968, 1507]. Spallation [924]. Spark [230]. Sparse [3449, 4406]. spatial [1776]. Special [3461, 82, 3560, 457, 2155]. specialised [3183]. Specialist [4110, 3310, 555, 869, 4120, 4078, 4060, 4106, 2686, 2883, 3157, 3395, 3333, 3112, 2433, 4225, 2756, 3098, 2296, 2357, 4241, 2555, 3022, 3506, 2843, 2593, 3203, 3298, 2335, 2334, 3313, 2795, 3116, 2545, 2565, 3091, 3089, 3010, 3229, 3261, 3171, 4154, 1241, 1247, 2754, 2757, 3349, 3371, 3137, 3038, 3095, 3035, 3247, 3442, 2984, 2798, 2838, 2785, 3194, 3179, 3353, 4272, 2735, 2872, 2755, 3502, 4155, 4248, 2791, 1229, 2841, 3036, 2806, 4072, 2429, 2660, 2887, 3205, 4068, 3184, 3081, 3074, 2996, 3190, 3309, 3394, 3479, 3477, 2313, 2330, 3131, 3093, 2771, 3050, 3392, 2898, 3034]. specialist [3398, 3492, 2431, 2524, 3273, 3399, 3128, 3080, 3005, 2437, 2808, 3230, 3210, 3445, 4186, 2435, 3340, 3026, 3003, 3424, 2680, 3078, 2983, 2288, 2466, 2973, 3058, 3246, 3145, 3232, 3307, 3449, 4200, 4205, 3065, 3430, 2839, 2652, 2449, 2651, 2784, 3086, 3213, 2415, 3094, 3332, 2624, 2837, 2834, 2760, 2876, 2750, 2880, 3119, 3172, 3409, 2521, 2846, 3256, 3347, 2441, 2440, 3402, 2546, 2443, 2695, 4255, 4219, 4301, 4276, 4350, 3346, 2725, 2322, 2310, 2314, 2390, 2701, 2896, 2741, 3228, 3461, 2315, 2316, 2376, 2378, 2879, 3148, 3404, 3391, 3410, 3440, 3482, 3434, 3505, 3439, 3504]. specialist [3433, 3503, 2349, 2793, 2400, 3217, 2799, 2822, 3087, 3032, 2665, 2666, 2832, 3484, 4058, 4067, 4215, 2598, 3397, 3039, 3067, 4168, 2882, 2769, 4062, 4069, 4113, 4061, 2057, 4347, 4348, 4346, 4165, 4336, 4310, 4172]. Specialist/postgraduate [555]. specialists [2921, 3318, 3333, 4211, 4286, 2989, 2851, 2843, 2828, 2955, 2912, 3056, 3349, 3137, 3095, 3035, 4307, 3150, 3104, 2988, 3444, 2907, 3152, 4155, 2930, 2684, 3496, 2542, 2954, 3407, 2737, 3113, 3007, 3057, 2993, 3044, 3016, 3268, 3266, 3219, 3364, 3422, 2387, 2519, 3262, 3163, 2340, 3121, 3128, 3018, 3210, 3415, 4186, 3283, 3423, 2505, 2688, 2916, 2748, 2923, 2649, 3432, 4200, 4205, 3234, 3218, 2908, 2778, 3066, 2943, 2909, 2833, 2711, 2745, 4311, 3014, 2717, 3225, 3461, 2925, 2582, 2910, 2913, 2468, 2501, 2789, 2826, 2986, 3178, 3134, 2964, 2919, 2650, 2610, 2769, 4254, 4355, 4303]. specialists [4338, 4342, 4340, 4320]. specialists/Scope [2501]. specific [1049, 935]. specifying [2463]. specimen [4114]. Speckle [1999, 1023]. spectra [314, 657, 135, 555, 2788, 3631]. Spectral [2561, 2468, 1442, 3686, 753, 3732]. Spectrographs [2504]. spectrology [495]. spectrometer [1076]. spectrometers [1020]. spectroscope [116]. Spectroscopic [2007, 1428, 1788, 204, 210]. Spectroscopies [3408]. Spectroscopy [3395, 2788, 2386, 2865, 4365, 2975, 2779, 2878, 1162, 658, 2678, 3424, 2295, 2750, 692, 1204, 3506, 1217, 1225, 103, 3646, 1161, 3223, 916, 1934, 760, 1762, 2963, 1349, 1881, 216, 1056, 693, 4315, 917, 386, 401, 1670, 4112, 1489, 447, 705, 4245, 1197, 1826, 698, 699, 2326, 915, 2788, 2555, 2696, 2541, 2680, 4112, 4169, 44, 3773, 3392, 2547, 3933]. spectrum [1090, 1488, 3615]. Speculations [2658, 2343, 3502]. Speed [2538, 652, 1012, 1176]. speeds [517]. spheres [4206]. Spin [1607, 1862, 912, 913, 3373, 1106, 3269, 3209, 3214, 3539, 2184, 4278, 1998, 4119, 4279, 839, 653, 1494, 1076, 3937, 3184, 3827]. Spin-down [1607]. spin-echo [1076]. spin-orbit [4119]. Spin-polarized [1862]. Spinning [1436, 3191]. Spinor [4228]. Spins [411, 2076, 4029]. Spintronics [3030]. spiral [1836]. spiritual [18]. Spontaneous [1922, 3822, 3820, 2102]. Spooky [1839]. Sports [2851, 2582]. Sprawls [3119]. spread [1347, 4219]. spreading [4300]. Springer [4149, 4110, 4254, 4018, 4154, 4191, 4118, 4155, 2806, 4148, 2794, 4031, 4047, 3993, 3957, 4091, 4115, 4159, 4160, 2288, 4321, 4218, 4208, 2378, 2804, 4220, 4171, 4059, 4145, 2344]. Springer-Praxis [4059]. Springford [1057]. sputtered [1162]. Spyridon [3581]. square [1334, 1190]. squares [2549]. Squeezing [1387]. SQUID [1038]. Squires [2955, 3813]. Sreenivasan [3143]. Sridevi [3351]. Srivastava [4259]. SSC [1600]. Stability [3241, 2601, 992, 1309, 1738, 3626]. stabilized [1413]. stable [1380]. Stacey [2646]. Stachel [3473]. stage [3494, 3450, 3452, 4081, 4259, 4269, 4295, 3258, 3429]. Stallard [4042]. Stamps [4009]. Standard [2744, 3316, 2595, 2322, 2631, 4177, 3769, 3783, 3430, 616, 3731, 3847, 4366, 1804, 2554]. Standards [874, 468, 687, 2038, 2087, 1401, 2327, 1554]. Stanescu [3785]. Stanisislav [3059]. Star [3400, 2505, 610, 3348, 3602, 3860, 3993, 3416, 1530, 4091, 2780, 3228]. star-forming [3860]. star-shine [3228]. Stargates [4411]. Stargazing [2420]. Starikov [2807]. starlight [4013]. starring [2236]. Starry [3472]. Stars [3456, 2748, 3106, 2970, 3039, 3267, 643, 609, 4213, 2122, 3477, 3891, 3970, 902, 1882, 984, 4114, 4398, 1512, 2501, 2433, 3152, 2923, 3234, 2501, 3001, 4327, 3553, 3364, 3456, 3018]. Starships [4411]. Started [2420]. Starzewski [4061, 2410]. State [1888, 2537, 2873, 2987, 3520, 381, 1049, 4009, 871, 1935, 4333, 3659, 1777, 58, 453, 678, 1742, 564, 3960, 466, 4051, 1528, 900, 1673, 1602, 820, 4337, 4096, 736, 39, 2865]. state-of-the-art [3520]. States [3460, 4402, 722, 1765, 2217, 3386, 1879, 4138, 1776, 2290, 2369]. Station [3276, 89, 168, 277, 264, 278, 334]. Statistical [2502, 1409, 1424, 3376, 3038, 2740, 2823, 2406, 608, 2351, 2940, 2587, 2657, 3183, 3132, 2416, 2510, 2368, 2619, 2454, 2665, 3178, 2475, 1483, 1437, 4217, 3934, 3715, 3663, 3507, 3591, 3820, 3819, 2208, 3584, 3864, 3831, 4188, 3760, 4095, 4265, 907, 4185, 2560, 4184, 2872, 2766, 4052, 2625, 3557, 2497, 4282, 3214, 2676, 3188, 4175]. statisticians [2354]. statistico [3798]. statistico-mechanical [3798]. Statistics [2465, 2888, 4148, 2259, 1199, 2260, 3577, 2888]. status [1475, 3174]. steady [240, 736]. steady-state [736]. Steane [3037, 3150, 3679, 3756]. Steckel [4273]. Steel [4199]. Steering [2061]. Stefan [3895, 4269, 3321, 4017, 3246, 3248, 2627, 3379]. Stefano [3659, 3075]. Stefanucci [3090]. Steiner [2443]. Steinmetz [2880]. Stellar [2621, 4378, 3007, 3422, 2898, 2440, 611, 755, 1801, 3970, 59, 436, 3928, 3897, 3990, 2468, 2626, 2907, 3364, 3422, 2620, 3586]. Stenger [2753, 3521]. Stenholm [2770]. step [4103, 2960]. Stepanov [3007]. Stephan [2757, 4081, 2805, 3265]. Stéphane [2929]. Stephen [3618, 3171, 3035, 3192, 3436, 4346, 3262, 3345, 2455, 2511, 2711, 2758, 2599, 2990, 4117]. Steps [2318, 2825]. Stern [3673]. Steur [4099]. Steve [2457, 4136]. Steven [4057, 3561, 3081, 4150]. Steves [2839]. Steward [2854]. Sticky [1010]. Stigmatic [4170]. Still [1936]. stimulated [832]. stimulation [3630]. Stishov [4055, 4223]. Stochastic [3123, 3449, 2821, 3132, 2419, 3014, 2361, 4406, 3765, 4364, 2699, 3514, 4141, 3489, 1844, 2285, 2931, 3486, 2782, 3256, 2419, 2745, 4310]. stochastics [3720]. Stockar [4269]. Stockholm [3088, 834]. Stocks [2419]. Stokes [2765, 3573, 2765]. Stone [3097, 3703, 3129]. Stoof [2525]. storage [3767, 680, 1896, 4049, 1271, 3174, 4272]. store.ioppublishing.org [4201]. Stories [2491]. Storm [156, 3066, 2663]. Storms [2721, 3787]. Stormy [3284]. Story [3318, 3403, 2867, 2506, 3088, 2694, 3106, 2645, 3609, 3844, 4155, 1651, 3477, 4037, 1681, 3777, 1916, 1917, 3662, 3981, 3268, 4256]. Stott [3099]. Stoyan [4345]. Stradivari [2532]. straight [2146, 240]. straightforward [2176]. strain [1012]. Strand [3119]. Strange [267, 3380, 722, 3662]. Strangeway [3627]. strategic [4234]. Strategies [2354, 2387]. Straten [3598]. stratosphere [3468, 3468]. streamer [2137].

Strength [409, 873, 24, 434]. Strengthening [3154]. stretching [2077, 805]. strike [1844]. String [2786, 2856, 3004, 3523, 3583, 2163, 3783, 4181, 4417, 2893, 3873, 4029, 3219, 4131, 4233]. Strings [1830, 2141]. Stripes [3214]. Strocchi [3125]. Strominger [4085]. Strong [272, 1372, 2221, 895, 3176, 680, 4050, 3659]. strong-field [2221]. Stroscio [2377]. Structural [992, 4217, 950, 3432, 4215]. Structure [1091, 2966, 2545, 355, 2621, 2539, 2680, 2487, 3265, 3048, 911, 4149, 19, 1204, 3, 600, 4278, 666, 1758, 874, 339, 1187, 1163, 1738, 670, 4081, 1665, 4085, 4346, 2514, 81, 1706, 1428, 3832, 4159, 4151, 901, 762, 3665, 4122, 651, 3903, 2053, 3707, 809, 74, 84, 2366]. Structures [2672, 2847, 3930, 2675, 3046, 2517, 3310, 2869, 2389, 782, 2195, 1547, 3872, 546, 629, 1727, 800, 3496, 295, 1712, 725, 1831, 1635, 1278, 1282, 2461, 2797]. Struggle [3253]. Stuart [4177, 2826, 2365]. Student [3212, 2944, 3258, 4385, 3411, 3868, 3946, 4273, 2350, 3692, 3955, 3793, 3936, 4105, 4292, 3189, 3511, 4161, 4079, 4169, 4170, 4234, 2422, 2503, 4296, 4079]. student/research [2503]. Students [3114, 4078, 3025, 2297, 2290, 2385, 2458, 2486, 2659, 2485, 2425, 2868, 2411, 2749, 2842, 16, 2352, 2341, 2342, 2421, 2563, 2294, 2366, 2567, 3438, 2457, 2597, 2489, 2566, 2609, 2528, 2816, 2706, 2918, 3013, 4187, 4283, 2402, 2396, 2473, 2351, 2461, 2354, 2438, 2902, 3187, 3186, 3182, 3454, 3206, 4114, 2517, 2543, 2518, 2653, 2499, 2512, 2743, 2739, 2803, 2940, 2884, 2946, 2998, 4326, 2295, 2467, 2825, 2418, 2416, 2419, 2510, 2511, 2833, 2711, 2745, 3064, 2895, 4265, 2579, 2399, 2683, 2831, 2894, 2454, 2453, 2388, 2475, 2613, 2646, 2591, 2663, 2885, 2956, 2900, 2710, 2969, 2345]. students [2358, 2704, 4117, 2451]. students/researchers [2831, 2894]. Studies [2333, 2354, 1149, 423, 2108, 791, 287, 1428, 1788, 2062, 768, 352, 879, 2512]. Study [40, 4417, 177, 782, 4317, 1137, 388, 675, 687, 587, 629, 744, 657, 30, 926, 74, 2358]. studying [2567, 385, 833]. Stuff [3403, 3182]. Styer [4268, 2768]. Sub [3550, 2500, 2195, 1233]. Sub-cycle [3550]. sub-millimetre [1233]. sub-Planck [2195]. Sub-Surface [2500]. Subatomic [3171, 2314]. Subhas [4251]. subject [584]. subjects [4211]. submicron [1737]. substitutions [3749]. substrate [1529]. substructure [2257, 1097, 2258]. Subtle [2661, 1165]. success [2384]. Sudipta [4169]. Suetens [2464]. Sugano [2417]. suggested [3727]. Suits [4286, 4307, 4352]. Sultana [2788]. sum [4311]. Suman [2585]. Sumeet [2943]. Sumigawa [2925]. Summer [3772, 40, 2760, 2846, 2322, 3836]. Summons [2387]. Sun [250, 2843, 1523, 1801, 2302, 3003, 2380, 2086, 3693, 2501, 1826, 2559, 3553, 4115, 4305, 3400, 3332]. sunsets [1561]. sunspots [3777]. sunstone [3251]. Super [1938, 2603, 3948, 2190, 2360, 3537, 2675, 3795]. Super-Earths [2603]. super-LHC [2360]. super-resolution [3537, 3795]. supercapacitors [4250]. supercollider [1600]. Superconducting [3371, 618, 3024, 4152, 2169, 1600, 1073]. Superconductivity [523, 1609, 1166, 825, 4262, 1009, 1715, 3728, 1769, 477, 1543, 1128, 1494, 376, 492, 3819, 2858, 2413]. superconductivity-ten [1769]. superconductors [1139, 873, 4226, 371, 1993, 3277, 4094, 3474]. Supercontinuum [2725]. Superdeformed [1348]. Superfluid [1172, 4402, 1532, 2573, 3635, 730, 1107, 1942, 2277, 1737, 3460]. Superfluidity [2330, 2185]. Superfluids [3187, 1942, 2276]. supergiants [3865]. supergravity [1335]. superheating [805]. Superheavy [1324]. Superionic [918]. Superlattice [2785]. superlattices [2361]. Supermassive [3963, 2013, 4335]. supernatural [2099, 2081]. supernova [1355]. Supernovae [3075, 2929, 3039, 1415]. Superradiance [1065]. Supersolidity [2185]. Superstars [2929]. superstrings [1434, 1374, 2509]. superstructures [1537]. Supersymmetric [2953, 3511, 2947]. Supersymmetry [1335, 2926, 1932, 2044, 3783]. supplement [2301]. supplementary [3311]. support [4057, 675, 4382, 3438]. Supramolecular [2795]. Suprathreshold [2419]. Surface [1420, 2340, 448, 1277, 2500, 2919, 728, 2228, 1143, 535, 1295, 768, 1057, 1517, 1028, 3890, 531, 1023, 33, 161, 174, 182, 190, 200, 3962, 4323]. Surface-functionalized [4323]. surfaces [1119, 1163, 717, 625, 1670, 1055, 915, 1395]. Surfing [3838]. surgery [2233]. Surges [3372]. Surprising [3400]. Surrow [3754]. Survey [869, 1169, 1242, 2294, 3577, 30, 3197, 119, 2022]. Susan [3590]. Sushkov [3492]. Suspension [2767]. Sussman [3264]. Sustainable [4049, 3892]. sustaining [2364]. Susumu [3409]. Sutton [4198]. Suzanne [2533]. Suzuki [3501]. Svelto [2736]. Sven [3320, 3119]. Sven-Erik [3119]. Svistunov [3460]. swampland [4181]. Swamy [2996, 2345]. Swapan [3543]. Swarup [2348]. Swendsen [3178]. Swimming [2337, 1249]. swing [3700]. switching [4213, 1594, 506]. Switzerland [4018, 4154, 4155, 4148, 3957, 4091, 4115, 4321, 4208, 4171, 4145]. Swope [2417]. symbolic [3996]. symbols [810]. symmetric [2138]. symmetrical [637]. Symmetries [2349, 3061, 4177, 1926, 4008, 354, 2244, 3358, 1569, 2301, 3801, 2243, 2535, 2758]. Symmetry [4365, 1938, 3354, 3424, 4226, 2540, 2793, 3971, 3758, 908, 4164, 2183, 1647, 2869, 2859, 3711, 3930, 3344, 3305, 3310, 2432, 3812, 2894]. symphony [3291]. Symposia [2737, 2467]. Symposium [2727, 3308, 834, 819]. Syms [3647]. Synchronisation [3638, 3820, 3466, 2285]. Synchronization [2060]. synchroton [2355]. synchrotron [884, 658, 437, 3467, 4215]. Synergetic [3095]. synoptic [2519]. Synthesis [4180, 1892, 700]. Synthetic [3418, 4119]. Syono [2472]. System [2737, 3443, 390, 4139, 387, 283, 4037, 218, 1849, 1572, 2219, 4412, 3165, 2832, 4173, 4054, 3136]. systematic [396]. système [560]. Systems [3047, 2476, 3340, 2869, 3090, 3311, 3033, 2760, 2745, 1140, 3362, 2536, 3129, 3766, 1622, 1317, 1469, 2136, 125, 2320, 3747, 3889, 2087, 999, 1008, 3496, 3920, 2325, 665, 3517, 2734, 731, 1474, 1871, 1807, 1911, 4280, 4019, 4249, 3735, 3833, 738, 2166, 3476, 2179, 1512, 4088, 1006, 3746, 4243, 3393, 3022, 2330, 3374, 3162, 2839, 2824, 2821, 3209, 2695, 2315, 3339, 2994, 2444, 2453, 4304, 3663, 3554, 3659, 3591, 3678, 4017, 4301, 4350, 3894, 3905, 3951]. Szabó [3247, 2786, 3247]. Szalma [3653]. Szwacka [2769]. Szwacki [2769].

’t [2397, 3379, 3757, 2346, 2893, 3301, 3699, 2385, 2485, 3203, 2334, 2350, 2751, 2981, 3373, 2617, 3626, 2394, 3054, 3999, 3043, 3705, 3577, 3627, 3689, 3754, 3797, 4037, 4042, 2313, 2243, 2353, 2691, 2476, 2644, 2525, 2773, 2685, 3789, 2935, 3209, 3613, 3402, 4129, 2720, 3813, 3812, 3236, 2900, 2957, 3085, 3067, 1580]. T.-Y [3999]. Ta-Pei [3071]. Tabakov [3119]. Tabar [2332]. table [3758]. Tables [3083, 3711]. tablets [1766]. Tadmor [2874]. Tagmatarchis [3051]. Tahir [4254]. tail [1921]. Tailored [4242]. Takahiro [2925]. Takashi [3184, 3103, 2925, 2789]. Takayuki [2925]. Take [361, 2246, 2245]. Takeuchi [2635]. Taking [239]. Tale [3244, 2163]. Tales [2548, 3739]. Tamar [2628]. Tamblyn [2708]. Tanaka [4015]. tandem [73]. Tang [3299]. Tang-Fong [3299]. tangled [2570]. Tanner [4006]. Tannoudji [2921]. Tanton [3689]. Tanury [3343]. targeted [3706]. targets [2307]. TASI [4417, 2322, 4028]. TASI-2006 [2322]. Tasker [4150]. Tatiana [3530]. Tatsu [2635]. Tatsuya [2789]. Taylor [4296, 4257, 4305, 4283, 4294, 3277, 4336, 4349, 4297, 4308, 4267, 2517, 2725]. teach [63]. Teacher [3178, 4120, 4078, 4060, 4106, 2686, 2883, 3395, 3318, 3211, 4225, 4275, 2851, 2843, 3107, 2593, 3298, 2565, 3089, 3056, 3237, 4123, 3349, 3137, 4113, 3038, 3095, 3035, 3247, 2823, 2785, 3353, 3408, 3488, 4272, 3092, 2621, 3168, 3139, 3448, 3036, 4072, 4178, 4190, 4198, 2693, 3407, 2887, 3184, 3317, 3435, 3270, 2612, 3233, 4165, 3398, 3492, 3128, 3018, 3274, 3210, 4186, 2881, 3428, 2685, 3084, 3145, 3259, 3232, 3307, 3351, 3449, 3486, 4200, 4205, 3231, 3243, 3430, 2874, 2760, 2876, 3066, 3119, 4054, 2722, 4071, 4208, 4188, 4311, 4199, 3197, 3225, 3461, 3148, 3141, 3391, 3151, 3285]. teacher [3286, 3087, 3032, 2665, 2666, 3306, 2650, 3164, 2769, 4328, 4307, 4084, 4322, 4321, 4171]. teachers [2921, 342, 2989, 2865, 2981, 2866, 2988, 2907, 2937, 2684, 2944, 14, 2933, 2899, 2649, 2940, 279, 2924, 2717, 2582, 2938, 2913, 2982, 79, 2967, 151]. Teaching [2146, 377, 69, 27, 167, 5, 307, 249, 177, 374, 237, 245, 302, 539, 290, 315, 4129, 25, 120, 231, 362]. technique [1076]. Techniques [3098, 2614, 2616, 3232, 2804, 1045, 1053, 630, 208, 3808, 4309, 4108, 1255, 1421, 2515, 1188, 378, 3890, 3743, 2326, 2557, 3286, 4165, 4007, 3395, 3321, 160]. technological [1694]. Technologies [2991, 3274, 2569, 2480, 4234, 3581, 1191, 3365]. Technology [3371, 221, 764, 1560, 2646, 3767, 1365, 1276, 1934, 4252, 3369, 825, 3531, 3571, 154, 1894, 3892, 1336, 1516, 2930, 2542, 2719, 3670, 4251, 3998, 3740]. technomimetics [3940]. Teitsworth [2711]. Teixeira [4270, 3507]. Telecommunication [3503]. Telecommunications [871]. Telepistemology [1951]. Teleportation [1840]. telescope [255, 1098, 1587, 1713, 1516, 1874, 1714, 3821, 4263, 1482]. telescopes [614, 1516, 2302, 3615, 3805]. Television [25, 69]. Teller [2428, 2600]. Temperature [2188, 2547, 1007, 817, 1769, 477, 1543, 208, 707, 2161, 1494, 1467, 94, 4102]. temperatures [1049, 229, 724, 3301, 2580]. Ten [3379, 1769, 3886, 4399, 3876, 3143]. Teng [2932]. Tennyson [2788, 4112]. tension [728, 424, 33]. tensor [3947, 3485, 4061]. Tensor-valued [4061]. Tensors [2944]. Terahertz [2930, 3060, 2386, 3117, 3564]. Terascale [2751]. Terawatt [1443]. tercentenary [64]. Terence [3622]. Terentjev [3505]. Teresa [2769]. terms [362]. ternary [1166]. Tero [3301]. terrain [2196]. Terrestrial [891, 1380, 4031, 690]. Terry [3433]. TESLA [3149]. test [917]. testimonies [4113]. testimony [3702]. Testing [1122, 4089, 1804, 2155, 2208]. Tests [1219, 2212, 4280, 1303, 1262, 4108]. tetraquarks [2159]. Tetsuro [2580]. Tetsuzo [2998]. text [2921, 3395, 3180, 3301, 4118, 2853, 2341, 3257, 3181, 3280, 3323, 3321, 3317, 3309, 3046, 3233, 2844, 3303, 3277, 3210, 2473, 2430, 3345, 2292, 2623, 2517, 3231, 3234, 3243, 2972, 3248, 3282, 3316, 4282, 3272, 3311, 3255, 3286, 3236, 3278, 3178, 3041, 3215, 3242, 3170, 3260, 4122, 3110, 3411]. text-book [2517, 2972]. Textbook [4253, 4120, 4078, 4060, 3122, 3393, 3368, 3025, 3211, 3112, 3199, 3350, 4317, 2502, 2788, 2829, 2346, 3061, 3073, 3480, 3498, 3201, 2357, 3509, 4217, 4284, 2936, 3115, 4232, 4268, 2290, 2911, 2927, 3097, 3107, 2385, 2386, 2458, 2828, 3365, 2375, 2966, 3491, 3485, 4055, 4210, 4273, 4271, 4304, 4296, 2425, 3072, 2955, 2865, 2981, 3089, 3037, 3237, 3383, 4182, 4177, 4184, 2320, 3095, 3035, 4333, 3123, 2939, 3023, 3507, 3396, 2838, 3493, 3488, 4183, 3076, 3144, 3212, 3193, 2931, 3092, 2907, 2937, 2352, 2766, 2878, 2992, 4325, 2533, 2655, 3036, 2901, 3117, 4072, 4052, 4178, 4190]. textbook [4198, 4252, 2965, 3344, 3494, 3501, 3452, 3451, 4081, 4204, 4203, 4162, 4222, 4269, 4281, 2944, 4242, 4250, 3297, 3081, 3043, 3190, 3377, 3341, 3315, 3435, 3438, 3479, 3437, 3436, 4105, 4339, 2330, 2348, 2820, 4318, 4319, 3258, 3366, 3429, 2396, 3235, 3398, 3492, 2300, 2365, 2817, 3382, 3114, 3200, 3080, 3005, 4157, 4292, 3446, 2292, 2889, 3186, 3026, 3424, 2355, 3015, 2950, 3447, 2881, 2929, 4112, 4114, 4161, 2497, 2309, 2465, 2464, 3140, 3058, 2649, 2653, 2512, 2940, 3259, 3351, 3216, 3053, 2689, 2713, 2873, 2875, 2729, 2824, 2914, 2821, 2855, 2935, 2818, 2890, 2903, 3155, 3000, 3045, 3090]. textbook [2971, 2985, 3019, 3238, 3166, 3500, 2460, 3311, 2472, 2432, 2487, 4197, 3214, 2415, 3332, 4337, 4354, 4244, 4249, 2874, 2997, 4169, 4170, 4189, 2436, 2677, 2676, 2728, 3028, 3188, 2440, 2500, 2520, 3096, 4306, 2443, 2418, 2510, 3064, 3476, 3489, 4071, 4188, 4301, 4310, 4070, 4066, 4077, 4166, 4167, 3014, 2310, 2314, 2390, 2701, 2896, 2741, 2895, 3228, 4265, 4332, 2317, 2377, 2568, 3148, 3141, 3264, 3265, 3391, 3483, 4063, 2938, 2910, 3151, 4086, 4171, 2579, 2978, 4216, 2442, 2349, 2951, 2952, 2859, 2947, 2982, 2401, 2370, 2399, 3191, 4290, 2488, 3147, 2905, 2994, 2444]. textbook [2831, 2953, 2894, 2291, 2475, 2373, 2646, 2507, 2656, 2484, 2885, 2900, 3068, 3153, 2969, 3021, 3411, 4207, 4176, 4175, 4185, 2319, 2344, 2469, 2917, 2920, 2967, 2358, 4083, 2773, 3975]. textbook/monograph [2831, 2894]. textbook/reference [2292]. textbook/supplementary [3311]. Texts [3142, 3026]. Th [2743, 2739]. Th.W [2649]. Thad [2783]. Their [2504, 961, 3003, 3033, 1055, 2826, 3379, 1265, 554, 790, 832, 1671, 618, 951, 2057, 2070, 4285, 845, 3204, 3861, 3891, 4090, 102, 1987, 4150, 4114, 2272, 56, 2843]. them [2549, 3517, 3964, 4322]. theme [2257, 1097, 2258]. Theodore [2868, 4347, 4340]. theorem [3639, 3726, 2160, 239, 2683, 3817, 4336]. theorems [2235]. Theoretical [2615, 2679, 4365, 4417, 3144, 2937, 2523, 3424, 2750, 2997, 2322, 2422, 2398, 2599, 3021, 752, 3747, 3784, 645, 2250, 3884, 2249, 954, 1529, 2487, 746, 4174, 3733, 3559, 3784, 3980, 3952]. Theories [3079, 2927, 3124, 2658, 3856, 3126, 2353, 3216, 2349, 3443, 3509, 553, 3502, 4138, 608, 4108, 1001, 1466, 2202, 1305, 1282, 4412, 3511]. theorists [2427, 2946, 1844]. Theory [3368, 2414, 2346, 2893, 2786, 3073, 2290, 2897, 2659, 2375, 2715, 2585, 2955, 3384, 2545, 2498, 2856, 3133, 2948, 3444, 2621, 2766, 2992, 2999, 2561, 2687, 2806, 3826, 3930, 2965, 3450, 3219, 4380, 2524, 3363, 3108, 3210, 3345, 2474, 2797, 3447, 2309, 2622, 3310, 2875, 2824, 2935, 3090, 3070, 3316, 2487, 2815, 2633, 2835, 3198, 3138, 3575, 2631, 2540, 2758, 2369, 2291, 2666, 2373, 3041, 3766, 1263, 2264, 4120, 82, 4324, 3523, 1342, 2170, 2078, 2138, 3699, 1317, 1424, 113, 3491, 3625, 3899, 3873, 3911, 4056, 4273, 3853, 2193, 3969, 4177, 1241, 1247, 3646, 1194, 1026]. theory [1250, 2382, 3585, 3900, 2163, 3518, 1122, 4417, 2174, 3758, 1850, 326, 1830, 1564, 4373, 2241, 1037, 3960, 4021, 4162, 2383, 1143, 3884, 3783, 2133, 1488, 1603, 2164, 3469, 1545, 4051, 1108, 2202, 4017, 984, 2466, 1138, 1494, 3768, 2455, 2499, 3486, 3819, 4089, 4181, 1566, 322, 330, 2242, 3665, 1943, 4131, 3703, 4122, 3879, 3880, 3959, 1467, 1601, 274, 2223, 297, 4306, 4255, 376, 4167, 3503, 3672, 3743, 1511, 1786, 970, 1431, 492, 259, 2454, 3792, 4233, 1483, 3685, 1988, 362, 4267, 2911, 3313, 2490, 2801, 3004, 3375, 3202, 3189, 2543, 2587, 3125, 3218, 2681, 2815]. Theory [2924, 3461, 2592, 3583, 3644, 4035, 3791, 3652, 3739, 4002, 4029, 4085, 2263, 4308, 2883, 3097, 4210, 4008, 2526, 4219, 2905, 3916, 3914, 3925, 3833, 3903]. Therapy [3442, 3706]. There [2733, 1506, 2411, 983, 2202, 2308]. Therese [3415]. Thermal [2955, 208, 2625, 3352, 537, 2511, 333, 1615, 834, 3901, 1049, 3961, 339, 355, 2198, 3966, 4282, 378, 830, 1289, 2866, 3064, 3760]. Thermally [832]. thermodynamic [1545, 2712]. Thermodynamics [2502, 3038, 2948, 2874, 4217, 3514, 3626, 4052, 665, 2605, 4097, 4074, 1840, 4310, 3549, 2740, 3178, 4269, 3526, 3766, 3663, 2708, 3679, 4026, 3563, 4249, 4189, 4188, 4094, 2969, 3495, 3983]. Thermoelectric [107, 1007, 638]. thermoelectricity [3580]. Thermography [582]. thermoluminescence [1442]. Thermonuclear [2928, 3642, 2129, 3578, 719, 2045]. thermostatistical [3221]. Thew [3680]. Thierry [2614]. Thijssen [3271]. Thin [3105, 484, 1886, 3028, 1511, 1666, 217, 2828]. Thin-Film [3028, 2828]. Things [2312, 2509, 4149, 662, 2034, 4276]. Thinking [1949, 2742]. Third [3200, 3063, 65, 1545, 3777]. Thirty [1981, 1131, 1079, 1026, 784]. Thirty-five [1131]. Thomas [3539, 2867, 3343, 4269, 2844, 3186, 3148, 3483, 4174, 3667, 3164, 3411]. Thompson [4135, 2566, 44]. Thomson [796, 2765, 3237, 2765, 3387]. Thorne [4184, 4190, 4176]. Thorpe [2191]. Thorsten [2465, 4169]. thought [4089, 3380]. Thoughtful [2886]. Thouless [3313, 3704]. threads [2570]. three [1552, 2021, 615, 1172, 60, 4192, 2318, 3218]. three-dimensional [615]. Three-Particle [3218]. three-party [4192]. threshold [1430]. Thron [3738]. Throughout [2780, 3907]. Thurber [4053]. Thurner [4017]. Thyssen [3758]. Tian [3441]. tidal [1127]. Tidemand [3619]. Tidemand-Lichtenberg [3619]. Tides [3372, 1127]. Tien [3693]. Tigner [3208, 3484]. tilings [3749]. Tilling [2437]. Tim [3107, 2668]. Timberlake [4037]. Time [2857, 3261, 3406, 2057, 2984, 3207, 4418, 1726, 132, 153, 4381, 2026, 3182, 1873, 2780, 2994, 1554, 1736, 4408, 1835, 3164, 2008, 1454, 1928, 4213, 2382, 1002, 1852, 1967, 1423, 4126, 395, 283, 3637, 1773, 1697, 3421, 3794, 1819, 4121, 2698, 3967, 1647, 1957, 262, 3654, 1373, 3815, 2545, 2981, 3471, 4346, 4208, 3654, 3137, 3023, 2965, 3106, 3379, 3164, 3861]. Time-Dependent [2965]. time-of-flight [283]. Time-resolved [1736]. time-travels [4208]. Timepix [3389]. Times [4400, 2718, 3984, 2057, 1783, 3810, 2648]. timescales [2282, 3379]. Timothy [3283, 3662]. Timour [3413]. Timpson [3210]. tiny [4155]. Tipler [3092]. Tipton [3027]. Tirion [2505]. tissue [4241, 872]. Tissues [3058]. Titanic [2363]. title [2075, 2152]. Tito [2771]. Tiwari [2988, 4162]. Tkalcic [3722]. TMTSF [1139]. Toal [2852]. Tobochnik [4282, 2625]. today [413, 4115]. Tokamak [1218]. Tokamaks [2882]. told [1916]. Tom [2774, 3345, 3305, 2546, 3693, 2952]. Tommy [2801]. Tomographic [2395]. Tomography [2465, 2030, 3098]. tomorrow [413]. Tompkins [2908]. Toms [2524]. Tones [2462]. Tony [3348]. Too [1192]. tool [2339, 1550, 3060, 192, 2641, 638]. Tools [2880, 2825, 2316, 4330, 4201, 3251, 3067]. Top [4367, 4351, 3690, 1679]. Top-Down [4367, 3690]. Topics [2875, 3172, 447, 4070, 848, 2422, 3658, 4297, 2911, 3144]. Topological [4247, 3277, 3189, 1823, 3785, 3872, 2271, 4166, 1728, 4098, 3277, 3836, 3704]. Topology [3241, 3540, 4024]. Tornari [2803]. Torner [2629]. Török [4020]. Toropov [3530]. Torres [4081, 2629, 3265]. Torus [869]. Totality [3751]. Tothill [3113]. Tour [3138, 2951, 3021, 3719, 3929, 4082]. tournament [3633]. Toussaint [3444]. Townsend [3670, 2669]. toys [4324]. Tracing [2013, 1875, 3644]. track [4051]. tracks [552, 827]. Tracy [3644]. tradition [1927]. traffic [2109, 348]. tragedian [2191]. Tragic [4404, 2191]. training [41, 14, 106]. Trajectory [2979]. Trangenstein [2653]. Transactional [3086]. Transfer [2905, 3068, 1922, 637, 653, 3760]. Transform [4353, 3768, 1489, 745]. Transformation [3130]. Transformed [3405]. transforms [4149, 2357]. transient [999, 1008, 398]. transients [3550]. transistors [2306, 3101]. Transition [3026, 2909, 777, 1385, 2209, 4272]. transition-metal-ion [1385]. Transitions [4371, 3280, 4379, 2729, 3539, 4055, 4329, 2282, 1805, 1824, 3452, 4098, 1089, 1366, 697, 1429, 2445, 934, 907, 2224, 1244, 2819, 3230, 2728, 4265, 3920]. translated [2794, 3299, 3103, 3066, 3575, 3379, 2547, 3680]. Translation [100]. Translational [3373]. translator [2371]. transmission [2216, 3198]. Transport [3301, 2402, 2556, 2416, 2745, 3439, 304, 4411, 1364, 1283, 3626, 4072, 4377, 2605, 3417, 1030, 2001, 3760, 4260, 2361, 531, 3118, 2711, 3265, 4081, 4111, 3672]. transuranic [119]. transverse [3539]. trap [1378]. Trapp [3211]. Trapped [1256, 1767, 4212, 2641]. trapped-ion [2641]. trapping [2763, 1536, 1027, 274]. traps [2339, 3161, 2640, 1942]. Travel [4408, 3164, 1967, 281, 305, 312, 329, 337, 344, 3815]. traveled [3473]. traveler [4327]. travels [4208]. tread [6]. Treatise [3296, 3380]. Treatment [850, 246, 531, 3738]. Treccani [4323]. tree [141, 1834]. Tremaine [4301]. Trend [293]. Trends [2760, 3063, 4254, 639, 3499, 906]. Treps [3772]. Treumann [3072]. Trew [3776]. Triangulation [3328]. Tribute [2934, 1459]. Trigger [763]. Trigonometric [2385]. Trimble [4285, 2934]. Trindade [2771]. triple [2165]. triple- [2165]. Tripon [3434]. Tripon-Canseliet [3434]. Trivedi [3230]. Troccoli [3333]. Trojan [1892]. Tropical [1091, 864, 1699, 789]. Troshkin [4231]. Trouton [545]. Trudeau [3823]. true [2133, 1171]. Trypsteen [3773]. Tsai [3299]. Tsai-Chien [3299]. Tsori [2443]. Tsu [2785]. Tsukasa [4284]. Tsunamis [465, 3284, 3372]. Tuan [4302, 3440]. tube [273]. tubes [2079, 673, 3957]. Tuchin [3246, 2671]. Tuck [4351]. Tucker [3741, 2342]. Tudor [3785]. Tuffin [3503]. tunable [1385, 3611]. tuned [3695, 4387]. Tuning [2753]. Tuniz [3120]. tunnelling [1049, 2076, 1502]. tunnelling-state [1049]. turbines [837]. Turbulence [3376, 2481, 3361, 2277, 3143, 2593, 2452, 2689, 2869]. Turbulent [2818]. Turcotte [3588]. Tureanu [3240]. Turin [2940]. Turing [3111]. turn [1902]. turned [3525]. Turner [2988]. Turning [3093, 972, 8]. tutorial [2604]. TVs [2668]. Tweezers [2598, 2005, 3634]. Twentieth [3121, 2682, 2548, 235]. Twentieth-Century [3121, 2682]. Twenty [3369, 2460, 3909, 1254, 1464]. Twenty-first [2460, 3909, 1464]. Twice [784]. twin [3981]. twinkling [902]. twist [1921]. Twisted [3779, 2629]. Twisting [2108]. Twistor [2193, 1564]. Twistors [2544]. Two [3079, 163, 2770, 3381, 2902, 2622, 2278, 3953, 2123, 640, 2163, 2021, 3626, 3358, 1429, 1993, 2259, 1199, 2260, 2555]. Two-and-a-half [2278]. two-dimensional [3358, 2555]. Type [3075, 1430, 1545]. tyre [6]. Tyson [2589, 2638, 3701]. Tyukin [2695].

U [3677, 3444, 3050, 4321, 2295, 2943, 2443]. Ubaldo [2634]. ubiquitous [1038]. Ueno [4284, 3630]. Ugo [3442]. UK [4286, 4202, 4232, 4140, 4139, 4038, 4035, 4100, 4328, 4307, 4333, 4272, 4084, 4303, 4072, 4178, 4021, 4081, 4130, 4203, 4222, 4259, 4269, 4309, 4281, 4270, 4274, 4187, 4283, 4165, 4157, 4197, 4337, 4354, 4079, 4144, 4341, 4306, 4071, 4219, 4276, 4320, 4077, 4199, 4207, 4233, 4335, 4185, 1587]. Ullrich [2965]. Ulrich [2889, 3162, 3310]. Ulrike [4168]. ultimate [1694, 1936]. Ultra [191, 463, 1594, 960, 1039, 2447, 1075, 3232]. Ultra-cold [1039]. Ultra-fast [1594]. Ultra-high [463, 960, 1075]. Ultra-realistic [3232]. Ultracold [3245, 4068, 4118, 2107, 4116, 3022, 2525, 2846]. Ultracondensed [3919]. Ultrafast [1179, 1742, 2233, 3506, 4202, 3683]. ultrahigh [1791]. Ultrasonic [203, 538]. Ultrasonics [1842, 798, 508]. Ultrasound [886]. Ultraviolet [643, 2295]. Umberto [3646]. Umran [2967]. unasked [4080]. unbound [3931]. Uncertainty [3112, 2935, 3414, 1705, 3497, 1658]. uncertainty-principle-limited [1658]. Uncollapsing [2338]. unconditional [3515]. Unconventional [1715]. Uncorked [3196]. Undergraduate [3314, 4253, 3178, 4120, 4078, 4060, 2883, 3122, 3318, 3393, 3025, 3211, 3333, 3052, 3199, 3350, 2502, 2788, 2829, 4062, 2781, 2809, 3073, 4343, 2667, 3509, 4241, 4217, 4284, 4275, 2491, 2560, 3115, 3180, 3458, 4202, 4268, 3301, 2851, 2843, 3107, 3011, 2966, 3485, 4296, 2425, 2411, 3072, 2544, 2471, 2459, 2508, 3037, 3383, 4123, 4184, 2320, 3349, 3137, 4113, 4356, 3035, 4307, 2740, 2867, 2801, 3055, 3142, 3442, 3123, 3296, 3023, 2708, 3507, 2838, 2836, 2853, 3104, 3275, 3194, 3353, 3408, 3493, 3488, 4183, 4240, 4272, 2975, 3144, 3212, 3193, 3092, 3152, 3168, 2735, 2790, 2755]. undergraduate [4084, 4155, 2992, 3245, 2791, 4325, 2541, 2533, 2655, 4173, 2561, 2845, 3117, 3679, 4072, 4178, 4190, 4198, 4252, 2529, 2625, 2537, 2777, 2794, 3181, 3279, 3343, 3344, 3494, 3450, 3452, 4204, 4203, 4162, 4222, 4257, 4259, 4269, 4309, 4281, 4270, 4295, 4323, 3323, 3407, 2887, 3184, 4242, 4250, 3297, 2726, 3081, 3341, 3317, 3394, 3435, 3437, 3436, 4339, 2457, 3099, 4187, 4283, 4294, 2318, 2372, 2348, 2519, 2538, 2675, 2765, 2783, 3046, 3224, 3262, 3163, 3287, 3244, 3196, 3369, 3335, 3405, 2690, 3270, 2612, 4322, 4318, 3431, 3233, 3258, 3366, 3429, 2844, 2724, 3135, 4165, 3034, 3235, 3398, 3492]. undergraduate [2891, 3399, 3128, 3018, 3274, 2365, 2817, 3114, 3200, 3277, 3401, 3005, 3210, 3345, 3415, 3446, 4186, 2371, 3100, 2474, 2476, 3186, 3423, 2682, 3026, 2950, 2881, 2738, 2594, 2497, 2773, 3428, 4321, 2464, 3140, 3027, 2395, 2455, 2517, 2499, 2512, 2940, 3084, 3259, 3232, 3307, 3351, 3486, 4200, 4205, 3049, 3231, 3234, 3243, 3304, 3065, 2587, 3053, 2839, 2908, 2689, 2890, 3155, 3000, 2971, 2985, 3248, 3238, 3282, 3500, 4282, 3272, 2460, 2664, 2877, 3240, 3311, 4326, 2784, 4349, 2415, 2595, 2700, 4201, 2835, 2874, 2837, 2834, 2760, 2880, 3062, 3066, 3119, 2997, 3063, 4054, 4083]. undergraduate [4053, 4080, 4079, 4169, 4170, 4189, 4234, 2436, 2391, 2657, 2677, 2676, 2854, 2728, 3028, 3109, 3183, 3188, 2440, 2611, 2607, 2556, 2513, 2705, 2860, 2722, 3096, 2976, 2418, 2695, 2510, 2511, 3064, 3414, 3380, 3476, 4071, 4255, 4208, 4188, 4311, 4276, 4350, 4070, 4066, 4077, 4167, 2310, 2314, 2390, 2608, 2620, 2896, 2895, 3197, 3228, 3461, 4265, 4266, 4332, 2317, 2377, 2331, 2568, 2719, 2805, 3148, 3255, 3264, 3265, 3482, 2531, 3151, 3285, 4171, 4174, 4216, 2442, 3286, 2540, 2951, 2952, 2758, 2947, 3236, 2982, 2401, 2619, 3191, 3185, 3278, 2417, 4290, 2488]. undergraduate [2618, 2826, 2905, 2683, 2492, 3087, 2669, 2536, 2475, 2373, 2613, 2646, 2507, 2591, 2656, 2484, 2802, 2956, 2721, 2900, 2710, 2736, 2832, 3082, 3153, 3110, 3041, 3215, 3242, 3170, 3260, 3411, 3338, 3372, 3475, 3495, 4215, 4335, 4175, 4185, 2319, 2917, 3378, 3355, 2882, 2852, 2358, 2768, 3164, 2769, 4232, 4304, 4293, 4333, 4327, 4156, 4105, 4353, 4319, 4292, 4112, 4161, 4337, 4354, 4244, 4249, 4351, 4341, 4352, 4306, 4066, 4308, 4086, 4175, 4258, 4243]. undergraduate/beginners [4353]. undergraduate/beginning [2442]. undergraduate/graduate [2683]. undergraduate/postgraduate [2502, 2788]. undergraduate/pre [2358]. undergraduate/pre-university [2358]. undergraduate/specialist [3399]. Undergraduates [3014, 2921, 2297, 2936, 2911, 2868, 2865, 3384, 2293, 2866, 2942, 2749, 2842, 2504, 2907, 2937, 2494, 2901, 2684, 3496, 2944, 3043, 2993, 3364, 3438, 3422, 2333, 2300, 2523, 2461, 2438, 2525, 2889, 3454, 3487, 2929, 2933, 2864, 4117, 2899, 2649, 3216, 2924, 2943, 3198, 2833, 2717, 3483, 2579, 2468, 2501, 2953, 2848, 2559, 3068, 3069, 3021, 4058, 4207, 2922, 2915, 2967, 2919, 4176]. underground [456]. Underlie [4367, 3690]. underlying [1124, 2208]. Understand [3034, 3763, 4005]. Understanding [3089, 1852, 2844, 2929, 2722, 1958, 2417, 3372, 33, 3711, 3290, 3739, 4200, 3867, 3802, 4324, 2819, 2537, 3177, 3243, 2513, 3979, 3928]. undoing [2338]. Undulators [971]. undulatory [2337]. unfinished [3291]. Unification [1490, 1486]. Unified [1001, 3021, 3839, 1305]. uniform [61]. Unifying [1566, 3967]. Union [3938, 3039, 3267, 70, 2008]. unique [3572]. Unitary [2793]. United [222, 358]. Units [4363, 390, 810, 3877, 387, 374, 1401, 176, 797, 560, 311]. unity [112, 1124]. Universal [4372, 3453, 3860]. universality [1490, 4276]. UNIverse [2814, 2350, 2495, 2742, 2707, 1799, 3302, 1824, 4374, 3455, 2690, 4384, 1811, 2714, 2097, 2493, 2753, 3049, 3243, 2778, 2815, 1647, 2576, 2895, 3197, 2423, 1505, 1680, 4407, 2637, 2345, 2917, 2450, 2162, 3602, 2156, 865, 3695, 4123, 3869, 1758, 1637, 3249, 3931, 3620, 3762, 2127, 1741, 4196, 1466, 1472, 4387, 1302, 894, 1100, 1183, 3532, 2068, 3636, 3967, 1015, 3963, 1925, 1392, 3222, 3378, 2166, 1121, 3241, 2747, 3081, 3099, 2318, 2473, 2430, 4151, 3151, 1562, 2970, 3594, 4275, 3635, 4143, 3650, 4200, 3533, 3956, 4243, 2459, 3834]. Universe- [1811]. universities [41]. University [4078, 4060, 4016, 4211, 4137, 4286, 4082, 4275, 4135, 4038, 4000, 4032, 4210, 4035, 53, 3425, 4005, 3945, 4045, 4073, 4182, 4177, 4184, 4027, 4100, 4356, 4328, 3035, 4307, 4061, 2057, 4143, 4084, 4248, 112, 4327, 4355, 4291, 4285, 4344, 4026, 4072, 4052, 4178, 4190, 4198, 2191, 4006, 4021, 4081, 4156, 4130, 4203, 4162, 4309, 4274, 237, 4025, 4002, 4001, 4046, 4043, 4041, 4085, 4141, 3459, 4037, 4040, 3994, 4048, 4042, 4003, 4015, 4105, 4146, 4345, 4338, 4342, 4346, 4339, 4340, 4187, 4322, 4165, 3421, 4097, 4107, 4157, 4292, 4007, 4017, 4158, 4093, 4200, 4151, 4205, 290]. University [4019, 4282, 4326, 4131, 4122, 4197, 4316, 4020, 4036, 4079, 4071, 4129, 4219, 4301, 4311, 4310, 4276, 4350, 4320, 4066, 4077, 4166, 4199, 4266, 4065, 4064, 4063, 4092, 4034, 4109, 4039, 4023, 4058, 4176, 4215, 4335, 2787, 4175, 4185, 4168, 2358, 4256, 270, 2809, 4343, 4232, 3349, 3137, 3152, 2944, 942, 245, 3315, 3405, 3135, 3121, 3128, 3415, 4186, 3100, 3423, 2682, 2881, 2899, 3234, 3243, 2908, 3066, 3461, 2938, 2913, 3285, 4086, 713, 2826, 3134, 3215, 3164, 766, 4317, 4163]. Unpublished [2674]. Unravelling [2101, 1374]. unreason [3641]. Unrolling [1454]. unruly [4143]. Unseen [4374, 3455]. unshackled [3525]. Unterborn [4150]. Unveiling [2813]. Unwin [291]. updated [3817]. upon [2008]. upper [923]. Urbani [4210]. Urs [4269]. US$140.00 [4168]. US$156.00 [4323]. US$159 [4027]. US$59.99 [4185]. US$88.12 [4294]. US$92.70/GBP70 [4353]. USA [4183, 4190, 4269, 4274, 4310, 4297, 4332, 4256]. USD [4061, 4333, 4190, 4207, 4243, 4256, 4267]. Use [2909, 2722, 4057, 69, 1048, 2235, 657, 807, 1752, 926, 962, 638, 209, 262, 3624, 3807]. used [340]. useful [1621, 3078]. useless [3525]. User [2380]. Users [2633]. Uses [884, 1314]. Using [3464, 2174, 2847, 4230, 2518, 3314, 3059, 2463, 631, 3559, 482, 791, 4373, 1020, 3450, 260, 4353, 1885, 909, 2062, 3770, 3351, 1271, 1693, 4219, 4236, 2155, 7]. usually [129]. Utke [3059]. Utracki [2767]. UV [1788]. Uzunov [2715].

V [2414, 2659, 3313, 2425, 4231, 2350, 3076, 3530, 3444, 2639, 3648, 2563, 2791, 326, 2541, 2806, 4044, 4008, 4098, 3007, 2609, 4187, 2348, 2402, 3998, 2437, 2343, 2379, 2916, 2671, 3246, 2803, 3218, 2914, 2821, 555, 3127, 2750, 3607, 4234, 4129, 4219, 3014, 1461, 2377, 2804, 3483, 3814, 3743, 2531, 3788, 2947, 2424, 2807, 4059, 3032, 2934, 3474, 3495, 4116, 2319, 3593, 2547, 3680, 337]. vacua [463]. Vacuum [2856, 1263, 2264, 3780, 4050, 960, 2263]. Vahid [3772]. Valenta [2878]. Valenzuela [3184]. Valeri [3576]. Valerio [2741, 4243]. Valery [2671, 3246, 2815]. Valia [4154]. Valles [3230]. Valley [879]. Valued [2766, 4061]. VanderPlas [3577]. Vandoren [3379]. Vangelis [2685, 2743, 2739]. Vanier [2917]. vanished [3981]. VanWyk [2457]. Vapour [625]. vapours [451, 2763]. Varadan [2836]. Varga [2749]. Variability [2898]. variables [2180]. variation [141]. variational [3819]. variations [1685]. Varieties [1129]. variety [2166]. Varlamov [2913]. Varying [2441]. Vasa [3683]. Vasilevich [3984]. vast [3935]. Vaughan [3259]. VCH [4295, 4323, 4022]. Vecchia [3004]. Vectors [2944]. Vedral [3180, 2495, 2655]. Vedres [3031]. vehicles [671, 2745]. Vekstein [3233]. Veljić [4322, 4326]. velocities [1737]. velocity [817, 61]. Velusamy [4324, 4319, 4297, 3657, 3658]. Velychko [3747]. Velzel [3605]. Venerus [4072]. Venkatesh [3041]. Venus [4237, 1022]. Verbiest [2614]. Verdaguer [4141, 3446]. Vergni [4208]. verification [2235]. Verlag [4022]. Veronica [4272, 3022]. versatile [2339]. versus [868]. vertebrates [590]. Very [2989, 3120, 3405, 3018, 3263, 2434, 2426, 2650, 2990, 3757, 3622, 3651, 3790, 3805, 4016, 375, 1833, 4073, 3599, 1101, 517, 3856, 981, 2413]. Veryaskin [4234]. Vespignani [2352]. Vesselin [4191]. vi [554, 1748, 2378]. via [3138]. vibration [253, 1258, 955]. vibrational [555]. Vibrations [4365, 3077, 3424, 2520, 2900, 88, 4071]. vibroacoustics [3715]. Victor [3308, 3114, 2753, 3521]. Victoria [248]. Victorian [3472, 1827]. video [2039]. Vienna [834]. Vietri [2415]. View [3167, 2619, 1137, 420, 4048, 1649, 944, 2030, 3394]. viewpoint [1749]. Views [2196, 1836, 2770]. Vignale [2375, 2679]. vii [555]. viii [2057]. Vijay [2836]. Viking [3251]. Viktoria [3496]. Vincent [2614, 2852]. Vineet [2556]. Vinod [2853]. violation [1978, 1647, 2586, 3061]. violin [1684, 1284]. Violins [2532]. viral [2893, 2880]. Virginia [4285]. Virtual [1636, 1774, 580, 4194, 4232]. Virtual-reality [1636]. virus [629, 4290]. Viscoelasticity [2927]. viscosity [2235]. Vishveshwara [2379]. visible [806, 3697, 4233, 3623]. Vision [3145, 3151, 590, 1227, 601]. Visions [1857, 2574]. visitor [36]. Visitors [3355]. Visual [775]. Visualisations [3314]. visualization [3432, 4413]. Vita [4115, 2380]. Vita-Finzi [4115, 2380]. Vital [4407, 3869]. Vitaly [2324]. Vivien [3481]. Vivo [4415, 2470, 3457]. Vladimir [2800, 3204, 2678, 4322, 3114, 3230, 3632, 2652, 3276, 4326, 4249, 4094]. Vladislav [2541]. Vlatko [3180, 2495, 2655]. Vlemmings [3203]. Vo [3440]. Vo-Dinh [3440]. Voicu [3319]. voids [4135]. Vol [380, 2858, 291, 2806, 2794, 3187, 2288, 2416, 2804, 3020, 379, 417, 3934, 4184, 317, 316, 4224]. volcanoes [2001]. Voldemar [4163]. Volker [4349]. Vologodskii [3606]. Volovik [2423]. Volpe [3634]. Vols [2765]. voltage [125, 476, 616]. Volume [2883, 2926, 2527, 2596, 2840, 2795, 3371, 3353, 3889, 4190, 3364, 3422, 3200, 2437, 3246, 2744, 2818, 2760, 3119, 3836, 3132, 4265, 2822, 3953, 3509, 2858, 4009, 3854, 946, 4306, 2535, 4095, 3976, 4176]. Volumes [3079, 366, 3216, 3172, 3440, 3558, 1483]. vortex [1737]. Vukmirović [4322, 4326]. Vulpiani [4208, 2351].

W [3052, 2296, 2357, 3796, 4000, 2774, 3376, 2347, 3371, 2995, 2636, 3738, 2257, 3275, 3524, 3444, 2847, 2791, 3620, 4026, 4052, 3054, 1069, 3126, 3999, 4068, 3754, 3804, 4346, 2489, 1359, 2886, 2249, 2243, 2247, 2354, 2505, 2773, 2940, 3653, 4224, 3231, 2681, 2778, 2391, 4351, 2510, 2711, 3996, 2316, 2805, 3567, 3149, 2952, 2409, 2253, 2632, 2454, 2599, 2646, 2663, 3484, 3675, 3684, 2534, 2787, 3039, 1301, 44, 3633, 3694]. Wa [3109]. Waals [718, 579]. Wakeham [4269]. Walczak [3613]. Walecka [4062, 2615, 2823, 3144, 2937]. Wales [315]. Wali [2532]. walk [2195, 4208]. Walker [3684, 3564, 3773, 3732, 2783, 2670]. Walks [2825, 2054, 2268, 2267]. wall [748]. Wallace [4037, 3469]. Waller [3804]. walls [2551]. Walsh [3647]. Walter [4132, 3015, 3053, 2994, 2889]. Walther [2296]. wandering [3942]. Wang [3629, 4086, 3542, 3633, 2756, 3542]. War [4357, 4358, 3322, 3335, 2329, 3226, 1041]. wardana [3167]. warm [2156]. Warp [3164]. wars [2010, 4401, 3875]. Wartak [3297, 3140]. wartime [2452]. Washington [4256]. Wasserman [3064, 3815]. waste [1501, 925]. Wataru [2657]. watcher [1059]. Water [594, 1826, 3699, 2216, 3867]. waters [2643, 2484]. Watson [2644]. Wave [2659, 1994, 2567, 3361, 1328, 2974, 2711, 1518, 1541, 3644, 4154, 1283, 1760, 3510, 1460, 2515, 1793, 2266, 3893, 3811, 3863, 2265, 2300]. Wave-Boundary-Layer [2974]. wave-particle [3510, 3863]. wavefront [3418]. Wavefronts [4293, 2776]. wavefunction [2338, 3346]. waveguide [2567]. Waveguides [1584, 2486]. wavelengths [3697]. wavelet [3768]. Wavepacket [1892]. wavepackets [3506]. Waves [3199, 1265, 3007, 2817, 332, 3214, 3191, 3284, 644, 4140, 912, 423, 3747, 191, 4329, 223, 3642, 3854, 49, 2307, 3060, 1013, 4101, 978, 234, 4121, 1468, 4244, 3572, 1812, 3812, 538, 3792, 3107, 3370, 2900, 3699, 4281, 4064, 4086, 4016, 2458, 2702, 3787, 3381, 3272, 4080, 3543, 2484]. Way [3781, 3804, 3758, 3933].

Wayne [4140]. Ways [2928, 3414]. Weaire [2259, 2041, 2334]. Weak [4299]. Weapons [3322]. Wear [2726, 162]. Weather [3035, 952, 2956, 3333, 2063, 572, 1310]. Web [3268, 3650]. Webb [3821, 3201]. webs [2361]. Webster [3407]. Wee [2773]. Wei [4119, 3003, 3118, 3537, 2799]. Weidberg [3613]. Weigand [3066]. weight [1059]. weight-watcher [1059]. Weikun [3030]. Weinberg [4063, 2999, 3561, 3045]. Weiner [3122, 1092, 3712]. Weinheim [4252, 4295, 4323, 4242, 4022, 2627]. Weintraub [4285, 2690]. weirdness [3694]. Weiren [3371]. Weiss [3689, 3162]. Weisskopf [3308]. Weitkamp [2288]. Weitz [3505]. welding [459, 624]. well [1727]. Weltman [2984]. Welzel [3050]. Wendell [3688]. Wendisch [2905]. were [4355]. Wermes [4158]. Werner [2492, 3510]. Wesson [4121, 2882, 2450]. Western [3163]. Westfall [2356]. Westmoreland [2536]. Weyl [2451]. Wheatstone [785]. Wheeler [1653]. Where [2305, 2080, 1392]. which [623, 637]. whimper [1904]. Whipple [2353]. Whisnant [3032]. Whispering [770]. Whitaker [3281]. white [1749]. Whitehouse [2919]. Whitney [4201]. Whittle [4204]. Who [63, 4400, 662, 3810, 3731]. whole [1171]. Wholefield [2557]. Wickramasinghe [3423]. wide [2737, 3113, 2467, 3084]. wide-ranging [2737, 3113, 2467]. Widrith [77]. width [753]. Wiesbaden [4218]. Wigmans [3818]. Wigner [2195]. Wijers [3057]. Wilcox [3738, 3053]. Wild [3809]. WILEY [4323, 4120, 4273, 4304, 4272, 4252, 4235, 4295, 4242, 4250, 4354, 4022, 3995, 4334, 4332, 2888, 4116]. Wiley-VCH [4295, 4022]. Wilk [3262]. Wilkinson [4040, 2126, 3195]. Will [2843, 4269, 3264, 3403, 1967, 63]. William [3052, 3116, 4027, 3023, 2906, 4269, 2567, 2489, 2765, 2783, 3093, 3080, 4007, 2860, 3096, 3402, 2720, 3536, 3595, 3082]. Williams [4222]. Willink [3112]. Willis [3236]. Wilson [3349, 2765]. wily [2002]. Wim [3002]. WIMP [2046]. win [4057]. wind [1887, 423, 2550, 1085, 1473]. wind-generated [423]. Window [2322, 3406]. Winitzki [2425]. Winn [4338]. winner [4200, 4151, 4205]. winning [3476]. Winter [2851]. Winterberg [2928]. Wio [3132, 3123]. Wireless [923]. Wires [1833]. Wisdom [3264]. Wiseman [2530]. Within [3472]. Without [2493, 3968, 3420, 1621, 653, 2399]. Wittgenstein [2770]. Wohl [4296]. Wohland [4169]. Wohlman [4096]. Wolf [3177, 4092, 3055, 3386, 3279, 2309, 3837]. Wolfenstein [2608]. Wolfgang [3072, 2945, 2057, 3202, 3066, 4418, 356]. Wolski [3350]. women [4285]. Wonderful [3150, 2683, 3817, 3756]. Wonders [2913, 2056, 3623, 3973]. Wong [2838, 3348, 3299, 3109]. Wood [4207, 141, 2816]. Woodruff [4215]. Woodstock [4274]. Woodworth [3372]. Wool [136]. Woolfson [4114, 3234, 3106, 2938, 3443, 3624]. Woosley [3057]. Wooten [2432]. words [4285]. work [2433, 211, 4154, 331, 1843, 2962, 807, 3568, 139, 3281]. Workbook [3043]. worker [2350]. workers [2395, 2469]. Working [2449, 177, 2355]. workings [1982]. Works [3079, 3885, 2602, 3432, 4413]. Workshop [3593, 2414, 765, 2387]. World [4132, 4106, 4010, 4057, 4324, 4062, 4225, 4268, 4056, 4055, 4231, 4293, 3133, 3875, 4028, 4123, 4069, 2082, 4113, 2867, 3150, 3322, 3308, 4030, 4138, 4163, 3999, 4029, 4044, 4008, 4024, 4108, 4068, 4119, 4098, 4196, 3998, 2371, 3423, 4112, 4114, 4099, 4164, 4121, 4223, 4083, 4080, 3239, 3996, 4255, 4070, 4167, 3337, 4095, 4096, 4265, 4086, 4401, 4004, 3956, 4033, 3997, 4067, 4172, 4243, 3326, 254, 3719, 1833, 3010, 2017, 3516, 2384, 3563, 1325, 3756, 1840, 3949, 83, 288, 1630, 4233, 614, 2742, 2856, 3287, 2971, 3134, 4274, 4340, 3838]. Worlds [2508, 4372, 1092, 3878, 2962, 3453, 3010]. Wormholes [2857, 4411]. Worster [2513]. Wouter [3203, 4018]. Wren [1251]. Wright [3771, 2373]. Writings [2674, 3798]. Wrong [2864, 2549]. WSPC [4351]. Wu [3999, 3208, 3299]. Wulfman [2859]. www.cambridge.org [4211]. www.elsevier.com [4186]. Wyart [2527, 2596]. Wyngaard [2593].

X [3629, 1139, 3047, 3131, 2680, 2295, 2633, 3675, 3775, 4182, 4274, 617, 314, 1204, 2013, 328, 802, 553, 2572, 3550, 629, 573, 48, 583, 938, 2642, 3601, 91, 1419, 4079, 911, 3387, 698, 699, 1736, 3015]. X-ray [3131, 2680, 2295, 2633, 3675, 3775, 617, 314, 1204, 328, 802, 553, 2572, 3550, 629, 48, 583, 938, 2642, 3601, 91, 1419, 4079, 911, 3387, 698, 699, 1736]. X-rays [2013, 573, 3015]. XAFS [2680]. xerography [962]. xi [4018]. Xia [3030]. Xiangqun [3113]. Xiao [3003, 3542]. Xiao-Dong [3542]. Xiaoyi [3098]. Xie [2836]. xii [662, 4139, 4295, 4337, 3337]. xiii [4218]. Xin [2756, 2784]. Xin-Nian [2756]. Xin-She [2784]. xiv [4123, 4321]. xix [4191]. Xun [2660]. xv [4140, 4154, 4257, 4309, 4281]. xvi [4177, 4155, 112, 4235, 4270]. xviii [4005, 4248, 4354]. xx [4045]. xxi [3618, 4221, 4259]. xxii [4184]. xxiv [4269]. XXIX [417]. XXIXth [3593]. XXVI [317]. XXVII [316]. XXVIII [379]. xxx [869].

Yabe [2789]. Yagi [2313]. Yair [4255]. Yakovlev [2541]. Yale [2787]. Yamaji [3103, 2789]. Yamamoto [3993, 3802]. Yan [3319, 3560]. Yanagisawa [3103]. Yang [2948, 4001, 4024, 2341, 3003, 2784, 2905]. Yanofsky [2317]. Yaroslav [3079]. Yaroslavsky [3559]. Yasuhiko [2472]. Year [1082, 3289, 3193, 2765, 3335, 2396, 2300, 2126, 2231]. Years [3313, 3171, 2755, 2850, 3267, 3885, 1026, 1981, 1131, 784, 1769, 3987, 1156, 3886, 2303, 1117, 1254, 2571, 1079, 3797, 3428]. Yeh [2486, 3998]. Yeseren [4290]. Yi [4068, 3003, 3542]. yields [1826]. Yin [3998, 2982]. Ying [4347]. Yitzhak [2622]. Ylinen [3710]. Yn [2927]. Yn-Hwang [2927]. Yoccoz [2315]. Yoel [3213]. Yor [3138]. York [4000, 4148, 3459, 3421, 4117, 4158, 144, 4171, 4116]. Yoshi [2657]. Yoshimura [2998]. Young [3318, 3609, 4322, 4326, 3633]. yttrium [791]. Yu [2485, 2792, 2783, 2822]. Yuan [2783]. Yuan-Yu [2783]. Yuen [3506]. Yuen-Zhou [3506]. Yuichi [2657]. Yuka [4150]. Yukawa [2134]. Yukio [3143]. Yuli [3076]. Yun [2930]. Yun-Shik [2930]. Yupapin [3534]. Yuri [3276]. Yurov [4044]. Yves [2834, 2610].

Z [2858, 4056, 3312, 2422, 2827]. Zaffaroni [4070]. Zafiropulos [2803]. Zagoskin [2797]. Zain [4188]. Zaitsev [3007]. Zamboni [3370]. Zamboni-Rached [3370]. Zamolodchikov [3079, 3079]. Zamponi [4210]. Zandvoort [3795]. Zanghi [4154]. Zangwill [4320, 3166]. Zanni [2696]. Zanotti [3278]. Zanotto [3254]. Zap [3262]. Zappe [3611]. Zaremba [2580]. Zavidovique [2884]. Zdzislaw [2399]. Zee [2666, 3625, 3155]. Zeljko [3577]. Zelnikov [3576, 2914]. Zenghu [3011]. Zero [843]. Zero-gravity [843]. Zettili [2579]. Zewail [3432]. Zhang [4068, 4119, 4107, 3674, 3145]. Zheng [4250, 3823, 3996, 3674]. Zhong [4167]. Zhong-Qi [4167]. Zhou [3506]. Zhu [2982, 3165]. Zichichi [3702]. Ziman [2935]. Zimmermann [3208, 3484]. Zipf [2142]. Zirker [3284]. ZnO [3003]. Zoltán [3591]. zone [525, 274]. zones [274, 282]. Zooniverse [4109]. Zuber [2701]. Zuckerman [4123]. Zwaan [2440].

References

Hinshelwood:1959:F

[1] C. N. Hinshelwood. Foreword. Contemporary Physics, 1(1):1, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1959:MB

[2] O. R. Frisch. Molecular beams. Contemporary Physics, 1(1):3–16, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blin-Stoyle:1959:SAN

[3] R. J. Blin-Stoyle. The structure of the atomic nucleus. Contemporary Physics, 1(1):17–34, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kellner:1959:AHO

[4] L. Kellner. Alexander von Humboldt and the organization of international collaboration in geophysical research. Contemporary Physics, 1 (1):35–48, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beck:1959:SDE

[5] H. V. Beck. The selection and design of experiments for alternating current teaching. Contemporary Physics, 1(1):49–55, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sabey:1959:SFE

[6] Barbara E. Sabey. Skidding friction the effect of hysteresis losses in tyre tread rubber. Contemporary Physics, 1(1):56–61, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1959:LSU

[7] E. J. Wilson and J. D. H. Hughes. Light sources using radioisotopes. Contemporary Physics, 1(1):62–69, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1959:RTP

[8] Anonymous. A review of: “Turning points in physics”. Contemporary Physics, 1(1):70–76, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Emms:1959:RRG

[9] D. A. Emms. A review of: “Reading German for scientists”. Contemporary Physics, 1(1):77–80, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1959:SR

[10] H. S. W. Massey. Space research. Contemporary Physics, 1(2):81–95, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tolansky:1959:DP

[11] S. Tolansky. Diamond. Part 1. Contemporary Physics, 1(2):96–111, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

TerHaar:1959:SRD

[12] D. Ter Haar. Some recent developments in the many-body problem. Contemporary Physics, 1(2):112–133, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosser:1959:ESO

[13] W. G. V. Rosser. Electromagnetism as a second-order effect (I). The definition of the ampere. Contemporary Physics, 1(2):134–141, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lester:1959:PST

[14] F. Lester. Physical science for teachers in training. Contemporary Physics, 1(2):142–146, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cockcroft:1959:CC

[15] Sir John Cockcroft. Churchill College. Contemporary Physics, 1(2): 147–150, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Elton:1959:CHP

[16] L. R. B. Elton. A clearing house for physics students. Contemporary Physics, 1(2):151–153, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1959:F

[17] Anonymous. Foundations. Contemporary Physics, 1(2):154–158, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1959:RSC

[18] Anonymous. A review of: “The spiritual crisis of the scientific age”. Contemporary Physics, 1(2):159–162, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1959:RES

[19] Anonymous. A review of: “Exploring the structure of matter”. Contemporary Physics, 1(2):163–168, 1959. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chick:1960:PR

[20] D. R. Chick. Plasma research. Contemporary Physics, 1(3):169–190, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1960:HNNa

[21] N. Feather. A history of neutrons and nuclei. Part 1. Contemporary Physics, 1(3):191–203, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1960:CRP

[22] J. G. Wilson. Cosmic radiation: The primary particles. Contemporary Physics, 1(3):204–219, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Crombie:1960:SRS

[23] A. C. Crombie. The scientific revolution of the seventeenth century and its consequences. Contemporary Physics, 1(3):220–229, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holloway:1960:SG

[24] D. G. Holloway. The strength of glass. Contemporary Physics, 1(3): 230–235, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scupham:1960:TTS

[25] John Scupham. Television and the teaching of science. Contemporary Physics, 1(3):236–241, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lockett:1960:SMA

[26] H. Lockett. The science masters’ association annual meeting. Contemporary Physics, 1(3):242–244, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:IPR

[27] Anonymous. The Institute of Physics report on the teaching of mathematics to physicists (1959). Contemporary Physics, 1(3):245–247, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:F

[28] Anonymous. Foundations. Contemporary Physics, 1(3):247, 1960. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:DRG

[29] Anonymous. D.S.I.R. grants. Contemporary Physics, 1(3):247, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lewis:1960:PSS

[30] J. L. Lewis. The physical science study committee (1) a survey of its publications. Contemporary Physics, 1(3):248–252, 1960. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1960:RHS

[31] G. R. Noakes. A review of: “Handbook of scientific instruments and apparatus”. Contemporary Physics, 1(3):253–256, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1960:HNNb

[32] N. Feather. A history of neutrons and nuclei. Part 2. Contemporary Physics, 1(4):257–266, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yarnold:1960:UST

[33] The Rev. G. D. Yarnold. Understanding surface tension. Contemporary Physics, 1(4):267–275, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tolansky:1960:DP

[34] S. Tolansky. Diamond. Part 2. Contemporary Physics, 1(4):276–286, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Naylor:1960:PC

[35] D. Naylor. Primary cells. Contemporary Physics, 1(4):287–303, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Siddall:1960:PSE

[36] W. Siddall. The 44th physical society exhibition, 1960: (1) a visitor’s impressions. Contemporary Physics, 1(4):304–305, 1960. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1960:PSE

[37] M. H. Wilson. The 44th physical society exhibition, 1960: (2) the physical society colour group’s exhibition. Contemporary Physics, 1(4):305– 306, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Essen:1960:PSE

[38] L. Essen. The 44th physical society exhibition, 1960: (3) atomic clocks. Contemporary Physics, 1(4):306–307, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wright:1960:PSE

[39] D. A. Wright. The 44th Physical Society Exhibition, 1960: (4) recent advances in solid-state physics. Contemporary Physics, 1(4):307–308, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Boulind:1960:PSS

[40] Henry F. Boulind. The Physical Science Study Committee. (2) The Summer Institute courses. Contemporary Physics, 1(4):309–312, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:PTP

[41] Anonymous. The postgraduate training of physicists in British universities. Contemporary Physics, 1(4):313–314, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1960:GFR

[42] G. R. Noakes. Gulbenkian foundation reports. Contemporary Physics, 1(4):315–317, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:SES

[43] Anonymous. Science education in a scientific age. Contemporary Physics, 1(4):318, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Woodward:1960:ERA

[44] L. A. Woodward. Essay review: Advances in spectroscopy [edited by H. W. Thompson]. Contemporary Physics, 1(4):319–321, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:BRa

[45] Anonymous. Book reviews. Contemporary Physics, 1(4):322–336, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salam:1960:EP

[46] Abdus Salam. Elementary particles. Contemporary Physics, 1(5):337– 354, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kip:1960:CRS

[47] Arthur F. Kip. Cyclotron resonance in solids. Contemporary Physics, 1(5):355–369, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lipson:1960:XRC

[48] H. Lipson. X-ray crystallography — a meeting place of the sciences. Contemporary Physics, 1(5):370–384, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jephcott:1960:MW

[49] D. F. Jephcott. Magnetohydrodynamic waves. Contemporary Physics, 1(5):385–389, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bragg:1960:AMF

[50] Sir Lawrence Bragg. Atoms and molecules: A Friday evening discourse delivered on November 20 1959. Contemporary Physics, 1(5):390–393, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Green:1960:PSC

[51] L. L. Green and H. Muirhead. The Physical Society Conference on nuclear physics at Liverpool, March 1960. Contemporary Physics, 1(5): 394–399, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hutcherson:1960:PSS

[52] S. W. Hutcherson. The preparatory schools science course at denstone college, 1960. Contemporary Physics, 1(5):400–402, 1960. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bullen:1960:RDR

[53] G. J. Bullen. Research day at the Royal Institution. The Department of Physics, Birkbeck College, University of London. Contemporary Physics, 1(5):403–405, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:ODR

[54] Anonymous. Optical design research at imperial college. Contemporary Physics, 1(5):406, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1960:CRa

[55] G. R. Noakes. Contemporary reading. Contemporary Physics, 1(5): 407–409, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plant:1960:ERE

[56] Sir Arnold Plant. Essay review: Electronic computers and their business applications [by A. J. Burton and R. J. Mills]. Contemporary Physics, 1(5):410–411, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:BRb

[57] Anonymous. Book reviews. Contemporary Physics, 1(5):412–416, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cottrell:1960:MS

[58] A. H. Cottrell. The metallic state. Contemporary Physics, 1(6):417– 432, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Johnson:1960:CSI

[59] Martin Johnson. The clue to a stellar interior from apse motion of a binary orbit. Contemporary Physics, 1(6):433–444, 1960. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1960:DTD

[60] C. A. Taylor. Diffraction in three dimensions. Contemporary Physics, 1(6):445–452, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosser:1960:ESO

[61] W. G. V. Rosser. Electromagnetism as a second-order effect (II). The electric and magnetic fields of a charge moving with uniform velocity. Contemporary Physics, 1(6):453–466, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henderson:1960:IF

[62] H. Henderson. Interference filters. Contemporary Physics, 1(6):467– 478, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Laybourn:1960:WWT

[63] K. Laybourn. Who will teach? Contemporary Physics, 1(6):479–482, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1960:RST

[64] L. F. Bates. The Royal Society tercentenary celebrations. Contemporary Physics, 1(6):483–485, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1960:MET

[65] V. T. Saunders. Medical electronics: Third international conference and exhibition. Contemporary Physics, 1(6):486–494, 1960. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1960:RPP

[66] N. Thompson. Reports on progress in physics. Contemporary Physics, 1(6):495–497, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:BRc

[67] Anonymous. Book reviews. Contemporary Physics, 1(6):498–500, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stevens:1960:M

[68] K. W. H. Stevens. Masers. Contemporary Physics, 2(1):1–13, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:CSU

[69] Anonymous. O.E.E.C. seminar on the use of television for the teaching of science. Contemporary Physics, 2(1):14, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:ISR

[70] Anonymous. The international scientific radio union. Contemporary Physics, 2(1):14, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Poole:1960:NB

[71] M. J. Poole. The neutron booster. Contemporary Physics, 2(1):15– 20, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tricker:1960:SP

[72] R. A. R. Tricker. The sceptical physicist. Contemporary Physics, 2(1): 21–30, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Allen:1960:TG

[73] W. D. Allen. The tandem generator. Contemporary Physics, 2(1):31– 38, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

White:1960:PSS

[74] Stephen White. The physical science study committee (3) the planning and structure of the course. Contemporary Physics, 2(1):39–54, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1960:PBA

[75] L. F. Bates. Physics at the British Association, Cardiff meeting, 1960. Contemporary Physics, 2(1):55–58, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1960:NPM

[76] W. S. C. Williams. The nuclear physics meeting of the Institute of Physics and the Physical Society, Glasgow, September, 1960. Contemporary Physics, 2(1):59–62, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1960:OAE

[77] V. T. Saunders. Opening of the Atomic Energy Establishment, Widrith. Contemporary Physics, 2(1):63–68, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:ICP

[78] Anonymous. International conference on physics education. Contemporary Physics, 2(1):69–70, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wenham:1960:CPS

[79] E. J. Wenham. Conference of the physical sciences section of the association of teachers in colleges and departments of education. Contemporary Physics, 2(1):71–72, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1960:CRb

[80] G. R. Noakes. Contemporary reading. Contemporary Physics, 2(1): 73–74, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linfoot:1960:RSE

[81] E. H. Linfoot. A review of: “La structure et l’evolution de l’univers”. Contemporary Physics, 2(1):75–77, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1960:RST

[82] Anonymous. A review of: “The special theory of relativity”. Contemporary Physics, 2(1):78–80, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rotherham:1960:EWN

[83] L. Rotherham. Energy for the world’s needs. Contemporary Physics, 2 (2):81–92, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1960:FSM

[84] L. Pearce Williams. Faraday and the structure of matter. Contemporary Physics, 2(2):93–105, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stephenson:1960:SS

[85] I. M. Stephenson. Science in space. Contemporary Physics, 2(2):106– 114, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1960:CA

[86] G. R. Noakes. Corso di aggiornamento. Contemporary Physics, 2(2): 115, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evans:1960:RA

[87] J. V. Evans. Radar astronomy. Contemporary Physics, 2(2):116–142, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lee:1960:NLV

[88] E. W. Lee. Non-linear vibrations. Contemporary Physics, 2(2):143– 156, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1960:POR

[89] V. T. Saunders. The Post Office Research Station, Dollis Hill, London, N.W.2 open days 22–23 September 1960. Contemporary Physics, 2(2): 157–164, 1960. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sutton:1961:AFM

[90] Sir Graham Sutton, C.B.E. Advancing fronts in meteorology. Contemporary Physics, 2(3):165–182, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nixon:1961:XRM

[91] W. C. Nixon. X-ray microscopy. Contemporary Physics, 2(3):183–197, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Freitag:1961:FS

[92] E. H. Freitag. The friction of solids. Contemporary Physics, 2(3):198– 216, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mitchell:1961:PD

[93] Dr. E. W. J. Mitchell. The physics of diamond. Contemporary Physics, 2(3):217–222, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shoenberg:1961:LTP

[94] D. Shoenberg. Low temperature physics. Contemporary Physics, 2(3): 223–230, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

N:1961:CRa

[95] G. R. N. Contemporary reading. Contemporary Physics, 2(3):231–235, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:OIC

[96] Anonymous. Optics at imperial college. Contemporary Physics, 2(3): 235, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:AEI

[97] Anonymous. The 45th annual exhibition of the Institute of Physics and the Physical Society. Contemporary Physics, 2(3):236–237, 1961. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lockett:1961:SMA

[98] H. Lockett. The science masters’ association annual meeting. Contemporary Physics, 2(3):238–240, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:DRG

[99] Anonymous. D.S.I.R. grants and awards, 1961. Contemporary Physics, 2(3):240, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cobbold:1961:RTR

[100] G. B. Cobbold. A review of: “Translation from Russian for scientists”. Contemporary Physics, 2(3):241–244, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mcaulay:1961:SCD

[101] J. Mcaulay. Science in civil defence. Contemporary Physics, 2(4):245– 252, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

King-Hele:1961:SOT

[102] D. G. King-Hele. Satellite orbits and their geophysical implications. Contemporary Physics, 2(4):253–267, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Braddick:1961:IS

[103] H. J. J. Braddick. Interference spectroscopy. Contemporary Physics, 2(4): 268–276, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jennings:1961:M

[104] R. E. Jennings. The microtron. Contemporary Physics, 2(4):277–285, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Birss:1961:EI

[105] R. R. Birss. Electromagnetic induction. Contemporary Physics, 2(4): 286–297, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wenham:1961:PST

[106] E. J. Wenham. Physical science in the training college. Contemporary Physics, 2(4):298–301, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bateman:1961:TPG

[107] P. J. Bateman. Thermoelectric power generation. Contemporary Physics, 2(4):302–311, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1961:NCL

[108] V. T. Saunders. The national chemical laboratory. Contemporary Physics, 2(4):312–317, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:MPE

[109] Anonymous. Mathematics for physicists and engineers. Contemporary Physics, 2(4):318, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:CR

[110] Anonymous. Contemporary reading. Contemporary Physics, 2(4):319– 320, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosser:1961:RCE

[111] W. G. V. Rosser. A review of: “Classical electricity and magnetism”. Contemporary Physics, 2(4):321–324, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1961:BRD

[112] Otto Robert Frisch. Book review: From dualism to unity in physics. By A. Landé (Cambridge University Press, 1960). Pp. xvi + 114. Contemporary Physics, 2(4):323, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blin-Stoyle:1961:PFT

[113] R. J. Blin-Stoyle. Particles, field theory and force. Contemporary Physics, 2(5):325–344, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harper:1961:EFC

[114] W. R. Harper. Electrification following the contact of solids. Contemporary Physics, 2(5):345–359, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vigoureux:1961:GRP

[115] P. Vigoureux. The gyromagnetic ratio of the proton. Contemporary Physics, 2(5):360–366, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1961:SS

[116] J. M. A. Lenihan. The scalpel and the spectroscope. Contemporary Physics, 2(5):367–380, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

N:1961:PGS

[117] G. R. N. Physics for grammar schools. Contemporary Physics, 2(5): 381, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:R

[118] Anonymous. On the record. Contemporary Physics, 2(5):382–384, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Waldron:1961:SRT

[119] M. B. Waldron. A survey of research on the transuranic elements. Contemporary Physics, 2(5):385–402, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

StJoanofArc:1961:TMP

[120] Sister St. Joan of Arc. The teaching of modern physics in schools. Contemporary Physics, 2(5):403–408, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

DeBarr:1961:ASC

[121] A. E. De Barr. The applications of single crystals. Contemporary Physics, 2(6):409–422, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lewis:1961:NAL

[122] G. M. Lewis. Neutrino and antineutrino losses in nuclear processes. Contemporary Physics, 2(6):423–427, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1961:REI

[123] R. D. Davies. Radio emission from interstellar neutral hydrogen. Contemporary Physics, 2(6):428–446, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

VonEngel:1961:NPP

[124] A. Von Engel. The nature and properties of a plasma. Contemporary Physics, 2(6):447–462, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:1961:AVR

[125] G. W. Brown. Accurate voltage reference systems. Contemporary Physics, 2(6):463–471, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chalmers:1961:CBC

[126] J. Alan Chalmers. The correspondence between a current-carrying coil and a magnetic dipole. Contemporary Physics, 2(6):472–474, 1961. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

S:1961:DRE

[127] V. T. S. DSIR research establishments. Contemporary Physics, 2(6): 475–478, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

N:1961:CRb

[128] G. R. N. Contemporary reading. Contemporary Physics, 2(6):479–481, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Powles:1961:ERN

[129] J. G. Powles. Essay review: The nucleus as it usually is. The Principles of Nuclear Magnetism, by A. Abragam. Contemporary Physics, 2(6): 482–483, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:BRa

[130] Anonymous. Book reviews. Contemporary Physics, 2(6):484–488, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:1961:LML

[131] J. W. Martin. Looking at metals: A lecture for schools. Contemporary Physics, 3(1):1–15, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1961:TRP

[132] O. R. Frisch. Time and relativity: Part I. Contemporary Physics, 3 (1):16–27, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosser:1961:ESO

[133] W. G. V. Rosser. Electromagnetism as a second order effect. Paper III. The Biot–Savart law. Contemporary Physics, 3(1):28–44, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jelley:1961:CRO

[134] J. V. Jelley. Čerenkov radiation: Its origin, properties and applications. Contemporary Physics, 3(1):45–57, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parkinson:1961:SAI

[135] D. H. Parkinson. Some aspects of the infra-red absorption spectra of semiconductors. Contemporary Physics, 3(1):58–74, 1961. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1961:WIR

[136] V. T. Saunders. The wool industries research association. Contemporary Physics, 3(1):75–79, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salam:1961:PAI

[137] A. Salam. The proceedings of the 1960 Annual International Conference on high energy physics at Rochester, N.Y. Contemporary Physics, 3 (1):80, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stevens:1961:CA

[138] S. H. G. Stevens. Clean air. Contemporary Physics, 3(2):81–90, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1961:CPW

[139] M. H. Wilson and R. W. Brocklebank. Colour and perception: The work of Edwin Land in the light of current concepts. Contemporary Physics, 3 (2):91–111, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ballentyne:1961:E

[140] D. W. G. Ballentyne. Electroluminescence. Contemporary Physics, 3(2): 112–124, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:DWD

[141] Anonymous. 1. Determination of wood density variation across tree growth rings by means of β-particle absorption. Contemporary Physics, 3(2):125–133, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1961:NPL

[142] V. T. Saunders. The national physical laboratory: Solids at extremely high pressures. Contemporary Physics, 3(2):134–137, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:IPP

[143] Anonymous. The Institute of Physics and the Physical Society conference on nuclear physics. Contemporary Physics, 3(2):138–139, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nightingale:1961:RFI

[144] A. Nightingale. Report on the fourth International Conference and Exhibition on Medical Electronics, New York, 1961. Contemporary Physics, 3(2):140–144, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:RHE

[145] Anonymous. The Rutherford High Energy Laboratory of the National Institute for Research in Nuclear Science. Contemporary Physics, 3(2): 145–149, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbs:1961:ER

[146] D. F. Gibbs. Essay review. Contemporary Physics, 3(2):150–151, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Griffith:1961:BAM

[147] M. V. Griffith. The British Association meeting at Norwich. Contemporary Physics, 3(2):152–154, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Geake:1961:RM

[148] J. E. Geake. Rutherford in Manchester. Contemporary Physics, 3(2): 155–158, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1961:BRb

[149] Anonymous. Book review. Contemporary Physics, 3(2):159–160, 1961. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aitken:1962:PAAa

[150] M. J. Aitken. Physics applied to archaeology — Part I. Contemporary Physics, 3(3):161–176, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:AAP

[151] Anonymous. The American Association of Physics Teachers: Demonstration experiments in physics. Contemporary Physics, 3(3):176, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blow:1962:MAB

[152] D. M. Blow. The molecular approach to biology. Contemporary Physics, 3(3):177–193, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1962:TRP

[153] O. R. Frisch. Time and relativity: Part II. Contemporary Physics, 3(3): 194–201, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pentland:1962:DTC

[154] N. Pentland. The diploma in technology: Courses in applied physics. Contemporary Physics, 3(3):202–211, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:ER

[155] Anonymous. Essay review. Contemporary Physics, 3(3):212–216, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Marsden:1962:ICC

[156] P. L. Marsden. The International Conference on Cosmic Rays and the Earth Storm, Kyoto, Japan. September 3–15, 1961. Contemporary Physics, 3(3):217–221, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lee:1962:KIC

[157] E. W. Lee. Kyoto 1961. The international conference on magnetism. Contemporary Physics, 3(3):222–238, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:AEI

[158] Anonymous. The 46th annual exhibition of the Institute of Physics and the Physical Society — 15–19 January, 1962. Contemporary Physics, 3 (3):239, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:HRS

[159] Anonymous. Harwell Reactor School: Science masters course no. 2. Contemporary Physics, 3(3):240, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pickavance:1962:RTH

[160] T. G. Pickavance. A review of: “Techniques of high energy physics”. Contemporary Physics, 3(3):240, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSa

[161] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 3(4):241–256, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Crook:1962:SPA

[162] A. W. Crook. Some physical aspects of lubrication and wear. Contemporary Physics, 3(4):257–271, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brocklebank:1962:TIC

[163] R. W. Brocklebank. Two international colour meetings. Contemporary Physics, 3(4):272–277, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chalmers:1962:LEE

[164] J. Alan Chalmers. The laws of electromagnetism and an electrostatic analogue. Contemporary Physics, 3(4):278–285, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ramanathan:1962:FLI

[165] K. G. Ramanathan. Faraday’s law of induction and the force on a body due to change in its magnetization in an electric field. Contemporary Physics, 3(4):286–289, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1962:NMM

[166] V. T. Saunders. The NPL “microwave mekometer”. Contemporary Physics, 3(4):290–292, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:TSM

[167] Anonymous. The teaching of science and mathematics. Contemporary Physics, 3(4):292, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1962:DRE

[168] V. T. Saunders. DSIR research establishments: The Building Research Station, Garston. Contemporary Physics, 3(4):293–296, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1962:BAG

[169] V. T. Saunders. The British Association Granada lectures 1961. Contemporary Physics, 3(4):297–298, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paul:1962:MN

[170] E. B. Paul. Masses of nuclei. Contemporary Physics, 3(4):299–302, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:HEI

[171] Anonymous. High energy instrumentation. Contemporary Physics, 3(4): 303–304, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:IS

[172] Anonymous. Isotope separation. Contemporary Physics, 3(4):304–306, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:BRa

[173] Anonymous. Book reviews. Contemporary Physics, 3(4):307–320, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSb

[174] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 3(5):321–333, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aitken:1962:PAAb

[175] M. J. Aitken. Physics applied to archaeology — part II. Contemporary Physics, 3(5):334–352, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hay:1962:ESB

[176] E. J. Hay. Electromagnetism for schools, based on M.K.S.A. units. Contemporary Physics, 3(5):353–367, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haber-Schaim:1962:PSS

[177] Uri Haber-Schaim. The Physical Science Study Committee (4): The working session on physics teaching at the Cavendish Laboratory, Cambridge. Contemporary Physics, 3(5):368–374, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frith:1962:OEA

[178] R. Frith. Ozone in the Earth’s atmosphere. Contemporary Physics, 3(5): 375–381, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Buchdahl:1962:PBP

[179] Gerd Buchdahl. The philosophical basis of physics. Contemporary Physics, 3(5):382–394, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harlow:1962:SMA

[180] D. W. Harlow. The Science Masters’ Association annual meeting, Imperial College, 2nd to 5th January, 1962. Contemporary Physics, 3(5): 395–397, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:BRb

[181] Anonymous. Book reviews. Contemporary Physics, 3(5):398–400, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSc

[182] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 3(6):401–414, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barrell:1962:M

[183] H. Barrell. The metre. Contemporary Physics, 3(6):415–434, 1962. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ingram:1962:L

[184] D. E. J. Ingram. Lasers. Contemporary Physics, 3(6):435–452, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tricker:1962:ACP

[185] R. A. R. Tricker. Ampère as a contemporary physicist. Contemporary Physics, 3(6):453–468, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1962:RIC

[186] V. T. Saunders. The Royal Institution Christmas lectures, 1961. Contemporary Physics, 3(6):469–473, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mendelssohn:1962:SCC

[187] K. Mendelssohn. Sciences in Communist China. Contemporary Physics, 3(6):474–476, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1962:ER

[188] J. G. Wilson. Essay review. Contemporary Physics, 3(6):477–479, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:BRc

[189] Anonymous. Book reviews. Contemporary Physics, 3(6):480, 1962. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSd

[190] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 4(1):1–14, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Comley:1962:UHF

[191] W. J. Comley. Ultra high frequency acoustic waves. Contemporary Physics, 4(1):15–26, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1962:CCT

[192] B. J. Mason. The cloud chamber as a tool in cloud physics. Contemporary Physics, 4(1):27–48, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jephcott:1962:SCP

[193] D. F. Jephcott. The Salzburg conference on plasma physics and controlled nuclear fusion research (4th to 9th September, 1961). Contemporary Physics, 4(1):49–54, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:RSR

[194] Anonymous. The Royal Society and research in schools. Contemporary Physics, 4(1):55–64, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1962:NPL

[195] V. T. Saunders. At the national physical laboratory open days, may 1962. Contemporary Physics, 4(1):65–72, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1962:RM

[196] N. Feather. Rutherford at McGill. Contemporary Physics, 4(1):73–78, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1962:EP

[197] J. M. A. Lenihan. Elementary particles. Contemporary Physics, 4(1): 76–78, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:E

[198] Anonymous. Erratum. Contemporary Physics, 4(1):78, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:BRd

[199] Anonymous. Book reviews. Contemporary Physics, 4(1):79–80, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ziman:1962:EMSe

[200] J. M. Ziman. Electrons in metals: A short guide to the Fermi surface. Contemporary Physics, 4(2):81–99, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:1962:PPS

[201] N. R. Howard. The physical principles of semiconductor power rectifier operation. Contemporary Physics, 4(2):100–112, 1962. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1962:AEM

[202] L. Pearce Williams. Ampère’s electrodynamic molecular model. Contemporary Physics, 4(2):113–123, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mackinnon:1962:UA

[203] L. Mackinnon. Ultrasonic attenuation. Contemporary Physics, 4(2): 124–138, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:1962:FSG

[204] D. H. Martin. Filling the spectroscopic gap between microwaves and the infra-red. Part I. Contemporary Physics, 4(2):139–153, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bright:1962:ERA

[205] A. W. Bright. Essay review: Applied electrostatics. Contemporary Physics, 4(2):154–155, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:SMC

[206] Anonymous. Science master course at the Harwell Reactor School. Contemporary Physics, 4(2):156, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1962:BRe

[207] Anonymous. Book reviews. Contemporary Physics, 4(2):156–160, 1962. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Glaser:1963:TIT

[208] Peter E. Glaser. Thermal imaging techniques for high-temperature research. Contemporary Physics, 4(3):161–175, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roberts:1963:CPU

[209] J. E. Roberts. Codes of practice for the use of ionizing radiations. Contemporary Physics, 4(3):176–186, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:1963:FSG

[210] D. H. Martin. Filling the spectroscopic gap between microwaves and the infra-red. Part II. Contemporary Physics, 4(3):187–205, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beck:1963:PCW

[211] H. V. Beck. Practical class work at the Cavendish Laboratory: A new approach. Contemporary Physics, 4(3):206–220, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parasnis:1963:PSS

[212] A. S. Parasnis. The production of a single slit diffraction pattern by reflection. Contemporary Physics, 4(3):221–222, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:DREa

[213] V. T. Saunders. DSIR research establishments. The British Leather Manufacturers’ Research Association. Contemporary Physics, 4(3):223–227, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:SER

[214] V. T. Saunders. Some exhibits at the Royal Society conversazione, May 1962. Contemporary Physics, 4(3):228–230, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:BRa

[215] Anonymous. Book reviews. Contemporary Physics, 4(3):231–240, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ghosh:1963:RS

[216] S. N. Ghosh and Sharda Nand. Rocket spectroscopy. Contemporary Physics, 4(4):241–255, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Silcox:1963:EMT

[217] J. Silcox. The electron microscopy of thin metal foils. Contemporary Physics, 4(4):256–277, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1963:OSS

[218] W. H. McCrea. The origin of the solar system. Contemporary Physics, 4(4):278–290, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:NPLa

[219] V. T. Saunders. At the national physical laboratory. Contemporary Physics, 4(4):291–292, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:DREb

[220] V. T. Saunders. DSIR research establishments. Contemporary Physics, 4(4):293–298, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harlow:1963:SMA

[221] D. W. Harlow. The Science Masters’ Association annual meeting Manchester College of Science and Technology, 2nd to 5th January, 1963. Contemporary Physics, 4(4):299–303, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:ESU

[222] Anonymous. Emigration of scientists from the United Kingdom. Contemporary Physics, 4(4):304–305, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dolder:1963:ERA

[223] K. Dolder. Essay review: The application of shock waves to research in chemistry and physics. Contemporary Physics, 4(4):306–308, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heywood:1963:SF

[224] Harold Heywood. Solar furnaces. Contemporary Physics, 4(4):308–312, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbs:1963:RPP

[225] D. F. Gibbs. Reports on progress in physics. Contemporary Physics, 4(4): 312–314, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:BRb

Karasz:1963:AC

[227] F. E. Karasz. Adiabatic calorimetry. Contemporary Physics, 4(5):321– 330, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thomas:1963:SAP

[228] D. K. Thomas. Some aspects of the physics of space research. Contemporary Physics, 4(5):331–337, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Powles:1963:NAT

[229] J. G. Powles. Negative absolute temperatures and rotating temperatures. Contemporary Physics, 4(5):338–355, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bacon:1963:SC

[230] D. F. Bacon and W. F. Nash. Spark chambers. Contemporary Physics, 4(5):356–372, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

StJoanofArc:1963:ETP

[231] Sister St. Joan of Arc. An enquiry into the teaching of physical science in girls’ grammar schools. Contemporary Physics, 4(5):373–391, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Geake:1963:PBM

[232] J. E. Geake. Physics at the B.A. in Manchester — 1962. Contemporary Physics, 4(5):392–398, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:BRc

[233] Anonymous. Book reviews. Contemporary Physics, 4(5):399–400, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pain:1963:SWP

[234] H. J. Pain. Shock waves in a plasma. Contemporary Physics, 4(6): 401–421, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mendelssohn:1963:PST

[235] K. A. G. Mendelssohn. Physical science in the twentieth century. Contemporary Physics, 4(6):422–434, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hilsum:1963:SDP

[236] C. Hilsum and P. G. R. King. Some demonstrations of the properties of optical masers. Contemporary Physics, 4(6):435–444, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holton:1963:TPOa

[237] Gerald Holton. On the teaching of physics as one of the basic sciences at university level — Part I: Objectives. Contemporary Physics, 4(6): 445–452, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lyness:1963:EAE

[238] R. C. Lyness. The equivalence of Ampère’s Electrodynamic Law and that of Biot and Savart. Contemporary Physics, 4(6):453–455, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chalmers:1963:TMG

[239] J. Alan Chalmers. Taking the mystery out of Gauss’s Theorem. Contemporary Physics, 4(6):456–458, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sansom:1963:MFA

[240] D. J. Sansom. The magnetic field arising from a steady current flowing in a long straight conductor of rectangular cross-section. Contemporary Physics, 4(6):459–461, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:HFA

[241] V. T. Saunders. “Harwell” — a film about the Atomic Energy Research Establishment. Contemporary Physics, 4(6):462, 1963. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:NCL

[242] V. T. Saunders. At the National Chemical Laboratory open days, September 1962. Contemporary Physics, 4(6):463–467, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:RG

[243] Anonymous. Relativity and gravity. Contemporary Physics, 4(6):468– 472, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:BRd

[244] Anonymous. Book reviews. Contemporary Physics, 4(6):473–480, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holton:1963:TPOb

[245] Gerald Holton, Professor of Physics. On the teaching of physics as one of the basic sciences at university level — part II: Curriculum development. Contemporary Physics, 5(1):1–14, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:1963:NTD

[246] G. Burniston Brown. A new treatment of diffraction. Contemporary Physics, 5(1):15–27, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1963:FDE

[247] L. Pearce Williams. Faraday’s discovery of electromagnetic induction. Contemporary Physics, 5(1):28–37, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caro:1963:PCV

[248] D. E. Caro. Physics courses in Victoria, Australia. Contemporary Physics, 5(1):38–49, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bunton:1963:LFP

[249] Helen Bunton. Looking forward in physics and chemistry teaching. Contemporary Physics, 5(1):50–51, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:QS

[250] Anonymous. The quiet sun. Contemporary Physics, 5(1):52–54, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:M

[251] Anonymous. Magnetism. Contemporary Physics, 5(1):55–58, 1963. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:EE

[252] Anonymous. The Earth’s environment. Contemporary Physics, 5(1): 58–61, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:1963:CLR

[253] R. E. D. Bishop and V. T. Saunders. The Christmas lectures at the Royal Institution on vibration. Contemporary Physics, 5(1):62–63, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:CMW

[254] Anonymous. ‘Communication in the modern world’. Contemporary Physics, 5(1):64, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:NRT

[255] Anonymous. The new radio telescope at the Mullard Observatory at Cambridge. Contemporary Physics, 5(1):65–66, 1963. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:NPLb

[256] V. T. Saunders. At the national physical laboratory open days, may 1963. Contemporary Physics, 5(1):67–70, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:BRe

[257] Anonymous. Book reviews. Contemporary Physics, 5(1):71–80, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schawlow:1963:OIR

[258] Arthur L. Schawlow. Optical and infra-red masers. Contemporary Physics, 5(2):81–102, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thomson:1963:RTP

[259] Sir George Thomson. Research in theory and practice. Contemporary Physics, 5(2):103–111, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hill:1963:DPP

[260] Nora E. Hill. Demonstration of the properties of polarized light using a Michelson interferometer. Contemporary Physics, 5(2):112–113, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:1963:RSD

[261] K. A. Pounds. The Royal Society discussion on satellite Ariel, May 1963. Contemporary Physics, 5(2):114–116, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:NPLc

[262] V. T. Saunders. At the national physical laboratory open days, may, 1963: The use of artificial satellites for relating time at distant points. Contemporary Physics, 5(2):117–119, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:RRR

[263] Anonymous. Recent research reports. Contemporary Physics, 5(2):120– 121, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:DREc

[264] V. T. Saunders. DSIR research establishments: New physics division laboratories at the building research station, Garston. Contemporary Physics, 5(2):122–131, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1963:NPLd

[265] V. T. Saunders. At the national physical laboratory open days, 1963. Contemporary Physics, 5(2):132–136, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:RPI

[266] Anonymous. A review of: “Proceedings of the International Conference on the Ionosphere”. Contemporary Physics, 5(2):137, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rochester:1963:SP

[267] G. D. Rochester. Strange particles. Contemporary Physics, 5(2):138– 140, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Grassie:1963:AD

[268] A. D. C. Grassie. Apparatus drawings. Contemporary Physics, 5(2): 140–143, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1963:CR

[269] G. R. Noakes. Contemporary reading. Contemporary Physics, 5(2): 144–147, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:PFU

[270] Anonymous. Problems facing university physics departments. Contemporary Physics, 5(2):148–150, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1963:BRf

[271] Anonymous. Book reviews. Contemporary Physics, 5(2):151–160, 1963. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Olsen:1964:SMF

[272] J. L. Olsen. Strong magnetic fields. Contemporary Physics, 5(3):161– 179, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bull:1964:FTD

[273] R. M. Bull, W. F. Nash, and B. C. Rastin. Flash tube detectors. Contemporary Physics, 5(3):180–197, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosser:1964:VARa

[274] W. G. V. Rosser. The Van Allen radiation zones I: The theory of trapping and the inner zone. Contemporary Physics, 5(3):198–211, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1964:SRR

[275] J. M. A. Lenihan. The Scottish Research Reactor Centre. Contemporary Physics, 5(3):212–216, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1964:SER

[276] V. T. Saunders. Some exhibits at the Royal Society conversazione, 16th May, 1963. Contemporary Physics, 5(3):217–221, 1964. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1964:POR

[277] V. T. Saunders. Post Office Research Station (Dollis Hill) open days, 12th and 13th September, 1963. Contemporary Physics, 5(3):222–226, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1964:HRS

[278] V. T. Saunders. Hydraulics research station open day, 11th September, 1963. Contemporary Physics, 5(3):227–232, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1964:FST

[279] G. R. Noakes. Films for science teachers. Contemporary Physics, 5(3): 233–234, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1964:BRa

[280] Anonymous. Book reviews. Contemporary Physics, 5(3):234–240, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1964:STEa

[281] Sir Harrie Massey. Space travel and exploration. Contemporary Physics, 5(4):241–254, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosser:1964:VARb

[282] W. G. V. Rosser. The Van Allen radiation zones. Contemporary Physics, 5(4):255–269, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ferguson:1964:IFN

[283] A. T. G. Ferguson. IBIS. A fast neutron time-of-flight system. Contemporary Physics, 5(4):270–279, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

King:1964:RDR

[284] P. G. R. King. Rotation detection with a ‘ring-laser’. Contemporary Physics, 5(4):280–283, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1964:EDQ

[285] G. R. Noakes. Elementary demonstrations: (1) quantitative electrostatics. Contemporary Physics, 5(4):284–291, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chalmers:1964:DEE

[286] J. Alan Chalmers. A difficulty in elementary electromagnetism. Contemporary Physics, 5(4):292, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henderson:1964:DSD

[287] J. C. Henderson. Diffraction studies of dislocations in silicon. Contemporary Physics, 5(4):293–295, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1964:CMW

[288] V. T. Saunders. Communication in the modern world. Contemporary Physics, 5(4):296–297, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harlow:1964:ASE

[289] D. W. Harlow. The association for science education annual meeting. Contemporary Physics, 5(4):298–301, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Oldham:1964:URC

[290] Frank Oldham. University reactions to changes in school physics teaching. Contemporary Physics, 5(4):302–303, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burcham:1964:RMC

[291] W. E. Burcham. Rutherford at Manchester, 1907–19 The Collected Papers of Lord Rutherford of Nelson, Vol. 2. (London: George Allen & Unwin, 1963.) [Pp. 590] 105s. Contemporary Physics, 5(4):304–308, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jennings:1964:IPR

[292] John C. Jennings, M.P. From inside Parliament (1): The Robbins report. Contemporary Physics, 5(4):308–310, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Albu:1964:IPR

[293] Austen Albu, M.P. From inside Parliament (2): The Robbins report and the Trend report. Contemporary Physics, 5(4):311–313, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1964:BRb

[294] Anonymous. Book reviews. Contemporary Physics, 5(4):314–320, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hume-Rothery:1964:SNS

[295] W. Hume-Rothery. Some notes on the structures of alloys of iron. Contemporary Physics, 5(5):321–347, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burhop:1964:PEP

[296] E. H. S. Burhop. The physics of elementary particles. Contemporary Physics, 5(5):348–358, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rowlinson:1964:TLE

[297] J. S. Rowlinson. The theory of liquids at equilibrium. Contemporary Physics, 5(5):359–376, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Felici:1964:CEP

[298] N. J. Felici. Contemporary electrostatics — i: Physical background and applications. Contemporary Physics, 5(5):377–390, 1964. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1964:PSA

[299] V. T. Saunders. Particle size analysis. Contemporary Physics, 5(5): 391–393, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Llowarch:1964:SM

[300] W. Llowarch. Schlieren methods. Contemporary Physics, 5(5):394–397, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1964:BRc

[301] Anonymous. Book reviews. Contemporary Physics, 5(5):398–400, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1964:TQP

[302] Sir Nevill Mott. On teaching quantum phenomena. Contemporary Physics, 5(6):401–418, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Felici:1964:CEI

[303] N. J. Felici. Contemporary electrostatics — II: Ionization charging and insulation. Contemporary Physics, 5(6):419–437, 1964. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Suddaby:1964:TPL

[304] Arthur Suddaby. Transport processes in liquids. Contemporary Physics, 5(6):438–456, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1964:STEb

[305] Sir Harrie Massey. Space travel and exploration. Contemporary Physics, 5(6):457–471, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1964:CR

[306] J. M. A. Lenihan. Contemporary reading. Contemporary Physics, 5(6): 472–475, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bolton:1964:ASE

[307] W. Bolton. The association for science education: Physics teaching aids. Contemporary Physics, 5(6):475, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1964:AEI

[308] J. M. A. Lenihan. The 48th annual exhibition of the Institute of Physics and the Physical Society, January 1964. Contemporary Physics, 5(6):

476–477, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1964:BRd

[309] Anonymous. Book reviews. Contemporary Physics, 5(6):478–480, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wright:1964:NLO

[310] J. K. Wright. Non-linear optics. Contemporary Physics, 6(1):1–14, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vigoureux:1964:EUA

[311] P. Vigoureux. Electric units and atomic constants. Contemporary Physics, 6(1):15–25, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1964:STEc

[312] Sir Harrie Massey. Space travel and exploration. Contemporary Physics, 6(1):26–48, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1964:MEC

[313] J. M. A. Lenihan. Macmillan–Eothen ‘Cinette’ film loops. Contemporary Physics, 6(1):49, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleton:1964:SXR

[314] A. Appleton. The soft X-ray emission spectra of metals and alloys. Contemporary Physics, 6(1):50–67, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parry:1964:EST

[315] L. G. Parry. Elementary science teaching in New South Wales. Contemporary Physics, 6(1):68–72, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keohane:1964:ERR

[316] K. W. Keohane. Essay review: Reports on progress in physics, vol. XXVII, 1964. Contemporary Physics, 6(1):73–75, 1964. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haworth:1964:RPP

[317] C. W. Haworth. Reports on progress in physics, vol. XXVI, 1963. Contemporary Physics, 6(1):75–76, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1964:BRe

[318] Anonymous. Book reviews. Contemporary Physics, 6(1):77–80, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1964:NNR

[319] O. R. Frisch. Nuclei and nuclear reactions. Contemporary Physics, 6 (2):81–93, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Postnikov:1964:FM

[320] S. N. Postnikov. The friction of metals. Contemporary Physics, 6(2): 94–111, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tabor:1964:LSP

[321] David Tabor. Large-scale properties of matter. Contemporary Physics, 6(2):112–128, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:1964:DQT

[322] R. E. Peierls. The development of quantum theory. Part I. Formulation and interpretation. Contemporary Physics, 6(2):129–139, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Arndt:1964:AL

[323] U. W. Arndt. Automation in the laboratory. Contemporary Physics, 6 (2):140–146, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cockcroft:1964:RGS

[324] Sir John Cockcroft. The role of government in scientific research and development. Contemporary Physics, 6(2):147–151, 1964. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heavens:1964:QE

[325] O. S. Heavens. Quantum electronics. Contemporary Physics, 6(2):152– 154, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1964:BRI

[326] O. R. Frisch. Book review: An introduction to the theory of relativity, by A. V. Rosser. Contemporary Physics, 6(2):155–156, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1964:BRf

[327] Anonymous. Book reviews. Contemporary Physics, 6(2):157–160, 1964. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bragg:1965:HXR

[328] Sir Lawrence Bragg. The history of X-ray analysis. Contemporary Physics, 6(3):161–171, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1965:STEa

[329] Sir Harrie Massey. Space travel and exploration. Contemporary Physics, 6(3):172–191, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:1965:DQT

[330] R. E. Peierls. The development of quantum theory. Contemporary Physics, 6(3):192–205, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hartley:1965:SWS

[331] Sir Harold Hartley. The scientific work of Sir Humphry Davy. Contemporary Physics, 6(3):206–216, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mendoza:1965:W

[332] E. Mendoza. Waves. Contemporary Physics, 6(3):217–224, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tye:1965:TCR

[333] R. P. Tye. Thermal conductivity: Recent developments. Contemporary Physics, 6(3):225–229, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1965:RRR

[334] V. T. Saunders. Radio research: A report on the 1964 open days at the radio research station, Ditton Park, Slough. Contemporary Physics, 6(3): 230–236, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1965:CR

[335] J. M. A. Lenihan. Contemporary reading. Contemporary Physics, 6(3): 237–239, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:BRa

[336] Anonymous. Book reviews. Contemporary Physics, 6(3):240, 1965. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1965:STEb

[337] Sir Harrie Massey. Space travel and exploration V. Lunar and planetary probes — meteorology and communications — geodesy and navigation. Contemporary Physics, 6(4):241–260, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lee:1965:NIC

[338] E. W. Lee. Nottingham 1964. The international conference on magnetism. Contemporary Physics, 6(4):261–273, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Egelstaff:1965:STMa

[339] P. A. Egelstaff and P. Schofield. The structure and thermal motion of simple liquids — I. Contemporary Physics, 6(4):274–284, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrence:1965:GUI

[340] A. W. Lawrence. Gyroscopes used in inertial navigation. Contemporary Physics, 6(4):285–298, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harlow:1965:ASE

[341] D. W. Harlow. The association for science education annual meeting. Contemporary Physics, 6(4):299–302, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:NCC

[342] Anonymous. A new concept in courses for teachers of science. Contemporary Physics, 6(4):303, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:BRb

[343] Anonymous. Book reviews. Contemporary Physics, 6(4):304–320, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1965:STEc

[344] Sir Harrie Massey. Space travel and exploration. Contemporary Physics, 6(5):321–337, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

TerHaar:1965:OCR

[345] D. Ter Haar. On the origin of cosmic rays. Contemporary Physics, 6(5): 338–348, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morley:1965:CGG

[346] J. G. Morley. Crystal growth in glass. Contemporary Physics, 6(5): 349–364, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tucker:1965:RPP

[347] J. W. Tucker. Relaxation processes in paramagnetic crystals. Contemporary Physics, 6(5):365–380, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1965:RTS

[348] V. T. Saunders. Road traffic and safety. Contemporary Physics, 6(5): 381–384, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1965:BR

[349] V. T. Saunders. Building research. Contemporary Physics, 6(5):385– 389, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1965:SER

[350] V. T. Saunders. Some exhibits at the Royal Society conversazione, 1964. Contemporary Physics, 6(5):390–395, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:BRc

[351] Anonymous. Book reviews. Contemporary Physics, 6(5):396–400, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palmer:1965:HRS

[352] H. P. Palmer. High resolution studies in radio astronomy. Contemporary Physics, 6(6):401–420, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frost:1965:ARF

[353] Albert D. Frost. Astronomy at radio frequencies. Contemporary Physics, 6(6):421–435, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kibble:1965:EPS

[354] T. W. B. Kibble. Elementary particle symmetries. Contemporary Physics, 6(6):436–452, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Egelstaff:1965:STMb

[355] P. A. Egelstaff and P. Schofield. Structure and thermal motion of simple liquids–II. Contemporary Physics, 6(6):453–464, 1965. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

TerHaar:1965:WP

[356] D. Ter Haar. Wolfgang Pauli 1900–1958. Contemporary Physics, 6(6): 465–471, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:BRd

[357] Anonymous. Book reviews. Contemporary Physics, 6(6):472–480, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:OCS

[358] Anonymous. The organization of civil science in the United Kingdom. Contemporary Physics, 7(1):1–12, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ingram:1965:NMRa

[359] D. J. E. Ingram. Nuclear magnetic resonance (part i). Contemporary Physics, 7(1):13–28, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1965:PBNa

[360] N. Feather. The physical basis of Newtonian mechanics (Part I). Contemporary Physics, 7(1):29–44, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1965:TPL

[361] O. R. Frisch. Take a photon …. Contemporary Physics, 7(1):45–53, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilkinson:1965:NST

[362] W. J. Wilkinson. The new science teaching as seen in terms of educational and psychological theory. Contemporary Physics, 7(1):54–56, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chalmers:1965:LSI

[363] J. Alan Chalmers. A logical sequence of ideas in electromagnetism. Contemporary Physics, 7(1):57–58, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1965:BPN

[364] V. T. Saunders. Basic physics at the national physical laboratory. Contemporary Physics, 7(1):59–65, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1965:NPL

[365] V. T. Saunders. At the national physical laboratory open days, 19th and 20th may, 1965. Contemporary Physics, 7(1):66–69, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blin-Stoyle:1965:CIP

[366] R. J. Blin-Stoyle. Congrès international de physique nucléaire (Paris). July 1964. volumes I and II. Contemporary Physics, 7(1):70–72, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:BRe

Grassie:1965:IBD

[368] A. D. C. Grassie, J. E. Spice, and A. Gerrard. Introducing the Boltzmann distribution. Contemporary Physics, 7(2):81–102, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ingram:1965:NMRb

[369] D. J. E. Ingram. Nuclear magnetic resonance (part II). Contemporary Physics, 7(2):103–121, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1965:PBNb

[370] N. Feather. The physical basis of Newtonian mechanics. Part II. Contemporary Physics, 7(2):122–134, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rose-Innes:1965:NS

[371] A. C. Rose-Innes. The new superconductors. Contemporary Physics, 7 (2):135–151, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:1965:PSS

[372] R. Brown and D. L. Gardiner. Physics at school. A small-scale investigation. Contemporary Physics, 7(2):152–157, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1965:BRf

[373] Anonymous. Book reviews. Contemporary Physics, 7(2):157–160, 1965. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Halligey:1966:EME

[374] P. Halligey. The evolution of MKS electrical units: Effects on the teaching of electricity. Contemporary Physics, 7(3):161–173, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Austin:1966:SUV

[375] I. G. Austin. Solids under very high pressure. Contemporary Physics, 7(3):174–191, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1966:EPT

[376] G. A. Saunders. The electron pair theory of superconductivity. Contemporary Physics, 7(3):192–209, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bolton:1966:TP

[377] W. Bolton. Teaching photoelectricity. Contemporary Physics, 7(3):210– 216, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Powell:1966:MTM

[378] R. F. Powell. Modern techniques for the measurement of the thermal conductivity of metals. Contemporary Physics, 7(3):217–227, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:RPP

[379] Anonymous. Reports on progress in physics, vol. XXVIII, 1965. Contemporary Physics, 7(3):228–229, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:CPL

[380] Anonymous. The collected papers of Lord Rutherford of Nelson. Vol. III — The Cavendish Laboratory. Contemporary Physics, 7(3):230–231, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:CS

[381] Anonymous. The crystalline state. Contemporary Physics, 7(3):232– 233, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weinstock:1966:IBD

[382] Robert Weinstock. On introducing the Boltzmann distribution. Contemporary Physics, 7(3):234, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:BRa

Hoyaux:1966:PPA

[384] M. F. Hoyaux. Plasma physics and its applications. Contemporary Physics, 7(4):241–260, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Odier:1966:SPM

[385] Marc Odier. A simple photographic method for studying motion. Contemporary Physics, 7(4):261–277, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lomer:1966:NSSa

[386] W. M. Lomer. Neutron spectroscopy of solids — Part I. Contemporary Physics, 7(4):278–293, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1966:FIR

[387] N. Feather. ‘Factors of ignorance’ and the rationalized M.K.S.A. system of units. Contemporary Physics, 7(4):294–303, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1966:SM

[388] L. F. Bates. The study of magnetism. Contemporary Physics, 7(4): 304–307, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Oldham:1966:EPP

[389] F. Oldham. The education of professional physicists. Contemporary Physics, 7(4):307–309, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:MSU

[390] Anonymous. The metric system and S.I. units. Contemporary Physics, 7(4):309–312, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1966:ARA

[391] G. R. Noakes. Annual report of the Advisory Council on Scientific Policy 1963–64. Contemporary Physics, 7(4):313, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:BRb

[392] Anonymous. Book reviews. Contemporary Physics, 7(4):314–320, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rawlins:1966:PPF

[393] F. I. G. Rawlins. The place of physics in the fine arts. Contemporary Physics, 7(5):321–330, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mendelssohn:1966:DRI

[394] K. Mendelssohn. Dewar at the Royal Institution. Contemporary Physics, 7(5):331–342, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Essen:1966:MTF

[395] L. Essen. The measurement of time and frequency. Contemporary Physics, 7(5):343–349, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vigoureux:1966:EOS

[396] P. Vigoureux. Errors of observation and systematic errors. Contemporary Physics, 7(5):350–357, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bragg:1966:RIL

[397] Sir Lawrence Bragg. The Royal Institution lectures in science for members of the administrative class of the civil service. Contemporary Physics, 7(5):358–361, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Oatley:1966:MEO

[398] C. W. Oatley. Mechanical and electrical oscillations; transient and alternating electric currents. Contemporary Physics, 7(5):362–391, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Macfadyen:1966:PMS

[399] K. A. Macfadyen. The physics of musical sounds. Contemporary Physics, 7(5):392–394, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:BRc

[400] Anonymous. Book reviews. Contemporary Physics, 7(5):395–400, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lomer:1966:NSSb

[401] W. M. Lomer. Neutron spectroscopy of solids — part II. Contemporary Physics, 7(6):401–418, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sandars:1966:MC

[402] P. G. H. Sandars. Magnetic charge. Contemporary Physics, 7(6):419– 429, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hall:1966:MFI

[403] S. H. Hall. The magnetic field of the ionospheric dynamo. Contemporary Physics, 7(6):430–439, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vigoureux:1966:RP

[404] P. Vigoureux. Radiation pressure. Contemporary Physics, 7(6):440– 446, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

TerHaar:1966:AIA

[405] D. Ter Haar. Adiabatic invariance and adiabatic invariants. Contemporary Physics, 7(6):447–458, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1966:NPL

[406] V. T. Saunders. At the national physical laboratory open days, 1966. Contemporary Physics, 7(6):459–467, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chalmers:1966:DBC

[407] J. Alan Chalmers. The difference between a current-carrying coil and a permanent magnet. Contemporary Physics, 7(6):468–469, 1966. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1966:BRd

Cottrell:1967:SMF

[409] A. H. Cottrell. Strength of materials: A foreword introducing the contemporary physics series. Contemporary Physics, 8(1):1–3, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Alexander:1967:SER

[410] W. O. Alexander. The simple economics of rivalries between materials. Contemporary Physics, 8(1):5–19, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kurti:1967:SC

[411] N. Kurti. Spins and cryogenics. Contemporary Physics, 8(1):21–40, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vick:1967:EAE

[412] F. A. Vick, O.B.E. Explorations in atomic energy. Contemporary Physics, 8(1):41–61, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1967:MOT

[413] B. J. Mason, D.Sc., F.R.S. The meteorological office — today and tomorrow. Contemporary Physics, 8(1):63–85, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hutchinson:1967:AME

[414] W. C. A. Hutchinson. Angular momentum — an experiment to illustrate its conservation. Contemporary Physics, 8(1):87–93, 1967. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nash:1967:PNI

[415] W. F. Nash. Proceedings of the ninth international conference on cosmic rays. Contemporary Physics, 8(1):95–100, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnes:1967:HEP

[416] K. J. Barnes. High energy physics and elementary particles. Contemporary Physics, 8(1):101–103, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:RPP

[417] Anonymous. Reports on progress in physics, vol. XXIX. Contemporary Physics, 8(1):105–107, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:BRa

[418] Anonymous. Book reviews. Contemporary Physics, 8(1):108–112, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:Ea

[419] Anonymous. Editorial. Contemporary Physics, 8(1):113, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Biggs:1967:PME

[420] W. D. Biggs. Properties of materials: The engineering designer’s point of view. Contemporary Physics, 8(2):113–126, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Littler:1967:EH

[421] T. S. Littler. The ear and hearing. Contemporary Physics, 8(2):127– 151, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ennos:1967:HA

[422] A. E. Ennos. Holography and its applications. Contemporary Physics, 8(2):153–170, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cartwright:1967:MSW

[423] D. E. Cartwright. Modern studies of wind-generated ocean waves. Contemporary Physics, 8(2):171–183, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trevena:1967:BLU

[424] D. H. Trevena. The behaviour of liquids under tension. Contemporary Physics, 8(2):185–195, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mansfield:1967:AAD

[425] R. Mansfield. Application of the Argand diagram to magnetoconductivity effects. Contemporary Physics, 8(2):197–202, 1967. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:BRb

[426] Anonymous. Book reviews. Contemporary Physics, 8(2):203–212, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:Eb

[427] Anonymous. Erratum. Contemporary Physics, 8(2):212, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Broom:1967:CFE

[428] T. Broom. Creep and fatigue of engineering materials. Contemporary Physics, 8(3):213–233, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parkinson:1967:HMF

[429] D. H. Parkinson. High magnetic fields. Contemporary Physics, 8(3): 235–249, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1967:ADA

[430] A. H. Cook. The absolute determination of the acceleration due to gravity. Contemporary Physics, 8(3):251–266, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hinde:1967:M

[431] B. J. Hinde and D. I. Gaunt. Microseisms. Contemporary Physics, 8(3): 267–283, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Malinowski:1967:PPH

[432] J. Malinowski. Properties of photoexcited holes in silver bromide crystals. Contemporary Physics, 8(3):285–301, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:BRc

[433] Anonymous. Book reviews. Contemporary Physics, 8(3):303–312, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kelly:1967:HSM

[434] A. Kelly. High strength materials. Contemporary Physics, 8(4):313– 329, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kopal:1967:REM

[435] Zdeněk Kopal. Recent explorations of the Moon by spacecraft. Contemporary Physics, 8(4):331–356, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Longair:1967:QSR

[436] M. S. Longair. Quasi-stellar radio sources. Contemporary Physics, 8(4): 357–372, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Merrison:1967:NGE

[437] A. W. Merrison. NINA — the 4 GeV electron synchrotron of the Science Research Council. Contemporary Physics, 8(4):373–384, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mee:1967:RMM

[438] C. D. Mee. Recent measurements of the magneto-optical properties of some garnets. Contemporary Physics, 8(4):385–400, 1967. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tredgold:1967:SCC

[439] R. H. Tredgold. Space-charge conduction. Contemporary Physics, 8(4): 401–411, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1967:ERE

[440] L. F. Bates. Essay review: An essay in exposition. Magnetism in Solids, by D. H. Martin. Contemporary Physics, 8(4):413–415, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lee:1967:PE

[441] E. W. Lee. The physics exhibition 1967. Contemporary Physics, 8(4): 417–420, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pike:1967:FPS

[442] E. R. Pike. The French Physical Society’s 1967 exhibition. Contemporary Physics, 8(4):420–425, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:BRd

[443] Anonymous. Book reviews. Contemporary Physics, 8(4):427–432, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wells:1967:FED

[444] A. A. Wells. Fracture in engineering design. Contemporary Physics, 8(5): 433–445, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hall:1967:MMF

[445] S. H. Hall. Measuring magnetic fields in space. Contemporary Physics, 8(5):447–458, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cracknell:1967:MG

[446] Arthur P. Cracknell. Magnetic groups. Contemporary Physics, 8(5): 459–473, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Popplewell:1967:STC

[447] R. J. L. Popplewell. Some topics of current interest in infrared spectroscopy. Contemporary Physics, 8(5):475–487, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kaischew:1967:SPE

[448] R. Kaischew and E. Budevski. Surface processes in electrocrystallization. Contemporary Physics, 8(5):489–516, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1967:BRe

[449] Anonymous. Book reviews. Contemporary Physics, 8(5):517–528, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbs:1967:EPM

[450] H. G. Gibbs and D. W. Saunders. Engineering with polymeric materials. Contemporary Physics, 8(6):529–560, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chamberlain:1967:RAV

[451] A. C. Chamberlain. Radioactive aerosols and vapours. Contemporary Physics, 8(6):561–581, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cusack:1967:PLM

[452] N. E. Cusack. The physics of liquid metals. Contemporary Physics, 8(6): 583–606, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dearnaley:1967:SSR

[453] G. Dearnaley. Solid-state radiation detectors. Contemporary Physics, 8 (6):607–626, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenwood:1967:BR

[454] D. A. Greenwood. Book reviews. Contemporary Physics, 8(6):627–628, 1967. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thomson:1968:EHE

[455] Sir George Thomson. The early history of electron diffraction. Contemporary Physics, 9(1):1–15, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thirlaway:1968:DUE

[456] H. I. S. Thirlaway. Diagnosing underground explosions and earthquakes. Contemporary Physics, 9(1):17–31, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McGill:1968:ASR

[457] N. C. McGill. The apparent shape of rapidly moving objects in special relativity. Contemporary Physics, 9(1):33–48, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stedman:1968:GF

[458] G. E. Stedman. Green’s functions. Contemporary Physics, 9(1):49–69, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Crossland:1968:FEW

[459] B. Crossland and A. S. Bahrani. Fundamentals, of explosive welding. Contemporary Physics, 9(1):71–87, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRa

[460] Anonymous. Book reviews. Contemporary Physics, 9(1):89–94, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRb

[461] Anonymous. Books received. Contemporary Physics, 9(1):95, 1968. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Betts:1968:HIR

[462] David S. Betts. Helium isotope refrigeration. Contemporary Physics, 9 (2):97–114, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hopkins:1968:UHV

[463] B. J. Hopkins. Ultra-high vacua. Contemporary Physics, 9(2):115–135, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1968:ADI

[464] L. F. Bates. Anisotropy, defects, impurities and ripple a short report on the international congress on magnetism, 1967. Contemporary Physics, 9(2):137–143, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

VanDorn:1968:T

[465] W. G. Van Dorn. Tsunamis. Contemporary Physics, 9(2):145–164, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hoyaux:1968:PPS

[466] M. F. Hoyaux. Plasma phenomena and the solid state. Contemporary Physics, 9(2):165–196, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRc

[467] Anonymous. Book reviews. Contemporary Physics, 9(2):197–202, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barrell:1968:SHM

[468] H. Barrell. A short history of measurement standards at the national physical laboratory. Contemporary Physics, 9(3):205–226, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1968:NDC

[469] A. H. Cook. A new determination of the constant of gravitation. Contemporary Physics, 9(3):227–238, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1968:OAM

[470] S. F. Mason. Optical activity and molecular dissymmetry. Contemporary Physics, 9(3):239–256, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palmer:1968:DG

[471] E. W. Palmer and J. F. Verrill. Diffraction gratings. Contemporary Physics, 9(3):257–276, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bewilogua:1968:CRG

[472] Ludwig Bewilogua and Christel Gladun. Condensed rare gases. Contemporary Physics, 9(3):277–300, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRd

[473] Anonymous. Book reviews. Contemporary Physics, 9(3):301–308, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRe

[474] Anonymous. Books received. Contemporary Physics, 9(3):309, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Littler:1968:PAP

[475] T. S. Littler. The physical and analytical properties of the ear. Contemporary Physics, 9(4):311–332, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cosslett:1968:HVE

[476] V. E. Cosslett. The high-voltage electron microscope. Contemporary Physics, 9(4):333–354, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ginzburg:1968:PHT

[477] V. L. Ginzburg. The problem of high temperature superconductivity. Contemporary Physics, 9(4):355–374, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1968:CPC

[478] M. W. Thompson. The channelling of particles in crystals. Contemporary Physics, 9(4):375–398, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Baker:1968:APC

[479] John R. Baker. American physics curriculum projects. Contemporary Physics, 9(4):399–415, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRf

[480] Anonymous. Book reviews. Contemporary Physics, 9(4):417–421, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRg

[481] Anonymous. Books received. Contemporary Physics, 9(4):422, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chesterman:1968:GEO

[482] W. Deryck Chesterman. Geophysical exploration of the ocean floor using acoustic methods. Contemporary Physics, 9(5):423–446, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Watkins:1968:SM

[483] M. T. Watkins. The shaping of metals. Contemporary Physics, 9(5): 447–474, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leaver:1968:TFF

[484] K. D. Leaver. Thin ferromagnetic films. Contemporary Physics, 9(5): 475–499, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lee:1968:PE

[485] E. W. Lee. The physics exhibition, 1968. Contemporary Physics, 9(5): 501–504, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tricker:1968:EM

[486] R. A. R. Tricker. Electricity and matter. Contemporary Physics, 9(5): 505–508, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Buchdahl:1968:NP

[487] Gerd Buchdahl. The nature of physics. Contemporary Physics, 9(5): 509–510, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRh

[488] Anonymous. Book reviews. Contemporary Physics, 9(5):511–518, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRi

[489] Anonymous. Book received. Contemporary Physics, 9(5):519, 1968. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

OConnor:1968:ME

[490] D. A. O’Connor. The Mössbauer effect. Contemporary Physics, 9(6): 521–535, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trevena:1968:EL

[491] D. H. Trevena. Elastic liquids. Contemporary Physics, 9(6):537–548, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1968:MTS

[492] A. W. B. Taylor. The microscopic theory of superconductivity. Contemporary Physics, 9(6):549–564, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1968:PSF

[493] D. Taylor Smith. Physics and sea-floor minerals. Contemporary Physics, 9(6):565–585, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Skinner:1968:EP

[494] N. J. Skinner. The electrojet phenomenon. Contemporary Physics, 9(6): 587–602, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fellgett:1968:FIR

[495] P. B. Fellgett. Far infra-red spectrology. Contemporary Physics, 9(6): 603–604, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRj

[496] Anonymous. Book reviews. Contemporary Physics, 9(6):605–614, 1968. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1968:BRk

[497] Anonymous. Book received. Contemporary Physics, 9(6):615, 1968. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Adams:1969:PTN

[498] J. B. Adams. Progress towards a nuclear fusion reactor. Contemporary Physics, 10(1):1–20, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Niewodniczanski:1969:INP

[499] H. Niewodniczański. The Institute of Nuclear Physics, Cracow, and its recent investigations in nuclear reactions. Contemporary Physics, 10(1): 21–47, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kumar:1969:CLM

[500] Dr. Rajendra Kumar. Clusters in liquid metals. Contemporary Physics, 10(1):49–58, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Walton:1969:KLS

[501] Alan J. Walton and Alan G. Woodruff. A kinetic liquid-simulator. Contemporary Physics, 10(1):59–70, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nixon:1969:SEM

[502] W. C. Nixon. Scanning electron microscopy. Contemporary Physics, 10 (1):71–96, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRa

[503] Anonymous. Book reviews. Contemporary Physics, 10(1):97–102, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRb

[504] Anonymous. Books received. Contemporary Physics, 10(1):103–104, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davidge:1969:MPC

[505] R. W. Davidge. Mechanical properties of ceramic materials. Contemporary Physics, 10(2):105–124, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1969:CSN

[506] N. F. Mott. Conduction and switching in non-crystalline materials. Contemporary Physics, 10(2):125–138, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Petley:1969:JE

[507] B. W. Petley. The Josephson effects. Contemporary Physics, 10(2): 139–158, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1969:MU

[508] R. G. F. Taylor and A. J. Pointon. Microwave ultrasonics. Contemporary Physics, 10(2):159–178, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fremlin:1969:RLP

[509] J. H. Fremlin. Relativity and the lever problem. Contemporary Physics, 10(2):179–180, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Walton:1969:APG

[510] Alan J. Walton. Archimedes’ principle in gases. Contemporary Physics, 10(2):181–186, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tricker:1969:CER

[511] R. A. R. Tricker. Classical electromagnetism and relativity. Contemporary Physics, 10(2):187–189, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRc

[512] Anonymous. Book reviews. Contemporary Physics, 10(2):191–203, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRd

Parsons:1969:ESC

[514] Roger Parsons. Electrolysis and simple cells. Contemporary Physics, 10 (3):205–220, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hill:1969:ECD

[515] R. M. Hill. Electrical conduction in discontinuous metal films. Contemporary Physics, 10(3):221–240, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Betts:1969:QF

[516] David S. Betts. Quantum fluids. Contemporary Physics, 10(3):241– 262, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1969:FTA

[517] L. Davies and J. L. Wilson. Flight through the air at very high speeds. Contemporary Physics, 10(3):263–276, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Large:1969:IINa

[518] L. N. Large. Ion implantation: A new method of doping semiconductors — I. Contemporary Physics, 10(3):277–298, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRe

[519] Anonymous. Book reviews. Contemporary Physics, 10(3):299–302, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRf

[520] Anonymous. Books received. Contemporary Physics, 10(3):303–304, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1969:CPP

[521] B. L. Smith. Critical point phenomena. Contemporary Physics, 10(4): 305–329, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bright:1969:MEG

[522] A. W. Bright and B. Makin. Modern electrostatic generators. Contemporary Physics, 10(4):331–353, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brandt:1969:SHP

[523] N. B. Brandt and N. I. Ginzburg. Superconductivity at high pressures. Contemporary Physics, 10(4):355–386, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mayne:1969:PEB

[524] K. I. Mayne. Polarized electron beams. Contemporary Physics, 10(4): 387–412, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ferrier:1969:ZPR

[525] W. G. Ferrier. The zone plate and its role in holography. Contemporary Physics, 10(4):413–436, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Melville:1969:PE

[526] D. Melville. The physics exhibition, 1969. Contemporary Physics, 10(4): 437–442, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1969:BR

[527] B. L. Smith. Book reviews. Contemporary Physics, 10(4):443–447, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRg

[528] Anonymous. Books received. Contemporary Physics, 10(4):448, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scott:1969:PMS

[529] Ronald F. Scott. The physics and mechanics of soil. Contemporary Physics, 10(5):449–472, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Turnbull:1969:UWC

[530] David Turnbull. Under what conditions can a glass be formed? Contemporary Physics, 10(5):473–488, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Walton:1969:BST

[531] Alan J. Walton. Bulk and surface transport in liquids — an introductory treatment. Contemporary Physics, 10(5):489–504, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Large:1969:IINb

[532] L. N. Large. Ion implantation: A new method of doping semiconductors — II. Contemporary Physics, 10(5):505–531, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRh

[533] Anonymous. Book reviews. Contemporary Physics, 10(5):533–534, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRi

[534] Anonymous. Books received. Contemporary Physics, 10(5):535–536, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khaikin:1969:MSL

[535] M. S. Khaikin and M. Kumar. Magnetic surface levels. Contemporary Physics, 10(6):537–557, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Webster:1969:HA

[536] Peter J. Webster. Heusler alloys. Contemporary Physics, 10(6):559– 577, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Powell:1969:TCR

[537] R. W. Powell. Thermal conductivity: A review of some important developments. Contemporary Physics, 10(6):579–600, 1969. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trevena:1969:UWL

[538] D. H. Trevena. Ultrasonic waves in liquids. Contemporary Physics, 10(6): 601–614, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Odier:1969:NTM

[539] Marc Odier. Neoncinegraphy in the teaching of mechanics. Contemporary Physics, 10(6):615–626, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stedman:1969:BR

[540] G. E. Stedman. Book reviews. Contemporary Physics, 10(6):627–634, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1969:BRj

[541] Anonymous. Books received. Contemporary Physics, 10(6):635–636, 1969. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pamplin:1970:NDC

[542] Brian R. Pamplin. Negative differential conductivity effects in semiconductors. Contemporary Physics, 11(1):1–19, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Simmons:1970:AAS

[543] J. G. Simmons. Applications of amorphous semiconductors in electronic devices. Contemporary Physics, 11(1):21–41, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Meredith:1970:FP

[544] R. Meredith. Fibrous polymers. Contemporary Physics, 11(1):43–58, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Strnad:1970:NTN

[545] J. Strnad. A note on the Trouton–Noble experiment. Contemporary Physics, 11(1):59–64, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Craik:1970:MDS

[546] D. J. Craik. The magnetic domain structures of small crystals. Contemporary Physics, 11(1):65–97, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Loudon:1970:BR

[547] R. Loudon. Book reviews. Contemporary Physics, 11(1):99–103, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRa

[548] Anonymous. Books received. Contemporary Physics, 11(1):104, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1970:RI

[549] D. R. Bates. Reactions in the ionosphere. Contemporary Physics, 11(2): 105–124, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1970:IGC

[550] B. L. Smith. Inert gas crystals. Contemporary Physics, 11(2):125–149, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Markland:1970:F

[551] E. Markland. Fluidics. Contemporary Physics, 11(2):151–172, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morgan:1970:CPT

[552] D. Vernon Morgan and Dirck Van Vliet. Charged particle tracks in solids. Contemporary Physics, 11(2):173–193, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Edwards:1970:RTX

[553] S. J. Edwards. A review of the theories of X-ray satellite lines. Contemporary Physics, 11(2):195–200, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1970:NDC

[554] Professor L. F. Bates. A new departure in conference reports. Conference Booklet No. 1. High Magnetic Fields and their Applications: Nottingham 1969. (The Institute of Physics and The Physical Society, 1969.) [Pp. vi + 168] 441, 5s.Od. Contemporary Physics, 11(2):201–203, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Popplewell:1970:EVS

[555] Richard J. L. Popplewell. Electronic and vibrational spectra of molecules. Edited by V. Skobel’tsyn. (Consultants Bureau (Plenum Press), 1968.)

[Pp. vii + 217.] $22.50. Scope: Library. Level: Specialist/postgraduate. Contemporary Physics, 11(2):205–207, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRb

[556] Anonymous. Book received. Contemporary Physics, 11(2):208, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Southworth:1970:FIM

[557] H. N. Southworth. Field-ion microscopy. Contemporary Physics, 11(3): 209–226, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Owen:1970:SGPa

[558] A. E. Owen. Semiconducting glasses part 1: Glass as an electronic conductor. Contemporary Physics, 11(3):227–255, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Owen:1970:SGPb

[559] A. E. Owen. Semiconducting glasses part II: Properties and interpretation. Contemporary Physics, 11(3):257–286, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1970:EMU

[560] D. H. Smith. Electrical and magnetic units in the système international. Contemporary Physics, 11(3):287–300, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1970:ISM

[561] L. F. Bates. The introduction of SI in magnetism. Contemporary Physics, 11(3):301–306, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRc

[562] Anonymous. Book reviews. Contemporary Physics, 11(3):307–310, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRd

[563] Anonymous. Books received. Contemporary Physics, 11(3):311–312, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gee:1970:GSP

[564] Geoffrey Gee. The glassy state in polymers. Contemporary Physics, 11 (4):313–334, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burhop:1970:EA

[565] E. H. S. Burhop. Exotic atoms. Contemporary Physics, 11(4):335–358, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Byrne:1970:RFN

[566] J. Byrne. The radioactivity of free neutrons. Contemporary Physics, 11 (4):359–386, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gechele:1970:FPA

[567] G. B. Gechele and F. Coccioli. Future polymer applications. Contemporary Physics, 11(4):387–402, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Melville:1970:PE

[568] D. Melville. The physics exhibition, 1970. Contemporary Physics, 11(4): 403–407, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRe

[569] Anonymous. Book reviews. Contemporary Physics, 11(4):409–422, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1970:REM

[570] K. N. R. Taylor. The rare-earth metals. Contemporary Physics, 11(5): 423–454, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Austin:1970:MS

[571] I. G. Austin and D. Elwell. Magnetic semiconductors. Contemporary Physics, 11(5):455–476, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1970:RSO

[572] B. J. Mason. The role of satellites in observing and forecasting the weather. Contemporary Physics, 11(5):477–496, 1970. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jope:1970:AXR

[573] J. A. Jope. The absorption of X-rays. Contemporary Physics, 11(5): 497–508, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Atkinson:1970:DOI

[574] P. A. Atkinson and S. J. Edwards. Doppler observations of ionospheric disturbances. Contemporary Physics, 11(5):509–515, 1970. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRf

[575] Anonymous. Book reviews. Contemporary Physics, 11(5):517–520, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRg

[576] Anonymous. Books received. Contemporary Physics, 11(5):521–522, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tite:1970:INS

[577] M. S. Tite. The impact of the natural sciences on archaeology. Contemporary Physics, 11(6):523–539, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilbody:1970:ACP

[578] H. B. Gilbody. Atomic collision phenomena in gases–come modern aspects. Contemporary Physics, 11(6):541–558, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Winterton:1970:VWF

[579] R. H. S. Winterton. Van der Waals forces. Contemporary Physics, 11(6): 559–574, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawson:1970:SAR

[580] J. D. Lawson. Some attributes of real and virtual photons. Contemporary Physics, 11(6):575–580, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wright:1970:ED

[581] P. G. Wright. Entropy and disorder. Contemporary Physics, 11(6):581– 588, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cade:1970:TIF

[582] C. Maxwell Cade and Boris V. Barlow. Thermography: Imaging in the far infrared. Contemporary Physics, 11(6):589–608, 1970. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lipson:1970:NRP

[583] H. Lipson. Notes on recent progress. X-ray and electron crystallography. Contemporary Physics, 11(6):609–612, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1970:MGP

[584] L. F. Bates. Magnetism as a growth-point subject. An essay review of magnetism and metallurgy. Edited by A. Berkowitz and Kneller (Academic Press, 1970). Contemporary Physics, 11(6):613–618, 1970. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRh

[585] Anonymous. Book reviews. Contemporary Physics, 11(6):619–624, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1970:BRi

[586] Anonymous. Books received. Contemporary Physics, 11(6):625, 1970. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Field:1971:BFS

[587] J. E. Field. Brittle fracture: Its study and application. Contemporary Physics, 12(1):1–31, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Treloar:1971:RE

[588] L. R. G. Treloar. Rubber elasticity. Contemporary Physics, 12(1):33– 56, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knott:1971:PD

[589] R. G. A. Knott. Plasma diagnostics. Contemporary Physics, 12(1):57– 74, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Crawford:1971:PVV

[590] B. H. Crawford and H. Ikeda. The physics of vision in vertebrates. Contemporary Physics, 12(1):75–97, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRa

[591] Anonymous. Book reviews. Contemporary Physics, 12(1):99–103, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRb

[592] Anonymous. Books received. Contemporary Physics, 12(1):104, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1971:AIG

[593] B. L. Smith. Applications of the inert gases. Contemporary Physics, 12 (2):105–132, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hasted:1971:WP

[594] J. B. Hasted. Water and ‘polywater’. Contemporary Physics, 12(2): 133–152, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Betts:1971:HIQ

[595] D. S. Betts. Helium isotope quantum fluid mixtures. Contemporary Physics, 12(2):153–178, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1971:GL

[596] E. A. Mason and P. G. Wright. Graham’s laws. Contemporary Physics, 12(2):179–186, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Strnad:1971:GIE

[597] J. Strnad. The gravity-induced electric field. Contemporary Physics, 12 (2):187–191, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRc

[598] Anonymous. Books received. Contemporary Physics, 12(2):192, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRd

[599] Anonymous. Book reviews. Contemporary Physics, 12(2):193–208, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:1971:SPH

[600] T. L. Blundell, E. Dodson, G. Dodson, and M. Vijayan. The structure of a protein hormone, insulin. Contemporary Physics, 12(3):209–228, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ruddock:1971:PCV

[601] K. H. Ruddock. The physics of colour vision. Contemporary Physics, 12 (3):229–256, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hughes:1971:ESF

[602] A. E. Hughes. Exploring solids with far infrared radiation. Contemporary Physics, 12(3):257–282, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Figgins:1971:ILS

[603] R. Figgins. Inelastic light scattering in liquids: Brillouin scattering. Contemporary Physics, 12(3):283–297, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beer:1971:PPA

[604] Tom Beer. Particle precipitation at auroral heights: Ground-based observations. Contemporary Physics, 12(3):299–310, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1971:BMP

[605] G. R. Noakes. The Bragg Medal and Prize of the Institute of Physics and the Physical Society, 1971. Contemporary Physics, 12(3):311, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRe

[606] Anonymous. Books received. Contemporary Physics, 12(3):312, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Meaden:1971:CES

[607] G. T. Meaden. Conduction electron scattering and the resistance of the magnetic elements. Contemporary Physics, 12(4):313–337, 1971. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hoover:1971:STM

[608] W. G. Hoover and M. Rossj. Statistical theories of melting. Contemporary Physics, 12(4):339–356, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:1971:MAD

[609] R. H. Brown. Measuring the angular diameters of stars. Contemporary Physics, 12(4):357–377, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Penston:1971:SF

[610] M. V. Penston. Star formation. Contemporary Physics, 12(4):379–394, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Woolley:1971:SM

[611] S. R. Woolley. Stellar movement. Contemporary Physics, 12(4):395– 409, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRf

[612] Anonymous. Book reviews. Contemporary Physics, 12(4):411–416, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRg

[613] Anonymous. Books received. Contemporary Physics, 12(4):417, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1971:HBW

[614] Boris V. Barlow. Hydrostatic bearings for the World’s great telescopes. Contemporary Physics, 12(5):419–436, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

DeRosier:1971:RTD

[615] David J. De Rosier. The reconstruction of three-dimensional images from electron micrographs. Contemporary Physics, 12(5):437–452, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Petley:1971:TQV

[616] B. W. Petley. Towards a quantum voltage standard. Contemporary Physics, 12(5):453–470, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Adams:1971:CXR

[617] D. J. Adams. Cosmic X-ray astronomy. Contemporary Physics, 12(5): 471–493, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Catterall:1971:SMT

[618] J. A. Catterall. Superconducting materials and their engineering applications. Contemporary Physics, 12(5):495–514, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Melville:1971:PE

[619] D. Melville. The physics exhibition, 1971. Contemporary Physics, 12(5): 515–520, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRh

[620] Anonymous. Book reviews. Contemporary Physics, 12(5):521, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRi

[621] Anonymous. Books received. Contemporary Physics, 12(5):522, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1971:LLS

[622] R. A. Smith. Lasers and light scattering. Contemporary Physics, 12(6): 523–536, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1971:AAA

[623] Sir Harrie Massey. Atom-atom and atom-ion collisions: I. Collisions between neutral atoms in which no excitation occurs. Contemporary Physics, 12(6):537–558, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Crossland:1971:FW

[624] B. Crossland. Friction welding. Contemporary Physics, 12(6):559–574, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lisgarten:1971:VPC

[625] N. D. Lisgarten, J. R. Sambles, and L. M. Skinner. Vapour pressure over curved surfaces — the Kelvin equation. Contemporary Physics, 12(6): 575–593, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Present:1971:NAI

[626] R. D. Present. Non-additive interactions. Contemporary Physics, 12(6): 595–602, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRj

[627] Anonymous. Book reviews. Contemporary Physics, 12(6):603–613, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1971:BRk

[628] Anonymous. Books received. Contemporary Physics, 12(6):614, 1971. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Finch:1972:SVS

[629] J. T. Finch. The study of virus structures by X-ray diffraction. Contemporary Physics, 13(1):1–21, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Diamond:1972:SCD

[630] R. Diamond. Some contemporary developments in the techniques of protein crystallography. Contemporary Physics, 13(1):23–43, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Baker:1972:DMA

[631] J. M. Baker, E. R. Davies, and T. Rs. Reddy. Detailed mapping of atomic positions using electron nuclear double resonance (endor). Contemporary Physics, 13(1):45–59, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chechev:1972:FCC

[632] V. P. Chechev and Ya M. Kramarovsky. Are the fundamental constants constant? Contemporary Physics, 13(1):61–74, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kenyon:1972:BCF

[633] I. R. Kenyon. Bubble chamber film analysis. Contemporary Physics, 13 (1):75–104, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRa

[634] Anonymous. Book reviews. Contemporary Physics, 13(1):105–111, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRb

[635] Anonymous. Books received. Contemporary Physics, 13(1):112, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lee:1972:RPA

[636] J. A. Lee and M. B. Waldron. Recent progress in actinide research. Contemporary Physics, 13(2):113–133, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1972:AAAa

[637] Sir Harrie Massey. Atom-atom and atom-ion collisions: II. Collisions between atoms and ions in which no excitation occurs-symmetrical charge transfer. Contemporary Physics, 13(2):135–158, 1972. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Powell:1972:RDU

[638] R. F. Powell. Recent developments in the use of the thermoelectric power of metals and alloys as a research tool. Contemporary Physics, 13(2): 159–177, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chamberlain:1972:APP

[639] A. C. Chamberlain and S. A. Penkett. Atmospheric pollution–present trends and future problems. Contemporary Physics, 13(2):179–198, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Christov:1972:ECT

[640] S. G. Christov. Electron currents through barriers between two metals. Contemporary Physics, 13(2):199–222, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRc

[641] Anonymous. Book reviews. Contemporary Physics, 13(2):223, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRd

[642] Anonymous. Books received. Contemporary Physics, 13(2):224, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1972:UES

[643] B. Bates and C. D. McKeith. Ultraviolet emission from the stars. Contemporary Physics, 13(3):225–246, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beer:1972:AWI

[644] Tom Beer. Atmospheric waves and the ionosphere. Contemporary Physics, 13(3):247–271, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hawking:1972:GRT

[645] S. W. Hawking. Gravitational radiation: The theoretical aspect. Contemporary Physics, 13(3):273–282, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aplin:1972:GRE

[646] Peter S. Aplin. Gravitational radiation experiments. Contemporary Physics, 13(3):283–293, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lenihan:1972:WBE

[647] J. M. A. Lenihan. What is bio-engineering? Contemporary Physics, 13 (3):295–309, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:SJC

[648] Anonymous. Sir James Chadwick, C.H., LL.D., F.R.S.: 80th birthday. Contemporary Physics, 13(3):310, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRe

[649] Anonymous. Book reviews. Contemporary Physics, 13(3):311–319, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRf

[650] Anonymous. Books received. Contemporary Physics, 13(3):320, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shoenberg:1972:MEM

[651] D. Shoenberg. Metallic electrons in magnetic fields: Part I. Electronic structure and electron dynamics. Contemporary Physics, 13(4):321–354, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1972:HSR

[652] Boris V. Barlow and R. S. Read. High-speed rotating devices. Contemporary Physics, 13(4):355–374, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1972:AAAb

[653] Sir Harrie Massey. Atom-atom and atom-ion collisions. III. Collisions involving excitation transfer, spin exchange or depolarization, without internal energy change. Contemporary Physics, 13(4):375–392, 1972. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paine:1972:MFS

[654] Thomas O. Paine. Man’s future in space. Contemporary Physics, 13(4): 393–401, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRg

[655] Anonymous. Books received. Contemporary Physics, 13(4):402, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRh

[656] Anonymous. Book review. Contemporary Physics, 13(4):403–418, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:1972:UFP

[657] W. J. Jones. On the use of a Fabry–Pérot interferometer for the study of Raman spectra of gases under high resolution. Contemporary Physics, 13(5):419–439, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hayes:1972:SSL

[658] W. Hayes. Spectroscopy with a synchrotron light source. Contemporary Physics, 13(5):441–456, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Canter:1972:SPR

[659] K. F. Canter. Slow positrons in rare gases. Contemporary Physics, 13(5): 457–478, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henderson:1972:MS

[660] C. Henderson. Molecules in space. Contemporary Physics, 13(5):479– 499, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRi

[661] Anonymous. Books received. Contemporary Physics, 13(5):500, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1972:GMW

[662] L. F. Bates. “A great man who did good things” [review of Rutherford: Recollections of the Cambridge Days, by Mark Oliphant (Elsevier, 1972). [Pp. xii + 158] Dfl 22.50; $7.00]. Contemporary Physics, 13(5): 501–502, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRj

[663] Anonymous. Book reviews. Contemporary Physics, 13(5):503–518, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Golub:1972:EDM

[664] R. Golub and J. M. Pendlebury. The electric dipole moment of the neutron. Contemporary Physics, 13(6):519–558, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Irani:1972:TMO

[665] R. S. Irani. The thermodynamics and mechanisms of ordering systems. Contemporary Physics, 13(6):559–583, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:1972:ISP

[666] G. H. A. Cole. The internal structure of planetary bodies. Contemporary Physics, 13(6):585–600, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1972:LBP

[667] L. F. Bates. A link between past and present in European magnetism. Contemporary Physics, 13(6):601–614, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRk

[668] Anonymous. Book reviews. Contemporary Physics, 13(6):615–617, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1972:BRl

[669] Anonymous. Books received. Contemporary Physics, 13(6):618, 1972. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Groves:1973:SAK

[670] G. V. Groves. The structure of the atmosphere up to 150 kilometre. Contemporary Physics, 14(1):1–24, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bright:1973:EPS

[671] A. W. Bright and B. Makin. The electrical propulsion of space vehicles. Contemporary Physics, 14(1):25–38, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fielder:1973:GM

[672] G. Fielder. The geology of the moon. Contemporary Physics, 14(1): 39–54, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1973:PIT

[673] R. G. F. Taylor and H. A. H. Boot. Pyroelectric image tubes. Contemporary Physics, 14(1):55–87, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRa

Bates:1973:ERA

[675] L. F. Bates. Essay review: An American study of research expenditure and support for physics. Atomic and Molecular Physics. Contemporary Physics, 14(1):89–90, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRb

[676] Anonymous. Book reviews. Contemporary Physics, 14(1):91–100, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berman:1973:HCN

[677] R. Berman. Heat conductivity of non-metallic crystals. Contemporary Physics, 14(2):101–117, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eagland:1973:CS

[678] D. Eagland. The colloidal state. Contemporary Physics, 14(2):119–148, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saunders:1973:SNG

[679] G. A. Saunders. Semimetals and narrow gap semiconductors. Contemporary Physics, 14(2):149–166, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cocconi:1973:CIS

[680] G. Cocconi. The CERN intersecting storage rings and strong interactions at high energies. Contemporary Physics, 14(2):167–186, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRc

[681] Anonymous. Book reviews. Contemporary Physics, 14(2):187–199, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRd

[682] Anonymous. Books received. Contemporary Physics, 14(2):200, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pearson:1973:AMO

[683] R. F. Pearson. Application of magneto-optic effects in magnetic materials. Contemporary Physics, 14(3):201–211, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frazer:1973:DPG

[684] M. C. Frazer. The dynamo problem and the geomagnetic field. Contemporary Physics, 14(3):213–228, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rishbeth:1973:PCI

[685] H. Rishbeth. Physics and chemistry of the ionosphere. Contemporary Physics, 14(3):229–249, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mitchell:1973:LEE

[686] K. A. R. Mitchell. Low-energy electron diffraction. Contemporary Physics, 14(3):251–271, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1973:SMP

[687] L. F. Bates. The study and maintenance of physical standards. Contemporary Physics, 14(3):273–277, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRe

[688] Anonymous. Book reviews. Contemporary Physics, 14(3):278–286, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRf

[689] Anonymous. Books received. Contemporary Physics, 14(3):287–288, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kim:1973:SQT

[690] Young S. Kim. The search for quarks in terrestrial matter. Contemporary Physics, 14(4):289–318, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beer:1973:EI

[691] Tom Beer. The equatorial ionosphere. Contemporary Physics, 14(4): 319–327, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bahr:1973:PS

[692] John L. Bahr. Photoelectron spectroscopy. Contemporary Physics, 14(4): 329–355, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hasted:1973:ESS

[693] J. B. Hasted. Electron scattering spectroscopy. Contemporary Physics, 14(4):357–387, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chandrasekhar:1973:PMD

[694] S. Chandrasekhar. P. A. M. Dirac on his seventieth birthday. Contemporary Physics, 14(4):389–394, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRg

[695] Anonymous. Book reviews. Contemporary Physics, 14(4):395–398, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRh

[696] Anonymous. Books received. Contemporary Physics, 14(4):399–400, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1973:MIT

[697] N. F. Mott. Metal-insulator transitions. Contemporary Physics, 14(5): 401–413, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Valkovic:1973:XREa

[698] V. Valković. X-ray emission spectroscopy — Part I. Contemporary Physics, 14(5):415–438, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Valkovic:1973:XREb

[699] V. Valković. X-ray emission spectroscopy — part II. Contemporary Physics, 14(5):439–462, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lemmon:1973:PS

[700] R. M. Lemmon. Prebiological synthesis. Contemporary Physics, 14(5): 463–477, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1973:NAC

[701] L. F. Bates. A new American compilation of physics data. Contemporary Physics, 14(5):479–481, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRi

[702] Anonymous. Book reviews. Contemporary Physics, 14(5):482–495, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRj

[703] Anonymous. Books received. Contemporary Physics, 14(5):496, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Massey:1973:IC

[704] Sir Harrie Massey. Inelastic collisions. Contemporary Physics, 14(6): 497–512, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riviere:1973:AES

[705] J. C. Rivière. Auger electron spectroscopy. Contemporary Physics, 14(6): 513–539, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Megaw:1973:BAA

[706] W. J. Megaw. The background atmospheric aerosol. Contemporary Physics, 14(6):541–557, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hamberger:1973:DHT

[707] S. M. Hamberger. The development of high temperature plasma physics in the last decade. Contemporary Physics, 14(6):559–573, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1973:BRk

Anonymous:1973:BRl

[709] Anonymous. Books received. Contemporary Physics, 14(6):588, 1973. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chandrasekhar:1974:BHA

[710] S. Chandrasekhar. The black hole in astrophysics: The origin of the concept and its role. Contemporary Physics, 15(1):1–24, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jackson:1974:CMC

[711] D. P. Jackson and D. V. Morgan. Computer modelling of collision processes in solids. Contemporary Physics, 15(1):25–48, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Series:1974:RC

[712] G. W. Series. The Rydberg constant. Contemporary Physics, 15(1): 49–68, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shott:1974:POU

[713] M. Shott. Physics in the open university. Contemporary Physics, 15 (1):69–85, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRa

[714] Anonymous. Books received. Contemporary Physics, 15(1):86, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRb

[715] Anonymous. Book reviews. Contemporary Physics, 15(1):87–96, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:1974:RPC

[716] J. J. Quenby. Recent progress in cosmic ray physics. Contemporary Physics, 15(2):97–120, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Geake:1974:PLP

[717] J. E. Geake. The physics of lunar and planetary surfaces. Contemporary Physics, 15(2):121–158, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Israelachvili:1974:NVW

[718] Jacob N. Israelachvili. The nature of van der Waals forces. Contemporary Physics, 15(2):159–178, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pease:1974:ICR

[719] R. S. Pease and D. R. Willson. International collaboration in research on controlled thermonuclear fusion. Contemporary Physics, 15(2):179– 192, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRc

[720] Anonymous. Book reviews. Contemporary Physics, 15(2):193–195, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRd

[721] Anonymous. Books received. Contemporary Physics, 15(2):196, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Backenstoss:1974:EBS

[722] G. Backenstoss and J. Zakrzewski. Exotic bound states of strange hadrons. Contemporary Physics, 15(3):197–225, 1974. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Betts:1974:PCA

[723] David S. Betts. Pomeranchuk cooling by adiabatic solidification of helium-3. Contemporary Physics, 15(3):227–247, 1974. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Swithenby:1974:MLH

[724] S. J. Swithenby. Magnetometry at liquid helium temperatures. Contemporary Physics, 15(3):249–269, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Little:1974:SLL

[725] V. I. Little, M. A. Cutter, P. Y. Key, and R. G. Harrison. Scattering of light from light induced periodic structures. Contemporary Physics, 15 (3):271–298, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRe

[726] Anonymous. Book reviews. Contemporary Physics, 15(3):299, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRf

[727] Anonymous. Books received. Contemporary Physics, 15(3):300, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:1974:STL

[728] R. C. Brown. The surface tension of liquids. Contemporary Physics, 15 (4):301–327, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beynon:1974:ISS

[729] W. J. G. Beynon. Incoherent scatter sounding of the ionosphere. Contemporary Physics, 15(4):329–352, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lounasmaa:1974:SPL

[730] O. V. Lounasmaa. The superfluid phases of liquid^{3} he. Contemporary Physics, 15(4):353–374, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jordan:1974:ESM

[731] R. G. Jordan. Electrotransport in solid metal systems. Contemporary Physics, 15(4):375–400, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charap:1974:PCN

[732] J. M. Charap. The physicist’s conception of nature. Contemporary Physics, 15(4):401–404, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRg

[733] Anonymous. Book reviews. Contemporary Physics, 15(4):405–407, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRh

Powles:1974:LAB

[735] J. G. Powles. Liquids — the awkward in-between. Contemporary Physics, 15(5):409–425, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Twiddy:1974:SSF

[736] N. D. Twiddy. The steady-state flowing afterglow method. Contemporary Physics, 15(5):427–444, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shaw:1974:PI

[737] M. J. Shaw. Penning ionization. Contemporary Physics, 15(5):445–464, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rotherham:1974:NPS

[738] L. Rotherham. Nuclear power systems — some principles. Contemporary Physics, 15(5):465–478, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lipson:1974:SPA

[739] H. S. Lipson. Scientific papers of Arthur Holly Compton. Contemporary Physics, 15(5):479–480, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:1974:SPC

[740] G. R. Bishop. Science policy and co-operative research in Europe. Contemporary Physics, 15(5):481–482, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRi

[741] Anonymous. Book reviews. Contemporary Physics, 15(5):483–499, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRj

[742] Anonymous. Books received. Contemporary Physics, 15(5):500, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chesterman:1974:OF

[743] W. Deryck Chesterman. The ocean floor. Contemporary Physics, 15(6): 501–516, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hardman:1974:SPC

[744] J. S. Hardman and B. A. Lilley. The study of powder compaction mechanisms. Contemporary Physics, 15(6):517–536, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1974:CFT

[745] R. P. Williams. A convolution and Fourier-transform approach to paraxial optics. Contemporary Physics, 15(6):537–552, 1974. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ruddock:1974:TPB

[746] K. H. Ruddock. From theoretical physics to biology. Contemporary Physics, 15(6):553–558, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1974:PD

[747] W. H. McCrea. Physics in debate. Contemporary Physics, 15(6):559– 561, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1974:DWP

[748] L. F. Bates. Domain wall pinning. Contemporary Physics, 15(6):562– 564, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1974:CN

[749] N. Feather. Chadwick’s neutron. Contemporary Physics, 15(6):565–572, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRk

[750] Anonymous. Book reviews. Contemporary Physics, 15(6):573–597, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1974:BRl

[751] Anonymous. Books received. Contemporary Physics, 15(6):598, 1974. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blandford:1975:EDR

[752] R. D. Blandford and M. J. Rees. Extragalactic double radio sources — the current observational and theoretical position. Contemporary Physics, 16(1):1–16, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peach:1975:WSL

[753] G. Peach. The width of spectral lines. Contemporary Physics, 16(1): 17–34, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Allen:1975:PM

[754] Geoffrey Allen. Polymeric materials. Contemporary Physics, 16(1):35– 49, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Browne:1975:ASE

[755] P. F. Browne. Aspects of stellar evolution reappraised. Contemporary Physics, 16(1):51–68, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hayes:1975:SLS

[756] W. Hayes. Scattering of light by solids. Contemporary Physics, 16(1): 69–91, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRa

[757] Anonymous. Books received. Contemporary Physics, 16(1):92, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRb

[758] Anonymous. Book review. Contemporary Physics, 16(1):93–100, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Putley:1975:HEA

[759] E. H. Putley. The hall effect and its applications. Contemporary Physics, 16(2):101–126, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dale:1975:MS

[760] B. W. Dale. Mössbauer spectroscopy. Contemporary Physics, 16(2): 127–146, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fluendy:1975:RSM

[761] Malcolm A. D. Fluendy. Reactive scattering of molecular beams. Contemporary Physics, 16(2):147–172, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Perkins:1975:NNS

[762] D. H. Perkins. Neutrinos and nucleon structure. Contemporary Physics, 16(2):173–205, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1975:TD

[763] W. H. McCrea. Trigger for disaster. Contemporary Physics, 16(2):207– 208, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hunt:1975:TAH

[764] S. E. Hunt. Technology arising from high energy physics. Contemporary Physics, 16(2):209–210, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1975:CLA

[765] Norman Feather. The Cavendish Laboratory — aristocrat’s playground to people’s workshop. Contemporary Physics, 16(2):211–213, 1975. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRR

[766] Anonymous. Book reviews: Reviews by Professor John Lenihan, Glasgow University. Contemporary Physics, 16(2):215–218, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRc

[767] Anonymous. Books received. Contemporary Physics, 16(2):219–220, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morgan:1975:RAS

[768] D. Vernon Morgan. Recent advances in surface studies: Ion beam analysis. Contemporary Physics, 16(3):221–241, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

TerHaar:1975:P

[769] D. Ter Haar. Pulsars. Contemporary Physics, 16(3):243–256, 1975. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beer:1975:WG

[770] Tom Beer. Whispering galleries. Contemporary Physics, 16(3):257–262, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bradley:1975:SRA

[771] D. J. Bradley. Some recent advances in lasers and opto-electronics. Contemporary Physics, 16(3):263–286, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:1975:ORS

[772] O. C. Jones. Optical radiation scales. Contemporary Physics, 16(3): 287–310, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRd

[773] Anonymous. Book reviews. Contemporary Physics, 16(3):311–315, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRe

[774] Anonymous. Books received. Contemporary Physics, 16(3):316, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ruddock:1975:VFP

[775] K. H. Ruddock. Visual form perception. Contemporary Physics, 16(4): 317–348, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1975:SPR

[776] Denzil Taylor Smith. The sea: A physical resource bonanza? Contemporary Physics, 16(4):349–374, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bell:1975:DTM

[777] A. E. Bell and A. D. Caplin. Dilute transition metal alloys and the Kondo problem. Contemporary Physics, 16(4):375–394, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1975:IPM

[778] A. H. Cook. The importance of precise measurement in physics. Contemporary Physics, 16(4):395–408, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRf

[779] Anonymous. Book reviews. Contemporary Physics, 16(4):409–411, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRg

[780] Anonymous. Books received. Contemporary Physics, 16(4):412, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parley:1975:AMM

[781] F. J. M. Parley. The anomalous magnetic moment of the muon. Contemporary Physics, 16(5):413–441, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anderson:1975:SLR

[782] J. S. Anderson and J. L. Hutchison. The study of long range order in hexagonal barium ferrite layer structures. Contemporary Physics, 16(5): 443–467, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Briggs:1975:IIR

[783] B. H. Briggs. Ionospheric irregularities and radio scintillations. Contemporary Physics, 16(5):469–488, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1975:TTY

[784] Norman Feather. Twice thirty years of physics. Contemporary Physics, 16(5):489–498, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowers:1975:CWE

[785] Brian Bowers. Charles Wheatstone — experimental philosopher. Contemporary Physics, 16(5):499–512, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRh

Anonymous:1975:BRi

[787] Anonymous. Book received. Contemporary Physics, 16(5):516, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hurd:1975:ASE

[788] C. M. Hurd. Asymmetric scattering of electrons in metals. Contemporary Physics, 16(6):517–532, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1975:GAT

[789] B. J. Mason. The GARP Atlantic tropical experiment. Contemporary Physics, 16(6):533–546, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1975:MMT

[790] C. A. Bates and P. H. Wood. Magnetic moments and their interaction in insulating crystals. Contemporary Physics, 16(6):547–560, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Enoch:1975:MDS

[791] R. D. Enoch. Magnetic domain studies on yttrium iron garnet using an infrared image converter. Contemporary Physics, 16(6):561–573, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stratton:1975:ORO

[792] A. Stratton. Operational research: (i) its origins and development in government service. Contemporary Physics, 16(6):575–588, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1975:GO

[793] W. H. McCrea, F.R.S. Greenwich Observatory. Contemporary Physics, 16(6):589–590, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRj

[794] Anonymous. Book reviews. Contemporary Physics, 16(6):591–610, 1975. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1975:BRk

Anonymous:1976:OSG

[796] Anonymous. Obituary: Sir George Paget Thomson, F.R.S. (1892–1975). Contemporary Physics, 17(1):i–iii, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mayneord:1976:SUM

[797] W. V. Mayneord. SI units in medical physics. Contemporary Physics, 17 (1):1–12, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cracknell:1976:AU

[798] Margaret F. Cracknell and Arthur P. Cracknell. Applications of ultrasonics. Contemporary Physics, 17(1):13–44, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evans:1976:D

[799] T. Evans. Diamonds. Contemporary Physics, 17(1):45–70, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Grimes:1976:CSA

[800] N. W. Grimes. Crystal structures with asymmetric and off-centre atoms. Contemporary Physics, 17(1):71–93, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rivett:1976:ORP

[801] B. H. P. Rivett. Operational research: (2) perspective and prologue. Contemporary Physics, 17(1):95–102, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bragg:1976:DXR

[802] Sir Lawrence Bragg. The development of X-ray analysis. Contemporary Physics, 17(1):103–104, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRa

[803] Anonymous. Book reviews. Contemporary Physics, 17(1):105–107, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRb

[804] Anonymous. Book received. Contemporary Physics, 17(1):108, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trevena:1976:SSL

[805] D. H. Trevena. The stretching and superheating of liquids. Contemporary Physics, 17(2):109–126, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garlick:1976:IVL

[806] G. F. J. Garlick. Infrared to visible light conversion. Contemporary Physics, 17(2):127–144, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Killingbeck:1976:UPE

[807] J. P. Killingbeck. The use of pocket electronic calculators in scientific work. Contemporary Physics, 17(2):145–168, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kent:1976:OBB

[808] Sir Peter Kent. Oil from beneath Britain’s seas. Contemporary Physics, 17(2):169–172, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weaire:1976:SAS

[809] D. Weaire. The structure of amorphous solids. Contemporary Physics, 17(2):173–191, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1976:BAO

[810] L. F. Bates. The bases of authority for our symbols and units. Contemporary Physics, 17(2):193–196, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feather:1976:BMF

[811] Noman Feather. Biographical memoirs of Fellows of the Royal Society. Contemporary Physics, 17(2):197–198, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Enoch:1976:MB

[812] R. D. Enoch. Magnetic bubbles. Contemporary Physics, 17(2):199–202, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRc

[813] Anonymous. Books received. Contemporary Physics, 17(2):203–204, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRd

[814] Anonymous. Book reviews. Contemporary Physics, 17(2):205–208, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coleman:1976:II

[815] C. I. Coleman and A. Boksenberg. Image intensifiers. Contemporary Physics, 17(3):209–236, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dolder:1976:CBE

[816] K. T. Dolder. Crossed beam experiments. Contemporary Physics, 17(3): 237–248, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Durao:1976:IVT

[817] D. F. G. Durão and J. H. Whitelaw. Instantaneous velocity and temperature measurements in oscillating diffusion flames. Contemporary Physics, 17(3):249–274, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1976:ODM

[818] J. J. Davies. Optically-detected magnetic resonance and its applications. Contemporary Physics, 17(3):275–294, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1976:QGO

[819] A. H. Cook. Quantum gravity: An Oxford symposium. Contemporary Physics, 17(3):295–296, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parker:1976:PNE

[820] K. Parker. Peaceful nuclear explosions — the state of the art. Contemporary Physics, 17(3):297–300, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRe

[821] Anonymous. Book reviews. Contemporary Physics, 17(3):301–305, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRf

[822] Anonymous. Book received. Contemporary Physics, 17(3):306, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shaw:1976:PPD

[823] Dennis Shaw. Physics in the prevention and detection of crime. Contemporary Physics, 17(4):307–330, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berlinsky:1976:OM

[824] A. J. Berlinsky. Organic metals. Contemporary Physics, 17(4):331–354, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Komarek:1976:ST

[825] P. Komarek. Superconductivity in technology. Contemporary Physics, 17(4):355–386, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Allen:1976:IHF

[826] P. S. Allen. Internal hyperfine fields in magnetic solids. Contemporary Physics, 17(4):387–402, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1976:NTS

[827] J. G. Wilson. Nuclear tracks in solids. Contemporary Physics, 17(4): 403–404, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRg

[828] Anonymous. Book reviews. Contemporary Physics, 17(4):405–409, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRh

Squires:1976:STN

[830] G. L. Squires. The scattering of thermal neutrons by crystals. Contemporary Physics, 17(5):411–441, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heddle:1976:ROE

[831] D. W. O. Heddle. Resonances in optical excitation functions. Contemporary Physics, 17(5):443–460, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowlt:1976:TSE

[832] C. Bowlt. Thermally stimulated effects in dielectrics and their application to radiation dosimetry. Contemporary Physics, 17(5):461–482, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rishbeth:1976:ENP

[833] H. Rishbeth. EISCAT: A new project for studying the high latitude ionosphere. Contemporary Physics, 17(5):483–489, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yapp:1976:ERN

[834] W. Brunsdon Yapp. Essay review: Nuclear energy and the environment. Combined Effects of Radioactive, Chemical and Thermal Rebases to the Environment. Proceedings of a Symposium held in Stockholm, 2–5 June 1975. Vienna, 1975. International Atomic Energy Agency. London, H.M.S.O., pp. 358. £10.00. Impacts of Nuclear Releases into the Aquatic Environment. Proceedings of an International Symposium held at Otaniemi, Finland, 30 June–4 July 1975. Vienna, 1975. International Atomic Energy Agency. London, H.M.S.O., pp. 524. £14.60. Contemporary Physics, 17(5):491–494, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRi

[835] Anonymous. Book review. Contemporary Physics, 17(5):495–505, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRj

Clarke:1976:RET

[837] M. E. Clarke and J. B. Gardner. Refrigeration with expansion turbines. Contemporary Physics, 17(6):507–528, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heavens:1976:SRD

[838] O. S. Heavens. Some recent developments in gas lasers. Contemporary Physics, 17(6):529–552, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mansfield:1976:PSI

[839] P. Mansfield. Proton spin imaging by nuclear magnetic resonance. Contemporary Physics, 17(6):553–576, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1976:MDM

[840] A. H. Cook. Measurement of distances to the Moon and artificial satellites. Contemporary Physics, 17(6):577–598, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRk

[841] Anonymous. Book reviews. Contemporary Physics, 17(6):599–609, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1976:BRl

[842] Anonymous. Book received. Contemporary Physics, 17(6):610, 1976. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morrell:1977:ZGP

[843] R. Morrell. ‘Zero-gravity’ physics — the Spacelab project. Contemporary Physics, 18(1):1–22, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1977:GBL

[844] B. V. Barlow. Ground based low light astronomy. Contemporary Physics, 18(1):23–46, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Grundy:1977:MBT

[845] P. J. Grundy. Magnetic bubbles and their observation in the electron microscope. Contemporary Physics, 18(1):47–72, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Croft:1977:SAB

[846] A. J. Croft. Scientific aids for the blind. Contemporary Physics, 18(1): 73–80, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:1977:PSLa

[847] D. H. Martin. Probing solids and liquids (part 1). Contemporary Physics, 18(1):81–98, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bates:1977:ERN

[848] L. F. Bates. Essay review: New growth points in magnetism: a review of Magnetism: Selected Topics, edited by Simon Foner. Contemporary Physics, 18(1):99–103, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRa

[849] Anonymous. Book reviews. Contemporary Physics, 18(1):104, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lambert:1977:ERI

[850] B. E. Lambert. Essay review: The insoluble plutonium hazard. The Diagnosis and Treatment of Radionuclides. Contemporary Physics, 18 (1):105–107, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRb

[851] Anonymous. Book reviews. Contemporary Physics, 18(1):108–110, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRc

[852] Anonymous. Books received. Contemporary Physics, 18(1):111–112, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pease:1977:PCN

[853] R. S. Pease. The potential of controlled nuclear fusion. Contemporary Physics, 18(2):113–135, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pitt:1977:DHP

[854] G. D. Pitt. Developments in high pressure physics. Contemporary Physics, 18(2):137–164, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ratcliffe:1977:M

[855] J. A. Ratcliffe. The magnetosphere. Contemporary Physics, 18(2):165– 182, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Warren:1977:SB

[856] C. H. E. Warren. Sonic bangs. Contemporary Physics, 18(2):183–192, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:1977:PSLb

[857] D. H. Martin. Probing solids and liquids (part 2). Contemporary Physics, 18(2):193–218, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRd

[858] Anonymous. Book reviews. Contemporary Physics, 18(2):219–222, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRe

[859] Anonymous. Books received. Contemporary Physics, 18(2):223–224, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1977:EG

[860] N. F. Mott. Electrons in glass. Contemporary Physics, 18(3):225–245, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brochard:1977:NFS

[861] L. F. Brochard. Nematic fluids: Some easy demonstration experiments. Contemporary Physics, 18(3):247–264, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Killingbeck:1977:SAP

[862] J. P. Killingbeck. Some applications of programmable calculators. Contemporary Physics, 18(3):265–278, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parker:1977:PMS

[863] Martin R. Parker. The physics of magnetic separation. Contemporary Physics, 18(3):279–306, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beer:1977:TM

[864] Tom Beer. Tropical meteorology. Contemporary Physics, 18(3):307– 314, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bondi:1977:NCK

[865] H. Bondi. Nigel Calder: ‘the key to the universe’. Contemporary Physics, 18(3):315–316, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rowlinson:1977:HB

[866] J. S. Rowlinson. The hydrogen bond. Contemporary Physics, 18(3): 317–318, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leslie:1977:PFR

[867] D. C. Leslie. Pulsed fusion reactors. Contemporary Physics, 18(3):319, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yapp:1977:SVP

[868] W. Brunsdon Yapp. Science versus pollution. Contemporary Physics, 18 (3):320, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1977:RJP

[869] Professor M. M. R. Williams. A review of: “The JET project: Design proposal for the Joint European Torus, Commission of the European Communities, Brussels 1976. [Pp. xxx + 615.] £9.80., Scope: Report/ Survey. Level: Specialist”. Contemporary Physics, 18(3):321, 1977. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRf

[870] Anonymous. Books received. Contemporary Physics, 18(3):322, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bray:1977:TSA

[871] W. J. Bray. Telecommunications — the state of the art. Contemporary Physics, 18(4):323–349, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gadian:1977:NMR

[872] D. G. Gadian. Nuclear magnetic resonance in living tissue. Contemporary Physics, 18(4):351–372, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haasen:1977:PHM

[873] P. Haasen. Precipitation hardening for mechanical strength, magnetic hardness and hardness of superconductors. Contemporary Physics, 18 (4):373–392, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1977:SMS

[874] A. H. Cook. Standards of measurement and the structure of physical knowledge. Contemporary Physics, 18(4):393–409, 1977. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hiley:1977:ERF

[875] B. J. Hiley. Essay review: Foundations of quantum mechanics. Conceptual Foundations of Quantum Mechanics, by Bernard D’Espagnat. Contemporary Physics, 18(4):411–414, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRg

[876] Anonymous. Book reviews. Contemporary Physics, 18(4):415–420, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRh

Pepper:1977:ISI

[878] M. Pepper. An introduction to silicon inversion layers. Contemporary Physics, 18(5):423–454, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Skinner:1977:RGS

[879] N. J. Skinner. Recent geophysical studies of the Kenya Rift Valley. Contemporary Physics, 18(5):455–470, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haselden:1977:LNG

[880] G. G. Haselden. The liquefaction of natural gas. Contemporary Physics, 18(5):471–488, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cooper:1977:CSE

[881] Malcolm J. Cooper. Compton scattering and electron momentum. Contemporary Physics, 18(5):489–517, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRi

[882] Anonymous. Book reviews. Contemporary Physics, 18(5):519–524, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRj

[883] Anonymous. Books received. Contemporary Physics, 18(5):525–526, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ansaldo:1977:USR

[884] E. J. Ansaldo. Uses of synchrotron radiation. Contemporary Physics, 18 (6):527–546, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Edwards:1977:ACS

[885] S. J. Edwards. Applications of colour science. Contemporary Physics, 18(6):547–554, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hurley:1977:UMD

[886] G. D. Hurley and J. F. Malone. Ultrasound in medical diagnosis. Contemporary Physics, 18(6):555–570, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matthews:1977:BMF

[887] J. R. Matthews. The behaviour of materials in fast reactors. Contemporary Physics, 18(6):571–591, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRk

[888] Anonymous. Book reviews. Contemporary Physics, 18(6):593–603, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1977:BRl

[889] Anonymous. Books received. Contemporary Physics, 18(6):604, 1977. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1978:EE

[890] P. M. Williams. The evolution of the elements. Contemporary Physics, 19(1):1–24, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Debney:1978:TSC

[891] B. T. Debney and J. R. Knight. Terrestrial solar cells — present and future. Contemporary Physics, 19(1):25–45, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1978:HPA

[892] B. V. Barlow. Hypersensitizing processes in astrophotography. Contemporary Physics, 19(1):47–64, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stwalley:1978:LRM

[893] William C. Stwalley. Long-range molecules. Contemporary Physics, 19 (1):65–80, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1978:FU

[894] W. H. McCrea. Faces of the universe. Contemporary Physics, 19(1): 81–82, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burhop:1978:SI

[895] E. H. S. Burhop. The strong interaction. Contemporary Physics, 19 (1):83–84, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1978:RSE

[896] Boris v. Barlow. Remote sensing of the environment. Contemporary Physics, 19(1):85–86, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRa

[897] Anonymous. Book reviews. Contemporary Physics, 19(1):87–94, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRb

[898] Anonymous. Books received. Contemporary Physics, 19(1):95–96, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:OPF

[899] Anonymous. Obituary Professor L. F. Bates. Contemporary Physics, 19 (2):97–98, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Malinowski:1978:PSS

[900] J. Malinowski and A. Buroff. Photostimulated solid state reactions. Contemporary Physics, 19(2):99–108, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ovshinsky:1978:LSB

[901] Stanford R. Ovshinsky and David Adler. Local structure, bonding, and electronic properties of covalent amorphous semiconductors. Contemporary Physics, 19(2):109–126, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jakeman:1978:TS

[902] E. Jakeman, G. Parry, E. R. Pike, and P. N. Pusey. The twinkling of stars. Contemporary Physics, 19(2):127–145, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hooper:1978:FIC

[903] A. Hooper. Fast ionic conductors. Contemporary Physics, 19(2):147– 168, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRc

[904] Anonymous. Books received. Contemporary Physics, 19(2):169–175, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRd

[905] Anonymous. Books received. Contemporary Physics, 19(2):176, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:1978:CT

[906] Ian Solomon and Robert Golub. Climatic trends. Contemporary Physics, 19(3):177–201, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1978:SMP

[907] Colin J. Thompson. The statistical mechanics of phase transitions. Contemporary Physics, 19(3):203–224, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henderson:1978:CSI

[908] B. Henderson. Colour, symmetry and imperfect crystals. Contemporary Physics, 19(3):225–267, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mapps:1978:MDD

[909] D. J. Mapps. Magnetic domain demonstrations using the Kerr magnetooptic effect. Contemporary Physics, 19(3):269–281, 1978. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRe

[910] Anonymous. Books reviews. Contemporary Physics, 19(3):283–286, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stern:1978:SDX

[911] E. A. Stern. Structure determination by X-ray absorption. Contemporary Physics, 19(4):289–310, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Borovik-Romanov:1978:SWA

[912] A. S. Borovik-Romanov and L. A. Prozorova. Spin waves in antiferromagnetics. Contemporary Physics, 19(4):311–326, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coqblin:1978:SFM

[913] B. Coqblin, J. R. Iglesias-Sicardi, and R. Jullien. Spin fluctuations in metals: Application to the actinides. Contemporary Physics, 19(4):327– 354, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dawe:1978:ROP

[914] R. A. Dawe and E. O. Egbogah. The recovery of oil from petroleum reservoirs. Contemporary Physics, 19(4):355–376, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1978:ESS

[915] R. H. Williams. Electron spectroscopy of surfaces. Contemporary Physics, 19(5):389–413, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Connerade:1978:NRS

[916] J. P. Connerade. The non-Rydberg spectroscopy of atoms. Contemporary Physics, 19(5):415–447, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lehmann:1978:ITM

[917] J. C. Lehmann. Iodine, a test molecule in modern spectroscopy. Contemporary Physics, 19(5):449–467, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hayes:1978:SC

[918] W. Hayes. Superionic conductors. Contemporary Physics, 19(5):469– 486, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bondi:1978:MCP

[919] Hermann Bondi. As much of cosmology as is possible. Contemporary Physics, 19(5):487–488, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRf

[920] Anonymous. Book reviews. Contemporary Physics, 19(5):489–492, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRg

[921] Anonymous. Books received. Contemporary Physics, 19(5):493–494, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:E

[922] Anonymous. Editorial. Contemporary Physics, 19(6):i, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ratcliffe:1978:WUA

[923] J. A. Ratcliffe. Wireless and the upper atmosphere 1900–1935. Contemporary Physics, 19(6):495–504, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Manning:1978:SNS

[924] G. Manning. Spallation neutron sources for neutron beam research. Contemporary Physics, 19(6):505–529, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Feates:1978:NPR

[925] F. S. Feates. Nuclear power and radioactive waste management. Contemporary Physics, 19(6):531–542, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nelson:1978:UIB

[926] R. S. Nelson. The use of ion beams to study radiation effects in metals. Contemporary Physics, 19(6):543–569, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burcham:1978:NP

[927] W. E. Burcham. Nuclei and particles. Contemporary Physics, 19(6): 571–572, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1978:HR

[928] B. V. Barlow. Hologram recording. Contemporary Physics, 19(6):573– 574, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1978:AC

[929] W. H. McCrea. Astronomy and cosmology. Contemporary Physics, 19(6): 575–577, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRh

[930] Anonymous. Book reviews. Contemporary Physics, 19(6):579–582, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1978:BRi

[931] Anonymous. Books received. Contemporary Physics, 19(6):583–584, 1978. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cottrell:1979:NPE

[932] Alan Cottrell. The natural philosophy of engines. Contemporary Physics, 20(1):1–10, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Colles:1979:OE

[933] M. J. Colles, N. R. Geddes, and E. Mehdizadeh. The optoacoustic effect. Contemporary Physics, 20(1):11–36, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:1979:LIM

[934] Stig Stenholm. Laser-induced multiphoton transitions. Contemporary Physics, 20(1):37–54, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cracknell:1979:SHM

[935] A. P. Cracknell and A. O. Tooke. The specific heats of magneticallyordered materials. Contemporary Physics, 20(1):55–82, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yapp:1979:CP

[936] W. Brunsdon Yapp. Contemporary pollution. Contemporary Physics, 20 (1):83–84, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1979:OID

[937] B. V. Barlow. Optical and infrared detectors. Contemporary Physics, 20 (1):85–86, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lipson:1979:EXR

[938] H. Lipson. Elements of X-ray diffraction. Contemporary Physics, 20(1): 87–88, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRa

[939] Anonymous. Book reviews. Contemporary Physics, 20(1):89–118, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRb

[940] Anonymous. Books received. Contemporary Physics, 20(1):119, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Migdal:1979:PSC

[941] A. B. Migdal. On the psychology of scientific creativity. Contemporary Physics, 20(2):121–148, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hockney:1979:LPC

[942] Roger W. Hockney. The large parallel computer and university research. Contemporary Physics, 20(2):149–185, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tanner:1979:AD

[943] B. K. Tanner. Antiferromagnetic domains. Contemporary Physics, 20(2): 187–210, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Murray:1979:AVP

[944] C. A. Murray. An astronomer’s view of the pole. Contemporary Physics, 20(2):211–226, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1979:NPS

[945] Sir Nevill Mott. Nobel prizes in science. Contemporary Physics, 20(2): 227–228, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1979:ECV

[946] G. R. Noakes. Einstein. a centenary volume. Contemporary Physics, 20 (2):229–230, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:SNBa

[947] Anonymous. Short notices of books. Contemporary Physics, 20(2):231– 232, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRc

[948] Anonymous. Book reviews. Contemporary Physics, 20(2):233–238, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rice:1979:EHL

[949] T. M. Rice. The electron-hole liquid. Contemporary Physics, 20(3): 241–255, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morley:1979:CMD

[950] J. G. Morley. Composite materials: Designing for structural integrity. Contemporary Physics, 20(3):257–292, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:1979:NPT

[951] F. E. Close. The new particles and their significance. Contemporary Physics, 20(3):293–313, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1979:WFP

[952] B. J. Mason. Weather forecasting as a problem in fluid dynamics. Contemporary Physics, 20(3):315–335, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1979:EHL

[953] Sir Nevill Mott. The electron-hole liquid. Contemporary Physics, 20(3): 337–338, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1979:OTA

[954] W. H. McCrea. Observational and theoretical astronomy. Contemporary Physics, 20(3):339–341, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1979:PV

[955] C. A. Taylor. The physics of vibration. Contemporary Physics, 20(3): 343–344, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRd

[956] Anonymous. Book reviews. Contemporary Physics, 20(3):345–368, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:SNBb

[957] Anonymous. Short notices of books. Contemporary Physics, 20(3):369– 371, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRe

[958] Anonymous. Books received. Contemporary Physics, 20(3):373–374, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blin-Stoyle:1979:NL

[959] R. J. Blin-Stoyle. The nucleus as a laboratory. Contemporary Physics, 20(4):377–399, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leck:1979:UHV

[960] J. H. Leck. Ultra-high vacuum measurement. Contemporary Physics, 20 (4):401–415, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kulcinski:1979:FRT

[961] G. L. Kulcinski. Fusion reactors: Their challenge to materials scientists. Contemporary Physics, 20(4):417–447, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pfister:1979:NAE

[962] G. Pfister. New aspects of electronic properties of amorphous selenium and its use in xerography. Contemporary Physics, 20(4):449–479, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Motz:1979:SRA

[963] M. Motz. Some recent advances in optics. Contemporary Physics, 20(4): 481–482, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pepper:1979:SP

[964] Michael Pepper. Semiconductor physics. Contemporary Physics, 20(4): 483–485, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gowing:1979:ERM

[965] Margaret Gowing. Essay review: My life: Recollections of a Nobel laureate, by Max Born. Contemporary Physics, 20(4):485–486, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRf

[966] Anonymous. Book reviews. Contemporary Physics, 20(4):487–498, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRg

[967] Anonymous. Books received. Contemporary Physics, 20(4):499–500, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garcia-Moliner:1979:PS

[968] F. García-Moliner. Physics in Spain. Contemporary Physics, 20(5):503– 513, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1979:MA

[969] C. A. Taylor. Musical acoustics. Contemporary Physics, 20(5):515–534, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stoneham:1979:TDS

[970] A. M. Stoneham. The theory of defects in solids. Contemporary Physics, 20(5):535–545, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Motz:1979:UFE

[971] H. Motz. Undulators and ‘free-electron lasers’. Contemporary Physics, 20(5):547–568, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nelson:1979:NPT

[972] R. L. Nelson. Nuclear power at a turning point. Contemporary Physics, 20(5):569–571, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRh

[973] Anonymous. Book reviews. Contemporary Physics, 20(5):573–576, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:SNBc

[974] Anonymous. Short notices of books. Contemporary Physics, 20(5):577– 578, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRi

[975] Anonymous. Books received. Contemporary Physics, 20(5):579–580, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:1979:O

[976] O. R. Frisch and Nevill Mott. Obituary. Contemporary Physics, 20 (6):583, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bondi:1979:GC

[977] H. Bondi. Gravitational capture. Contemporary Physics, 20(6):585– 591, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nagpal:1979:SAG

[978] O. P. Nagpal. The sources of atmospheric gravity waves. Contemporary Physics, 20(6):593–609, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Klein:1979:SDP

[979] J. Klein. The self-diffusion of polymers. Contemporary Physics, 20(6): 611–629, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Latimer:1979:REI

[980] C. J. Latimer. Recent experiments involving highly excited atoms. Contemporary Physics, 20(6):631–653, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gowing:1979:LAL

[981] Margaret Gowing. “… I have always lived very much in the present … ”. Contemporary Physics, 20(6):655–657, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tanner:1979:EM

[982] B. K. Tanner. Experimental magnetism. Contemporary Physics, 20(6): 657–659, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Goldsmith:1979:THL

[983] Maurice Goldsmith. Is there hope … ? Contemporary Physics, 20(6): 659–660, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1979:TRS

[984] W. H. McCrea. The theory of rotating stars. Contemporary Physics, 20 (6):660–661, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRj

[985] Anonymous. Books reviews. Contemporary Physics, 20(6):663–675, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:SNBd

[986] Anonymous. Short notices of books. Contemporary Physics, 20(6):677– 678, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1979:BRk

[987] Anonymous. Books received. Contemporary Physics, 20(6):679–680, 1979. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:1980:MMP

[988] Sir Rudolf Peierls. Model-making in physics. Contemporary Physics, 21 (1):3–17, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harrison:1980:MPI

[989] R. G. Harrison and S. R. Butcher. Multiple photon infrared processes in polyatomic molecules. Contemporary Physics, 21(1):19–41, 1980. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cahn:1980:MG

[990] R. W. Cahn. Metallic glasses. Contemporary Physics, 21(1):43–75, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Araujo:1980:PGC

[991] R. J. Araujo. Photochromism in glasses containing silver halides. Contemporary Physics, 21(1):77–84, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cottrell:1980:SSP

[992] Sir Alan Cottrell, F.R.S. Structural stability in physics. Contemporary Physics, 21(1):85–86, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burcham:1980:FN

[993] W. E. Burcham, F.R.S. The forces of nature. Contemporary Physics, 21 (1):86–87, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caplin:1980:CC

[994] A. D. Caplin. Clear as crystal. Contemporary Physics, 21(1):87–90, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRa

[995] Anonymous. Book reviews. Contemporary Physics, 21(1):91–93, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:SNBa

[996] Anonymous. Short notices of books. Contemporary Physics, 21(1):94, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRb

[997] Anonymous. Books received. Contemporary Physics, 21(1):95–96, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rees:1980:GCC

[998] Martin J. Rees. Gravitational collapse and cosmology. Contemporary Physics, 21(2):99–120, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gutzow:1980:ICGa

[999] I. Gutzow. Induced crystallization of glass-forming systems: A case of transient heterogeneous nucleation, part 1. Contemporary Physics, 21(2): 121–137, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1980:AA

[1000] C. A. Taylor. Architectural acoustics. Contemporary Physics, 21(2): 139–158, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Iliopoulos:1980:UTE

[1001] J. Iliopoulos. Unified theories of elementary particle interactions. Contemporary Physics, 21(2):159–183, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:1980:NST

[1002] G. H. A. Cole. Notions of space and time. Contemporary Physics, 21(2): 185–186, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRc

Anonymous:1980:SNBb

[1004] Anonymous. Short notices of books. Contemporary Physics, 21(2):191– 193, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRd

[1005] Anonymous. Books received. Contemporary Physics, 21(2):195, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vorreiter:1980:CSS

[1006] John W. Vorreiter. Cryogenic systems for spacecraft. Contemporary Physics, 21(3):201–217, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bhandari:1980:SGA

[1007] C. M. Bhandari and D. M. Rowe. Silicon–germanium alloys as hightemperature thermoelectric materials. Contemporary Physics, 21(3):219– 242, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gutzow:1980:ICGb

[1008] I. Gutzow. Induced crystallization of glass-forming systems: A case of transient heterogeneous nucleation (part II). Contemporary Physics, 21 (3):243–263, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1980:SLI

[1009] T. F. Smith and T. R. Finlayson. Superconductivity and lattice instability. Contemporary Physics, 21(3):265–275, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kendall:1980:SS

[1010] K. Kendall. Sticky solids. Contemporary Physics, 21(3):277–297, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burcham:1980:MP

[1011] W. E. Burcham, F.R.S. The matter of physics. Contemporary Physics, 21(3):299–301, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cottrell:1980:EHS

[1012] Sir Alan Cottrell, F.R.S. Effects of high-speed strain. Contemporary Physics, 21(3):301–302, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1980:GW

[1013] W. H. McCrea, F.R.S. Gravitation in waves. Contemporary Physics, 21 (3):302–305, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vick:1980:NCP

[1014] Sir Arthur Vick. Need for contemporary physics. Contemporary Physics, 21(3):305–308, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saslaw:1980:LGN

[1015] William C. Saslaw. A layman’s guide to the new universe. Contemporary Physics, 21(3):308–310, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRe

[1016] Anonymous. Book reviews. Contemporary Physics, 21(3):311–322, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:SNBc

[1017] Anonymous. Short notices of books. Contemporary Physics, 21(3):323– 325, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:LBRa

[1018] Anonymous. List of books received. Contemporary Physics, 21(3):327– 328, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

French:1980:ICP

[1019] A. P. French. The international commission on physics education. Contemporary Physics, 21(4):331–344, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hall:1980:ACD

[1020] E. T. Hall. Advances in carbon dating using high energy mass spectrometers. Contemporary Physics, 21(4):345–358, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kopal:1980:PES

[1021] Zdeněk Kopal. Planetary exploration by spacecraft. Contemporary Physics, 21(4):359–380, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1980:AVJ

[1022] B. J. Mason. The atmospheres of Venus and Jupiter. Contemporary Physics, 21(4):381–399, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Welford:1980:LSS

[1023] W. T. Welford. Laser speckle and surface roughness. Contemporary Physics, 21(4):401–412, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRf

[1024] Anonymous. Book reviews. Contemporary Physics, 21(4):413–424, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:LBRb

[1025] Anonymous. List of books received. Contemporary Physics, 21(4):425– 426, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chandrasekhar:1980:GTR

[1026] S. Chandrasekhar. The general theory of relativity: The first thirty years. Contemporary Physics, 21(5):429–449, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hayes:1980:STE

[1027] W. Hayes. Self-trapping of electronic excitations and radiolysis of ionic solids. Contemporary Physics, 21(5):451–461, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schuurmans:1980:FAN

[1028] M. F. H. Schuurmans. The fluorescence of atoms near a glass surface. Contemporary Physics, 21(5):463–482, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lynn:1980:HNH

[1029] J. E. Lynn. Helios: The new Harwell electron linear accelerator, and its scientific programme. Contemporary Physics, 21(5):483–500, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicholas:1980:QTS

[1030] R. J. Nicholas. Quantum transport in semiconductors. Contemporary Physics, 21(5):501–521, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRg

[1031] Anonymous. Book reviews. Contemporary Physics, 21(5):523–533, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:SNBd

[1032] Anonymous. Short notices of books. Contemporary Physics, 21(5):533– 535, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRh

[1033] Anonymous. Books received. Contemporary Physics, 21(5):536, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vick:1980:DKM

[1034] Arthur Vick. Dr. Kurt Mendelssohn, F.R.S. Contemporary Physics, 21 (6):539–540, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Svanberg:1980:LPA

[1035] S. Svanberg. Lasers as probes for air and sea. Contemporary Physics, 21(6):541–576, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Domb:1980:PC

[1036] C. Domb, E. Stoll, and T. Schneider. Percolation clusters. Contemporary Physics, 21(6):577–592, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hammersley:1980:PTR

[1037] J. M. Hammersley and D. J. A. Welsh. Percolation theory and its ramifications. Contemporary Physics, 21(6):593–605, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Petley:1980:US

[1038] B. W. Petley. The ubiquitous SQUID. Contemporary Physics, 21(6): 607–630, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1980:ERU

[1039] K. F. Smith. Essay review: Ultra-cold neutrons. Contemporary Physics, 21(6):631–650, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casimir:1980:ERP

[1040] H. B. G. Casimir. Essay review: Peter Kapitza, by H. B. G. Casimir. Contemporary Physics, 21(6):651–653, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gowing:1980:AW

[1041] Margaret Gowing. Atoms of war. Contemporary Physics, 21(6):653– 656, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1980:LD

[1042] Charles A. Taylor. Light delight. Contemporary Physics, 21(6):656– 657, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRi

[1043] Anonymous. Book reviews. Contemporary Physics, 21(6):659–665, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1980:BRj

[1044] Anonymous. Books received. Contemporary Physics, 21(6):667–668, 1980. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cosslett:1981:DEMa

[1045] V. E. Cosslett. The development of electron microscopy and related techniques at the Cavendish Laboratory, Cambridge, 1946–79 part 1 1946– 60. Contemporary Physics, 22(1):3–36, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Swift:1981:ED

[1046] D. A. Swift. The electrical discharge. Contemporary Physics, 22(1): 37–60, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carroll:1981:LPP

[1047] P. K. Carroll and E. T. Kennedy. Laser-produced plasmas. Contemporary Physics, 22(1):61–96, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowlt:1981:URM

[1048] C. Bowlt. The use of radioisotopes in medical diagnosis. Contemporary Physics, 22(1):97–115, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bhattacharyya:1981:TSM

[1049] Anjan Bhattacharyya. A tunnelling-state model for the specific heat capacity and thermal conductivity of amorphous materials at low temperatures. Contemporary Physics, 22(1):117–127, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Penkett:1981:SA

[1050] Dr. S. A. Penkett. Something in the air. Contemporary Physics, 22(1): 129–130, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRa

[1051] Anonymous. Book reviews. Contemporary Physics, 22(1):131–141, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRb

[1052] Anonymous. Books received. Contemporary Physics, 22(1):143, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cosslett:1981:DEMb

[1053] V. E. Cosslett. The development of electron microscopy and related techniques at the Cavendish Laboratory, Cambridge, 1946–79 (part II). Contemporary Physics, 22(2):147–182, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1981:DCL

[1054] E. J. N. Wilson. The design and commissioning of large particle accelerators. Contemporary Physics, 22(2):183–213, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tabor:1981:SST

[1055] David Tabor. Solid surfaces: Their atomic, electronic and macroscopic properties. Contemporary Physics, 22(2):215–234, 1981. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Goldsmith:1981:OS

[1056] J. E. M. Goldsmith and J. E. Lawler. Optogalvanic spectroscopy. Contemporary Physics, 22(2):235–248, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1981:ERE

[1057] Sir Nevill Mott. Essay review: Electrons at the Fermi surface. A Festschrift to honour Professor David Shoenberg, edited by M. Springford. Contemporary Physics, 22(2):249–250, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1981:AS

[1058] Boris V. Barlow. The art of sensorship. Contemporary Physics, 22(2): 250–252, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carr:1981:WWG

[1059] B. J. Carr. A weight-watcher’s guide to the galaxy. Contemporary Physics, 22(2):252–254, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRc

[1060] Anonymous. Book reviews. Contemporary Physics, 22(2):255–258, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:SNBa

[1061] Anonymous. Short notices of books. Contemporary Physics, 22(2):259, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRd

[1062] Anonymous. Books received. Contemporary Physics, 22(2):261, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pease:1981:EBL

[1063] R. F. W. Pease. Electron beam lithography. Contemporary Physics, 22 (3):265–290, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Landshoff:1981:LEP

[1064] Peter Landshoff. The large electron positron collider (LEP) for particle physics. Contemporary Physics, 22(3):291–298, 1981. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Macgillivray:1981:SAM

[1065] J. C. Macgillivray and M. S. Feld. Superradiance in atoms and molecules. Contemporary Physics, 22(3):299–310, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lightenrerg:1981:QC

[1066] D. B. Lightenrerg. Quantum chromodynamics. Contemporary Physics, 22(3):311–334, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blutndell:1981:EHR

[1067] D. J. Blutndell. Earthquake hazard and risk mitigation. Contemporary Physics, 22(3):335–347, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Farmer:1981:ARR

[1068] F. R. Farmer. The assessment of risk in relation to major hazards, with particular reference to nuclear reactors. Contemporary Physics, 22(3): 349–360, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1981:ERH

[1069] P. E. Hodgson. Essay review: Heavy ions. Nuclear Reactions with Heavy Ions, by Reiner Bass. Heavy Ion Collisions, edited by R. Bock. Deep Inelastic and Fusion Reactions with Heavy Ions, edited by W. von Oertzen. Contemporary Physics, 22(3):361–363, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1981:CC

[1070] P. C. W. Davies. On clusters and conferences. Contemporary Physics, 22(3):364–366, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRe

[1071] Anonymous. Book reviews. Contemporary Physics, 22(3):367–370, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRf

[1072] Anonymous. Books received. Contemporary Physics, 22(3):371–372, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weisskopf:1981:FCP

[1073] Victor F. Weisskopf. The formation of Cooper pairs and the nature of superconducting currents. Contemporary Physics, 22(4):375–395, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:1981:APP

[1074] G. H. A. Cole. Aspects of the physics of planetary interiors. Contemporary Physics, 22(4):397–424, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rochester:1981:CRU

[1075] G. D. Rochester and K. E. Turver. Cosmic rays of ultra-high energy. Contemporary Physics, 22(4):425–450, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicholson:1981:NSE

[1076] L. Kay Nicholson. The neutron spin-echo spectrometer: A new high resolution technique in neutron scattering. Contemporary Physics, 22 (4):451–475, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRg

Anonymous:1981:BRh

[1078] Anonymous. Books received. Contemporary Physics, 22(4):483–484, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Series:1981:TYO

[1079] G. W. Series. Thirty years of optical pumping. Contemporary Physics, 22(5):487–509, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1981:NCM

[1080] P. E. Hodgson. Nuclear charge and matter distributions. Contemporary Physics, 22(5):511–532, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1981:NPA

[1081] F. Graham Smith. New prospects for astronomy. Contemporary Physics, 22(5):533–544, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Allibone:1981:PYS

[1082] T. E. Allibone. Power in the year 2000. Some reminiscences and forecasts. Contemporary Physics, 22(5):545–558, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:1981:ERI

[1083] G. H. A. Cole. Essay review: Inside information. The Interiors of the Planets, by A. H. Cook. Contemporary Physics, 22(5):559–560, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:1981:SF

[1084] Dr. B. J. T. Jones. The spacious firmament. Contemporary Physics, 22 (5):560–562, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilchrist:1981:WS

[1085] Dr. A. Gilchrist. A wind in season. Contemporary Physics, 22(5):562– 564, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRi

[1086] Anonymous. Book reviews. Contemporary Physics, 22(5):565–589, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:SNBb

[1087] Anonymous. Short notices of books. Contemporary Physics, 22(5):590– 591, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRj

[1088] Anonymous. Books received. Contemporary Physics, 22(5):593–594, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Huang:1981:LRM

[1089] Kun Huang. Lattice relaxation and multiphonon transitions. Contemporary Physics, 22(6):599–612, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Johari:1981:SI

[1090] U. P. Johari. The spectrum of ice. Contemporary Physics, 22(6):613– 642, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anthes:1981:TCS

[1091] Richard A. Anthes. Tropical cyclones: Structure, computer simulation models, and operational prediction. Contemporary Physics, 22(6):643– 680, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gowing:1981:ERO

[1092] Margaret Gowing. Essay review: Oppenheimer and Los Alamos. Robert Oppenheimer: Letters and Recollections, edited by Alice Kimball Smith and Charles Weiner. J. Robert Oppenheimer, “Shatterer of Worlds”, by Peter Goodchild. Reminiscences of Los Alamos, 1943–1945, edited by Lawrence Badash, Joseph O. Hirschfelder and Herbert P. Broida. Contemporary Physics, 22(6):681–686, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1981:IR

[1093] G. R. Noakes. Interpretation and review. Contemporary Physics, 22(6): 687–689, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRk

[1094] Anonymous. Book reviews. Contemporary Physics, 22(6):691–703, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1981:BRl

[1095] Anonymous. Books received. Contemporary Physics, 22(6):705–706, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frank:1982:SC

[1096] F. C. Frank. Snow crystals. Contemporary Physics, 23(1):3–22, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Panofsky:1982:PSC

[1097] W. K. H. Panofsky. Particle substructure. A common theme of discovery in this century. Contemporary Physics, 23(1):23–44, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1982:STA

[1098] Boris V. Barlow. Space for the telescope. An account of NASA’s projected space telescope. Contemporary Physics, 23(1):45–63, 1982. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shanks:1982:PDA

[1099] Ian A. Shanks. The physics and display applications of liquid crystals. Contemporary Physics, 23(1):65–91, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1982:ERU

[1100] W. H. McCrea. Essay review: The universe in a nutshell. Cosmology, by Edward R. Harrison. Contemporary Physics, 23(1):93–94, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:1982:VED

[1101] G. H. A. Cole. The very earth did shake. Contemporary Physics, 23 (1):95–96, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRa

[1102] Anonymous. Book reviews. Contemporary Physics, 23(1):97–107, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRb

[1103] Anonymous. Books received. Contemporary Physics, 23(1):109–110, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kinchin:1982:NRS

[1104] G. H. Kinchin. Nuclear reactor safety: Physics and engineering aspects. Contemporary Physics, 23(2):113–128, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dose:1982:MH

[1105] V. Dose. Metastable hydrogen. Contemporary Physics, 23(2):129–140, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ford:1982:SG

[1106] Peter J. Ford. Spin glasses. Contemporary Physics, 23(2):141–168, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salomaa:1982:MNI

[1107] M. Salomaa. The mobility of negative ions in superfluid ^{3}He. Contemporary Physics, 23(2):169–188, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Llewellyn-smith:1982:PPT

[1108] C. H. Llewellyn-smith. Particle physics theory. Contemporary Physics, 23(2):189–190, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRc

[1109] Anonymous. Book reviews. Contemporary Physics, 23(2):191–202, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRd

[1110] Anonymous. Books received. Contemporary Physics, 23(2):203–204, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Radhakrishnan:1982:NP

[1111] V. Radhakrishnan. On the nature of pulsars. Contemporary Physics, 23 (3):207–231, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Geddes:1982:CRA

[1112] J. Geddes. Chemical reactions of atoms with molecules in colliding beams. Contemporary Physics, 23(3):233–255, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hughes:1982:C

[1113] David W. Hughes. Comets. Contemporary Physics, 23(3):257–283, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibson:1982:LCF

[1114] A. F. Gibson. Lasers for compression and fusion. Contemporary Physics, 23(3):285–297, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRe

[1115] Anonymous. Book reviews. Contemporary Physics, 23(3):299–306, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRf

[1116] Anonymous. Books received. Contemporary Physics, 23(3):307–308, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1982:PRY

[1117] B. J. Mason. Personal reflections on 35 years of cloud seeding. Contemporary Physics, 23(4):311–327, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1982:SBP

[1118] R. H. Williams. The Schottky barrier problem. Contemporary Physics, 23(4):329–351, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Conlon:1982:DSR

[1119] T. W. Conlon. Doping surfaces with radioactive atoms — for research and industry. Contemporary Physics, 23(4):353–369, 1982. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pennycook:1982:HRE

[1120] S. J. Pennycook. High resolution electron microscopy and microanalysis. Contemporary Physics, 23(4):371–400, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:ERI

[1121] Anonymous. Essay review: The Isotropic Universe, by D. R. Raine. Contemporary Physics, 23(4):401–402, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1982:TT

[1122] P. C. W. Davies. Testing the theory. Contemporary Physics, 23(4): 402–403, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paton:1982:PN

[1123] J. E. Paton. The permissiveness of nature. Contemporary Physics, 23(4): 403–405, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kibble:1982:UU

[1124] T. W. B. Kibble. The underlying unity. Contemporary Physics, 23(4): 406–408, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRg

[1125] Anonymous. Book reviews. Contemporary Physics, 23(4):409–414, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRh

[1126] Anonymous. Books received. Contemporary Physics, 23(4):415–416, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Webb:1982:TTE

[1127] D. J. Webb. Tides and tidal energy. Contemporary Physics, 23(5):419– 442, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishikawa:1982:CMS

[1128] M. Ishikawa. Can magnetism and superconductivity coexist? Contemporary Physics, 23(5):443–468, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hurd:1982:VMO

[1129] C. M. Hurd. Varieties of magnetic order in solids. Contemporary Physics, 23(5):469–493, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1982:HIR

[1130] P. E. Hodgson. Heavy ion reactions. Contemporary Physics, 23(5):495– 512, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cosslett:1982:ERT

[1131] V. E. Cosslett. Essay review: Thirty-five years on. Electron Microscopy and Analysis, 1981, edited by M. J. Goringe. Contemporary Physics, 23 (5):513–514, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1982:ECG

[1132] W. H. McCrea. Ether its comings and goings. Contemporary Physics, 23(5):515–516, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joffe:1982:FI

[1133] A. D. Joffe. Foreign invaders. Contemporary Physics, 23(5):517–519, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRi

[1134] Anonymous. Book reviews. Contemporary Physics, 23(5):521–526, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRj

[1135] Anonymous. Books received. Contemporary Physics, 23(5):527–528, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1982:RSA

[1136] W. H. McCrea. Recollections of Sir Arthur Eddington O.M., F.R.S., 28 December 1882–22 November 1944. Contemporary Physics, 23(6): 531–540, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bacon:1982:NBS

[1137] G. E. Bacon. Neutron beams for the study of condensed matter: A view of the first half-century. Contemporary Physics, 23(6):541–552, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moriyasu:1982:RGT

[1138] K. Moriyasu. The renaissance of gauge theory. Contemporary Physics, 23(6):553–581, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Friedel:1982:OST

[1139] J. Friedel and D. Jérome. Organic superconductors: The (TMTSF)_{2} x family. Contemporary Physics, 23(6):583–624, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Series:1982:ERR

[1140] G. W. Series. Essay review: Recycling radiation. Dissipative Systems in Quantum Optics, edited by R. Bonifacio. Contemporary Physics, 23(6): 625–626, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRk

[1141] Anonymous. Book review. Contemporary Physics, 23(6):627–654, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1982:BRl

[1142] Anonymous. Books received. Contemporary Physics, 23(6):655–656, 1982. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hoch:1983:KCE

[1143] P. K. Hoch. A key concept from the electron theory of metals: history of the Fermi surface 1933–60. Contemporary Physics, 24(1):3–23, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pieranski:1983:CC

[1144] Pawel Pieranski. Colloidal crystals. Contemporary Physics, 24(1):25– 73, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hill:1983:DBC

[1145] R. M. Hill and A. K. Jonscher. The dielectric behaviour of condensed matter and its many-body interpretation. Contemporary Physics, 24(1): 75–110, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wohlfarth:1983:ERH

[1146] E. P. Wohlfarth. Essay review: High pressure physics. Physics of Solids under High Pressure, edited by J. S. Schilling and R. N. Shelton. Contemporary Physics, 24(1):111–114, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRa

[1147] Anonymous. Book review. Contemporary Physics, 24(1):115–123, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRb

[1148] Anonymous. Books received. Contemporary Physics, 24(1):125–126, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Robson:1983:ESB

[1149] J. M. Robson. Experimental studies of the beta decay of the neutron: a historical review. Contemporary Physics, 24(2):129–141, ???? 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812

(electronic). URL http://www.tandfonline.com/doi/abs/10.1080/
00107518308210669.

Spetser:1983:NP

[1150] D. Spetser. Newton’s Principia. Contemporary Physics, 24(2):143–162, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCarthy:1983:OME

[1151] I. E. McCarthy and E. Weigold. Observing the motion of electrons in atoms and molecules. Contemporary Physics, 24(2):163–184, 1983. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hamzeht:1983:NMO

[1152] F. M. Hamzeht. Neutrino masses and oscillations. Contemporary Physics, 24(2):185–194, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1983:KPP

[1153] Sir Nevill Mott, F.R.S. From Karl Popper’s ‘postscript’. Contemporary Physics, 24(2):195–197, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paton:1983:AD

[1154] J. E. Paton. Accident or design? Contemporary Physics, 24(2):198– 199, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burcham:1983:NBD

[1155] W. E. Burcham, F.R.S. Nuclear beta-decay. Contemporary Physics, 24 (2):199–203, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hoch:1983:BDY

[1156] Paul Hoch. Bohr’s decisive years. Contemporary Physics, 24(2):203– 204, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:1983:CP

[1157] G. H. A. Cole. Comparative planetology. Contemporary Physics, 24(2): 205–206, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRc

[1158] Anonymous. Book reviews. Contemporary Physics, 24(2):207–216, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRd

[1159] Anonymous. Books received. Contemporary Physics, 24(2):217–218, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:1983:EDN

[1160] Rudolf Peierls. The early days of neutrino physics. Contemporary Physics, 24(3):221–228, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Challis:1983:PS

[1161] L. J. Challis. Phonon spectroscopy. Contemporary Physics, 24(3):229– 250, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hannaford:1983:SSA

[1162] P. Hannaford. Spectroscopy with sputtered atoms. Contemporary Physics, 24(3):251–270, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Forty:1983:SCC

[1163] A. J. Forty. The structure and chemistry of crystal surfaces. Contemporary Physics, 24(3):271–299, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1983:SF

[1164] Boris V. Barlow. Space flight. Contemporary Physics, 24(3):301–302, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Polkinghorne:1983:ERS

[1165] Revd. Dr. J. C. Polkinghorne, F.R.S. Essay review: The science and the life of Albert Einstein: Subtle Is The Lord: the Science and the Life of Albert Einstein, by A. Pais. Contemporary Physics, 24(3):302– 305, ???? 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishikawa:1983:STC

[1166] M. Ishikawa. Superconductivity in ternary compounds. Contemporary Physics, 24(3):305–307, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRe

Anonymous:1983:BRf

[1168] Anonymous. Books received. Contemporary Physics, 24(3):315–316, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blackman:1983:LSH

[1169] M. Blackman. The lodestone: A survey of the history and the physics. Contemporary Physics, 24(4):319–331, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lugiato:1983:OB

[1170] Luigi A. Lugiato. Optical bistability. Contemporary Physics, 24(4):333– 371, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

March:1983:QMA

[1171] N. H. March. Quantum mechanics has accounted for a large part of physics and the whole of chemistry. True or false? Contemporary Physics, 24(4):373–387, ???? 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobbs:1983:SHT

[1172] E. Roland Dobbs. Superfluid helium three. Contemporary Physics, 24(4): 389–413, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRg

[1173] Anonymous. Book reviews. Contemporary Physics, 24(4):415–422, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRh

[1174] Anonymous. Books received. Contemporary Physics, 24(4):423–424, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Irvine:1983:CLP

[1175] J. M. Irvine. The constancy of the laws of physics in the light of prehistoric nuclear reactors. Contemporary Physics, 24(5):427–437, September 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812

(electronic). URL http://www.tandfonline.com/doi/full/10.1080/
00107518308210695.

Field:1983:HSP

[1176] J. E. Field. High-speed photography. Contemporary Physics, 24(5): 439–459, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Boyd:1983:RLB

[1177] Ian W. Boyd. A review of laser beam applications for processing silicon. Contemporary Physics, 24(5):461–490, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1983:EMN

[1178] P. E. Hodgson. Effective mass in nuclei. Contemporary Physics, 24(5): 491–503, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Doust:1983:UP

[1179] T. A. M. Doust. Ultrafast phenomena. Contemporary Physics, 24(5): 505–507, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1983:HR

[1180] Sir Nevill Mott, F.R.S. Haphazard reality. Contemporary Physics, 24(5): 508–509, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1983:MKP

[1181] Sir Nevill Mott, F.R.S. More from Karl Popper’s ‘postscript’. Contemporary Physics, 24(5):509–510, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heavens:1983:LTR

[1182] O. S. Heavens. Longer than red. Contemporary Physics, 24(5):510– 511, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1983:OFU

[1183] W. H. McCrea, F.R.S. The once and future universe. Contemporary Physics, 24(5):511–512, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRi

[1184] Anonymous. Book reviews. Contemporary Physics, 24(5):513–529, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRj

[1185] Anonymous. Books received. Contemporary Physics, 24(5):531–532, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rhead:1983:MM

[1186] G. E. Rhead. Metal monolayers. Contemporary Physics, 24(6):535– 559, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Enderby:1983:NSS

[1187] J. E. Enderby. Neutron scattering and the structure of liquids. Contemporary Physics, 24(6):561–575, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Potter:1983:DTC

[1188] A. R. Potter. Digital techniques in communications. Contemporary Physics, 24(6):577–589, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Halzen:1983:SMQ

[1189] Francis Halzen. The search for matter in its quark-gluon phase. Contemporary Physics, 24(6):591–622, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weinstock:1983:NPI

[1190] Robert Weinstock. Newton’s Principia and inverse-square orbits. Contemporary Physics, 24(6):623–625, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barlow:1983:ERO

[1191] B. V. Barlow. Essay review: Optical rotation sensors. Fiber-Optic Rotation Sensors and Related Technologies, edited by S. Ezekiel and H. J. Arditty. Contemporary Physics, 24(6):627–628, 1983. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kilmister:1983:TCM

[1192] C. W. Kilmister. “too complex to model?”. Contemporary Physics, 24(6): 628–629, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1983:IO

[1193] C. A. Taylor. Integrated optics. Contemporary Physics, 24(6):629–631, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casimir:1983:HDQ

[1194] H. B. G. Casimir. The historical development of quantum theory. Contemporary Physics, 24(6):631–634, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRk

[1195] Anonymous. Book reviews. Contemporary Physics, 24(6):635–638, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1983:BRl

[1196] Anonymous. Books received. Contemporary Physics, 24(6):639–641, 1983. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Series:1984:LS

[1197] G. W. Series. Laser spectroscopy. Contemporary Physics, 25(1):3–29, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kennedy:1984:PIL

[1198] E. T. Kennedy. Plasmas and intense laser light. Contemporary Physics, 25(1):31–58, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weaire:1984:SCS

[1199] D. Weaire and N. Rivier. Soap, cells and statistics — random patterns in two dimensions. Contemporary Physics, 25(1):59–99, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRa

Anonymous:1984:BRb

[1201] Anonymous. Books received. Contemporary Physics, 25(1):105–106, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roberts:1984:LBF

[1202] G. G. Roberts. Langmuir–Blodgett films. Contemporary Physics, 25(2): 109–128, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nikoonahad:1984:RAH

[1203] Mehrdad Nikoonahad. Recent advances in high resolution acoustic microscopy. Contemporary Physics, 25(2):129–158, 1984. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Banna:1984:ESE

[1204] M. Salim Banna. The electronic structure of the elements from gasphase X-ray photoelectron spectroscopy. Contemporary Physics, 25(2): 159–176, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1984:PSH

[1205] Nevill Mott. Professor Sir Harrie Massey. Contemporary Physics, 25(2): 177–178, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burcham:1984:DNR

[1206] W. E. Burcham, F.R.S. Direct nuclear reactions. Contemporary Physics, 25(2):179–183, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paton:1984:IC

[1207] J. E. Paton. Initial conditions. Contemporary Physics, 25(2):183–186, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRc

Anonymous:1984:BRd

[1209] Anonymous. Books received. Contemporary Physics, 25(2):207–208, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wiser:1984:ERS

[1210] Nathan Wiser. The electrical resistivity of the simple metals. Contemporary Physics, 25(3):211–249, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ourmazd:1984:EPD

[1211] A. Ourmazd. The electrical properties of dislocations in semiconductors. Contemporary Physics, 25(3):251–268, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gubbins:1984:EMF

[1212] David Gubbins. The Earth’s magnetic field. Contemporary Physics, 25 (3):269–290, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wohlfarth:1984:HFM

[1213] Professor E. P. Wohlfarth. High field magnetism. Contemporary Physics, 25(3):291–294, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rochester:1984:EPP

[1214] G. D. Rochester. Early particle physics. Contemporary Physics, 25(3): 294–300, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hammersley:1984:RS

[1215] J. M. Hammersley, F.R.S. The rough with the smooth. Contemporary Physics, 25(3):300–302, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRe

[1216] Anonymous. Book reviews. Contemporary Physics, 25(3):303–307, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bleaney:1984:MSOa

[1217] B. Bleaney. Microwave spectroscopy in Oxford: The first decade. Part I: Microwave gas spectroscopy. Contemporary Physics, 25(4):315–329, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haines:1984:TP

[1218] M. G. Haines. Tokamak physics. Contemporary Physics, 25(4):331– 353, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vessot:1984:TGR

[1219] Robert F. C. Vessot. Tests of gravitation and relativity. Contemporary Physics, 25(4):355–380, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

North:1984:EBM

[1220] A. C. T. North and J. E. Lydon. The evolution of biological macromolecules. Contemporary Physics, 25(4):381–393, 1984. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vick:1984:CA

[1221] Sir Arthur Vick. Cockcroft and the atom. Contemporary Physics, 25(4): 395–399, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRf

[1222] Anonymous. Book reviews. Contemporary Physics, 25(4):401–412, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRg

[1223] Anonymous. Books received. Contemporary Physics, 25(4):413–414, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:E

[1224] Anonymous. Erratum. Contemporary Physics, 25(4):415, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bleaney:1984:MSOb

[1225] B. Bleaney. Microwave spectroscopy in Oxford: The first decade. Part II. Electron magnetic resonance. Contemporary Physics, 25(5):419–440, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailes:1984:NI

[1226] D. R. Bailes and D. I. Bryant. NMR imaging. Contemporary Physics, 25(5):441–475, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Foster:1984:CV

[1227] David H. Foster. Colour vision. Contemporary Physics, 25(5):477–497, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1984:DS

[1228] G. R. Noakes. Into the depths of space. Contemporary Physics, 25 (5):499, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frs:1984:NSG

[1229] G. D. Rochester Frs. A non-specialist’s guide to particle physics. Contemporary Physics, 25(5):500–501, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRh

[1230] Anonymous. Book reviews. Contemporary Physics, 25(5):503–511, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRi

[1231] Anonymous. Books received. Contemporary Physics, 25(5):513–514, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Noakes:1984:GE

[1232] G. R. Noakes. Guest editorial. Contemporary Physics, 25(6):517, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Webster:1984:MSM

[1233] A. S. Webster and M. S. Longair. Millimetre and sub-millimetre astronomy. Contemporary Physics, 25(6):519–534, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berge:1984:RBC

[1234] P. Bergé and M. Dubois. Rayleigh–Bénard convection. Contemporary Physics, 25(6):535–582, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:1984:SCH

[1235] A. J. Leggett. Schrödinger’s cat and her laboratory cousins. Contemporary Physics, 25(6):583–598, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowden:1984:OP

[1236] K. F. Bowden. Ocean physics. Contemporary Physics, 25(6):599–602, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yapp:1984:SL

[1237] W. Brunsdon Yapp. On size and life. Contemporary Physics, 25(6): 603–606, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRj

[1238] Anonymous. Book reviews. Contemporary Physics, 25(6):607–616, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1984:BRk

[1239] Anonymous. Books received. Contemporary Physics, 25(6):617–618, 1984. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eisner:1985:IPB

[1240] D. A. Eisner and S. C. Wray. Ion pumping in biological membranes. Contemporary Physics, 26(1):3–21, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Callaway:1985:LLNa

[1241] David J. E. Callaway. Lattices for laymen: A non-specialist’s introduction to lattice gauge theory part 1. Contemporary Physics, 26(1): 23–48, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Domb:1985:CPB

[1242] C. Domb. Critical phenomena: A brief historical survey. Contemporary Physics, 26(1):49–72, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paton:1985:QSF

[1243] J. E. Paton. From QED to sci-fi solid crud. Contemporary Physics, 26 (1):73–75, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Woodruff:1985:DIE

[1244] D. P. Woodruff. Desorption induced by electronic transitions. Contemporary Physics, 26(1):76–78, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRa

[1245] Anonymous. Book reviews. Contemporary Physics, 26(1):79–91, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRb

[1246] Anonymous. Books received. Contemporary Physics, 26(1):92, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Callaway:1985:LLNb

[1247] David J. E. Callaway. Lattices for laymen: A non-specialist’s introduction to lattice gauge theory part II. Contemporary Physics, 26(2): 95–123, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scofield:1985:DPN

[1248] Christopher L. Scofield and Leon N. Cooper. Development and properties of neural networks. Contemporary Physics, 26(2):125–145, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kessler:1985:COC

[1249] John O. Kessler. Co-operative and concentrative phenomena of swimming micro-organisms. Contemporary Physics, 26(2):147–166, 1985. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charap:1985:QTG

[1250] Professor J. M. Charap. Quantum theory of gravity. Contemporary Physics, 26(2):167–168, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Webster:1985:CW

[1251] Charles Webster. Christopher Wren. Contemporary Physics, 26(2):169– 170, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRc

[1252] Anonymous. Book reviews. Contemporary Physics, 26(2):171–197, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRd

[1253] Anonymous. Books received. Contemporary Physics, 26(2):199–200, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matt:1985:ENC

[1254] Nevill Matt. Electrons in non-crystalline materials. The last twenty five years. Contemporary Physics, 26(3):203–215, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kalmus:1985:ETP

[1255] P. I. P. Kalmus. Experimental techniques in particle physics. Contemporary Physics, 26(3):217–239, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Werth:1985:TI

[1256] G. Werth. Trapped ions. Contemporary Physics, 26(3):241–256, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pepper:1985:LQS

[1257] M. Pepper. Localization and quantization in silicon inversion iayers. Contemporary Physics, 26(3):257–293, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gough:1985:PV

[1258] Dr. W. Gough. The physics of vibration. Contemporary Physics, 26(3): 295–296, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1985:ME

[1259] W. H. McCrea. Maxwell’s equations. Contemporary Physics, 26(3): 297, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRe

[1260] Anonymous. Book reviews. Contemporary Physics, 26(3):299–306, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRf

[1261] Anonymous. Books received. Contemporary Physics, 26(3):307–308, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stedman:1985:RIT

[1262] G. E. Stedman. Ring interferometric tests of classical and quantum gravity. Contemporary Physics, 26(4):311–332, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aitchison:1985:NPV

[1263] I. J. R. Aitchison. Nothing’s plenty the vacuum in modern quantum field theory. Contemporary Physics, 26(4):333–391, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mccrea:1985:CEB

[1264] W. H. Mccrea. Chandrasekhar on Eddington and on black holes. Contemporary Physics, 26(4):393–396, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbour:1985:ERG

[1265] Dr. Julian Barbour. Essay review: Galileo, free fall, and the law of inertia. From Falling Bodies to Radio Waves: Classical Physicists and their Discoveries, by Emilio Segrè. Contemporary Physics, 26(4):397– 405, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRg

[1266] Anonymous. Book reviews. Contemporary Physics, 26(4):407–414, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRh

[1267] Anonymous. Books received. Contemporary Physics, 26(4):415–417, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:CP

[1268] Anonymous. Conducting polymers. Contemporary Physics, 26(5):419– 420, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hayes:1985:CP

[1269] W. Hayes. Conducting polymers. Contemporary Physics, 26(5):421– 441, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Camparo:1985:DLA

[1270] J. C. Camparo. The diode laser in atomic physics. Contemporary Physics, 26(5):443–477, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pappu:1985:ISU

[1271] S. V. Pappu. Information storage using alkali halide crystals. Contemporary Physics, 26(5):479–497, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1985:OPI

[1272] Professor Charles Taylor. The optical papers of Isaac Newton. Contemporary Physics, 26(5):499–501, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1985:UNS

[1273] Sir William McCrea. Under Newton’s shadow. Contemporary Physics, 26(5):503–504, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRi

[1274] Anonymous. Book reviews. Contemporary Physics, 26(5):505–515, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRj

[1275] Anonymous. Books received. Contemporary Physics, 26(5):517–518, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Conlon:1985:NPM

[1276] T. W. Conlon. Nuclear physics for materials technology accelerator based methods? Contemporary Physics, 26(6):521–558, 1985. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rieder:1985:SIA

[1277] K. H. Rieder. Surface investigations with atomic beams. Contemporary Physics, 26(6):559–578, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Venkataraman:1985:CSA

[1278] G. Venkataraman and Debendranath Sahoo. Curved space and amorphous structures. Part I. Geometric models. Contemporary Physics, 26 (6):579–615, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Irvine:1985:KPP

[1279] Professor J. M. Irvine. Key problems in physics and astrophysics. Contemporary Physics, 26(6):617–620, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRk

[1280] Anonymous. Book reviews. Contemporary Physics, 26(6):621–630, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1985:BRl

[1281] Anonymous. Books received. Contemporary Physics, 26(6):631–632, 1985. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Venkataraman:1986:CSA

[1282] G. Venkataraman and Debendranath Sahoo. Curved space and amorphous structures. Part II. Gauge theories. Contemporary Physics, 27 (1):3–35, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gill:1986:CDW

[1283] J. C. Gill and H. H. Wills. Charge-density wave transport. Contemporary Physics, 27(1):37–59, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Woodhouse:1986:PV

[1284] J. Woodhouse. The physics of the violin. Contemporary Physics, 27 (1):61–62, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kelly:1986:HES

[1285] Dr. M. J. Kelly. Hot electrons in semiconductors. Contemporary Physics, 27(1):63–67, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRa

[1286] Anonymous. Book reviews. Contemporary Physics, 27(1):68–86, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRb

[1287] Anonymous. Book received. Contemporary Physics, 27(1):87–88, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davis:1986:H

[1288] D. H. Davis. Hypernuclei. Contemporary Physics, 27(2):91–116, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vandersande:1986:TCI

[1289] J. W. Vandersande and C. Wood. The thermal conductivity of insulators and semiconductors. Contemporary Physics, 27(2):117–144, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duesbery:1986:MPD

[1290] M. S. Duesbery. The mechanical properties of the dislocation core. Contemporary Physics, 27(2):145–168, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Irvine:1986:PDM

[1291] Professor J. M. Irvine. Physics of dense matter. Contemporary Physics, 27(2):169–172, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yoffe:1986:ODC

[1292] Dr. A. D. Yoffe. One-dimensional compounds. Contemporary Physics, 27(2):173–175, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRc

[1293] Anonymous. Book reviews. Contemporary Physics, 27(2):176–196, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRd

[1294] Anonymous. Books received. Contemporary Physics, 27(2):197–200, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Margoninski:1986:PDS

[1295] Y. Margoninski. Photoelectron diffraction and surface science. Contemporary Physics, 27(3):203–240, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Townsend:1986:III

[1296] P. D. Townsend. Ion implantation — an introduction. Contemporary Physics, 27(3):241–256, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Atkins:1986:ERC

[1297] Dr. P. W. Atkins. Entropy in relation to complete knowledge. Contemporary Physics, 27(3):257–259, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tuck:1986:DCS

[1298] Dr. Brian Tuck. Diffusion in crystalline solids. Contemporary Physics, 27(3):261–263, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRe

[1299] Anonymous. Book reviews. Contemporary Physics, 27(3):265–284, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRf

[1300] Anonymous. Books received. Contemporary Physics, 27(3):285–288, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Watkins:1986:DWB

[1301] Peter M. Watkins. Discovery of the w and 2 bosons. Contemporary Physics, 27(4):291–324, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Longair:1986:UPP

[1302] M. S. Longair. The universe — present, past and future. Contemporary Physics, 27(4):325–343, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rauch:1986:NIT

[1303] H. Rauch. Neutron interferometric tests of quantum mechanics. Contemporary Physics, 27(4):345–360, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1986:DS

[1304] Dr. R. C. Smith. Distances in space. Contemporary Physics, 27(4): 361–362, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1986:GUT

[1305] Professor J. C. Taylor. Grand unified theories. Contemporary Physics, 27(4):363–365, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRg

[1306] Anonymous. Book reviews. Contemporary Physics, 27(4):367–388, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRh

[1307] Anonymous. Books received. Contemporary Physics, 27(4):389–392, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keldysh:1986:EHL

[1308] L. V. Keldysh. The electron-hole liquid in semiconductors. Contemporary Physics, 27(5):395–428, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Exerowa:1986:HMS

[1309] D. Exerowa and D. Kashchiev. Hole-mediated stability and permeability of bilayers. Contemporary Physics, 27(5):429–461, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1986:NWP

[1310] B. J. Mason. Numerical weather prediction. Contemporary Physics, 27 (5):463–472, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Russell:1986:PS

[1311] Dr. G. J. Russell. Polycrystalline semiconductors. Contemporary Physics, 27(5):473–477, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRi

[1312] Anonymous. Book reviews. Contemporary Physics, 27(5):479–491, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRj

[1313] Anonymous. Books received. Contemporary Physics, 27(5):493–496, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Browning:1986:URM

[1314] K. A. Browning. Uses of radar in meteorology. Contemporary Physics, 27(6):499–517, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hayes:1986:P

[1315] W. Hayes. Premelting. Contemporary Physics, 27(6):519–532, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Maher:1986:DFN

[1316] K. P. Maher and J. F. Malone. Digital fluoroscopy: A new development in medical imaging. Contemporary Physics, 27(6):533–552, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:1986:QTF

[1317] Dr. R. F. Bishop. Quantum theory of finite systems. Contemporary Physics, 27(6):553–555, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Colley:1986:QLN

[1318] Professor D. C. Colley. Quarks and leptons: The new elementary particles? Contemporary Physics, 27(6):557–559, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRk

[1319] Anonymous. Book reviews. Contemporary Physics, 27(6):561–586, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:SN

[1320] Anonymous. Short notices. Contemporary Physics, 27(6):587–590, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:E

[1321] Anonymous. Erratum. Contemporary Physics, 27(6):591, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1986:BRl

[1322] Anonymous. Books received. Contemporary Physics, 27(6):593–597, 1986. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pendlebury:1987:MB

[1323] J. M. Pendlebury and K. I. Smith. Molecular beams. Contemporary Physics, 28(1):3–32, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Seaborg:1987:SE

[1324] G. T. Seaborg. Superheavy elements. Contemporary Physics, 28(1): 33–48, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1987:WCR

[1325] B. J. Mason. The world climate research programme. Contemporary Physics, 28(1):49–57, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Horan:1987:ONN

[1326] Paul Horan and Werner Blau. Optical nonlinearity near the bandgap in semiconductors. Contemporary Physics, 28(1):59–68, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vick:1987:LS

[1327] Sir Arthur Vick. A life in science. Contemporary Physics, 28(1):69– 70, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riley:1987:WIF

[1328] Professor N. Riley. Wave interactions and fluid flows. Contemporary Physics, 28(1):71–73, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRa

[1329] Anonymous. Book reviews. Contemporary Physics, 28(1):75–104, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:SNa

[1330] Anonymous. Short notices. Contemporary Physics, 28(1):105–108, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRb

[1331] Anonymous. Books received. Contemporary Physics, 28(1):109–113, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lipson:1987:HC

[1332] S. G. Lipson. Helium crystals. Contemporary Physics, 28(2):117–142, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coles:1987:HFI

[1333] B. R. Coles. Heavy-fermion intermetallic compounds. Contemporary Physics, 28(2):143–157, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1987:ISL

[1334] A. H. Cook. The inverse square law of gravitation. Contemporary Physics, 28(2):159–175, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:1987:ERS

[1335] David Bailin. Essay reviews supersymmetry and supergravity. Contemporary Physics, 28(2):177–181, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1987:NPT

[1336] Professor H. W. Wilson. Nuclear power technology. Contemporary Physics, 28(2):183–185, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRc

[1337] Anonymous. Book reviews. Contemporary Physics, 28(2):187–211, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRd

[1338] Anonymous. Books received. Contemporary Physics, 28(2):213–216, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gratias:1987:Q

[1339] Denis Gratias. Quasicrystals. Contemporary Physics, 28(3):219–239, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beling:1987:LEP

[1340] C. D. Beling and M. Charlton. Low-energy positron beams-origins, developments and applications. Contemporary Physics, 28(3):241–266, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Markov:1987:MEG

[1341] I. Markov and S. Stoyanov. Mechanisms of epitaxial growth. Contemporary Physics, 28(3):267–320, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbour:1987:MPT

[1342] Dr. Julian B. Barbour. Mach on the principles of the theory of heat. Contemporary Physics, 28(3):321–322, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Learner:1987:HDC

[1343] Dr. R. Learner. Horace Darwin and the Cambridge Scientific Instrument Company. Contemporary Physics, 28(3):323–324, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRe

[1344] Anonymous. Book reviews. Contemporary Physics, 28(3):325–336, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:SNb

[1345] Anonymous. Short notices. Contemporary Physics, 28(3):337–341, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRf

[1346] Anonymous. Books received. Contemporary Physics, 28(3):343–345, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cazabat:1987:HDD

[1347] Anne-Marie Cazabat. How does a droplet spread? Contemporary Physics, 28(4):347–364, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1987:SN

[1348] P. E. Hodgson. Superdeformed nuclei. Contemporary Physics, 28(4): 365–382, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ferguson:1987:LSA

[1349] A. I. Ferguson and J. M. Tolchard. Laser spectroscopy of atomic hydrogen. Contemporary Physics, 28(4):383–405, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joanny:1987:PD

[1350] Dr. Jean François Joanny. Polymer dynamics. Contemporary Physics, 28(4):407–410, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barrow:1987:AP

[1351] Dr. John D. Barrow. An astrophysical primer. Contemporary Physics, 28(4):411–412, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRg

[1352] Anonymous. Book reviews. Contemporary Physics, 28(4):413–434, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:SNc

[1353] Anonymous. Short notices. Contemporary Physics, 28(4):433–434, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRh

[1354] Anonymous. Books received. Contemporary Physics, 28(4):435–438, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Murdin:1987:SLM

[1355] Paul Murdin. The supernova in the Large Magellanic Cloud. Contemporary Physics, 28(5):441–455, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blair:1987:RBG

[1356] D. G. Blair. Resonant-bar gravitational radiation antennas. Contemporary Physics, 28(5):457–475, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jullien:1987:APF

[1357] R. Jullien. Aggregation phenomena and fractal aggregates. Contemporary Physics, 28(5):477–493, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:1987:LSM

[1358] Dr. David Bailin. The large, the small and the miniscule. Contemporary Physics, 28(5):495–499, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jaynes:1987:CRP

[1359] Professor E. T. Jaynes. Comment on a review by P. W. Atkins. Contemporary Physics, 28(5):501–502, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRi

[1360] Anonymous. Book reviews. Contemporary Physics, 28(5):503–515, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:SNd

[1361] Anonymous. Short notices. Contemporary Physics, 28(5):515–516, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRj

[1362] Anonymous. Books received. Contemporary Physics, 28(5):517–519, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Spain:1987:SHP

[1363] Ian L. Spain. Semiconductors at high pressure: New physics with the diamond-anvil cell. Contemporary Physics, 28(6):523–546, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dugdale:1987:ETM

[1364] J. S. Dugdale. Electron transport in metallic glasses. Contemporary Physics, 28(6):547–572, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowler:1987:PPC

[1365] K. C. Bowler and R. D. Kenway. Physics on parallel computers part 1: The new technology. Contemporary Physics, 28(6):573–598, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:1987:PT

[1366] Dr. I. D. Lawrie. Phase transitions. Contemporary Physics, 28(6):599– 601, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRk

[1367] Anonymous. Book reviews. Contemporary Physics, 28(6):603–612, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:SNe

[1368] Anonymous. Short notices. Contemporary Physics, 28(6):613, 1987. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1987:BRl

[1369] Anonymous. Books received. Contemporary Physics, 28(6):615–616, 1987. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Batchelder:1988:CCC

[1370] D. N. Batchelder. Colour and chromism of conjugated polymers. Contemporary Physics, 29(1):3–31, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowler:1988:PPC

[1371] K. C. Bowler and R. D. Kenway. Physics on parallel computers part 2: Applications. Contemporary Physics, 29(1):33–55, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Agostini:1988:PEU

[1372] P. Agostini and G. Petite. Photoelectric effect under strong irradiation. Contemporary Physics, 29(1):57–77, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1988:LDT

[1373] Professor R. J. Taylor. Light from the depths of time. Contemporary Physics, 29(1):79–82, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Olive:1988:US

[1374] Professor D. Olive. Unravelling superstrings. Contemporary Physics, 29 (1):83–85, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRa

[1375] Anonymous. Book reviews. Contemporary Physics, 29(1):87–96, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:SNa

[1376] Anonymous. Short notices. Contemporary Physics, 29(1):96–98, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRb

[1377] Anonymous. Books received. Contemporary Physics, 29(1):99–102, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:1988:LFP

[1378] Stig Stenholm. Light forces put a handle on the atom: To cool and trap atoms by laser light. Contemporary Physics, 29(2):105–123, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashfold:1988:MPM

[1379] M. N. R. Ashfold and J. D. Prince. Multiphoton processes in molecular gases. Contemporary Physics, 29(2):125–157, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1988:TSN

[1380] P. F. Smith. Terrestrial searches for new stable particles. Contemporary Physics, 29(2):159–186, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1988:NP

[1381] Dr. P. E. Hodgson. Nuclear physics. Contemporary Physics, 29(2):187– 191, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRc

[1382] Anonymous. Book reviews. Contemporary Physics, 29(2):193–223, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:SNb

[1383] Anonymous. Short notices. Contemporary Physics, 29(2):224–228, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRd

[1384] Anonymous. Books received. Contemporary Physics, 29(2):229–231, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henderson:1988:OPT

[1385] B. Henderson and G. F. Imbusch. Optical processes in tunable transitionmetal-ion lasers. Contemporary Physics, 29(3):235–272, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Djabourov:1988:AGG

[1386] Madeleine Djabourov. Architecture of gelatin gels. Contemporary Physics, 29(3):273–297, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leuchs:1988:SQF

[1387] Gerd Leuchs. Squeezing the quantum fluctuations of light. Contemporary Physics, 29(3):299–314, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aitchison:1988:EB

[1388] Dr. I. J. R. Aitchison. Either ‘and’ or ‘or’ but not both. Contemporary Physics, 29(3):315–320, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRe

[1389] Anonymous. Book reviews. Contemporary Physics, 29(3):321–334, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRf

[1390] Anonymous. Books received. Contemporary Physics, 29(3):335–337, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harrison:1988:DIC

[1391] Robert G. Harrison. Dynamical instabilities and chaos in lasers. Contemporary Physics, 29(4):341–371, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trimble:1988:DMU

[1392] Virginia Trimble. Dark matter in the universe: Where, what, and why? Contemporary Physics, 29(4):373–392, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pippard:1988:IIS

[1393] A. B. Pippard. The invincible ignorance of science. Contemporary Physics, 29(4):393–405, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1988:QC

[1394] Professor C. H. Llewellyn Smith. Quantum chromodynamics. Contemporary Physics, 29(4):407–409, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yofe:1988:PS

[1395] Dr. A. D. Yofe. Physics at surfaces. Contemporary Physics, 29(4):411– 414, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRg

[1396] Anonymous. Book reviews. Contemporary Physics, 29(4):415–426, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:SNc

[1397] Anonymous. Short notices. Contemporary Physics, 29(4):426–427, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRh

[1398] Anonymous. Books received. Contemporary Physics, 29(4):429–431, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Land:1988:OAE

[1399] Michael F. Land. The optics of animal eyes. Contemporary Physics, 29 (5):435–455, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1988:MPN

[1400] P. E. Hodgson. Multistep processes in nuclear reactions. Contemporary Physics, 29(5):457–476, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hartland:1988:QSE

[1401] A. Hartland. Quantum standards for electrical units. Contemporary Physics, 29(5):477–498, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:1988:PIC

[1402] Dr. R. F. Bishop. Path integrals and condensed matter physics. Contemporary Physics, 29(5):499–502, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRi

[1403] Anonymous. Book reviews. Contemporary Physics, 29(5):503–521, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRj

[1404] Anonymous. Books received. Contemporary Physics, 29(5):523–524, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chandrasekhar:1988:RDP

[1405] S. Chandrasekhar. Recent developments in the physics of liquid crystals. Contemporary Physics, 29(6):527–558, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Adams:1988:IRA

[1406] Nigel G. Adams and David Smith. Ionic reactions in atmospheric, interstellar and laboratory plasmas. Contemporary Physics, 29(6):559–578, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leger:1988:EP

[1407] L. Léger and J. L. Viovy. Entangled polymers. Contemporary Physics, 29(6):579–595, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pickering:1988:AC

[1408] Dr. John Pickering. An anthropic cosmogeny? Contemporary Physics, 29(6):597–599, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barton:1988:SMN

[1409] Dr. G. Barton. Statistical mechanics need not be dull. Contemporary Physics, 29(6):601–602, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRk

[1410] Anonymous. Book reviews. Contemporary Physics, 29(6):603–610, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1988:BRl

[1411] Anonymous. Books received. Contemporary Physics, 29(6):611–612, 1988. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berry:1989:HGN

[1412] R. Stephen Berry. How good is Niels Bohr’s atomic model? Contemporary Physics, 30(1):1–19, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hamilton:1989:ISL

[1413] M. W. Hamilton. An introduction to stabilized lasers. Contemporary Physics, 30(1):21–33, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1989:PNS

[1414] W. G. Williams and S. W. Lovesey. Pulsed neutron sources and condensed matter research. Contemporary Physics, 30(1):35–53, 1989. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lattimer:1989:GCS

[1415] James M. Lattimer. Gravitational collapse and supernovae. Contemporary Physics, 30(1):55–64, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:1989:GD

[1416] Gerry Gilmore. Galactic dynamics. Contemporary Physics, 30(1):65– 67, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRa

[1417] Anonymous. Book reviews. Contemporary Physics, 30(1):69–74, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRb

[1418] Anonymous. Books received. Contemporary Physics, 30(1):75–76, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Patchett:1989:XRA

[1419] B. E. Patchett. X-ray astronomy. Contemporary Physics, 30(2):77–88, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dash:1989:SM

[1420] J. G. Dash. Surface melting. Contemporary Physics, 30(2):89–100, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pegg:1989:SNT

[1421] D. T. Pegg. Some new techniques in nuclear magnetic resonance. Contemporary Physics, 30(2):101–112, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Croquette:1989:CPDa

[1422] Vincent Croquette. Convective pattern dynamics at low Prandtl number: Part I. Contemporary Physics, 30(2):113–133, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1989:BHT

[1423] P. C. W. Davies. A brief history of time. Contemporary Physics, 30(2): 135–136, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:1989:SFT

[1424] R. F. Bishop. Statistical field theory. Contemporary Physics, 30(2): 137–140, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRc

[1425] Anonymous. Book reviews. Contemporary Physics, 30(2):141–150, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRd

[1426] Anonymous. Books received. Contemporary Physics, 30(2):151–152, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Croquette:1989:CPDb

[1427] Vincent Croquette. Convective pattern dynamics at low Prandtl number: Part II. Contemporary Physics, 30(3):153–171, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martinson:1989:SSS

[1428] I. Martinson and L. J. Curtis. Spectroscopic studies of the structure of highly ionised atoms. Contemporary Physics, 30(3):173–185, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Naumovets:1989:PTT

[1429] A. G. Naumovets. Phase transitions in two dimensions. Contemporary Physics, 30(3):187–201, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Adam:1989:PTG

[1430] M. Adam and M. Delsanti. Percolation-type growth of branched polymers near gelation threshold. Contemporary Physics, 30(3):203–218, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sutton:1989:MTM

[1431] A. P. Sutton. Modern theory of metals. Contemporary Physics, 30(3): 219–222, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRe

Anonymous:1989:BRf

[1433] Anonymous. Books received. Contemporary Physics, 30(3):235–236, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:1989:WS

[1434] D. Bailin. Why superstrings? Contemporary Physics, 30(4):237–250, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chernov:1989:FCS

[1435] A. A. Chernov. Formation of crystals in solutions. Contemporary Physics, 30(4):251–276, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Newton:1989:SN

[1436] J. O. Newton. Spinning nuclei. Contemporary Physics, 30(4):277–299, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barton:1989:FSM

[1437] G. Barton. Foundations of statistical mechanics. Contemporary Physics, 30(4):301–302, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parries:1989:OF

[1438] M. C. Parries. Optical fibre. Contemporary Physics, 30(4):303–304, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRg

[1439] Anonymous. Book reviews. Contemporary Physics, 30(4):305–317, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRh

[1440] Anonymous. Books received. Contemporary Physics, 30(4):318–320, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brock:1989:HOL

[1441] Joel D. Brock, Robert J. Birgeneau, David Litster, and Amnon Aharony. Hexatic ordering in liquid crystal films. Contemporary Physics, 30(5):

321–335, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Townsend:1989:SMD

[1442] P. D. Townsend and Y. Kirsh. Spectral measurement during thermoluminescence — an essential requirement. Contemporary Physics, 30(5): 337–354, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hutchinson:1989:TL

[1443] M. H. R. Hutchinson. Terawatt lasers. Contemporary Physics, 30(5): 355–365, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gifford:1989:DLI

[1444] D. Gifford. Dose limits for ionising radiation. Contemporary Physics, 30 (5):367–376, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:1989:CPR

[1445] David Bailin. Of conservatives, pragmatists, revolutionaries and cranks. Contemporary Physics, 30(5):377–380, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRi

[1446] Anonymous. Book reviews. Contemporary Physics, 30(5):381–397, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:SN

[1447] Anonymous. Short notices. Contemporary Physics, 30(5):397–399, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRj

[1448] Anonymous. Books received. Contemporary Physics, 30(5):400–402, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salaneck:1989:CDC

[1449] W. R. Salaneck. Conformational defects in a conducting polymer. Contemporary Physics, 30(6):403–416, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1989:GE

[1450] B. J. Mason. The greenhouse effect. Contemporary Physics, 30(6):417– 432, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Backenstoss:1989:AA

[1451] G. Backenstoss. Antiprotonic atoms. Contemporary Physics, 30(6):433– 448, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1989:LAE

[1452] N. J. Mason. Laser-assisted electron-atom collisions. Contemporary Physics, 30(6):449–463, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burnett:1989:NP

[1453] K. Burnett. The new physics. Contemporary Physics, 30(6):465–466, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbour:1989:UT

[1454] J. B. Barbour. Unrolling time. Contemporary Physics, 30(6):467–468, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRk

[1455] Anonymous. Book reviews. Contemporary Physics, 30(6):469–482, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1989:BRl

[1456] Anonymous. Books received. Contemporary Physics, 30(6):483–484, 1989. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charvolin:1990:CFF

[1457] Jean Charvolin. Crystals of fluid films. Contemporary Physics, 31(1): 1–17, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bransden:1990:CEI

[1458] B. H. Bransden. Charge exchange in ion-atom collisions. Contemporary Physics, 31(1):19–33, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bagguley:1990:FRC

[1459] D. M. S. Bagguley and B. Bleaney. Ferromagnetic resonance at the Clarendon Laboratory, Oxford. A tribute to J. H. E. Griffiths (1908– 1981). Contemporary Physics, 31(1):35–42, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCrea:1990:OWM

[1460] Sir William McCrea. Origin of wave mechanics. Contemporary Physics, 31(1):43–48, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Series:1990:JLL

[1461] G. W. Series. Journey into light. Life and science of C. V. Raman. Contemporary Physics, 31(1):49–51, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRa

[1462] Anonymous. Book reviews. Contemporary Physics, 31(1):53–70, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRb

[1463] Anonymous. Books received. Contemporary Physics, 31(1):71–72, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vvedensky:1990:QMM

[1464] D. D. Vvedensky, S. Crampin, M. E. Eberhart, and J. M. Maclaren. Quantum mechanics and mechanical properties: Towards twenty-first century materials. Contemporary Physics, 31(2):73–97, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1990:ACN

[1465] P. E. Hodgson. Alpha clustering in nuclei. Contemporary Physics, 31(2): 99–111, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kaiser:1990:BFUa

[1466] Nick Kaiser. Bulk flows in the universe: Part 1 — current cosmogonical theories. Contemporary Physics, 31(2):113–125, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rivers:1990:FTF

[1467] R. J. Rivers. Finite-temperature field theory. Contemporary Physics, 31 (2):127–128, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Power:1990:DW

[1468] E. A. Power. Diffusion and waves. Contemporary Physics, 31(2):129– 130, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:1990:QMP

[1469] R. F. Bishop. Quantum many-particle systems. Contemporary Physics, 31(2):131–134, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRc

[1470] Anonymous. Book reviews. Contemporary Physics, 31(2):135–145, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRd

[1471] Anonymous. Books received. Contemporary Physics, 31(2):147–148, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kaiser:1990:BFUb

[1472] Nick Kaiser. Bulk flows in the universe: Part 2 — observations and implications. Contemporary Physics, 31(3):149–163, 1990. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milborrow:1990:EW

[1473] D. J. Milborrow. Energy from the wind. Contemporary Physics, 31(3): 165–178, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mannella:1990:NND

[1474] R. Mannella and P. V. E. McClintock. Noise in nonlinear dynamical systems. Contemporary Physics, 31(3):179–194, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joyce:1990:MBE

[1475] B. A. Joyce. Molecular beam epitaxy-fundamentals and current status. Contemporary Physics, 31(3):195–197, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kilmister:1990:ARM

[1476] C. W. Kilmister. Absolute or relative motion? Contemporary Physics, 31(3):199–202, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRe

[1477] Anonymous. Book reviews. Contemporary Physics, 31(3):203–215, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRf

[1478] Anonymous. Books received. Contemporary Physics, 31(3):217–218, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ponomarev:1990:MCF

[1479] L. I. Ponomarev. Muon catalysed fusion. Contemporary Physics, 31(4): 219–245, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

She:1990:RMA

[1480] C. Y. She. Remote measurement of atmospheric parameters: New applications of physics with lasers. Contemporary Physics, 31(4):247–260, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weise:1990:PCP

[1481] W. Weise. The pion as a composite particle. Contemporary Physics, 31 (4):261–270, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1990:HST

[1482] Sir Francis Graham Smith. The Hubble Space Telescope. Contemporary Physics, 31(4):271–272, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wallace:1990:SFT

[1483] D. J. Wallace. Statistical field theory. Volumes 1 and 2. Contemporary Physics, 31(4):273–274, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRg

[1484] Anonymous. Book reviews. Contemporary Physics, 31(4):275–290, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRh

[1485] Anonymous. Books received. Contemporary Physics, 31(4):291–293, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1990:UNO

[1486] P. E. Hodgson. The unification of the nuclear optical potential. Contemporary Physics, 31(5):295–308, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bjornholm:1990:CCM

[1487] S. Bjørnholm. Clusters, condensed matter in embryonic form. Contemporary Physics, 31(5):309–324, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kenway:1990:PSH

[1488] R. D. Kenway. Particles from scratch: The hadron spectrum from lattice gauge theory. Contemporary Physics, 31(5):325–333, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parker:1990:DFT

[1489] T. J. Parker. Dispersive Fourier transform spectroscopy. Contemporary Physics, 31(5):335–353, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:1990:UU

[1490] I. D. Lawrie. Unification and universality. Contemporary Physics, 31(5): 355–359, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRi

[1491] Anonymous. Book reviews. Contemporary Physics, 31(5):361–368, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:SNa

[1492] Anonymous. Short notices. Contemporary Physics, 31(5):368–369, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRj

[1493] Anonymous. Books received. Contemporary Physics, 31(5):371–372, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mott:1990:HTS

[1494] N. F. Mott. High temperature superconductivity; the spin polaron theory. Contemporary Physics, 31(6):373–385, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myers:1990:DCC

[1495] S. Myers and E. Picasso. The design, construction and commissioning of the CERN large electron–positron collider. Contemporary Physics, 31(6): 387–403, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:1990:LIS

[1496] P. T. Greenland. Laser isotope separation. Contemporary Physics, 31(6): 405–424, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vick:1990:GRN

[1497] Arthur Vick. G. R. Noakes (1906–1990). Contemporary Physics, 31 (6):425, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRk

Anonymous:1990:SNb

[1499] Anonymous. Short notices. Contemporary Physics, 31(6):436–437, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1990:BRl

[1500] Anonymous. Books received. Contemporary Physics, 31(6):439–440, 1990. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chua:1991:NRH

[1501] Christopher Chua and Kosta Tsipis. Nuclear reactor high-level waste: Origin and safe disposal. Contemporary Physics, 32(1):1–9, 1991. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shen:1991:STM

[1502] J. Shen, R. G. Pritchard, and R. E. Thurstans. Scanning tunnelling microscopy. Contemporary Physics, 32(1):11–20, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joyce:1991:MBE

[1503] B. A. Joyce. Molecular beam epitaxy of semiconductor films — atomic dimension control and the evaluation of crystal growth dynamics. Contemporary Physics, 32(1):21–34, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Copeland:1991:FE

[1504] Ed Copeland. The fifth essence. Contemporary Physics, 32(1):35–38, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tayler:1991:U

[1505] R. J. Tayler. Universe. Contemporary Physics, 32(1):39–41, 1991. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barton:1991:CYG

[1506] G. Barton. Can you go there from here? Contemporary Physics, 32 (1):42–43, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hochheimer:1991:IS

[1507] H. D. Hochheimer. Ian L. Spain (1940–1990). Contemporary Physics, 32 (1):45–46, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRa

[1508] Anonymous. Book reviews. Contemporary Physics, 32(1):47–70, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRb

[1509] Anonymous. Books received. Contemporary Physics, 32(1):71–72, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Celnikier:1991:PMGa

[1510] L. M. Celnikier and N. Meyer. A plain man’s guide to the cometary environment. Contemporary Physics, 32(2):73–88, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smilauer:1991:TMF

[1511] Pavel Šmilauer. Thin metal films and percolation theory. Contemporary Physics, 32(2):89–102, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trimble:1991:NSB

[1512] Virginia Trimble. Neutron stars and black holes in binary systems. Contemporary Physics, 32(2):103–119, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRc

[1513] Anonymous. Book reviews. Contemporary Physics, 32(2):121–139, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRd

[1514] Anonymous. Books received. Contemporary Physics, 32(2):140–142, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mathon:1991:EIG

[1515] J. Mathon. Exchange interactions and giant magnetoresistance in magnetic multilayers. Contemporary Physics, 32(3):143–156, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:1991:AON

[1516] R. N. Wilson. Active optics and the new technology telescope (NTT): The key to improved optical quality at lower cost in large astronomical telescopes. Contemporary Physics, 32(3):157–172, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sambles:1991:OES

[1517] J. R. Sambles, G. W. Bradbery, and Fuzi Yang. Optical excitation of surface plasmons: An introduction. Contemporary Physics, 32(3):173– 183, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Alber:1991:LIEa

[1518] G. Alber and P. Zoller. Laser-induced excitation of electronic Rydberg wave packets. Contemporary Physics, 32(3):185–189, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRe

[1519] Anonymous. Book reviews. Contemporary Physics, 32(3):191–215, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRf

[1520] Anonymous. Books received. Contemporary Physics, 32(3):216–218, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aitchison:1991:A

[1521] I. J. R. Aitchison and N. E. Mavromatos. Anyons. Contemporary Physics, 32(4):219–233, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowley:1991:PEE

[1522] S. W. H. Cowley. The plasma environment of the Earth. Contemporary Physics, 32(4):235–250, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Elliott:1991:DNS

[1523] S. R. Elliott and R. G. H. Robertson. Detecting neutrinos from the sun. Contemporary Physics, 32(4):251–263, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rae:1991:QR

[1524] Dr. A. I. M. Rae. Quantics rudiments. Contemporary Physics, 32(4): 265–266, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRg

[1525] Anonymous. Book reviews. Contemporary Physics, 32(4):267–286, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRh

[1526] Anonymous. Books received. Contemporary Physics, 32(4):287–289, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wambach:1991:NS

[1527] J. Wambach. Nuclear sound. Contemporary Physics, 32(5):291–303, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Malcolm:1991:DPS

[1528] G. P. A. Malcolm and A. I. Ferguson. Diode-pumped solid-state lasers. Contemporary Physics, 32(5):305–319, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milchev:1991:EPF

[1529] Alexander Milchev. Electrochemical phase formation on a foreign substrate — basic theoretical concepts and some experimental results. Contemporary Physics, 32(5):321–332, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mestel:1991:NSM

[1530] L. Mestel. Neutron star magnetospheres. Contemporary Physics, 32(5): 333–334, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:1991:EA

[1531] S. M. Barnett. Exciting atoms. Contemporary Physics, 32(5):335–337, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tilley:1991:SH

[1532] John Tilley. Superfluid helium 3. Contemporary Physics, 32(5):339– 340, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRi

[1533] Anonymous. Book reviews. Contemporary Physics, 32(5):341–362, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRj

[1534] Anonymous. Books received. Contemporary Physics, 32(5):363–365, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Celnikier:1991:PMGb

[1535] L. M. Celnikier and N. Meyer. A plain man’s guide to the cometary environment. Contemporary Physics, 32(5):367, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Foot:1991:LCT

[1536] C. J. Foot. Laser cooling and trapping of atoms. Contemporary Physics, 32(6):369–381, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bauer:1991:MS

[1537] Günther Bauer and Heinz Krenn. Multilayered superstructures. Contemporary Physics, 32(6):383–402, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keyes:1991:LCE

[1538] Robert W. Keyes. Limits and challenges in electronics. Contemporary Physics, 32(6):403–419, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRk

[1539] Anonymous. Book reviews. Contemporary Physics, 32(6):421–442, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1991:BRl

[1540] Anonymous. Books received. Contemporary Physics, 32(6):443–444, 1991. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Alber:1991:LIEb

[1541] G. Alber and P. Zoller. Laser-induced excitation of electronic Rydberg wave packets. Contemporary Physics, 32(6):445, 1991. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hall:1992:MCP

[1542] G. Hall. Modern charged particle detectors. Contemporary Physics, 33 (1):1–14, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ginzburg:1992:OAA

[1543] V. L. Ginzburg. Once again about high-temperature superconductivity. Contemporary Physics, 33(1):15–23, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Viovy:1992:PDE

[1544] Jean Louis Viovy, Thomas Duke, and François Caron. The physics of DNA electrophoresis. Contemporary Physics, 33(1):25–40, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lebon:1992:QAA

[1545] G. Lebon, D. Jou, and J. Casas-Vazquez. Questions and answers about a thermodynamic theory of the third type. Contemporary Physics, 33 (1):41–51, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Richards:1992:C

[1546] D. Richards. Chaos. Contemporary Physics, 33(1):53–54, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Banks:1992:GS

[1547] R. Banks. Geological structures. Contemporary Physics, 33(1):55–56, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRa

[1548] Anonymous. Book reviews. Contemporary Physics, 33(1):57–78, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRb

[1549] Anonymous. Books received. Contemporary Physics, 33(1):79–80, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hewett:1992:GLN

[1550] P. C. Hewett. Gravitational lensing as a new cosmological tool. Contemporary Physics, 33(2):81–90, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dogiel:1992:GRA

[1551] V. A. Dogiel. Gamma-ray astronomy. Contemporary Physics, 33(2): 91–109, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Challis:1992:PLT

[1552] L. J. Challis. Physics in less than three dimensions. Contemporary Physics, 33(2):111–127, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Richards:1992:M

[1553] D. Richards. Mechanics. Contemporary Physics, 33(2):129–130, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1992:TNS

[1554] R. C. Thompson. Time for new standards? Contemporary Physics, 33 (2):131–134, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Copeland:1992:MC

[1555] E. Copeland. Modern cosmology. Contemporary Physics, 33(2):135– 136, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRc

[1556] Anonymous. Book reviews. Contemporary Physics, 33(2):137–159, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRd

[1557] Anonymous. Books received. Contemporary Physics, 33(2):160–161, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:E

[1558] Anonymous. Editorial. Contemporary Physics, 33(3):163, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Graham-Smith:1992:BMP

[1559] Francis Graham-Smith. The binary and millisecond pulsars. Contemporary Physics, 33(3):165–173, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lucas:1992:TRH

[1560] J. S. Lucas. Technology review: High hopes carried in orbit. Contemporary Physics, 33(3):175–176, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1992:SS

[1561] F. W. Taylor. Of seaweed and sunsets. Contemporary Physics, 33(3): 177–179, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Squires:1992:ERC

[1562] E. Squires. Essay review: Cosmogenesis — The Growth of Order in the Universe, by D. Lazer. ISBN 0-19-506908-0. Contemporary Physics, 33 (3):181–184, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Steer:1992:SM

[1563] W. S. Steer. Scanning microscopy. Contemporary Physics, 33(3):185– 186, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:1992:TT

[1564] G. W. Gibbons. Twistor theory. Contemporary Physics, 33(3):187–188, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lucas:1992:PA

[1565] J. S. Lucas. Physics for all. Contemporary Physics, 33(3):189–190, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Paton:1992:UFT

[1566] J. Paton. Unifying field theory. Contemporary Physics, 33(3):191–195, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRe

[1567] Anonymous. Book reviews. Contemporary Physics, 33(3):197–218, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRf

[1568] Anonymous. Books received. Contemporary Physics, 33(3):219–220, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:1992:BSD

[1569] R. Peierls. Broken symmetries: (Dirac memorial lecture, Cambridge, 15 June 1992). Contemporary Physics, 33(4):221–226, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1992:SCN

[1570] B. J. Mason. Snow crystals, natural and man made. Contemporary Physics, 33(4):227–243, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evesque:1992:SDP

[1571] P. Evesque. Shaking dry powders and grains. Contemporary Physics, 33 (4):245–261, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1992:RSS

[1572] F. W. Taylor. Remote sensing of the solar system. Contemporary Physics, 33(4):263–266, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1992:NP

[1573] P. E. Hodgson. Nuclear physics. Contemporary Physics, 33(4):267–270, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garrett:1992:E

[1574] A. J. M. Garrett. Entropy. Contemporary Physics, 33(4):271–274, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Farid:1992:GB

[1575] B. Farid. From groups to bands. Contemporary Physics, 33(4):275– 277, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRg

[1576] Anonymous. Book reviews. Contemporary Physics, 33(4):279–301, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRh

[1577] Anonymous. Books received. Contemporary Physics, 33(4):303–304, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Abragam:1992:NFA

[1578] Anatole Abragam. Nuclear ferromagnetism and ant ferromagnetism. Contemporary Physics, 33(5):305–312, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:1992:CF

[1579] Peter W. Milonni and Mei-Li Shih. Casimir forces. Contemporary Physics, 33(5):313–322, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caplin:1992:HC

[1580] D. Caplin. High T_{c} chemistry. Contemporary Physics, 33(5):323, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Richards:1992:MM

[1581] D. Richards. Motion and mechanics. Contemporary Physics, 33(5):325, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Guy:1992:BMF

[1582] C. Guy. Biology and magnetic fields. Contemporary Physics, 33(5): 327–328, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jordan:1992:SP

[1583] C. Jordan. Solar physics. Contemporary Physics, 33(5):329–330, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Russell:1992:W

[1584] P. St. J. Russell. Waveguides. Contemporary Physics, 33(5):331–333, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRi

[1585] Anonymous. Book reviews. Contemporary Physics, 33(5):335–353, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRj

[1586] Anonymous. Books received. Contemporary Physics, 33(5):355–356, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1992:IAU

[1587] J. K. Davies. Infared astronomy from the UK infrared telescope. Contemporary Physics, 33(6):357–368, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whall:1992:PMG

[1588] T. E. Whall. A plain man’s guide to delta-doped semiconductors. Contemporary Physics, 33(6):369–381, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowley:1992:SPP

[1589] S. W. H. Cowley. Space plasma physics. Contemporary Physics, 33(6): 383–384, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Firth:1992:EM

[1590] I. M. Firth. Exploring music. Contemporary Physics, 33(6):385–387, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lucas:1992:MP

[1591] J. S. Lucas. Massively parallel. Contemporary Physics, 33(6):389–392, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRk

[1592] Anonymous. Book reviews. Contemporary Physics, 33(6):393–408, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1992:BRl

[1593] Anonymous. Books received. Contemporary Physics, 33(6):409–410, 1992. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ironside:1993:UFA

[1594] C. N. Ironside. Ultra-fast all-optical switching. Contemporary Physics, 34(1):1–18, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1993:ROG

[1595] B. J. Mason. Role of the oceans in global climate. Contemporary Physics, 34(1):19–30, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henshaw:1993:REH

[1596] D. L. Henshaw. Radon exposure in the home: Its occurrence and possible health effects. Contemporary Physics, 34(1):31–48, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

New:1993:CP

[1597] G. H. C. New. Is creation purposeless. Contemporary Physics, 34(1): 49–51, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRa

[1598] Anonymous. Book reviews. Contemporary Physics, 34(1):53–65, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRb

[1599] Anonymous. Books received. Contemporary Physics, 34(1):67–68, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kane:1993:PGS

[1600] G. L. Kane. Physics goals of the superconducting supercollider (SSC). Contemporary Physics, 34(2):69–78, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ross:1993:TE

[1601] G. G. Ross. A ‘theory of everything’? Contemporary Physics, 34(2): 79–88, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matthew:1993:SQA

[1602] J. A. D. Matthew. Some quasi-atomic effects in the solid state. Contemporary Physics, 34(2):89–98, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kibble:1993:FT

[1603] Tom Kibble. The final theory. Contemporary Physics, 34(2):99–101, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRc

[1604] Anonymous. Book reviews. Contemporary Physics, 34(2):103–115, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRd

[1605] Anonymous. Books received. Contemporary Physics, 34(2):117–118, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rempe:1993:AOC

[1606] Gerhard Rempe. Atoms in an optical cavity: Quantum electrodynamics in confined space. Contemporary Physics, 34(3):119–129, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beskin:1993:SRP

[1607] V. S. Beskin. Spin-down of radio pulsars. Contemporary Physics, 34(3): 131–145, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Satterthwaite:1993:AD

[1608] G. E. Satterthwaite. Astronomical discovery. Contemporary Physics, 34 (3):147–150, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caplin:1993:S

[1609] A. D. Caplin. Superconductivity. Contemporary Physics, 34(3):151– 152, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1993:MM

[1610] R. C. Thompson. Modern metrology. Contemporary Physics, 34(3): 153–155, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRe

Anonymous:1993:SN

[1612] Anonymous. Short notices. Contemporary Physics, 34(3):173–174, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRf

[1613] Anonymous. Book received. Contemporary Physics, 34(3):175–176, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hughes:1993:EP

[1614] R. J. Hughes. The equivalence principle. Contemporary Physics, 34(4): 177–191, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

White:1993:STE

[1615] G. K. White. Solids: Thermal expansion and contraction. Contemporary Physics, 34(4):193–204, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1993:HS

[1616] F. W. Taylor. A history of science. Contemporary Physics, 34(4):205– 207, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Liddle:1993:MC

[1617] A. R. Liddle. Modern cosmology. Contemporary Physics, 34(4):209– 211, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:1993:CB

[1618] P. V. E. McClintock. Chaos for beginners. Contemporary Physics, 34(4): 213–214, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRg

[1619] Anonymous. Book reviews. Contemporary Physics, 34(4):215–232, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRh

[1620] Anonymous. Books received. Contemporary Physics, 34(4):233–234, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mandel:1993:LIU

[1621] Paul Mandel. Lasing without inversion: A useful concept? Contemporary Physics, 34(5):235–246, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ben-Jacob:1993:SFG

[1622] Eshel Ben-Jacob. From snowflake formation to growth of bacterial colonies Part I. Diffusive patterning in azoic systems. Contemporary Physics, 34(5):247–273, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Elitzur:1993:OL

[1623] A. C. Elitzur. The origin of life. Contemporary Physics, 34(5):275– 278, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hassard:1993:NF

[1624] J. Hassard. Nuclear fixes. Contemporary Physics, 34(5):279–280, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRi

[1625] Anonymous. Book review. Contemporary Physics, 34(5):281–294, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRj

[1626] Anonymous. Books received. Contemporary Physics, 34(5):295–296, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smilansky:1993:HSD

[1627] U. Smilansky. On hearing the shape of a drum, and related problems. Contemporary Physics, 34(6):297–302, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Copeland:1993:IAC

[1628] Edmund J. Copeland. Inflation after the cosmic background explorer (COBE). Contemporary Physics, 34(6):303–316, 1993. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1993:AIE

[1629] S. P. Thompson. The 2175 å interstellar extinction feature. Contemporary Physics, 34(6):317–327, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Squires:1993:CQW

[1630] E. J. Squires. Consciousness in the quantum world. Contemporary Physics, 34(6):329–331, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:1993:GRC

[1631] L. H. Ryder. General relativity comes of age. Contemporary Physics, 34 (6):333–336, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRk

[1632] Anonymous. Book reviews. Contemporary Physics, 34(6):337–358, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1993:BRl

[1633] Anonymous. Books received. Contemporary Physics, 34(6):359–360, 1993. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Adams:1994:AO

[1634] C. S. Adams. Atom optics. Contemporary Physics, 35(1):1–19, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shi:1994:MSM

[1635] Shu Shi. Molecular structures, macroscopic nonlinear optical effects, and photonic devices. Contemporary Physics, 35(1):21–36, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kenway:1994:VRS

[1636] R. D. Kenway. Virtual-reality science. Contemporary Physics, 35(1): 37–39, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Copeland:1994:DMU

[1637] E. Copeland. Dark matter in the universe. Contemporary Physics, 35 (1):41–43, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRa

[1638] Anonymous. Book reviews. Contemporary Physics, 35(1):45–72, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRb

[1639] Anonymous. Books received. Contemporary Physics, 35(1):73–74, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hughes:1994:CA

[1640] David W. Hughes. Comets and asteroids. Contemporary Physics, 35 (2):75–93, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Broglia:1994:CMC

[1641] R. A. Broglia. The colour of metal clusters and of atomic nuclei. Contemporary Physics, 35(2):95–104, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tennyson:1994:HFP

[1642] Jonathan Tennyson and Steven Miller. H_{3}^{+}: from first principles to Jupiter. Contemporary Physics, 35(2):105–116, 1994. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keesing:1994:NC

[1643] R. G. Keesing. Newton in context. Contemporary Physics, 35(2):117– 118, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wallis:1994:FCS

[1644] Max K. Wallis. Fallible cometary scientists. Contemporary Physics, 35 (2):119–122, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRc

[1645] Anonymous. Book reviews. Contemporary Physics, 35(2):123–147, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRd

[1646] Anonymous. Books received. Contemporary Physics, 35(2):149–150, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:1994:MAA

[1647] Lewis H. Ryder. Matter-antimatter asymmetry in the early universe and violation of time reversal symmetry. Contemporary Physics, 35(3): 151–163, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bryant:1994:EAA

[1648] Duncan A. Bryant. Electron acceleration in the aurora. Contemporary Physics, 35(3):165–179, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mingos:1994:CVC

[1649] D. Michael P. Mingos. A chemist’s view of clusters. Contemporary Physics, 35(3):181–189, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yerozolimsky:1994:FND

[1650] B. G. Yerozolimsky. Free neutron decay. Contemporary Physics, 35(3): 191–208, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:1994:SCF

[1651] P. T. Greenland. The story of cold fusion. Contemporary Physics, 35(3): 209–211, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:1994:SC

[1652] R. C. Smith. The significance of COBE. Contemporary Physics, 35(3): 213–215, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Power:1994:ERI

[1653] E. A. Power. Essay review: The insights of John Wheeler. Contemporary Physics, 35(3):217–219, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRe

Anonymous:1994:BRf

[1655] Anonymous. Books received. Contemporary Physics, 35(3):229–230, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Drury:1994:ACR

[1656] L. O’C. Drury. Acceleration of cosmic rays. Contemporary Physics, 35(4): 231–242, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Codling:1994:CES

[1657] K. Codling and L. J. Frasinski. Coulomb explosion of simple molecules in intense laser fields. Contemporary Physics, 35(4):243–255, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thomas:1994:MPM

[1658] J. E. Thomas. Measuring the position of moving atoms with uncertaintyprinciple-limited precision. Contemporary Physics, 35(4):257–267, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:1994:CCM

[1659] David A. Williams. Cosmic chemistry: Molecules in space. Contemporary Physics, 35(4):269–283, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:1994:AP

[1660] D. J. Miller. Accelerator physics. Contemporary Physics, 35(4):285– 287, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:1994:CA

[1661] M. J. Thompson. Conceptual astronomy. Contemporary Physics, 35(4): 289–290, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRg

Anonymous:1994:BRh

[1663] Anonymous. Books received. Contemporary Physics, 35(4):309–310, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Monteiro:1994:CAP

[1664] T. S. Monteiro. Chaos in atomic physics. Contemporary Physics, 35(5): 311–327, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1994:NS

[1665] P. E. Hodgson. Nuclear structure. Contemporary Physics, 35(5):329– 346, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howson:1994:MTF

[1666] Mark A. Howson. Magnetism of thin films and multilayers. Contemporary Physics, 35(5):347–359, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wherrett:1994:PG

[1667] B. S. Wherrett. The physics of golf. Contemporary Physics, 35(5):361– 362, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRi

[1668] Anonymous. Book reviews. Contemporary Physics, 35(5):363–374, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRj

[1669] Anonymous. Books received. Contemporary Physics, 35(5):375–376, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCarthy:1994:EMS

[1670] I. E. McCarthy and E. Weigold. Electron momentum spectroscopy of solid surfaces. Contemporary Physics, 35(6):377–384, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowlt:1994:ERM

[1671] C. Bowlt. Environmental radioactivity: Measurements and their meaning. Contemporary Physics, 35(6):385–398, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:1994:OMM

[1672] D. Bailin. One man’s myth. Contemporary Physics, 35(6):399–405, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

March:1994:LS

[1673] N. H. March. The liquid state. Contemporary Physics, 35(6):407–408, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:1994:BLF

[1674] Stephen M. Barnett. A backward look at the future. Contemporary Physics, 35(6):409–411, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRk

[1675] Anonymous. Book reviews. Contemporary Physics, 35(6):413–431, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1994:BRl

[1676] Anonymous. Books received. Contemporary Physics, 35(6):433–434, 1994. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Macgillivray:1995:RRL

[1677] W. R. Macgillivray. The role of resonant laser radiation in completely determining electron-atom collision parameters. Contemporary Physics, 36(1):1–14, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:1995:ECI

[1678] B. W. Shore. Examples of counter-intuitive physics. Contemporary Physics, 36(1):15–28, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Albrow:1995:TQ

[1679] Michael G. Albrow. The top quark. Contemporary Physics, 36(1):29– 36, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tayler:1995:HU

[1680] Roger J. Tayler. Helium in the universe. Contemporary Physics, 36(1): 37–48, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Maartens:1995:SBH

[1681] R. Maartens. The story of black holes. Contemporary Physics, 36(1): 49–51, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRa

[1682] Anonymous. Book reviews. Contemporary Physics, 36(1):53–76, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRb

[1683] Anonymous. Books received. Contemporary Physics, 36(1):77–78, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schumacher:1995:CMV

[1684] R. T. Schumacher and J. Woodhouse. Computer modelling of violin playing. Contemporary Physics, 36(2):79–92, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holstein:1995:ABE

[1685] Barry R. Holstein. The Aharonov–Bohm effect and variations. Contemporary Physics, 36(2):93–102, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hammond:1995:NFG

[1686] Richard T. Hammond. New fields in general relativity. Contemporary Physics, 36(2):103–114, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Griffin:1995:SL

[1687] M. J. Griffin. Science for the layman. Contemporary Physics, 36(2): 115–116, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keesing:1995:NA

[1688] R. G. Keesing. Newton’s alchemy. Contemporary Physics, 36(2):117– 119, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRc

[1689] Anonymous. Book reviews. Contemporary Physics, 36(2):121–136, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRd

[1690] Anonymous. Books received. Contemporary Physics, 36(2):137–138, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wodkiewicz:1995:NLL

[1691] Krzysztof Wódkiewicz. Non-local and local ghost fields in quantum correlations. Contemporary Physics, 36(3):139–147, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hughes:1995:QC

[1692] Richard J. Hughes, D. M. Alde, P. Dyer, G. G. Luther, G. L. Morgan, and M. Schauer. Quantum cryptography. Contemporary Physics, 36(3): 149–163, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Phoenix:1995:QCH

[1693] Simon J. D. Phoenix and Paul D. Townsend. Quantum cryptography: How to beat the code breakers using quantum mechanics. Contemporary Physics, 36(3):165–195, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:1995:UTF

[1694] P. E. Hodgson. The ultimate technological fix? Contemporary Physics, 36(3):197–200, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRe

Anonymous:1995:BRf

[1696] Anonymous. Books received. Contemporary Physics, 36(3):213–214, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harris:1995:LFF

[1697] Michael Harris. Light-field fluctuations in space and time. Contemporary Physics, 36(4):215–233, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Letokhov:1995:LMD

[1698] V. S. Letokhov. Laser Maxwell’s demon. Contemporary Physics, 36(4): 235–243, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anderson:1995:TOG

[1699] David L. T. Anderson. The tropical ocean global atmosphere programme. Contemporary Physics, 36(4):245–265, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jacobs:1995:EMF

[1700] J. A. Jacobs. The Earth’s magnetic field. Contemporary Physics, 36(4): 267–277, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRg

[1701] Anonymous. Book reviews. Contemporary Physics, 36(4):279–295, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRh

[1702] Anonymous. Books received. Contemporary Physics, 36(4):297–298, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1995:PCC

[1703] B. J. Mason. Predictions of climate changes caused by man-made emissions of greenhouse gases: A critical assessment. Contemporary Physics, 36(5):299–319, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Waxman:1995:ICA

[1704] D. Waxman. Impact of computer algebra on physics. Contemporary Physics, 36(5):321–327, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kerner:1995:HCU

[1705] Edward H. Kerner. How certain is the uncertainty principle? Contemporary Physics, 36(5):329–334, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:1995:SP

[1706] A. D. Martin. The structure of the proton. Contemporary Physics, 36(5): 335–353, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:1995:KQG

[1707] G. W. Gibbons. Knots and quantum gravity. Contemporary Physics, 36 (5):355–356, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:1995:BBB

[1708] R. S. Ellis. The big bang for beginners. Contemporary Physics, 36 (5):357, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRi

[1709] Anonymous. Book reviews. Contemporary Physics, 36(5):359–372, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRj

[1710] Anonymous. Books received. Contemporary Physics, 36(5):373–374, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knight:1995:RP

[1711] Professor P. L. Knight. Rudolf Peierls (1907–1995). Contemporary Physics, 36(6):375, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Johnson:1995:ECM

[1712] Neil F. Johnson. Electron correlations in mesoscopic structures. Contemporary Physics, 36(6):377–387, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Willstrop:1995:NDT

[1713] R. V. Willstrop. New developments in telescope design. Contemporary Physics, 36(6):389–410, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:1995:HST

[1714] R. S. Ellis. Hubble Space Telescope observations of distant galaxies. Contemporary Physics, 36(6):411–421, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Annett:1995:US

[1715] James F. Annett. Unconventional superconductivity. Contemporary Physics, 36(6):423–437, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:1995:RHE

[1716] J. J. Quenby. The review of high-energy astrophysics. Contemporary Physics, 36(6):439–442, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Isenberg:1995:PB

[1717] Cyril Isenberg. The physics of bubbles. Contemporary Physics, 36(6): 443–444, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRk

[1718] Anonymous. Book reviews. Contemporary Physics, 36(6):445–456, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1995:BRl

[1719] Anonymous. Books received. Contemporary Physics, 36(6):457–458, 1995. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burnett:1996:BEC

[1720] Keith Burnett. Bose–Einstein condensation with evaporatively cooled atoms. Contemporary Physics, 37(1):1–14, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Guy:1996:SRM

[1721] C. N. Guy. The second revolution in medical imaging. Contemporary Physics, 37(1):15–45, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yerozolimsky:1996:EPN

[1722] B. G. Yerozolimsky. Electromagnetic properties of the neutron. Contemporary Physics, 37(1):47–69, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:1996:RNG

[1723] F. E. Close. Releasing the nuclear genie. Contemporary Physics, 37(1): 71–73, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRa

[1724] Anonymous. Book reviews. Contemporary Physics, 37(1):75–91, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRb

[1725] Anonymous. Books received. Contemporary Physics, 37(1):93–94, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fink:1996:TRA

[1726] M. Fink. Time reversal in acoustics. Contemporary Physics, 37(2):95– 109, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:1996:OQW

[1727] A. M. Fox. Optoelectronics in quantum well structures. Contemporary Physics, 37(2):111–125, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Watson:1996:HCT

[1728] Greg Watson. Hall conductance as a topological invariant. Contemporary Physics, 37(2):127–143, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coley:1996:GA

[1729] D. A. Coley. Genetic algorithms. Contemporary Physics, 37(2):145– 154, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:1996:CA

[1730] A. H. Harker. Computer algebra. Contemporary Physics, 37(2):155– 157, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Copeland:1996:OCD

[1731] E. J. Copeland. Observing cosmic distances. Contemporary Physics, 37 (2):159–161, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRc

[1732] Anonymous. Book reviews. Contemporary Physics, 37(2):163–171, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRd

[1733] Anonymous. Books received. Contemporary Physics, 37(2):173–174, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wolfenstein:1996:NMP

[1734] L. Wolfenstein. Neutrino mass in particle physics and astrophysics. Contemporary Physics, 37(3):175–182, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bracco:1996:NR

[1735] A. Bracco, F. Alasia, S. Leoni, E. Vigezz, and R. A. Broglia. Nuclear rotations. Contemporary Physics, 37(3):183–204, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wark:1996:TRX

[1736] Justin Wark. Time-resolved X-ray diffraction. Contemporary Physics, 37(3):205–218, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zlmmermann:1996:FST

[1737] W. Zlmmermann. The flow of superfluid^{4} he through submicron apertures: Phase slip and critical velocities due to quantum vortex motion.

Contemporary Physics, 37(3):219–234, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fowler:1996:FSS

[1738] Patrick W. Fowler. Fullerene stability and structure. Contemporary Physics, 37(3):235–247, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRe

[1739] Anonymous. Book reviews. Contemporary Physics, 37(3):249–260, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRf

[1740] Anonymous. Books received. Contemporary Physics, 37(3):261–262, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hendry:1996:AU

[1741] Martin A. Hendry and Roger J. Tayler. The age of the universe. Contemporary Physics, 37(4):263–281, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

French:1996:USS

[1742] P. M. W. French. Ultrafast solid-state lasers. Contemporary Physics, 37 (4):283–301, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schenzle:1996:IRM

[1743] Axel Schenzle. Illusion or reality: The measurement process in quantum optics. Contemporary Physics, 37(4):303–320, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mousseau:1996:GL

[1744] Normand Mousseau. The ‘game of life’. Contemporary Physics, 37(4): 321–323, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Celnikier:1996:SHC

[1745] L. M. Celnikier. Space and the human condition. Contemporary Physics, 37(4):325–328, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRg

[1746] Anonymous. Book reviews. Contemporary Physics, 37(4):329–341, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRh

[1747] Anonymous. Books received. Contemporary Physics, 37(4):343–344, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Prior:1996:DIV

[1748] K. Prior. The development of II–VI semiconductors for blue diode lasers. Contemporary Physics, 37(5):345–358, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barstow:1996:HWD

[1749] M. A. Barstow. Hot white dwarfs: An observational viewpoint. Contemporary Physics, 37(5):359–374, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barenco:1996:QPC

[1750] Adriano Barenco. Quantum physics and computers. Contemporary Physics, 37(5):375–389, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:1996:RSE

[1751] F. W. Taylor. Remote sensing of the earth from space. Contemporary Physics, 37(5):391–405, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ledingham:1996:DDU

[1752] K. W. D. Ledingham. The development of detectors for use at low light levels. Contemporary Physics, 37(5):407–408, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McMurry:1996:BM

[1753] S. M. McMurry. The beauty of mathematics. Contemporary Physics, 37 (5):409–411, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bat:1996:PNS

[1754] D. J. Bat. Physics for non-scientists. Contemporary Physics, 37(5): 413–415, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRi

[1755] Anonymous. Book reviews. Contemporary Physics, 37(5):417–424, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRj

[1756] Anonymous. Book received. Contemporary Physics, 37(5):425–426, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coles:1996:PSN

[1757] Professor B. R. Coles. Professor Sir Nevill Mott (1905–1996). Contemporary Physics, 37(6):427, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coles:1996:LSS

[1758] Peter Coles. The large-scale structure of the universe. Contemporary Physics, 37(6):429–440, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Midgley:1996:QED

[1759] P. A. Midgley and M. Saunders. Quantitative electron diffraction: From atoms to bonds. Contemporary Physics, 37(6):441–456, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hellings:1996:GWD

[1760] Ronald W. Hellings. Gravitational wave detectors in space. Contemporary Physics, 37(6):457–469, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Phoenix:1996:SR

[1761] S. J. D. Phoenix. Science and religion. Contemporary Physics, 37(6): 471–473, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dholakia:1996:LSC

[1762] K. Dholakia. Laser spectroscopy comes of age. Contemporary Physics, 37(6):475–477, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRk

[1763] Anonymous. Book reviews. Contemporary Physics, 37(6):479–499, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1996:BRl

[1764] Anonymous. Book reviews. Contemporary Physics, 37(6):501–502, 1996. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:1997:GHN

[1765] F. E. Close. Glueballs and hybrids: New states of matter. Contemporary Physics, 38(1):1–12, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morrison:1997:CGB

[1766] L. V. Morrison and F. R. Stephenson. Contemporary geophysics from Babylonian clay tablets. Contemporary Physics, 38(1):13–23, 1997. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Horvath:1997:FPT

[1767] G. Zs. K. Horvath, R. C. Thompson, and P. L. Knight. Fundamental physics with trapped ions. Contemporary Physics, 38(1):25–48, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moss:1997:DGC

[1768] David Moss. Dynamo generation of cosmic magnetic fields. Contemporary Physics, 38(1):49–62, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ford:1997:HTS

[1769] P. J. Ford and G. A. Saunders. High-temperature superconductivity-ten years on. Contemporary Physics, 38(1):63–81, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:1997:ERP

[1770] P. R. Hobson. Essay review popular particle physics. Contemporary Physics, 38(1):83–85, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRa

Anonymous:1997:BRb

[1772] Anonymous. Books received. Contemporary Physics, 38(1):101, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Halliday:1997:RDT

[1773] Alex N. Halliday. Radioactivity, the discovery of time and the earliest history of the Earth. Contemporary Physics, 38(2):103–114, 1997. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gillan:1997:VML

[1774] M. J. Gillan. The virtual matter laboratory. Contemporary Physics, 38 (2):115–130, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Maher:1997:CSC

[1775] K. P. Maher and J. F. Malone. Computerized scatter correction in diagnostic radiology. Contemporary Physics, 38(2):131–148, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Makdisi:1997:PSD

[1776] Y. Y. Makdisi, K. S. Bhatia, and J. P. Connerade. Probing the spatial distribution of excited atomic states. Contemporary Physics, 38(2): 149–155, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1997:ERS

[1777] A. H. Cook. Essay review the state of gravity. Contemporary Physics, 38(2):157–160, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbour:1997:ERR

[1778] J. B. Barbour. Essay review reflections on the aims, methods, and criteria of the history of science. Contemporary Physics, 38(2):161–166, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRc

[1779] Anonymous. Book reviews. Contemporary Physics, 38(2):167–177, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRd

Greenland:1997:A

[1781] P. T. Greenland. Antimatter. Contemporary Physics, 38(3):181–203, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ben-Jacob:1997:SFG

[1782] Eshel Ben-Jacob. From snowflake formation to growth of bacterial colonies II: Cooperative formation of complex colonial patterns. Contemporary Physics, 38(3):205–241, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Drolias:1997:ERI

[1783] B. F. Drolias. Essay review interesting times for astronomy. Contemporary Physics, 38(3):243–245, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRe

[1784] Anonymous. Book reviews. Contemporary Physics, 38(3):247–253, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRf

[1785] Anonymous. Books received. Contemporary Physics, 38(3):255–256, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:1997:JQT

[1786] Stig Stenholm and Martin Wilkens. Jumps in quantum theory. Contemporary Physics, 38(4):257–268, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Glatzmaier:1997:SG

[1787] Gary A. Glatzmaier and Paul H. Roberts. Simulating the geodynamo. Contemporary Physics, 38(4):269–288, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:1997:SSO

[1788] Nigel J. Mason and S. K. Pathak. Spectroscopic studies of ozone- the Earth’s UV filter. Contemporary Physics, 38(4):289–299, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRg

[1789] Anonymous. Book reviews. Contemporary Physics, 38(4):301–311, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRh

[1790] Anonymous. Books received. Contemporary Physics, 38(4):313–314, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ditmire:1997:ACU

[1791] T. Ditmire. Atomic clusters in ultrahigh intensity light fields. Contemporary Physics, 38(5):315–328, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sholtz:1997:SPS

[1792] Paul Sholtz, Michael Bretz, and Franco Nori. Sound-producing sand avalanches. Contemporary Physics, 38(5):329–342, 1997. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Raymer:1997:MQM

[1793] M. G. Raymer. Measuring the quantum mechanical wave function. Contemporary Physics, 38(5):343–355, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRi

[1794] Anonymous. Book reviews. Contemporary Physics, 38(5):357–367, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRj

[1795] Anonymous. Books received. Contemporary Physics, 38(5):369, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bier:1997:BRP

[1796] Martin Bier. Brownian ratchets in physics and biology. Contemporary Physics, 38(6):371–379, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ross:1997:FEP

[1797] Marc Ross. Fuel efficiency and the physics of automobiles. Contemporary Physics, 38(6):381–394, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jameson:1997:BD

[1798] R. F. Jameson and S. T. Hodgkin. Brown dwarfs. Contemporary Physics, 38(6):395–407, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dunlop:1997:CFI

[1799] James Dunlop. Cosmic fireworks: Illuminating the history of our universe. Contemporary Physics, 38(6):409–427, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barton:1997:ER

[1800] G. Barton. Essay review. Contemporary Physics, 38(6):429–430, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gough:1997:ERS

[1801] D. O. Gough. Essay review the sun as seen by a stellar physicist. Contemporary Physics, 38(6):431–432, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRk

[1802] Anonymous. Book reviews. Contemporary Physics, 38(6):433–446, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1997:BRl

[1803] Anonymous. Books received. Contemporary Physics, 38(6):447, 1997. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riles:1998:TSM

[1804] Keith Riles. Testing the Standard Model. Contemporary Physics, 39 (1):1–11, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gill:1998:DDP

[1805] Alasdair J. Gill. Defects and the dynamics of phase transitions. Contemporary Physics, 39(1):13–47, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fearn:1998:SCT

[1806] H. Fearn and A. M. Arroyo. Seeing clearly through the atmosphere. Contemporary Physics, 39(1):49–66, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:1998:CQN

[1807] G. J. Milburn and He Bi Sun. Classical and quantum noise in electronic systems. Contemporary Physics, 39(1):67–79, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:1998:ERR

[1808] S. M. Barnett. Essay review reflections on the EPR paradox. Contemporary Physics, 39(1):81–82, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRa

[1809] Anonymous. Book reviews. Contemporary Physics, 39(1):83–92, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRb

[1810] Anonymous. Books received. Contemporary Physics, 39(1):93–94, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Liddle:1998:FUF

[1811] Andrew R. Liddle. Fingerprinting the universe- future cosmic microwave background experiments. Contemporary Physics, 39(2):95–105, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ricci:1998:SGW

[1812] Fulvio Ricci. The search for gravitational waves: An experimental physics challenge. Contemporary Physics, 39(2):107–135, 1998. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plenio:1998:ERQ

[1813] Martin Plenio. Essay review quantum optics. Contemporary Physics, 39 (2):137–139, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thurstans:1998:ERE

[1814] R. Thurstans. Essay review electron microscopy for beginners. Contemporary Physics, 39(2):141–143, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRc

Anonymous:1998:BRd

[1816] Anonymous. Books received. Contemporary Physics, 39(2):161–162, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilets:1998:FMD

[1817] Lawrence Wilets and James S. Cohen. Fermion molecular dynamics in atomic, molecular, and optical physics. Contemporary Physics, 39(3): 163–175, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shaposhnikov:1998:EB

[1818] M. E. Shaposhnikov. Electroweak baryogenesis. Contemporary Physics, 39(3):177–194, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pegg:1998:ERR

[1819] David T. Pegg. Essay review reflections on time. Contemporary Physics, 39(3):195–198, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bennion:1998:ERO

[1820] Ian Bennion. Essay review optical fibre sensors. Contemporary Physics, 39(3):199–202, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRe

[1821] Anonymous. Book reviews. Contemporary Physics, 39(3):203–222, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRf

[1822] Anonymous. Books received. Contemporary Physics, 39(3):223–224, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vachaspati:1998:TDC

[1823] Tanmay Vachaspati. Topological defects in the cosmos and lab. Contemporary Physics, 39(4):225–237, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gleiser:1998:PTU

[1824] Marcelo Gleiser. Phase transitions in the universe. Contemporary Physics, 39(4):239–253, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Todd:1998:PMC

[1825] T. N. Todd and C. G. Windsor. Progress in magnetic confinement fusion research. Contemporary Physics, 39(4):255–282, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tennyson:1998:WSS

[1826] Jonathan Tennyson and Oleg L. Polyansky. Water on the sun: The sun yields more secrets to spectroscopy. Contemporary Physics, 39(4): 283–294, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weaire:1998:ERP

[1827] D. Weaire. Essay review: Pride and precedence in Victorian science. Contemporary Physics, 39(4):295–297, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRg

[1828] Anonymous. Book reviews. Contemporary Physics, 39(4):299–314, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRh

[1829] Anonymous. Books received. Contemporary Physics, 39(4):315–316, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gauntlett:1998:MTS

[1830] Jerome P. Gauntlett. M-theory: Strings, duality and branes. Contemporary Physics, 39(5):317–328, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Meacher:1998:OLC

[1831] David R. Meacher. Optical lattices-crystalline structures bound by light. Contemporary Physics, 39(5):329–350, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dholakia:1998:AH

[1832] K. Dholakia. Atom hosepipes. Contemporary Physics, 39(5):351–369, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Broglia:1998:WSA

[1833] Ricardo A. Broglia. Wires of seven atoms — Feynman’s very, very small world. Contemporary Physics, 39(5):371–376, September/October 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keesing:1998:HNA

[1834] R. G. Keesing. The history of Newton’s apple tree. Contemporary Physics, 39(5):377–391, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:1998:ERT

[1835] Ken Zetie. Essay review time’s quantum arrow revisited. Contemporary Physics, 39(5):393–395, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Binney:1998:ERC

[1836] J. J. Binney. Essay review conflicting views of spiral galaxies. Contemporary Physics, 39(5):397–398, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRi

[1837] Anonymous. Book reviews. Contemporary Physics, 39(5):399–413, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRj

[1838] Anonymous. Books received. Contemporary Physics, 39(5):415–417, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hardy:1998:SAD

[1839] Lucien Hardy. Spooky action at a distance in quantum mechanics. Contemporary Physics, 39(6):419–429, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plenio:1998:TET

[1840] Martin B. Plenio and Vlatko Vedral. Teleportation, entanglement and thermodynamics in the quantum world. Contemporary Physics, 39(6):

431–446, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eisenberg:1998:ICB

[1841] Bob Eisenberg. Ionic channels in biological membranes- electrostatic analysis of a natural nanotube. Contemporary Physics, 39(6):447–466, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Povey:1998:UF

[1842] Malcolm J. W. Povey. Ultrasonics of food. Contemporary Physics, 39(6): 467–478, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:1998:ERM

[1843] Harry Jones. Essay review magnets at work and play. Contemporary Physics, 39(6):479–482, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wallis:1998:ERF

[1844] M. K. Wallis. Essay review: Field theorists strike back — stochastic electrodynamics. The Quantum Dice, by Luis de la Peña Ana-Maria Cetto. ISBN 0-7923-3818-9. Contemporary Physics, 39(6):483–486, 1998. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRk

[1845] Anonymous. Book reviews. Contemporary Physics, 39(6):487–508, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1998:BRl

[1846] Anonymous. Books received. Contemporary Physics, 39(6):509–510, 1998. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collins:1999:SPG

[1847] C. A. Collins. Searches for primeval galaxies. Contemporary Physics, 40 (1):1–10, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:1999:S

[1848] P. T. Greenland. Sonoluminescence. Contemporary Physics, 40(1):11– 30, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stefanovska:1999:PHC

[1849] Aneta Stefanovska. Physics of the human cardiovascular system. Contemporary Physics, 40(1):31–55, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fort:1999:ITB

[1850] J. Fort. Information theory and blackbody radiation. Contemporary Physics, 40(1):57–70, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mcclintock:1999:ERP

[1851] P. V. E. Mcclintock. Essay review puzzles of evolution. Contemporary Physics, 40(1):71–72, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:1999:ERU

[1852] A. Cook. Essay review understanding the flow of time. Contemporary Physics, 40(1):73–76, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRa

[1853] Anonymous. Book reviews. Contemporary Physics, 40(1):77–91, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRb

[1854] Anonymous. Books received. Contemporary Physics, 40(1):93–94, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schofield:1999:NFL

[1855] A. J. Schofield. Non-Fermi liquids. Contemporary Physics, 40(2):95– 115, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ward:1999:LL

[1856] A. J. Ward. Light on a lattice. Contemporary Physics, 40(2):117–137, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Freyberger:1999:ERV

[1857] M. Freyberger. Essay review visions. Contemporary Physics, 40(2):139– 141, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rivers:1999:ERA

[1858] R. Rivers. Essay review artful dodging. Contemporary Physics, 40(2): 143–145, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRc

[1859] Anonymous. Book reviews. Contemporary Physics, 40(2):147–168, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:SNa

[1860] Anonymous. Short notices. Contemporary Physics, 40(2):169–171, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRd

[1861] Anonymous. Books received. Contemporary Physics, 40(2):173, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:1999:SPM

[1862] S. J. Blundell. Spin-polarized muons in condensed matter physics. Contemporary Physics, 40(3):175–192, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:1999:FCE

[1863] Derryck T. Reid. Few cycle EM pulses. Contemporary Physics, 40(3): 193–204, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Winterton:1999:NLC

[1864] R. H. S. Winterton. Newton’s law of cooling. Contemporary Physics, 40 (3):205–212, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbour:1999:SH

[1865] J. Barbour. Science and history. Contemporary Physics, 40(3):213–218, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hassard:1999:NS

[1866] John Hassard. Nuclear safety? Contemporary Physics, 40(3):219–221, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRe

[1867] Anonymous. Book reviews. Contemporary Physics, 40(3):223–231, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:SNb

[1868] Anonymous. Short notices. Contemporary Physics, 40(3):233–234, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRf

[1869] Anonymous. Books received. Contemporary Physics, 40(3):235, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chandrasekhar:1999:RDD

[1870] S. Chandrasekhar and S. Krishna Prasad. Recent developments in discotic liquid crystals. Contemporary Physics, 40(4):237–245, 1999. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mesquita:1999:CBS

[1871] Marcus V. Mesquita, Aurea R. Vasconcellos, and Roberto Luzzi. Complexity in biological systems. Contemporary Physics, 40(4):247–256, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hey:1999:RFC

[1872] Tony Hey. Richard Feynman and computation. Contemporary Physics, 40(4):257–265, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pegg:1999:TR

[1873] D. T. Pegg. Time as a river. Contemporary Physics, 40(4):267–269, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:1999:OHS

[1874] John K. Davies. Observing with the Hubble Space Telescope. Contemporary Physics, 40(4):271–274, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bullock:1999:THP

[1875] F. W. Bullock. Tracing the history of particle physics. Contemporary Physics, 40(4):275–278, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRg

[1876] Anonymous. Book reviews. Contemporary Physics, 40(4):279–293, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:SNc

[1877] Anonymous. Short notices. Contemporary Physics, 40(4):295–296, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRh

[1878] Anonymous. Books received. Contemporary Physics, 40(4):297–298, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fulton:1999:CNN

[1879] B. R. Fulton. Clustering in nuclei: Nuclear chains, nuclear molecules and other exotic states of nuclear matter. Contemporary Physics, 40(5): 299–311, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Calogeracos:1999:HPK

[1880] A. Calogeracos and N. Dombey. History and physics of the Klein paradox. Contemporary Physics, 40(5):313–321, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fujita:1999:RDM

[1881] F. E. Fujita. Recent developments in Mössbauer spectroscopy. Contemporary Physics, 40(5):323–337, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Liddle:1999:FSL

[1882] Andrew R. Liddle. It’s full of stars …. Contemporary Physics, 40(5): 339–341, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Polkinghorne:1999:SCM

[1883] J. C. Polkinghorne. Science for the common man. Contemporary Physics, 40(5):343–345, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRi

Ledingham:1999:NPM

[1885] K. W. D. Ledingham and P. A. Norreys. Nuclear physics merely using a light source. Contemporary Physics, 40(6):367–383, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lieb:1999:TFA

[1886] Klaus-Peter Lieb. Thin film analysis with nuclear methods. Contemporary Physics, 40(6):385–413, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Campbell:1999:NDM

[1887] D. M. Campbell. Nonlinear dynamics of musical reed and brass wind instruments. Contemporary Physics, 40(6):415–431, 1999. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garraway:1999:QSC

[1888] B. M. Garraway. Quantum state conclusion? Contemporary Physics, 40 (6):433–435, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRj

[1889] Anonymous. Book reviews. Contemporary Physics, 40(6):437–459, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:SNd

[1890] Anonymous. Short notices. Contemporary Physics, 40(6):461–463, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:1999:BRk

[1891] Anonymous. Books received. Contemporary Physics, 40(6):465, 1999. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Uzer:2000:SCA

[1892] T. Uzer, Ernestine A. Lee, David Farrelly, and Andrea F. Brunello. Synthesis of a classical atom: Wavepacket analogues of the Trojan asteroids. Contemporary Physics, 41(1):1–14, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lunkenheimer:2000:GD

[1893] P. Lunkenheimer, U. Schneider, R. Brand, and A. Loid. Glassy dynamics. Contemporary Physics, 41(1):15–36, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sluckin:2000:LCP

[1894] T. J. Sluckin. The liquid crystal phases: Physics and technology. Contemporary Physics, 41(1):37–56, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRa

[1895] Anonymous. Books received. Contemporary Physics, 41(1):57–59, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kirk:2000:NSD

[1896] K. J. Kirk. Nanomagnets for sensors and data storage. Contemporary Physics, 41(2):61–78, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salje:2000:F

[1897] Ekhard K. H. Salje. Ferroelasticity. Contemporary Physics, 41(2):79– 91, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Herr:2000:MPI

[1898] Ulrich Herr. Metastable phases in interface controlled materials. Contemporary Physics, 41(2):93–104, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:2000:QP

[1899] P. E. Hodgson. Quantum philosophy. Contemporary Physics, 41(2): 105–108, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRb

[1900] Anonymous. Books received. Contemporary Physics, 41(2):109–110, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRc

[1901] Anonymous. Book reviews. Contemporary Physics, 41(2):111–127, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kalmus:2000:PPT

[1902] P. I. P. Kalmus. Particle physics at the turn of the century. Contemporary Physics, 41(3):129–142, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caves:2000:PFD

[1903] Carlton M. Caves. Predicting future duration from present age: A critical assessment. Contemporary Physics, 41(3):143–153, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Drolias:2000:ERB

[1904] B. Drolias. Essay reviews: Not with a bang but a whimper. Contemporary Physics, 41(3):155–157, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2000:NOM

[1905] P. J. Dobson. Nanotechnology: Opportunities missed. Contemporary Physics, 41(3):159–161, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charap:2000:RG

[1906] J. M. Charap. Relativistic gravitation. Contemporary Physics, 41(3): 163–165, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:2000:SRP

[1907] S. M. Barnett. Scientific revolutions, paradoxes and paradigms. Contemporary Physics, 41(3):167–169, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRd

[1908] Anonymous. Book reviews. Contemporary Physics, 41(3):171–185, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:SNa

[1909] Anonymous. Short notices. Contemporary Physics, 41(3):187–188, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRe

[1910] Anonymous. Books received. Contemporary Physics, 41(3):189–190, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Millar:2000:ESS

[1911] T. J. Millar. The evolution of solar systems. Contemporary Physics, 41 (4):191–202, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sander:2000:DLA

[1912] Leonard M. Sander. Diffusion-limited aggregation: A kinetic critical phenomenon? Contemporary Physics, 41(4):203–218, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Amusia:2000:ANL

[1913] M. Ya. Amusia and Y. Kornyushin. Application of the nuclear liquid drop model to atomic and molecular physics problems. Contemporary Physics, 41(4):219–229, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2000:YIM

[1914] D. Bailin. And yet it moves. Contemporary Physics, 41(4):231–233, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Goodlad:2000:CS

[1915] S. Goodlad. Communicating science. Contemporary Physics, 41(4):235– 237, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2000:SGS

[1916] F. W. Taylor. The second greatest story ever told. Contemporary Physics, 41(4):239–241, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thurstans:2000:ROS

[1917] R. Thurstans. Retelling the old story. Contemporary Physics, 41(4): 243–245, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Syroka:2000:MMP

[1918] J. Syroka and R. Toumi. Making money out of physics. Contemporary Physics, 41(4):247–250, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRf

Anonymous:2000:BRg

[1920] Anonymous. Books received. Contemporary Physics, 41(4):273–274, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Padgett:2000:LTT

[1921] Miles Padgett and L. Allen. Light with a twist in its tail. Contemporary Physics, 41(5):275–285, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnes:2000:SEE

[1922] W. L. Barnes. Spontaneous emission and energy transfer in the optical microcavity. Contemporary Physics, 41(5):287–300, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leonhardt:2000:LMM

[1923] U. Leonhardt and P. Piwnicki. Light in moving media. Contemporary Physics, 41(5):301–308, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harries:2000:PER

[1924] J. E. Harries. Physics of the Earth’s radiative energy balance. Contemporary Physics, 41(5):309–322, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2000:MU

[1925] F. W. Taylor. Measuring the universe. Contemporary Physics, 41(5): 323–325, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:2000:DSN

[1926] P. E. Hodgson. The deep symmetries of nature. Contemporary Physics, 41(5):327–330, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sillitto:2000:TE

[1927] R. M. Sillitto. In a tradition of excellence. Contemporary Physics, 41(5): 331–333, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:2000:AT

[1928] H. R. Brown. The arrow of time. Contemporary Physics, 41(5):335– 336, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRh

[1929] Anonymous. Book reviews. Contemporary Physics, 41(5):337–350, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:SNb

[1930] Anonymous. Short notices. Contemporary Physics, 41(5):351–353, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRi

[1931] Anonymous. Books received. Contemporary Physics, 41(5):355–357, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kane:2000:SWW

[1932] Gordon L. Kane. Supersymmetry: What? why? when? Contemporary Physics, 41(6):359–367, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Batelaan:2000:KDE

[1933] H. Batelaan. The Kapitza–Dirac effect. Contemporary Physics, 41(6): 369–381, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cummins:2000:NMR

[1934] H. K. Cummins and A. Jones. Nuclear magnetic resonance: A quantum technology for computation and spectroscopy. Contemporary Physics, 41 (6):383–399, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chefles:2000:QSD

[1935] Anthony Chefles. Quantum state discrimination. Contemporary Physics, 41(6):401–424, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2000:ERS

[1936] K. Zetie. Essay review: Still seeking the ultimate in physics. Contemporary Physics, 41(6):425–427, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rutter:2000:ERC

[1937] George H. Rutter. Essay review: Count Rumford — a forgotten genius. Contemporary Physics, 41(6):429–431, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2000:ERS

[1938] R. Gilmore. Essay review: Symmetry not so super. Contemporary Physics, 41(6):433–435, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRj

[1939] Anonymous. Book reviews. Contemporary Physics, 41(6):437–452, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:SNc

[1940] Anonymous. Short notices. Contemporary Physics, 41(6):455–457, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2000:BRk

[1941] Anonymous. Books received. Contemporary Physics, 41(6):459–460, 2000. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Timmermans:2001:SSM

[1942] Eddy Timmermans. Superfluids and superfluid mixtures in atom traps. Contemporary Physics, 42(1):1–24, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plenio:2001:PFL

[1943] M. B. Plenio and V. Vitelli. The physics of forgetting: Landauer’s erasure principle and information theory. Contemporary Physics, 42(1): 25–60, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kibble:2001:WM

[1944] B. P. Kibble. What is mass? Contemporary Physics, 42(1):61–63, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRa

[1945] Anonymous. Book reviews. Contemporary Physics, 42(1):65–74, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRb

[1946] Anonymous. Books received. Contemporary Physics, 42(1):75, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Godun:2001:PAI

[1947] R. M. Godun, M. B. D’Arcy, G. S. Summy, and K. Burnett. Prospects for atom interferometry. Contemporary Physics, 42(2):77–95, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sandars:2001:EDM

[1948] P. G. H. Sandars. Electric dipole moments of charged particles. Contemporary Physics, 42(2):97–111, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bugg:2001:TM

[1949] David Bugg. Thinking machines. Contemporary Physics, 42(2):113– 115, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:2001:ER

[1950] P. E. Hodgson. Einstein’s religion. Contemporary Physics, 42(2):117– 119, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davies:2001:TKA

[1951] E. R. Davies. Telepistemology — knowledge in the age of the Internet. Contemporary Physics, 42(2):121–124, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clarke:2001:SSH

[1952] N. M. Clarke. Science as social history. Contemporary Physics, 42(2): 125–127, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRc

Anonymous:2001:BRd

[1954] Anonymous. Books received. Contemporary Physics, 42(2):147–150, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanders:2001:SMR

[1955] T. M. Sanders, Jr. Semiclassical mechanics of rotons. Contemporary Physics, 42(3):151–157, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kornyushin:2001:FCE

[1956] Y. Kornyushin and M. Ya. Amusia. On the frequencies of collective electron oscillations in nanotubes. Contemporary Physics, 42(3):159– 165, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Satterthwaite:2001:LMG

[1957] G. E. Satterthwaite. A large man in a grand time. Contemporary Physics, 42(3):167–170, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2001:UDN

[1958] E. Sheldon. Understanding the decay of nuclei. Contemporary Physics, 42(3):171–173, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Poole:2001:PBO

[1959] M. W. Poole. Particle beam optics and the physics of accelerators. Contemporary Physics, 42(3):175–178, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Poon:2001:PGS

[1960] W. C. K. Poon and P. B. Warren. Physics going soft: The emergence of soft condensed matter physics. Contemporary Physics, 42(3):179–183, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRe

Anonymous:2001:SNa

[1962] Anonymous. Short notice. Contemporary Physics, 42(3):205, 2001. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRf

[1963] Anonymous. Books received. Contemporary Physics, 42(3):207, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hands:2001:PDQ

[1964] Simon Hands. The phase diagram of QCD. Contemporary Physics, 42(4): 209–225, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

James:2001:PF

[1965] Jessica James. Physics in finance. Contemporary Physics, 42(4):227– 238, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2001:QC

[1966] P. T. Greenland. Quantum computing. Contemporary Physics, 42(4): 239–241, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cullerne:2001:FWR

[1967] J. P. Cullerne. Free will and the resolution of time travel paradoxes. Contemporary Physics, 42(4):243–245, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Steel:2001:CSP

[1968] Duncan Steel. Carl Sagan: Practitioner, popularizer and proponent of science. Contemporary Physics, 42(4):247–249, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plenio:2001:PI

[1969] M. B. Plenio. The physics of information. Contemporary Physics, 42(4): 251–253, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRg

Anonymous:2001:BRh

[1971] Anonymous. Books received. Contemporary Physics, 42(4):273–274, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gelletly:2001:SRB

[1972] W. Gelletly. Science with radioactive beams: The alchemist’s dream. Contemporary Physics, 42(5):285–314, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thurstans:2001:LIP

[1973] R. Thurstans. Looking inside people: The physics of medical imaging. Contemporary Physics, 42(5):315–317, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Balogh:2001:FSE

[1974] André Balogh. The future of space exploration: Is it asking for the moon? Contemporary Physics, 42(5):319–321, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Galbraith:2001:BC

[1975] S. Galbraith. Beauty from complexity. Contemporary Physics, 42(5): 323–325, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRi

[1976] Anonymous. Book reviews. Contemporary Physics, 42(5):327–336, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRj

[1977] Anonymous. Books received. Contemporary Physics, 42(5):337, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Green:2001:CVB

[1978] M. G. Green and T. R. McMahon. CP violation in b meson decays. Contemporary Physics, 42(6):339–351, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keitel:2001:RQO

[1979] Christoph H. Keitel. Relativistic quantum optics. Contemporary Physics, 42(6):353–363, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cahn:2001:INP

[1980] Robert W. Cahn. Intermetallics: New physics. Contemporary Physics, 42(6):365–375, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coleman:2001:TYG

[1981] P. G. Coleman. Thirty years of gaseous positronics. Contemporary Physics, 42(6):377–379, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2001:WS

[1982] Geoffrey Duxbury. The workings of science. Contemporary Physics, 42 (6):381–382, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:2001:SR

[1983] P. E. Hodgson. Science and religion. Contemporary Physics, 42(6):383– 386, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garrett:2001:RC

[1984] Anthony Garrett. Reality cheque. Contemporary Physics, 42(6):387– 389, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:BRk

[1985] Anonymous. Book reviews. Contemporary Physics, 42(6):391–403, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2001:SNb

[1986] Anonymous. Short notices. Contemporary Physics, 42(6):405–407, 2001. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2002:RCR

[1987] B. J. Mason. The role of clouds in the radiative balance of the atmosphere and their effects on climate. Contemporary Physics, 43(1):1–11, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Waxman:2002:MPG

[1988] David Waxman. A model of population genetics and its mathematical relation to quantum theory. Contemporary Physics, 43(1):13–20, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

James:2002:PCP

[1989] Jessica James. Politically correct, but physically confusing. Contemporary Physics, 43(1):21–23, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRa

[1990] Anonymous. Book reviews. Contemporary Physics, 43(1):25–45, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRb

[1991] Anonymous. Books received. Contemporary Physics, 43(1):47–49, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khalil:2002:EDM

[1992] Shaaban Khalil and Carlos Muñoz. The enigma of the dark matter. Contemporary Physics, 43(2):51–61, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Singleton:2002:QTD

[1993] John Singleton and Charles Mielke. Quasi-two-dimensional organic superconductors: A review. Contemporary Physics, 43(2):63–96, 2002. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garraway:2002:WPD

[1994] Barry M. Garraway and Kalle-Antti Suominen. Wave packet dynamics in molecules. Contemporary Physics, 43(2):97–114, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2002:MMC

[1995] E. Sheldon. The mystique of mesons and cosmic radiation. Contemporary Physics, 43(2):115–120, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRc

[1996] Anonymous. Book review. Contemporary Physics, 43(2):121–142, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRd

[1997] Anonymous. Books received. Contemporary Physics, 43(2):143–145, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garraway:2002:DFE

[1998] B. M. Garraway and S. Stenholm. Does a flying electron spin? Contemporary Physics, 43(3):147–160, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sirohi:2002:SI

[1999] Rajpal S. Sirohi. Speckle interferometry. Contemporary Physics, 43(3): 161–180, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Watson:2002:CRH

[2000] A. A. Watson. Cosmic rays of the highest energies. Contemporary Physics, 43(3):181–195, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pinkerton:2002:TEM

[2001] Harry Pinkerton, Lionel Wilson, and Ray MacDonald. The transport and eruption of magma from volcanoes: A review. Contemporary Physics, 43(3):197–210, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2002:WCB

[2002] R. Gilmore. The wily colonial boy. Contemporary Physics, 43(3):211– 213, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hatherly:2002:LA

[2003] P. A. Hatherly. The legacy of Apollo. Contemporary Physics, 43(3): 215–217, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRe

[2004] Anonymous. Book reviews. Contemporary Physics, 43(3):219–240, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Molloy:2002:LAO

[2005] Justin E. Molloy and Miles J. Padgett. Lights, action: Optical tweezers. Contemporary Physics, 43(4):241–258, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ralph:2002:SII

[2006] J. F. Ralph and M. Bernhardt. Smart imaging in the infrared. Contemporary Physics, 43(4):259–272, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Henderson:2002:SED

[2007] Brian Henderson. Spectroscopic effects of disorder in laser materials. Contemporary Physics, 43(4):273–300, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2002:OUT

[2008] D. Bailin. Once upon a time, in the Soviet Union …. Contemporary Physics, 43(4):301–302, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pert:2002:FP

[2009] G. J. Pert. Fusion in prospect. Contemporary Physics, 43(4):303–306, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Altmann:2002:SW

[2010] S. L. Altmann. The science wars. Contemporary Physics, 43(4):307– 310, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRf

[2011] Anonymous. Book reviews. Contemporary Physics, 43(4):311–324, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blair:2002:FLC

[2012] Grahame A. Blair. The future linear collider. Contemporary Physics, 43 (5):325–338, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barger:2002:TAD

[2013] A. J. Barger. Tracing activity in distant supermassive black holes with X-rays. Contemporary Physics, 43(5):339–349, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Warburton:2002:SAS

[2014] Richard J. Warburton. Self-assembled semiconductor quantum dots. Contemporary Physics, 43(5):351–364, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zheludev:2002:NON

[2015] N. I. Zheludev. Nonlinear optics on the nanoscale. Contemporary Physics, 43(5):365–377, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Luchinsky:2002:DPN

[2016] D. G. Luchinsky. Deterministic patterns of noise and the control of chaos. Contemporary Physics, 43(5):379–395, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Byrne:2002:IRQ

[2017] J. Byrne. Illusion and reality in the quantum world. Contemporary Physics, 43(5):397–400, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanders:2002:HGB

[2018] T. M. Sanders. Heisenberg and the German bomb. Contemporary Physics, 43(5):401–403, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jacobs:2002:HWG

[2019] D. J. Jacobs. How we got here. Contemporary Physics, 43(5):405–407, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRg

[2020] Anonymous. Book reviews. Contemporary Physics, 43(5):409–425, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dash:2002:MOT

[2021] J. G. Dash. Melting from one to two to three dimensions. Contemporary Physics, 43(6):427–436, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Loveday:2002:SDS

[2022] Jon Loveday. The Sloan Digital Sky Survey. Contemporary Physics, 43 (6):437–449, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McKay:2002:WLG

[2023] Timothy A. McKay. When light goes astray: Gravitational lensing in astrophysics. Contemporary Physics, 43(6):451–460, 2002. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hodgson:2002:APC

[2024] P. E. Hodgson. Alpha-particle clustering in nuclear reactions. Contemporary Physics, 43(6):461–472, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Arruda-Neto:2002:POP

[2025] J. D. T. Arruda-Neto, S. S. Dutra, J. Arrais, and J. Mesa. Physics in organizations: A percolating approach. Contemporary Physics, 43(6): 473–479, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2002:TLA

[2026] D. G. C. Jones. Time to leave the alchemist’s cave? Contemporary Physics, 43(6):481–484, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2002:LNF

[2027] John Mason. From lightning to nuclear fusion. Contemporary Physics, 43(6):485–491, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRh

[2028] Anonymous. Book reviews. Contemporary Physics, 43(6):493–505, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2002:BRi

[2029] Anonymous. Books received. Contemporary Physics, 43(6):507–510, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ott:2003:PET

[2030] Bob Ott. Positron emission tomography: A view of the body’s function. Contemporary Physics, 44(1):1–15, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Townsend:2003:PPP

[2031] Peter D. Townsend. Photocathodes — past performance and future potential. Contemporary Physics, 44(1):17–34, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Arnold:2003:MFL

[2032] Kam S. Arnold and C. Y. She. Metal fluorescence lidar (light detection and ranging) and the middle atmosphere. Contemporary Physics, 44(1): 35–49, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kieu:2003:CNC

[2033] Tien D. Kieu. Computing the non-computable. Contemporary Physics, 44(1):51–71, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benson:2003:ERF

[2034] Andrew J. Benson. Essay review — a few of my favourite things. Contemporary Physics, 44(1):73–75, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2003:ERN

[2035] Derry W. Jones. Essay review — nuclear radiation, radioactivity and safety. Contemporary Physics, 44(1):77–79, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRa

[2036] Anonymous. Book reviews. Contemporary Physics, 44(1):81–95, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRb

Conroy:2003:FSM

[2038] R. S. Conroy. Frequency standards, metrology and fundamental constants. Contemporary Physics, 44(2):99–135, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leach:2003:FPV

[2039] Jonathan Leach, Miles Padgett, and Johannes Courtial. Fractals in pixellated video feedback. Contemporary Physics, 44(2):137–143, 2003. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Polturak:2003:IQS

[2040] Emil Polturak and Nir Gov. Inside a quantum solid. Contemporary Physics, 44(2):145–151, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:2003:ERS

[2041] J. I. B. Wilson. Essay review — scholarly froth and engineering skeletons. The Physics of Foams, by D. Weaire and S. Hutzler. ISBN 0-19-851097-7 (paperback), 0-19-850551-5 (hardback). Contemporary Physics, 44(2): 153–155, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRc

[2042] Anonymous. Book reviews. Contemporary Physics, 44(2):157–190, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRd

[2043] Anonymous. Books received. Contemporary Physics, 44(2):191–192, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khalil:2003:SSL

[2044] Shaaban Khalil. Search for supersymmetry at LHC. Contemporary Physics, 44(3):193–201, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Samm:2003:CTF

[2045] Ulrich Samm. Controlled thermonuclear fusion at the beginning of a new era. Contemporary Physics, 44(3):203–217, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morgan:2003:DDW

[2046] Ben Morgan. Direct detection of WIMP dark matter. Contemporary Physics, 44(3):219–235, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Farhadi:2003:RRC

[2047] Alessio Farhadi and Dimitri Vvedensky. Risk, randomness, crashes and quants. Contemporary Physics, 44(3):237–257, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vourdas:2003:IMJ

[2048] A. Vourdas. Interaction of mesoscopic Josephson devices with nonclassical microwaves. Contemporary Physics, 44(3):259–264, 2003. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2003:SL

[2049] Cm Linton. Signs of life. Contemporary Physics, 44(3):265–267, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2003:FRIa

[2050] Dw Jones. Fusion research and international scientific collaboration. Contemporary Physics, 44(3):269–272, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRe

[2051] Anonymous. Book reviews. Contemporary Physics, 44(3):273–284, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRf

[2052] Anonymous. Book received. Contemporary Physics, 44(3):285–287, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Squires:2003:DSD

[2053] Gl Squires. The discovery of the structure of DNA. Contemporary Physics, 44(4):289–305, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kempe:2003:QRW

[2054] J. Kempe. Quantum random walks: An introductory overview. Contemporary Physics, 44(4):307–327, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sergienko:2003:QIP

[2055] Av Sergienko and Gs Jaeger. Quantum information processing and precise optical measurement with entangled-photon pairs. Contemporary Physics, 44(4):341–356, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2003:NAW

[2056] Peter Bussey. Nature in all its wonders. Contemporary Physics, 44(4): 357–360, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cook:2003:ERT

[2057] Alan Cook. Essay review: The times of their lives, Pauli and company [No Time to Be Brief. A Scientific Biography of Wolfgang Pauli. By C. P. Enz. (Oxford University Press, 2002.). Pp. viii + 573. £35.00 (hardback). ISBN 0-19-856479-1. Scope: biography. Level: non-specialist]. Contemporary Physics, 44(4):361–364, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRg

[2058] Anonymous. Book reviews. Contemporary Physics, 44(4):365–387, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jm:2003:MGM

[2059] Valverde Jm, Castellanos A., and Quintanilla Mas. The memory of granular materials. Contemporary Physics, 44(5):389–399, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosenblum:2003:SPC

[2060] Michael Rosenblum and Arkady Pikovsky. Synchronization: From pendulum clocks to chaotic lasers and chemical oscillators. Contemporary Physics, 44(5):401–416, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chin:2003:SCS

[2061] J. Chin, J. Harting, S. Jha, Pv Coveney, Ar Porter, and Sm Pickles. Steering in computational science: Mesoscale modelling and simulation. Contemporary Physics, 44(5):417–434, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mij:2003:MDS

[2062] Probert Mij. Molecular dynamics studies of liquids using a Beowulf computer. Contemporary Physics, 44(5):435–450, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fletcher:2003:CBS

[2063] Lyndsay Fletcher. The costs and benefits of space weather. Contemporary Physics, 44(5):451–453, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2003:FRIb

[2064] Derry Jones. Fusion research and international scientific collaboration. Contemporary Physics, 44(5):455–458, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRh

[2065] Anonymous. Book reviews. Contemporary Physics, 44(5):459–467, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRi

[2066] Anonymous. Book received. Contemporary Physics, 44(5):469–471, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anderson:2003:NAO

[2067] Brian P. Anderson and Pierre Meystre. Nonlinear atom optics. Contemporary Physics, 44(6):473–483, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rajantie:2003:DFE

[2068] Arttu Rajantie. Defect formation in the early universe. Contemporary Physics, 44(6):485–502, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2003:QGO

[2069] Graham M. Shore. Quantum gravitational optics. Contemporary Physics, 44(6):503–521, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Engel:2003:GRB

[2070] Ar Engel. Have gamma-ray bursts kept their secrets? Contemporary Physics, 44(6):523–526, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hartley:2003:EM

[2071] J. Hartley. Entanglement for the masses. Contemporary Physics, 44(6): 527–528, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRj

[2072] Anonymous. Book reviews. Contemporary Physics, 44(6):529–545, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRk

[2073] Anonymous. Books received. Contemporary Physics, 44(6):547–548, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRAa

[2074] Anonymous. Books reviewed: author index. Contemporary Physics, 44 (6):549–552, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2003:BRT

[2075] Anonymous. Books reviewed: title index. Contemporary Physics, 44(6): 553–555, 2003. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Durkan:2004:DSE

[2076] C. Durkan. Detection of single electronic spins by scanning tunnelling microscopy. Contemporary Physics, 45(1):1–10, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harris:2004:PDS

[2077] Sarah Anne Harris. The physics of DNA stretching. Contemporary Physics, 45(1):11–30, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bekenstein:2004:BHI

[2078] Jacob D. Bekenstein. Black holes and information theory. Contemporary Physics, 45(1):31–43, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertram:2004:FPF

[2079] Cd Bertram. Flow phenomena in floppy tubes. Contemporary Physics, 45 (1):45–60, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dr:2004:WE

[2080] Pa Hatherly Dr. Where is everybody? Contemporary Physics, 45(1): 61–64, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbour:2004:NS

[2081] Julian Barbour. Is nature supernatural? Contemporary Physics, 45(1): 65–69, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cahn:2004:NIR

[2082] Robert W. Cahn. Nobel idiosyncracies: a review of Nobel Lectures: Physics (1996–2000). Edited by Gösta Ekspong. (World Scientific, 2002). Pp. ix + 485. ISBN 981-238-004-3. Contemporary Physics, 45(1):71– 73, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Maxwell:2004:HM

[2083] C. J. Maxwell. The heart of the matter. Contemporary Physics, 45(1): 75–77, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:BRa

[2084] Anonymous. Book reviews. Contemporary Physics, 45(1):79–92, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fantz:2004:OPO

[2085] U. Fantz. Optical phenomena in the open air. Contemporary Physics, 45 (2):93–108, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rose:2004:SCH

[2086] Sj Rose. Set the controls for the heart of the sun. Contemporary Physics, 45(2):109–121, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Flowers:2004:HSF

[2087] Jl Flowers, Ha Klein, and Hs Margolis. Hydrogenic systems, frequency standards and fundamental constants. Contemporary Physics, 45(2):123– 145, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Parrondo:2004:BMG

[2088] Jmr Parrondo and Luis Dinís. Brownian motion and gambling: from ratchets to paradoxical games. Contemporary Physics, 45(2):147–157, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2004:ARN

[2089] Richard S. Ellis. Astrobiology: a respectable new science. Contemporary Physics, 45(2):159–162, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garrett:2004:QMC

[2090] Anthony Jm Garrett. Quantum mechanics: cleaning up metaphysical baggage. Contemporary Physics, 45(2):163–165, 2004. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:SNa

[2091] Anonymous. Short notices. Contemporary Physics, 45(2):167–168, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:BRb

[2092] Anonymous. Book reviews. Contemporary Physics, 45(2):169–189, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pendry:2004:NR

[2093] Jb Pendry. Negative refraction. Contemporary Physics, 45(3):191–202, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bakule:2004:GAS

[2094] Pavel Bakule and Elvezio Morenzoni. Generation and applications of slow polarized muons. Contemporary Physics, 45(3):203–225, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brax:2004:CCP

[2095] Ph Brax. The cosmological constant problem. Contemporary Physics, 45(3):227–236, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kusmartsev:2004:EMS

[2096] Fv Kusmartsev. Electronic molecules in semiconductors and other novel materials. Contemporary Physics, 45(3):237–250, 2004. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2004:DUM

[2097] Dr. Matravers. Does the universe have a mask? Contemporary Physics, 45(3):251–254, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Freyberger:2004:NF

[2098] M. Freyberger. The new Fundamentals. Contemporary Physics, 45(3): 255–256, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Altmann:2004:CER

[2099] Simon L. Altmann. Comments on ‘essay review is nature supernatural?’. Contemporary Physics, 45(3):257–258, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:BRc

[2100] Anonymous. Book reviews. Contemporary Physics, 45(3):259–276, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Conroy:2004:UD

[2101] Rs Conroy and C. Danilowicz. Unravelling DNA. Contemporary Physics, 45(4):277–302, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Angelakis:2004:PCI

[2102] Dg Angelakis, Pl Knight, and E. Paspalakis. Photonic crystals and inhibition of spontaneous emission: an introduction. Contemporary Physics, 45(4):303–318, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCubbin:2004:BPL

[2103] Norman McCubbin. Beauty in physics: the legacy of Paul Dirac. Contemporary Physics, 45(4):319–333, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2004:PGA

[2104] E. Sheldon. Per gravitas ad astra. Contemporary Physics, 45(4):335– 339, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:SNb

[2105] Anonymous. Short notices. Contemporary Physics, 45(4):341–342, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:BRd

[2106] Anonymous. Book reviews. Contemporary Physics, 45(4):343–366, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jaksch:2004:OLU

[2107] D. Jaksch. Optical lattices, ultracold atoms and quantum information processing. Contemporary Physics, 45(5):367–381, 2004. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charvin:2004:TDS

[2108] G. Charvin, J.-F. Allemand, Tr Strick, D. Bensimon, and V. Croquette. Twisting DNA: single molecule studies. Contemporary Physics, 45(5): 383–403, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Helbing:2004:PTR

[2109] Dirk Helbing and Kai Nagel. The physics of traffic and regional development. Contemporary Physics, 45(5):405–426, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2004:PE

[2110] Peter J. Bussey. Pressing to the extremes. Contemporary Physics, 45(5): 427–430, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:SNc

[2111] Anonymous. Short notice. Contemporary Physics, 45(5):431, 2004. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:BRe

[2112] Anonymous. Book reviews. Contemporary Physics, 45(5):433–454, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evans:2004:CN

[2113] Ts Evans. Complex networks. Contemporary Physics, 45(6):455–474, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2004:FM

[2114] Randall Peters. Friction at the mesoscale. Contemporary Physics, 45(6): 475–490, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zuber:2004:DBD

[2115] K. Zuber. Double beta decay. Contemporary Physics, 45(6):491–502, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sagues:2004:PN

[2116] F. Sagués and Jm Sancho. Patterns out of noise. Contemporary Physics, 45(6):503–514, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilkins:2004:PCA

[2117] D. R. Wilkins. A problem-centred approach to the history of mathematics. Contemporary Physics, 45(6):515–517, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:SNd

[2118] Anonymous. Short notices. Contemporary Physics, 45(6):519–520, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2004:BRf

[2119] Anonymous. Book reviews. Contemporary Physics, 45(6):521–541, 2004. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Maricic:2005:KAN

[2120] Jelena Maricic and John Learned. The KamLAND anti-neutrino oscillation experiment. Contemporary Physics, 46(1):1–14, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McGloin:2005:BBD

[2121] D. McGloin and K. Dholakia. Bessel beams: Diffraction in a new light. Contemporary Physics, 46(1):15–28, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chrysostomou:2005:FS

[2122] Antonio Chrysostomou and Phil Lucas. The formation of stars. Contemporary Physics, 46(1):29–40, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bier:2005:MPM

[2123] Martin Bier. Modelling processive motor proteins: Moving on two legs in the microscopic realm. Contemporary Physics, 46(1):41–51, 2005. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Altmann:2005:CE

[2124] Sl Altmann. Complexity in everything. Contemporary Physics, 46(1): 53–56, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRa

[2125] Anonymous. Book reviews. Contemporary Physics, 46(1):57–75, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peiris:2005:FYW

[2126] Hiranya Peiris. First year Wilkinson Microwave Anisotropy Probe results: Implications for cosmology and inflation. Contemporary Physics, 46(2): 77–91, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Halliwell:2005:HQU

[2127] Jj Halliwell. How the quantum universe became classical. Contemporary Physics, 46(2):93–104, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Baldwin:2005:MHA

[2128] Kenneth Baldwin. Metastable helium: Atom optics with nano-grenades. Contemporary Physics, 46(2):105–120, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Janev:2005:AMP

[2129] Rk Janev. Atomic and molecular physics issues in current thermonuclear fusion research. Contemporary Physics, 46(2):121–140, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRb

[2130] Anonymous. Book reviews. Contemporary Physics, 46(2):141–159, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Renzoni:2005:CAR

[2131] Ferruccio Renzoni. Cold atom realizations of Brownian motors. Contemporary Physics, 46(3):161–171, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Oxborrow:2005:SPS

[2132] Mark Oxborrow and Alastair G. Sinclair. Single-photon sources. Contemporary Physics, 46(3):173–206, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Isham:2005:ITI

[2133] Cj Isham. Is it true; or is it false; or somewhere in between? the logic of quantum theory. Contemporary Physics, 46(3):207–219, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2005:RPG

[2134] Derry W. Jones. Remarkable physicists: from Galileo to Yukawa. Contemporary Physics, 46(3):221–224, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRc

[2135] Anonymous. Book reviews. Contemporary Physics, 46(3):225–247, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blencowe:2005:NS

[2136] Mp Blencowe. Nanoelectromechanical systems. Contemporary Physics, 46(4):249–264, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Arrayas:2005:IPB

[2137] Manuel Arrayás and José L. Trueba. Investigations of pre-breakdown phenomena: streamer discharges. Contemporary Physics, 46(4):265–276, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bender:2005:ISQ

[2138] Carl M. Bender. Introduction to PT -symmetric quantum theory. Contemporary Physics, 46(4):277–292, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:SNa

[2139] Anonymous. Short notice. Contemporary Physics, 46(4):293, 2005. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRd

[2140] Anonymous. Book reviews. Contemporary Physics, 46(4):295–312, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davis:2005:FCS

[2141] A.-C. Davis and Twb Kibble. Fundamental cosmic strings. Contemporary Physics, 46(5):313–322, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Newman:2005:PLP

[2142] M. E. J. Newman. Power laws, Pareto distributions and Zipf’s law. Contemporary Physics, 46(5):323–351, September 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2005:CRP

[2143] Lawrence W. Jones. Cosmic-ray physics in the space age. Contemporary Physics, 46(5):353–371, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:SNb

[2144] Anonymous. Short notices. Contemporary Physics, 46(5):373, 2005. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRe

[2145] Anonymous. Book reviews. Contemporary Physics, 46(5):375–393, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bachor:2005:TLB

[2146] Hans-A. Bachor, Claude Fabre, Ping Koy Lam, and Nicolas Treps. Teaching a laser beam to go straight. Contemporary Physics, 46(6):395–405, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Spiller:2005:IQI

[2147] Timothy P. Spiller, William J. Munro, Sean D. Barrett, and Pieter Kok. An introduction to quantum information processing: applications and realizations. Contemporary Physics, 46(6):407–436, 2005. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Omar:2005:IPQ

[2148] Yasser Omar. Indistinguishable particles in quantum mechanics: an introduction. Contemporary Physics, 46(6):437–448, 2005. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2005:RGR

[2149] E. Sheldon. The regimen of general relativity. Contemporary Physics, 46(6):449–452, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRf

[2150] Anonymous. Book reviews. Contemporary Physics, 46(6):453–470, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRAa

[2151] Anonymous. Books reviewed: author index. Contemporary Physics, 46 (6):471–476, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2005:BRT

[2152] Anonymous. Books reviewed: title index. Contemporary Physics, 46(6): 477–480, 2005. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dowker:2006:CSD

[2153] Fay Dowker. Causal sets as discrete spacetime. Contemporary Physics, 47(1):1–9, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pilkuhn:2006:KGE

[2154] Hartmut Pilkuhn. Klein–Gordon equation for atoms. Contemporary Physics, 47(1):11–24, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Varcoe:2006:TSR

[2155] B. T. H. Varcoe. Testing special relativity using slow light. Contemporary Physics, 47(1):25–32, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berera:2006:WIU

[2156] Arjun Berera. The warm inflationary universe. Contemporary Physics, 47 (1):33–49, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2006:WDP

[2157] E. Sheldon. We delight in physics. Contemporary Physics, 47(1):51– 56, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2006:BRa

[2158] Anonymous. Book reviews. Contemporary Physics, 47(1):57–65, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:2006:QDT

[2159] F. E. Close. Quarks, diquarks, tetraquarks, and pentaquarks. Contemporary Physics, 47(2):67–78, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wiseman:2006:ETB

[2160] H. M. Wiseman. From Einstein’s theorem to Bell’s theorem: a history of quantum non-locality. Contemporary Physics, 47(2):79–88, 2006. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hartmann:2006:MLS

[2161] Michael Hartmann. Minimal length scales for the existence of local temperature. Contemporary Physics, 47(2):89–102, 2006. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ambjorn:2006:US

[2162] J. Ambjørn, J. Jurkiewicz, and R. Loll. The universe from scratch. Contemporary Physics, 47(2):103–117, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Conlon:2006:STL

[2163] Joseph P. Conlon. The string theory landscape: a tale of two hydras. Contemporary Physics, 47(2):119–129, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Klimchitskaya:2006:ETC

[2164] G. L. Klimchitskaya and V. M. Mostepanenko. Experiment and theory in the Casimir effect. Contemporary Physics, 47(3):131–144, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tout:2006:TPO

[2165] Christopher A. Tout. The triple-α process and the origin of the elements. Contemporary Physics, 47(3):145–155, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanchez-Lavega:2006:PVP

[2166] Agustín Sánchez-Lavega. ‘Planetodiversity’: the variety of planets and planetary systems in the Universe. Contemporary Physics, 47(3):157– 165, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2006:HSP

[2167] D. R. Matravers. A history skirmish in physics. Contemporary Physics, 47(3):167–170, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2006:BRb

[2168] Anonymous. Book reviews. Contemporary Physics, 47(3):171–180, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rahman:2006:SDA

[2169] F. Rahman. Superconducting detectors in astronomy. Contemporary Physics, 47(4):181–194, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Battaglia:2006:OTQ

[2170] D. A. Battaglia and L. Stella. Optimization through quantum annealing: theory and some applications. Contemporary Physics, 47(4):195– 208, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haan:2006:PF

[2171] Patrick Voss-De Haan. Physics and fingerprints. Contemporary Physics, 47(4):209–230, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2006:BRc

[2172] Anonymous. Book reviews. Contemporary Physics, 47(4):231–238, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wouters:2006:PBF

[2173] Fred S. Wouters. The physics and biology of fluorescence microscopy in the life sciences. Contemporary Physics, 47(5):239–255, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dunningham:2006:UQT

[2174] Jacob A. Dunningham. Using quantum theory to improve measurement precision. Contemporary Physics, 47(5):257–267, 2006. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Compagno:2006:EPB

[2175] G. Compagno, R. Passante, and F. Persico. Edwin Power and the birth of dressed atoms. Contemporary Physics, 47(5):269–278, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jacobs:2006:SIC

[2176] Kurt Jacobs and Daniel A. Steck. A straightforward introduction to continuous quantum measurement. Contemporary Physics, 47(5):279– 303, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lieb:2006:LII

[2177] Klaus-Peter Lieb and Juhani Keinonen. Luminescence of ion-irradiated α-quartz. Contemporary Physics, 47(5):305–331, 2006. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2006:BRd

Shore:2006:OCD

[2179] B. W. Shore and N. V. Vitanov. Overdamping of coherently driven quantum systems. Contemporary Physics, 47(6):341–362, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2006:CV

[2180] Robert Connon Smith. Cataclysmic variables. Contemporary Physics, 47(6):363–386, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bekenstein:2006:MND

[2181] Jacob D. Bekenstein. The modified Newtonian dynamics — MOND and its implications for new physics. Contemporary Physics, 47(6):387– 403, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2006:BRe

[2182] Anonymous. Book reviews. Contemporary Physics, 47(6):405–411, 2006. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quigg:2007:HBE

[2183] Chris Quigg. Higgs bosons, electroweak symmetry breaking, and the physics of the Large Hadron Collider. Contemporary Physics, 48(1):1– 11, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bose:2007:QCT

[2184] Sougato Bose. Quantum communication through spin chain dynamics: an introductory overview. Contemporary Physics, 48(1):13–30, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Balibar:2007:SS

[2185] Sébastien Balibar. Supersolidity and superfluidity. Contemporary Physics, 48(1):31–39, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nieto:2007:SSP

[2186] Michael Martin Nieto and John D. Anderson. Search for a solution of the Pioneer anomaly. Contemporary Physics, 48(1):41–54, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2007:BRa

[2187] Anonymous. Book reviews. Contemporary Physics, 48(1):55–61, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Alfe:2007:TCE

[2188] D. Alfè, M. J. Gillan, and G. D. Price. Temperature and composition of the Earth’s core. Contemporary Physics, 48(2):63–80, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Echenique:2007:IPF

[2189] Pablo Echenique. Introduction to protein folding for physicists. Contemporary Physics, 48(2):81–108, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cochran:2007:HNS

[2190] William D. Cochran and Michael Endl. Hot Neptunes and super Earths. Contemporary Physics, 48(2):109–117, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Guy:2007:ERT

[2191] Christopher Guy. Essay review: A tragic intellect or a tragedian? [review of Oppenheimer: The Tragic Intellect, by C. Thorpe, The University of Chicago Press, Chicago & London, 2006, pp. 384, £24.00, hardback (ISBN 0-226-79845-3). Scope: sociological biography. Level: nonspecialist]. Contemporary Physics, 48(2):119–121, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2007:BRb

[2192] Anonymous. Book reviews. Contemporary Physics, 48(2):123–130, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brandhuber:2007:TIM

[2193] Andreas Brandhuber, Bill Spence, and Gabriele Travaglini. Twistor inspired methods in gauge theory and gravity. Contemporary Physics, 48 (3):131–142, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Silberhorn:2007:DQL

[2194] Christine Silberhorn. Detecting quantum light. Contemporary Physics, 48(3):143–156, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Banerji:2007:EFR

[2195] J. Banerji. Exploring fractional revivals and sub-Planck structures: a walk in phase space with Wigner and Kirkwood. Contemporary Physics, 48(3):157–171, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2007:VTP

[2196] Peter Bussey. Views over the terrain of physics. Contemporary Physics, 48(3):173–176, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2007:BRc

[2197] Anonymous. Book reviews. Contemporary Physics, 48(3):177–182, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greffet:2007:CTR

[2198] Jean-Jacques Greffet and Carsten Henkel. Coherent thermal radiation. Contemporary Physics, 48(4):183–194, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

King:2007:NM

[2199] S. F. King. Neutrino mass. Contemporary Physics, 48(4):195–211, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thomas:2007:M

[2200] J. Thomas and P. Vahle. Minos. Contemporary Physics, 48(4):213– 226, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2007:BMP

[2201] D. G. C. Jones. The birth of modern physics. Contemporary Physics, 48(4):227–229, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2007:TTH

[2202] David Matravers. Is there a theory of how to construct theories? Contemporary Physics, 48(4):231–234, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2007:BRd

Maher:2007:EMC

[2204] Barbara A. Maher. Environmental magnetism and climate change. Contemporary Physics, 48(5):247–274, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2007:MM

[2205] Stephen J. Blundell. Molecular magnets. Contemporary Physics, 48(5): 275–290, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Melia:2007:BFS

[2206] Tom Melia. Beethoven’s fifth ‘sine’-phony: the science of harmony and discord. Contemporary Physics, 48(5):291–295, 2007. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2007:BRe

[2207] Anonymous. Book reviews. Contemporary Physics, 48(5):297–306, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palus:2007:NCS

[2208] Milan Paluš. From nonlinearity to causality: statistical testing and inference of physical mechanisms underlying complex dynamics. Contemporary Physics, 48(6):307–348, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sudheendra:2007:MIT

[2209] L. Sudheendra, V. Moshnyaga, and K. Samwer. Metal–insulator transition and colossal magnetoresistance: relevance of electron–lattice coupling and electronic phase separation. Contemporary Physics, 48(6): 349–364, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2007:BRf

[2210] Anonymous. Book reviews. Contemporary Physics, 48(6):365–374, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Georgiou:2008:PSO

[2211] Savas Georgiou, Demetrios Anglos, and Costas Fotakis. Photons in the service of our past: lasers in the preservation of cultural heritage. Contemporary Physics, 49(1):1–27, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Charlton:2008:APT

[2212] M. Charlton, S. Jonsell, L. V. Jørgensen, N. Madsen, and D. P. van der Werf. Antihydrogen for precision tests in physics. Contemporary Physics, 49(1):29–41, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Siewert:2008:MEF

[2213] Jens Siewert and Carmelo Ferlito. Mechanical erosion by flowing lava. Contemporary Physics, 49(1):43–54, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2008:BRa

[2214] Anonymous. Book reviews. Contemporary Physics, 49(1):55–69, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Trotta:2008:BSB

[2215] Roberto Trotta. Bayes in the sky: Bayesian inference and model selection in cosmology. Contemporary Physics, 49(2):71–104, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Godin:2008:STT

[2216] Oleg A. Godin. Sound transmission through water–air interfaces: new insights into an old problem. Contemporary Physics, 49(2):105–123, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dowling:2008:QOM

[2217] Jonathan P. Dowling. Quantum optical metrology — the lowdown on high-N00N states. Contemporary Physics, 49(2):125–143, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2008:BRb

[2218] Anonymous. Book reviews. Contemporary Physics, 49(2):145–156, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

VanderWeele:2008:GGD

[2219] Ko Van der Weele. Granular gas dynamics: how Maxwell’s demon rules in a non-equilibrium system. Contemporary Physics, 49(3):157–178, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Agostini:2008:AHI

[2220] P. Agostini and L. F. DiMauro. Atoms in high intensity mid-infrared pulses. Contemporary Physics, 49(3):179–197, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Becker:2008:MES

[2221] W. Becker and H. Rottke. Many-electron strong-field physics. Contemporary Physics, 49(3):199–223, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2008:BRc

[2222] Anonymous. Book reviews. Contemporary Physics, 49(3):225–235, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roussev:2008:EES

[2223] Ilia I. Roussev. Eruptive events in the solar atmosphere: new insights from theory and 3-D numerical modelling. Contemporary Physics, 49(4): 237–254, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Townsend:2008:LDP

[2224] P. D. Townsend, B. Yang, and Y. Wang. Luminescence detection of phase transitions, local environment and nanoparticle inclusions. Contemporary Physics, 49(4):255–280, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carosi:2008:SA

[2225] G. Carosi, K. van Bibber, and M. Pivovaroff. The search for axions. Contemporary Physics, 49(4):281–297, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2008:BRd

[2226] Anonymous. Book reviews. Contemporary Physics, 49(4):299–311, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berman:2008:LS

[2227] Paul R. Berman. Light scattering. Contemporary Physics, 49(5):313– 330, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowburn:2008:LSA

[2228] Russell Cowburn. Laser surface authentication — reading nature’s own security code. Contemporary Physics, 49(5):331–342, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:2008:RH

[2229] F. E. Close. Rumsfeld hadrons. Contemporary Physics, 49(5):343–350, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whitehouse:2008:N

[2230] D. J. Whitehouse. Nanometrology. Contemporary Physics, 49(5):351– 374, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2008:OAY

[2231] E. Sheldon. An overview of almost 20 years’ research on cold fusion. Contemporary Physics, 49(5):375–378, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2008:BRe

[2232] Anonymous. Book reviews. Contemporary Physics, 49(5):379–390, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gabel:2008:FLB

[2233] Christopher V. Gabel. Femtosecond lasers in biology: nanoscale surgery with ultrafast optics. Contemporary Physics, 49(6):391–411, 2008. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2008:QN

[2234] G. J. Milburn and M. J. Woolley. Quantum nanoscience. Contemporary Physics, 49(6):413–433, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jacobs:2008:QFC

[2235] Kurt Jacobs and Alireza Shabani. Quantum feedback control: how to use verification theorems and viscosity solutions to find optimal protocols. Contemporary Physics, 49(6):435–448, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2008:SRN

[2236] Eric Sheldon. A starring role for nuclear physics. Contemporary Physics, 49(6):449–453, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2008:BRf

[2237] Anonymous. Book reviews. Contemporary Physics, 49(6):455–467, 2008. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Editor:2009:IAI

[2238] Peter Knight Editor. Introduction to the 50th anniversary issue of Contemporary Physics. Contemporary Physics, 50(1):1, 2009. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:2009:EPA

[2239] Professor F. Close. Elementary particles (1960) by Abdus Salam. Contemporary Physics, 50(1):3, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salam:2009:EP

[2240] Abdus Salam. Elementary particles. Contemporary Physics, 50(1):5– 22, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2009:DQT

[2241] Dr. P. T. Greenland. The development of quantum theory. Part 1. Formulation and interpretation (1964) by R. E. Peierls. Contemporary Physics, 50(1):23, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peierls:2009:DQT

[2242] R. E. Peierls. The development of quantum theory. Contemporary Physics, 50(1):25–35, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2009:EPS

[2243] Professor A. J. Leggett. Elementary particle symmetries (1965) by T. W. B. Kibble. Contemporary Physics, 50(1):37, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kibble:2009:EPS

[2244] T. W. B. Kibble. Elementary particle symmetries. Contemporary Physics, 50(1):39–55, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2009:TPL

[2245] Dr. P. W. Milonni. Take a photon … (1965) by O. R. Frisch. Contemporary Physics, 50(1):57, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frisch:2009:TPL

[2246] O. R. Frisch. Take a photon …. Contemporary Physics, 50(1):59–67, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2009:JEB

[2247] Professor P. V. E. McClintock. The Josephson effects (1969) by B. W. Petley. Contemporary Physics, 50(1):69, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Petley:2009:JE

[2248] B. W. Petley. The Josephson effects. Contemporary Physics, 50(1): 71–90, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2009:GRT

[2249] Professor R. Joseph. Gravitational radiation: the theoretical aspect (1972) by S. W. Hawking. Contemporary Physics, 50(1):91, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hawking:2009:GRT

[2250] S. W. Hawking. Gravitational radiation: the theoretical aspect. Contemporary Physics, 50(1):93–102, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2009:BHA

[2251] Professor R. Joseph. The black hole in astrophysics: the origin of the concept and its role (1974) by S. Chandrasekhar. Contemporary Physics, 50(1):103, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chandrasekhar:2009:BHA

[2252] S. Chandrasekhar. The black hole in astrophysics: the origin of the concept and its role. Contemporary Physics, 50(1):105–128, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:2009:RCG

[2253] Professor R. Thompson. The Rydberg constant (1974) by G. W. Series. Contemporary Physics, 50(1):129, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Series:2009:RC

[2254] G. W. Series. The Rydberg constant. Contemporary Physics, 50(1): 131–150, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fisher:2009:SCF

[2255] Professor A. Fisher. Snow crystals (1982) by F. C. Frank. Contemporary Physics, 50(1):151, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frank:2009:SC

[2256] F. C. Frank. Snow crystals. Contemporary Physics, 50(1):153–172, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:2009:PSC

[2257] Professor F. Close. Particle substructure. A common theme of discovery in this century (1982) by W. K. H. Panofsky. Contemporary Physics, 50 (1):173, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Panofsky:2009:PSC

[2258] W. K. H. Panofsky. Particle substructure. A common theme of discovery in this century. Contemporary Physics, 50(1):175–196, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stefanovska:2009:SCS

[2259] Professor A. Stefanovska. Soap, cells and statistics — random patterns in two dimensions (1984) by D. Weaire and N. Rivier. Contemporary Physics, 50(1):197, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weaire:2009:SCS

[2260] D. Weaire and N. Rivier. Soap, cells and statistics — random patterns in two dimensions. Contemporary Physics, 50(1):199–239, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eberly:2009:SCH

[2261] Professor J. H. Eberly. Schrödinger’s cat and her laboratory cousins (1984) by A. J. Leggett. Contemporary Physics, 50(1):241, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2009:SCH

[2262] A. J. Leggett. Schrödinger’s cat and her laboratory cousins. Contemporary Physics, 50(1):243–258, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2009:NPV

[2263] Dr. P. W. Milonni. Nothing’s plenty. The vacuum in modern quantum field theory (1985) by I. J. R. Aitchison. Contemporary Physics, 50(1): 259, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aitchison:2009:NPV

[2264] I. J. R. Aitchison. Nothing’s plenty. The vacuum in modern quantum field theory. Contemporary Physics, 50(1):261–319, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:2009:MQM

[2265] Professor R. Thompson. Measuring the quantum mechanical wave function, by M. G. Raymer (1997). Contemporary Physics, 50(1):321, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Raymer:2009:MQM

[2266] M. G. Raymer. Measuring the quantum mechanical wave function. Contemporary Physics, 50(1):323–335, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plenio:2009:QRW

[2267] Professor Martin Plenio. Quantum random walks: an introductory overview (2003) by J. Kempe. Contemporary Physics, 50(1):337, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kempe:2009:QRW

[2268] J. Kempe. Quantum random walks: an introductory overview. Contemporary Physics, 50(1):339–359, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2009:NRJ

[2269] Dr. P. W. Milonni. Negative refraction (2004) by J. B. Pendry. Contemporary Physics, 50(1):361, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pendry:2009:NR

[2270] J. B. Pendry. Negative refraction. Contemporary Physics, 50(1):363– 374, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pachos:2009:MTE

[2271] Jiannis K. Pachos. Manifestations of topological effects in graphene. Contemporary Physics, 50(2):375–389, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Murdin:2009:FIF

[2272] B. N. Murdin. Far-infrared free-electron lasers and their applications. Contemporary Physics, 50(2):391–406, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Park:2009:OAP

[2273] Q-Han Park. Optical antennas and plasmonics. Contemporary Physics, 50(2):407–423, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2009:BRa

[2274] Anonymous. Book reviews. Contemporary Physics, 50(2):425–435, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2009:AMM

[2275] M. Vogel. The anomalous magnetic moment of the electron. Contemporary Physics, 50(3):437–452, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roberts:2009:WSD

[2276] David C. Roberts. When superfluids are a drag. Contemporary Physics, 50(3):453–461, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tsubota:2009:QTS

[2277] Makoto Tsubota. Quantum turbulence: from superfluid helium to atomic Bose–Einstein condensates. Contemporary Physics, 50(3):463–475, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2009:THM

[2278] Eric Sheldon. Two-and-a-half millennia of scientific endeavour and enlightenment. Contemporary Physics, 50(3):477–478, 2009. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2009:BP

[2279] Peter J. Bussey. Beauty and physics. Contemporary Physics, 50(3): 479–482, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2009:BRb

[2280] Anonymous. Book reviews. Contemporary Physics, 50(3):483–494, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beck:2009:GIE

[2281] Christian Beck. Generalised information and entropy measures in physics. Contemporary Physics, 50(4):495–510, 2009. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ditlevsen:2009:CTL

[2282] Peter D. Ditlevsen. Climate transitions on long timescales. Contemporary Physics, 50(4):511–532, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Normand:2009:FFM

[2283] B. Normand. Frontiers in frustrated magnetism. Contemporary Physics, 50(4):533–552, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2009:BRc

Ward:2009:PNS

[2285] Lawrence M. Ward. Physics of neural synchronisation mediated by stochastic resonance. Contemporary Physics, 50(5):563–574, 2009. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hillery:2009:QM

[2286] Mark Hillery and Vladimír Bužek. Quantum machines. Contemporary Physics, 50(5):575–586, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Havey:2009:FLC

[2287] Mark D. Havey. Freezing light with cold atoms. Contemporary Physics, 50(5):587–599, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2009:LRR

[2288] Dr. Peter W. Milonni. Lidar. Range-Resolved Optical Remote Sensing of the Atmosphere, in the Springer Series in Optical Sciences, Vol. 102, edited by Claus Weitkamp, Scope: review. Level: specialist. Contemporary Physics, 50(5):601–602, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

James:2009:BEI

[2289] Dr. B. W. James. Bright Earth — The Invention of Colour, by P. Ball, Scope: monograph. Level: general reader. Contemporary Physics, 50 (5):602, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benedict:2009:QFT

[2290] Dr. Keith Benedict. Quantum Field Theory of Non-Equilibrium States, by J. Rammer, Scope: textbook. Level: students. Contemporary Physics, 50(5):602–603, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Virmani:2009:FTP

[2291] Dr. Shashank Virmani. Field Theory — A Path Integral Approach, by A. Das, Scope: textbook. Level: graduate. Contemporary Physics, 50 (5):604, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2009:CMB

[2292] Professor David Matravers. The Cosmic Microwave Background, by R. Durrer, Scope: textbook/reference text. Level: advanced. Contemporary Physics, 50(5):604–605, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cairns:2009:PPF

[2293] Professor R. A. Cairns. Plasma Physics and Fusion Energy, by J. Freidberg, Scope: monograph. Level: senior undergraduates, beginning postgraduates and researchers. Contemporary Physics, 50(5):605–606, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hall:2009:PLP

[2294] Professor Geoff Hall. Perspectives on LHC Physics, edited by G. Kane and A. Pierce, Scope: survey of possible discoveries at the LHC, the detectors and experimental challenges. Level: graduate students and scientists interested in the LHC. Contemporary Physics, 50(5):606–607, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2009:UXR

[2295] Professor Ken Pounds. Ultraviolet and X-ray Spectroscopy of the Solar Atmosphere, in Cambridge Astrophysics Series, by K. J. H. Phillips, U. Feldman and E. Landi, Scope: monograph. Level: students and senior researchers. Contemporary Physics, 50(5):607–608, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:2009:EQI

[2296] Professor Stephen M. Barnett. Elements of Quantum Information, edited by W. P. Schleich and H. Walther, Scope: overview and progress report. Level: specialist and non-specialist. Contemporary Physics, 50(5): 608–609, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:2009:LQI

[2297] Professor Stephen M. Barnett. Lectures on Quantum Information, edited by D. Bruß and G. Leuchs, Scope: lectures. Level: students and undergraduates. Contemporary Physics, 50(5):609–610, 2009. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2009:QGC

[2298] Professor S. Stenholm. Quantum Gravity, by C. Rovelli, Scope: report. Level: advanced graduate. Contemporary Physics, 50(5):610–611, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McFarland:2009:OMM

[2299] John McFarland. Observing the Moon: The Modern Astronomer’s Guide, 2nd edition, by G. North, Scope: reference guide. Level: general. Contemporary Physics, 50(5):611, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2009:WPE

[2300] Professor Chris Linton. Wave Propagation. From Electrons to Photonic Crystals and Left-Handed Materials, by P. Markoš and C. M. Soukoulis, Scope: textbook. Level: final year undergraduates and first year postgraduates. Contemporary Physics, 50(5):611–612, 2009. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2009:FIA

[2301] Professor G. W. Gibbons. Functional Integration: Action and Symmetries, by P. Cartier and C. DeWitt-Morette, Scope: supplement. Level: graduate level and above. Contemporary Physics, 50(5):612–613, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2009:OSC

[2302] Professor Peter R. Hobson. Observing the Sun with Coronado^{T M} Telescopes, by P. Pugh, Scope: introductory guide. Level: general. Contemporary Physics, 50(5):613, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knight:2009:YFP

[2303] Peter Knight. 50 years at the forefront of physics. Contemporary Physics, 50(6):615–616, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Baker:2009:IPB

[2304] Matthew A. B. Baker and Richard M. Berry. An introduction to the physics of the bacterial flagellar motor: a nanoscale rotary electric motor. Contemporary Physics, 50(6):617–632, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dimopoulos:2009:WGR

[2305] Konstantinos Dimopoulos. Where galaxies really come from. Contemporary Physics, 50(6):633–646, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keyes:2009:PTC

[2306] Robert W. Keyes. Physics, and transistors in computers. Contemporary Physics, 50(6):647–651, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2009:LTW

[2307] Derry W. Jones. Locating targets with waves: radar and sonar. Contemporary Physics, 50(6):653–655, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tennyson:2009:TAT

[2308] Jonathan Tennyson. Is there anybody out there? Contemporary Physics, 50(6):657–659, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2009:ITC

[2309] Dr. Peter R. T. Munro. Introduction to the Theory of Coherence and Polarization of Light, by E. Wolf, Scope: textbook and reference. Level: graduate and practising scientist/engineer. Contemporary Physics, 50(6): 661–662, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2009:INA

[2310] Professor Eric Sheldon. An Introduction to Nuclear Astrophysics, by R. N. Boyd, Scope: textbook. Level: advanced undergraduate, postgraduate and specialist. Contemporary Physics, 50(6):662–663, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2009:QRE

[2311] Dr. P. J. Bussey. Quantum Reflections, edited by J. Ellis and D. Amati, Scope: essays. Level: graduate and general academic physicist. Contemporary Physics, 50(6):663–664, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Altmann:2009:ETY

[2312] Dr. Simon Altmann. 100 Essential Things You Didn’t Know You Didn’t Know, by J. D. Barrow, Scope: general knowledge. Level: general reader.

Contemporary Physics, 50(6):664, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2009:QGP

[2313] Dr. M. Jamil and Professor J. T. Rhee. Quark–Gluon Plasma, by K. Yagi, T. Hatsuda and Y. Miake, Scope: monograph. Level: postgraduate and specialist. Contemporary Physics, 50(6):665–666, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2009:SPE

[2314] Eric Sheldon. Subatomic Physics, 3rd edition, by E. M. Henley and A. García, Scope: textbook. Level: advanced undergraduate, postgraduate and specialist. Contemporary Physics, 50(6):666–667, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2009:DSE

[2315] Professor K. Alan Shore. Dynamical Systems, edited by A. Fathi and J.- C. Yoccoz, Scope: monograph. Level: specialist. Contemporary Physics, 50(6):667, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2009:MTO

[2316] Professor K. Alan Shore. Mathematical Tools for One-Dimensional Dynamics, E. de Faria and W. de Melo, Scope: monograph. Level: specialist. Contemporary Physics, 50(6):667–668, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2009:QCC

[2317] Professor K. Alan Shore. Quantum Computing for Computer Scientists, N. S. Yanofsky and M. A. Manucci, Scope: textbook. Level: undergraduate. Contemporary Physics, 50(6):668–669, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2009:TSU

[2318] Dr. D. G. C. Jones. Three Steps to the Universe, D. Garfinkle and R. Garfinkle, Scope: introductory. Level: general reader and undergraduate. Contemporary Physics, 50(6):669, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ward:2009:PCP

[2319] Professor Ian M. Ward. Polymers: Chemistry and Physics of Modern Materials, 3rd edition, by J. M. G. Cowie and V. Arrighi, Scope: textbook. Level: undergraduate. Contemporary Physics, 50(6):670, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carroll:2009:EPE

[2320] Professor John Carroll. Essentials of Photonics, 2nd edition, by A. Rogers, Scope: textbook covering classical photonic devices and systems. Level: undergraduate or introductory postgraduate courses. Contemporary Physics, 50(6):670–671, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2009:MMM

[2321] Professor F. W. Taylor. Mars: From Myth and Mystery to Recent Discoveries, by M. Hotakainen, Scope: monograph. Level: general reader. Contemporary Physics, 50(6):671, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2009:CNW

[2322] Professor Eric Sheldon. Colliders and Neutrinos: The Window into Physics beyond the Standard Model. Proceedings of the 2006 Theoretical Advanced Summer Institute (TASI-2006) Lectures, edited by S. Dawson and R. N. Mohapatra, Scope: advanced summer school proceedings. Level: postgraduate and specialist. Contemporary Physics, 50(6): 671–673, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Foreman:2009:SN

[2323] Matthew Foreman. Short notices. Contemporary Physics, 50(6):673, 2009. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knight:2010:ONV

[2324] Peter Knight. Obituary notice of Vitaly Ginzburg. Contemporary Physics, 51(1):1, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hutt:2010:OAE

[2325] Axel Hutt. Oscillatory activity in excitable neural systems. Contemporary Physics, 51(1):3–16, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vamivakas:2010:PAA

[2326] A. N. Vamivakas and M. Atatüre. Photons and (artificial) atoms: an overview of optical spectroscopy techniques on quantum dots. Contemporary Physics, 51(1):17–36, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Margolis:2010:OFS

[2327] H. S. Margolis. Optical frequency standards and clocks. Contemporary Physics, 51(1):37–58, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coecke:2010:QP

[2328] Bob Coecke. Quantum picturalism. Contemporary Physics, 51(1):59– 83, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dombey:2010:ERF

[2329] Norman Dombey. Essay review: The first war of physics: Atomic: the First War of Physics and the Secret History of the Atomic Bomb, 1939– 49, by J. Baggott. ISBN 978-1-84831-044-5. Contemporary Physics, 51 (1):85–90, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Johnson:2010:NSP

[2330] Professor R. C. Johnson and Dr M. Oi. Nuclear Superfluidity. Pairing in Finite Systems, by D. M. Brink and R. A. Broglia, Scope: textbook. Level: specialist, postdoctoral and postgraduate. Contemporary Physics, 51(1):91–92, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2010:NPM

[2331] Professor K. Alan Shore. The Noisy Pendulum, by M. Gitterman, Scope: monograph. Level: undergraduate, graduate and general reader. Contemporary Physics, 51(1):92, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2010:CPC

[2332] Dr. Matt Probert. Computational Physics of Carbon Nanotubes, by H. Rafii-Tabar, Scope: introduction/reference. Level: researchers and postgraduate. Contemporary Physics, 51(1):92–93, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Juel:2010:MFD

[2333] Dr. Anne Juel. Magneto-Fluid Dynamics. Fundamentals and Case Studies of Natural Phenomena, by P. Lorrain, F. Lorrain and S. Houle, Scope: introduction. Level: advanced undergraduates and postgraduates. Contemporary Physics, 51(1):93–94, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2010:PPP

[2334] Professor S. J. Blundell. The Pursuit of Perfect Packing, by T. Aste and D. Weaire, Scope: monograph. Level: general reader, postgraduate and specialist. Contemporary Physics, 51(1):94–95, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2010:EOH

[2335] Professor S. J. Blundell. The Earth: Its Origin, History and Physical Constitution, 6th edition, by Sir Harold Jeffreys, Scope: monograph. Level: specialist. Contemporary Physics, 51(1):95–96, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Farmelo:2010:DDP

[2336] Graham Farmelo. Did Dirac predict the positron? Contemporary Physics, 51(2):97–101, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cohen:2010:SLR

[2337] Netta Cohen and Jordan H. Boyle. Swimming at low Reynolds number: a beginners guide to undulatory locomotion. Contemporary Physics, 51(2): 103–123, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jordan:2010:UWU

[2338] Andrew N. Jordan and Alexander N. Korotkov. Uncollapsing the wavefunction by undoing quantum measurements. Contemporary Physics, 51 (2):125–147, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blaum:2010:PTV

[2339] K. Blaum, Yu. N. Novikov, and G. Werth. Penning traps as a versatile tool for precise experiments in fundamental physics. Contemporary Physics, 51(2):149–175, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Juel:2010:FSF

[2340] Dr. Anne Juel. Free Surface Flows Under Gravity Compensated Conditions, by M. E. Dreyer, Scope: review. Level: postgraduates, researchers and specialists. Contemporary Physics, 51(2):177, 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2010:FPC

[2341] Professor Mark Fox. Fundamentals of Photonic Crystal Guiding, by M. Skorobogatiy and J. Yang, Scope: text. Level: graduate students and researchers. Contemporary Physics, 51(2):177–178, 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2010:SLS

[2342] Professor Mark Fox. Slow Light. Science and Applications, edited by J. B. Khurgin and R. S. Tucker, Scope: research monograph. Level: graduate students and researchers. Contemporary Physics, 51(2):178, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2010:FSC

[2343] Professor R. S. MacKay. Facts and Speculations in Cosmology, by J. V. Narlikar and G. Burbidge, Scope: interplay between observation and cosmological models. Level: generally accessible. Contemporary Physics, 51(2):178–180, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ward:2010:EMP

[2344] Professor Ian M. Ward. Electron Microscopy of Polymers, in the series Springer Laboratory, by G. H. Michler, Scope: textbook. Level: postgraduate. Contemporary Physics, 51(2):180–181, 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2010:DUS

[2345] Professor N. C. Wickramasinghe. Dust in the Universe — Similarities and Differences, by K. S. Krishna Swamy, Scope: introduction. Level: graduate students. Contemporary Physics, 51(2):181–182, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2010:MQF

[2346] Professor D. Bailin. Modern Quantum Field Theory: A Concise Introduction, by T. Banks, Scope: textbook. Level: postgraduate. Contemporary Physics, 51(2):182–183, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carter:2010:MME

[2347] Professor Richard Carter. The Method of Moments in Electromagnetics, by W. C. Gibson, Scope: modelling of electromagnetic radiation and scattering problems. Level: postgraduate and professional. Contemporary Physics, 51(2):183–184, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:CQM

[2348] Dr. D. G. C. Jones. Concepts in Quantum Mechanics, by V. Swarup Mathur and S. Singh, Scope: comprehensive textbook. Level: advanced undergraduate and postgraduate. Contemporary Physics, 51(2):184, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2010:BBS

[2349] Professor Allan Solomon. Boulevard of Broken Symmetries: Effective Field Theories of Condensed Matter, by A. M. J. Schakel, Scope: textbook and reference. Level: postgraduate and specialist. Contemporary Physics, 51(2):184–185, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2010:QLD

[2350] Dr. P. J. Bussey. Quantum Leap: From Dirac and Feynman, Across the Universe to Human Body and Mind, by V. Ivancevic and T. Ivancevic, Scope: monograph collection. Level: graduate student and professional research worker. Contemporary Physics, 51(2):185–186, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:CCG

[2351] Professor Peter V. E. McClintock. Chaos and Coarse Graining in Statistical Mechanics, by P. Castiglione, M. Falcioni, A. Lesne and A. Vulpiani, Scope: monograph. Level: professional physicists, researchers and Ph.D. students. Contemporary Physics, 51(2):186–187, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evans:2010:DPC

[2352] Dr. Tim Evans. Dynamical Processes on Complex Networks, by A. Barrat, M. Barthélemy and A. Vespignani, Scope: textbook. Level: graduate students in any scientific disciplines. Contemporary Physics, 51(2): 187–188, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2010:HPT

[2353] Professor Chris Linton. A History of Physical Theories of Comets, From Aristotle to Whipple, by T. Heidarzadeh, Scope: monograph. Level: educated lay person and science historians. Contemporary Physics, 51(2):

188, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:MDL

[2354] Professor Peter V. E. McClintock. Missing Data in Longitudinal Studies: Strategies for Bayesian Modeling and Sensitivity Analysis, by M. J. Daniels and J. W. Hogan. Scope: research monograph. Level: professional statisticians and Ph.D. students. Contemporary Physics, 51(2): 188–189, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Michette:2010:IPN

[2355] Professor Alan Michette. Interactions of Photons and Neutrons with Matter, by S.-H. Chen and M. Kotlarchyk, Scope: textbook. Level: beginning graduates and above, scientists working at synchroton and neutron facilities. Contemporary Physics, 51(2):189, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2010:FPF

[2356] Dr. P. J. Bussey. Fermilab: Physics, The Frontier and Megascience, by L. Hoddeson, A. Kolb and C. Westfall, Scope: historical. Level: general reader. Contemporary Physics, 51(2):190, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:2010:HTF

[2357] Professor Stephen M. Barnett. Hilbert Transforms, by F. W. King, Scope: reference/textbook. Level: specialist. Contemporary Physics, 51(2):190– 191, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2010:LPF

[2358] Dr. Ken Zetie. The Language of Physics: A Foundation for University Study, by J. P. Cullerne and A. Machacek, Scope: undergraduate/preuniversity textbook. Level: physics and engineering students. Contemporary Physics, 51(2):191–192, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2010:DEM

[2359] Salvatore Esposito. The disappearance of Ettore Majorana: an analytic examination. Contemporary Physics, 51(3):193–209, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic). URL

http://adsabs.harvard.edu/abs/2010ConPh..51..193E;   http:/
/www.tandfonline.com/doi/abs/10.1080/00107510903510196.

Mangano:2010:SL

[2360] Michelangelo L. Mangano. The super-LHC. Contemporary Physics, 51(3):211–231, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Soskin:2010:SWQ

[2361] S. M. Soskin, P. V. E. McClintock, T. M. Fromhold, I. A. Khovanov, and R. Mannella. Stochastic webs and quantum transport in superlattices: an introductory review. Contemporary Physics, 51(3):233–248, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palima:2010:GPC

[2362] Darwin Palima and Jesper Glückstad. Generalised phase contrast: microscopy, manipulation and more. Contemporary Physics, 51(3):249–265, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:THR

[2363] Derry W. Jones. Titanic Hungarian refugee physicists. Contemporary Physics, 51(3):267–271, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:LSL

[2364] Derry W. Jones. Lovelock: a sustaining life. Contemporary Physics, 51(3):273–276, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2010:FCC

[2365] Professor Chris Linton. A First Course in Continuum Mechanics, by O. Gonzalez and A. Stuart, Scope: textbook. Level: graduate and advanced undergraduate. Contemporary Physics, 51(3):277, May/June 2010. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2010:EIH

[2366] Professor Peter R. Hobson. Electromagnetic Interactions and Hadronic Structure, edited by F. Close, S. Donnachie and G. Shore, Scope: reference. Level: graduate students and researchers in the field of particle physics. Contemporary Physics, 51(3):277–278, May/June 2010. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matthewman:2010:HSP

[2367] J. Matthewman. A History of the Science and Politics of Climate Change, by B. Bolin, Scope: documentary. Level: general reader. Contemporary Physics, 51(3):278–279, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:JTS

[2368] Professor Stig Stenholm. A Journey through Statistical Physics. Selecta of Jürg Fröhlich, edited by G. Felder, G. M. Graf and K. Hepp, Scope: monograph. Level: research. Contemporary Physics, 51(3):279– 280, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:QFT

[2369] Professor Stig Stenholm. Quantum Field Theory of Non-equilibrium States, by J. Rammer, Scope: monograph. Level: graduate research. Contemporary Physics, 51(3):280, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:HSO

[2370] Professor Stig Stenholm. Hilbert Space Operators in Quantum Physics, 2nd edition, by J. Blank, P. Exner and M. Havlicek, Scope: textbook. Level: graduate. Contemporary Physics, 51(3):281, May/June 2010. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2010:EWF

[2371] Professor David Matravers. Entangled World — The Fascination of Quantum Information and Computation, edited by J. Audretsch (translator R. Ehrlich), Scope: popular science. Level: undergraduate with many parts for a general reader. Contemporary Physics, 51(3):281–282, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:CIH

[2372] Professor D. G. C. Jones. Cambridge Illustrated Handbook of Optoelectronics and Photonics, by S. Kasap, H. Ruda and Y. Boucher, Scope: comprehensive reference. Level: undergraduate and above. Contemporary Physics, 51(3):282–283, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:ITN

[2373] Dr. Manuel Vogel. An Introduction to the Theory of Numbers, 6th edition, by G. H. Hardy and E. M. Wright, Scope: textbook. Level: advanced undergraduate and above. Contemporary Physics, 51(3):283, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barton:2010:AMQ

[2374] Dr. G. Barton. Angular Momentum in Quantum Physics, by L. C. Biedenharn and J. D. Louck, Scope: monograph. Level: postgraduate. Contemporary Physics, 51(3):283–284, May/June 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2010:QTE

[2375] Professor S. J. Blundell. Quantum Theory of the Electron Liquid, by G. F. Giuliani and G. Vignale, Scope: textbook. Level: graduate. Contemporary Physics, 51(3):284–285, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2010:MDE

[2376] Professor K. Alan Shore. Monsoon Dynamics, edited by J. Lighthill and R. Pearce, Scope: monograph. Level: specialist. Contemporary Physics, 51(3):285, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2010:INV

[2377] Professor K. Alan Shore. Introduction to Nanoelectronics, by V. V. Mitin, V. A. Kochelap and M. A. Stroscio, Scope: textbook. Level: undergraduate and graduate. Contemporary Physics, 51(3):285–286, May/ June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2010:PNE

[2378] Professor K. Alan Shore. Progress in Nano-Electro-Optics VI, Springer Series in Optical Sciences, edited by M. Ohtsu, Scope: monograph. Level:specialist. Contemporary Physics, 51(3):286–287, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:CPE

[2379] Peter V. E. McClintock. Cosmic Perspectives, edited by S. K. Biswas, D. C. V. Mallik and C. V. Vishveshwara, Scope: review. Level: general reader. Contemporary Physics, 51(3):287, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2010:SUM

[2380] Dr. J. Rogel-Salazar. The Sun: A User’s Manual, By C. Vita-Finzi, Scope: popular science. Level: general public. Contemporary Physics, 51(3):287–288, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kuhn:2010:CBS

[2381] Axel Kuhn and Daniel Ljunggren. Cavity-based single-photon sources. Contemporary Physics, 51(4):289–313, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cheng:2010:IRQ

[2382] T. Cheng, Q. Su, and R. Grobe. Introductory review on quantum field theory with space–time resolution. Contemporary Physics, 51(4):315– 330, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heckman:2010:FTE

[2383] Jonathan J. Heckman. F-theory and experiment. Contemporary Physics, 51(4):331–348, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ion:2010:WNF

[2384] Sue Ion. The world’s nuclear future — built on material success. Contemporary Physics, 51(4):349–364, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2010:GTO

[2385] Professor Mario Bertolotti. Geometrical and Trigonometric Optics, by E. L. Dereniak and T. D. Dereniak, Scope: textbook. Level: students. Contemporary Physics, 51(4):365, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2010:TSP

[2386] Professor Mario Bertolotti. Terahertz Spectroscopy: Principles and Applications, edited by S. L. Dexheimer, Scope: textbook. Level: beginners. Contemporary Physics, 51(4):366, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:SLD

[2387] Professor Barrie W. Jones. Strategies of Life Detection, edited by O. Botta, J. L. Bada, J. Gómez-Elvira, E. Javaux, F. Selsis and R Summons, Scope: workshop papers. Level: postgraduates/specialists/ advanced hobbyists. Contemporary Physics, 51(4):366–367, 2010. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:HCP

[2388] Dr. Manuel Vogel. Handbook of Charged Particle Optics, 2nd edition, edited by J. Orloff, Scope: handbook/reference. Level: experts/advanced students. Contemporary Physics, 51(4):367, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sambles:2010:ASC

[2389] Professor J. R. Sambles. Aperiodic Structures in Condensed Matter, by E. M. Barber, Scope: monograph. Level: postgraduate. Contemporary Physics, 51(4):367–369, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2010:NPP

[2390] Professor Eric Sheldon. Nuclear and Particle Physics: An Introduction, 2nd edition, by B. R. Martin, Scope: textbook. Level: advanced undergraduate, postgraduate and specialist. Contemporary Physics, 51(4): 369–370, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2010:CMR

[2391] Dr. J. Rogel-Salazar. Classical Mechanics and Relativity, by H. J. W. Müller-Kirsten, Scope: introductory. Level: undergraduate. Contemporary Physics, 51(4):370–371, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2010:AFC

[2392] Professor G. W. Gibbons. Antimatter, by F. Close, Scope: popular account. Level: general readership. Contemporary Physics, 51(4):371– 372, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riley:2010:DPJ

[2393] Professor N. Riley. The Detonation Phenomenon, by J. H. S. Lee, Scope: monograph. Level: postgraduate. Contemporary Physics, 51(4):372–373, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2010:PAN

[2394] Dr. A. H. Harker. Physics and Applications of Negative Refractive Index Materials, by S. Anantha Ramakrishna and T. M. Grzegorczyk, Scope: introduction/reference. Level: newcomer to the field and above. Contemporary Physics, 51(4):373–374, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2010:ATM

[2395] Assistant Professor Nikolaos E. Myridis. Advanced Tomographic Methods in Materials Research and Engineering, edited by J. Banhart, Scope: monograph. Level: undergraduate, postgraduate and postdoctoral workers. Contemporary Physics, 51(4):374, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kolosov:2010:BBM

[2396] Dr. Oleg Kolosov. Biophotonics: Biological and Medical Physics, Biomedical Engineering, edited by L. Pavesi and P. M. Fauchet, Scope: textbook. Level: graduate students in final year to academics. Contemporary Physics, 51(4):375, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kalmus:2010:PPG

[2397] Professor Peter Kalmus. Playing with Planets, by Gerard ’t Hooft, Scope: popular science. Level: general reader. Contemporary Physics, 51(4): 375–376, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:GTA

[2398] Professor Stig Stenholm. A Group-Theoretical Approach to Quantum Optics, by A. B. Klimov and S. M. Chumakov, Scope: monograph. Level: advanced lecturer or researcher. Contemporary Physics, 51(4): 376–377, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:QCM

[2399] Professor Stig Stenholm. Quantum Computing without Magic, by Zdzislaw Meglicki, Scope: textbook. Level: researchers and advanced students. Contemporary Physics, 51(4):377–378, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:IPC

[2400] Professor Stig Stenholm. Information, Physics and Computation, by Marc Mézard and Andrea Montanari, Scope: monograph. Level: specialist researcher. Contemporary Physics, 51(4):378, 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:ACM

[2401] Professor Stig Stenholm. Advanced Condensed Matter Physics, by L. M. Sander, Scope: textbook. Level: advanced undergraduate and beginning graduate. Contemporary Physics, 51(4):378–379, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:QTI

[2402] Professor D. G. C. Jones. Quantum Transport. Introduction to Nanoscience, by Y. V. Nazarov and Y. M. Blanter, Scope: general introduction. Level: postgraduate students and physicists. Contemporary Physics, 51(4):379–380, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2010:PCM

[2403] Professor S. J. Blundell. A Paradigm Called Magnetism, by S. Dattagupta, Scope: general introduction. Level: postgraduate. Contemporary Physics, 51(4):380, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Williams:2010:AES

[2404] Jonathan P. Williams. The astrophysical environment of the solar birthplace. Contemporary Physics, 51(5):381–396, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stoneham:2010:BFA

[2405] Doreen Stoneham and Marshall Stoneham. Beating the forger: authenticating ceramic antiquities. Contemporary Physics, 51(5):397–411, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evans:2010:SPS

[2406] R. M. L. Evans. Statistical physics of shear flow: a non-equilibrium problem. Contemporary Physics, 51(5):413–427, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:SSS

[2407] Derry W. Jones. Sailing: saga and science. Contemporary Physics, 51(5): 429–431, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:MMM

[2408] Derry W. Jones. Magnetic mines, molecular models and the mind. Contemporary Physics, 51(5):433–437, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stefanovska:2010:BBE

[2409] Dr. Aneta Stefanovska. The Biophysical Basis of Excitability, by H. G. Fereira and M. W. Marshall, Scope: research. Level: postgraduate. Contemporary Physics, 51(5):439–440, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sambles:2010:MRS

[2410] Professor J. R. Sambles. Microstructural Randomness and Scaling in Mechanics of Materials, by M. Ostoja-Starzewski, Scope: monograph. Level: postgraduate. Contemporary Physics, 51(5):440, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Boyd:2010:LTL

[2411] Professor T. J. M. Boyd. Let there be Light, by A. Montwill and A. Breslin, Scope: monograph. Level: general reader, undergraduate, graduate physics students. Contemporary Physics, 51(5):440–441, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carroll:2010:MCA

[2412] Professor John E. Carroll. Metamaterials: Critique and Alternatives, by B. A. Munk, Scope: research monograph. Level: masters degree and above. Contemporary Physics, 51(5):441–444, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:SVS

[2413] Dr. Manuel Vogel. Superconductivity — A very short introduction, by S. Blundell, Scope: scientific essay. Level: interested reader. Contemporary Physics, 51(5):444, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Athorne:2010:ATD

[2414] Dr. Chris Athorne. Algebraic Theory of Differential Equations, edited by M. A. H. MacCallum and A. V. Mikhailov, Scope: workshop proceedings. Level: postgraduate/research level. Contemporary Physics, 51(5): 444–445, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:2010:FHE

[2415] Professor John J. Quenby. Foundations of High-Energy Astrophysics, by M. Vietri, Scope: textbook and reference. Level: postgraduate, advanced undergraduate and specialist. Contemporary Physics, 51(5):445– 446, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2010:MCF

[2416] Professor Miguel A. F. Sanjuán. Microscopic Chaos, Fractals and Transport in Nonequilibrium Statistical Mechanics, Advanced Series in Nonlinear Dynamics, Vol. 24, by R. Klages, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 51(5):446–447, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sudbery:2010:PAD

[2417] Dr. Tony Sudbery. Physics and the Art of Dance: Understanding Movement, by K. Laws and A. Sugano, photographs by M. Swope and G. Schiavone, Scope: monograph. Level: general reader and undergraduate. Contemporary Physics, 51(5):447–448, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2010:ILM

[2418] Professor Miguel A. F. Sanjuán. An Introduction to Lagrangian Mechanics, by A. J. Brizard, Scope: textbook. Level: undergraduate and graduate students. Contemporary Physics, 51(5):448, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2010:SRS

[2419] Professor Miguel A. F. Sanjuán. Stochastic Resonance. From Suprathreshold Stochastic Resonance to Stochastic Signal Quantization, by M. D. McDonnell, N. G. Stocks, C. E. M. Pearce and D. Abbott, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 51(5):448–449, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2010:SBG

[2420] Dr. K. P. Zetie. Stargazing Basics: Getting Started in Recreational Astronomy, by P. E. Kinzer, Scope: introductory guide. Level: general reader. Contemporary Physics, 51(5):449–450, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2010:SQB

[2421] Professor Mark Fox. Semiconductor Quantum Bits, edited by F. Henneberger and O. Benson, Scope: research monograph. Level: graduate students and researchers. Contemporary Physics, 51(5):450, 2010. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2010:QHE

[2422] Professor Allan Solomon. Quantum Hall Effects. Field Theoretical Approach and Related Topics, 2nd ed., by Z. F. Ezawa, Scope: encyclopedic compendium. Level: student and researcher. Contemporary Physics, 51 (5):450–451, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:UHD

[2423] Professor Stig Stenholm. The Universe in a Helium Droplet, by G. E. Volovik, Scope: monograph. Level: research level. Contemporary Physics, 51(5):451–452, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:ALV

[2424] Professor Stig Stenholm. Astrophysical Lasers, by V. Letokhov and S. Johansson, Scope: monograph. Level: research level. Contemporary Physics, 51(5):452–453, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bohmer:2010:IQE

[2425] Dr. C. Böhmer. Introduction to Quantum Effects in Gravity, by V. F. Mukhanov and S. Winitzki, Scope: textbook. Level: undergraduate and graduate students, researchers. Contemporary Physics, 51(5):453, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:NVS

[2426] Dr. Manuel Vogel. Nothing: A Very Short Introduction, by F. Close. Scope: scientific essay. Level: — interested reader. Contemporary Physics, 51(5):453–454, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2010:RBF

[2427] Assistant Professor Nikolaos E. Myridis. Radial Basis Functions, by M. D. Buhmann, Scope: monograph. Level: general audience, mathematics theorists. Contemporary Physics, 51(5):454, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2010:JTE

[2428] Professor Geoffrey Duxbury. The Jahn–Teller Effect, by I. B. Bersuker, Scope: monograph. Level: advanced/research. Contemporary Physics, 51(5):455, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2010:PHN

[2429] Professor Peter Hobson. Polarization Holography, by L. Nikolova and P. S. Ramanujam, Scope: monograph. Level: specialist. Contemporary Physics, 51(5):455–456, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:2010:YHP

[2430] Professor Alan Martin. You are Here: A Portable History of the Universe, by C. Potter, Scope: plain text. Level: general reader. Contemporary Physics, 51(5):456, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:2010:AQG

[2431] Professor Ian D. Lawrie. Approaches to Quantum Gravity, edited by Daniele Oriti, Scope: monograph. Level: specialist. Contemporary Physics, 51(5):457, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2010:SCM

[2432] Dr. Matt Probert. Symmetry and Condensed Matter Physics — A Computational Approach, by M. El-Batanouny and F. Wooten, Scope: textbook. Level: active researcher at postgraduation and above. Contemporary Physics, 51(5):457–458, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2010:BCE

[2433] Dr. Stephen H. Ashworth. The Behaviour of Chemical Elements in Stars, by C. Jascheck and M. Jasheck, Scope: reference work. Level: postgraduate/specialist. Contemporary Physics, 51(5):458–459, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2010:GVS

[2434] Professor Ken Pounds. Galaxies: A Very Short Introduction, by J. Gribbin, Scope: monograph. Level: general reader. Contemporary Physics, 51(5):459–460, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2010:FBB

[2435] Professor David Matravers. Finding the Big Bang, edited by P. J. E. Peebles, L. A. Page, Jr. and R. B. Partridge, Scope: history of science. Level: general reader with an interest in cosmology to specialist in history or philosophy of science. Contemporary Physics, 51(5):460–463, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2010:BEC

[2436] Dr. J. Rogel-Salazar. Bose–Einstein Condensation in Dilute Gases, 2nd ed., by C. J. Pethick and H. Smith. Scope: textbook. Level: advanced undergraduate and postgraduate. Contemporary Physics, 51(5):463, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2010:CIN

[2437] Professor R. S. MacKay. The Correspondence of Isaac Newton, Volume V, 1709–1713, edited by A. Rupert Hall and L. Tilling, Scope: reference. Level: specialist in history of science. Contemporary Physics, 51(5): 463–464, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:SMB

[2438] Professor Peter V. E. McClintock. Sirius Matters, by N. Brosch, Scope: monograph. Level: undergraduates, graduate students, amateur astronomers, educated public. Contemporary Physics, 51(5):464–465, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2010:ONS

[2439] Professor S. J. Blundell. Observing the Night Sky with Binoculars, by S. J. O’Meara, Scope: popular astronomy. Level: general. Contemporary Physics, 51(5):465, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roussev:2010:SSM

[2440] Professor Ilia I. Roussev. Solar and Stellar Magnetic Activity, by C. Schrijver and C. Zwaan, Scope: textbook. Level: specialist, postgraduate, graduate and undergraduate. Contemporary Physics, 51(5):466, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roussev:2010:CDH

[2441] Professor Ilia I. Roussev. Cool Discs, Hot Flows: The Varying Faces of Accreting Compact Objects, edited by M. Axelsson, Scope: conference proceedings. Level: specialist, postgraduate and graduate. Contemporary Physics, 51(5):466–467, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2010:APP

[2442] Dr. Robert C. Smith. Astrophysics Processes: The Physics of Astronomical Phenomena, by H. Bradt, Scope: textbook. Level: senior undergraduate/beginning postgraduate. Contemporary Physics, 51(5):467, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sangoro:2010:PLC

[2443] Joshua R. Sangoro. Polymers, Liquids and Colloids in Electric Fields, edited by Y. Tsori and U. Steiner, Scope: textbook. Level: specialist and non-specialist. Contemporary Physics, 51(5):467–468, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2010:QCI

[2444] Dr. S. C. Tiwari. From Quantum Cohomology to Integrable Systems, by M. A. Guest, Scope: textbook. Level: graduate, researchers. Contemporary Physics, 51(5):468–469, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palffy:2010:NEA

[2445] Adriana Pálffy. Nuclear effects in atomic transitions. Contemporary Physics, 51(6):471–496, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gerry:2010:POQ

[2446] Christopher C. Gerry and Jihane Mimih. The parity operator in quantum optical metrology. Contemporary Physics, 51(6):497–511, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zavrtanik:2010:UHE

[2447] Danilo Zavrtanik. Ultra high energy cosmic rays. Contemporary Physics, 51(6):513–529, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2010:QQ

[2448] P. J. Bussey. Quantum quicksands. Contemporary Physics, 51(6):531– 534, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2010:SEL

[2449] Professor Ken Pounds. Space Enterprise — Living and Working Offworld in the 21st Century, by Philip R. Harris, Scope: source book. Level: general reader and specialist. Contemporary Physics, 51(6):535, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

vanderWeele:2010:LDU

[2450] Professor Ko van der Weele. The Light/Dark Universe. Light from Galaxies, Dark Matter and Dark Energy, by J. M. Overduin and P. S. Wesson, Scope: monograph. Level: informed reader interested in astronomy. Contemporary Physics, 51(6):535–536, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2010:PMN

[2451] Professor Stig Stenholm. Philosophy of Mathematics and Natural Science, by H. Weyl. Source: monograph. Level. researchers and advanced students. Contemporary Physics, 51(6):537–539, 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2010:TGF

[2452] Emeritus Professor Derry W. Jones. Turbulence, by Giles Foden, Scope: wartime meteorology and fiction. Level: general reader. Contemporary Physics, 51(6):539–540, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:FSP

[2453] Dr. Manuel Vogel. Future Spacecraft Propulsion Systems, 2nd ed., by P. A. Czysz and C. Bruno, Scope: electric rockets explained, IAC, propulsion space, spacecraft engines. Level: advanced students and experts on the field. Contemporary Physics, 51(6):540–541, 2010. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vedral:2010:QSM

[2454] Professor Vlatko Vedral. Quantum Statistical Mechanics, by W. C. Schieve and L. P. Horwitz, Scope: basic theory, with applications from quantum optics to quantum gravity. Level: researchers and advanced students. Contemporary Physics, 51(6):541–542, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2010:QIO

[2455] Assistant Professor Nikolaos E. Myridis. Quantum Information, Oxford Master Series in Physics, by Stephen M. Barnett, Scope: quantum information, quantum computing and the foundations of quantum theory. Level: undergraduate, graduate and postgraduate. Contemporary Physics, 51(6):542–543, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wallis:2010:CIK

[2456] Dr. Max Wallis. Cosmic Imagery: Key Images in the History of Science, by John D. Barrow, Scope: key concepts. Level: general reader. Contemporary Physics, 51(6):543–544, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2010:CSP

[2457] Dr. M. Jamil. Computer Solutions in Physics: With Applications in Astrophysics, Biophysics, Differential Equations, and Engineering, by Steve VanWyk, Scope: monograph. Level: undergraduate students. Contemporary Physics, 51(6):544–545, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2010:WMS

[2458] Professor M. Bertolotti. Waves in Metamaterials, by L. Solymar and E. Shamonina, Scope: textbook. Level: researchers and advanced students. Contemporary Physics, 51(6):545–546, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2010:UMB

[2459] Dr. Peter J. Bussey. Universe or Multiverse?, by Bernard Carr, Scope: review. Level: physicists and philosophers advanced undergraduate and above. Contemporary Physics, 51(6):546–550, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pierri-Galvao:2010:NPT

[2460] Dr. Monica Pierri-Galvao. The New Physics: For the Twenty-first Century, by Gordon Fraser, Scope: textbook. Level: undergraduate. Contemporary Physics, 51(6):550–551, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:DSO

[2461] Professor Peter V. E. McClintock. Dynamics of Self-Organized and Self-Assembled Structures, by Rashmi C. Desai and Raymond Kapral, Scope: monograph. Level: professional physicists, researchers, Ph.D. students, advanced undergraduates. Contemporary Physics, 51(6):551–552, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2010:MTI

[2462] Dr. Stephen H. Ashworth. Measured Tones: The Interplay of Physics and Music, 3rd edition, by Ian Johnston, Scope: monograph. Level: general reader. Contemporary Physics, 51(6):552, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2010:BSA

[2463] Dr. Stephen H. Ashworth. Building Scientific Apparatus, 4th edition, by John H. Moore, Christopher C. Davis, Michael A. Coplan and Sandra E. Greer, Scope: reference. Level: anyone using, designing, modifying or specifying scientific apparatus. Contemporary Physics, 51(6): 552–553, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2010:FMI

[2464] Dr. Peter R. T. Munro. Fundamentals of Medical Imaging, by Paul Suetens, Scope: textbook and reference. Level: undergraduate, graduate and practicing scientist/engineer. Contemporary Physics, 51(6): 553–554, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2010:CTP

[2465] Dr. Peter R. T. Munro. Computed Tomography: From Photon Statistics to Modern Cone-Beam CT, by Thorsten M. Buzug, Scope: textbook and reference. Level: graduate and practising scientist/engineer. Contemporary Physics, 51(6):554–555, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2010:ENF

[2466] Dr. Peter W. Milonni. Electronic Noise and Fluctuations in Solids, by Sh. Kogan, Scope: general reader, postgraduate and specialist. Level: anyone interested in the theory of fluctuation phenomena. Contemporary Physics, 51(6):555, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2010:FEG

[2467] Professor Ken Pounds. Formation and Evolution of Galaxy Bulges (IAU S245), Proceedings of the IAU Symposia and Colloquia, edited by Martin Bureau, E. Athanassoula, and Beatriz Barbuy Scope: wide-ranging review. Level: graduate students and active researchers. Contemporary Physics, 51(6):556, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2010:SSC

[2468] Dr. Robert C. Smith. Stellar Spectral Classification, by Richard O. Gray and Christopher J. Corbally, Scope: monograph and reference book. Level: advanced undergraduates, postgraduates and specialists. Contemporary Physics, 51(6):556–557, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2010:IPI

[2469] Professor N. C. Wickramasinghe. An Introduction to the Physics of Interstellar Dust, by E. Krugel, Scope: textbook. Level: workers in the field of interstellar dust. Contemporary Physics, 51(6):557, 2010. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:VOI

[2470] Dr. Manuel Vogel. In Vivo Optical Imaging of Brain Function, 2nd ed., edited by Ron D. Frostig, Scope: monograph. Level: advanced graduate and postgraduate researchers. Contemporary Physics, 51(6):557– 558, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2010:QFL

[2471] Dr. Peter J. Bussey. The Quantum Frontier: The Large Hadron Collider, by D. Lincoln, Scope: general interest. Level: general readership, undergraduate. Contemporary Physics, 51(6):558–559, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2010:PMM

[2472] Dr. Matt Probert. Physics Meets Mineralogy: Condensed-Matter Physics in Geosciences, edited by Hideo Aoki, Yasuhiko Syono and Russell J. Hemley, Scope: geophysicists, mineralogists, condensed matter physicists. Level: graduate level textbook. Contemporary Physics, 51(6): 559, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:2010:OLU

[2473] Professor Alan Martin. Origins of Life in the Universe, by R. Jastrow and M. Rampino, Scope: illustrated text for non-science students. Level: general reader. Contemporary Physics, 51(6):559–560, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:EGT

[2474] Professor Peter V. E. McClintock. Engineering the Guitar: Theory and Practice, by Richard Mark French, Scope: monograph. Level: undergraduate physicists and engineers, musicologists, guitar builders and players. Contemporary Physics, 51(6):560–561, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2010:ISP

[2475] Dr. Manuel Vogel. Introduction to Statistical Physics, 2nd ed., by Kerson Huang, Scope: textbook. Level: undergraduate students. Contemporary Physics, 51(6):561, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2010:PSD

[2476] Professor Peter V. E. McClintock. Planetary Systems: Detection, Formation and Habitability of Extrasolar Planets, by M. Ollivier, T. Encrenaz, F. Roques, F. Selsis and F. Casoli, Scope: monograph. Level: undergraduate and postgraduate physicists and researchers. Contemporary Physics, 51(6):561–562, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Foreman:2010:SN

[2477] Matthew R. Foreman. Short notices. Contemporary Physics, 51(6):562– 563, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2010:EB

[2478] Anonymous. Editorial board. Contemporary Physics, 51(6):ebi, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Centrella:2011:MBH

[2479] Joan Centrella, John Baker, Bernard Kelly, and James van Meter. Merging black holes. Contemporary Physics, 52(1):1–14, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carter:2011:ATC

[2480] Richard G. Carter. Acceleration technologies for charged particles: an introduction. Contemporary Physics, 52(1):15–41, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eling:2011:GGT

[2481] Christopher Eling, Itzhak Fouxon, and Yaron Oz. Gravity and a geometrisation of turbulence: an intriguing correspondence. Contemporary Physics, 52(1):43–54, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nugent:2011:MPT

[2482] Keith A. Nugent. The measurement of phase through the propagation of intensity: an introduction. Contemporary Physics, 52(1):55–69, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2011:QAL

[2483] Peter V. E. McClintock. Quantum aspects of life. Contemporary Physics, 52(1):71–73, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:WOC

[2484] Dr. Manuel Vogel. Waves in Oceanic and Coastal Waters, by Leo H. Holthuijsen, Scope: textbook. Level: advanced undergraduate and above. Contemporary Physics, 52(1):75, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2011:OPR

[2485] Professor Mario Bertolotti. Optical Pattern Recognition, edited by F. T. S. Yu and S. Jutamulia, Scope: review. Level: graduate students and researchers. Contemporary Physics, 52(1):75–76, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2011:EDW

[2486] Professor Mario Bertolotti. The Essence of Dielectric Waveguides, by C. Yeh and F. Shimabukuro, Scope: overview. Level: researchers and students in optical communications, photonics and applied electromagnetics. Contemporary Physics, 52(1):76–77, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2011:ESB

[2487] Dr. Matt Probert. Electronic Structure: Basic Theory and Practical Methods, by Richard M. Martin, Scope: graduate level textbook. Level: theoretical materials scientists/condensed matter physicists/computational chemists. Contemporary Physics, 52(1):77, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tarleton:2011:ASM

[2488] Dr. Edmund Tarleton. Applied Solid Mechanics, by P. Howell, G. Kozyreff and J. Ockendon, Scope: textbook. Level: graduate and advanced undergraduate. Contemporary Physics, 52(1):77–79, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2011:IPP

[2489] Dr. M. Jamil. An Introduction to the Physics of Particle Accelerators, by Mario Conte and William W. MacKay, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 52(1):79–80, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:QMB

[2490] Dr. Peter J. Bussey. Quantum Mechanics with Basic Field Theory, by Bipin R. Desai, Scope: course book. Level: graduate. Contemporary Physics, 52(1):80, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2011:MPE

[2491] Professor Michael Belsley. Mrs. Perkins’s Electric Quilt and Other Intriguing Stories of Mathematical Physics, by Paul J. Nahin, Scope: general. Level: undergraduate. Contemporary Physics, 52(1):80–81, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2011:HSM

[2492] Dr. S. Tiwari. Hilbert Space Methods in Quantum Mechanics, by Werner O. Amrein, Scope: handbook. Level: undergraduate. Contemporary Physics, 52(1):81–82, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2011:UOD

[2493] Professor Peter V. E. McClintock. The Universe — Order Without Design, by Carlos I. Calle, Scope: monograph. Level: general reader. Contemporary Physics, 52(1):82–83, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2011:MMM

[2494] Dr. A. M. Fox. Magnetism and Magnetic Materials, by J. M. D. Coey, Scope: handbook. Level: advanced undergraduates, postgraduates and researchers. Contemporary Physics, 52(1):83–84, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:DRU

[2495] Dr. Peter J. Bussey. Decoding Reality: The Universe as Quantum Information, by Vlatko Vedral, Scope: monograph. Level: general reader. Contemporary Physics, 52(1):84–85, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2011:AAN

[2496] Professor Eric Sheldon. Alex’s Adventures in Numberland, by Alex Bellos, Scope: general interest. Level: general readership. Contemporary Physics, 52(1):85–86, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2011:SPP

[2497] Dr. Peter W. Milonni. Statistical Physics of Particles, by Mehran Kardar, Scope: textbook. Level: advanced undergraduate and graduate. Contemporary Physics, 52(1):86–87, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:QRT

[2498] Dr. Peter J. Bussey. Quantum Reality: Theory and Philosophy, by Jonathan Allday, Scope: general interest. Level: general readership. Contemporary Physics, 52(1):87, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:PCR

[2499] Assistant Professor Nikolaos E. Myridis. Perspectives in Computation, by Robert Geroch, Scope: theory of computation. Level: graduate and advanced undergraduate students, scientists. Contemporary Physics, 52 (1):87–88, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2011:GFD

[2500] Dr. Amgad Salama. Geological Fluid Dynamics: Sub-Surface Flow and Reactions, by O. M. Phillips, Scope: textbook. Level: graduate. Contemporary Physics, 52(1):88–89, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2011:RSS

[2501] Dr. Robert C. Smith. The Rotation of Sun and Stars, by J.-P. Rozelot and C. Neiner/Physics, Formation and Evolution of Rotating Stars, by A. Maeder, Scope: monograph and reference. Level: postgraduates and specialists/scope: monograph. Level: postgraduates, specialists and advanced undergraduates. Contemporary Physics, 52(1):89–91, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2011:MTS

[2502] Dr. Stephen H. Ashworth. Modern Thermodynamics with Statistical Mechanics, by Carl S. Helrich, Scope: textbook. Level: advanced undergraduate/postgraduate. Contemporary Physics, 52(1):91–92, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stoneham:2011:CNE

[2503] Emeritus Massey Professor Marshall Stoneham. Chirality at the Nanoscale, edited by David B. Amabilino, Scope: monograph. Level: graduate student/research scientist. Contemporary Physics, 52(1):92– 93, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2011:AST

[2504] Richard Ellis. Astronomical Spectrographs and Their History, by John J. Hearnshaw, Scope: monograph. Level: science historians, general scientific, undergraduates. Contemporary Physics, 52(1):93, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McFarland:2011:CDS

[2505] John McFarland. The Cambridge Double Star Atlas, by J. Mullaney and W. Tirion, Scope: handbook. Level: general reader and specialists. Contemporary Physics, 52(1):93–94, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:SFS

[2506] Emeritus Professor Derry W. Jones. Seeing Further: The Story of Science and the Royal Society, edited by Bill Bryson, Scope: articles on history, philosophy and a range of science. Level: general reader. Contemporary Physics, 52(1):94–95, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:INS

[2507] Dr. Manuel Vogel. Introduction to Nuclear Science, by J. C. Bryan, Scope: textbook. Level: undergraduate. Contemporary Physics, 52(1): 95–96, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:MWH

[2508] Dr. Peter J. Bussey. The Many Worlds of Hugh Everett III, by Peter Byrne, Scope: biography. Level: undergraduate physicist and philosopher and above. Contemporary Physics, 52(1):96–97, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carroll:2011:SOT

[2509] Professor John E. Carroll. Superstrings and Other Things: A Guide to Physics, 2nd ed., by Carlos I. Calle, Scope: manual. Level: general scientific reader. Contemporary Physics, 52(1):97, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2011:BSP

[2510] Professor Miguel A. F. Sanjuán. Basics of Statistical Physics, by Harald J. W. Müller-Kirsten, Scope: textbook, monograph. Level: undergraduate and graduate students. Contemporary Physics, 52(1):97–98, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2011:CTP

[2511] Professor Miguel A. F. Sanjuán. Concepts in Thermal Physics, 2nd ed., by Stephen J. Blundell and Katherine M. Blundell, Scope: handbook. Level: undergraduate and graduate students. Contemporary Physics, 52 (1):98, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:SCC

[2512] Assistant Professor Nikolaos E. Myridis. Scientific Computing with Case Studies, by Dianne P. O’Leary, Scope: textbook. Level: undergraduate and early graduate students. Contemporary Physics, 52(1):98–99, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2011:UFF

[2513] Dr. Amgad Salama. Understanding Fluid Flow, by Grae Worster, Scope: general interest. Level: undergraduate. Contemporary Physics, 52(1): 99–100, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Krucken:2011:ISS

[2514] Reiner Krücken. Introduction to shell structure in exotic nuclei. Contemporary Physics, 52(2):101–120, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pettinelli:2011:RWT

[2515] Elena Pettinelli, Pier Matteo Barone, Elisabetta Mattei, and Sebastian Emanuel Lauro. Radio wave techniques for non-destructive archaeological investigations. Contemporary Physics, 52(2):121–130, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keeling:2011:EPC

[2516] Jonathan Keeling and Natalia G. Berloff. Exciton–polariton condensation. Contemporary Physics, 52(2):131–151, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:DMP

[2517] Assistant Professor Nikolaos E. Myridis. Discrete Mathematics: Proof, Structures, and Applications, 3rd edn., by Rowan Garnier and John Taylor, Scope: text-book. Level: undergraduate computer science students. Contemporary Physics, 52(2):153, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:MMU

[2518] Assistant Professor Nikolaos E. Myridis. The Mathematical Mechanic: Using Physical Reasoning to Solve Problems, by Mark Levi, Scope: handbook. Level: students, scientists, physicists, mathematicians. Contemporary Physics, 52(2):154, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:EPD

[2519] D. G. C. Jones. The Edge of Physics: Dispatches from the Frontiers of Cosmology, by Anil Ananthaswamy, Scope: synoptic picture of experimental modern astronomy, cosmology and particle physics. Level: general readers, a level and undergraduate physicists, specialists in other fields. Contemporary Physics, 52(2):155, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2011:GVP

[2520] Dr. Amgad Salama. Good Vibrations: The Physics of Music, by Barry Parker, Scope: textbook. Level: general reader. Contemporary Physics, 52(2):155–156, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:IRC

[2521] Professor J. Rogel-Salazar. Interest Rates and Coupon Bonds in Quantum Finance, by Belal E. Baaquie, Scope: monograph. Level: specialist. Contemporary Physics, 52(2):156, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2011:MMF

[2522] Professor S. J. Blundell. Micromagnetism and the Microstructure of Ferromagnetic Solids, by Helmut Kronmüller and Manfred Fähnle, Scope: monograph. Level: researcher. Contemporary Physics, 52(2):157, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2011:AMM

[2523] Professor Chris Linton. Applied Mathematical Methods in Theoretical Physics, 2nd edn., by Michio Masujima, Scope: reference. Level: undergraduates, postgraduates, engineers. Contemporary Physics, 52(2): 157–158, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:2011:QFT

[2524] Professor Ian D. Lawrie. Quantum Field Theory in Curved Spacetime: Quantized Fields and Gravity, by L. Parker and D. Toms, Scope: monograph. Level: postgraduate, researcher, specialist. Contemporary Physics, 52(2):158–159, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2011:UQF

[2525] Professor Peter V. E. McClintock. Ultracold Quantum Fields, by H. T. C. Stoof, K. B. Gubbels and D. B. M. Dickerscheid, Scope: monograph. Level: researchers, postgraduate physicists and senior undergraduates. Contemporary Physics, 52(2):159, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McFarland:2011:TOD

[2526] John McFarland. Theory of Orbit Determination, by Andrea Milani and Giovanni F. Gronchi, Scope: reference. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 52(2):160, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benhamou:2011:PGGa

[2527] Mabrouk Benhamou. P. G. de Gennes’ Impact on Science: Volume I, edited by Julien Bok, Jacques Prost and Françoise Brochard-Wyart, Scope: handbook. Level: general readership. Contemporary Physics, 52 (2):160–161, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2011:SCG

[2528] Dr. M. Jamil. Scientific Computation, by Gaston H. Gonnet and Ralf Scholl, Scope: monograph. Level: graduate level students. Contemporary Physics, 52(2):161–162, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2011:CNL

[2529] Dr. A. H. Harker. Concepts of Nanochemistry, by Ludovico Cademartiri and Geoffrey A. Ozin, Scope: monograph. Level: undergraduate and postgraduate. Contemporary Physics, 52(2):162, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2011:QMC

[2530] Professor Allan Solomon. Quantum Measurement and Control, by H. M. Wiseman and G. J. Milburn, Scope: monograph. Level: postgraduate and researcher. Contemporary Physics, 52(2):162–163, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2011:IRJ

[2531] Dr. R. S. Shorter. An Introduction to Relativity, by Jayant V. Narlikar, Scope: review. Level: undergraduate and postgraduate. Contemporary Physics, 52(2):163–164, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2011:CVP

[2532] Dr. R. S. Shorter. Cremona Violins: A Physicist’s Quest for Secrets of Stradivari, by Kameshwar C. Wali, Scope: general interest. Level: general readership. Contemporary Physics, 52(2):164, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garrett:2011:IPM

[2533] Dr. Natalie L. Garrett. Introduction to Physics in Modern Medicine, 2nd edn., by Suzanne Amador Kane, Scope: textbook. Level: undergraduate. Contemporary Physics, 52(2):164–165, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2011:DME

[2534] Dr. Giovanni Volpe. Darwin Meets Einstein. On the Meaning of Science, by Frans W. Saris, Scope: popular account. Level: general reader. Contemporary Physics, 52(2):165–166, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sambles:2011:SPP

[2535] Professor J. R. Sambles. Symmetries in Physics: Philosophical Reflections, edited by Katherine Brading and Elena Castellani, Scope: edited volume. Level: postgraduate. Contemporary Physics, 52(2):166–168, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vedral:2011:QPS

[2536] Professor Vlatko Vedral. Quantum Processes, Systems, and Information, by Benjamin Schumacher and Michael Westmoreland, Scope: review. Level: undergraduate and postgraduate. Contemporary Physics, 52(2): 168, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2011:USS

[2537] Dr. A. H. Harker. Understanding Solid State Physics, by Sharon Ann Holgate, Scope: monograph. Level: undergraduate. Contemporary Physics, 52(2):169, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:SDH

[2538] Dr. D. G. C. Jones. Semiconductor Devices for High-Speed Optoelectronics, by Giovanni Ghione, Scope: monograph and review. Level: undergraduate, postgraduate and early career researcher. Contemporary Physics, 52(2):169–170, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2011:PDP

[2539] Professor David Matravers. The Primordial Density Perturbation: Cosmology, Inflation and the Origin of Structure, by David H. Lyth and Andrew R. Liddle, Scope: monograph. Level: postgraduate, early career researcher. Contemporary Physics, 52(2):170–171, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2011:NRQ

[2540] Professor Allan Solomon. Non-relativistic Quantum Theory: Dynamics, Symmetry and Geometry, by Kai S. Lam, Scope: monograph. Level: undergraduate and postgraduate. Contemporary Physics, 52(2):171, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garrett:2011:BAN

[2541] Dr. Natalie L. Garrett. Biochemical Applications of Nonlinear Optical Spectroscopy, edited by Vladislav V. Yakovlev, Scope: edited book. Level: advanced undergraduate, postgraduate and researcher. Contemporary Physics, 52(2):171–172, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2011:LHC

[2542] Professor Peter R. Hobson. The Large Hadron Collider: A Marvel of Technology, edited by Lyndon Evans, Scope: edited book. Level: general reader and non-specialists. Contemporary Physics, 52(2):172–173, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:LPD

[2543] Assistant Professor Nikolaos E. Myridis. Linear Partial Differential Equations and Fourier Theory, by Marcus Pivato, Scope: monograph. Level: students in mathematics and physics, lecturers. Contemporary Physics, 52(2):173–174, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2011:SIT

[2544] Professor Thomas J. Bridges. Solitons, Instantons, and Twistors, by Maciej Dunajski, Scope: review. Level: undergraduate, postgraduate, early career researcher. Contemporary Physics, 52(2):174–175, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:CMT

[2545] Dr. Peter J. Bussey. Causality, Measurement Theory and the Differentiable Structure of Space–Time, by R. N. Sen, Scope: monograph. Level: specialist. Contemporary Physics, 52(2):175, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sambles:2011:EAB

[2546] Professor J. R. Sambles. Electromagnetic Anisotropy and Bianisotropy: A Field Guide, by Tom G. Mackay and Akhlesh Lakhtakia, Scope: monograph. Level: postgraduate, early career researcher, specialist. Contemporary Physics, 52(2):176–177, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:2011:SLT

[2547] Professor J. I. B. Wilson. Spectroscopy of Low Temperature Plasma, by V. N. Ochkin translated by S. Kittell, Scope: monograph. Level: researcher. Contemporary Physics, 52(2):177, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:FFC

[2548] Emeritus Professor Derry W. Jones. On Fact and Fraud: Cautionary Tales from the Front Lines of Science, by David Goodstein, Scope: twentieth century physics history and advice for scientists and administrators. Level: general reader. Contemporary Physics, 52(2):178–179, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Banerji:2011:MOL

[2549] J. Banerji. The method of (not so) ordinary least squares: what can go wrong and how to fix them. Contemporary Physics, 52(3):181–195, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2011:GSR

[2550] G. W. Gibbons and C. M. Warnick. The geometry of sound rays in a wind. Contemporary Physics, 52(3):197–209, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Redondo:2011:LST

[2551] Javier Redondo and Andreas Ringwald. Light shining through walls. Contemporary Physics, 52(3):211–236, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:ERI

[2552] J. Rogel-Salazar. Essay review: Improving science literacy: from atoms to galaxies. From Atoms to Galaxies by Sadri Hassani. Contemporary Physics, 52(3):237, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Foreman:2011:OM

[2553] Matthew R. Foreman. Opus musicum. Contemporary Physics, 52(3): 239–242, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2011:PRB

[2554] S. C. Tiwari. Physical reality beyond the Standard Models: revolution in fundamental physics? Contemporary Physics, 52(3):243–248, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2011:TDO

[2555] Michael Belsley. Two-dimensional Optical Spectroscopy, by Minhaeng Cho, Scope: monograph. Level: postgraduate, researcher, specialist. Contemporary Physics, 52(3):249–250, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2011:IMT

[2556] Amgad Salama. An Introduction to Modeling of Transport Processes, by Ashim Datta and Vineet Rakesh, Scope: reference. Level: undergraduate. Contemporary Physics, 52(3):250–251, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shamir:2011:OMM

[2557] Joseph Shamir. Optical Methods of Measurement: Wholefield Techniques, 2nd edn., by R. S. Sirohi, Scope: monograph and reference. Level: early career researcher and engineers. Contemporary Physics, 52(3):251, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sudbery:2011:QEB

[2558] Tony Sudbery. Quantum: Einstein, Bohr and the Great Debate About the Nature of Reality, by Manjit Kumar, Scope: general interest. Level: general readership. Contemporary Physics, 52(3):251–254, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:PSF

[2559] Manuel Vogel. Physics of the Sun: A First Course, by D. J. Mullan, Scope: monograph. Level: undergraduates. Contemporary Physics, 52 (3):254, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2011:SME

[2560] Michael Belsley. Statistical Mechanics: An Elementary Outline, revised edition, by Lahiri Avijit, Scope: monograph. Level: undergraduate. Contemporary Physics, 52(3):254–255, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gotte:2011:IST

[2561] Jörg Bernhard Götte. Intermediate Spectral Theory and Quantum Dynamics, by C. R. de Oliveira, Scope: monograph. Level: undergraduate and postgraduate. Contemporary Physics, 52(3):255–256, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Laberge:2011:CEC

[2562] Yves Laberge. Cracking the Einstein Code: Relativity and the Birth of Black Hole Physics, by F. Melia, Scope: monograph. Level: general readership. Contemporary Physics, 52(3):256, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2011:INS

[2563] Mark Fox. Introduction to Nanophotonics, Sergey V. Gaponenko, Scope: manual. Level: graduate students and researchers. Contemporary Physics, 52(3):257, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:MHG

[2564] Peter J. Bussey. Massive: The Hunt for the God Particle, by Ian Sample, Scope: general interest. Level: general readership and postgraduate. Contemporary Physics, 52(3):257–258, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:KLO

[2565] Peter J. Bussey. Kinematics: The Lost Origins of Einstein’s Relativity, by Alberto A. Martínez, Scope: biography and reference. Level: general readership, teacher, specialist. Contemporary Physics, 52(3):258, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2011:NRA

[2566] M. Jamil. Nuclear Reactions for Astrophysics: Principles, Calculation and Applications of Low-Energy Reactions, by Ian J. Thompson and Filomena M. Nunes, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 52(3):259, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2011:FGW

[2567] Peter R. Hobson. Fundamentals of Guided-Wave Optoelectronic Devices, by William S. C. Chang, Scope: reference. Level: graduate level students studying optoelectronic waveguide devices. Contemporary Physics, 52(3):

259–260, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2011:ACA

[2568] K. Alan Shore. Algorithmic Cryptanalysis, by Antoine Joux, Scope: handbook/textbook. Level: undergraduate, graduate. Contemporary Physics, 52(3):260–261, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2011:ATC

[2569] Anonymous. Acceleration technologies for charged particles: an introduction. Contemporary Physics, 52(3):263, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Padgett:2011:KTT

[2570] Miles J. Padgett, Kevin O’Holleran, Robert P. King, and Mark R. Dennis. Knotted and tangled threads of darkness in light beams. Contemporary Physics, 52(4):265–279, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

New:2011:NOF

[2571] G. H. C. New. Nonlinear optics: the first 50 years. Contemporary Physics, 52(4):281–292, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Falcone:2011:NDX

[2572] Roger Falcone, Chris Jacobsen, Janos Kirz, Stefano Marchesini, David Shapiro, and John Spence. New directions in X-ray microscopy. Contemporary Physics, 52(4):293–318, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berman:2011:INP

[2573] Oleg L. Berman, Roman Ya. Kezerashvili, and German V. Kolmakov. Introduction to nonlinear phenomena in superfluid liquids and Bose– Einstein condensates: helium, semiconductors and graphene. Contemporary Physics, 52(4):319–340, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:VD

[2574] D. G. C. Jones. Visions of discovery. Contemporary Physics, 52(4): 341–343, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2011:OP

[2575] F. W. Taylor. Origin of planets. Contemporary Physics, 52(4):345– 347, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2011:DMU

[2576] Eric Sheldon. Dark mysteries of the universe. Contemporary Physics, 52 (4):349–354, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holland:2011:QH

[2577] Peter Holland. A quantum of history. Contemporary Physics, 52(4): 355–358, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:PP

[2578] Peter J. Bussey. Physics and poetry. Contemporary Physics, 52(4):359– 362, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sivakumar:2011:QMC

[2579] S. Sivakumar. Quantum Mechanics: Concepts and Applications, 2nd edn., by Nouredine Zettili, Scope: textbook. Level: undergraduates and advanced students. Contemporary Physics, 52(4):363, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:BCG

[2580] J. Rogel-Salazar. Bose-Condensed Gases at Finite Temperatures, by Allan Griffin, Tetsuro Nikuni and Eugene Zaremba, Scope: review. Level: graduate and researcher. Contemporary Physics, 52(4):364, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:EEA

[2581] J. Rogel-Salazar. Everywhere and Everywhen: Adventures in Physics and Philosophy, by Nick Huggett, Scope: reference. Level: popular science. Contemporary Physics, 52(4):364–365, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2011:GMP

[2582] R. S. Shorter. Gold Medal Physics: The Science of Sports, by John Eric Goff, Scope: review. Level: general readership, non-specialists, teachers.

Contemporary Physics, 52(4):365, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bengtsson:2011:FMM

[2583] Ingemar Bengtsson. Feynman Motives, by Matilde Marcolli, Scope: monograph. Level: researcher. Contemporary Physics, 52(4):366, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2011:PAR

[2584] Peter R. T. Munro. Polarisation: Applications in Remote Sensing, by Shane Cloude, Scope: reference. Level: graduate and practising scientist/engineer. Contemporary Physics, 52(4):366–367, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Briggs:2011:CQA

[2585] Andrew Briggs. Crafting the Quantum: Arnold Sommerfeld and the Practice of Theory, 1890–1926, by Suman Seth, Scope: general interest. Level: general readership. Contemporary Physics, 52(4):367, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zavrtanik:2011:CVE

[2586] Danilo Zavrtanik. CP Violation, 2nd edn., by I. I. Bigi and A. I. Sanda, Scope: monograph. Level: early career researcher. Contemporary Physics, 52(4):368, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ogilvie:2011:ISF

[2587] M. Ogilvie. Introduction to Statistical Field Theory, by E. Brézin, Scope: reference. Level: undergraduate. Contemporary Physics, 52(4):368–369, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2011:PSJ

[2588] R. S. Shorter. Physics of Sailing, by John Kimball, Scope: review. Level: general readership. Contemporary Physics, 52(4):369, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2011:LSA

[2589] R. S. Shorter. Lighter Side of Adaptive Optics, by Robert K. Tyson, Scope: review. Level: general readership. Contemporary Physics, 52 (4):370, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:HSE

[2590] Manuel Vogel. Handbook of Space Engineering, Archaeology, and Heritage, edited by A. G. Darrin and B. L. O’Leary, Scope: handbook. Level: experts. Contemporary Physics, 52(4):370–371, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:OCS

[2591] Manuel Vogel. Observational Cosmology, by S. Serjeant, Scope: monograph. Level: advanced undergraduate students. Contemporary Physics, 52(4):371, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2011:NRA

[2592] K. Alan Shore. Nonlinear Resonance Analysis: Theory, Computation, Applications, by Elena Kartashova, Scope: research monograph. Level: postgraduate and beyond. Contemporary Physics, 52(4):371–372, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2011:TAJ

[2593] Mario Bertolotti. Turbulence in the Atmosphere, by John C. Wyngaard, Scope: monograph. Level: postgraduate, early career researcher, researcher, teacher, specialist. Contemporary Physics, 52(4):372–373, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2011:APA

[2594] Peter W. Milonni. Astrophysics for Physicists, by Arnab Rai Choudhuri, Scope: monograph. Level: advanced undergraduate, beginning postgraduate. Contemporary Physics, 52(4):373–374, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:2011:IPP

[2595] John J. Quenby. An Introduction to Particle Physics and the Standard Model, by Robert Mann, Scope: review. Level: undergraduate, postgraduate. Contemporary Physics, 52(4):374–377, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benhamou:2011:PGGb

[2596] Mabrouk Benhamou. P. G. de Gennes’ Impact on Science: Volume II, edited by Julien Bok, Jacques Prost and Françoise Brochard-Wyart, Scope: review. Level: researchers. Contemporary Physics, 52(4):377– 378, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2011:HER

[2597] M. Jamil. High Energy Radiation from Black Holes: Gamma Rays, Cosmic Rays, and Neutrinos, by Charles D. Dermer and Govind Menon, Scope: monograph. Level: graduate students and advanced researchers. Contemporary Physics, 52(4):378–379, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2011:OTM

[2598] Giovanni Volpe. Optical Tweezers: Methods and Applications, edited by Miles J. Padgett, Justin Molloy and David McGloin, Scope: edited book. Level: specialist. Contemporary Physics, 52(4):379–380, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Verma:2011:FTP

[2599] Murli Manohar Verma. The Future of Theoretical Physics and Cosmology: Celebrating Stephen Hawking’s Contributions to Physics, edited by G. W. Gibbons, E. P. S. Shellard and S. J. Rankin, Scope: monograph. Level: general readership. Contemporary Physics, 52(4):380–382, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:JET

[2600] Peter J. Bussey. Judging Edward Teller, by Istvan Hargittai, Scope: general interest. Level: general readership. Contemporary Physics, 52 (4):382–383, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holland:2011:SMQ

[2601] Peter Holland. The Stability of Matter in Quantum Mechanics, by E. H. Lieb and R. Seiringer, Scope: monograph. Level: postgraduate. Contemporary Physics, 52(4):383–384, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brookes:2011:OPH

[2602] Jennifer C. Brookes. Olfaction: the physics of how smell works? Contemporary Physics, 52(5):385–402, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Haghighipour:2011:SEN

[2603] Nader Haghighipour. Super-Earths: a new class of planetary bodies. Contemporary Physics, 52(5):403–438, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garcia-Ojalvo:2011:PAD

[2604] Jordi Garcia-Ojalvo. Physical approaches to the dynamics of genetic circuits: a tutorial. Contemporary Physics, 52(5):439–464, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jou:2011:MTE

[2605] D. Jou and L. Restuccia. Mesoscopic transport equations and contemporary thermodynamics: an introduction. Contemporary Physics, 52(5): 465–474, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2011:DBD

[2606] David Bailin. Double beta decay and all that. Contemporary Physics, 52(5):475–478, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2011:HPA

[2607] Amgad Salama. Hydroclimatology: Perspectives and Applications, by Marlyn L. Shelton, Scope: reference. Level: undergraduate. Contemporary Physics, 52(5):479–480, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2011:EFP

[2608] Eric Sheldon. Exploring Fundamental Particles, by Lincoln Wolfenstein and João P. Silva, Scope: monograph, general interest. Level: undergraduate, advanced undergraduate, scientist. Contemporary Physics, 52(5): 480–481, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2011:QCP

[2609] M. Jamil. Quantum Chromodynamics: Perturbative and Nonperturbative Aspects, by B. L. Ioffe, V. S. Fadin and L. N. Lipatov, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 52(5): 481–482, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2011:QAG

[2610] Ken Zetie. A Question and Answer Guide to Astronomy, by Pierre-Yves Bely, Scope: general interest. Level: general readership, non-specialists. Contemporary Physics, 52(5):482, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:2011:RGC

[2611] Lewis Ryder. Relativity, Gravitation and Cosmology, by Robert J. A. Lambourne, Scope: review. Level: undergraduate. Contemporary Physics, 52(5):482–483, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kalmus:2011:NFC

[2612] Peter Kalmus. Neutrino, by Frank Close, Scope: review. Level: general readership, undergraduate, postgraduate, teacher, scientist. Contemporary Physics, 52(5):483–484, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:CHM

[2613] Manuel Vogel. Cosmochemistry, by Harry Y. McSween, Jr. and Gary R. Huss, Scope: monograph. Level: advanced undergraduate students. Contemporary Physics, 52(5):484, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2011:SON

[2614] Michael Belsley. Second-Order Nonlinear Optical Characterization Techniques: An Introduction, by Thierry Verbiest, Koen Clays and Vincent Rodriguez, Scope: review. Level: postgraduate and researcher. Contemporary Physics, 52(5):484–485, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:AMP

[2615] Peter J. Bussey. Advanced Modern Physics: Theoretical Foundations, by John Dirk Walecka, Scope: manual. Level: postgraduate. Contemporary Physics, 52(5):485–486, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2011:ECT

[2616] Peter V. E. McClintock. Experimental and Computational Techniques in Soft Condensed Matter Physics, edited by Jeffrey Olafsen, Scope: monograph. Level: researchers and postgraduate physicists. Contemporary Physics, 52(5):486, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dragoman:2011:GFF

[2617] Mircea Dragoman. Gravitation: Foundations and Frontiers, by T. Padmanabhan, Scope: review. Level: early career researcher. Contemporary Physics, 52(5):486–487, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2011:IPA

[2618] F. W. Taylor. An Introduction to Planetary Atmospheres, by A. Sanchez-Lavega, Scope: monograph. Level: undergraduate. Contemporary Physics, 52(5):487, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2011:KVS

[2619] Stig Stenholm. A Kinetic View of Statistical Physics, by Pavel L. Krapivsky, Sidney Redner and Eli Ben-Naim, Scope: monograph. Level: advanced undergraduate, postgraduate, researcher. Contemporary Physics, 52(5):488, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2011:SEN

[2620] Eric Sheldon. Stellar Evolution and Nucleosynthesis, by Sean G. Ryan and Andrew J. Norton, Scope: review. Level: undergraduate, advanced undergraduate, scientist. Contemporary Physics, 52(5):489, 2011. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2011:ITS

[2621] R. Ellis. An Introduction to the Theory of Stellar Structure and Evolution, 2nd edn., by Dina Prialnik, Scope: monograph. Level: early career researcher, postgraduate, teacher. Contemporary Physics, 52(5): 490, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ogilvie:2011:NPF

[2622] M. Ogilvie. Non-Perturbative Field Theory — From Two Dimensional Conformal Field Theory to QCD in Four Dimensions, by Yitzhak Frishman and Jacob Sonnenschein, Scope: review. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 52(5):490– 491, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2011:BLD

[2623] David Matravers. Bayesian Logical Data Analysis for the Physical Sciences, edited by P. C. Gregory, Scope: text book, handbook. Level: postgraduate. Contemporary Physics, 52(5):491–492, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2011:NMO

[2624] Andrew Resnick. Nanoscopy and Multidimensional Optical Fluorescence Microscopy, edited by Alberto Diaspro, Scope: edited book, guide, review. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 52(5):492, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2011:STP

[2625] A. H. Harker. Statistical and Thermal Physics with Computer Applications, by Harvey Gould and Jan Tobochnik, Scope: monograph. Level: undergraduate. Contemporary Physics, 52(5):492–493, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2011:SPP

[2626] R. Ellis. Stellar Populations: Planning for the Next Decade, edited by G. Bruzual and S. Charlot. Contemporary Physics, 52(5):493, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2011:HPO

[2627] K. Alan Shore. Handbook of Plastic Optics, edited by Stefan Baumer, Weinheim. Contemporary Physics, 52(5):494, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2011:CDP

[2628] K. Alan Shore. Current-Driven Phenomena in Nanoelectronics, edited by Tamar Seideman. Contemporary Physics, 52(5):494–495, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2011:TPA

[2629] K. Alan Shore. Twisted Photons: Application of Light with Orbital Angular Momentum, edited by Juan P. Torres and Lluis Torner. Contemporary Physics, 52(5):495, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sumner:2011:PDM

[2630] Timothy J. Sumner. Particle Dark Matter, edited by Gianfranco Bertone. Contemporary Physics, 52(5):496, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2011:GTS

[2631] Allan I. Solomon. Group Theory for the Standard Model of Particle Physics and Beyond, by Ken J. Barnes. Contemporary Physics, 52(5):

496–497, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:2011:NHM

[2632] Ian Thompson. NIST Handbook of Mathematical Functions, edited by Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, Charles W. Clark. Contemporary Physics, 52(5):497–498, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Redfern:2011:EST

[2633] Simon Redfern. Elementary Scattering Theory For X-ray and Neutron Users, by D. S. Sivia. Contemporary Physics, 52(5):498, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:FPM

[2634] J. Rogel-Salazar. Finitary Probabilistic Methods in Econophysics, by Ubaldo Garibaldi and Enrico Scalas. Contemporary Physics, 52(5):499, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Laberge:2011:IGR

[2635] Yves Laberge. An Illustrated Guide to Relativity, by Tatsu Takeuchi. Contemporary Physics, 52(5):499–500, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chia:2011:IOQ

[2636] Andy Chia. Introduction to Optical Quantum Information Processing, by Pieter Kok and Brendon W. Lovett. Contemporary Physics, 52(5): 500–501, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2011:IBI

[2637] Giovanni Volpe. The Inverted Bowl: Introductory Accounts of the Universe and its Life, by George H. A. Cole. Contemporary Physics, 52 (5):501, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2011:PAO

[2638] R. S. Shorter. Principles of Adaptive Optics, 3rd edn., by Robert K. Tyson. Contemporary Physics, 52(5):501–502, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2011:PEB

[2639] Salvatore Esposito. Physics Over Easy: Breakfasts with Beth and Physics, 2nd edn., by Leonid V. Azároff. Contemporary Physics, 52(5): 502–503, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hughes:2011:MIT

[2640] Marcus D. Hughes, Bjoern Lekitsch, Jiddu A. Broersma, and Winfried K. Hensinger. Microfabricated ion traps. Contemporary Physics, 52(6): 505–529, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ozeri:2011:TIQ

[2641] Roee Ozeri. The trapped-ion qubit tool box. Contemporary Physics, 52 (6):531–550, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Marangos:2011:INS

[2642] J. P. Marangos. Introduction to the new science with X-ray free electron lasers. Contemporary Physics, 52(6):551–569, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Slunyaev:2011:RW

[2643] Alexey Slunyaev, Ira Didenkulova, and Efim Pelinovsky. Rogue waters. Contemporary Physics, 52(6):571–590, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2011:ERR

[2644] Peter V. E. McClintock. Essay review: Revolutions that made the Earth, by T. Lenton and A. Watson. ISBN 978-0-19-958704-9, Scope: monograph. Level: general reader with a science education. Contemporary Physics, 52(6):591–593, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sudbery:2011:ASQ

[2645] Tony Sudbery. The Amazing Story of Quantum Mechanics, Scope: general interest. Level: general readership. Contemporary Physics, 52(6): 595, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:FIP

[2646] Manuel Vogel. Fusion: An Introduction to the Physics and Technology of Magnetic Confinement Fusion, 2nd edn., by W. M. Stacey, Scope: textbook. Level: advanced undergraduate students. Contemporary Physics, 52(6):595–596, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:LEA

[2647] Nikolaos E. Myridis. At the Leading Edge: The ATLAS and CMS LHC Experiments, edited by Dan Green, Scope: reference. Level: physicists, graduates in particle physics, engineers. Contemporary Physics, 52(6): 596–597, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bengtsson:2011:CMC

[2648] Ingemar Bengtsson. Classical Measurements in Curved Space–Times, by Fernando de Felice, Scope: monograph. Level: early career researcher, researcher. Contemporary Physics, 52(6):597, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:NRS

[2649] Nikolaos E. Myridis. Numerical Relativity: Solving Einstein’s Equations on the Computer, by Th.W. Baumgarte and S. L. Shapiro, Scope: textbook. Level: advanced undergraduates, researchers, teachers, specialists, scientists. Contemporary Physics, 52(6):598–599, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2011:MFV

[2650] Ken Zetie. Michael Faraday: A Very Short Introduction, by Frank A. J. L. James, Scope: biography, general interest. Level: general readership, non-specialists, teacher. Contemporary Physics, 52(6):599, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2011:CBH

[2651] K. Pounds. Coevolution of Black Holes and Galaxies, edited by Luis C. Ho, Scope: conference proceedings. Level: postgraduate, researcher, specialist. Contemporary Physics, 52(6):599–600, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Phelps:2011:CDP

[2652] Alan Phelps. Complex and Dusty Plasmas: From Laboratory to Space, edited by Vladimir E. Fortov and Gregor E. Morfill, Scope: edited monograph. Level: postgraduate, researcher, specialist. Contemporary Physics, 52(6):600–601, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:NSH

[2653] Nikolaos E. Myridis. Numerical Solution of Hyperbolic Partial Differential Equations, by John A. Trangenstein, Scope: review. Level: textbook for engineering students; engineering researchers, computer scientists. Contemporary Physics, 52(6):601–602, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tennyson:2011:EAP

[2654] Jonathan Tennyson. Exoplanet Atmospheres: Physical Processes, by Sara Seager, Scope: reference. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 52(6):602–603, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gerry:2011:IQP

[2655] Christopher Gerry. Introduction to Quantum Physics and Relativity, by Jacob Dunningham and Vlatko Vedral, Scope: textbook. Level: undergraduate. Contemporary Physics, 52(6):603–604, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:QCQ

[2656] Manuel Vogel. Quantum Computation and Quantum Information, by M. A. Nielsen and I. L. Chuang, Scope: textbook. Level: advanced undergraduate and above. Contemporary Physics, 52(6):604–605, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:ECS

[2657] J. Rogel-Salazar. Econophysics and Companies: Statistical Life and Death in Complex Business Networks, by Hideaki Aoyama, Yoshi Fijiwara, Yuichi Ikeda, Hirishi Iyetomi and Wataru Souma, Scope: review. Level: advanced undergraduate and postgraduate. Contemporary Physics, 52(6):605–606, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:HSG

[2658] P. J. Bussey. Higher Speculations: Grand Theories and Failed Revolutions in Physics and Cosmology, by Helge Kragh, Scope: general interest. Level: general reader. Contemporary Physics, 52(6):606, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2011:MWT

[2659] Mario Bertolotti. Methods of Wave Theory in Dispersive Media, by M. V. Kuzelev and A. A. Rukhadze, Scope: monograph. Level: researchers and experts, students and post-graduate. Contemporary Physics, 52(6): 607, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2011:ODD

[2660] Peter Hobson. Optoelectronic Devices: Design, Modeling, and Simulation, by Xun Li, Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist, scientist and engineers. Contemporary Physics, 52(6):607–608, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2011:ESC

[2661] Stephen H. Ashworth. Energy, the Subtle Concept, by Jennifer Coopersmith, Scope: monograph. Level: general reader with some scientific background. Contemporary Physics, 52(6):608–609, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:PPE

[2662] P. J. Bussey. Particle Physics Experiments at High Energy Colliders, by John Hauptman, Scope: review. Level: postgraduate. Contemporary Physics, 52(6):609–610, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2011:SCD

[2663] Manuel Vogel. Storm and Cloud Dynamics, 2nd edn., by W. R. Cotton, G. H. Bryan and S. C. van den Heever, Scope: monograph. Level: advanced students and experts. Contemporary Physics, 52(6):610–611, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pierri-Galvao:2011:DMP

[2664] Monica Pierri-Galvao. The Dark Matter Problem, by Robert H. Sanders, Scope: review. Level: advanced undergraduate. Contemporary Physics, 52(6):611, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Virmani:2011:IQS

[2665] S. Virmani. Introduction to Quantum Statistical Mechanics, 2nd edn., by N. N. Bogolubov and N. N. Bogolubov, Jr., Scope: monograph. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 52(6):611–612, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Virmani:2011:QFT

[2666] S. Virmani. Quantum Field Theory in a Nutshell, 2nd edn., by A. Zee, Princeton, Scope: edited book. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 52(6):612, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnes:2011:OPS

[2667] Bill Barnes. Optical Properties of Solids, 2nd edn, by Mark Fox, Scope: monograph. Level: undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 52(6):612–613, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:SSF

[2668] Derry W. Jones. Soap, Science and Flat-Screen TVs: A History of Liquid Crystals, by David Dunmur and Tim Sluckin, Scope: history of science with biographies. Level: general reader. Contemporary Physics, 52(6): 613–614, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vedral:2011:QPF

[2669] Vlatko Vedral. Quantum Physics: A Fundamental Approach to Modern Physics, by John S. Townsend, Scope: reference. Level: undergraduate. Contemporary Physics, 52(6):614–615, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kibble:2011:RSR

[2670] B. P. Kibble. Reliability in Scientific Research: Improving the Dependability of Measurements, Calculations, Equipment, and Software, by I. R. Walker, Scope: review. Level: postgraduate, early career researcher, researcher, scientist, engineers. Contemporary Physics, 52(6):615–616, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2011:HPB

[2671] Peter R. T. Munro. Handbook of Photonics for Biomedical Science, edited by Valery V. Tuchin, Scope: handbook. Level: graduate and practising scientist/engineer. Contemporary Physics, 52(6):616–617, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Counihan:2011:MCI

[2672] Martin Counihan. Meteorite Craters and Impact Structures of the Earth, by Paul Hodge, Scope: reference. Level: general scientific readership. Contemporary Physics, 52(6):617, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2011:HBB

[2673] Peter J. Dobson. Handbook of Biosensors and Biosensor Kinetics, by A. Sadana and N. Sadana, Scope: monograph. Level: general readership. Contemporary Physics, 52(6):617–618, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2011:NPN

[2674] M. Jamil. Neutron Physics for Nuclear Reactors: Unpublished Writings by Enrico Fermi, edited by S. Esposito and O. Pisanti, Scope: monograph. Level: nuclear physicists, nuclear engineers and researchers. Contemporary Physics, 52(6):618–619, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:SSS

[2675] Derry W. Jones. Super Structures: The Science of Bridges, Buildings, Dams and Other Feats of Engineering, by Mark Denny, Scope: review. Level: undergraduate or general reader. Contemporary Physics, 52(6): 619, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:SME

[2676] J. Rogel-Salazar. Statistical Mechanics, 3rd edn., by R. K. Pathria and P. D. Beale, Scope: textbook. Level: postgraduate or advanced undergraduate. Contemporary Physics, 52(6):619–620, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2011:HQM

[2677] J. Rogel-Salazar. Heisenberg’s Quantum Mechanics, by Mohsen Razavy, Scope: advanced textbook. Level: advanced undergraduate. Contemporary Physics, 52(6):620–621, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2011:PLS

[2678] D. G. C. Jones. Principles of Laser Spectroscopy and Quantum Optics, by Paul R. Berman and Vladimir S. Malinovsky, Scope: review. Level: postgraduate. Contemporary Physics, 52(6):621–622, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2011:BIC

[2679] P. J. Bussey. The Beautiful Invisible: Creativity, Imagination and Theoretical Physics, by Giovanni Vignale, Scope: general interest. Level: general readership. Contemporary Physics, 52(6):622–623, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Michette:2011:IXP

[2680] Alan Michette. Introduction to XAFS. A Practical Guide to X-ray Absorption Fine Structure Spectroscopy, by Grant Bunker, Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 52(6):623, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2011:MAQ

[2681] Thomas Peters. Mathematical Aspects of Quantum Field Theory, by E. de Faria and W. de Melo, Scope: monograph. Level: graduate. Contemporary Physics, 52(6):623–624, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McFarland:2011:ATC

[2682] John McFarland. Astrophysics and Twentieth-Century Astronomy to 1950: Part A, edited by O. Gingerich, Scope: edited book. Level: pre-university, undergraduate, postgraduate, researcher. Contemporary Physics, 52(6):624–625, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2011:ENW

[2683] S. C. Tiwari. Emmy Noether’s Wonderful Theorem, by D. E. Neuenschwander, Scope: monograph. Level: senior undergraduate/graduate physics students. Contemporary Physics, 52(6):625–626, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gursoy:2011:SII

[2684] Doğa Gürsoy. The Science of Imaging: An Introduction, 2nd edn., by Graham Saxby, Scope: general interest. Level: general readership, nonspecialists, undergraduates, teachers. Contemporary Physics, 52(6):626, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2011:PCO

[2685] Nikolaos E. Myridis. Paradigms of Combinatorial Optimization: Problems and New Approaches, edited by Vangelis T. Paschos, Scope: edited book. Level: postgraduate, researcher, teacher, scientist. Contemporary Physics, 52(6):626–627, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2011:IQO

[2686] David L. Andrews. Introduction to Quantum Optics: From the Semi-Classical Approach to Quantized Light, by Gilbert Grynberg, Alain Aspect and Claude Fabre, Scope: monograph. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 52(6):627–628, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gotte:2011:QMQ

[2687] Jörg Götte. Quantum Mechanics and Quantum Field Theory: A Mathematical Primer, by Jonathan Dimock, Scope: reference. Level: postgraduate. Contemporary Physics, 52(6):628–629, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2011:WGR

[2688] D. J. Miller. What are Gamma-Ray Bursts?, by Joshua S. Bloom, Scope: review. Level: non-specialists. Contemporary Physics, 52(6):629–630, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2011:TSM

[2689] Thomas Peters. Turbulence and Shell Models, by P. D. Ditlevsen, Scope: textbook. Level: undergraduate. Contemporary Physics, 52(6):630–631, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2011:HOU

[2690] Robert D. Joseph. How Old is the Universe?, by David A. Weintraub, Scope: review. Level: undergraduate. Contemporary Physics, 52(6):631– 632, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:2011:QPP

[2691] Alan Martin. Quantum Physics for Poets, by Leon M. Lederman and Christopher T. Hill, Scope: general interest. Level: general readership.

Contemporary Physics, 52(6):632, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2011:CQI

[2692] Gerard Milburn. Classical and Quantum Information, by D. C. Marinescu and G. M. Marinescu, Scope: monograph. Level: postgraduate. Contemporary Physics, 52(6):632–633, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hendry:2011:LGA

[2693] Martin Hendry. Lectures on Gravitation, by Ashok Das, Scope: review. Level: postgraduate, early career researcher, researcher, teacher. Contemporary Physics, 52(6):633–634, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palffy:2011:QSH

[2694] Adriana Pálffy. The Quantum Story: A History in 40 Moments, by Jim Baggott, Scope: general interest. Level: general readership. Contemporary Physics, 52(6):634–635, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2011:ADS

[2695] Miguel A. F. Sanjuán. Adaptation in Dynamical Systems, by Ivan Tyukin, Scope: monograph, reference. Level: general reader, undergraduate, postgraduate, specialist. Contemporary Physics, 52(6):635, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2011:CMI

[2696] Geoffrey Duxbury. Concepts and Methods of 2D Infrared Spectroscopy, by Peter Hamm and Martin Zanni, Scope: monograph. Level: postgraduate. Contemporary Physics, 52(6):636, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2011:EB

[2697] Anonymous. Editorial board. Contemporary Physics, 52(6):ebi, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ralph:2012:RQI

[2698] Timothy C. Ralph and Tony G. Downes. Relativistic quantum information and time machines. Contemporary Physics, 53(1):1–16, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Balenzuela:2012:GSR

[2699] Pablo Balenzuela, Holger Braun, and Dante R. Chialvo. The ghost of stochastic resonance: an introductory review. Contemporary Physics, 53 (1):17–38, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Raine:2012:PBL

[2700] Derek Raine and Sarah Symons. Problem-based learning: undergraduate physics by research. Contemporary Physics, 53(1):39–51, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2012:ERN

[2701] Eric Sheldon. Essay review: Neutrino Physics, 2nd edn., by K. Zuber. ISN 978-1-4200-6471-1, Scope: textbook, review level: postgraduate, researcher, specialist. Contemporary Physics, 53(1):53–54, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harris:2012:WRE

[2702] David Harris. Waves and Rays in Elastic Continua, by Michael A. Slawinski, Scope: monograph. Level: postgraduate, scientist, engineers. Contemporary Physics, 53(1):55, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2012:OIS

[2703] S. W. Lovesey. Optical Interactions in Solids, 2nd edn, Scope: monograph. Level: researcher. Contemporary Physics, 53(1):56, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

delCampo:2012:FNF

[2704] Adolfo del Campo. Fundamentals and New Frontiers of Bose–Einstein Condensation, Scope: monograph. Level: advanced graduate and postgraduate students and researchers. Contemporary Physics, 53(1):56–57, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2012:CMI

[2705] Amgad Salama. Classical Mechanics Illustrated by Modern Physics: 42 Problems with Solutions, Scope: reference. Level: undergraduate. Contemporary Physics, 53(1):57, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2012:MAN

[2706] M. Jamil. Modern Atomic and Nuclear Physics (Revised Edition): Problems and Solutions Manual, Scope: manual. Level: graduate students and researchers. Contemporary Physics, 53(1):58, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:LUS

[2707] P. J. Bussey. Lawless Universe: Science and the Hunt for Reality, Scope: general interest. Level: general reader. Contemporary Physics, 53(1): 58–59, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowan:2012:THJ

[2708] B. Cowan. Thermodynamics, by H. J. Kreuzer and Isaac Tamblyn, Scope: manual. Level: undergraduate. Contemporary Physics, 53(1): 59–60, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:FEE

[2709] Barrie W. Jones. Formation and Evolution of Exoplanets, edited by Rory Barnes, Scope: monograph. Level: scientist. Contemporary Physics, 53 (1):60–61, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:PHB

[2710] Manuel Vogel. Physics of the Human Body, by Richard P. McCall, Scope: monograph/essay. Level: undergraduate students. Contemporary Physics, 53(1):69, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2012:NWM

[2711] Miguel A. F. Sanjuán. Nonlinear Wave Methods for Charge Transport, by Luis L. Bonilla and Stephen W. Teitsworth, Scope: monograph. Level: graduate students and specialists. Contemporary Physics, 53(1):61–62, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2012:TPS

[2712] Matt Probert. Thermodynamic Properties of Solids: Experiment and Modeling, edited by Samrath L. Chaplot, Ranjan Mittal and Narayani Choudhury, Scope: review. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 53(1):62–63, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:AMG

[2713] Thomas Peters. Advanced Mechanics and General Relativity, by J. Franklin, Scope: textbook. Level: graduate. Contemporary Physics, 53 (1):63–64, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2012:DUM

[2714] R. S. MacKay. The Dark Universe: Matter, Energy, and Gravity, edited by Mario Livio, Scope: review. Level: general readership. Contemporary Physics, 53(1):64–65, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2012:ITC

[2715] S. J. Blundell. Introduction to the Theory of Critical Phenomena: Mean Field, Fluctuations and Renormalization, 2nd edn., by Dimo I. Uzunov, Scope: research monograph. Level: early career researcher, researcher. Contemporary Physics, 53(1):65–66, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:2012:CGA

[2716] Ian Lawrie. Canonical Gravity and Applications: Cosmology, Black Holes, and Quantum Gravity, by M. Bojowald, Scope: monograph. Level: postgraduate, early career researcher. Contemporary Physics, 53(1):66– 67, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Senekerimyan:2012:OSB

[2717] Vahan Senekerimyan. Optical Sensors: Basics and Applications, by J. Haus, Scope: general interest, monograph. Level: general readership, non-specialists, undergraduates, teachers. Contemporary Physics, 53(1): 67, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Siklos:2012:EST

[2718] S. T. C. Siklos. Exact Space–Times in Einstein’s General Relativity, by Jerry B. Griffiths and Jiří Podolský, Scope: reference. Level: postgraduate. Contemporary Physics, 53(1):67–69, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:FOP

[2719] K. Alan Shore. Fiber Optics: Physics and Technology, by Fedor Mitschke, Scope: manual. Level: advanced undergraduate and postgraduate. Contemporary Physics, 53(1):69–70, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:NEB

[2720] K. Alan Shore. Nanoimaging, edited by Beth A. Goins and William T. Phillips, Scope: research monograph. Level: postgraduate and beyond. Contemporary Physics, 53(1):70–71, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:HSS

[2721] Manuel Vogel. Heliophysics: Space Storms and Radiation: Causes and Effects, edited by C. J. Schrijver and G. L. Siscoe, Scope: monograph. Level: advanced undergraduate and above. Contemporary Physics, 53 (1):71, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2012:FEP

[2722] Amgad Salama. Foundations of Environmental Physics: Understanding Energy Use and Human Impacts, by Kyle Forinash, Scope: monograph, reference. Level: undergraduate, teacher, scientist, engineers. Contemporary Physics, 53(1):71–72, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sivakumar:2012:IQO

[2723] S. Sivakumar. An Introduction to Quantum Optics: Photon and Biphoton Physics, Scope: monograph. Level: researchers and lecturers. Contemporary Physics, 53(1):72–73, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kourkoumelis:2012:RMS

[2724] Nikolaos Kourkoumelis. Research Methods for Science, by Michael P. Marder, Scope: handbook. Level: undergraduate. Contemporary Physics, 53(1):73, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Seferyan:2012:SGO

[2725] Hrant Ye. Seferyan. Supercontinuum Generation in Optical Fibers, edited by J. M. Dudley and J. R. Taylor, Scope: edited book, handbook, reference, review. Level: postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 53(1):74–75, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Isenberg:2012:CMA

[2726] C. Isenberg. Classical Mechanics with Applications, by Porter Wear Johnson, Scope: general interest. Level: undergraduate and postgraduate. Contemporary Physics, 53(1):75–76, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:RFA

[2727] Peter J. Bussey. Relativity in Fundamental Astronomy: Dynamics, Reference Frames, and Data Analysis, IAU Symposium 261, edited by S. A. Klioner, P. K. Seidelmann and M. H. Soffel, Scope: conference proceedings. Level: researcher. Contemporary Physics, 53(1):76–77, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:QPT

[2728] J. Rogel-Salazar. Quantum Phase Transitions, 2nd edn., by S. Sachdev, Scope: textbook. Level: postgraduate or advanced undergraduate. Contemporary Physics, 53(1):77, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:EPT

[2729] Thomas Peters. Elements of Phase Transitions and Critical Phenomena, by H. Nishimori and G. Ortiz, Scope: textbook. Level: graduate. Contemporary Physics, 53(1):78, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Garcia-Morales:2012:CGL

[2730] Vladimir García-Morales and Katharina Krischer. The complex Ginzburg–Landau equation: an introduction. Contemporary Physics, 53 (2):79–95, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Faccio:2012:LPA

[2731] Daniele Faccio. Laser pulse analogues for gravity and analogue Hawking radiation. Contemporary Physics, 53(2):97–112, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Learned:2012:CGI

[2732] John G. Learned, R.-P. Kudritzki, Sandip Pakvasa, and A. Zee. The Cepheid galactic Internet. Contemporary Physics, 53(2):113–118, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gammaitoni:2012:TPE

[2733] Luca Gammaitoni. There’s plenty of energy at the bottom (micro and nano scale nonlinear noise harvesting). Contemporary Physics, 53(2): 119–135, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Janson:2012:NLD

[2734] Natalia B. Janson. Non-linear dynamics of biological systems. Contemporary Physics, 53(2):137–168, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2012:FHN

[2735] S. Esposito. The fisherman and his net: is Physics discovered or invented?, Scope: monograph. Level: advanced undergraduate, postgraduate, specialist researcher. Contemporary Physics, 53(2):169–172, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:PLE

[2736] Manuel Vogel. Principles of Lasers, 5th edn., by O. Svelto, Scope: monograph. Level: advanced undergraduate and above. Contemporary Physics, 53(2):173, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2012:IBS

[2737] B. Ishak and J. McFarland. Icy Bodies of the Solar System (IAU S263). Proceedings of the IAU Symposia and Colloquia, edited by J. A. Fernández, D. Lazarro, D. Prialnik and R. Schulz. Scope: wide-ranging review. Level: postgraduates and specialists. Contemporary Physics, 53 (2):173–174, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:HEA

[2738] D. J. Miller. High Energy Astrophysics, 3rd edn., by Malcolm S. Longair, Scope: monograph. Level: undergraduate. Contemporary Physics, 53(2): 174–175, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2012:CCO

[2739] Nikolaos E. Myridis. Concepts of Combinatorial Optimization, edited by Vangelis Th. Paschos, Scope: edited book. Level: researchers, practitioners, MSc or Ph.D. students. Contemporary Physics, 53(2):175–176, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2012:IST

[2740] Aniruddha Chakraborty. Introductory Statistical Thermodynamics, by Nils Dalarsson, Mariana Dalarsson and Leonardo Golubovic, Scope: review. Level: undergraduate. Contemporary Physics, 53(2):176, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2012:PCG

[2741] Eric Sheldon. Primordial Cosmology, by Giovanni Montani, Marco Valerio Battisti, Riccardo Benini and Giovanni Imponente, Scope: review, textbook. Level: researcher, specialist, scientist. Contemporary Physics, 53(2):176–177, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:KHD

[2742] Peter J. Bussey. Knocking on Heaven’s Door: How Physics and Scientific Thinking Illuminate The Universe and the Modern World, by Lisa Randall, Scope: general interest. Level: general reader. Contemporary Physics, 53(2):178, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2012:ACO

[2743] Nikolaos E. Myridis. Applications of Combinatorial Optimization, edited by Vangelis Th. Paschos, Scope: edited book. Level: researchers, practitioners, MSc or Ph.D. students. Contemporary Physics, 53(2):178– 179, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:APP

[2744] Thomas Peters. Advanced Particle Physics, Volume I: Particles, Fields, and Quantum Electrodynamics, by O. M. Boyarkin. Advanced Particle Physics, Volume II: The Standard Model and Beyond, by O. M. Boyarkin. Scope: monograph. Level: graduate, Scope: monograph. Level graduate. Contemporary Physics, 53(2):179–180, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2012:STC

[2745] Miguel A. F. Sanjuán. Stochastic Transport in Complex Systems: From Molecules to Vehicles, by Andreas Schadschneider, Scope: review. Level: graduate students and specialists. Contemporary Physics, 53(2):180– 181, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:PFQ

[2746] D. G. C. Jones. Physical Foundations of Quantum Electronics, by David Klyshko, edited by Maria Chekhova and Sergey Kulik, Scope: edited book. Level: postgraduate. Contemporary Physics, 53(2):181–182, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dragoman:2012:PUM

[2747] Mircea Dragoman. Physics of the Universe, by Mendel Sachs, Scope: review. Level: scientist. Contemporary Physics, 53(2):182, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:HDF

[2748] D. J. Miller. How did the First Stars and Galaxies Form?, by Abraham Loeb, Scope: review. Level: non-specialists. Contemporary Physics, 53 (2):183, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dimakis:2012:CNA

[2749] Nicholas Dimakis. Computational Nanoscience: Applications for Molecules, Clusters, and Solids, by Kálmán Varga and Joseph A. Driscoll, Scope: review. Level: advanced undergraduates and graduate students in computational physics, material science, and engineering. Contemporary Physics, 53(2):183–184, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:RAS

[2750] Andrew Resnick. Recent Advances in Spectroscopy: Theoretical, Astrophysical and Experimental Perspectives, edited by R. K. Chaudhuri, M. V. Mekkaden, A. V. Raveendran, A. Satya Narayanan, Scope: conference proceedings. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 53(2):184–185, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:PTE

[2751] Peter J. Bussey. Physics at the Terascale, edited by I. Brock and T. Schörner-Sadenius, Scope: reference. Level: researcher. Contemporary Physics, 53(2):185, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whitaker:2012:INO

[2752] Benjamin J. Whitaker. Introduction to Nonlinear Optics, by Geoffrey New, Scope: reference. Level: postgraduate. Contemporary Physics, 53 (2):186, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2012:FFT

[2753] Peter R. T. Munro. The Fallacy of Fine-Tuning: Why The Universe is Not Designed for Us, by Victor J. Stenger, Scope: general interest. Level: general readership. Contemporary Physics, 53(2):186–187, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carroll:2012:FCI

[2754] J. E. Carroll. Fractional Calculus: An Introduction For Physicists, by Richard Herrmann, Scope: review. Level: specialist. Contemporary Physics, 53(2):187–188, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2012:YAL

[2755] Salvatore Esposito. 50 Years of Anderson Localization, edited by Elihu Abrahams, Scope: edited book. Level: undergraduate, postgraduate, researcher, specialist. Contemporary Physics, 53(2):188–189, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2012:QGP

[2756] David Bailin. Quark–Gluon Plasma 4, edited by Rudolph C. Hwa and Xin-Nian Wang, Scope: edited book, monograph. Level: specialist. Contemporary Physics, 53(2):189–190, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carter:2012:FCA

[2757] R. G. Carter. Field Computation for Accelerator Magnets: Analytical and Numerical Methods for Electromagnetic Design and Optimization, by Stephan Russenschuck, Scope: monograph, reference. Level: postgraduate, researcher, specialist. Contemporary Physics, 53(2):190–191, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2012:SCL

[2758] Allan Solomon. Symmetries and Conservation Laws in Particle Physics: An Introduction to Group Theory for Particle Physicists, by Stephen Haywood, Scope: monograph. Level: undergraduate, postgraduate, early career researcher. Contemporary Physics, 53(2):191–192, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2012:NHQ

[2759] G. M. Schütz. Non-Hermitian Quantum Mechanics, by Nimrod Moiseyev, Scope: monograph, reference. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 53(2):192, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:NTP

[2760] Andrew Resnick. New Trends in the Physics and Mechanics of Biological Systems: Lecture Notes of the Les Houches Summer School, Volume 92, edited by Martine Ben Amar, Alain Goriely, Martin Michael Müller and Leticia Cugliandolo, Scope: conference proceedings. Level: undergraduate, postgraduate, early career researcher, teacher, specialist. Contemporary Physics, 53(2):193, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rajantie:2012:IMM

[2761] Arttu Rajantie. Introduction to magnetic monopoles. Contemporary Physics, 53(3):195–211, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

House:2012:MEN

[2762] Thomas House. Modelling epidemics on networks. Contemporary Physics, 53(3):213–225, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chevrollier:2012:RTL

[2763] Martine Chevrollier. Radiation trapping and lévy flights in atomic vapours: an introductory review. Contemporary Physics, 53(3):227– 239, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dombey:2012:IPF

[2764] Norman Dombey. The Infinity Puzzle, by F. Close, Scope: reference. Level: general reader. Contemporary Physics, 53(3):241–245, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:ERP

[2765] Derry W. Jones. Essay review: Physics letters: the Stokes–Thomson effect. The Correspondence between Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs: Vols. 1 & 2, edited by D. B.

Wilson, Scope: review. Level: final-year science undergraduate. Contemporary Physics, 53(3):247–250, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Foreman:2012:SSP

[2766] Matthew R. Foreman. Statistical Signal Processing of Complex-Valued Data: The Theory of Improper and Noncircular Signals, by Peter J. Schreier and Louis L. Scharf, Scope: textbook. Level: postgraduate, researcher. Contemporary Physics, 53(3):251–252, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benhamou:2012:PPS

[2767] Mabrouk Benhamou. Polymer Physics: From Suspension to Nanocomposites and Beyond, edited by Leszek A. Utracki and Alexander M. Jamieson, Scope: review. Level: researcher. Contemporary Physics, 53 (3):252–253, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2012:RQM

[2768] Ken Zetie. Relativity for the Questioning Mind, by Daniel F. Styer, Scope: general interest. Level: undergraduate. Contemporary Physics, 53(3):253, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zhu:2012:BEC

[2769] Xuetao Zhu. Basic Elements of Crystallography, by Nevill Gonzalez Szwacki and Teresa Szwacka, Scope: handbook, reference. Level: nonspecialists, undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 53(3):253–254, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:QRB

[2770] Peter J. Bussey. The Quest for Reality: Bohr and Wittgenstein — Two Complementary Views, by S. Stenholm, Scope: general interest. Level: scientist. Contemporary Physics, 53(3):254–255, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kourkoumelis:2012:NPB

[2771] Nikolaos Kourkoumelis. Nanocomposite Particles for Bio-Applications: Materials and Bio-Interfaces, by Tito Trindade and Ana L. Daniel da Silva, Scope: edited book. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 53(3):255–256, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2012:ERG

[2772] David Bailin. Essay review: The geese that never cackled. Secret Days: Codebreaking in Bletchley Park, by Asa Briggs. ISBN 978-1-84832-615-6, Scope: review. Level: general readership. Contemporary Physics, 53(3): 256–262, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2012:SNI

[2773] Peter W. Milonni. Science at the Nanoscale: An Introductory Textbook, by C. W. Shong, S. C. Haur and A. T. S. Wee, Scope: monograph. Level: undergraduate. Contemporary Physics, 53(3):262, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2012:PPO

[2774] S. J. Blundell. Properties of Perovskites and Other Oxides, by K. Alex Müller and Tom W. Kool, Scope: monograph. Level: researcher. Contemporary Physics, 53(3):263, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2012:AIP

[2775] Geoffrey Duxbury. Advances in Infrared Photodetectors, edited by Sarath D. Gunapala, David R. Rhiger and Chennupatti Jagadish, Scope: monograph. Level: early career researcher. Contemporary Physics, 53(3):263– 264, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2012:WRC

[2776] A. H. Harker. Wavefronts and Rays as Characteristics and Asymptotics, by A. Bóna and M. A. Slawinski, Scope: monograph. Level: researcher. Contemporary Physics, 53(3):265, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2012:NCP

[2777] A. H. Harker. Nanocomputing: Computational Physics for Nanoscience and Nanotechnology, by James J. Y. Hsu, Scope: monograph. Level: undergraduate, postgraduate. Contemporary Physics, 53(3):265–266, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2012:EUA

[2778] K. A. Pounds. The Expanding Universe: Astronomy’s ‘Great Debate’, 1900–1931, by Robert W. Smith, Scope: general interest, monograph. Level: general readership, non-specialists. Contemporary Physics, 53(3): 266–267, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2012:SAE

[2779] Geoffrey Duxbury. 3D Spectroscopy in Astronomy, edited by E. Mediavilla, S. Arribas, M. Roth, J. Cepa-Nogué and F. Cepa-Nogué, Scope: guide. Level: researcher. Contemporary Physics, 53(3):267–268, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reipurth:2012:SCB

[2780] Bo Reipurth. Star Clusters: Basic Galactic Building Blocks Throughout Time and Space, IAU S266, edited by Richard de Grijs and Jacques Lépine, Scope: conference proceedings. Level: postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 53(3):268– 269, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Athorne:2012:SPC

[2781] Chris Athorne. Solved Problems in Classical Mechanics: Analytical and Numerical Solutions with Comments, by O. L. de Lange and J. Pierrus, Scope: review. Level: undergraduate. Contemporary Physics, 53 (3):269, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:SCM

[2782] J. Rogel-Salazar. Stochastic Control and Mathematical Modeling: Applications in Economics, by H. Morimoto, Scope: review. Level: advanced postgraduate or scientist. Contemporary Physics, 53(3):269–270, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:OPA

[2783] D. G. C. Jones. Optically Pumped Atoms, by William Happer, Yuan-Yu Jau and Thad G. Walker, Scope: monograph. Level: undergraduate, postgraduate. Contemporary Physics, 53(3):270–271, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2012:EOI

[2784] Matt Probert. Engineering Optimisation: An Introduction with Metaheuristic Applications, by Xin-She Yang, Scope: reference. Level:undergraduate, postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 53(3):271–272, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2012:SNR

[2785] Peter J. Dobson. Superlattice to Nanoelectronics, by Raphael Tsu, Scope: reference. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 53(3):272– 273, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2012:IST

[2786] David Bailin. An Introduction to String Theory and D-Brane Dynamics: With Problems and Solutions, 2nd edn., by Richard J. Szabo, Scope: monograph. Level: postgraduate. Contemporary Physics, 53(3):273–274, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2012:OPM

[2787] Giovanni Volpe. Optical Processes in Microparticles and Nanostructures, A Festschrift, dedicated to Richard Kounai Chang on his Retirement from Yale University, edited by Ali Serpengüzel and Andrew W. Poon, Scope: review. Level: practitioners. Contemporary Physics, 53(3):274– 275, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2012:ASI

[2788] Stephen H. Ashworth. Astronomical Spectroscopy: An Introduction to the Atomic and Molecular Physics of Astronomical Spectra, 2nd edn., by Jonathan Tennyson. Atomic Astrophysics and Spectroscopy, by Anil K. Pradhan and Sultana N. Nahar. Scope: textbook. Level: undergraduate. Scope: textbook. Level: advanced undergraduate/postgraduate. Contemporary Physics, 53(3):275–276, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sykes:2012:MCA

[2789] J. M. Sykes. The Magnesium Civilisation: An Alternative New Source of Energy to Oil, by Takashi Yabe and Tatsuya Yamaji, Scope: general interest. Level: general readership and non-specialists. Contemporary Physics, 53(3):276–277, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2012:FMS

[2790] Salvatore Esposito. Fluid Mechanics: A Short Course for Physicists, by Gregory Falkovich, Scope: handbook. Level: undergraduate and postgraduate. Contemporary Physics, 53(3):277–278, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Freise:2012:OMF

[2791] A. Freise. Optical Metamaterials: Fundamentals and Applications, by W. Cai and V. Shalaev, Scope: handbook, review. Level: undergraduate, postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 53(3):278–279, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2012:FSP

[2792] A. M. Fox. Fundamentals of Semiconductors: Physics and Materials Properties, 4th edn., by Peter Y. Yu and Manuel Cardona, Scope: manual. Level: postgraduate. Contemporary Physics, 53(3):279–280, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2012:AUS

[2793] Allan I. Solomon. Applications of Unitary Symmetry and Combinatorics, by James D. Louck, Scope: monograph. Level: specialist. Contemporary Physics, 53(3):280, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2012:NMM

[2794] A. H. Harker. Numerical Modeling in Materials Science and Engineering, Springer Series in Computational Mathematics, Vol. 32, by M. Rappaz, M. Bellet, and M. Deville, translated by R. Snyder, Scope: monograph. Level: undergraduate, postgraduate. Contemporary Physics, 53(3):280– 281, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bucknall:2012:FSA

[2795] D. G. Bucknall. Functional Supramolecular Architectures: For Organic Electronics and Nanotechnology, 2 Volume set, edited by P. Samorì and F. Cacialli, Scope: edited book. Level: postgraduate, early career researcher, researcher, specialist. Contemporary Physics, 53(3):281–282, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palffy:2012:QEP

[2796] Adriana Pálffy. Quantum Enigma: Physics Encounters Consciousness, 2nd edn., by Bruce Rosenblum and Fred Kuttner, Scope: general interest. Level: general readership. Contemporary Physics, 53(3):282–283, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2012:QET

[2797] Gerard Milburn. Quantum Engineering: Theory and Design of Quantum Coherent Structures, by A. M. Zagoskin, Scope: monograph. Level: postgraduate. Contemporary Physics, 53(3):283–284, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2012:ANE

[2798] Peter J. Dobson. Advances in Nanotechnology and the Environment, edited by Juyoung Kim, Scope: review. Level: researcher, specialist and scientist. Contemporary Physics, 53(3):284–285, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sykes:2012:ENP

[2799] John Sykes. Electrochemical Nanofabrication: Principles and Applications, edited by Di Wei, Scope: edited book. Level: early career researcher, researcher, specialist and scientist. Contemporary Physics, 53 (3):285–286, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2012:ONV

[2800] Peter R. Hobson. Optics of Nanomaterials, by Vladimir I. Gavrilenko, Scope: monograph. Level: postgraduate, early career researcher, scientist and engineers. Contemporary Physics, 53(3):286–287, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2012:RQP

[2801] Aniruddha Chakraborty. Relativistic Quantum Physics: From Advanced Quantum Mechanics to Introductory Quantum Field Theory, by Tommy Ohlsson, Scope: review. Level: undergraduate. Contemporary Physics, 53(3):287, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:EME

[2802] Manuel Vogel. Electricity and Magnetism, 2nd ed., by Edward Purcell, Scope: monograph. Level: undergraduate. Contemporary Physics, 53(3): 287–288, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2012:LPC

[2803] Nikolaos E. Myridis. Lasers in the Preservation of Cultural Heritage: Principles and Applications, by C. Fotakis, D. Anglos,V. Zafiropulos, S. Georgiou, and V. Tornari, Scope: edited book. Level: researchers, practitioners, postgraduate students. Contemporary Physics, 53(3):288– 289, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:OMT

[2804] K. Alan Shore. Optical Measurement Techniques: Innovations for Industry and the Life Sciences, Springer Series in Optical Sciences, Vol. 136, by Kai-Erik Peiponen, Risto Mullylä and Alexander V. Priezzhev, Scope: monograph. Level: postgraduate and beyond. Contemporary Physics, 53 (3):289, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:SQO

[2805] K. Alan Shore. Semiconductor Quantum Optics, by Mackillo Kira and Stephan W. Koch, Scope: monograph, reference. Level: advanced undergraduate and postgraduate. Contemporary Physics, 53(3):289–290, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2012:MTP

[2806] P. T. Greenland. The Mesoscopic Theory of Polymer Dynamics, 2nd edn., Springer Series in Chemical Physics, Vol. 95, by V. N. Pokrovskii, Scope: monograph. Level: specialist. Contemporary Physics, 53(3):290– 291, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2012:CLB

[2807] F. W. Taylor. Collisional Line Broadening and Shifting of Atmospheric Gases: A Practical Guide for Line Shape Modelling by Current Semi-Classical Approaches, by J. Buldyreva, N. Lavrentiveva and V. Starikov, Scope: reference. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 53(3):291, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2012:PGR

[2808] R. S. MacKay. Probability on Graphs: Random Processes on Graphs and Lattices, by G. Grimmett, Scope: monograph. Level: postgraduate, researcher, specialist. Contemporary Physics, 53(3):291–292, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2012:FCP

[2809] David Bailin. Fast Car Physics, by Chuck Edmondson, Scope: general interest. Level: pre-university, undergraduate. Contemporary Physics, 53(3):292–293, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:2012:HBB

[2810] Frank Close. Higgs boson: beginning of the end or end of the beginning? Contemporary Physics, 53(4):295–300, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

OConnell:2012:RRG

[2811] R. F. O’Connell. Radiation reaction: general approach and applications, especially to electrodynamics. Contemporary Physics, 53(4):301– 313, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

DiBari:2012:ILN

[2812] Pasquale Di Bari. An introduction to leptogenesis and neutrino properties. Contemporary Physics, 53(4):315–338, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gomboc:2012:USG

[2813] Andreja Gomboc. Unveiling the secrets of gamma ray bursts. Contemporary Physics, 53(4):339–355, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2012:OSU

[2814] Eric Sheldon. Our single UNIverse or a plethora of MULTIverses?, Scope: general interest. Level: general readership, scientist. Contemporary Physics, 53(4):357–360, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:2012:ITE

[2815] John Quenby. Introduction to the Theory of the Early Universe: Cosmological Perturbations and Inflation Theory, by Dimitry S. Gorbunov and Valery A. Rubakov, Scope: monograph. Level: postgraduates and early career researchers. Contemporary Physics, 53(4):361–362, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2012:FNM

[2816] M. Jamil. Fundamentals of Nuclear Models: Foundational Models, by David J. Rowe and John L. Wood, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 53(4):363, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2012:IMW

[2817] Chris Linton. An Introduction to Metamaterials and Waves in Composites, by B. Banerjee, Scope: textbook. Level: graduate and advanced undergraduate. Contemporary Physics, 53(4):363–364, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:MPP

[2818] Thomas Peters. Modern Plasma Physics, Volume 1: Physical Kinetics of Turbulent Plasmas, by P. H. Diamond, S.-I. Itoh and K. Itoh, Scope: textbook. Level: graduate. Contemporary Physics, 53(4):364–365, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benedict:2012:UQP

[2819] Keith Benedict. Understanding Quantum Phase Transitions, edited by Lincoln Carr, Scope: edited book. Level: postgraduate, early career researcher and researcher. Contemporary Physics, 53(4):365–366, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:LPG

[2820] D. G. C. Jones. Light Propagation in Gain Media Media: Optical Amplifiers, by Malin Premaratne and Govind P. Agrawal, Scope: textbook. Level: postgraduate and early career researcher. Contemporary Physics, 53(4):366, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:LDS

[2821] Thomas Peters. Lectures on Dynamics of Stochastic Systems, by V. I. Klyatskin, Scope: textbook. Level: graduate. Contemporary Physics, 53 (4):366–367, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2012:GMV

[2822] S. C. Tiwari. Gribov-80 Memorial Volume: Quantum Chromodynamics and Beyond, edited by Yu. L. Dokshitzer, P. Lévai and J. Nyíri, Scope: conference proceedings. Level: specialist. Contemporary Physics, 53(4): 367–368, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cole:2012:ISM

[2823] Milton W. Cole. Introduction to Statistical Mechanics, by John Dirk Walecka, Scope: monograph. Level: postgraduate and teacher. Contemporary Physics, 53(4):368, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:FTN

[2824] Thomas Peters. Field Theory of Non-Equilibrium Systems, by A. Kamenev, Scope: textbook. Level: graduate. Contemporary Physics, 53(4):368–369, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:FSR

[2825] J. Rogel-Salazar. First Steps in Random Walks: From Tools to Applications, by J. Klafter and I. M. Sokolov, Scope: review. Level: graduate students or scientists. Contemporary Physics, 53(4):369–370, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2012:PCT

[2826] Fredric W. Taylor. Planetary Crusts: Their Composition, Origin and Evolution, by Stuart Ross Taylor and Scott M. McLennan, Scope: monograph. Level: non-specialists, pre-university, undergraduate, postgraduate and early career researcher. Contemporary Physics, 53(4):370–371, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:QQQ

[2827] Manuel Vogel. Quips, Quotes and Quanta: An Anecdotal History of Physics, 2nd edn, by Anton Z. Capri, Scope: monograph. Level: general reader. Contemporary Physics, 53(4):371, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2012:TFO

[2828] Mario Bertolotti. Thin-Film Optical Filters, 4th edition, by H. Angus Macleod, Scope: textbook. Level: researchers, specialists, engineers.

Contemporary Physics, 53(4):371–372, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2012:MQM

[2829] Stephen H. Ashworth. Molecular Quantum Mechanics, 5th edn, by Peter Atkins and Ronald Friedman, Scope: textbook. Level: undergraduate. Contemporary Physics, 53(4):372–373, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blamire:2012:HMM

[2830] M. Blamire. Handbook of Magnetic Materials, edited by K. H. J. Buschow, Scope: handbook. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 53(4):373–374, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2012:FPQ

[2831] S. C. Tiwari. Foundations of Perturbative QCD, by John Collins, Scope: textbook/monograph. Level: graduate students/researchers in high energy physics. Contemporary Physics, 53(4):374–375, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:RCM

[2832] Manuel Vogel. Relativistic Celestial Mechanics of the Solar System, by Sergei Kopeikin, Michael Efroimsky and George Kaplan, Scope: monograph. Level: undergraduate, postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 53(4):375–376, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2012:NIP

[2833] Miguel A. F. Sanjuán. Noise-Induced Phenomena in the Environmental Sciences, by Luca Ridolfi, Paolo D’Odorico and Francesco Laio, Scope: monograph. Level: graduate students and specialists, general readers, undergraduates, postgraduates, specialists. Contemporary Physics, 53 (4):376, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:LNA

[2834] Andrew Resnick. Life at the Nanoscale: Atomic Force Microscopy of Live Cells, edited by Yves Dufrêne, Scope: guide, handbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 53(4):377, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:AOE

[2835] Andrew Resnick. Astronomical Optics and Elasticity Theory: Active Optics Methods, by G. R. Lemaitre, Scope: reference. Level: postgraduate, advanced undergraduate, early career researcher, researcher, scientist. Contemporary Physics, 53(4):377–378, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2012:NDA

[2836] Peter J. Dobson. Nanomedicine: Design and Applications of Magnetic Nanomaterials, Nanosensors and Nanosystems, by Vijay K. Varadan, Linfeng Chen and Jining Xie, Scope: monograph, reference. Level: postgraduate, advanced undergraduate, early career researcher, researcher, engineers. Contemporary Physics, 53(4):378–379, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:IEB

[2837] Andrew Resnick. Introduction to Experimental Biophysics: Biological Methods for Physical Scientists, by Jay Nadeau, Scope: handbook, manual, reference. Level: undergraduate, postgraduate, advanced undergraduate, early career researcher, specialist, scientist. Contemporary Physics, 53(4):379, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2012:CNG

[2838] Peter J. Dobson. Carbon Nanotube and Graphene Device Physics, by H.-S. Philip Wong and Deji Akinwande, Scope: monograph, reference, textbook. Level: non-specialist, postgraduate, advanced undergraduate, researcher. Contemporary Physics, 53(4):379–380, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pearce:2012:ESP

[2839] Frazer R. Pearce. Extra-Solar Planets: The Detection, Formation, Evolution and Dynamics of Planetary Systems, edited by Bonnie Steves, Martin Hendry and Andrew C. Cameron, Scope: edited book. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist. Contemporary Physics, 53(4):380, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2012:RND

[2840] Frank Berkshire. Reviews of Nonlinear Dynamics and Complexity, Volume 3, edited by Heinz Georg Schuster, Scope: review. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 53(4):

380–381, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gajjar:2012:HCA

[2841] Jitesh S. B. Gajjar. Handbook of Compressible Aerodynamics, by Jean Délery, Scope: handbook. Level: postgraduate, early career researcher, researcher, specialist, engineers. Contemporary Physics, 53(4):381–382, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dimakis:2012:IMI

[2842] Nicholas Dimakis. Introduction to Medical Imaging — Physics, Engineering and Clinical Applications, by Nicholas Dimakis, Scope: medical imaging. Level: advanced undergraduates and graduate students in medical imaging and related disciplines. Contemporary Physics, 53(4): 382–383, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2012:TAW

[2843] Frank Berkshire. Their Arrows Will Darken the Sun — The Evolution and Science of Ballistics, by Mark Denny, Scope: general interest, guide. Level: general readership, non-specialists, undergraduate, teacher, specialist, scientist. Contemporary Physics, 53(4):383–384, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kourkoumelis:2012:QUB

[2844] Nikolaos Kourkoumelis. Quantitative Understanding of Biosystems: An Introduction to Biophysics, by Thomas M. Nordlund, Scope: text book. Level: undergraduate, advanced undergraduate. Contemporary Physics, 53(4):384, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gotte:2012:AMR

[2845] Jörg Götte. Analytical Mechanics for Relativity and Quantum Mechanics, 2nd edition, by Oliver Davies Johns, Scope: monograph. Level: postgraduate, advanced undergraduate. Contemporary Physics, 53(4):385, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:UGQ

[2846] J. Rogel-Salazar. Ultracold Gases and Quantum Information: Lecture Notes of the Les Houches Summer School in Singapore, edited by C. Miniatura, L-C. Kwek, M. Ducloy, B. Grémaud, B-G. Englert, L. F.

Cugliandolo, A. Ekert and K. K. Phua, Scope: specialist. Level: postgraduate. Contemporary Physics, 53(4):385–386, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2012:MQS

[2847] A. M. Fox. Manipulating Quantum Structures Using Laser Pulses, by Bruce W. Shore, Scope: monograph. Level: postgraduate. Contemporary Physics, 53(4):386–387, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tucker:2012:CEC

[2848] R. Tucker. Covariant Electrodynamics: A Concise Guide, by John M. Charap, Scope: review. Level: advanced undergraduates. Contemporary Physics, 53(4):387, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:DCW

[2849] Derry W. Jones. Drive and Curiosity: What Fuels the Passion, by Istvan Hargittai, Scope: recent history of scientists. Level: general reader and scientists. Contemporary Physics, 53(4):387–389, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ironside:2012:MDF

[2850] Charlie Ironside. More is Different: Fifty Years of Condensed Matter Physics, edited by N. Phuan Ong and Ravin N. Bhatt, Scope: review. Level: researcher. Contemporary Physics, 53(4):389, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2012:GGP

[2851] Frank Berkshire. Gliding for Gold: The Physics of Winter Sports, by Mark Denny, Scope: general interest, guide, monograph. Level: general readership, non-specialists, undergraduate, teacher, scientist. Contemporary Physics, 53(4):389–390, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wykes:2012:IHV

[2852] Catherine M. Wykes. Introduction to Holography, by Vincent Toal, Scope: monograph. Level: postgraduate, advanced undergraduate, researcher. Contemporary Physics, 53(4):390–391, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2012:NPC

[2853] Peter J. Dobson. Nanosensors: Physical, Chemical, and Biological, by Vinod Kumar Khanna, Scope: reference, review, text book. Level: general readership, undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 53(4):391– 392, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:QME

[2854] J. Rogel-Salazar. Quantum Mechanics: Its Early Development and the Road to Entanglement and Beyond, by E. G. Steward, Scope: general science. Level: undergraduate. Contemporary Physics, 53(4):392, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:MIP

[2855] Thomas Peters. A Modern Introduction to Particle Physics, 3rd edition, by Fayyazuddin and Riazuddin, Scope: textbook. Level: postgraduate. Contemporary Physics, 53(4):393, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:QVS

[2856] Peter J. Bussey. The Quantum Vacuum: A Scientific and Philosophical Concept, from Electrodynamics to String Theory and the Geometry of the Microscopic World, by Luciano Boi, Scope: monograph. Level: researcher. Contemporary Physics, 53(4):393–394, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:BHW

[2857] Peter J. Bussey. Black Holes, Wormholes and Time Machines (2nd edition), by Jim Al-Khalili, Scope: general interest. Level: general readership. Contemporary Physics, 53(4):394, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2012:SNM

[2858] S. J. Blundell. Superconductivity in New Materials, Vol. 4, volume edited by Z. Fisk and H. R. Ott, Scope: edited book. Level: researcher. Contemporary Physics, 53(4):394–395, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2012:DSC

[2859] Allan I. Solomon. Dynamical Symmetry, by Carl E. Wulfman, Scope: textbook. Level: post-graduate. Contemporary Physics, 53(4):395–396, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2012:EMC

[2860] Amgad Salama. Exploring Monte Carlo Methods, by William L. Dunn and J. Kenneth Shultis, Scope: review. Level: undergraduate. Contemporary Physics, 53(4):396, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Messager:2012:HPP

[2861] Valérie Messager, Robert Gilmore, and Christophe Letellier. Henri Poincaré and the principle of relativity. Contemporary Physics, 53(5): 397–415, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kim:2012:LSE

[2862] Jeong San Kim, Gilad Gour, and Barry C. Sanders. Limitations to sharing entanglement. Contemporary Physics, 53(5):417–432, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2012:WLS

[2863] Peter V. E. McClintock. What is life?, Scope: review. Level: general reader. Contemporary Physics, 53(5):433–435, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:YWM

[2864] D. J. Miller. You are Wrong Mr Einstein! Newton, Einstein, Heisenberg and Feynman Discussing Quantum Mechanics, by Harald Fritzsch, Scope: general interest. Level: undergraduates and/or general readers. Contemporary Physics, 53(5):437–439, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bucknall:2012:SSS

[2865] David Bucknall. Solid-State Spectroscopy: An Introduction, 2nd edn., by Hans Kuzmany, Scope: textbook. Level: postgraduates, advanced undergraduates, early career researchers, researchers, teachers, scientists and engineers. Contemporary Physics, 53(5):439–440, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2012:TPE

[2866] Marco Casazza. Thermal Physics: Entropy and Free Energies, 2nd edn., by Joon Chang Lee, Scope: review. Level: undergraduates, postgraduates, advanced undergraduates, early career researchers and teachers.

Contemporary Physics, 53(5):440, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2012:PWS

[2867] Aniruddha Chakraborty. The Physicist’s World: The Story of Motion and the Limits to Knowledge, by Thomas Grissom, Scope: review. Level: undergraduate. Contemporary Physics, 53(5):440–441, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bohmer:2012:GPI

[2868] Christian G. Böhmer. The Geometry of Physics: An Introduction, 3rd edn., by Theodore Frankel, Scope: review. Level: undergraduates and graduate students, researchers. Contemporary Physics, 53(5):441–442, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:TCS

[2869] Thomas Peters. Turbulence, Coherent Structures, Dynamical Systems and Symmetry, 2nd edn., by P. Holmes, Scope: monograph. Level: researchers. Contemporary Physics, 53(5):442, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:CAA

[2870] P. J. Bussey. Cosmic Anger: Abdus Salam — The First Muslim Nobel Scientist, by Gordon Fraser, Scope: biography. Level: general readership. Contemporary Physics, 53(5):442–443, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:ODA

[2871] Manuel Vogel. Oxford Dictionary of Astronomy, by Ian Ridpath, Scope: reference. Level: general readership. Contemporary Physics, 53(5):443– 444, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2012:SPL

[2872] Salvatore Esposito. Statistical Physics of Liquids at Freezing and Beyond, by S. P. Das, Scope: monograph. Level: postgraduate, researcher and specialist. Contemporary Physics, 53(5):444–445, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:ASS

[2873] Thomas Peters. Advanced Solid State Physics, 2nd edn., by P. Phillips, Scope: textbook. Level: postgraduate. Contemporary Physics, 53(5):

445, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:CMT

[2874] Andrew Resnick. Continuum Mechanics and Thermodynamics: From Fundamental Concepts to Governing Equations, by Ellad B. Tadmor, Ronald E. Miller and Ryan S. Elliott, Scope: textbook. Level: postgraduate, advanced undergraduate, early career researcher, teacher, scientist, engineers. Contemporary Physics, 53(5):445–446, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:ATQ

[2875] Thomas Peters. Advanced Topics in Quantum Field Theory: A Lecture Course, by M. Shifman, Scope: textbook. Level: graduate. Contemporary Physics, 53(5):446–447, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:PCO

[2876] Andrew Resnick. Physics for Clinical Oncology, edited by Amen Sibtain, Scope: handbook. Level: teacher, specialist. Contemporary Physics, 53 (5):447, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pierri-Galvao:2012:NFC

[2877] Monica Pierri-Galvao. Neutrino, by Frank Close, Scope: general interest. Level: general reader, undergraduate through professionals. Contemporary Physics, 53(5):447–448, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2012:LSS

[2878] Mark Fox. Luminescence Spectroscopy of Semiconductors, Ivan Pelant and Jan Valenta, Scope: textbook. Level: postgraduate. Contemporary Physics, 53(5):448, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:PNC

[2879] K. Alan Shore. Plasmonic Nanoguides and Circuits, edited by Sergei I. Bozhevolnyi, Scope: monograph. Level: researcher, specialist. Contemporary Physics, 53(5):448–449, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:VNT

[2880] Andrew Resnick. Viral Nanoparticles: Tools for Materials Science and Biomedicine, by Nicole Steinmetz and Marianne Manchester, Scope: guide, monograph, review. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 53(5):449, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:CRP

[2881] D. J. Miller. Cosmic Rays for Particle and Astroparticle Physics, by S. Giani, C. Leroy and P. G. Rancoita, Scope: general interest, textbook. Level: pre-university, undergraduate, advanced undergraduate, teacher, scientist. Contemporary Physics, 53(5):450, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:2012:TEJ

[2882] John I. B. Wilson. Tokamaks, 4th edn., by John Wesson, Scope: monograph. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 53(5):450–451, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2012:TMF

[2883] David L. Andrews. Theory of Molecular Fluids Volume 2: Applications, by Christopher G. Gray, Keith E. Gubbins, and Christopher G. Joslin, Scope: monograph. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 53(5):451–452, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2012:SIA

[2884] Nikolaos E. Myridis. Science: Image in Action (The Science and Culture Series: Astrophysics), edited by Bertrand Zavidovique and Giosuè Lo Bosco, Scope: conference proceedings. Level: students, researchers and academics. Contemporary Physics, 53(5):452–453, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:APE

[2885] Manuel Vogel. Accelerator Physics, 3rd edn., by S. Y. Lee, Scope: textbook. Level: graduate students. Contemporary Physics, 53(5):453–454, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:MDN

[2886] Derry W. Jones. More and Different: Notes from a Thoughtful Curmudgeon, by Philip W. Anderson, Scope: biography, edited book, review. Level: general readership but parts need physics background. Contemporary Physics, 53(5):454–455, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holland:2012:HOG

[2887] Peter Holland. A History of Optics from Greek Antiquity to the Nineteenth Century, by Olivier Darrigol, Scope: monograph. Level: Postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist. Contemporary Physics, 53(5):455–456, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:SIO

[2888] K. Alan Shore. Statistics for Imaging, Optics, and Photonics, (Wiley Series in Probability and Statistics), by Peter Bajorski, Scope: monograph. Level: postgraduate and professional. Contemporary Physics, 53 (5):456, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2012:APS

[2889] Peter V. E. McClintock. Astronautics: The Physics of Space Flight, 2nd edn., by Ulrich Walter, Scope: textbook. Level: space industry professionals, postgraduates, senior undergraduates. Contemporary Physics, 53(5):456–457, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:PAC

[2890] Thomas Peters. Physics of the Atmosphere and Climate, 2nd edn., by M. L. Salby, Scope: textbook. Level: undergraduate. Contemporary Physics, 53(5):457–458, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2012:PSC

[2891] Tony Leggett. The Physics of Solar Cells, by Jenny Nelson, Scope: monograph. Level: advanced undergraduate, scientist. Contemporary Physics, 53(5):458–459, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Becker:2012:NKD

[2892] Peter Becker. The new kilogram definition based on counting the atoms in a^{28} si crystal. Contemporary Physics, 53(6):461–479, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2012:ERA

[2893] David Bailin. Essay review: AdS/CFT goes viral. String Theory and its Applications: From meV to the Planck Scale, edited by M. Dine, T. Banks, and S. Sachdev. ISBN 978-981-4350-51-8. Contemporary Physics, 53(6):481–483, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2012:SGP

[2894] S. C. Tiwari. Symmetry and geometry: pursuit of beauty in physics, Scope: textbook/monograph. level Ph.D. students/researchers. Contemporary Physics, 53(6):485–490, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2012:UAD

[2895] Eric Sheldon. The Universe in all its dimensions, Scope: review, textbook. Level: undergraduate and graduate students, scientists, general readership. Contemporary Physics, 53(6):491–494, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2012:IGS

[2896] Eric Sheldon. The imprint of genius, Scope: textbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 53(6):495–498, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benedict:2012:BAQ

[2897] Keith A. Benedict. Basic Aspects of the Quantum Theory of Solids: Order and Elementary Excitations, by Daniel I. Khomskii, Scope: monograph. Level: postgraduate. Contemporary Physics, 53(6):499–500, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kuhn:2012:SSV

[2898] Jeff Kuhn. Solar and Stellar Variability: Impact on Earth and Planets, IAU S264, by A. G. Kosovichev, A. H. Andrei and J-P Rozelot. Scope: conference proceedings. Level: postgraduate, researcher, specialist. Contemporary Physics, 53(6):500–501, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2012:EQP

[2899] Peter W. Milonni. Exploring Quantum Physics through Hands-On Projects, by D. Prutchi and S. R. Prutchi, Scope: general interest, guide, monograph. Level: pre-university, undergraduates, advanced undergraduates, teachers. Contemporary Physics, 53(6):501–502, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:MEV

[2900] Manuel Vogel. Mechanical and Electromagnetic Vibrations and Waves, by T. Bécherrawy, Scope: textbook. Level: undergraduate students. Contemporary Physics, 53(6):502–503, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2012:DAP

[2901] P. T. Greenland. Data Analysis for Physical Scientists Featuring Excel(R), 2nd edn., by Les Kirkup, Scope: textbook. Level: undergraduates, advanced undergraduates. Contemporary Physics, 53(6):503, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2012:GCT

[2902] Peter V. E. McClintock. Graphene: Carbon in Two Dimensions, by Mikhail I. Katsnelson, Scope: monograph. Level: postgraduate students, researchers. Contemporary Physics, 53(6):503–504, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:RCG

[2903] Thomas Peters. Relativistic Cosmology, by G. F. R. Ellis, R. Maartens and M. A. H. MacCallum, Scope: textbook. Level: graduates. Contemporary Physics, 53(6):504–505, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:QPI

[2904] Peter J. Bussey. Quantum Physics: Illusion or Reality?, by Alastair M. Rae, Scope: review. Level: general readership. Contemporary Physics, 53(6):505–506, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2012:TAR

[2905] F. W. Taylor. Theory of Atmospheric Radiative Transfer: A Comprehensive Introduction, by Manfred Wendisch and Ping Yang, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 53(6):

506, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Evans:2012:MIA

[2906] Mike Evans. Meteorites, Ice and Antarctica: A Personal Account, by William A. Cassidy, Scope: monograph/general interest. Level: general reader. Contemporary Physics, 53(6):506–507, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2012:SME

[2907] R. Ellis. Stellar Magnetism, 2nd edn., by Leon Mestel, Scope: textbook. Level: postgraduates, advanced undergraduates, early career researchers, researchers, teachers, specialists, scientists. Contemporary Physics, 53 (6):507–508, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pedder:2012:MTP

[2908] Chris Pedder. Mr Tompkins in Paperback, by George Gamow, Scope: general interest. Level: general readership, non-specialists, preuniversity, undergraduate. Contemporary Physics, 53(6):508–509, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:OLE

[2909] J. Rogel-Salazar. Organic Light-Emitting Diodes: The Use of Rare-Earth Transition Metals, by Luiz Pereira, Scope: review. Level: advanced graduate and postgraduate, researchers, specialists. Contemporary Physics, 53(6):509, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2012:PPK

[2910] R. S. Shorter. Plant Physics, by Karl Niklas, Scope: textbook. Level: specialists, engineers. Contemporary Physics, 53(6):509–510, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bengtsson:2012:TFG

[2911] Ingemar Bengtsson. Topics in the Foundations of General Relativity and Newtonian Gravitation Theory, by David B. Malament, Scope: monograph, textbook. Level: postgraduates, advanced undergraduates, early career researcher. Contemporary Physics, 53(6):510–511, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2012:AEM

[2912] Peter J. Bussey. Albert Einstein Memorial Lectures, by Jacob D. Bekenstein and Raphael Mechoulam. Scope: edited book. Level: nonspecialists. Contemporary Physics, 53(6):511–512, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2012:WPE

[2913] R. S. Shorter. The Wonders of Physics, 3rd edn., by Andrey Varlamov and Lev Aslamazov, Scope: review. Level: general readership, non-specialists, pre-university, teachers. Contemporary Physics, 53(6): 512, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:IBH

[2914] Thomas Peters. Introduction to Black Hole Physics, by V. P. Frolov and A. Zelnikov, Scope: textbook. Level: postgraduate. Contemporary Physics, 53(6):512–513, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:2012:PMJ

[2915] Lionel Wilson. The Physics of Microdroplets, by Jean Berthier and Kenneth A. Brakke, Scope: review. Level: postgraduates, advanced undergraduates, researchers, scientists, engineers. Contemporary Physics, 53 (6):513–514, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:BCB

[2916] David J. Miller. From Black Clouds to Black Holes, 3rd edn., by Jayant V. Narlikar, Scope: general interest. Level: general readership, nonspecialists. Contemporary Physics, 53(6):514–515, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2012:UCM

[2917] Chandra Wickramasinghe. The Universe: A Challenge to the Mind, by Jacques Vanier, Scope: general interest, textbook. Level: undergraduate. Contemporary Physics, 53(6):515, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2012:NMP

[2918] M. Jamil. Nuclear Medicine Physics, edited by J. J. De Lima, Scope: monograph. Level: graduate students and researchers. Contemporary Physics, 53(6):516, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wykes:2012:HSN

[2919] Catherine M. Wykes. Handbook of Surface and Nanometrology, 2nd edn., by David J. Whitehouse, Scope: review. Level: postgraduates, advanced undergraduates, early career researchers, researchers, specialists, engineers. Contemporary Physics, 53(6):516–517, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:2012:PRF

[2920] John I. B. Wilson. Physics of Radio-Frequency Plasmas, by Pascal Chabert and Nicholas Braithwaite, Scope: monograph, textbook. Level: postgraduates, early career researchers, researchers, scientists, engineers. Contemporary Physics, 53(6):517–519, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Agio:2012:AAP

[2921] Mario Agio. Advances in Atomic Physics: An Overview, by Claude Cohen-Tannoudji and David Guéry-Odelin, Scope: monograph, text book. Level: postgraduates, advanced undergraduates, early career researchers, researchers, teachers, specialists, scientists. Contemporary Physics, 53(6):519–520, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2012:ACD

[2922] Chandra Wickramasinghe. Adventures in Cosmology, by David Goodstein, Scope: edited book. Level: advanced undergraduates. Contemporary Physics, 53(6):520, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:NSE

[2923] David J. Miller. From Nuclei to Stars, edited by Sabine Lee, Scope: edited book. Level: specialists. Contemporary Physics, 53(6):520–521, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2012:FLM

[2924] Andrew Resnick. Femtosecond Laser–Matter Interactions: Theory, Experiments, and Applications, by Eugene G. Gamaly, Scope: monograph. Level: postgraduates, advanced undergraduates, early career researchers, teachers, scientists, engineers. Contemporary Physics, 53(6):521–522, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2012:FNT

[2925] K. Alan Shore. Fracture Nanomechanics, by Takayuki Kitamura, Hiroyuki Hirakata, Takashi Sumigawa and Takahiro Shimada, Scope: monograph. Level: researchers, specialists. Contemporary Physics, 53 (6):522, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2012:PSI

[2926] David Bailin. Perspectives on Supersymmetry II, Advanced Series on Directions in High Energy Physics: Volume 21, edited by Gordon L. Kane, Scope: edited book. Level: researcher. Contemporary Physics, 53 (6):522–524, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benhamou:2012:PVB

[2927] Mabrouk Benhamou. Polymer Viscoelasticity — Basics, Molecular Theories, Experiments and Simulations, 2nd edn, by Yn-Hwang Lin, Scope: textbook. Level: researchers. Contemporary Physics, 53(6):524, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:RTE

[2928] Manuel Vogel. The Release of Thermonuclear Energy by Inertial Confinement: Ways Towards Ignition, by Friedwardt Winterberg, Scope: monograph. Level: scientists, engineers. Contemporary Physics, 53(6): 525, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:ESU

[2929] David J. Miller. Exploding Superstars: Understanding Supernovae and Gamma Ray Bursts, by Alain Mazure and Stéphane Basa, Scope: general interest, textbook. Level: general readership, advanced undergraduates, early career researchers, scientists. Contemporary Physics, 53(6): 525–526, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2012:PTS

[2930] P. T. Greenland. Principles of Terahertz Science and Technology, by Yun-Shik Lee, Scope: monograph. Level: researchers, specialists. Contemporary Physics, 53(6):526–527, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eberly:2012:SPR

[2931] J. H. Eberly. Stochastic Processes, by Richard F. Bass, Scope: textbook. Level: postgraduates, researchers. Contemporary Physics, 53(6): 527, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2012:CSP

[2932] D. G. C. Jones. Compound Semiconductor Photonics: Materials, Devices and Integration, by Chua Soo-Jin, Teng Jinghua and Aaron Danner, Scope: conference proceedings. Level: researchers. Contemporary Physics, 53(6):527–528, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2012:QAD

[2933] David J. Miller. The Quantum Adventure: Does God Play Dice?, by Alex Montwill and Ann Breslin, Scope: general interest. Level: general readership, undergraduates, teachers. Contemporary Physics, 53(6): 528–529, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:FFB

[2934] Manuel Vogel. Fluid Flows to Black Holes: A Tribute to S. Chandrasekhar on His Birth Centenary, by D. J. Saikia and V. Trimble, Scope: biography. Level: general readership. Contemporary Physics, 53(6):529– 530, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2012:MLQ

[2935] Thomas Peters. The Mathematical Language of Quantum Theory: From Uncertainty to Entanglement, by T. Heinosaari and M. Ziman, Scope: textbook. Level: postgraduates. Contemporary Physics, 53(6):530, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2012:LQM

[2936] Michael Belsley. Lectures on Quantum Mechanics, 2nd edn, by Ashok Das, Scope: textbook. Level: postgraduates, advanced undergraduates. Contemporary Physics, 53(6):530–531, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2012:IMP

[2937] Salvatore Esposito. Introduction to Modern Physics: Theoretical Foundations, by John Dirk Walecka, Scope: textbook. Level: undergraduates, teachers. Contemporary Physics, 53(6):531–532, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2012:FIP

[2938] R. S. Shorter. The Fundamentals of Imaging: From Particles to Galaxies, by Michael Mark Woolfson, Scope: textbook. Level: General readership, pre-university, teachers. Contemporary Physics, 53(6):532, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coule:2012:QGE

[2939] David H. Coule. Quantum Gravity, 3rd edn, by Claus Kiefer, Scope: monograph, textbook. Level: postgraduates, early career researchers, researchers. Contemporary Physics, 53(6):532–533, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2012:PRP

[2940] Nikolaos E. Myridis. Probability, Random Processes, and Statistical Analysis, by H. Kobayashi, B. L. Mark and W. Turin, Scope: textbook. Level: undergraduate students, researchers, teachers. Contemporary Physics, 53(6):533–534, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

El-Amin:2012:RFP

[2941] M. F. El-Amin. Rotating Flow, by Peter Childs, Scope: review. Level: postgraduates, scientists, engineers. Contemporary Physics, 53(6):534, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2012:DSP

[2942] Aniruddha Chakraborty. Disordered Semiconductors — Physics and Applications, by Anatoly Popov, Scope: review. Level: advanced undergraduates. Contemporary Physics, 53(6):534–535, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2012:CAH

[2943] J. Rogel-Salazar. Computational Analysis of the Human Eye with Applications, by Sumeet Dua, Rajendra Acharya U. and E. Y. K. Ng, Scope: review. Level: postgraduates, advanced undergraduates, researchers, specialists. Contemporary Physics, 53(6):535–536, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2012:SGV

[2944] Peter R. Hobson. A Student’s Guide to Vectors and Tensors, by Daniel Fleisch, Scope: textbook. Level: pre-university, undergraduates, teachers, scientists, engineers. Contemporary Physics, 53(6):536, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2012:QBI

[2945] Aniruddha Chakraborty. Quantum Bio-Informatics IV: From Quantum Information to Bio-Informatics, 1st edn, by Luigi Accardi Wolfgang Freudenberg and Masanori Ohya, Scope: review. Level: researchers in quantum information, bio-informatics and life science. Contemporary Physics, 53(6):537, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2012:SQB

[2946] Nikolaos E. Myridis. Semiconductor Quantum Bits, edited by F. Henneberger and O. Benson, Scope: edited book. Level: researchers, theorists, postgraduate students. Contemporary Physics, 53(6):537–538, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2012:SQM

[2947] Allan I. Solomon. Supersymmetric Quantum Mechanics: An Introduction, by A. Gangopadhyaya, J. V. Mallow and C. Rasinariu, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 53(6):538–539, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2012:TQY

[2948] Aniruddha Chakraborty. The Thermodynamics of Quantum Yang–Mills Theory: Theory And Applications, by Ralf Hofmann, Scope: review. Level: researchers. Contemporary Physics, 53(6):539, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2012:AIL

[2949] Mario Bertolotti. Atoms in Intense Laser Fields, by C. J. Joachain, N. J. Kylstra and R. M. Potvliege, Scope: monograph. Level: postgraduate, researcher, scientist. Contemporary Physics, 53(6):539–540, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2012:EHD

[2950] Gerard Milburn. Essentials of Hamiltonian Dynamics, by J. H. Lowenstein, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 53(6):541, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2012:BTM

[2951] Allan I. Solomon. A Brief Tour of Modern Quantum Mechanics, by Moshe Gitterman, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 53(6):541–542, 2012. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Solomon:2012:CME

[2952] Allan I. Solomon. Classical Mechanics, 5th edn., by Tom W. B. Kibble and Frank H. Berkshire, Scope: textbook. Level: undergraduate. Contemporary Physics, 53(6):542–543, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2012:SDE

[2953] Suresh Chandra Tiwari. The Supersymmetric Dirac Equation: The Application to Hydrogenic Atoms, by Allen Hirshfeld, Scope: textbook. Level: undergraduates. Contemporary Physics, 53(6):543–544, 2012. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2012:PRI

[2954] Peter R. Hobson. Principles of Radiation Interaction in Matter and Detection, 3rd edn., by Claude Leroy and Pier-Giorgio Rancoita, Scope: monograph, reference. Level: postgraduates, early career researchers, researchers, specialists, scientists, engineers. Contemporary Physics, 53 (6):544, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bucknall:2012:ITT

[2955] David Bucknall. Introduction to the Theory of Thermal Neutron Scattering, 3rd edn., by G. L. Squires, Scope: textbook. Level: early career researchers, researchers, specialists, scientists. Contemporary Physics, 53 (6):544–545, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:FNW

[2956] Manuel Vogel. Fundamentals of Numerical Weather Prediction, by Jean Coiffier, Scope: monograph. Level: advanced undergraduate students.

Contemporary Physics, 53(6):545–546, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2012:PGC

[2957] Manuel Vogel. A Philosophical Guide to Chance, by T. Handfield, Scope: monograph. Level: general readership. Contemporary Physics, 53(6): 546–547, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2012:EB

[2958] Anonymous. Editorial board. Contemporary Physics, 53(6):ebi, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lyons:2013:BFP

[2959] Louis Lyons. Bayes and frequentism: a particle physicist’s perspective. Contemporary Physics, 54(1):1–16, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schinckus:2013:IET

[2960] Christophe Schinckus. Introduction to econophysics: towards a new step in the evolution of physical sciences. Contemporary Physics, 54(1):17– 32, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Green:2013:IGG

[2961] A. G. Green. An introduction to gauge-gravity duality and its application in condensed matter. Contemporary Physics, 54(1):33–48, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2013:MWM

[2962] D. G. C. Jones. Many worlds make hard work. Contemporary Physics, 54 (1):49–51, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2013:RMS

[2963] Geoffrey Duxbury. Reinventing molecular spectroscopy. Contemporary Physics, 54(1):52–54, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2013:CPJ

[2964] Chandra Wickramasinghe. Cosmic Paradoxes, by Julio A. Gonzalo, Scope: general interest. Level: non-specialists. Contemporary Physics, 54(1):55, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2013:TDD

[2965] A. H. Harker. Time-Dependent Density-Functional Theory: Concepts and Applications, by Carsten A. Ullrich, Scope: monograph, textbook. Level: postgraduate, early career researcher. Contemporary Physics, 54 (1):55–56, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2013:OAE

[2966] S. J. Blundell. Orbital Approach to the Electronic Structure of Solids, by E. Canadell, M.-L. Doublet and C. Iung, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 54(1): 56–57, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wilson:2013:PPP

[2967] John I. B. Wilson. Principles of Plasma Physics for Engineers and Scientists, by Umran S. Inan and Marek Gołkowski, Scope: textbook. Level: postgraduates, advanced undergraduates, early career researchers, teachers, scientists, engineers. Contemporary Physics, 54(1):57–58, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2013:GJH

[2968] F. W. Taylor. Galileo, by J. L. Heilbron, Scope: monograph. Level: generalreaders. Contemporary Physics, 54(1):58, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:TKG

[2969] Manuel Vogel. Thermodynamics, by S. Klein and G. Nellis, Scope: textbook. Level: students and engineers. Contemporary Physics, 54(1):59, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:BSO

[2970] Manuel Vogel. Beyond the Stars: Our Origins and the Search for Life in The Universe, by P. Saraceno, Scope: general interest. Level:general readers. Contemporary Physics, 54(1):59–60, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:PCM

[2971] Thomas Peters. Physics of Continuous Matter: Exotic and Everyday Phenomena in the Macroscopic World, 2nd ed., by B. Lautrup, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(1):60, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:CCA

[2972] Thomas Peters. Chaos and Complexity in Astrophysics, by O. Regev, Scope: text-book. Level: graduate. Contemporary Physics, 54(1):60– 61, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2013:HMM

[2973] Peter R. T. Munro. Handbook of Mathematical Methods in Imaging, edited by Otmar Scherzer, Scope: handbook. Level: specialist. Contemporary Physics, 54(1):61–62, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riley:2013:SWB

[2974] N. Riley. Shock Wave-Boundary-Layer Interactions, edited by Holger Babinsky and John K. Harvey, Scope: monograph. Level: postgraduate. Contemporary Physics, 54(1):62–63, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dorai:2013:SPM

[2975] Kavita Dorai. Signal Processing in Magnetic Resonance Spectroscopy with Biomedical Applications, by Dzevad Belkic and Karen Belkic, Scope: review. Level:undergraduate. Contemporary Physics, 54(1):63–64, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2013:IGF

[2976] Amgad Salama. Introduction to Geophysical Fluid Dynamics: Physical and Numerical Aspects, 2nd edn., by B. Cushman-Roisin and J.-M. Beckers, Scope: review. Level: undergraduate. Contemporary Physics, 54(1): 64–65, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2013:PGG

[2977] Marco Casazza. Pierre-Gilles de Gennes: A Life in Science by Laurence Plévert, Scope: review. Level: general readership. Contemporary Physics, 54(1):65–66, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sivakumar:2013:CEN

[2978] Srinivasan Sivakumar. Classical Electromagnetism in a Nutshell, by Anupam Garg, Scope: textbook. Level: postgraduate. Contemporary Physics, 54(1):66, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holland:2013:TDQ

[2979] Peter Holland. A Trajectory Description of Quantum Processes. I. Fundamentals: A Bohmian Perspective, by Á.S. Sanz and Á.S. Miret-Artés, Scope: monograph. Level: postgraduate. Contemporary Physics, 54(1): 66–67, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2013:PFR

[2980] Derry W. Jones. Prize Fight: The Race and the Rivalry to be the First in Science, by Morton A. Myers, Scope: review. Level: general readers and scientists. Contemporary Physics, 54(1):67–68, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:PPS

[2981] Peter J. Bussey. Philosophy of Physics: Space and Time, by T. Maudlin, Scope: textbook. Level: teachers. Contemporary Physics, 54(1):68–69, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Starkey:2013:EBG

[2982] Natalie A. Starkey. The Earth: Its Birth and Growth, 2nd edn., by Minoru Ozima, Jun Korenaga and Qing-Zhu Yin, Scope: general interest, textbook. Level: undergraduate,graduate, general readership, teachers. Contemporary Physics, 54(1):69–70, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2013:INC

[2983] David J. Miller. Interpreting Newton: Critical Essays, edited by Andrew Janiak and Eric Schliesser, Scope:biography. Level: specialist. Contemporary Physics, 54(1):70, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coule:2013:FST

[2984] David H. Coule. Foundations of Space and Time: Reflections on Quantum Gravity, edited by Jeff Murugan, Amanda Weltman and George F. R. Ellis, Scope:conference proceedings. Level: researcher, specialist.

Contemporary Physics, 54(1):70–71, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:QIC

[2985] Thomas Peters. Quantum Information, Computation and Communication, by J. A. Jones and D. Jaksch, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(1):71–72, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2013:HID

[2986] S. C. Tiwari. Higgs: The Invention and Discovery of the ‘God Particle’, by J. Baggott, Scope: general interest. Level: generalreaders, non-specialists. Contemporary Physics, 54(1):72–73, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sofo:2013:PSS

[2987] Jorge O. Sofo. Problems in Solid State Physics with Solutions, by Fuxiang Han, Scope: guide. Level: postgraduate. Contemporary Physics, 54(1):73, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2013:INE

[2988] Peter J. Dobson. Intelligent Nanomaterials, edited by A. Tiwari, A. K. Mishra, H. Kobayashi and A. P. F. Turner. General interest. Level: nonspecialists, early career researchers, teachers. Contemporary Physics, 54 (1):74, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2013:APV

[2989] David Bailin. ABC of Physics: A Very Brief Guide, by Lev B. Okun, Scope: general interest. Level: non-specialists, early career researchers, teachers. Contemporary Physics, 54(1):74–75, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

vanKampen:2013:MVS

[2990] Paul van Kampen. Magnetism: A Very Short Introduction, by Stephen Blundell, Scope: general interest. Level: general readership. Contemporary Physics, 54(1):75–76, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2013:BMT

[2991] Peter J. Dobson. Biosensors and Molecular Technologies for Cancer Diagnostics, edited by Keith E. Herold and Avraham Rasooly. Contemporary Physics, 54(1):76–77, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2013:QMT

[2992] Mark Fox. Quantum Mechanics: Theory and Experiment, by Mark Beck, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 54(1):77, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:HES

[2993] B. Ishak. Heliophysics: Evolving Solar Activity and the Climates of Space and Earth, edited by C. J. Schrijver and G. L. Siscoe, Scope: advanced undergraduates, graduates, post-docs, specialists. Contemporary Physics, 54(1):77–78, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2013:ETE

[2994] F. W. Taylor. Entropy and the Time Evolution of Macroscopic Systems, by Walter Grandy, Jr., Scope: textbook. Level: scientist. Contemporary Physics, 54(1):78–79, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2013:IQM

[2995] Aniruddha Chakraborty. Introduction to Quantum Mechanics: Schrödinger Equation and Path Integral, 2nd edn., by Harald J. W. Müller-Kirsten, Scope: review. Level: postgraduate. Contemporary Physics, 54(1):79, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:PCE

[2996] B. Ishak and J. McFarland. Physics of Comets, 3rd edn., by K. S. Krishna Swamy, Scope: monograph. Level: specialist. Contemporary Physics, 54 (1):79–80, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2013:ITC

[2997] Andrew Resnick. Introduction to Theoretical and Computational Fluid Dynamics, 2nd edn., by Constantine Pozrikidis, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 54(1): 80–81, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2013:OES

[2998] Nikolaos E. Myridis. Optical Electronics: Self-Organized Integration and Applications, by Tetsuzo Yoshimura, Scope: monograph. Level: researchers, scientists, postgraduate students. Contemporary Physics, 54 (1):81–82, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2013:CSQ

[2999] G. W. Gibbons. Classical Solutions in Quantum Field Theory: Solitons and Instantons in High Energy Physics, by E. J. Weinberg, Scope: monograph. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(1):82–83, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:ICM

[3000] Thomas Peters. An Introduction to Celestial Mechanics, by R. Fitzpatrick, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(1):83–84, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:ESG

[3001] Manuel Vogel. Europe to the Stars, by G. Schilling and L. L. Christensen, Scope: general interest. Level: general readership. Contemporary Physics, 54(1):84, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2013:DHG

[3002] Andrew Resnick. Dynamical Heterogeneities in Glasses, Colloids, and Granular Media, edited by Ludovic Berthier, Giulio Biroli, Jean-Philippe Bouchaud, Luca Cipelletti, Wim van Saarloos, Scope: general interest. Level: general readership. Contemporary Physics, 54(1):84–85, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McGlynn:2013:ZNT

[3003] Enda McGlynn. ZnO Nanostructures and Their Applications, edited by Xiao Wei Sun and Yi Yang, Scope: review, monograph. level: postgraduate, early career researcher, researcher,specialist, scientist, engineers. Contemporary Physics, 54(1):85–86, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2013:BST

[3004] G. W. Gibbons. The Birth of String Theory, edited by Andrea Cappelli, Elena Castellani, Filippo Colomo and Paolo Di Vecchia, Scope: review. Level: general readers, postgraduates,researchers. Contemporary Physics, 54(1):86–87, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ma:2013:MCD

[3005] Angie Ma. Mechanics of the Cell, by David Boal, Scope: textbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 54(1): 87–88, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palffy:2013:PEE

[3006] Adriana Palffy. The Practical Einstein: Experiments, Patents, and Inventions, by József Illy, Scope: biography, general interest. Level: general readers. Contemporary Physics, 54(1):88–89, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:CSW

[3007] B. Ishak. Coronal Seismology: Waves and Oscillations in Stellar Coronae, by A. V. Stepanov, V. V. Zaitsev and V. M. Nakariakov, Scope: monograph. Level: early career researchers, researchers, specialists. Contemporary Physics, 54(1):89, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mindlin:2013:PBP

[3008] G. B. Mindlin. The physics of birdsong production. Contemporary Physics, 54(2):91–96, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2013:SWE

[3009] P. T. Greenland. Should we eat Einstein’s brain?, Scope: general interest. Level: general readership. Contemporary Physics, 54(2):97–99, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:WMW

[3010] Peter J. Bussey. The world of many worlds, Scope: edited book, monograph. Level: specialist. Contemporary Physics, 54(2):100–103, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2013:FAO

[3011] Mario Bertolotti. Fundamentals of Attosecond Optics, by Zenghu Chang, Scope: monograph. Level: postgraduate, advanced undergraduate, researchers, engineers. Contemporary Physics, 54(2):104–105, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2013:HPP

[3012] Matt Probert. High-Pressure Physics, edited by John Loveday. Scope: conference proceedings, edited book, reference. Level: postgraduate, early career researcher. Contemporary Physics, 54(2):105, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2013:CSR

[3013] M. Jamil. Compound Semiconductor Radiation Detectors: Series in Sensors, by Alan Owens, Scope: monograph. Level: graduate students, researchers. Contemporary Physics, 54(2):105–106, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2013:IPS

[3014] Gunter M. Schütz. Introduction to Probability and Stochastic Processes with Applications, by L. B. Castañeda, V. Arunachalam and S. Dharmaraja, Scope: textbook. Level: non-specialists, advanced undergraduates. early career researchers, engineers. Contemporary Physics, 54(2): 106–107, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Michette:2013:XRN

[3015] Alan Michette. X-Rays, Neutrons and Muons: Photons and Particles for Material Characterization, by Walter E. Fischer and Rudolf Morf, Scope: textbook. Level: postgraduate, early career researcher. Contemporary Physics, 54(2):107, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:PAE

[3016] B. Ishak. Pulsar Astronomy, 4th edn., by A. Lyne and F. Graham-Smith, Scope: monograph. Level: Ph.D.s, post-docs and, specialists. Contemporary Physics, 54(2):107–108, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:RJT

[3017] Peter J. Bussey. Reminiscences: A Journey Through Particle Physics, by Adrian Melissinos, Scope: biography. Level: postgraduate, researcher.

Contemporary Physics, 54(2):108–109, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leslie:2013:SVS

[3018] Dawn E. Leslie. Stars: A Very Short Introduction, by Andrew King, Scope: guide, general interest. Level: general readers, non-specialists, pre-undergraduate, undergraduate, teacher. Contemporary Physics, 54 (2):109, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:QOG

[3019] Thomas Peters. Quantum Optics, by G. S. Agarwal, Scope: textbook. Level: graduate. Contemporary Physics, 54(2):109–110, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:IAM

[3020] Manuel Vogel. The International Atlas of Mars Exploration: The First Five Decades, Vol. 1, 1953 to 2003, Scope: general interest. Level: general readers. Contemporary Physics, 54(2):110–111, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:UGT

[3021] Manuel Vogel. A Unified Grand Tour of Theoretical Physics, 3rd edn., by Ian D. Lawrie, Scope: textbook. Level: advanced undergraduates and above. Contemporary Physics, 54(2):111–112, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2013:UAO

[3022] Michael Belsley. Ultracold Atoms in Optical Lattices: Simulating Quantum Many-Body Systems, by Maciej Lewenstein, Anna Sanpera and Verònica Ahufinger, Scope: monograph. Level: postgraduate, early career researcher, specialist. Contemporary Physics, 54(2):112–113, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coule:2013:TRC

[3023] David H. Coule. Time Reversibility, Computer Simulation, Algorithms, 2nd edn., by William G. Hoover and Carol G. Hoover, Scope: textbook. Level: postgraduate, advanced undergraduate, researcher. Contemporary Physics, 54(2):113, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Korotkov:2013:FNO

[3024] Alexander Korotkov. Fluctuating Nonlinear Oscillators: From Nano-Mechanics to Quantum Superconducting Circuits, edited by Mark I. Dykman, Scope: edited book. Level: postgraduate, researcher. Contemporary Physics, 54(2):113–114, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2013:FGF

[3025] Juan A. Añel. Fundamentals of Geophysical Fluid Dynamics, by James C. McWilliams, Scope: textbook. Level: undergraduate students. Contemporary Physics, 54(2):114–115, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McGlynn:2013:EPR

[3026] Enda McGlynn. Electron Paramagnetic Resonance of Transition Ions, Oxford Classic Texts in the Physical Sciences, by A. Abragam and B. Bleaney, Scope: textbook, monograph. Level: advanced undergraduate, postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 54(2):115–116, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Murdin:2013:QCG

[3027] B. N. Murdin. Quantum Computing from the Ground Up, by Riley Tipton Perry, Scope: monograph. Level: undergraduate. Contemporary Physics, 54(2):116–117, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2013:ATF

[3028] J. Rogel-Salazar. Advances in Thin-Film Solar Cells, by I. M. Dharmadasa, Scope: textbook. Level: undergraduate, postgraduate. Contemporary Physics, 54(2):117, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2013:CEQ

[3029] S. J. Blundell. Correlated Electrons in Quantum Matter, by Peter Fulde, Scope: monograph. Level: postgraduate. Contemporary Physics, 54(2): 117–118, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Murdin:2013:SSJ

[3030] B. N. Murdin. Semiconductor Spintronics, by Jianbai Xia, Weikun Ge, Kai Chang, Scope: reference. Level: researcher. Contemporary Physics, 54(2):118–119, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2013:NSP

[3031] Nikolaos E. Myridis. Networks in Social Policy Problems, edited by Balázs Vedres and Marco Scotti, Scope: edited book. Level: general readers, researcher, scientist. Contemporary Physics, 54(2):119–120, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2013:PNV

[3032] Suresh Chandra Tiwari. The Physics of Neutrinos, by V. Barger, D. Marfatia and K. Whisnant, Scope: review, monograph. Level: specialist, researcher, teacher. Contemporary Physics, 54(2):120–121, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2013:CBM

[3033] Matt Probert. Computer-Based Modelling of Novel Carbon Systems and Their Properties: Beyond Nanotubes, edited by L. Colombo and A. Fasolino, Scope: edited book, reference. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(2):121, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2013:DWR

[3034] Tom Lancaster. Do We Really Understand Quantum Mechanics?, by Franck Laloë, Scope: monograph. Level: advanced undergraduate to specialist. Contemporary Physics, 54(2):121–122, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2013:WOH

[3035] Marco Casazza. The Weather Observer’s Handbook, by Stephen Burt, Scope: handbook, manual, textbook. Level: general readership, nonspecialists, pre-university, undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist. engineers. Contemporary Physics, 54(2):122–123, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gotte:2013:PNO

[3036] Jörg Götte. Principles of Nano-Optics, 2nd edn., by Lukas Novotny and Bert Hecht, Scope: textbook. Level: postgraduate, researcher, teacher, specialist. Contemporary Physics, 54(2):123–124, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:RMR

[3037] Peter J. Bussey. Relativity Made Relatively Easy, by Andrew M. Steane, Scope: textbook. Level: postgraduate. advanced undergraduate. Contemporary Physics, 54(2):124, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2013:NET

[3038] Marco Casazza. Non-Equilibrium Thermodynamics and Statistical Mechanics: Foundations and Applications, by Phyl Attard, Scope: monograph, reference. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 54(2): 124–125, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2013:DMS

[3039] Chih-Yueh Wang. Death of Massive Stars (IAU S279): Supernovae and Gamma-Ray Bursts, Proceedings of the International Astronomical Union (IAU S279), edited by P. W. A. Roming, N. Kawai and E. Pian., Scope: conference proceeding. Level: postgraduate/specialist. Contemporary Physics, 54(2):125–126, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2013:BHC

[3040] G. W. Gibbons. Black Holes, Cosmology and Extra Dimensions, by Kirill A. Bronnikov and Sergey G. Rubin, Scope: review. Level: postgraduate. Contemporary Physics, 54(2):126–127, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:TPS

[3041] Manuel Vogel. The Theory of Probability, by Santosh S. Venkatesh, Scope: text book. Level: advanced undergraduate. Contemporary Physics, 54(2):127, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:WYH

[3042] Manuel Vogel. Why You Hear What You Hear, Eric J. Heller, Scope: general interest, monograph. Level: general readership. Contemporary Physics, 54(2):127–128, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:GRW

[3043] B. Ishak. A General Relativity Workbook, by T. A. Moore, Scope: textbook. Level: advanced undergraduates, lecturers. Contemporary Physics, 54(2):128–129, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:IPC

[3044] B. Ishak. Illustrating the Phaenomena: Celestial Cartography in Antiquity and the Middle Ages, by E. Dekker, Scope: monograph. Level: specialists. Contemporary Physics, 54(2):129–130, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:LQM

[3045] Thomas Peters. Lectures on Quantum Mechanics, by S. Weinberg, Scope: textbook. Level: graduate. Contemporary Physics, 54(2):130–131, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2013:MSQ

[3046] Garth Jones. Mathematical Structures of Quantum Mechanics, by Kow Lung Chang, Scope: text book. Level: postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 54 (2):131–132, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2013:ABM

[3047] P. T. Greenland. Applied Bohmian Mechanics from Nanoscale Systems to Cosmology, edited by X. Oriols, J. Mompart, Scope: edited book. Level: researcher. Contemporary Physics, 54(2):132–133, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2013:NBQ

[3048] R. S. Shorter. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925, by Helge Kragh, Scope: biography, general interest. Level: general readership. Contemporary Physics, 54(2): 133, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2013:MUM

[3049] Mark Nicol. Measuring the Universe: A Multiwavelength Perspective, by George H. Rieke, Scope: review. Level: postgraduate, advanced undergraduate, scientist, engineers. Contemporary Physics, 54(2):133–134, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kourkoumelis:2013:MDM

[3050] Nikolaos Kourkoumelis. Modern Diffraction Methods, edited by E. J. Mittemeijer and U. Welzel, Scope: edited book, handbook. Level: postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 54(2):134–135, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benhamou:2013:ACN

[3051] Mabrouk Benhamou. Advances in Carbon Nanomaterials: Science and Applications, by Nikos Tagmatarchis, Scope: edited book. Level: researchers. Contemporary Physics, 54(2):135–136, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2013:APE

[3052] Robert Appleby. Accelerator Physics: Example Problems with Solutions, by William W. MacKay and Mario Conte, Scope: review. Level: undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 54(2):136, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palffy:2013:QPP

[3053] Adriana Palffy. Quantum Principles and Particles, by Walter M. Wilcox, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(2): 136–137, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2013:LPE

[3054] Peter R. Hobson. LHC Physics, edited by T. Binoth, C. Buttar, P. J. Clark and E. W. N. Glover, Scope: conference proceedings. Level: postgraduate, early career researcher. Contemporary Physics, 54(2):137– 138, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2013:EEE

[3055] Aniruddha Chakraborty. Essentials of Electromagnetics for Engineering, by David A. De Wolf, Scope: review. Level: undergraduate. Contemporary Physics, 54(2):138, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:QDW

[3056] Peter J. Bussey. The Quantum Divide: Why Schrödinger’s Cat is either Dead or Alive, by Christopher C. Gerry and Kimberley M. Bruno, Scope: general interest. Level: non-specialists, teacher. Contemporary Physics, 54(2):138–139, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:GRB

[3057] B. Ishak. Gamma-Ray Bursts, by C. Kouveliotou, R. A. M. J. Wijers and S. Woosley, Scope: monograph. Level: Ph.D.s, post-docs and specialists. Contemporary Physics, 54(2):139, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2013:PDL

[3058] Peter R. T. Munro. Principles of Diffuse Light Propagation: Light Propagation in Tissues with Applications in Biology and Medicine, by Jorge Ripoll Lorenzo, Scope: textbook. Level: postgraduate and specialist. Contemporary Physics, 54(2):139–140, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2013:NUF

[3059] K. Alan Shore. Nanofabrication Using Focused Ion and Electron Beams: Principles and Applications, edited by Ivo Utke, Stanisislav Moshkalev and Phillip Russell, Scope: monograph. Level: researcher. Contemporary Physics, 54(2):140–141, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kawano:2013:TWT

[3060] Yukio Kawano. Terahertz waves: a tool for condensed matter, the life sciences and astronomy. Contemporary Physics, 54(3):143–165, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2013:DSC

[3061] David Bailin. Discrete symmetries and CP violation, Scope: textbook. Level: postgraduate and early career researcher. Contemporary Physics, 54(3):166–168, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2013:BNE

[3062] Andrew Resnick. Biomedical Nanosensors, edited by Joseph Irudayaraj, Scope: edited book, reference, review. Level: postgraduate, advanced undergraduate, early career researcher, researcher, scientist, engineers.

Contemporary Physics, 54(3):169, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2013:MTP

[3063] Andrew Resnick. Modern Trends in Physics Research. Third International Conference on Modern Trends in Physics Research. Cairo, Egypt 6–10 April 2008, edited by Lotfia M. El Nadi, Scope: conference proceedings. Level: postgraduate, advanced undergraduate, researcher, scientist. Contemporary Physics, 54(3):169–170, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2013:TPC

[3064] Miguel A. F. Sanjuán. Thermal Physics: Concepts and Practice, by Allen L. Wasserman, Scope: textbook. Level: advanced undergraduate, graduate students, lecturers. Contemporary Physics, 54(3):170–171, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nikolopoulos:2013:QCM

[3065] Lampros Nikolopoulos. Quantum Control of Molecular Processes, (Second, Revised and Enlarged Edition), by M. Shapiro and P. Brumer, Scope: review. Level: postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 54(3):171–172, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2013:CPM

[3066] Andrew Resnick. The Cambridge Photographic Moon Atlas, by Alan Chu, Wolfgang Paech, Mario Weigand, translated by Storm Dunlop, Scope: general interest. Level: general readership, non-specialists, preuniversity, undergraduate, advanced undergraduate, teacher. Contemporary Physics, 54(3):172, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2013:IBE

[3067] Chih-Yueh Wang. From Interacting Binaries to Exoplanets (IAU S282): Essential Modelling Tools, edited by M. T. Richards and I. Hubeny. Scope: conference proceedings. Level: postgraduate, researcher, specialist, scientist. Contemporary Physics, 54(3):172–173, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:EHT

[3068] Manuel Vogel. Essentials of Heat Transfer: Principles, Materials, and Applications, by Massoud Kaviany, Scope: textbook. Level: undergraduates, advanced undergraduates. Contemporary Physics, 54(3):173–174, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:FAJ

[3069] Manuel Vogel. Fundamentals of Astrometry, by Jean Kovalevsky and P. Kenneth Seidelmann, Scope: monograph. Level: undergraduates and above. Contemporary Physics, 54(3):174–175, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:SOC

[3070] Thomas Peters. Self-Organised Criticality: Theory, Models and Characterisation, by Gunnar Pruessner, Scope: monograph. Level: researcher. Contemporary Physics, 54(3):175, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2013:EPA

[3071] Tom Lancaster. Einstein’s Physics: Atoms, Quanta, and Relativity — Derived, Explained, and Appraised, by Ta-Pei Cheng, Scope: monograph. Level: researcher. Contemporary Physics, 54(3):175–176, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Boyd:2013:BSP

[3072] T. J. M. Boyd. Basic Space Plasma Physics, by Wolfgang Baumjohann and Rudolf A. Treumann, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 54(3):176–177, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2013:GTH

[3073] David Bailin. Group Theory for High Energy Physicists, by Mohammad Saleem and Muhammad Rafique, Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 54(3):177–178, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:NCJ

[3074] B. Ishak. Neutrino Cosmology, by J. Lesgourgues, G. Mangano, G. Miele and S. Pastor, Scope: monograph. Level: early career researcher, researcher, specialist. Contemporary Physics, 54(3):178–179, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2013:BPT

[3075] R. Ellis. Binary Paths to Type Ia Supernovae Explosions, edited by Rosanne Di Stefano, Marina Orio and Maxwell Moe, Scope: conference proceedings. Level: researcher. Contemporary Physics, 54(3):179–180, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2013:AQM

[3076] Charles Downing. Advanced Quantum Mechanics: A Practical Guide, by Yuli V. Nazarov and Jeroen Danon, Scope: textbook. Level: postgraduate. Contemporary Physics, 54(3):180, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Huelga:2013:VQB

[3077] S. F. Huelga and M. B. Plenio. Vibrations, quanta and biology. Contemporary Physics, 54(4):181–207, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2013:ERA

[3078] David J. Miller. Essay review: The anatomy of a paradigm-shift. The Continental Drift Controversy, by Henry R. Frankel. ISBN 978-0-521-87507-3. Scope: monograph. Level: accessible and useful to all levels of reader, from general interest to specialist. Contemporary Physics, 54(4): 208–210, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2013:QFT

[3079] David Bailin. Quantum Field Theories in Two Dimensions: Collected Works of Alexei Zamolodchikov (in 2 Volumes), edited by Alexander Belavin, Yaroslav Pugai and Alexander Zamolodchikov, Scope: edited book. Level: researcher. Contemporary Physics, 54(4):211–212, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ma:2013:BSP

[3080] Angie Ma. Biophysics: Searching for Principles, by William Bialek, Scope: textbook. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 54(4):212–213, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:FGU

[3081] B. Ishak. The First Galaxies in the Universe, by Abraham Loeb and Steven R. Furlanetto, Scope: textbook. Level: advanced undergraduate, postgraduate, early career researcher, specialist. Contemporary Physics, 54(4):213–214, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:AAM

[3082] Manuel Vogel. Astrometry for Astrophysics: Methods, Models, and Applications, edited by William F. van Altena, Scope: edited book. Level: undergraduate, advanced undergraduate. Contemporary Physics, 54(4): 214, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2013:HOC

[3083] Mark Fox. The Handbook on Optical Constants of Metals: In Tables and Figures & The Handbook on Optical Constants of Semiconductors: In Tables and Figures, by Sadao Adachi, Scope: reference. Level: researcher. Contemporary Physics, 54(4):215, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2013:AHS

[3084] Nikolaos E. Myridis. Artificial Human Sensors: Science and Applications, by Peter Wide, Scope: monograph. Level: general readership, advanced undergraduate, postgraduate,researcher, teacher, engineers. Contemporary Physics, 54(4):215–216, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:OME

[3085] Manuel Vogel. Our Magnetic Earth: The Science of Geomagnetism, by Ronald T. Merrill, Scope: monograph. Level: general readership. Contemporary Physics, 54(4):216–217, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2013:TIQ

[3086] Matt Probert. The Transactional Interpretation of Quantum Mechanics: The Reality of Possibility, by Ruth E. Kastner, Scope: reference. Level: specialist. Contemporary Physics, 54(4):217–218, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2013:AML

[3087] S. C. Tiwari. The Angular Momentum of Light, by D. L. Andrews and M. Babiker, Scope: reference, edited book. Level: advanced undergraduate, researcher,teacher, specialist. Contemporary Physics, 54(4):218– 219, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2013:LRS

[3088] Derry W. Jones. The Long Road to Stockholm, The Story of MRI: An Autobiography, by Peter Mansfield, Scope: autobiography. Level: scientist, general reader. Contemporary Physics, 54(4):219–220, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:QCP

[3089] Peter J. Bussey. Quantum Concepts in Physics: An Alternative Approach to the Understanding of Quantum Mechanics, by Malcolm Longair, Scope: textbook. Level: teacher, specialist. Contemporary Physics, 54(4):220–221, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:NMB

[3090] Thomas Peters. Nonequilibrium Many-Body Theory of Quantum Systems: A Modern Introduction, by G. Stefanucci and R. van Leeuwen, Scope: textbook. Level: postgraduate. Contemporary Physics, 54(4): 221, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:OHP

[3091] Peter J. Bussey. The Oxford Handbook of Philosophy of Physics, edited by R. Batterman, Scope: review, edited book. Level: postgraduate, researcher, specialist, scientist. Contemporary Physics, 54(4):221–223, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eberly:2013:RMP

[3092] Joseph Eberly. Review of Modern Physics, by Paul A. Tipler and Ralph A. Llewellyn, Scope: reference, textbook. Level: undergraduate, advanced undergraduate, teacher. Contemporary Physics, 54(4):223–224, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2013:INS

[3093] Robert D. Joseph. Isaac Newton’s Scientific Method: Turning Data into Evidence about Gravity and Cosmology, by William L. Harper. Scope: review. Level: specialist. Contemporary Physics, 54(4):224–225, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:2013:RJA

[3094] John Quenby. Relativistic Jets from Active Galactic Nuclei, edited by M. Boetttcher, D. E. Harris, and H. Krawczynski, Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist. Contemporary Physics, 54(4):225–226, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2013:SAH

[3095] Marco Casazza. Synergetic Agents, by Hermann Haken and Paul Levi, Scope: reference, monograph, manual, textbook. Level: non-specialists, postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 54(4):226–227, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2013:CME

[3096] Amgad Salama. Continuum Mechanics in the Earth Sciences, by William I. Newman, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 54(4):227, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benhamou:2013:TIF

[3097] Mabrouk Benhamou. Theory of Intermolecular Forces, by Anthony Stone, Scope: textbook. Level: postgraduate, researcher. Contemporary Physics, 54(4):227–228, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2013:CTE

[3098] D. C. Barber. Correction Techniques in Emission Tomography, edited by Mohammad Dawood, Xiaoyi Jiang and Klaus Schäfers, Scope: review. Level: postgraduate, early career researcher, specialist. Contemporary Physics, 54(4):228–229, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2013:FEU

[3099] M. Jamil. Fusion, the Energy of the Universe, by Garry McCracken and Peter Stott, Scope: monograph. Level: postgraduate, advanced undergraduate,early career researcher, researcher. Contemporary Physics, 54 (4):229–230, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Matravers:2013:BPM

[3100] David R. Matravers. The Ballet of the Planets: A Mathematician’s Musings on the Elegance of Planetary Motion, by Donald C. Benson, Scope: general interest. Level: pre-university, undergraduate. Contemporary Physics, 54(4):230–231, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yung:2013:IGE

[3101] K. C. Yung, W. M. Wu, M. P. Pierpoint, and F. V. Kusmartsev. Introduction to graphene electronics — a new era of digital transistors and devices. Contemporary Physics, 54(5):233–251, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dombey:2013:ERO

[3102] Norman Dombey. Essay review: Oppenheimer, the bomb and beyond, Scope: biography. Level: general readership [Review of Inside the Centre: The Life of J. Robert Oppenheimer, by Ray Monk, London, Jonathon Cape, 2012, 818 pp., £30.00 (hardback), ISBN 978-0-224-06262-6]. Contemporary Physics, 54(5):252–253, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Phillips:2013:PDM

[3103] Philip Phillips. The Physics of Dilute Magnetic Alloys, by Jun Kondo, translated by Shigeru Koikegami, Kosuke Odagiri, Kunihiko Yamaji and Takashi Yanagisawa, Scope: monograph. Level: postgraduate. Contemporary Physics, 54(5):254, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2013:BPE

[3104] Peter J. Dobson. Biomimetics in Photonics, edited by Olaf Karthaus, Scope: review. Level: general readership, non-specialists, undergraduate, postgraduate, scientist. Contemporary Physics, 54(5):254–255, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2013:PGO

[3105] Mark Fox. A Practical Guide to Optical Metrology for Thin Films, by Michael Quinten, Scope: monograph. Level: postgraduate. Contemporary Physics, 54(5):255, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pounds:2013:TSS

[3106] Ken Pounds. Time, Space, Stars and Man: The Story of the Big Bang, by Michael M. Woolfson. Scope: review. Level: general readership. Contemporary Physics, 54(5):255–256, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2013:IPW

[3107] Frank Berkshire. Introduction to the Physics of Waves, by Tim Freegarde, Scope: textbook. Level: general readership, undergraduate, advanced undergraduate, postgraduate, early career researcher, teacher. Contemporary Physics, 54(5):256–257, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2013:MCQ

[3108] S. W. Lovesey. A Modern Course in the Quantum Theory of Solids, edited by Fuxiang Han, Scope: review. Level: postgraduate. Contemporary Physics, 54(5):257, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2013:IMP

[3109] J. Rogel-Salazar. Introduction to Mathematical Physics: Methods and Concepts, 2nd edn., by Chun Wa Wong, Scope: review. Level: undergraduate. Contemporary Physics, 54(5):257–258, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:MPI

[3110] Manuel Vogel. Modern Physics: An Introductory Text, by Jeremy I. Pfeffer and Shlomo Nir, Scope: monograph. Level: undergraduate. Contemporary Physics, 54(5):258–259, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2013:TPI

[3111] Juan A. Añel. Turing: Pioneer of the Information Age, by B. Jack Copeland, Scope: general interest. Level: general reader. Contemporary Physics, 54(5):259, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2013:MUP

[3112] Robert Appleby. Measurement Uncertainty and Probability, by Robin Willink, Scope: textbook. Level: specialist. Contemporary Physics, 54 (5):259–260, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2013:AAE

[3113] B. Ishak and J. McFarland. Astrophysics from Antarctica, edited by Michael G. Burton, Xiangqun Cui and Nicholas F. H. Tothill, Scope: wide-ranging review. Level: postgraduates and specialists. Contemporary Physics, 54(5):260–261, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2013:EQM

[3114] S. W. Lovesey. Exploring Quantum Mechanics: A Collection of 700+ Solved Problems for Students, Lecturers, and Researchers, by Victor Galitski, Boris Karnakov, Vladimir Kogan, and Victor Galitski, Jr., Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 54 (5):261, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2013:EME

[3115] Michael Belsley. Electricity and Magnetism, by Edward M. Purcell and David J. Morin, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(5):261–262, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bucknall:2013:EOP

[3116] David Bucknall. Electronic and Optical Properties of Conjugated Polymers, 2nd edn., International Series of Monographs on Physics 159, by William Barford, Scope: review. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 54(5): 262–263, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Greenland:2013:TPR

[3117] P. T. Greenland. Terahertz Physics, by R. A. Lewis, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(5):263, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2013:CME

[3118] A. H. Harker. Computational Methods for Electromagnetic Phenomena: Electrostatics in Solvation, Scattering and Electron Transport, by Cai Wei, Scope: monograph. Level: postgraduate, researcher. Contemporary Physics, 54(5):263–264, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2013:EMP

[3119] Andrew Resnick. Encyclopaedia of Medical Physics (2 Volume Set), edited by Slavik Tabakov, Franco Milano, Sven-Erik Strand, Cornelius Lewis and Perry Sprawls, Scope: reference, encyclopedia. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 54(5): 264–265, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2013:RVS

[3120] Derry W. Jones. Radioactivity: A Very Short Introduction, by Claudio Tuniz, Scope: general interest. Level: general reader, preferably with physical science background. Contemporary Physics, 54(5):265– 266, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2013:RTC

[3121] A. J. Leggett. Revolutions in Twentieth-Century Physics, by David J. Griffiths, Scope: general interest. Level: general readership, nonspecialists, pre-university. Contemporary Physics, 54(5):266, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2013:LMI

[3122] David L. Andrews. Light–Matter Interaction: Physics and Engineering at the Nanoscale, by John Weiner and Frederico Nunes, Scope: textbook. Level: undergraduate, advanced undergraduate, scientist, engineers. Contemporary Physics, 54(5):267–268, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coffey:2013:PIS

[3123] William T. Coffey. Path Integrals for Stochastic Processes: An Introduction, by Horacio S. Wio, Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 54(5):268, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2013:FTC

[3124] S. J. Blundell. Field Theories of Condensed Matter Physics, 2nd edn., by Eduardo Fradkin, Scope: monograph. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(5):268–269, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Osborn:2013:INP

[3125] H. Osborn. An Introduction to Non-Perturbative Foundations of Quantum Field Theory, by Franco Strocchi, Scope: monograph. Level: scientist. Contemporary Physics, 54(5):269, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holland:2013:GTG

[3126] Peter Holland. Gauge Theories of Gravitation: A Reader with Commentaries, edited by Milutin Blagojević and Friedrich W. Hehl, Scope: edited book, reference, review. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(5):269–270, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2013:ACS

[3127] Matt Probert. Atomistic Computer Simulations: A Practical Guide, by V. Brázdová and D. R. Bowler, Scope: reference. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(5): 270–271, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leslie:2013:IPP

[3128] Dawn E. Leslie. The Infinity Puzzle: The Personalities, Politics and Extraordinary Science Behind the Higgs Boson, by Frank Close. Scope: general interest. Level: general readership, non-specialists, pre-university, undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 54(5):271, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:PSO

[3129] Manuel Vogel. Planetary Systems and the Origins of Life, Cambridge Astrobiology (No. 3), edited by Ralph Pudritz, Paul Higgs and Jonathon Stone, Scope: edited book. Level: general readership. Contemporary Physics, 54(5):271–272, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McCall:2013:TOC

[3130] Martin McCall. Transformation optics and cloaking. Contemporary Physics, 54(6):273–286, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2013:CCC

[3131] Derry W. Jones. Crystallographic centenary celebration: the emergence of X-ray analysis, Scope: general interest, monograph. Level: general reader, specialist. Contemporary Physics, 54(6):287–290, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rowlands:2013:ISP

[3132] George Rowlands. An Introduction to Stochastic Processes and Nonequilibrium Statistical Physics, revised edition, Series on Advances in Statistical Mechanics: Volume 19, edited by Horacio S. Wio, Roberto R. Deza and Juan M. López, Scope: review. Level: scientist. Contemporary Physics, 54(6):291, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2013:TQW

[3133] Peter J. Bussey. The Theory of the Quantum World: Proceedings of the 25th Solvay Conference on Physics, edited by David Gross, Marc Henneaux and Alexander Sevrin, Scope: conference proceedings. Level: postgraduate. Contemporary Physics, 54(6):291–292, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:EEH

[3134] Manuel Vogel. Earth: Evolution of a Habitable World, 2nd edn., by J. I. Lunine, Scope: monograph. Level: general readership, non-specialists, pre-university. Contemporary Physics, 54(6):292, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Krotov:2013:FMB

[3135] Sergey Krotov. First Magnitude: A Book of the Bright Sky, by James B. Kaler, Scope: edited book, general interest. Level: general readership, pre-university, undergraduate. Contemporary Physics, 54(6):293, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sumner:2013:DMN

[3136] T. J. Sumner. Dark Matter, Neutrinos and Our Solar System, Scope: monograph. Level: general readership, postgraduate, researcher. Contemporary Physics, 54(6):293–294, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2013:STS

[3137] Francisco Caruso. On Space and Time, Scope: edited book. Level: general readership, non-specialists, pre-university, undergraduate, advanced undergraduate, postgraduate, early career research, research, teacher, specialist. Contemporary Physics, 54(6):294–295, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2013:EPG

[3138] Gunter M. Schütz. Exercises in Probability: A Guided Tour from Measure Theory to Random Processes, via Conditioning, 2nd edn., by L. Chaumont and M. Yor, Scope: guide. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(6):295–296, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2013:ODS

[3139] Salvatore Esposito. Order and Disorder: Science Essentials for the Non-Scientist, by Myron Kaufman, Scope: general interest. Level: general readership, teacher. Contemporary Physics, 54(6):296–297, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2013:CPI

[3140] Peter R. T. Munro. Computational Photonics: An Introduction With MATLAB, by Marek S. Wartak, Scope: textbook. Level: undergraduate, postgraduate, practising scientists and engineers. Contemporary Physics, 54(6):297, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2013:SPF

[3141] K. Alan Shore. Silicon Photonics: Fundamentals and Devices, by M. Jamal Deen and P. K. Basu, Scope: textbook. Level: postgraduate, early career researcher, researcher, teacher. Contemporary Physics, 54(6):297– 298, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2013:SMM

[3142] Aniruddha Chakraborty. Simple Models of Magnetism, Oxford Graduate Texts, by Ralph Skomski, Scope: review. Level: advanced undergraduate. Contemporary Physics, 54(6):298–299, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khare:2013:TCT

[3143] Prashant Khare. Ten Chapters in Turbulence, edited by Peter A. Davidson, Yukio Kaneda and Katepalli R. Sreenivasan, Scope: edited book, reference, review. Level: postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 54(6):299–300, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2013:TMP

[3144] Charles A. Downing. Topics in Modern Physics: Theoretical Foundations, by John Dirk Walecka, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 54(6):300, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2013:HMV

[3145] Nikolaos E. Myridis. Handbook of 3D Machine Vision: Optical Metrology and Imaging, Series in Optics and Optoelectronics, edited by Song Zhang, Scope: edited book. Level: postgraduate, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 54(6):300–301, 2013. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2013:QCL

[3146] K. Alan Shore. Quantum Cascade Lasers, by Jérôme Faist, Scope: monograph. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 54(6):301–302, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tarleton:2013:DMS

[3147] Edmund Tarleton. Dislocations, Mesoscale Simulations and Plastic Flow, Oxford Series on Materials Modelling 5, by Ladislas Kubin, Scope: textbook. Level: postgraduate, researcher. Contemporary Physics, 54(6): 302–303, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2013:EMT

[3148] K. Alan Shore. Electromechanics and MEMS, by Thomas B. Jones and Nenad G. Nenadic, Scope: textbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist. Contemporary Physics, 54(6):303, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2013:TIE

[3149] Richard Shorter. TESLA: Inventor of the Electrical Age, by W. Bernard Carlson, Scope: biography. Level: general readership. Contemporary Physics, 54(6):303–304, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chakraborty:2013:WWR

[3150] Aniruddha Chakraborty. The Wonderful World of Relativity: A Precise Guide for the General Reader, by Andrew M. Steane, Scope: review.

Level: non-specialists. Contemporary Physics, 54(6):304, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2013:NVE

[3151] Richard Shorter. Night Vision: Exploring the Infrared Universe, by Michael Rowan-Robinson, Scope: general interest, textbook. Level: general readership, advanced undergraduate, teacher. Contemporary Physics, 54(6):304–305, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2013:LDS

[3152] R. Ellis. The Life and Death of Stars, by Kenneth R. Lang, Scope: monograph. Level: general readership, non-specialists, pre-university, undergraduate. Contemporary Physics, 54(6):305–306, 2013. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2013:GSE

[3153] Manuel Vogel. Gravitation and Spacetime, 3rd edn., by Hans C. Ohanian and Remo Ruffini, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 54(6):306, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:2013:SMC

[3154] L. M. Brown. Strengthening Mechanisms in Crystal Plasticity, by Ali Argon, Scope: review. Level: scientist, engineers. Contemporary Physics, 54(6):306–307, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2013:EGN

[3155] Thomas Peters. Einstein Gravity in a Nutshell, by A. Zee, Scope: textbook. Level: undergraduate. Contemporary Physics, 54(6):307–308, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riis:2013:OMS

[3156] Erling Riis. Optical Magnetometry, Scope: reference. Level: postgraduate. Contemporary Physics, 54(6):308–309, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2013:QPL

[3157] David L. Andrews. Quantum Physics and Linguistics: A Compositional, Diagrammatic Discourse, Scope: edited book, monograph. Level: postgraduate, early career researcher, researcher, specialist. Contemporary Physics, 54(6):309, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2013:EB

[3158] Anonymous. Editorial board. Contemporary Physics, 54(6):ebi, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fernandez-Garcia:2014:DCN

[3159] Roberto Fernández-García, Yannick Sonnefraud, Antonio I. Fernández-Domínguez, Vincenzo Giannini, and Stefan A. Maier. Design considerations for near-field enhancement in optical antennas. Contemporary Physics, 55(1):1–11, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knuth:2014:IBP

[3160] Kevin H. Knuth. Information-based physics: an observer-centric foundation. Contemporary Physics, 55(1):12–32, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harter:2014:CAI

[3161] A. Härter and J. Hecker Denschlag. Cold atom–ion experiments in hybrid traps. Contemporary Physics, 55(1):33–45, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2014:QDS

[3162] G. J. Milburn. Quantum Dissipative Systems, (Fourth edition), by Ulrich Weiss, Scope: monograph. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 55(1):46–47, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:NLH

[3163] D. G. C. Jones. Nature Loves to Hide: Quantum Physics and the Nature of Reality, a Western Perspective (Revised Edition), by Shimon Malin, Scope: review. Level: non-specialists, undergraduate. Contemporary Physics, 55(1):47, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2014:TTW

[3164] K. P. Zetie. Time Travel and Warp Drives: A Scientific Guide to Shortcuts through Time and Space, by Allen Everett and Thomas Roman, Scope: general interest. Level: undergraduate, pre-university, teacher, scientist. Contemporary Physics, 55(1):47–48, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2014:GOA

[3165] Mark Fox. Generalised Optomechanics and its Applications: Quantum Optical Properties of Generalised Optomechanical System, by Jin-Jin Li and Ka-Di Zhu, Scope: monograph. Level: researcher. Contemporary Physics, 55(1):48, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2014:MEA

[3166] Thomas Peters. Modern Electrodynamics, by Andrew Zangwill, Scope: textbook. Level: postgraduate. Contemporary Physics, 55(1):49, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2014:PVM

[3167] Peter J. Bussey. A Physicist’s View of Matter and Mind, by Chandre Dharma-wardana, Scope: general interest. Level: scientist. Contemporary Physics, 55(1):49–50, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2014:CCH

[3168] Richard Ellis. The Cosmic Century: A History of Astrophysics and Cosmology, by Malcolm S. Longair, Scope: monograph, reference, review. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher. Contemporary Physics, 55(1):50–51, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2014:PDE

[3169] Tom Lancaster. Polarons, by David Emin, Scope: monograph. Level: postgraduate, researcher. Contemporary Physics, 55(1):51, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2014:PMK

[3170] Manuel Vogel. Physical Mathematics, by Kevin Cahill, Scope: reference, text book. Level: advanced undergraduate, researcher. Contemporary Physics, 55(1):51–52, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2014:YSP

[3171] Peter J. Bussey. 100 Years of Subatomic Physics, edited by Ernest M. Henley and Stephen D. Ellis, Scope: edited book. Level: specialist. Contemporary Physics, 55(1):52–53, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2014:MTF

[3172] Andrew Resnick. Mathematical Topics in Fluid Mechanics — Volumes 1 & 2, by Pierre-Louis Lions, Scope: monograph, review. Level: postgraduate, researcher, specialist, scientist. Contemporary Physics, 55(1): 53, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Owen:2014:HAR

[3173] Hywel Owen, Ranald MacKay, Ken Peach, and Susan Smith. Hadron accelerators for radiotherapy. Contemporary Physics, 55(2):55–74, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2014:NGM

[3174] L. Wang and S. Gai. The next generation mass storage devices — physical principles and current status. Contemporary Physics, 55(2):75–93, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fedorov:2014:SME

[3175] M. V. Fedorov and N. I. Miklin. Schmidt modes and entanglement. Contemporary Physics, 55(2):94–109, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burton:2014:AER

[3176] David A. Burton and Adam Noble. Aspects of electromagnetic radiation reaction in strong fields. Contemporary Physics, 55(2):110–121, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2014:UNR

[3177] A. H. Harker. Understanding the Nanotechnology Revolution, by Edward L. Wolf and Manasa Medikonda, Scope: general interest. Level: general readership. Contemporary Physics, 55(2):122, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Virmani:2014:ISM

[3178] S. Virmani. An Introduction to Statistical Mechanics and Thermodynamics, by Robert H. Swendsen, Scope: text book. Level: non-specialists. undergraduate. postgraduate. advanced undergraduate. teacher. scientist. engineers. Contemporary Physics, 55(2):122–123, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2014:ONE

[3179] Peter J. Dobson. Optochemical Nanosensors, edited by Andrea Cusano, Francisco J. Arregui, Michele Giordano and Antonello Cutolo, Scope: monograph, reference. Level: postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 55(2): 123–124, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2014:IQI

[3180] Michael Belsley. Introduction to Quantum Information Science, by Vlatko Vedral, Scope: text book. Level: advanced undergraduate, postgraduate, early career researcher. Contemporary Physics, 55(2):124–125, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2014:CMP

[3181] A. H. Harker. Computational Methods for Physics, by Joel Franklin, Scope: text book. Level: undergraduate. Contemporary Physics, 55(2): 125, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2014:STS

[3182] Peter V. E. McClintock. Space, Time, & Stuff, by Frank Arntzenius, Scope: monograph. Level: postgraduate students, researchers. Contemporary Physics, 55(2):126, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2014:MSM

[3183] J. Rogel-Salazar. Modern Statistical Methods for Astronomy With R Applications, by Eric D. Feigelson and G. Jogesh Babu, Scope: reference. Level: advanced undergraduate, graduate, specialised. Contemporary Physics, 55(2):126–127, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Huan:2014:SCE

[3184] Chao Huan. Spin Current, edited by Sadamichi Maekawa, Sergio O. Valenzuela, Eiji Saitoh, and Takashi Kimura, Scope: monograph, review. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 55(2): 127–128, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stewart:2014:INR

[3185] John M. Stewart. Introduction to 3 + 1 Numerical Relativity, by Miguel Alcubierre, Scope: review. Level: advanced undergraduate, postgraduate, early career researcher. Contemporary Physics, 55(2):128–129, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2014:SBD

[3186] Peter V. E. McClintock. Simple Brownian Diffusion, by Daniel Thomas Gillespie and Effrosyni Seitaridou, Scope: textbook. Level: senior undergraduate students, postgraduates, researchers. Contemporary Physics, 55(2):129, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2014:NSV

[3187] Peter V. E. McClintock. Novel Superfluids: Vol. 1, edited by K. H. Bennemann and J. B. Ketterson, Scope: monograph. Level: postgraduate students, researchers. Contemporary Physics, 55(2):130–132, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2014:SDA

[3188] J. Rogel-Salazar. Statistical Data Analysis for the Physical Sciences, by Adrian Bevan, Scope: textbook. Level: undergraduate. Contemporary Physics, 55(2):132–133, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McOrist:2014:GTM

[3189] Jock McOrist. Geometric and Topological Methods for Quantum Field Theory, edited by Alexander Cardona, Iván Contreras and Andrés F. Reyes-Lega, Scope: conference proceedings, review. Level: postgraduate student, early career researcher, researcher. Contemporary Physics, 55 (2):133, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2014:FQG

[3190] B. Ishak. Foundations of Quantum Gravity, by James Lindesay, Scope: textbook. level: specialist. Contemporary Physics, 55(2):134, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stewart:2014:AGR

[3191] John M. Stewart. Advanced General Relativity: Gravity Waves, Spinning Particles and Black Holes, by Claude Barrabès and Peter A. Hogan, Scope: textbook, review. Level: advanced undergraduate, postgraduate, early career researcher. Contemporary Physics, 55(2):134–135, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2014:HOG

[3192] Salvatore Esposito. Herschel 400 Observing Guide, by Stephen James O’Meara, Scope: general interest. Level: general readership. Contemporary Physics, 55(2):135, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2014:FMP

[3193] C. A. Downing. Foundation Mathematics for the Physical Sciences, by K. F. Riley and M. P. Hobson, Scope: textbook, solution manual. Level: first year undergraduate. Contemporary Physics, 55(2):136, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2014:NEC

[3194] Peter J. Dobson. Nanotechnology for the Energy Challenge, edited by Javier Garcia-Martinez, Scope: monograph, reference, review. Level: undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 55(2):136–137, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2014:SRS

[3195] Peter J. Bussey. Science, Religion and the Search for Extraterrestrial Intelligence, by David Wilkinson, Scope: general interest. Level: general readership. Contemporary Physics, 55(2):137–138, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:USC

[3196] D. W. Jones. Uncorked: The Science of Champagne, Revised ed. by Gerard Liger-Belair, Scope: general interest, monograph. Level: postgraduate, advanced undergraduate, researcher, scientist. Contemporary Physics, 55(2):138–139, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2014:BSU

[3197] Eric Sheldon. Behind the Scenes of the Universe: From the Higgs to Dark Matter, by Gianfranco Bertone. Scope: popular survey, general interest. Level: general readership, undergraduate, teacher, scientist. Contemporary Physics, 55(2):139–140, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sambles:2014:TLE

[3198] J. R. Sambles. Transmission Lines: Equivalent Circuits, Electromagnetic Theory and Photons, by Richard Collier, Scope: review. Level: undergraduates. Contemporary Physics, 55(2):140–141, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2014:OWI

[3199] Rob Appleby. Oscillations and Waves: An Introduction, by Richard Fitzpatrick, Scope: textbook. Level: undergraduate. Contemporary Physics, 55(2):141, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2014:EMV

[3200] S. W. Lovesey. Electricity and Magnetism, Volume 1: Third Edition, Electricity and Magnetism, Volume 2: Third Edition, by B. I. Bleaney and B. Bleaney, Scope: textbook, level: undergraduate. Contemporary Physics, 55(2):141–142, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2014:WPM

[3201] D. C. Barber. Webb’s Physics of Medical Imaging, (Second Edition), edited by M. Flower, Scope: textbook, review. Level: postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 55(2): 142–143, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2014:CDG

[3202] S. W. Lovesey. Complex Dynamics of Glass-forming Liquids: A Modecoupling Theory, by Wolfgang Götze, Scope: review. Level: general readership. Contemporary Physics, 55(2):143–144, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2014:CMO

[3203] Mario Bertolotti. Cosmic Masers — From OH to Ho (IAU S287), edited by Roy S. Booth, Wouter H. T. Vlemmings and Elizabeth M. L. Humphreys. Scope: edited book. Level: specialist. Contemporary Physics, 55(2):144–145, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2014:IGD

[3204] Peter R. Hobson. Imaging Gaseous Detectors and their Applications, by Eugenio Nappi and Vladimir Peskov, Scope: monograph. Level: postgraduate, researcher. Contemporary Physics, 55(2):145, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Holland:2014:PFC

[3205] Peter Holland. Physical Foundations of Continuum Mechanics, by A. Ian Murdoch, Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 55(2):145–146, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2014:NPE

[3206] D. J. Miller. Nuclear Physics: Exploring the Heart of Matter, by National Research Council, Scope: report, review. Level: administrators of nuclear physics research programmes, politicians, interested non-experts, physics students looking for an overview, postgraduate, researcher. Contemporary Physics, 55(2):146–147, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duxbury:2014:LBM

[3207] Geoffrey Duxbury. Laser-Based Measurements for Time and Frequency Domain Applications: A Handbook, by Pasquale Maddaloni, Marco Bellini, and Paolo De Natale, Scope: monograph, reference, review. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 55(2):147–148, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2014:HAP

[3208] David J. Miller. Handbook of Accelerator Physics and Engineering, 2nd edition, by Alexander Wu Chao, Karl Hubert Mes, Maury Tigner and Frank Zimmermann, Scope: handbook. Level: postgraduate, researcher. Contemporary Physics, 55(2):148, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Professor:2014:FSS

[3209] S. J. Blundell Professor. Frustrated Spin Systems, (2nd ed.), edited by H. T. Diep, Scope: monograph. Level: researcher. Contemporary Physics, 55(2):148–149, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Marnham:2014:QIT

[3210] Lachlan Marnham. Quantum Information Theory and the Foundations of Quantum Mechanics, by Christopher G. Timpson, Scope: review, text book. Level: non-specialists, postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 55(2):149, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2014:MCP

[3211] Juan A. Añel. Mesoscale-Convective Processes in the Atmosphere, by Robert J. Trapp, Scope: textbook. Level: undergraduate, postgraduate, advanced undergraduate, researcher, teacher. Contemporary Physics, 55 (2):149–150, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2014:EMM

[3212] Charles Downing. Essential Mathematical Methods for the Physical Sciences & Student Solution Manual for Essential Mathematical Methods for the Physical Sciences, by K. F. Riley and M. P. Hobson, Scope: textbook, solution manual. Level: undergraduate. Contemporary Physics, 55 (2):150–151, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Professor:2014:GMB

[3213] N. Riley Professor. Granular Media; Between Fluid and Solid, by Bruno Andreotti, Yoe̋l Forterre and Olivier Pouliquen, Scope: monograph, reference. Level: postgraduate, researcher and specialist. Contemporary Physics, 55(2):151–152, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Professor:2014:SMM

[3214] S. J. Blundell Professor. Statistical Mechanics of Magnetic Excitations, From Spin Waves to Stripes and Checkerboards, by Enrico Rastelli, Scope: textbook. Level: postgraduate. Contemporary Physics, 55(2): 152–153, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2014:CAN

[3215] Manuel Vogel. The Cosmos: Astronomy in the New Millennium, 4th edition, by Jay M. Pasachoff and Alex Filippenko, Scope: general interest, text book. Level: pre-university, undergraduate. Contemporary Physics, 55(2):153, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Osborn:2014:GTP

[3216] H. Osborn. Gauge Theories in Particle Physics: A Practical Introduction, Fourth Edition, Volumes 1 & 2, by Ian Aitchison and Anthony Hey, Scope: textbook. Level: advanced undergraduates, graduates. Contemporary Physics, 55(2):154, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stenholm:2014:MQQ

[3217] Stig Stenholm. Mathematics of Quantization and Quantum Fields, by Jan Dereziński and Christian Gérard, Scope: monograph. Level: specialist. Contemporary Physics, 55(2):154–155, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Osborn:2014:TPP

[3218] H. Osborn. Three-Particle Physics and Dispersion Relation Theory, by A. V. Anisovich, V. V. Anisovich, M. A. Matveev, V. A. Nikonov, J. Nyiri, and A. V. Sarantsev, Scope: monograph. Level: specialists. Contemporary Physics, 55(2):155–156, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2014:STS

[3219] B. Ishak. String Theory and the Scientific Method, by Richard Dawid, Scope: general interest. Level: specialists, scientists. Contemporary Physics, 55(2):156, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palmer:2014:LGP

[3220] T. N. Palmer. Lorenz, Gödel and Penrose: new perspectives on determinism and causality in fundamental physics. Contemporary Physics, 55(3): 157–178, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tsallis:2014:INE

[3221] Constantino Tsallis. An introduction to nonadditive entropies and a thermostatistical approach to inanimate and living matter. Contemporary Physics, 55(3):179–197, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Walker:2014:UST

[3222] Matthew G. Walker and Abraham Loeb. Is the universe simpler than ΛCDM? Contemporary Physics, 55(3):198–211, 2014. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Child:2014:QMM

[3223] M. S. Child. Quantum monodromy and molecular spectroscopy. Contemporary Physics, 55(3):212–221, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:GBA

[3224] Derry W. Jones. The genesis of the Bohr atom, Scope: review. Level: general readership, undergraduate, postgraduate, advanced undergraduate. Contemporary Physics, 55(3):222–225, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2014:HBQ

[3225] Eric Sheldon. The Higgs boson, its quantum-field interaction mechanism and its aftermath. Scope: general interest. Level: general readership, non-specialists, teacher. Contemporary Physics, 55(3):226–230, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dombey:2014:CBH

[3226] Norman Dombey. Churchill’s Bomb: a hidden history of science, war and politics, Scope: history of science. Level: general reader. Contemporary Physics, 55(3):231–234, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Licata:2014:SQO

[3227] Ignazio Licata. In search of the quantum oneness, Scope: review. Level: scientist. Contemporary Physics, 55(3):235–237, 2014. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2014:NAS

[3228] Eric Sheldon. The nuclear astrophysics of star-shine, Scope: textbook. Level: advanced undergraduate, graduate, professional and specialist researcher, early career researcher. Contemporary Physics, 55(3):238–241, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2014:CCS

[3229] Peter J. Bussey. Complexity is complex, Scope: edited book. Level: general reader, specialist. Contemporary Physics, 55(3):242–245, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Marnham:2014:CIQ

[3230] Lachlan Marnham. Conductor–Insulator Quantum Phase Transitions, edited by Vladimir Dobrosavljevic, Nandini Trivedi, and James M. Valles, Jr., Scope: reference, review. Level: postgraduate, specialist, early career researcher, researcher, scientist. Contemporary Physics, 55 (3):246, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2014:PPR

[3231] Mark Nicol. Physical Principles of Remote Sensing, (3rd Edition), by W. G. Rees, Scope: handbook, text book. Level: undergraduate, postgraduate, advanced undergraduate, teacher, engineers. Contemporary Physics, 55(3):246–247, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2014:URI

[3232] Nikolaos E. Myridis. Ultra-realistic Imaging: Advanced Techniques in Analogue and Digital Colour Holography, by Hans Bjelkhagen and David Brotherton-Ratcliffe, Scope: handbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 55(3):247–248, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khare:2014:PCM

[3233] Prashant Khare. Physics of Continuous Media: Problems and Solutions in Electromagnetism, Fluid Mechanics and MHD, by Grigory Vekstein, Scope: reference, text book. Level: postgraduate, advanced undergraduate, early career researcher, teacher, engineers. Contemporary Physics, 55(3):248–249, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2014:PSS

[3234] Mark Nicol. Planetary Science: The Science of Planets around Stars, Second Edition, by George H. A. Cole and Michael M. Woolfson, Scope: general interest, text book. Level: general readership, non-specialists, pre-university, undergraduate, postgraduate, advanced undergraduate, scientist, engineers. Contemporary Physics, 55(3):249–250, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2014:AME

[3235] Tom Lancaster. Advanced Mechanics: From Euler’s Determinism to Arnold’s Chaos, by S. G. Rajeev, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 55(3):250, 2014. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Spurgeon:2014:ENS

[3236] Steven R. Spurgeon. Experimental Neutron Scattering, by B. T. M. Willis and C. J. Carlile, Scope: text book. Level: undergraduate. Contemporary Physics, 55(3):250–251, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2014:MPP

[3237] Peter J. Bussey. Modern Particle Physics, by Mark Thomson, Scope: textbook. Level: postgraduate, teacher. Contemporary Physics, 55(3): 251–252, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2014:FPS

[3238] Thomas Peters. Fundamental Planetary Science: Physics, Chemistry and Habitability, by Jack J. Lissauer and Imke de Pater, Scope: textbook, level: undergraduate. Contemporary Physics, 55(3):252, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salamin:2014:SKA

[3239] Y. I. Salamin. Schrödinger’s Killer App: Race to Build the World’s First Quantum Computer, by Jonathan P. Dowling, Scope: general interest. Level: general readership. Contemporary Physics, 55(3):252–254, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pierri-Galvao:2014:BCP

[3240] Dr. Monica Pierri-Galvao. Basic Concepts in Physics: From the Cosmos to Quarks, by Masud Chaichian, Hugo Perez Rojas and Anca Tureanu, Scope: review. Level: undergraduate. Contemporary Physics, 55 (3):254, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bengtsson:2014:TFS

[3241] Ingemar Bengtsson. On the Topology and Future Stability of the Universe, by Hans Ringström, Scope: monograph. Level: researcher. Contemporary Physics, 55(3):254–255, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2014:EPP

[3242] Manuel Vogel. Extreme physics: properties and behaviour of matter at extreme conditions, by Jeff Colvin and Jon Larsen, Scope: text book.

Level: advanced undergraduate. Contemporary Physics, 55(3):256, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2014:UUI

[3243] Mark Nicol. Understanding the Universe: An Inquiry Approach to Astronomy and the Nature of Scientific Research, by George Greenstein, Scope: text book. Level: pre-university, undergraduate, teacher. Contemporary Physics, 55(3):256–257, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:TSE

[3244] D. W. Jones. A Tale of Seven Elements, by Eric Scerri, Scope: accessible monograph. Level: undergraduate, general readership. Contemporary Physics, 55(3):257–258, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2014:APP

[3245] Mark Fox. Atomic Physics: Precise Measurements and Ultracold Matter, by Massimo Inguscio and Leonardo Fallani, Scope: monograph. Level: postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 55(3):258–259, 2014. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2014:HBV

[3246] Peter R. T. Munro. Handbook of Biophotonics: 3 Volume Set, edited by Jürgen Popp, Valery V. Tuchin, Arthur Chiou and Stefan H. Heinemann, Scope: handbook, reference. Level: specialist, scientist, engineers. Contemporary Physics, 55(3):259, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2014:PCC

[3247] Marco Casazza. The Principle of the Common Cause, by Gábor Hofer-Szabó, Miklós Rédei and Lázló E. Szabó, Scope: monograph. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 55(3):260, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2014:EPE

[3248] Thomas Peters. Econophysics and Physical Economics, by Peter Richmond, Jürgen Mimkes and Stefan Hutzler, Scope: text book. Level: advanced undergraduate. Contemporary Physics, 55(3):260–261, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Edmunds:2014:AMM

[3249] M. G. Edmunds. The Antikythera mechanism and the mechanical universe. Contemporary Physics, 55(4):263–285, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Battista:2014:SPS

[3250] F. Battista. Single particle sources and quantum heat fluctuations. Contemporary Physics, 55(4):286–301, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ropars:2014:SPS

[3251] Guy Ropars, Vasudevan Lakshminarayanan, and Albert Le Floch. The sunstone and polarised skylight: ancient Viking navigational tools? Contemporary Physics, 55(4):302–317, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Biondo:2014:MMB

[3252] A. E. Biondo, A. Pluchino, and A. Rapisarda. Micro and macro benefits of random investments in financial markets. Contemporary Physics, 55 (4):318–334, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:SRS

[3253] Derry W. Jones. Serving the Reich: The Struggle for the Soul of Physics under Hitler, by Philip Ball, Scope: general interest. Level: general readership. Contemporary Physics, 55(4):335–336, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:CGH

[3254] Derry W. Jones. Crystals in Glass: A Hidden Beauty, by E. D. Zanotto, Scope: general interest. Level: general readership. Contemporary Physics, 55(4):336–337, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2014:EPN

[3255] K. Alan Shore. Electronic Processes in Non-crystalline Materials, (Second Edition), by N. F. Mott and E. A. Davis, Scope: text book. Level: advanced undergraduate. Contemporary Physics, 55(4):337, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2014:SCD

[3256] J. Rogel-Salazar. Stochastic Calculus and Differential Equations for Physics and Finance, by Joseph L. McCauley, Scope: review. Level: postgraduate, researcher, specialist. Contemporary Physics, 55(4):338, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fox:2014:QPS

[3257] Mark Fox. Quantum Processes in Semiconductors, (Fifth Edition), by Brian K. Ridley, Scope: text book. Level: postgraduate. Contemporary Physics, 55(4):338–339, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khare:2014:SGE

[3258] Prashant Khare. A Student’s Guide to Entropy, by Don S. Lemons, Scope: textbook. Level: undergraduate, early stage postgraduate. Contemporary Physics, 55(4):339, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2014:SIL

[3259] Nikolaos E. Myridis. Scientific Inference: Learning from Data, by Simon Vaughan, Scope: textbook. Level: undergraduate, advanced undergraduate, researcher, teacher, scientist, engineers. Contemporary Physics, 55 (4):339–340, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2014:PQM

[3260] Manuel Vogel. The Physics of Quantum Mechanics, by James Binney and David Skinner, Scope: text book. Level: undergraduate, advanced undergraduate. Contemporary Physics, 55(4):340–341, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2014:PRS

[3261] Peter J. Bussey. The Physics of Reality: Space, Time, Matter, Cosmos, edited by Richard L. Amoroso, Louis H. Kauffman and Peter Rowlands, Scope: conference proceedings. Level: specialist. Contemporary Physics, 55(4):341–342, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:HRG

[3262] D. G. C. Jones. How the Ray Gun Got Its Zap: Odd Excursions into Optics, by Stephen R. Wilk, Scope: review. Level: general readership, non-specialists, undergraduate. Contemporary Physics, 55(4):342, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2014:AVS

[3263] Peter V. E. McClintock. Astrobiology: A Very Short Introduction, by David C. Catling, Scope: monograph. Level: general readership. Contemporary Physics, 55(4):342–344, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2014:FDG

[3264] K. Alan Shore. Functional Differential Geometry, by Gerald Jay Sussman and Jack Wisdom with Will Farr, Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 55(4):344, 2014. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2014:IGB

[3265] K. Alan Shore. Introduction to Graphene-Based Nanomaterials: From Electronic Structure to Quantum Transport, by Luis E. F. Foa Torres, Stephan Roche and Jean-Christophe Charlier, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 55(4): 344–345, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2014:PEA

[3266] B. Ishak. The Physics and Evolution of Active Galactic Nuclei, by Hagai Netzer, Scope: monograph. Level: specialists. Contemporary Physics, 55 (4):345, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2014:NSP

[3267] Chih-Yueh Wang. Neutron Stars and Pulsars (IAU S291): Challenges and Opportunities after 80 Years, Proceedings of the International Astronomical Union (IAU S279), edited by Joeri van Leeuwen., Scope: conference proceedings. Level: researcher, scientist. Contemporary Physics, 55(4):346–347, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2014:DMC

[3268] B. Ishak. Dark Matter and Cosmic Web Story, by Jaan Einasto, Scope: monograph. Level: general readership, specialists. Contemporary Physics, 55(4):347, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2014:PSG

[3269] S. W. Lovesey. Perspectives on Spin Glasses, by Pierluigi Contucci and Cristian Giardinà, Scope: reference. Level: postgraduate. Contemporary Physics, 55(4):347–348, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2014:MPE

[3270] Robert D. Joseph. Magnificent Principia: exploring Isaac Newton’s masterpiece, by Colin Pask, Scope: general interest, monograph. Level: general readership, undergraduate, postgraduate, advanced undergraduate, teacher. Contemporary Physics, 55(4):348–349, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nikolopoulos:2014:CPE

[3271] Lampros Nikolopoulos. Computational Physics, 2nd Edition, by Jos Thijssen, Scope: review. Level: general readership. Contemporary Physics, 55(4):349–350, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Phelps:2014:WOP

[3272] Alan D. R. Phelps. Waves and Oscillations in Plasmas, by Hans L. Pécseli, Scope: text book. Level: undergraduate, postgraduate, advanced undergraduate, early career researcher. Contemporary Physics, 55(4): 350, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:2014:BHN

[3273] Ian D. Lawrie. Black Holes — New Horizons, edited by Sean Hayward, Scope: edited book. Level: specialist. Contemporary Physics, 55(4): 350–351, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Licata:2014:ESP

[3274] Ignazio Licata. Energy Science, Principles, Technologies and Impacts: Second Edition, by John Andrews and Nick Jelley, Scope: review. Level: general readership, undergraduate, early career researcher, teacher. Contemporary Physics, 55(4):351–352, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2014:IMN

[3275] Peter J. Dobson. Introduction to Metal-nanoparticle Plasmonics, by Matthew Pelton and Garnett W. Bryant, Scope: review. Level: postgraduate, advanced undergraduate, researcher, scientist, engineers. Contemporary Physics, 55(4):352–353, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Phelps:2014:PRL

[3276] Alan D. R. Phelps. Plasma Research at the Limit: From the International Space Station to Applications on Earth, (with DVD-ROM), by Gregor Morfill, Yuri Baturin, and Vladimir Fortov. Scope: monograph. Level: general readership. Contemporary Physics, 55(4):353, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2014:TIT

[3277] S. W. Lovesey. Topological Insulators and Topological Superconductors, by B. Andrei Bernevig with Taylor L. Hughes, Scope: text book. Level: advanced undergraduate. Contemporary Physics, 55(4):353–354, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stewart:2014:RHL

[3278] John M. Stewart. Relativistic Hydrodynamics, by Luciano Rezzolla and Olindo Zanotti, Scope: monograph, text book. Level: postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 55(4):354–355, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2014:GNP

[3279] A. H. Harker. Graphene: A New Paradigm in Condensed Matter and Device Physics, by E. L. Wolf, Scope: monograph. Level: postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 55(4):355, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2014:PTM

[3280] A. H. Harker. Phase Transitions: Modern Applications, (2nd edition), by Moshe Gitterman, Scope: text book. Level: postgraduate. Contemporary Physics, 55(4):355–356, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2014:JCM

[3281] Richard Shorter. James Clerk Maxwell: Perspectives on His Life and Work, edited by Raymond Flood, Mark McCartney, and Andrew Whitaker, Scope: Review. Level: general readership. Contemporary Physics, 55(4):356, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2014:QMG

[3282] Thomas Peters. Quantum Mechanics, by Gennaro Auletta, Mauro Fortunato and Giorgio Parisi, Scope: text book. Level: advanced undergraduate. Contemporary Physics, 55(4):357, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2014:HBB

[3283] Peter V. E. McClintock. How Big is Big and How Small is Small: The Sizes of Everything and Why, by Timothy Paul Smith, Scope: monograph. Level: general readership, non-specialists, scientists, engineers. Contemporary Physics, 55(4):357–358, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2014:SOW

[3284] Manuel Vogel. The Science of Ocean Waves: Ripples, Tsunamis and Stormy Seas, by J. B. Zirker, Scope: general interest. Level: general readership. Contemporary Physics, 55(4):358–359, 2014. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2014:IME

[3285] Richard Shorter. An Introduction to Mechanics, (2nd Edition), by Daniel Kleppner and Robert Kolenkow, Scope: review. Level: pre-university, undergraduate, teacher. Contemporary Physics, 55(4):359, 2014. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2014:ATS

[3286] Robert C. Smith. Astrophysical Techniques, Sixth Edition, by C. R. Kitchin, Scope: general interest, text book. Level: general readership, postgraduate, advanced undergraduate, teacher. Contemporary Physics, 55(4):359–360, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2014:QWM

[3287] D. G. C. Jones. The Quantum World, by Michel Le Bellac and Alain Aspect, Scope: review. Level: advanced undergraduate. Contemporary Physics, 55(4):360–361, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2014:EB

[3288] Anonymous. Editorial board. Contemporary Physics, 55(4):ebi, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knight:2015:EIY

[3289] Peter Knight. Editorial — international year of light. Contemporary Physics, 56(1):1, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berry:2015:NOO

[3290] M. V. Berry. Nature’s optics and our understanding of light. Contemporary Physics, 56(1):2–16, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Treu:2015:GLE

[3291] Tommaso Treu and Richard S. Ellis. Gravitational lensing: Einstein’s unfinished symphony. Contemporary Physics, 56(1):17–34, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duff:2015:HFF

[3292] M. J. Duff. How fundamental are fundamental constants? Contemporary Physics, 56(1):35–47, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Neukirch:2015:NOO

[3293] Levi P. Neukirch and A. Nick Vamivakas. Nano-optomechanics with optically levitated nanoparticles. Contemporary Physics, 56(1):48–62, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Thompson:2015:ICC

[3294] Richard C. Thompson. Ion Coulomb crystals. Contemporary Physics, 56 (1):63–79, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2015:MEM

[3295] David Bailin. Mathematics is everything, maybe, Scope: general interest. Level: general readership. Contemporary Physics, 56(1):80–84, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coffey:2015:CTQ

[3296] William T. Coffey. A Concise Treatise on Quantum Mechanics in Phase Space, Scope: review. Level: postgraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 56(1):85– 87, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ironside:2015:CPI

[3297] Charlie Ironside. Computational Photonics: An Introduction with MAT-LAB, by Marek S. Wartak, Scope: textbook. Level: undergraduate. Contemporary Physics, 56(1):88, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2015:HOH

[3298] Mario Bertolotti. High-order Harmonic Generation in Laser Plasma Plumes, by Rashid Ganeev, Scope: review. Level: early career researcher, researcher, teacher, specialist. Contemporary Physics, 56(1):88–89, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2015:MWC

[3299] Mark Nicol. Madame Wu Chien-Shiung: The First Lady of Physics Research, by Tsai-Chien Chiang/translated by Tang-Fong Wong, Scope: review. Level: general readership. Contemporary Physics, 56(1):89–90, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2015:IDR

[3300] Chih-Yueh Wang. Intelligible Design: A Realistic Approach to the Philosophy and History of Science, by Julio A. Gonzalo and Manuel M. Carreira, Scope: general interest, monograph. Level: general readership, researcher, scientist, engineers. Contemporary Physics, 56(1):90, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benedict:2015:PNT

[3301] Keith A. Benedict. The Physics of Nanoelectronics: Transport and Fluctuation Phenomena at Low Temperatures, by Tero T. Heikkilä, Scope: text book. Level: postgraduate, advanced undergraduate. Contemporary Physics, 56(1):90–91, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2015:ERS

[3302] Salvatore Esposito. Einstein Relatively Simple: Our Universe Revealed in Everyday Language, by Ira Mark Egdall, Scope: general interest. Level: general readership. Contemporary Physics, 56(1):91, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2015:IMQ

[3303] Tom Lancaster. An Introduction to Macroscopic Quantum Phenomena and Quantum Dissipation, A. O. Caldeira, Scope: text book. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 56 (1):92, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2015:CMM

[3304] Mark Nicol. A Course in Mathematical Methods for Physicists, Scope: review. Level: undergraduate. Contemporary Physics, 56(1):92–93, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Osborn:2015:SFP

[3305] H. Osborn. Symmetry and Fundamental Physics: Tom Kibble at 80, edited by Jerome Gauntlett, Scope: review. Level: scientist. Contemporary Physics, 56(1):93–94, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2015:QSS

[3306] Manuel Vogel. Quantum Social Science, by Emmanuel Haven and Andrei Khrennikov, Scope: monograph. Level: postgraduate, early career researcher, researcher, teacher. Contemporary Physics, 56(1):94, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2015:FPE

[3307] Nikolaos E. Myridis. Fundamentals of Picoscience, edited by Klaus D. Sattler, Scope: edited book. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 56(1):95, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dombey:2015:POW

[3308] Norman Dombey. Physics and Our World: Reissue of the Proceedings of a Symposium in Honor of Victor F Weisskopf, Scope: history of science. Level: general reader. Contemporary Physics, 56(1):95–96, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:PCC

[3309] B. Ishak. Physics and Chemistry of Circumstellar Dust Shells, by Hans-Peter Gail and Erwin Sedlmayr, Scope: text book. Level: specialist. Contemporary Physics, 56(1):97, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Osborn:2015:SRB

[3310] H. Osborn. Symmetry Relationships between Crystal Structures: Applications of Crystallographic Group Theory in Crystal Chemistry, by Ulrich Müller, Scope: monograph. Level: postgraduate, specialist. Contemporary Physics, 56(1):97–98, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pikovski:2015:BEQ

[3311] Alexander Pikovski. Back-of-the-envelope Quantum Mechanics: With Extensions to Many-body Systems and Integrable PDEs, by Maxim Olshanii, Scope: textbook/supplementary text. Level: undergraduate and postgraduate. Contemporary Physics, 56(1):98, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2015:OHH

[3312] Bob Joseph. The Oxford Handbook of the history of physics, edited by Jed Z. Buchwald and Robert Fox, Scope: reference. Level: general readership. Contemporary Physics, 56(1):98–99, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2015:YBK

[3313] S. J. Blundell. 40 Years of Berezinskii–Kosterlitz–Thouless Theory, edited by Jorge V. José, Scope: monograph. Level: specialist. Contemporary Physics, 56(1):99–100, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2015:OHP

[3314] J. Rogel-Salazar. One Hundred Physics Visualisations Using Matlab, (With DVD-ROM), by Dan Green. Scope: Review. Level: Undergraduate. Contemporary Physics, 56(1):100, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:EPJ

[3315] B. Ishak. Essential Physics, by John Matolyak and Ajawad Haija, Scope: textbook. Level: pre-university. Contemporary Physics, 56(1):101, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2015:QFT

[3316] Thomas Peters. Quantum Field Theory and the Standard Model, by Matthew D. Schwartz, Scope: text book. Level: postgraduate. Contemporary Physics, 56(1):101–102, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:IEP

[3317] B. Ishak. Introduction to Elementary Particle Physics, (2nd Edition), by Alessandro Bettini, Scope: text book. Level: undergraduate, teacher.

Contemporary Physics, 56(1):102, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2015:PHS

[3318] David L. Andrews. From Photons to Higgs: A Story of Light, (2nd Edition), by Moo-Young Han. Scope: general interest. Level: non-specialists, postgraduate, advanced undergraduate, teacher, scientist. Contemporary Physics, 56(1):102–103, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2015:CLA

[3319] K. Alan Shore. Ceramic Lasers, by Akio Ikesue, Yan Lin Aung, and Voicu Lupei, Scope: monograph. Level: researcher. Contemporary Physics, 56 (1):103–104, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2015:PMQ

[3320] Peter J. Bussey. Probing the Meaning of Quantum Mechanics: Physical, Philosophical and Logical Perspectives, edited by Diederik Aerts, Sven Aerts and Christian de Ronde, Scope: edited book. Level: researcher. Contemporary Physics, 56(1):104, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hull:2015:THP

[3321] Stephen Hull. Techniques in High Pressure Neutron Scattering, by Stefan Klotz, Scope: text book. Level: postgraduate. Contemporary Physics, 56(1):104–105, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowan:2015:NDF

[3322] Brian Cowan. Nuclear Dawn: F. E. Simon and the race for Atomic Weapons in World War II, by Kenneth D. McRae, Scope: review. Level: general readership. Contemporary Physics, 56(1):105, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2015:LEO

[3323] Peter R. Hobson. Lasers and Electro-Optics: Fundamentals and Engineering, (2nd Edition), by Christopher C. Davis, Scope: text book. Level: undergraduate, advanced undergraduate, engineers. Contemporary Physics, 56(1):105–106, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2015:Ca

[3324] Anonymous. Corrigendum. Contemporary Physics, 56(1):107, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2015:Cb

[3325] Anonymous. Corrigendum. Contemporary Physics, 56(1):108, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Al-Khalili:2015:AOM

[3326] Jim Al-Khalili. Advances in optics in the medieval Islamic world. Contemporary Physics, 56(2):109–122, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Weisenburger:2015:LMO

[3327] Siegfried Weisenburger and Vahid Sandoghdar. Light microscopy: an ongoing contemporary revolution. Contemporary Physics, 56(2):123–143, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ettl:2015:IRF

[3328] Svenja Ettl. Introductory review on ‘flying triangulation’: a motionrobust optical 3D measurement principle. Contemporary Physics, 56(2): 144–158, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Erlich:2015:IHQ

[3329] Joshua Erlich. An introduction to holographic QCD for nonspecialists. Contemporary Physics, 56(2):159–171, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schuld:2015:IQM

[3330] Maria Schuld, Ilya Sinayskiy, and Francesco Petruccione. An introduction to quantum machine learning. Contemporary Physics, 56(2):172– 185, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kappeler:2015:CBM

[3331] Natascha Kappeler. Cantilevers for biological monitoring: opportunities for classical and quantum physics. Contemporary Physics, 56(2): 186–201, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:2015:MSS

[3332] John Quenby. Magnetohydrodynamics of the Sun, Scope: monograph, textbook. Level: postgraduate, researcher, specialist. Contemporary Physics, 56(2):202–205, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2015:IWC

[3333] Juan A. Añel. On the importance of weather and climate change for our present and future energy needs, edited by Troccoli, Dubus and Ellen Haupt, Scope: edited book. Level: undergraduate, non-specialists, advanced undergraduate, postgraduate, specialist, scientist, engineers, researcher. Contemporary Physics, 56(2):206–208, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2015:EP

[3334] Chih-Yueh Wang and Yuxiang Peng. Extrasolar planets. Contemporary Physics, 56(2):209–213, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2015:PWA

[3335] Derry W. Jones. The Physics of War: From Arrows to Atoms, Scope: history reviews. Level: first-year science undergraduate or general reader. Contemporary Physics, 56(2):214–219, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tiwari:2015:DRQ

[3336] S. C. Tiwari. Demystifying the riddle of quantum physics. Contemporary Physics, 56(2):220–224, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2015:ERN

[3337] Eric Sheldon. Essay review: Nobel lectures: Physics (2006–2010) [review of Nobel Lectures: Physics (2006–2010), edited by Lars Brink, Singapore, NJ, and London, World Scientific 2014, xii + 371 pp. ISBN 978-981-4612-67-8 (hardback), 978-9-81-461268-5 (paperback), 978-9-81-461270-08 (ebook)]. Contemporary Physics, 56(2):225–232, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2015:LPP

[3338] Manuel Vogel. Light — The Physics of the Photon, by Ole Keller, Scope: monograph. Level: postgraduate, advanced undergraduate. Contemporary Physics, 56(2):233, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2015:ISF

[3339] K. Alan Shore. Isochronous Systems, by Francesco Calogero, Scope: monograph. Level: researcher. Contemporary Physics, 56(2):233–234, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2015:NDS

[3340] Peter V. E. McClintock. Nonautonomous Dynamical Systems in the Life Sciences, edited by Peter E. Kloeden and Christian Pötzsche, Scope: edited book, monograph. Level: postgraduate, early career researcher, researcher, specialist. Contemporary Physics, 56(2):234–235, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:EMP

[3341] B. Ishak. Elements of Mechanics, by P. F. Kelly, Scope: textbook. Level: undergraduate. Contemporary Physics, 56(2):235–236, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2015:AP

[3342] Salvatore Esposito. Auroral physics. Contemporary Physics, 56(2):236, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2015:NSI

[3343] A. H. Harker. Nonlinear Science: An Interactive Mathematica^{T M} Notebook, by David K. Campbell, Sebastian M. Marotta and Thomas A. Tanury, Scope: guide. Level: undergraduate, postgraduate, advanced undergraduate. Contemporary Physics, 56(2):236–237, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2015:SCM

[3344] A. H. Harker. Symmetry of Crystals and Molecules, by Mark Ladd, Scope: textbook. Level: undergraduate, postgraduate, advanced undergraduate. Contemporary Physics, 56(2):237–238, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:2015:QFT

[3345] Alan D. Martin. Quantum Field Theory for the Gifted Amateur, by Tom Lancaster and Stephen J. Blundell, Scope: review, text book. Level: general readership, undergraduate, postgraduate. Contemporary Physics, 56(2):238–239, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2015:BWM

[3346] Gunter M. Schütz. The Bethe Wavefunction, by Michel Gaudin, Scope: monograph. Level: postgraduate, early career researcher, researcher, scientist, specialist. Contemporary Physics, 56(2):239–240, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2015:FCI

[3347] J. Rogel-Salazar. Fractional Calculus: An Introduction for Physicists, (2nd Edition) by Richard Herrmann, Scope: reference. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 56(2):240, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2015:MGD

[3348] Chih-Yueh Wang. Molecular Gas, Dust, and Star Formation in Galaxies (IAU S292), edited by Tony Wong and Jürgen Ott. Scope: conference proceeding. Level: researcher/scientist. Contemporary Physics, 56(2): 240–241, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carter:2015:EDC

[3349] Richard Carter. Engines of Discovery: A Century of Particle Accelerators, (Revised and Expanded Edition), by Andrew Sessler and Edmund Wilson, Scope: general interest. Level: general readership, nonspecialists, pre-university, undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 56(2):241–242, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2015:BDH

[3350] Rob Appleby. Beam Dynamics in High Energy Particle Accelerators, by Andrzej Wolski, Scope: textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 56(2):242–243, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2015:IPA

[3351] Nikolaos E. Myridis. Image processing and acquisition using Python, by Ravishankar Chityala and Sridevi Pudipeddi, Scope: textbook. Level: general readership, undergraduate, teacher. Contemporary Physics, 56 (2):243, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Modi:2015:TEN

[3352] Kavan Modi. Thermal energy at the nanoscale. Contemporary Physics, 56(2):243–244, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2015:HNO

[3353] Peter J. Dobson. Handbook of Nano-optics and Nanophotonics — 2 Volume Set, edited by Motoichi Ohtsu, Scope: reference, monograph, handbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 56(2):244–245, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2015:LAP

[3354] Robert S. MacKay. Local Activity Principle: The Cause of Complexity and Symmetry Breaking, Scope: monograph. Level: general reader. Contemporary Physics, 56(2):245–246, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2015:CVD

[3355] Chandra Wickramasinghe. COMETS! Visitors from Deep Space, by David J. Eicher, Scope: review, general interest. Level: general readership, undergraduate. Contemporary Physics, 56(2):246–247, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Narozhny:2015:ELP

[3356] N. B. Narozhny and A. M. Fedotov. Extreme light physics. Contemporary Physics, 56(3):249–268, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Townsend:2015:RPS

[3357] Peter Townsend. The role of physics in shaping music. Contemporary Physics, 56(3):269–291, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lutken:2015:IRM

[3358] C. A. Lütken. Introduction to the role of modular symmetries in graphene and other two-dimensional materials. Contemporary Physics, 56(3):292– 318, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hanson:2015:NPM

[3359] David E. Hanson and John L. Barber. A new paradigm for the molecular basis of rubber elasticity. Contemporary Physics, 56(3):319–337, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Valverde:2015:PFU

[3360] Jose Manuel Valverde. Pattern-formation under acoustic driving forces. Contemporary Physics, 56(3):338–358, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nazarenko:2015:WT

[3361] Sergey Nazarenko. Wave turbulence. Contemporary Physics, 56(3):359– 373, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2015:FEO

[3362] K. Alan Shore. Fundamentals of Electro-optic Systems Design: Communications, Lidar, and Imaging, Scope: monograph. Level: researcher. Contemporary Physics, 56(3):374–376, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2015:EOP

[3363] A. J. Leggett. Einstein’s Opponents: The Public Controversy about the Theory of Relativity in the 1920s, Scope: general interest, monograph. Level: general readership. Contemporary Physics, 56(3):377–378, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:SEP

[3364] B. Ishak. Stellar Evolution Physics: Volume 2: Advanced Evolution of Single Stars, Scope: monograph. Level: advanced undergraduates and specialists. Contemporary Physics, 56(3):379–380, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bishop:2015:FQO

[3365] Raymond F. Bishop. Foundations of quantum optics: a basis for the new quantum technologies, Scope: textbook. Level: postgraduate, earlycareer researcher, researcher. Contemporary Physics, 56(3):381–384, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khare:2015:FDP

[3366] Prashant Khare. Fluid Dynamics: Part 1: Classical Fluid Dynamics, Scope: textbook. Level: postgraduate, advanced undergraduate, early career researcher, scientist, engineers. Contemporary Physics, 56(3):385– 387, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2015:OAE

[3367] Peter R. T. Munro. Optical Antennas, edited by Mario Agio and Andrea Alù, Scope: handbook, reference. Level: postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 56(3):388, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2015:QTN

[3368] David L. Andrews. The Quantum Theory of Nonlinear Optics, by Peter D. Drummond and Mark Hillery, Scope: reference, monograph, textbook. Level: postgraduate, researcher, scientist. Contemporary Physics, 56(3):388–389, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2015:PTT

[3369] D. G. C. Jones. Photonics: technology for the Twenty First Century?, Scope: monograph. Level: undergraduate, postgraduate. Contemporary Physics, 56(3):389–392, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eberly:2015:NDW

[3370] Joseph Eberly. Non-diffracting Waves, edited by Hugo E. Hernandez-Figueroa, Michel Zamboni-Rached, and Erasmo Recami, Scope: edited book. Level: early career researcher. Contemporary Physics, 56(3):392, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Carter:2015:RAS

[3371] Richard Carter. Reviews of Accelerator Science and Technology: Volume 5 — Applications of Superconducting Technology to Accelerators, edited by Alexander W. Chao and Weiren Chou, Scope: review. Level: specialist, scientists, engineers. Contemporary Physics, 56(3):392–393, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2015:SLS

[3372] Manuel Vogel. Sea-level Science: Understanding Tides, Surges, Tsunamis and Mean Sea-level Changes, by David Pugh and Philip Woodworth, Scope: monograph. Level: advanced undergraduate. Contemporary Physics, 56(3):394, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowan:2015:TDM

[3373] Brian Cowan. Translational Dynamics and Magnetic Resonance: Principles of Pulsed Gradient Spin Echo NMR, by Paul T. Callaghan, Scope: monograph. Level: researcher. Contemporary Physics, 56(3):394–395, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knight:2015:QMM

[3374] Peter Knight. Quantum Machines: Measurement and Control of Engineered Quantum Systems, edited by Michel Devoret, Benjamin Huard, Robert Schoelkopf, and Leticia F. Cugliandolo, Scope: monograph. Level: postgraduate. Contemporary Physics, 56(3):395–396, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2015:NHB

[3375] Derry W. Jones. The Nature of the Hydrogen Bond: Outline of a Comprehensive Hydrogen Bond Theory, by Gastone Gilli and Paola Gilli, Scope: general interest. Level: scientist. Contemporary Physics, 56(3): 396–397, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2015:HIT

[3376] Thomas J. Bridges. Homogeneous, Isotropic Turbulence: Phenomenology, Renormalization, and Statistical Closures, by W. David McComb, Scope: monograph. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 56(3):397–398, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:DGE

[3377] B. Ishak. Dynamics of Galaxies, (2nd Edition), by Giuseppe Bertin, Scope: textbook. Level: postgraduate. Contemporary Physics, 56(3): 398–399, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wickramasinghe:2015:AUE

[3378] Chandra Wickramasinghe. Alien universe: Extraterrestrial Life in Our Minds and in the Cosmos, by Don Lincoln, Scope: general interest. Level: general readership, undergraduate. Contemporary Physics, 56(3):399– 400, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2015:TPT

[3379] Manuel Vogel. Time in Powers of Ten: Natural Phenomena and Their Timescales, by Gerard ’t Hooft and Stefan Vandoren, translated by Saskia Eisberg-’t Hooft, Scope: general interest. Level: general readership. Contemporary Physics, 56(3):400–401, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2015:SSC

[3380] Miguel A. F. Sanjuán. From Strange Simplicity to Complex Familiarity: A Treatise on Matter, Information, Life and Thought, by Manfred Eigen, Scope: monograph. Level: advanced undergraduate. Contemporary Physics, 56(3):401–402, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2015:WPT

[3381] Derry W. Jones. Waves and Particles: Two Essays on Fundamental Physics, by Roger G. Newton, Scope: historical review. Level: general readership. Contemporary Physics, 56(3):402, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Linton:2015:ESP

[3382] Chris Linton. Electromagnetic Scattering by Particles and Particle Groups: An Introduction, Scope: monograph, textbook. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 56(3): 402–403, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2015:WDB

[3383] Peter J. Bussey. What Does a Black Hole Look Like?, by Charles D. Bailyn, Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 56(3):403–404, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burns:2015:TMF

[3384] Stephen J. Burns. The Theory of Materials Failure, by Richard M. Christensen, Scope: monograph. Level: postgraduate, advanced undergraduates. Contemporary Physics, 56(3):404, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2015:Cc

[3385] Anonymous. Corrigendum. Contemporary Physics, 56(3):405, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eberly:2015:SWS

[3386] J. H. Eberly. Shimony–Wolf states and hidden coherences in classical light. Contemporary Physics, 56(4):407–416, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Umstadter:2015:ALD

[3387] Donald P. Umstadter. All-laser-driven Thomson X-ray sources. Contemporary Physics, 56(4):417–431, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wizinowich:2015:AOA

[3388] Peter Wizinowich. Adaptive optics in astronomy. Contemporary Physics, 56(4):432–450, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whyntie:2015:CDP

[3389] T. Whyntie, H. Bithray, J. Cook, A. Coupe, D. Eddy, R. L. Fickling, J. McKenna, B. Parker, A. Paul, and N. Shearer. CERN@school: demonstrating physics with the timepix detector. Contemporary Physics, 56(4): 451–467, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moss:2015:HBC

[3390] Ian G. Moss. Higgs boson cosmology. Contemporary Physics, 56(4): 468–476, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2015:NOP

[3391] K. Alan Shore. Nonlinear optics and photonics, Scope: monograph, textbook. Level: postgraduate, early career researcher, researcher, teacher, specialist. Contemporary Physics, 56(4):477–479, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kourkoumelis:2015:SBS

[3392] Nikolaos Kourkoumelis. Spectroscopy for biosciences, Scope: edited book. Level: specialist. Contemporary Physics, 56(4):480–482, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2015:QPL

[3393] David L. Andrews. Quantum Physics of Light and Matter: A Modern Introduction to Photons, Atoms and Many-Body Systems, by Luca Salasnich, Scope: reference, textbook. Level: general readership, undergraduate, advanced undergraduate. Contemporary Physics, 56(4):483, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:PRP

[3394] B. Ishak. Planetary rings: a post-equinox view, Scope: monograph. Level: general readership, undergraduate and specialist. Contemporary Physics, 56(4):484–486, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2015:LSB

[3395] David L. Andrews. Laser Spectroscopy 1: Basic Principles & Laser Spectroscopy 2: Experimental Techniques (5th edition), Scope: handbook, monograph, reference, text book. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 56(4):487–488, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobrynin:2015:SLM

[3396] A. N. Dobrynin and S. W. Lovesey. Shedding light on matter, Scope: textbook. Level: postgraduate. Contemporary Physics, 56(4):489–492, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2015:IPL

[3397] Giovanni Volpe. An Introduction to Practical Laboratory Optics, by J. F. James, Scope: monograph, manual, guide. Level: early career researcher, researcher, specialist. Contemporary Physics, 56(4):493–495, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2015:LCE

[3398] Tom Lancaster. Lectures on Classical Electrodynamics, by Berthold-Georg Englert, Scope: textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist. Contemporary Physics, 56(4):496, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2015:PNR

[3399] A. J. Leggett. Physics and Necessity: Rationalist Pursuits from the Cartesian Past to the Quantum Present, Scope: monograph. Level: undergraduate/specialist. Contemporary Physics, 56(4):497–499, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2015:NSS

[3400] B. Ishak. Nearest Star: The Surprising Science of our Sun, (2nd Edition), by Leon Golub and Jay M. Pasachoff, Scope: general interest. Level: general readership. Contemporary Physics, 56(4):500, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lovesey:2015:SMQ

[3401] S. W. Lovesey. Solution Manual for Quantum Mechanics (2nd edition), by Ahmed Ishtiaq, Fayyazuddin and Riazuddin, Scope: manual. Level: undergraduate. Contemporary Physics, 56(4):501, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2015:NPM

[3402] Amgad Salama. Novel Porous Media Formulation for Multiphase Flow Conservation Equations, by William T Sha, Scope: monograph. Level: postgraduate, researcher, specialist. Contemporary Physics, 56(4):501– 502, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2015:LHC

[3403] Peter J. Bussey. The Large Hadron Collider: The Extraordinary Story of the Higgs Boson and Other Stuff that will Blow Your Mind, by Don Lincoln. Scope: general interest. Level: general reader. Contemporary Physics, 56(4):502, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2015:MSA

[3404] K. Alan Shore. Modern Small Antennas, by Kyohei Fujimoto and Hisashi Morishita, Scope: monograph. Level: specialist, engineers, early career researchers. Contemporary Physics, 56(4):502–503, 2015. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2015:QAH

[3405] Derry W. Jones. The Quantum Age: How the Physics of the Very Small Transformed Our Lives, by Brian Clegg, Scope: general interest. Level: pre-university, undergraduate, general readership. Contemporary Physics, 56(4):503–504, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2015:MRS

[3406] Peter J. Bussey. Mind and Reality. The Space–Time Window, Scope: general interest. Level: general reader, researcher. Contemporary Physics, 56(4):504–505, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2015:MIS

[3407] Peter R. Hobson. Measurement Instrumentation and Sensors Handbook Electromagnetic, Optical, Radiation, Chemical, and Biomedical Measurement, by J. G. Webster and H. Eren, Scope: reference, handbook. Level: non-specialists, undergraduate, advanced undergraduate, postgraduate, early career researcher, teacher, scientist, engineers. Contemporary Physics, 56(4):505–506, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2015:NPE

[3408] Peter J. Dobson. Nanoantenna Plasmon-Enhanced Spectroscopies for Biotechnological Applications, edited by Marc Lamy de la Chapelle and Annemarie Pucci, Scope: monograph, reference. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, scientist, engineers. Contemporary Physics, 56(4):506, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2015:LGL

[3409] Andrew Resnick. Lie groups and Lie algebras for physicists, by Ashok Das and Susumu Okubo, Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 56(4):506–507, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2015:NPE

[3410] K. Alan Shore. Nonlinear Physics of Ecosystems, by Ehud Meron, Scope: monograph. Level: specialist, early career researchers. Contemporary Physics, 56(4):507–508, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2015:BHS

[3411] Manuel Vogel. Black Holes — A Student Text, (3rd edition), by Derek Rain and Edwin Thomas, Scope: textbook. Level: postgraduate, advanced undergraduate. Contemporary Physics, 56(4):508, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2015:PCH

[3412] Peter J. Bussey. Proceedings of the Conference in Honour of the 90th Birthday of Freeman Dyson, edited by K. K. Phua, L. C. Kwek, N. P. Chang, and A. H. Chang, Scope: conference proceedings. Level: postgraduate, researcher. Contemporary Physics, 56(4):508–509, 2015. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riley:2015:DDC

[3413] Norman Riley. Double-diffusive Convection, by Timour Radko, Scope: monograph. Level: postgraduate. Contemporary Physics, 56(4):509–510, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2015:CUN

[3414] Miguel A. F. Sanjuán. A Certain Uncertainty. Nature’s Random Ways, by Mark P. Silverman, Scope: monograph. Level: advanced undergraduate, postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 56(4):510, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2015:IPO

[3415] Nigel John Mason. An Introduction to Planets. Ours and Others; From Earth to Exoplanets, by Therese Encrenaz, Scope: guide, general interest. Level: general readership, non-specialists, pre-university, undergraduate, advanced undergraduate. Contemporary Physics, 56(4):511, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kirk:2016:NFB

[3416] A. Kirk. Nuclear fusion: bringing a star down to Earth. Contemporary Physics, 57(1):1–18, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morr:2016:CQC

[3417] Dirk K. Morr. Crossover from quantum to classical transport. Contemporary Physics, 57(1):19–45, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jesacher:2016:SHM

[3418] Alexander Jesacher and Monika Ritsch-Marte. Synthetic holography in microscopy: opportunities arising from advanced wavefront shaping. Contemporary Physics, 57(1):46–59, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jennings:2016:NRC

[3419] David Jennings and Matthew Leifer. No return to classical reality. Contemporary Physics, 57(1):60–82, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Duff:2016:H

[3420] M. J. Duff. M-history without the M. Contemporary Physics, 57(1): 83–85, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Licata:2016:ERA

[3421] Ignazio Licata. Essay review: The architecture of time. Principles of Discrete Time Mechanics, by George Jaroszkiewicz, Cambridge Monographs on Mathematical Physics, New York, NY, Cambridge University Press, 2014, 366 + 14 pp., £85.00 (hardback), ISBN: 978-1-107-03429-7. Scope: monograph. Level: scientist. Contemporary Physics, 57(1):86–89, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:SEP

[3422] B. Ishak. Stellar Evolution Physics — Volume 1. Physical Processes in Stellar Interiors, Scope: monograph. Level: advanced undergraduates and specialists. Contemporary Physics, 57(1):90–92, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2016:ERO

[3423] Peter V. E. McClintock. Essay review: The origins of life on Earth. [The Search for our Cosmic Ancestry, by Chandra Wickramasinghe, Singapore, World Scientific, 2014, 216 pp. £25.00 (paperback), ISBN 978-981-4616-97-3, (hardcover) ISBN 978-981-4616-96-6, (e-book) 978-981-4616-99-7], Scope: review, monograph, general interest. Level: general readership, non-specialists, pre-university, undergraduate, advanced undergraduate, postgraduate, early career researcher. Contemporary Physics, 57(1):93–95, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McGlynn:2016:CSL

[3424] Enda McGlynn. Crystal Symmetry, Lattice Vibrations, and Optical Spectroscopy of Solids: A Group Theoretical Approach, Scope: textbook. Level: postgraduate, early career researcher, researcher, specialist, scientist & engineers. Contemporary Physics, 57(1):96–99, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2016:ERI

[3425] Peter J. Bussey. Essay review: Illumination on information. Information and the Nature of Reality: from Physics to Metaphysics, edited by Paul Davies and Niels Henrik Gregersen, Cambridge, Cambridge University Press, 2014,. 488 pp. £12.99 (paperback). ISBN 978-1-107-68453-9. Scope: edited book. Level: general readership, researcher. Contemporary Physics, 57(1):100–103, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2016:FSF

[3426] Gunter M. Schütz. Faith and science: Friend or foe? Contemporary Physics, 57(1):104–109, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:DSC

[3427] B. Ishak. Deep-sky Companions: The Messier Objects, (2nd Edition), Scope: reference. Level: general readership. Contemporary Physics, 57 (1):110–112, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milonni:2016:LFY

[3428] Peter W. Milonni. Lasers: the first 50 years, Scope: review, general interest. Level: general readership, teacher, undergraduate. Contemporary Physics, 57(1):113–116, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Khare:2016:IFE

[3429] Prashant Khare. Incompressible Flow, 4th Edition, by Ronald L. Panton, Scope: textbook. Level: advanced undergraduate, early stage postgraduate. Contemporary Physics, 57(1):117–118, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ogilvie:2016:DSM

[3430] Michael C. Ogilvie. Dynamics of the standard model, 2nd edition, John F. Donoghue, Eugene Golowich and Barry R. Holstein, Scope: monograph, reference. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 57(1):118–120, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keesing:2016:DCQ

[3431] Richard Keesing. Discrete or Continuous? The Quest for Fundamental Length in Physics, by Amit Hagar, Scope: monograph. Level: advanced undergraduate, postgraduate. early career researcher, researcher. Contemporary Physics, 57(1):120–121, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2016:VMR

[3432] Nikolaos E. Myridis. 4D visualization of matter: recent collected works of Ahmed H. Zewail, Nobel Laureate, by Nikolaos E. Myridis, Scope: anthology (article collection). Level: academics, researchers, postgraduates and specialists in: physical chemistry, biochemistry, chemical biology, biophysics, bioengineering, imaging, structural biology, cancer research. Contemporary Physics, 57(1):122, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:TND

[3433] K. Alan Shore. Towards a new dimension: Exploring the Nanoscale, by Anne Marcovich and Terry Shinn, Scope: monograph. Level: specialist. Contemporary Physics, 57(1):122–124, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:NME

[3434] K. Alan Shore. Nanoscale microwave engineering: Optical control of nanodevices, by Charlotte Tripon-Canseliet and Jean Chazelas, Scope: monograph. Level: specialist. Contemporary Physics, 57(1):124–125, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:APP

[3435] B. Ishak. Astrophysics Processes: The Physics of Astronomical Phenomena, by Hale Bradt, Scope: textbook. Level: postgraduate, advanced undergraduate, teacher. Contemporary Physics, 57(1):125–126, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:IGC

[3436] B. Ishak. An Introduction to Galaxies and Cosmology, (2nd Edition), edited by Mark H. Jones, Robert J. A. Lambourne, and Stephen Serjeant, Scope: textbook. Level: general readership, undergraduate. Contemporary Physics, 57(1):126–128, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:GSE

[3437] B. Ishak. Gravitation and Spacetime, (3rd Edition), by Hans C. Ohanian and Remo Ruffini, Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 57(1):128, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:ATC

[3438] B. Ishak. Astrophysics Through Computation: With Mathematica^{(R)} Support, by Brian Koberlein and David Meisel, Scope: textbook. Level: advanced undergraduates, graduate students. Contemporary Physics, 57 (1):129, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:QTC

[3439] K. Alan Shore. Quantum Transport Calculations for Nanosystems, by Kenji Hirose and Nobuhiko Kobayashi, Scope: monograph. Level: specialist, researcher, early career researcher. Contemporary Physics, 57(1): 129–130, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BPH

[3440] K. Alan Shore. Biomedical Photonic Handbook, (3 Volumes), edited by Tuan Vo-Dinh, Scope: handbook. Level: specialist, engineers, early career researchers. Contemporary Physics, 57(1):130–131, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:HMO

[3441] K. Alan Shore. History of Modern Optics and Optoelectronics Development in China, edited by Fuxi Gan and Shouyun Tian, Scope: monograph. Level: general readership, engineers, early career researchers. Contemporary Physics, 57(1):131–132, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Close:2016:PAB

[3442] Frank Close. Particle Accelerators: From Big Bang Physics to Hadron Therapy, by Ugo Amaldi, Scope: monograph, review, postgraduate. Level: general reader, undergraduate, postgraduate, specialist. Contemporary Physics, 57(1):132–134, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:FSS

[3443] Manuel Vogel. The Formation of the Solar System: Theories Old and New, (2nd Edition), by Michael Woolfson, Scope: monograph, level: general readership. Contemporary Physics, 57(1):134–135, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eberly:2016:BPT

[3444] Lynn E. Eberly and Joseph H. Eberly. Bayesian Probability Theory: Applications in the Physical Science, by W. von der Linden, V. Dose and U. von Toussaint. Scope: review. Level: general readership, non-specialists. Contemporary Physics, 57(1):135–136, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2016:SIS

[3445] Nigel John Mason. The Science of Interstellar, Scope: general interest. Level: general reader, non specialist. Contemporary Physics, 57(1): 136–137, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2016:EMB

[3446] Nigel John Mason. Electrons in Molecules: From Basic Principles to Molecular Electronics, by Jean Pierre Launay and Michel Verdaguer, Scope: reference, textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 57(1): 137–138, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2016:QMT

[3447] G. J. Milburn. Quantum Measurement Theory and its Applications, by Kurt Jacobs, Scope: textbook. Level: postgraduate, early career researcher. Contemporary Physics, 57(1):138, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2016:CBO

[3448] Salvatore Esposito. A Chorus of Bells and Other Scientific Inquiries, by Jeremy Bernstein, Scope: article collection. Level: postgraduate, teacher, scientist. Contemporary Physics, 57(1):139, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2016:ISS

[3449] Nikolaos E. Myridis. An Introduction to Sparse Stochastic Processes, Scope: reference. Level: postgraduate, (early career) researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 57(1):139–140, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:MMU

[3450] A. H. Harker. Materials Modelling using Density Functional Theory: Properties and Predictions, by Giustino Feliciano, Scope: monograph. Level: advanced undergraduate, postgraduate, early stage researcher. Contemporary Physics, 57(1):140–141, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:SBB

[3451] A. H. Harker. Semiconductors: Bonds and Bands, by David K. Ferry, Scope: textbook. Level: postgraduate. Contemporary Physics, 57(1): 141, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:PTM

[3452] A. H. Harker. Phase transitions in materials, Scope: textbook. Level: advanced undergraduate; postgraduate, early stage researcher. Contemporary Physics, 57(1):141–142, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Joseph:2016:UFG

[3453] Robert D. Joseph. The Universal Force: Gravity — Creator of Worlds, Scope: monograph. Level: general reader. Contemporary Physics, 57 (1):142, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2016:MOR

[3454] Peter V. E. McClintock. Mid-Ocean Ridges, Scope: monograph. Level: undergraduates, graduate students, interdisciplinary researchers. Contemporary Physics, 57(1):143, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:UCU

[3455] B. Ishak. Unseen Cosmos: The Universe in Radio, Scope: reference, general interest. Level: general readership. Contemporary Physics, 57 (1):143–144, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:LSH

[3456] B. Ishak. Living with the Stars: how the human body is connected to the life cycles of the Earth, the Planets, and the Stars, Scope: general interest. Level: general readership. Contemporary Physics, 57(1):144, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2016:EVB

[3457] Peter V. E. McClintock. Essentials of in vivo Biomedical Imaging, edited by Simon R. Cherry, Ramsey D. Badawi and Jinyi Qi, Scope: edited book. Level: postgraduates, research scientists. Contemporary Physics, 57(1):144–145, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2016:EDP

[3458] Michael Belsley. Experiments and Demonstrations in Physics Bar-llan Physics Laboratory (second edition), Scope: guide or manual. Level: undergraduate. Contemporary Physics, 57(1):145–146, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:WWN

[3459] B. Ishak. Why we need nuclear power: the environmental case, Michael H. Fox, New York, NY, Oxford University Press, 2014, 320 pp., £22.99 (Hardback), ISBN 978-0-199-34457-4. Scope: general interest. Level: general readership. Contemporary Physics, 57(1):146–147, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2016:SSM

[3460] Peter V. E. McClintock. Superfluid States of Matter, by Boris Egor Nikolay Svistunov, Egor Nikolay Babaev and Nikolay Prokof’ev, Scope: monograph. Level: postgraduates, research scientists. Contemporary Physics, 57(1):147, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2016:RSG

[3461] Eric Sheldon. Relativity: The Special and the General Theory, 100th Anniversary Edition, by Albert Einstein, Scope: general interest, edited book. Level: general readership, non-specialists, pre-university, undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist. Contemporary Physics, 57(1):147–149, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2016:BPP

[3462] Peter J. Bussey. The beauty of physics: patterns, principles and perspectives, Scope: general interest. Level: general reader. Contemporary Physics, 57(1):149, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zhou:2016:RPS

[3463] Hai-Jun Zhou. The rock–paper–scissors game. Contemporary Physics, 57(2):151–163, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Burrage:2016:UAI

[3464] Clare Burrage and Edmund J. Copeland. Using atom interferometry to detect dark energy. Contemporary Physics, 57(2):164–176, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quigg:2016:PPA

[3465] Chris Quigg. Particle physics after the Higgs boson discovery: opportunities for the Large Hadron Collider. Contemporary Physics, 57(2): 177–187, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nakao:2016:PRA

[3466] Hiroya Nakao. Phase reduction approach to synchronisation of nonlinear oscillators. Contemporary Physics, 57(2):188–214, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schippers:2016:PIS

[3467] Stefan Schippers, A. L. David Kilcoyne, Ronald A. Phaneuf, and Alfred Müller. Photoionisation of ions with synchrotron radiation: from ions in space to atoms in cages. Contemporary Physics, 57(2):215–229, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2016:SHF

[3468] Juan A. Añel. The stratosphere: history and future a century after its discovery: A review of The Physics of the Stratosphere, by Goody. Contemporary Physics, 57(2):230–233, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:2016:EMQ

[3469] Ian D. Lawrie. The emergent multiverse — quantum theory according to the Everett interpretation: A review of The Emergent Multiverse — Quantum Theory According to the Everett Interpretation, by David Wallace. Contemporary Physics, 57(2):234–237, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Athorne:2016:FPC

[3470] Chris Athorne. Fundamental principles of classical mechanics: a geometrical perspective: A review of Fundamental Principles of Classical Mechanics: A Geometrical Perspective, by Kai S. Lam. Contemporary Physics, 57(2):238–241, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2016:MDR

[3471] Miguel A. F. Sanjuán. Modern dynamics: A review of Introduction to Modern Dynamics: Chaos, Networks, Space and Time by David D. Nolte. Contemporary Physics, 57(2):242–245, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2016:SAS

[3472] Stephen J. Blundell. The science and art of seeing: A review of The Starry Sky Within: Astronomy and the Reach of the Mind in Victorian Literature, by Anna Henchman. Contemporary Physics, 57(2):246– 249, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sheldon:2016:AER

[3473] Eric Sheldon. Albert Einstein’s relativistic road ‘less traveled’: A review The road to relativity: the history and meaning of Einstein’s ‘The Foundation of General Relativity’, by Albert Einstein (Facsimile German Manuscript) and foreword by John Stachel, with annotations and commentary by Hanoch Gutfreund and Jürgen Renn. Contemporary Physics, 57(2):250–254, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:SVN

[3474] Manuel Vogel. Superconductors, by A. V. Narlikar. Scope: monograph. level: postgraduates, researcher, scientist, engineers. Contemporary Physics, 57(2):255, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:MLS

[3475] Manuel Vogel. Masers and lasers, second edition: an historical approach, by Mario Bertolotti. Scope: monograph. level: advanced undergraduate, postgraduate, researcher. Contemporary Physics, 57(2):255–256, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2016:BRS

[3476] Miguel A. F. Sanjuán. Book review: Simulations of oscillatory systems: with award-winning software, physics of oscillations, edited by Eugene I. Butikov Scope: textbook. Level: undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 57(2):256–257, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:SCS

[3477] B. Ishak. The story of collapsing stars: black holes, naked singularities, and the cosmic play of quantum gravity, by Pankaj S. Joshi. Scope: general interest. level: specialist. Contemporary Physics, 57(2):257, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:AGO

[3478] B. Ishak. An amateur’s guide to observing and imaging the heavens, by Ian Morrison. Scope: guidebook. level: general readership. Contemporary Physics, 57(2):258, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:DEG

[3479] B. Ishak. Dynamics and evolution of galactic nuclei, by David Merritt. Scope: textbook. level: specialist. Contemporary Physics, 57(2): 258, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2016:DMB

[3480] David Bailin. Differential manifolds: a basic approach for experimental physicists, by Paul Baillon. Scope: textbook. level: postgraduate. Contemporary Physics, 57(2):259, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:HSP

[3481] K. Alan Shore. Handbook of silicon photonics, edited by Laurent Vivien and Lorenzo Pavesi. Scope: handbook, monograph, reference. level: researchers, early career researchers, engineers. Contemporary Physics, 57 (2):260–261, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:DQD

[3482] K. Alan Shore. Dynamics of quantum dot lasers: effects of optical feedback and external optical injection, by Christian Otto. Scope: monograph. level: advanced undergraduate, postgraduate, early career researchers, researchers, specialist, engineers. Contemporary Physics, 57 (2):261–262, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:NMP

[3483] K. Alan Shore. Numerical methods in photonics, by Andrei V. Lavrinenko, Jesper Laegsgaard, Niles Gregersen, Frank Schmidt, and Thomas Sondergaard. Scope: textbook, monograph. level: researchers, engineers, early career researchers, advanced undergraduates. Contemporary Physics, 57(2):262, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:HAP

[3484] Manuel Vogel. Handbook of accelerator physics and engineering, by A. W. Chao, K. H. Mess, M. Tigner and F. Zimmermann. Scope: reference, handbook. level: researcher, specialist, scientist, engineers. Contemporary Physics, 57(2):263, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2016:TCP

[3485] Stephen J. Blundell. Tensor calculus for physics, by Dwight E. Neuenschwander. Scope: textbook. level: undergraduate, advanced undergraduate, postgraduate. Contemporary Physics, 57(2):263–264, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2016:SPT

[3486] Nikolaos E. Myridis. Stochastic processes: theory for applications, by Robert G. Gallager. Scope: handbook. level: undergraduate, researcher, teacher. Contemporary Physics, 57(2):264–265, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2016:PBM

[3487] Peter V. E. McClintock. Physics and biology: from molecules to life, edited by Jean François Allemand and Pierre Desbiolles. Scope: edited book. level: general physicists, undergraduates, postgraduates. Contemporary Physics, 57(2):265, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2016:LCN

[3488] Peter J. Dobson. A laboratory course in nanoscience and nanotechnology, by Dr Gerrard Eddy Jai Poinern. Scope: manual, textbook. level: general readership, undergraduate, advanced undergraduate, postgraduate, teacher, scientist, engineer. Contemporary Physics, 57(2):266, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2016:ISD

[3489] Miguel A. F. Sanjuán. An introduction to stochastic dynamics, by Jinqiao Duan. Scope: textbook. level: postgraduate, early career researcher, researcher. Contemporary Physics, 57(2):266–267, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riley:2016:HPB

[3490] N. Riley. Henri Poincaré: a biography through the daily papers, by Jean-Marc Ginoux and Christian Gerini. Scope: monograph. level: postgraduate. Contemporary Physics, 57(2):267–268, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2016:LNF

[3491] Stephen J. Blundell. Lecture notes on field theory in condensed matter physics, by Christopher Mudry. Scope: textbook. level: graduate. Contemporary Physics, 57(2):268, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2016:PYF

[3492] Tom Lancaster. Physics on your feet: Berkeley graduate exam questions, by Dmitry Budker and Alexander O. Sushkov. Scope: textbook. level: advanced undergraduate to specialist (advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist). Contemporary Physics, 57(2):269, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2016:IQM

[3493] Peter J. Dobson. Introductory Quantum Mechanics for Applied Nanotechnology, by Dae Mann Kim. Scope: textbook. level: undergraduate and postgraduate. Contemporary Physics, 57(2):269–270, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:BCD

[3494] A. H. Harker. The basics of crystallography and diffraction, (Fourth Edition), by Christopher Hammond. Scope: textbook. level: undergraduate, postgraduate, early stage researcher. Contemporary Physics, 57 (2):270, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:TKM

[3495] Manuel Vogel. Thermodynamics, kinetics, and microphysics of clouds, by V. I. Khvorostyanov and J. A. Curry. Scope: monograph. level: advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 57(2):271, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:BPO

[3496] A. H. Harker. Bioinspired photonics: optical structures and systems inspired by nature, by Viktoria Greanya. Scope: guide. level: nonspecialists, undergraduates, postgraduates. Contemporary Physics, 57 (2):271–272, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Licata:2016:QFH

[3497] Ignazio Licata. Quantum fractals: from Heisenberg’s uncertainty to Barnsley’s fractality, by Arkadiusz Jadczyk. Scope: monograph. level: postgraduate, researcher. Contemporary Physics, 57(2):272–274, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2016:GGD

[3498] David Bailin. Gauge/gravity duality: foundations and applications, by Martin Ammon and Johanna Erdmenger. Scope: textbook. level: postgraduate. Contemporary Physics, 57(2):274–275, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2016:CTO

[3499] Peter J. Dobson. The current trends of optics and photonics, edited by Cheng-Chung Lee. Scope: monograph. level: postgraduate; early career researcher; researcher; scientist; engineer. Contemporary Physics, 57 (2):275, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2016:PHE

[3500] Thomas Peters. Physical hydrodynamics, by Etienne Guyon, Jean-Pierre Hulin, Luc Petit, and Catalin D Mitescu. Scope: textbook. level: undergraduate. Contemporary Physics, 57(2):276, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:ECG

[3501] A. H. Harker. Electrical conduction in graphene and nanotubes, by Shigeji Fujita and Akira Suzuki. Scope: textbook. level: postgraduate; early career researcher. Contemporary Physics, 57(2):276–277, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2016:HSG

[3502] Salvatore Esposito. Higher speculations: grand theories and failed revolutions in physics and cosmology, by Helge Kragh. Scope: monograph. level: general readership, researcher, specialist, scientist. Contemporary Physics, 57(2):277–278, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:TNE

[3503] K. Alan Shore. Telecommunication network economics: from theory to applications, by Patrick Maillé and Bruno Tuffin. Scope: monograph. level: specialist, economists, engineers, early career researchers. Contemporary Physics, 57(2):278–279, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:SCS

[3504] K. Alan Shore. Soliton coding for secured optical communication link, by Iraj Sadegh Amiri, Sayed Ehsan Alavi and Sevia Mahdaliza Idrus. Scope: monograph. level: specialist, engineers, early career researchers. Contemporary Physics, 57(2):279–280, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:OHS

[3505] K. Alan Shore. The Oxford handbook of soft condensed matter, edited by Eugene M. Terentjev and David A. Weitz. Scope: handbook. level: specialist, physicists, engineers. Contemporary Physics, 57(2):280–281, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2016:USQ

[3506] Michael Belsley. Ultrafast spectroscopy: quantum information and wavepackets, by Joel Yuen-Zhou, Jacob J. Krich, Ivan Kassal, Allan S. Johnson and Alán Aspuru-Guzik. Scope: monograph. level: postgraduate, early career researcher, researcher, specialist. Contemporary Physics, 57(2):281–282, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cowan:2016:ISP

[3507] Brian Cowan. Introduction to statistical physics, by João Paulo Casquilho and Paulo Ivo Cortez Teixeira. Scope: textbook. level: undergraduate, advanced undergraduate. Contemporary Physics, 57(2):282, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stewart:2016:GRG

[3508] John M. Stewart. General relativity and gravitation: a centennial perspective, by Abhay Ashtekar, Beverly K. Berger, James Isenberg, Malcolm MacCallum. Scope: monograph. level: postgraduate, early career researcher, researcher. Contemporary Physics, 57(2):282–283, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:2016:GTP

[3509] Stephen M. Barnett. Gauge theories in particle physics: a practical introduction, fourth edition — 2 volume set, by Ian J. R. Aitchison and Anthony J. G. Hey. Scope: textbook. level: advanced undergraduate or postgraduate. Contemporary Physics, 57(2):283–284, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Leggett:2016:NIL

[3510] Anthony J. Leggett. Neutron interferometry: lessons in experimental quantum mechanics, wave-particle duality, and entanglement (2nd edition), by Helmut Rauch and Samuel A. Werner. Scope: monograph. level: researcher. Contemporary Physics, 57(2):284–285, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McOrist:2016:SFT

[3511] Jock McOrist. Supersymmetric field theories, by Sergio Cecotti. Scope: review. level: postgraduate student, early career researcher, researcher. Contemporary Physics, 57(2):285–286, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Knight:2016:PBJ

[3512] Peter Knight. Professor Bob Joseph. Contemporary Physics, 57(3):287– ??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hollowood:2016:CQM

[3513] Timothy J. Hollowood. Copenhagen quantum mechanics. Contemporary Physics, 57(3):289–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ford:2016:MDM

[3514] Ian J. Ford. Maxwell’s demon and the management of ignorance in stochastic thermodynamics. Contemporary Physics, 57(3):309–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Genoni:2016:CUG

[3515] Marco G. Genoni, Ludovico Lami, and Alessio Serafini. Conditional and unconditional Gaussian quantum dynamics. Contemporary Physics, 57 (3):331–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gavai:2016:QCP

[3516] Rajiv V. Gavai. The QCD critical point: an exciting odyssey in the femto-world. Contemporary Physics, 57(3):350–??, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jain:2016:APQ

[3517] Nitin Jain, Birgit Stiller, Imran Khan, Dominique Elser, Christoph Marquardt, and Gerd Leuchs. Attacks on practical quantum key distribution systems (and how to prevent them). Contemporary Physics, 57(3): 366–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dalmonte:2016:LGT

[3518] M. Dalmonte and S. Montangero. Lattice gauge theory simulations in the quantum information era. Contemporary Physics, 57(3):388–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2016:BRI

[3519] Chih-Yueh Wang. Book review: Ideal MHD, by Jeffrey P. Freidberg. Contemporary Physics, 57(3):413–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2016:ERA

[3520] Juan A. Añel. Essay review: Atmospheric ozone: historical background and state-of-the-art: a review of Ozone in the atmosphere: basic principles, natural and human impacts, by P. Fabian and M. Dameris. Contemporary Physics, 57(3):417–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quenby:2016:ERG

[3521] John Quenby. Essay review: God and the multiverse by Victor J. Stenger. Contemporary Physics, 57(3):421–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2016:ERQ

[3522] Mario Bertolotti. Essay review: Quantum optics, by John Garrison and Raymond Chiao. Contemporary Physics, 57(3):426–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2016:ERW

[3523] David Bailin. Essay review: Why string theory? by Joseph Conlon. Contemporary Physics, 57(3):433–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dombey:2016:ERB

[3524] Norman Dombey. Essay review: Building the H Bomb: a personal history, by Kenneth W Ford. Contemporary Physics, 57(3):439–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McOrist:2016:BRS

[3525] Jock McOrist. Book review: Science unshackled: how obscure, abstract, seemingly useless scientific research turned out to be the basis for modern life, by C. Renee James. Contemporary Physics, 57(3):443–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salama:2016:BRP

[3526] Amgad Salama. Book review: The principles of thermodynamics, by N. D. Hari Dass. Contemporary Physics, 57(3):443–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nikolopoulos:2016:BRP

[3527] L. A. A. Nikolopoulos. Book review: Path integrals and Hamiltonians: principles and methods by Belal E. Baaquie. Contemporary Physics, 57 (3):444–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2016:BRPa

[3528] R. S. Shorter. Book review: The physics companion, 2nd edition, by Anthony C. Fischer-Cripps. Contemporary Physics, 57(3):445–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2016:BRPb

[3529] Richard Sebastian Shorter. Book review: Physics project lab, by Paul Gluck and John King. Contemporary Physics, 57(3):445–??, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2016:BRP

[3530] Charles Downing. Book review: Plasmonic effects in metal-semiconductor nanostructures, by Alexey A. Toropov and Tatiana V. Shubina. Contemporary Physics, 57(3):446–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2016:BRL

[3531] Nikolaos E. Myridis. Book review: Laser measurement technology: fundamentals and applications, by Axel Donges and Reinhard Noll. Contemporary Physics, 57(3):446–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2016:BRK

[3532] Mark Nicol. Book review: Kepler and the universe: how one man revolutionized astronomy, by David K. Love. Contemporary Physics, 57(3): 447–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2016:BRA

[3533] Mark Nicol. Book review: Astrophysics and the evolution of the universe, by Leonard S. Kisslinger. Contemporary Physics, 57(3):448–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BRS

[3534] K. Alan Shore. Book review: Small scale optics, by Preecha Yupapin and Jalil Ali. Contemporary Physics, 57(3):449–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:BRPa

[3535] Manuel Vogel. Book review: Properties of materials, by P. F. Kelly. Contemporary Physics, 57(3):449–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BRP

[3536] K. Alan Shore. Book review: Principles of optics for engineers, by William S. C. Chang. Contemporary Physics, 57(3):450–??, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BRNa

[3537] K. Alan Shore. Book review: Nonlinear super-resolution nano-optics and applications, by Jingsong Wei. Contemporary Physics, 57(3):451– ??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Davidson:2016:BRR

[3538] Mark Davidson. Book review: Relativistic quantum mechanics, by Lawrence P. Horwitz. Contemporary Physics, 57(3):452–??, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2016:BRQ

[3539] Stephen J. Blundell. Book review: Quantum phase transitions in transverse field spin models, by Amit Dutta, Gabriel Aeppli, Bikas K. Chakrabarti, Uma Divakaran, Thomas F. Rosenbaum and Diptiman Sen. Contemporary Physics, 57(3):453–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2016:BRM

[3540] Tom Lancaster. Book review: Many-body physics, topology and geometry, by Siddhartha Sen and Kumar Sankar Gupta. Contemporary Physics, 57 (3):454–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2016:BRE

[3541] D. C. Barber. Book review: Essential MATLAB and Octave, by Jesus Rogel-Salazar. Contemporary Physics, 57(3):455–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:BRD

[3542] Manuel Vogel. Book review: Dark energy, by Miao Li, Xiao-Dong Li, Shuang Wang and Yi Wang. Contemporary Physics, 57(3):455–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2016:BRW

[3543] Miguel A. F. Sanjuán. Book review: Waves and oscillations in nature. An introduction, by A. Satya Narayanan and Swapan K. Saha. Contemporary Physics, 57(3):456–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2016:BRE

[3544] Thomas Collier. Book review: Elastic and plastic deformation of carbon nanotubes, by Hiroyuki Shima and Motohiro Sato. Contemporary Physics, 57(3):457–??, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Proctor:2016:HQC

[3545] Timothy J. Proctor and Viv Kendon. Hybrid quantum computing with ancillas. Contemporary Physics, 57(4):459–476, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Plato:2016:GEQ

[3546] A. D. K. Plato, C. N. Hughes, and M. S. Kim. Gravitational effects in quantum mechanics. Contemporary Physics, 57(4):477–495, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gaskins:2016:RIS

[3547] Jennifer M. Gaskins. A review of indirect searches for particle dark matter. Contemporary Physics, 57(4):496–525, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kok:2016:PQI

[3548] Pieter Kok. Photonic quantum information processing. Contemporary Physics, 57(4):526–544, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vinjanampathy:2016:QT

[3549] Sai Vinjanampathy and Janet Anders. Quantum thermodynamics. Contemporary Physics, 57(4):545–579, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fattahi:2016:SCL

[3550] Hanieh Fattahi. Sub-cycle light transients for attosecond, X-ray, fourdimensional imaging. Contemporary Physics, 57(4):580–595, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jones:2016:BRB

[3551] Derry W. Jones. Book review: Birds and frogs: selected papers, 1990– 2014, by Freeman J. Dyson. Contemporary Physics, 57(4):596–598, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:BRE

[3552] B. Ishak. Book review: Essential dynamics & relativity, by Peter J. O’Donnell. Contemporary Physics, 57(4):599–601, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:BRI

[3553] B. Ishak. Book review: An introduction to the sun and stars, 2nd Edition, by Simon F. Green and Mark H. Jones (eds.). Contemporary Physics, 57(4):602–604, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2016:BRD

[3554] Thomas Collier. Book review: Differentiability and fractality in dynamics of physical systems, by Ioan Merches and Maricel Agop. Contemporary Physics, 57(4):605, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2016:BRE

[3555] R. S. Shorter. Book review: Experimental optics: a manual for the laboratory, by G. F. C. Searle. Contemporary Physics, 57(4):605–606, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:BRPb

[3556] Manuel Vogel. Book review: Plasma physics: an introduction, by Richard Fitzpatrick. Contemporary Physics, 57(4):606, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:BRS

[3557] A. H. Harker. Book review: Statistical mechanics and applications in condensed matter, by Carlo Di Castro and Roberto Raimondi. Contemporary Physics, 57(4):606–607, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2016:BRP

[3558] Andrew Resnick. Book review: Photons in Fock space and beyond (3 volumes), by Reinhard Honegger and Alfred Rieckers. Contemporary Physics, 57(4):607–608, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2016:BRT

[3559] D. C. Barber. Book review: Theoretical foundations of digital imaging using Matlab, by Leonid P. Yaroslavsky. Contemporary Physics, 57(4):

608–609, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2016:BRS

[3560] David Bailin. Book review: De Sitter invariant special relativity, by Mu-Lin Yan. Contemporary Physics, 57(4):609–610, 2016. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2016:BRL

[3561] A. H. Harker. Book review: Lectures on quantum mechanics (2nd edition), by Steven Weinberg. Contemporary Physics, 57(4):610, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2016:BRF

[3562] Nikolaos E. Myridis. Book review: Foundations of linear and generalized linear models, by Alan Agresti. Contemporary Physics, 57(4): 610–611, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2016:BRT

[3563] Robert S. MacKay. Book review: Thermodynamics kept simple — a molecular approach: what is the driving force in the world of molecules?, by Roland Kjellander. Contemporary Physics, 57(4):611–613, 2016. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2016:BRT

[3564] B. Ishak. Book review: Terahertz astronomy, by Christopher K. Walker. Contemporary Physics, 57(4):613–614, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BRQa

[3565] K. Alan Shore. Book review: Quantitative biomedical optics, by Irving J. Bigio and Sergio Fantini. Contemporary Physics, 57(4):614–615, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BRNb

[3566] K. Alan Shore. Book review: Nonlinear resonances, by S. Rajasekar and M. A. F. Sanjuan. Contemporary Physics, 57(4):615–616, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2016:BRQb

[3567] K. Alan Shore. Book review: Quantum optomechanics, by W. P. Bowen and G. J. Milburn. Contemporary Physics, 57(4):616–617, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2016:BRL

[3568] Manuel Vogel. Book review: Laser physics: from principles to practical work in the lab, by Marc Eichhorn. Contemporary Physics, 57(4): 617, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2016:EB

[3569] Anonymous. Editorial Board. Contemporary Physics, 57(4):ebi, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fairbairn:2017:MNL

[3570] Malcolm Fairbairn and James L. Pinfold. MoEDAL — a new light on the high-energy frontier. Contemporary Physics, 58(1):1–24, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Oi:2017:NQS

[3571] Daniel K. L. Oi, Alex Ling, James A. Grieve, Thomas Jennewein, Aline N. Dinkelaker, and Markus Krutzik. Nanosatellites for quantum science and technology. Contemporary Physics, 58(1):25–52, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Residori:2017:RWU

[3572] S. Residori, M. Onorato, U. Bortolozzo, and F. T. Arecchi. Rogue waves: a unique approach to multidisciplinary physics. Contemporary Physics, 58(1):53–69, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gorban:2017:BNS

[3573] A. N. Gorban and I. V. Karlin. Beyond Navier–Stokes equations: capillarity of ideal gas. Contemporary Physics, 58(1):70–90, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2017:BRDa

[3574] David Bailin. Book review: Dear Professor Dyson, edited by Dwight E. Neuenschwander. Contemporary Physics, 58(1):91–92, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRG

[3575] K. Alan Shore. Book review: Goethe’s ‘Exposure of Newton’s Theory’: a polemic on Newton’s theory of light and colour, translated by Michael Duck and Michael Petry. Contemporary Physics, 58(1):93–95, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRIa

[3576] B. Ishak. Book review: Introduction to black hole physics, by Valeri P. Frolov and Andrei Zelnikov. Contemporary Physics, 58(1):96–98, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRSa

[3577] B. Ishak. Book review: Statistics, data mining, and machine learning in astronomy: a practical Python guide for the analysis of survey data, by Zeljko Ivezić, Andrew J. Connolly, Jacob T. VanderPlas and Alexander Gray. Contemporary Physics, 58(1):99, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2017:BRC

[3578] Peter V. E. McClintock. Book review: Controlled thermonuclear fusion, by Jean Louis Bobin. Contemporary Physics, 58(1):100, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:BRN

[3579] Peter J. Bussey. Book review: The new cosmos, by David J. Eicher. Contemporary Physics, 58(1):100–101, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2017:BRF

[3580] Stephen J. Blundell. Book review: Fundamentals of thermoelectricity, by Kamran Behnia. Contemporary Physics, 58(1):101–102, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRL

[3581] Manuel Vogel. Book review: Light sources, second edition: basics of lighting technologies and applications, by Spyridon Kitsinelis. Contemporary Physics, 58(1):102, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRF

[3582] Andrew Resnick. Book review: Fourier optics and computational imaging, by Kedar Khare. Contemporary Physics, 58(1):102–103, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2017:BRI

[3583] David Bailin. Book review: Inflation and string theory, by Daniel Baumann and Liam McAllister. Contemporary Physics, 58(1):103–105, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRB

[3584] Andrew Resnick. Book review: Brownian ratchets: from statistical physics to bio and nano-motors, by David Cubero and Ferruccio Renzoni. Contemporary Physics, 58(1):105, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2017:BRP

[3585] John D. Clayton. Book review: Principles of physics: from quantum field theory to classical mechanics, by Jun Ni. Contemporary Physics, 58(1): 105–106, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRSa

[3586] Manuel Vogel. Book review: Stellar explosions: hydrodynamics and nucleosynthesis, by Jordi Jose. Contemporary Physics, 58(1):106–107, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2017:BRFa

[3587] Peter V. E. McClintock. Book review: The foundations of physical law, by Peter Rowlands. Contemporary Physics, 58(1):107–108, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRGa

[3588] B. Ishak. Book review: Geodynamics (3rd Edition), by Donald L. Turcotte. Contemporary Physics, 58(1):108, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2017:BRQ

[3589] Nikolaos E. Myridis. Book review: Quantum models of cognition and decision, by Jerome R. Busemeyer and Peter D. Bruza. Contemporary Physics, 58(1):109, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whyntie:2017:BRNa

[3590] Tom Whyntie. Book review: Next generation experiments to measure the neutron lifetime, edited by Susan J. Seestrom. Contemporary Physics, 58(1):109–110, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2017:BRI

[3591] Charles Downing. Book review: Introduction to the statistical physics of integrable many-body systems, by Ladislav Samaj & Zoltán Bajnok. Contemporary Physics, 58(1):110–111, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lawrie:2017:BRI

[3592] Ian D. Lawrie. Book review: Introduction to the AdS/CFT correspondence, by H. Năstase. Contemporary Physics, 58(1):111–112, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whyntie:2017:BRNb

[3593] Tom Whyntie. Book review: New results and actual problems in particle & astroparticle physics and cosmology: Proceedings of XXIXth International Workshop on High Energy Physics, edited by R. Ryutin, V. Petrov, and V. Kiselev. Contemporary Physics, 58(1):112, 2017. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Whyntie:2017:BRH

[3594] Tom Whyntie. Book review: The High Luminosity Large Hadron Collider: the new machine for illuminating the mysteries of Universe, edited by O. Brüning & L. Rossi. Contemporary Physics, 58(1):112–113, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRMa

[3595] K. Alan Shore. Book review: Mathematical methods for geophysics and space physics, by William I. Newman. Contemporary Physics, 58(1): 113–114, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRR

[3596] K. Alan Shore. Book review: Research on the physics of cancer, edited by Bernard S. Gertsman. Contemporary Physics, 58(1):114–115, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2017:BRO

[3597] Peter V. E. McClintock. Book review: The origin and nature of life on earth: the emergence of the fourth geosphere, by Eric Smith and Harold J. Morowitz. Contemporary Physics, 58(1):115–116, 2017. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRAa

[3598] Manuel Vogel. Book review: Atoms and molecules interacting with light, by Peter van der Straten and Harold Metcalf. Contemporary Physics, 58(1):116, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2017:BRC

[3599] John D. Clayton. Book review: Crystallography: a very short introduction, by A. M. Glazer. Contemporary Physics, 58(1):117, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Yin:2017:BQM

[3600] Zhang qi Yin and Tongcang Li. Bringing quantum mechanics to life: from Schrödinger’s cat to Schrödinger’s microbe. Contemporary Physics, 58(2):119–139, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Munro:2017:CXR

[3601] P. R. T. Munro. Coherent X-ray imaging across length scales. Contemporary Physics, 58(2):140–159, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barger:2017:HSF

[3602] A. J. Barger. The history of star formation in the universe. Contemporary Physics, 58(2):160–175, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2017:PCN

[3603] Miguel A. F. Sanjuán. Physics of cancer: the new adventure of physicists against cancer. Contemporary Physics, 58(2):176–178, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Riley:2017:BRF

[3604] Norman Riley. Book review: Fluid dynamics, by Peter S. Bernard. Contemporary Physics, 58(2):179, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Torok:2017:BRC

[3605] Peter Török. Book review: A course in lens design, by Chris Velzel. Contemporary Physics, 58(2):179–180, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRBa

[3606] Manuel Vogel. Book review: Biophysics of DNA, by Alexander Vologodskii. Contemporary Physics, 58(2):180–181, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRH

[3607] Andrew Resnick. Book review: Handbook of granular materials, edited by Scott V. Franklin and Mark D. Shattuck. Contemporary Physics, 58 (2):181, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2017:BRFb

[3608] Peter V. E. McClintock. Book review: Fundamentals of fluorescence microscopy: exploring life with light, by Partha Pratim Mondal and Alberto Diaspro. Contemporary Physics, 58(2):181–182, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:BRP

[3609] Peter J. Bussey. Book review: From photons to Higgs: a story of light (2nd edition), by Moo-Young Han. Contemporary Physics, 58(2):182– 183, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRMb

[3610] K. Alan Shore. Book review: Microphysics of atmospheric phenomena, by Boris M. Smirnov. Contemporary Physics, 58(2):183, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRTa

[3611] K. Alan Shore. Book review: Tunable micro-optics, edited by Hans Zappe and Claudia Duppe. Contemporary Physics, 58(2):183–184, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRMc

[3612] K. Alan Shore. Book review: Maxwell’s enduring legacy: a scientific history of the Cavendish laboratory, by Malcolm Longair. Contemporary Physics, 58(2):184–186, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:2017:BRP

[3613] Lewis Ryder. Book review: Particle physics in the LHC era, by G. Barr, R. Devenish, R. Walczak and T. Weidberg. Contemporary Physics, 58 (2):186, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Medel-Vera:2017:BRS

[3614] Carlos Medel-Vera. Book review: Seismic amplitude: an interpreter’s handbook, by Rob Simm and Mike Bacon. Contemporary Physics, 58(2): 186–187, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BREa

[3615] Manuel Vogel. Book review: Eyes on the sky: a spectrum of telescopes, by Francis Graham-Smith. Contemporary Physics, 58(2):187–188, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2017:BRP

[3616] A. H. Harker. Book review: The physics of solids, by J. B. Ketterson. Contemporary Physics, 58(2):188, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2017:BRF

[3617] A. H. Harker. Book review: Fundamentals and applications of magnetic materials, by Kannan M. Krishnan. Contemporary Physics, 58(2): 188–189, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2017:BRC

[3618] Robert Appleby. Book review: Challenges and goals for accelerators in the xxi century, edited by Oliver Bruning and Stephen Myers. Contemporary Physics, 58(2):189–190, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2017:BRN

[3619] Mario Bertolotti. Book review: Nonlinear optics: principles and applications, by Karsten Rottwitt and Peter Tidemand-Lichtenberg. Contemporary Physics, 58(2):190, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Frs:2017:BRE

[3620] Robert Sinclair Mackay (Frs). Book review: The expanding universe: a primer on relativistic cosmology, by W. D. Heacox. Contemporary Physics, 58(2):190–193, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRAb

[3621] Manuel Vogel. Book review: Astrophysics in a nutshell (2nd edition), by Dan Maoz. Contemporary Physics, 58(2):193, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Al-Tabbakh:2017:BRM

[3622] Ahmed A. Al-Tabbakh. Book review: Microscopy: a very short introduction, by Terence Allen. Contemporary Physics, 58(2):193–196, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lakshminarayanan:2017:BRV

[3623] Vasudevan Lakshminarayanan. Book review: Visible and invisible — the wonders of light phenomena, by Olmes Bisi. Contemporary Physics, 58 (2):196, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRCa

[3624] Manuel Vogel. Book review: Colour: how we see it and how we use it, by Michael Mark Woolfson. Contemporary Physics, 58(2):197, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2017:BRG

[3625] Stephen J. Blundell. Book review: Group theory in a nutshell for physicists, by A. Zee. Contemporary Physics, 58(2):197–198, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2017:BRH

[3626] M. P. Gururajan. Book review: Handbook of crystal growth, 2nd edition: Fundamentals IA: thermodynamics and kinetics & Fundamentals IB: transport and stability (two book set), edited by T. Nishinaga. Contemporary Physics, 58(2):198, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRSb

[3627] B. Ishak. Book review: Space physics: an introduction, by C. T. Russell, J. G. Luhmann and R. J. Strangeway. Contemporary Physics, 58 (2):199, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2017:BRI

[3628] Peter J. Dobson. Book review: An introduction to graphene plasmonics, by P. A. D. Gonçalves and N. M. R. Peres. Contemporary Physics, 58(2): 199–200, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2017:BRN

[3629] Peter J. Dobson. Book review: Novel optical nanoprobes for chemical and biological analysis, by L. Chen, Y. Wang, X. Fu, and L. Chen. Contemporary Physics, 58(2):200, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRB

[3630] B. Ishak. Book review: Biomagnetics: principles and applications of biomagnetic stimulation and imaging, edited by Shoogo Ueno and Masaki Sekino. Contemporary Physics, 58(2):200–201, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRSc

[3631] B. Ishak. Book review: Spectra of atoms and molecules (3rd edition), by Peter F. Bernath. Contemporary Physics, 58(2):201–202, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Menghrajani:2017:BRE

[3632] Kishan Menghrajani. Book review: Excitations in organic solids, by Vladimir M. Agranovich. Contemporary Physics, 58(2):202, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2017:BRI

[3633] Ken Zetie. Book review: International young physicists tournament: problems and solutions 2014, edited by S. Wang and W. Gao. Contemporary Physics, 58(2):202–203, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRO

[3634] Andrew Resnick. Book review: Optical tweezers: principles and applications, by Philip H. Jones, Onofrio M. Maragò, and Giovanni Volpe. Contemporary Physics, 58(2):203, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2017:BRS

[3635] Stephen J. Blundell. Book review: A superfluid universe, by K. Huang. Contemporary Physics, 58(2):203–204, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2017:BRF

[3636] J. Rogel-Salazar. Book review: Fashion, faith and fantasy in the new physics of the universe, by Roger Penrose. Contemporary Physics, 58(2): 204–205, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2017:BRI

[3637] G. W. Gibbons. Book review: Images of time: mind, science and reality, by George Jaroszkiewicz. Contemporary Physics, 58(2):205–206, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Eroglu:2017:SCA

[3638] Deniz Eroglu, Jeroen S. W. Lamb, and Tiago Pereira. Synchronisation of chaos and its applications. Contemporary Physics, 58(3):207–243, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ford:2017:FDT

[3639] G. W. Ford. The fluctuation–dissipation theorem. Contemporary Physics, 58(3):244–252, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Waxman:2017:DER

[3640] David Waxman. The diffusion equation of random genetic drift — biology’s analogue of the Schrödinger equation? Contemporary Physics, 58 (3):253–261, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:FU

[3641] Peter J. Bussey. Flight from unreason. Contemporary Physics, 58(3): 262–264, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:REW

[3642] B. Ishak. A review of Electromagnetic waves for thermonuclear fusion research, by Ernesto Mazzucato. Contemporary Physics, 58(3):265–267, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:RMS

[3643] B. Ishak. A review of To measure the sky: an introduction to observational astronomy, 2nd edition, by F. R. Chromey. Contemporary Physics, 58(3):268–270, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2017:BRR

[3644] Thomas J. Bridges. Book review: Ray tracing and beyond: phase space methods in plasma wave theory, by E. R. Tracy, A. J. Brizard, A. S. Richardson and A. N. Kaufman. Contemporary Physics, 58(3):271–272, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2017:BRE

[3645] Miguel A. F. Sanjuán. Book review: Entropy demystified: the second law reduced to plain common sense, 2nd edition, by Arieh Ben-Naim. Contemporary Physics, 58(3):272, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caridi:2017:BRL

[3646] Francesco Caridi. Book review: Laser-induced breakdown spectroscopy: theory and applications, edited by Sergio Musazzi and Umberto Perini. Contemporary Physics, 58(3):273, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Falcone:2017:BRE

[3647] Francisco Falcone. Book review: Electrical properties of materials, ninth edition, by Laszlo Solymar, Donald Walsh and Richard R. A. Syms. Contemporary Physics, 58(3):273, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2017:BRA

[3648] Salvatore Esposito. Book review: Atlas of meteorites, by M. M. Grady, G. Pratesi and V. Moggi Cecchi. Contemporary Physics, 58(3):273–274, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRSa

[3649] Andrew Resnick. Book review: Book review: Scientific explanation and methodology of science, by Guichun Guo and Chuang Liu. Contemporary Physics, 58(3):274–275, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2017:BRC

[3650] G. W. Gibbons. Book review: The cosmic web: mysterious architecture of the universe, by J. Richard Gott. Contemporary Physics, 58(3): 275–276, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2017:BRC

[3651] Juan A. Añel. Book review: Copernicus: a very short introduction, by Owen Gingerich. Contemporary Physics, 58(3):276, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Downing:2017:BRE

[3652] C. A. Downing. Book review: Exactly solvable models in many-body theory, by N. H. March and G. G. N. Angilella. Contemporary Physics, 58(3):276–277, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2017:BRCa

[3653] Nikolaos E. Myridis. Book review: The Cambridge handbook of applied perception research, edited by R. R. Hoffman, P. A. Hancock, M. W. Scerbo, R. Parasuraman, and J. L. Szalma. Contemporary Physics, 58 (3):277, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2017:BRB

[3654] Gunter M. Schütz. Book review: The briefest history of time: the history of histories of time and the misconstrued association between entropy and time, by Arieh Ben-Naim. Contemporary Physics, 58(3):278, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2017:BRE

[3655] Chih-Yueh Wang. Book review: Essential radio astronomy, by James J. Condon and Scott M. Ransom. Contemporary Physics, 58(3):278– 279, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2017:BRPa

[3656] Thomas J. Bridges. Book review: Particles in the coastal ocean, by D. R. Lynch, D. A. Greenberg, A. Bilgili, D. J. McGillicuddy Jr, J. P. Manning, and A. L. Aretxabaleta. Contemporary Physics, 58(3):279–280, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2017:BRQa

[3657] Thomas Collier. Book review: Quantum mechanics I, the fundamentals, by Shanmuganathan Rajasekar and Ramiah Velusamy. Contemporary Physics, 58(3):280, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2017:BRQb

[3658] Thomas Collier. Book review: Quantum mechanics II, advanced topics, by Shanmuganathan Rajasekar and Ramiah Velusamy. Contemporary Physics, 58(3):280–281, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2017:BRS

[3659] Thomas Collier. Book review: Strong light-matter coupling: from atoms to solid-state systems, edited by Alexia Auffèves, Dario Gerace, Maxime Richard, Stefano Portolan, Marcelo França Santos, Leong Chuan Kwek, and Christian Miniatura. Contemporary Physics, 58(3):281–282, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lakshminarayanan:2017:BRL

[3660] Vasudevan Lakshminarayanan. Book review: Learning scientific programming with Python, by Christian Hill. Contemporary Physics, 58(3): 282–284, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Johnson:2017:BRA

[3661] R. Barry Johnson. Book review: Advanced geometrical optics, by P. D. Lin. Contemporary Physics, 58(3):284–285, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRSb

[3662] Manuel Vogel. Book review: Strange glow: the story of radiation, by Timothy J. Jorgensen. Contemporary Physics, 58(3):285, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2017:BRT

[3663] Marco Casazza. Book review: Thermodynamics and statistical mechanics of macromolecular systems, by Michael Bachmann. Contemporary Physics, 58(3):285–286, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tanner:2017:AAS

[3664] B. K. Tanner. All about science: philosophy, history, sociology & communication, edited by Maria burguete and lui lam. Contemporary Physics, 58(3):286, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2017:BRS

[3665] Thomas Peters. Book review: The structure of complex networks: theory and applications, by Ernesto Estrada. Contemporary Physics, 58 (3):287, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2017:BRM

[3666] Matt Probert. Book review: 2D materials for nanoelectronics, edited by Michel Houssa, Athanasios Dimoulas, and Alessandro Molle. Contemporary Physics, 58(3):287–288, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRCb

[3667] Manuel Vogel. Book review: Cosmic rays and particle physics, 2nd edition, by Thomas K. Gaisser, Ralph Engel, and Elisa Resconi. Contemporary Physics, 58(3):288, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRD

[3668] B. Ishak. Book review: Deep-sky companions: the Caldwell objects, 2nd edition, by S. J. O’Meara. Contemporary Physics, 58(3):288–289, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRC

[3669] B. Ishak. Book review: Cosmic magnetic fields, by Philip P. Kronberg. Contemporary Physics, 58(3):289–290, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2017:BRD

[3670] Miriam R. Aczel. Book review: The dark side of technology, by Peter Townsend. Contemporary Physics, 58(3):290–291, 2017. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRIb

[3671] B. Ishak. Book review: Introduction to modern magnetohydrodynamics, by Sébastien Galtier. Contemporary Physics, 58(3):291–292, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRTb

[3672] K. Alan Shore. Book review: Transport in multilayered nanostructures: the dynamic mean-field theory approach, 2nd edition, by James K. Freericks. Contemporary Physics, 58(3):292–293, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRO

[3673] K. Alan Shore. Book review: Optical compressive imaging, edited by Adrian Stern. Contemporary Physics, 58(3):293, 2017. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2017:BRN

[3674] K. Alan Shore. Book review: Nanowires: building blocks for nanoscience and nanotechnology, by A. Zhang, G. Zheng, and C. Lieber. Contemporary Physics, 58(3):294, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRBb

[3675] Manuel Vogel. Book review: Basic X-ray scattering for soft matter, by W. H. de Jeu. Contemporary Physics, 58(3):294–295, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2017:BRI

[3676] Giovanni Volpe. Book review: The impact of discovering life beyond Earth, edited by S. J. Dick. Contemporary Physics, 58(3):295–296, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:BRE

[3677] Peter J. Bussey. Book review: Electroweak physics at the LHC, by M. U. Mozer. Contemporary Physics, 58(3):296, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2017:BRN

[3678] Peter V. E. McClintock. Book review: Nonlinear dynamics in biological systems, edited by Jorge Carballido-Landeira and Bruno Escribano. Contemporary Physics, 58(3):296–297, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2017:BRT

[3679] M. P. Gururajan. Book review: Thermodynamics: a complete undergraduate course, by A. M. Steane. Contemporary Physics, 58(3):297, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2017:BRQa

[3680] K. P. Zetie. Book review: Quantum physics: a first encounter, by V. Scarani, and translated by Rachael Thew. Contemporary Physics, 58(3): 297–298, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2017:BRP

[3681] Miguel A. F. Sanjuán. Book review: Peregrinations from physics to phylogeny: essays on the occasion of Hao Bailin’s 80th birthday, edited by K. K. Phua and M. Ge. Contemporary Physics, 58(3):298–299, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRSb

[3682] Andrew Resnick. Book review: Singular optics, by G. J. Gbur. Contemporary Physics, 58(3):299–300, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2017:BRU

[3683] Andrew Resnick. Book review: Ultrafast biophotonics, by P. Vasa and D. Mathur. Contemporary Physics, 58(3):300, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRM

[3684] Manuel Vogel. Book review: From matter to life: information and causality, edited by S. I. Walker, P. C. W. Davies and G. F. R. Ellis. Contemporary Physics, 58(3):300–301, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2017:BRC

[3685] Chih-Yueh Wang. Book review: Celestial mechanics and astrodynamics: theory and practice, by P. Gurfil and P. K. Seidelmann. Contemporary Physics, 58(3):301, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2017:PRP

[3686] Roy Howard. Pervasive randomness in physics: an introduction to its modelling and spectral characterisation. Contemporary Physics, 58(4):

303–330, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clements:2017:IPM

[3687] David L. Clements. An introduction to the Planck mission. Contemporary Physics, 58(4):331–348, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2017:BRI

[3688] Chih-Yueh Wang. Book review: ITER physics, by C. Wendell Horton, Jr. and Sadruddin Benkadda. Contemporary Physics, 58(4):349–352, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:RPE

[3689] B. Ishak. A review of Planetesimals: early differentiation and consequences for planets, edited by L. T. Elkins-Tanton and B. P. Weiss. Contemporary Physics, 58(4):353–355, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanchez-Canizares:2017:BRH

[3690] Javier Sánchez-Cañizares. Book review: How can physics underlie the mind? Top-down causation in the human context, by G. Ellis. Contemporary Physics, 58(4):356, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2017:BRCb

[3691] Nikolaos E. Myridis. Book review: Cosmology with MATLAB, by D. Green. Contemporary Physics, 58(4):357, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:BRS

[3692] Peter J. Bussey. Book review: A student’s manual for a first course in general relativity, by Robert B. Scott. Contemporary Physics, 58(4): 357–358, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2017:BRI

[3693] Miguel A. F. Sanjuán. Book review: An introduction to space plasma complexity, by Tom Tien Sun Chang. Contemporary Physics, 58(4):358– 359, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zetie:2017:BRQb

[3694] K. P. Zetie. Book review: Quantum weirdness, by W. J. Mullin. Contemporary Physics, 58(4):359, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:BRF

[3695] Peter J. Bussey. Book review: A fortunate universe: life in a finely tuned cosmos, by G. F. Lewis and L. A. Barnes. Contemporary Physics, 58(4): 359–360, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2017:BRA

[3696] Mario Bertolotti. Book review: Attosecond experiments on plasmonic nanostructures: principles and experiments, by J. Schötz. Contemporary Physics, 58(4):360–361, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2017:BRP

[3697] Mark Nicol. Book review: Practical optical interferometry: imaging at visible and infrared wavelengths, by David F. Buscher. Contemporary Physics, 58(4):361, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2017:BRH

[3698] Frank Berkshire. Book review: How do you find an exoplanet?, by J. A. Johnson. Contemporary Physics, 58(4):362, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2017:BRL

[3699] Frank Berkshire. Book review: Lectures on the theory of water waves, edited by T. J. Bridges, M. D. Groves and D. P. Nicholls. Contemporary Physics, 58(4):362–363, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2017:BRS

[3700] Frank Berkshire. Book review: The science of the perfect swing, by P. Dewhurst. Contemporary Physics, 58(4):363, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2017:BRP

[3701] Dr. Richard Sebastian Shorter. Book review: Principles of adaptive optics, 4th edition, by Robert K. Tyson. Contemporary Physics, 58(4):

364, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2017:BRL

[3702] R. S. Shorter. Book review: A lesson for the future of our science: my testimony on Lord Patrick M. S. Blackett, by A. Zichichi. Contemporary Physics, 58(4):364–365, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2017:BRT

[3703] Matt Probert. Book review: The theory of intermolecular forces, 2nd edition, by A. Stone. Contemporary Physics, 58(4):365, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:2017:BRT

[3704] Lewis Ryder. Book review: Topological quantum numbers in nonrelativistic physics, by D. J. Thouless. Contemporary Physics, 58(4):365– 366, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRA

[3705] B. Ishak. Book review: Asteroids: astronomical and geological bodies, by T. H. Burbine. Contemporary Physics, 58(4):366–368, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2017:BRP

[3706] Peter J. Dobson. Book review: Physics at the biomolecular interface: fundamentals for molecular targeted therapy, by A. Fernández. Contemporary Physics, 58(4):369, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BRS

[3707] Manuel Vogel. Book review: The structure and dynamics of cities, by M. Barthelemy. Contemporary Physics, 58(4):369–370, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRQ

[3708] B. Ishak. Book review: A question and answer guide to astronomy, 2nd edition, by C. Christian and J-R. Roy. Contemporary Physics, 58 (4):370, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2017:BRGb

[3709] B. Ishak. Book review: Gravitational lensing, by S. Dodelson. Contemporary Physics, 58(4):370–371, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2017:BRQ

[3710] Stephen J. Blundell. Book review: Quantum measurement, by P. Busch, P. J. Lahti, J.-P. Pellonpää, and K. Ylinen. Contemporary Physics, 58 (4):371, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Al-Tabbakh:2017:BRS

[3711] Ahmed A. Al-Tabbakh. Book review: Symmetry in crystallography: understanding the international tables, by Paolo G. Radaelli. Contemporary Physics, 58(4):371–373, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ooi:2017:BRL

[3712] Kelvin J. A. Ooi. Book review: Light-matter interaction: physics and engineering at the nanoscale, 2nd edition, by J. Weiner and F. Nunes. Contemporary Physics, 58(4):373–374, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2017:BRP

[3713] Fredric W. Taylor. Book review: Physics of radiation and climate, by Michael A. Box and Gail P. Box. Contemporary Physics, 58(4):374, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zubair:2017:BRF

[3714] Muhammad Zubair. Book review: Fundamental and applied nanoelectromagnetics, edited by A. Maffucci and S. A. Maksimenko. Contemporary Physics, 58(4):374–375, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2017:BRF

[3715] Marco Casazza. Book review: Foundation of statistical energy analysis in vibroacoustics, by A. Le Bot. Contemporary Physics, 58(4):375, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2017:BRPb

[3716] Thomas J. Bridges. Book review: A primer on quantum fluids, by C. F. Barenghi and N. G. Parker. Contemporary Physics, 58(4):375–376, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Medel-Vera:2017:BRC

[3717] Carlos Medel-Vera. Book review: Computational seismology, by H. Igel. Contemporary Physics, 58(4):376, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2017:BRF

[3718] Matt Probert. Book review: Fundamentals of condensed matter physics, by M. L. Cohen and S. G. Louie. Contemporary Physics, 58(4):377, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2017:BRP

[3719] David Bailin. Book review: The physical world: an inspirational tour of fundamental physics, by N. Manton and N. Mee. Contemporary Physics, 58(4):377–379, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Modi:2017:BRQ

[3720] Kavan Modi. Book review: Quantum stochastics, by Mou-Hsiung Chang. Contemporary Physics, 58(4):379, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2017:BRDb

[3721] David Bailin. Book review: A dictionary of science, 7th edition, by J. Law. Contemporary Physics, 58(4):379–380, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2017:BREb

[3722] Manuel Vogel. Book review: The Earth’s inner core: revealed by observational seismology, by H. Tkalcic. Contemporary Physics, 58(4):380, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2017:BRI

[3723] Peter J. Bussey. Book review: Introduction to cosmology, 2nd edition, by B. Ryden. Contemporary Physics, 58(4):381, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:2017:BRB

[3724] Lewis Ryder. Book review: Budapest scientific: a guidebook, by István Hargittai and Magdolna Hargittai. Contemporary Physics, 58(4):381, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2017:EB

[3725] Anonymous. Editorial board. Contemporary Physics, 58(4):ebi, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jack:2018:T

[3726] Ian Jack. The a-theorem. Contemporary Physics, 59(1):1–15, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2018:RSR

[3727] Art Hobson. Review and suggested resolution of the problem of Schrödinger’s cat. Contemporary Physics, 59(1):16–30, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cavalleri:2018:PIS

[3728] Andrea Cavalleri. Photo-induced superconductivity. Contemporary Physics, 59(1):31–46, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Villeneuve:2018:AS

[3729] D. M. Villeneuve. Attosecond science. Contemporary Physics, 59(1): 47–61, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bailin:2018:CB

[3730] David Bailin. Clear as a bell. Contemporary Physics, 59(1):62–67, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ryder:2018:RWC

[3731] Lewis Ryder. A review of Who cares about particle physics? Making sense of the Higgs Boson, the Large Hadron Collider and CERN, by Pauline Gagnon, and Modern elementary particle physics: explaining and extending the standard model, 2nd edition, by Gordon Kane. Contemporary Physics, 59(1):68–69, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:RSA

[3732] B. Ishak. A review of Spectral atlas for amateur astronomers, by R. Walker. Contemporary Physics, 59(1):70–72, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2018:BRE

[3733] Fredric W. Taylor. Book review: Exoplanetary atmospheres: theoretical concepts and foundations, by K. Heng. Contemporary Physics, 59 (1):73, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRQa

[3734] Andrew Resnick. Book review: Quantum mechanics, by Richard Fitzpatrick. Contemporary Physics, 59(1):73–74, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRA

[3735] Andrew Resnick. Book review: Advances in disordered systems, random processes and some applications, edited by P. Contucci and C. Giardina. Contemporary Physics, 59(1):74, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRC

[3736] Andrew Resnick. Book review: From classical to quantum fields, by L. Baulieu, J. Iliopoulos and R. Sénéor. Contemporary Physics, 59(1): 74–75, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRIa

[3737] Manuel Vogel. Book review: Introduction to planetary photometry, by M. K. Shepard. Contemporary Physics, 59(1):75–76, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2018:BRM

[3738] John D. Clayton. Book review: Macroscropic electrodynamics: an introductory graduate treatment, by W. Wilcox and C. Thron. Contemporary Physics, 59(1):76–77, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2018:BRT

[3739] M. P. Gururajan. Book review: Tales of the quantum: understanding physics’ most fundamental theory, by A. Hobson. Contemporary Physics, 59(1):77–78, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2018:BRL

[3740] Nikolaos E. Myridis. Book review: Lab-on-fiber technology, edited by A. Cusano, M. Consales, A. Crescitelli, and A. Ricciardi. Contemporary Physics, 59(1):78–79, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2018:BRA

[3741] Juan A. Añel. Book review: Advanced computational fluid and aerodynamics, by P. G. Tucker. Contemporary Physics, 59(1):79–80, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRH

[3742] K. Alan Shore. Book review: Hybrid phonons in nanostructures, by B. K. Ridley. Contemporary Physics, 59(1):80, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRO

[3743] K. Alan Shore. Book review: Optical microring resonators: theory, techniques, and applications, by V. Van. Contemporary Physics, 59(1):80– 81, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2018:BREa

[3744] Thomas Collier. Book review: Electromagnetic interactions, by S. D. Bosanac. Contemporary Physics, 59(1):81–82, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRIa

[3745] K. Alan Shore. Book review: An introduction to distributed optical fibre sensors, by A. H. Hartog. Contemporary Physics, 59(1):82–83, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zubair:2018:BRA

[3746] Muhammad Zubair. Book review: Advanced free space optics (FSO): a systems approach, by Arun K. Majumdar. Contemporary Physics, 59 (1):83, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Collier:2018:BREb

[3747] Thomas Collier. Book review: Electromagnetic waves in complex systems: selected theoretical and applied problems, edited by Y. Sirenko and L. Velychko. Contemporary Physics, 59(1):84, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRO

[3748] Peter J. Bussey. Book review: The Oxford handbook of the history of physics, by J. Buchwald and R. Fox. Contemporary Physics, 59(1):84– 85, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2018:BRC

[3749] Thomas Peters. Book review: Chaotic dynamics: fractals, tilings, and substitutions, by G. R. Goodson. Contemporary Physics, 59(1):85–86, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2018:BRA

[3750] Frank Berkshire. Book review: Advanced classical mechanics, by B. K. Bagchi. Contemporary Physics, 59(1):86, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRT

[3751] B. Ishak. Book review: Totality — the great American eclipses of 2017 and 2024, by M. Littmann and F. Espenak. Contemporary Physics, 59 (1):87, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRQ

[3752] B. Ishak. Book review: Quantum physics: what everybody needs to know, by M. G. Raymer. Contemporary Physics, 59(1):87–88, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2018:BRM

[3753] A. H. Harker. Book review: Materials for the 21st century, by D. Segal. Contemporary Physics, 59(1):88, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRFa

[3754] B. Ishak. Book review: Foundations of nuclear and particle physics, by T. W. Donnelly, J. A. Formaggio, B. R. Holstein, R.G. Milner, and B. Surrow. Contemporary Physics, 59(1):89, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2018:BRPa

[3755] A. H. Harker. Book review: Practical Bayesian inference: a primer for physical scientists, by C. A. L. Bailer-Jones. Contemporary Physics, 59 (1):89–90, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ooi:2018:BRW

[3756] Kelvin J. A. Ooi. Book review: The wonderful world of relativity: a precise guide for the general reader, by A. Steane. Contemporary Physics, 59(1):90, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRGa

[3757] Miriam R. Aczel. Book review: Gravity: a very short introduction, by T. Clifton. Contemporary Physics, 59(1):90–91, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2018:BRS

[3758] Salvatore Esposito. Book review: Shattered symmetry: group theory from the eightfold way to the periodic table, by P. Thyssen and A. Ceulemans. Contemporary Physics, 59(1):91–92, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRQa

[3759] Peter J. Bussey. Book review: The quantum revolution in philosophy, by R. Healey. Contemporary Physics, 59(1):92–94, 2018. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schutz:2018:BRT

[3760] Gunter M. Schütz. Book review: Thermal transport in low dimensions: from statistical physics to nanoscale heat transfer, edited by S. Lepri. Contemporary Physics, 59(1):94, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2018:BRG

[3761] Chih-Yueh Wang. Book review: The gas dynamics of explosions, by J. H. S. Lee. Contemporary Physics, 59(1):94–95, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2018:BRL

[3762] M. P. Gururajan. Book review: The lazy universe: an introduction to the principle of least action, by J. Coopersmith. Contemporary Physics, 59 (1):95–96, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ang:2018:BRM

[3763] Yee Sin Ang. Book review: Mass: the quest to understand matter from Greek atoms to quantum fields, by J. Baggott. Contemporary Physics, 59(1):96, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ang:2018:BRA

[3764] Yee Sin Ang. Book review: Amorphous semiconductors, by Sándor Kugler and Koichi Shimakawa. Contemporary Physics, 59(1):96–97, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRS

[3765] Manuel Vogel. Book review: Sound propagation through the Stochastic Ocean, by J. A. Colosi. Contemporary Physics, 59(1):97–98, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRT

[3766] Miriam R. Aczel. Book review: Thermodynamics of natural systems: theory and applications in geochemistry and environmental science, 3rd edition, by G. Anderson. Contemporary Physics, 59(1):98–99, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRH

[3767] Miriam R. Aczel. Book review: Hydrogen storage technology: materials and applications, by L. Klebanoff. Contemporary Physics, 59(1):99, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moya-Cessa:2018:BRI

[3768] Héctor Manuel Moya-Cessa. Book review: The illustrated wavelet transform handbook: introductory theory and applications in science, engineering, medicine and finance, 2nd edition, by P. S. Addison. Contemporary Physics, 59(1):100, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRSa

[3769] B. Ishak. Book review: The standard model in a nutshell, by D. Goldberg. Contemporary Physics, 59(1):100–101, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moya-Cessa:2018:BRL

[3770] Héctor Manuel Moya-Cessa. Book review: Lectures on light: nonlinear and quantum optics using the density matrix, 2nd edition, by S. C. Rand. Contemporary Physics, 59(1):101, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hobson:2018:BRP

[3771] Peter R. Hobson. Book review: The photomultiplier handbook, by A. G. Wright. Contemporary Physics, 59(1):101–102, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2018:BRQ

[3772] Mario Bertolotti. Book review: Quantum optics and nanophotonics: lecture notes of the Les Houches Summer School, edited by Claude Fabre, Vahid Sandoghdar, Nicolas Treps, and Leticia F. Cugliandolo. Contemporary Physics, 59(1):102–103, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRSb

[3773] B. Ishak. Book review: Spectroscopy for amateur astronomers, by M. F. M. Trypsteen and R. Walker. Contemporary Physics, 59(1):103, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2018:BRW

[3774] Stephen H. Ashworth. Book review: What goes up … gravity and scientific method, by P. Kosso. Contemporary Physics, 59(1):103–104, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRIb

[3775] Manuel Vogel. Book review: An introduction to X-ray physics, optics, and applications, by C. A. MacDonald. Contemporary Physics, 59(1): 104–105, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRGb

[3776] Miriam R. Aczel. Book review: Get funded: an insider’s guide to building an academic research program, by R. J. Trew. Contemporary Physics, 59(1):105–106, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2018:BRNa

[3777] Miguel A. F. Sanjuán. Book review: Nature’s third cycle: a story of sunspots, by A. R. Choudhuri. Contemporary Physics, 59(1):106–107, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

White:2018:DCG

[3778] C. D. White. The double copy: gravity from gluons. Contemporary Physics, 59(2):109–125, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Larocque:2018:TE

[3779] Hugo Larocque, Ido Kaminer, Vincenzo Grillo, Gerd Leuchs, Miles J. Padgett, Robert W. Boyd, Mordechai Segev, and Ebrahim Karimi. ‘Twisted’ electrons. Contemporary Physics, 59(2):126–144, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnett:2018:VFP

[3780] Stephen M. Barnett and Matthias Sonnleitner. The vacuum friction paradox and related puzzles. Contemporary Physics, 59(2):145–154, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2018:GDC

[3781] Gerard Gilmore. Gaia: 3-dimensional census of the Milky Way Galaxy. Contemporary Physics, 59(2):155–173, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Venegas-Andraca:2018:CDI

[3782] Salvador E. Venegas-Andraca, William Cruz-Santos, Catherine Mc-Geoch, and Marco Lanzagorta. A cross-disciplinary introduction to quantum annealing-based algorithms. Contemporary Physics, 59(2):174–197, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:RSS

[3783] B. Ishak. A review of Supersymmetry and string theory: beyond the standard model, 2nd edition, by Michael Dine. Contemporary Physics, 59(2):198–200, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2018:BRT

[3784] G. W. Gibbons. Book review: Theoretical frontiers in black holes and cosmology: theoretical perspectives in high energy physics, edited by R. Kallosh and E. Orazi. Contemporary Physics, 59(2):201, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benedict:2018:BRI

[3785] Keith A. Benedict. Book review: Introduction to topological quantum matter and quantum computation, by Tudor D. Stanescu. Contemporary Physics, 59(2):202, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sitek:2018:BRS

[3786] Marta Anna Sitek. Book review: Soil mechanics fundamentals, by Muni Budhu. Contemporary Physics, 59(2):202–203, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRW

[3787] B. Ishak. Book review: Waves, particles, and storms in geospace: a complex interplay, edited by G. Balasis, I. Daglis, and I. R. Mann. Contemporary Physics, 59(2):203, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sivakumar:2018:BRM

[3788] Srinivasan Sivakumar. Book review: Microcavities, by Alexey V. Kavokin, Jeremy J. Baumberg, Guillaume Maupuech, and Fabrice P. Laussy. Contemporary Physics, 59(2):203–204, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2018:BRG

[3789] Nikolaos E. Myridis. Book review: Generating random networks and graphs, by T. Coolen, A. Annibale, and E. Roberts. Contemporary Physics, 59(2):204–205, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2018:BRA

[3790] Stephen H. Ashworth. Book review: The atmosphere: a very short introduction, by P. I. Palmer. Contemporary Physics, 59(2):205, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRC

[3791] Peter J. Bussey. Book review: The conceptual framework of quantum field theory, by A. Duncan. Contemporary Physics, 59(2):205–206, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BROa

[3792] Manuel Vogel. Book review: An overview of gravitational waves: theory, sources and detection, edited by G. Auger and E. Plagnol. Contemporary Physics, 59(2):206–207, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2018:BRS

[3793] M. P. Gururajan. Book review: A student’s guide to dimensional analysis, by D. S. Lemons. Contemporary Physics, 59(2):207, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2018:BRB

[3794] Robert S. MacKay. Book review: Beyond peaceful co-existence: the emergence of space, time and quantum, by I. Licata. Contemporary Physics, 59(2):208, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRSa

[3795] Andrew Resnick. Book review: Super-resolution imaging in biomedicine, edited by A. Diaspro and M. A. M. J. van Zandvoort. Contemporary Physics, 59(2):208, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2018:BRF

[3796] Michael Belsley. Book review: Fundamentals of nonlinear optics, 2nd edition, by P. E. Powers and J. W. Haus. Contemporary Physics, 59(2): 208–209, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRP

[3797] B. Ishak. Book review: Precision cosmology: the first half million years, by B. J. T. Jones. Contemporary Physics, 59(2):209–210, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRA

[3798] Manuel Vogel. Book review: Atoms, mechanics, and probability: Ludwig Boltzmann’s statistico-mechanical writings — an exegesis, by O. Darrigol.

Contemporary Physics, 59(2):210–211, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zubair:2018:BRP

[3799] Muhammad Zubair. Book review: Physics of electronic materials: principles and applications, by J. Rammer. Contemporary Physics, 59(2): 211, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRAa

[3800] Peter J. Bussey. Book review: Atom projects: events and people, by B. Ioffe. Contemporary Physics, 59(2):211–212, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRO

[3801] B. Ishak. Book review: The origin of mass: elementary particles and fundamental symmetries, by J. Iliopoulos. Contemporary Physics, 59(2): 212–213, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRPa

[3802] Manuel Vogel. Book review: The pull of history: human understanding of magnetism and gravity through the ages, by Y. Yamamoto. Contemporary Physics, 59(2):213, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2018:BRNb

[3803] Miguel A. F. Sanjuán. Book review: Natural complexity: a modelling handbook by P. Charbonneau. Contemporary Physics, 59(2):214, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRM

[3804] B. Ishak. Book review: The Milky Way: an insider’s guide, by W. H. Waller. Contemporary Physics, 59(2):214–215, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2018:BRT

[3805] Stephen H. Ashworth. Book review: Telescopes: a very short introduction, by Geoff Cottrell. Contemporary Physics, 59(2):215–216, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRC

[3806] Miriam R. Aczel. Book review: A concise guide to communication in science & engineering, by D. H. Foster. Contemporary Physics, 59(2): 216–217, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRS

[3807] Miriam R. Aczel. Book review: The simple physics of energy use, by P. Rez. Contemporary Physics, 59(2):217–218, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2018:BRD

[3808] A. H. Harker. Book review: Data analysis techniques for physical scientists, by C. Pruneau. Contemporary Physics, 59(2):218, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2018:BRPb

[3809] A. H. Harker. Book review: Physical methods for materials characterisation, 3rd edition, by P. E. J. Flewitt and R. K. Wild. Contemporary Physics, 59(2):218–219, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2018:BRL

[3810] Tom Lancaster. Book review: The last man who knew everything: the life and times of Enrico Fermi, father of the nuclear age, by D. N. Schwartz. Contemporary Physics, 59(2):219–220, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosas-Ortiz:2018:BRM

[3811] Oscar Rosas-Ortiz. Book review: The meaning of the wave function: in search of the ontology of quantum mechanics, by S. Gao. Contemporary Physics, 59(2):220–221, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRS

[3812] K. Alan Shore. Book review: Symmetry, phase modulation and nonlinear waves, by T. J. Bridges. Contemporary Physics, 59(2):221–222, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRMa

[3813] K. Alan Shore. Book review: Microrheology, by E. M. Furst and T. M. Squires. Contemporary Physics, 59(2):222–223, 2018. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRC

[3814] K. Alan Shore. Book review: Complex networks: principles, methods and applications, by V. Latora, V. Nicosia, and G. Russo. Contemporary Physics, 59(2):223–224, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2018:BRP

[3815] F. W. Taylor. Book review: The paradoxes of time travel, by Ryan Wasserman. Contemporary Physics, 59(2):224, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRD

[3816] B. Ishak. Book review: Deconstructing cosmology, by R. H. Sanders. Contemporary Physics, 59(2):224–225, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRE

[3817] Miriam R. Aczel. Book review: Emmy Noether’s wonderful theorem, revised and updated edition, by D. E. Neunenschwander. Contemporary Physics, 59(2):225–226, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRCa

[3818] Manuel Vogel. Book review: Calorimetry — energy measurements in particle physics, 2nd edition, by R. Wigmans. Contemporary Physics, 59 (2):226–227, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nation:2018:BRQ

[3819] Charlie Nation. Book review: Quantum statistical field theory: an introduction to Schwinger’s variational method with Green’s function nanoapplications, graphene and superconductivity, by N. J. M. Horing. Contemporary Physics, 59(2):227, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gherardini:2018:SSN

[3820] Stefano Gherardini, Shamik Gupta, and Stefano Ruffo. Spontaneous synchronisation and nonequilibrium statistical mechanics of coupled phase oscillators. Contemporary Physics, 59(3):229–250, 2018. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kalirai:2018:SDJ

[3821] Jason Kalirai. Scientific discovery with the James Webb Space Telescope. Contemporary Physics, 59(3):251–290, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Couteau:2018:SPC

[3822] Christophe Couteau. Spontaneous parametric down-conversion. Contemporary Physics, 59(3):291–304, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2018:RHI

[3823] Peter V. E. McClintock and Igor Kh. Kaufman. A review of Handbook of Ion Channels, by Jie Zheng and Matthew C. Trudeau. Contemporary Physics, 59(3):305–307, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:HED

[3824] B. Ishak. High-energy-density physics: foundation of inertial fusion and experimental astrophysics. Contemporary Physics, 59(3):308–310, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Martin:2018:BRB

[3825] Alan D. Martin. Book review: The black book of quantum chromodynamics: a primer for the LHC ERA. Contemporary Physics, 59(3): 311–312, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2018:TDE

[3826] M. P. Gururajan. Theory of dislocations, 3rd edition. Contemporary Physics, 59(3):312–313, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRSb

[3827] Andrew Resnick. Book review: Spin current, 2nd edition. Contemporary Physics, 59(3):313–314, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2018:BRH

[3828] Frank Berkshire. Book review: A half-century of physical asymptotics and other diversions. Contemporary Physics, 59(3):314–315, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRI

[3829] Manuel Vogel. Book review: Introduction to nuclear reactor physics. Contemporary Physics, 59(3):315–316, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2018:BRG

[3830] D. C. Barber. Book review: A guide to mathematical methods for physicists with problems and solutions. Contemporary Physics, 59(3):316, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2018:BRN

[3831] Miguel A. F. Sanjuán. Book review: Nonequilibrium statistical physics: a modern perspective. Contemporary Physics, 59(3):316–317, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2018:BRM

[3832] Peter V. E. McClintock. Book review: Methods in molecular biophysics: structure, dynamics, function for biology and medicine, 2nd edition. Contemporary Physics, 59(3):317–318, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRI

[3833] Andrew Resnick. Book review: Interacting systems far from equilibrium: quantum kinetic theory. Contemporary Physics, 59(3):318, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2018:BRD

[3834] Miriam Aczel. Book review: The dialogues: conversations about the nature of the universe. Contemporary Physics, 59(3):318–319, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2018:BRE

[3835] Stephen J. Blundell. Book review: Elements of slow-neutron scattering. Contemporary Physics, 59(3):319–320, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRT

[3836] Andrew Resnick. Book review: Topological aspects of condensed matter physics: lecture notes of the Les Houches summer school: Volume 103, August 2014. Contemporary Physics, 59(3):320–321, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2018:BRW

[3837] Dr R. S. Shorter. Book review: Wolf prize in physics. Contemporary Physics, 59(3):321, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2018:BRSa

[3838] R. S. Shorter. Book review: Surfing the quantum world. Contemporary Physics, 59(3):321–322, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2018:BRC

[3839] Francisco Caruso. Book review: Can the laws of Physics be unified? Contemporary Physics, 59(3):322–323, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2018:BRM

[3840] Thomas J. Bridges. Book review: The Many Faces of Maxwell, Dirac and Einstein Equations: A Clifford Bundle Approach. Contemporary Physics, 59(3):323, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2018:BRB

[3841] Juan A. Añel. Book review: Blockbuster Science: The Real Science in Science Fiction. Contemporary Physics, 59(3):324, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2018:BRC

[3842] Matt Probert. Book review: Computer simulation of liquids, 2nd edition. Contemporary Physics, 59(3):324–325, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moya-Cessa:2018:BRN

[3843] Héctor Manuel Moya-Cessa. Book review: Non-relativistic quantum mechanics. Contemporary Physics, 59(3):325, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2018:BRO

[3844] Salvatore Esposito. Book review: Origins — the scientific story of Creation. Contemporary Physics, 59(3):325–326, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2018:BRCa

[3845] G. W. Gibbons. Book review: Cosmic paradoxes, 2nd edition. Contemporary Physics, 59(3):326–327, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2018:BRG

[3846] Thomas Peters. Book review: Gravitation. Contemporary Physics, 59 (3):327, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRCb

[3847] Manuel Vogel. Book review: Concepts in particle physics: a concise introduction to the standard model. Contemporary Physics, 59(3):327– 328, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRQb

[3848] Peter J. Bussey. Book review: Quantum sense and nonsense. Contemporary Physics, 59(3):328–329, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2018:BRM

[3849] Mark Nicol. Book review: Mathematics for physics: an illustrated handbook. Contemporary Physics, 59(3):329, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRPb

[3850] Manuel Vogel. Book review: Plasma physics — an introduction to laboratory, space, and fusion plasmas, 2nd edition. Contemporary Physics, 59(3):329–330, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heras:2018:DQC

[3851] Ricardo Heras. Dirac quantisation condition: a comprehensive review. Contemporary Physics, 59(4):331–355, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tarbutt:2018:LCM

[3852] M. R. Tarbutt. Laser cooling of molecules. Contemporary Physics, 59(4): 356–376, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowler:2018:BBM

[3853] David R. Bowler. Building bridges: matching density functional theory with experiment. Contemporary Physics, 59(4):377–390, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRGa

[3854] B. Ishak. Book review: Gravitational waves volume 2: astrophysics and cosmology. Contemporary Physics, 59(4):391–393, 2018. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRQc

[3855] Peter J. Bussey. Book review: Quantum Reality. Contemporary Physics, 59(4):394–397, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2018:BRH

[3856] Gerry Gilmore. Book review: The history of physics — a very short introduction / Theories of everything — ideas in profile. Contemporary Physics, 59(4):398–399, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRH

[3857] Manuel Vogel. Book review: The heroic age, the creation of quantum mechanics, 1925–1940. Contemporary Physics, 59(4):399–400, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2018:BRA

[3858] Gary Gibbons. Book review: Astrophysical applications of gravitational lensing. Contemporary Physics, 59(4):400, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gibbons:2018:BRCb

[3859] Gary Gibbons. Book review: The cosmological singularity. Contemporary Physics, 59(4):400–401, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRI

[3860] B. Ishak. Book review: From interstellar clouds to star-forming galaxies: universal processes? Contemporary Physics, 59(4):401–402, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRG

[3861] B. Ishak. Book review: Galaxies at high redshift and their evolution over cosmic time. Contemporary Physics, 59(4):402, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Menghrajani:2018:BRF

[3862] Kishan Menghrajani. Book review: Fibre-optic Fabry–Perot sensors: an introduction. Contemporary Physics, 59(4):402–403, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosas-Ortiz:2018:BRN

[3863] Oscar Rosas-Ortiz. Book review: Neutron interferometry: lessons in experimental quantum mechanics, wave-particle duality, and entanglement, Second edition. Contemporary Physics, 59(4):403–405, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2018:BRM

[3864] Miguel A. F. Sanjuán. Book review: Modern classical physics: optics, fluids, plasmas, elasticity, relativity, and statistical physics. Contemporary Physics, 59(4):405–406, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRA

[3865] B. Ishak. Book review: Astrophysics of red supergiants. Contemporary Physics, 59(4):406, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zubair:2018:BRF

[3866] Muhammad Zubair. Book review: Fractional diffusion equations and anomalous diffusion. Contemporary Physics, 59(4):406–407, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRG

[3867] Manuel Vogel. Book review: Geophysical fluid dynamics: understanding (almost) everything with rotating shallow water models. Contemporary Physics, 59(4):407–408, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2018:BRS

[3868] Stephen H. Ashworth. Book review: A student’s guide to dimensional analysis. Contemporary Physics, 59(4):408, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2018:BRS

[3869] Marco Casazza. Book review: Signatures of the artist: the vital imperfections that make our universe habitable. Contemporary Physics, 59 (4):409, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2018:BRSb

[3870] R. S. Shorter. Book review: Shoot the Moon: a complete guide to lunar imaging. Contemporary Physics, 59(4):409–410, 2018. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2018:BRQ

[3871] R. S. Shorter. Book review: Quantum physics: what everyone needs to know. Contemporary Physics, 59(4):410, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2018:BRSa

[3872] Stephen J. Blundell. Book review: Skyrmions: topological structures, properties, and applications. Contemporary Physics, 59(4):410–411, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2018:BRSb

[3873] Stephen J. Blundell. Book review: String theory methods for condensed matter physics. Contemporary Physics, 59(4):411, 2018. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2018:BRAa

[3874] Nikolaos E. Myridis. Book review: Advances in imaging and electron physics. Contemporary Physics, 59(4):411–412, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRAb

[3875] Peter J. Bussey. Book review: The age of innocence. Nuclear physics between the First and Second World Wars. Contemporary Physics, 59 (4):413, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BROb

[3876] Manuel Vogel. Book review: One of ten billion Earths — how we learn about our planet’s past and future from distant exoplanets. Contemporary Physics, 59(4):413–414, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2018:BRP

[3877] Francisco Caruso. Book review: Precise dimensions: a history of units from 1791–2018. Contemporary Physics, 59(4):414–415, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2018:BRP

[3878] David L. Andrews. Book review: Priest of nature: the religious worlds of Isaac Newton. Contemporary Physics, 59(4):415–416, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRQb

[3879] Andrew Resnick. Book review: Quantum field theory and condensed matter, an introduction. Contemporary Physics, 59(4):416–417, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2018:BRQc

[3880] Andrew Resnick. Book review: Quantum field theory approach to condensed matter physics. Contemporary Physics, 59(4):417, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2018:BRQd

[3881] Peter J. Bussey. Book review: The quantum measurement problem. Contemporary Physics, 59(4):417–418, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2018:BRAb

[3882] Nikolaos E. Myridis. Book review: Advances in imaging and electron physics. Contemporary Physics, 59(4):418–419, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2018:BRA

[3883] M. P. Gururajan. Book review: Applied computational physics. Contemporary Physics, 59(4):419–420, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2018:BRR

[3884] Eric Howard. Book review: Relativistic many-body theory: a new fieldtheoretical approach. Contemporary Physics, 59(4):420–421, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Athorne:2018:BRFa

[3885] Chris Athorne. Book review: Fifty years of mathematical physics: selected works of Ludwig Fadeev. Contemporary Physics, 59(4):421, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRFb

[3886] B. Ishak. Book review: First ten years of Hinode solar on-orbit observatory. Contemporary Physics, 59(4):422, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRE

[3887] K. Alan Shore. Book review: Essential classical mechanics. Contemporary Physics, 59(4):422–423, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRIb

[3888] K. Alan Shore. Book review: An introduction to organic lasers. Contemporary Physics, 59(4):423–424, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2018:BRH

[3889] Peter J. Dobson. Book review: Handbook on the physics and chemistry of rare earths, Volume 52: Alloy systems and compounds containing rare earth elements and carbon. Contemporary Physics, 59(4):424, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRMb

[3890] K. Alan Shore. Book review: Modern techniques of surface science, 3rd edition. Contemporary Physics, 59(4):424–425, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2018:BRH

[3891] B. Ishak. Book review: Heliophysics: active stars, their astrospheres, and impacts on planetary environments. Contemporary Physics, 59(4): 425, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRPc

[3892] Manuel Vogel. Book review: Physics and technology of sustainable energy. Contemporary Physics, 59(4):425–426, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosas-Ortiz:2018:BRF

[3893] Oscar Rosas-Ortiz. Book review: Fundamentals of interferometric gravitational wave detectors. Contemporary Physics, 59(4):426–428, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2018:BRG

[3894] A. H. Harker. Book review: Geometric and Electronic Properties of Graphene-Related Systems. Contemporary Physics, 59(4):428, 2018. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2018:BRC

[3895] A. H. Harker. Book review: The classical Stefan problem: basic concepts, modelling and analysis with quasi-analytical solutions and methods. Contemporary Physics, 59(4):428–429, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Athorne:2018:BRFb

[3896] Chris Athorne. Book review: Fractals in probability and analysis. Contemporary Physics, 59(4):429, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2018:BRPd

[3897] Manuel Vogel. Book review: Planetary habitability and stellar activity. Contemporary Physics, 59(4):430, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2018:BRR

[3898] Stephen J. Blundell. Book review: A review of magnetism in carbon nanostructures. Contemporary Physics, 59(4):430–431, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2018:BRQ

[3899] Stephen J. Blundell. Book review: Quantum field theory for economics and finance. Contemporary Physics, 59(4):431, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2018:BRG

[3900] John D. Clayton. Book review: Geometric methods in elastic theory of membranes in liquid crystal phases, 2nd ed. Contemporary Physics, 59 (4):431–433, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2018:BRF

[3901] Mario Bertolotti. Book review: Finn’s thermal physics. Contemporary Physics, 59(4):433, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRL

[3902] K. Alan Shore. Book review: The Lattice Boltzmann equation. Contemporary Physics, 59(4):433–434, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRB

[3903] K. Alan Shore. Book review: Berry phases in electronic structure theory. Contemporary Physics, 59(4):434–435, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2018:BRF

[3904] K. Alan Shore. Book review: On the future: prospects for humanity. Contemporary Physics, 59(4):435–436, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2018:BRM

[3905] Thomas Peters. Book review: A modern introduction to dynamical systems. Contemporary Physics, 59(4):436–437, 2018. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Konrad:2019:QMC

[3906] Thomas Konrad and Andrew Forbes. Quantum mechanics and classical light. Contemporary Physics, 60(1):1–22, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mattila:2019:EBL

[3907] Kalevi Mattila and Petri Väisänen. Extragalactic background light: inventory of light throughout the cosmic history. Contemporary Physics, 60 (1):23–44, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Coles:2019:RSE

[3908] Peter Coles. A revolution in science: the eclipse expeditions of 1919. Contemporary Physics, 60(1):45–59, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2019:ERR

[3909] Juan A. Añel. Essay review: Reflections on the Scientific Method at the beginning of the twenty-first century. Contemporary Physics, 60(1): 60–62, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:ERE

[3910] B. Ishak. Essay review: The exoplanet handbook, 2nd edition. Contemporary Physics, 60(1):63–65, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2019:ERQ

[3911] Stephen J. Blundell. Essay review: Quantum field theory lectures of Sidney Coleman. Contemporary Physics, 60(1):66–68, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2019:BRQ

[3912] Mario Bertolotti. Book review: Quantum photonics. Contemporary Physics, 60(1):69, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRF

[3913] Peter J. Bussey. Book review: Facts and mysteries in elementary particle physics. Contemporary Physics, 60(1):69–70, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRPa

[3914] Peter J. Bussey. Book review: A philosophical approach to quantum field theory. Contemporary Physics, 60(1):70–71, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRH

[3915] Eric Howard. Book review: Holographic quantum matter. Contemporary Physics, 60(1):71–72, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRT

[3916] Manuel Vogel. Book review: Theory and experiment in gravitational physics, 2nd edition. Contemporary Physics, 60(1):72, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRIa

[3917] Andrew Resnick. Book review: Introductory quantum physics and relativity, 2nd edition. Contemporary Physics, 60(1):73, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Berkshire:2019:BRB

[3918] Frank Berkshire. Book review: A bouquet of Dyson — and other reflections on science and scientists. Contemporary Physics, 60(1):73–74, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2019:BRU

[3919] John D. Clayton. Book review: Ultracondensed matter by dynamic compression. Contemporary Physics, 60(1):74–75, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRQa

[3920] Eric Howard. Book review: Quantum scaling in many-body systems: an approach to quantum phase transitions, 2nd edition. Contemporary Physics, 60(1):75–76, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2019:BRI

[3921] Robert Sinclair MacKay. Book review: Introduction to accelerator dynamics. Contemporary Physics, 60(1):76–77, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRF

[3922] Andrew Resnick. Book review: Fractional calculus: an introduction for physicists, 3rd edition. Contemporary Physics, 60(1):77, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2019:BRH

[3923] J. Rogel-Salazar. Book review: Hyperspace. Contemporary Physics, 60 (1):77–78, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRMa

[3924] Manuel Vogel. Book review: Marcel Grossmann — for the love of mathematics. Contemporary Physics, 60(1):78–79, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRI

[3925] Eric Howard. Book review: An introduction to quantum theory. Contemporary Physics, 60(1):79, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nicol:2019:BRA

[3926] Mark Nicol. Book review: The astronomers’ magic envelope: an introduction to astrophysics emphasizing general principles and orders of magnitude. Contemporary Physics, 60(1):80, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Benedict:2019:BRQ

[3927] Keith A. Benedict. Book review: Quantum Monte Carlo methods: algorithms for lattice models. Contemporary Physics, 60(1):80–81, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2019:BRU

[3928] Robert C. Smith. Book review: Understanding stellar evolution. Contemporary Physics, 60(1):81–82, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2019:BRSa

[3929] Miguel A. F. Sanjuán. Book review: Some critical questions in biological physics: a guided tour around the bugbears. Contemporary Physics, 60 (1):82–83, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2019:BRS

[3930] M. P. Gururajan. Book review: Symmetry Relationships between Crystal Structures: Applications of Crystallographic Group Theory in Crystal Chemistry. Contemporary Physics, 60(1):83–84, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2019:BRGa

[3931] Salvatore Esposito. Book review: Galileo unbound — a path across life, the universe and everything. Contemporary Physics, 60(1):84–85, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosas-Ortiz:2019:BRL

[3932] Oscar Rosas-Ortiz. Book review: Lagrangian and Hamiltonian dynamics. Contemporary Physics, 60(1):85–86, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2019:BRSb

[3933] Miguel A. F. Sanjuán. Book review: The semiclassical way to dynamics and spectroscopy. Contemporary Physics, 60(1):86–87, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blakely:2019:BRR

[3934] Jonathan Blakely. Book review: Review of application-driven quantum and statistical physics, vol. 1: foundations. Contemporary Physics, 60 (1):87, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRM

[3935] Peter J. Bussey. Book review: From my vast repertoire … Guido Altarelli’s legacy. Contemporary Physics, 60(1):88, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRSa

[3936] B. Ishak. Book review: A student’s guide to Python for physical modeling. Contemporary Physics, 60(1):88–89, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tabish:2019:BRE

[3937] Tanveer A. Tabish. Book review: Electron spin interactions in chemistry and biology. Contemporary Physics, 60(1):89, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRSa

[3938] Peter J. Bussey. Book review: The Soviet atomic project: how the Soviet Union obtained the atomic bomb. Contemporary Physics, 60(1):90, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2019:BRF

[3939] Professor K. Alan Shore. Book review: Friction at the atomic level. Contemporary Physics, 60(1):90–91, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2019:BRM

[3940] K. Alan Shore. Book review: Molecular devices: an introduction to technomimetics and its biological applications. Contemporary Physics, 60(1):91–92, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2019:BRG

[3941] K. Alan Shore. Book review: Graphene photonics. Contemporary Physics, 60(1):92–93, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2019:BRM

[3942] Miriam R. Aczel. Book review: From micro to macro: adventures of a wandering physicist. Contemporary Physics, 60(1):93–94, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2019:BRA

[3943] Thomas J. Bridges. Book review: Analytical mechanics. Contemporary Physics, 60(1):94–95, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2019:BRC

[3944] Stephen J. Blundell. Book review: Collective classical and quantum fields. Contemporary Physics, 60(1):95–96, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BREa

[3945] Peter J. Bussey. Book review: Einstein for the 21st century. His legacy in science, art and modern culture, edited by Peter L. Galison, Gerald James Holton, and S. S. (Silvan S.) Schweber, Princeton, NJ, Princeton University Press, 2018, ISBN 978-0-691-17790-8. Contemporary Physics, 60(1):96–97, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2019:BRS

[3946] Stephen H. Ashworth. Book review: A student’s guide to atomic physics. Contemporary Physics, 60(1):97, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2019:BRA

[3947] John D. Clayton. Book review: Applications of tensor analysis in continuum mechanics, 1st edition. Contemporary Physics, 60(1):97–99, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRCa

[3948] Eric Howard. Book review: Chern–Simons (Super)gravity. Contemporary Physics, 60(1):99–100, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRW

[3949] Andrew Resnick. Book review: The world according to Quantum Mechanics; why the laws of physics make perfect sense after all, 2nd edition. Contemporary Physics, 60(1):100, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2019:BRP

[3950] Nikolaos E. Myridis. Book review: Physical perspectives on computation, computational perspectives on physics. Contemporary Physics, 60 (1):100–101, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zubair:2019:BRM

[3951] Muhammad Zubair. Book review: Medical and biological microwave sensors and systems. Contemporary Physics, 60(1):101–102, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Moya-Cessa:2019:BRT

[3952] Héctor Manuel Moya-Cessa. Book review: Theoretical and mathematical physics: problems and solutions. Contemporary Physics, 60(1):102, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Al-Tabbakh:2019:BRP

[3953] Ahmed A. Al-Tabbakh. Book review: Principles of electron optics: volume one: basic geometrical optics & volume two: applied geometrical optics. Contemporary Physics, 60(1):103–104, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRQb

[3954] Eric Howard. Book review: Quantum gravity and the functional renormalization group: the road towards asymptotic safety. Contemporary Physics, 60(1):104–105, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRSb

[3955] Peter J. Bussey. Book review: A student’s guide to general relativity/ General relativity, a concise introduction. Contemporary Physics, 60(1): 105–106, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRFa

[3956] Manuel Vogel. Book review: Fireworks in a dark universe, by Amir Levinson, London, World Scientific Europe, 2018, 316 pp., £70.00 (hardback), ISBN: 978-1-78634-511-0. Contemporary Physics, 60(1):106, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRP

[3957] B. Ishak. Book review: Physics of magnetic flux tubes, 2nd edition, by M. Ryutova, Switzerland, Springer, 2018, 747 pp., £159.99 (hardback), ISBN: 978-3-319-96360-0. Contemporary Physics, 60(1):106–107, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2019:BRO

[3958] Gerry Gilmore. Book review: Our Cosmic Habitat. Contemporary Physics, 60(1):107–108, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRG

[3959] Andrew Resnick. Book review: Group theory in physics: a practitioner’s guide. Contemporary Physics, 60(1):108, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2019:BRG

[3960] A. H. Harker. Book review: Group theory in solid state physics and photonics: problem solving with Mathematica. Contemporary Physics, 60(1):108–109, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2019:BRA

[3961] Peter J. Dobson (retired). Book review: Applied thermal measurements at the nanoscale: a beginners guide to electrothermal methods. Contemporary Physics, 60(1):109, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2019:BRS

[3962] A. H. Harker. Book review: Surface and interfacial forces, 2nd edition. Contemporary Physics, 60(1):109–110, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Smith:2019:SBH

[3963] Aaron Smith and Volker Bromm. Supermassive black holes in the early universe. Contemporary Physics, 60(2):111–126, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schnack:2019:LMM

[3964] Jürgen Schnack. Large magnetic molecules and what we learn from them. Contemporary Physics, 60(2):127–144, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wise:2019:CR

[3965] John H. Wise. Cosmic reionisation. Contemporary Physics, 60(2):145– 163, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mitchison:2019:QTA

[3966] Mark T. Mitchison. Quantum thermal absorption machines: refrigerators, engines and clocks. Contemporary Physics, 60(2):164–187, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosas-Ortiz:2019:BDU

[3967] Oscar Rosas-Ortiz. Beyond the dynamical universe. Unifying block universe physics and time as experienced. Contemporary Physics, 60(2): 188–190, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BS

[3968] Peter J. Bussey. Believing without seeing? Contemporary Physics, 60(2): 191–194, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:PPQ

[3969] Peter J. Bussey. Philosophy of physics: quantum theory. Contemporary Physics, 60(2):195–196, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:IBS

[3970] B. Ishak. The impact of binary stars on stellar evolution. Contemporary Physics, 60(2):196, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2019:PSQ

[3971] Mario Bertolotti. PT symmetry in quantum and classical physics. Contemporary Physics, 60(2):196–197, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Modi:2019:QMB

[3972] Kavan Modi. Quantum many-body physics in a nutshell. Contemporary Physics, 60(2):197, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ooi:2019:BRW

[3973] Kelvin J. A. Ooi. Book review: The wonders of physics, 4th edition. Contemporary Physics, 60(2):197–198, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2019:BRI

[3974] Michael Belsley. Book review: An introduction to quantum physics: a first course for physicists, chemists, materials scientists, and engineers. Contemporary Physics, 60(2):198–199, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRP

[3975] Eric Howard. Book review: Principles of quantum computation and information: a comprehensive textbook. Contemporary Physics, 60(2): 199–200, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tabish:2019:BRM

[3976] Tanveer A. Tabish. Book review: Memorial volume on Abdus Salam’s 90th birthday. Contemporary Physics, 60(2):200–201, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tabish:2019:BRA

[3977] Tanveer A. Tabish. Book review: Applied nanophotonics. Contemporary Physics, 60(2):201, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Neukirch:2019:BRM

[3978] Thomas Neukirch. Book review: Magnetohydrodynamics of laboratory and astrophysical plasmas. Contemporary Physics, 60(2):202, 2019. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2019:BRU

[3979] R. S. Shorter. Book review: Understanding lasers: an entry level guide, 4th ed. Contemporary Physics, 60(2):202–203, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2019:BRT

[3980] Tom Lancaster. Book review: Theoretical fluid mechanics. Contemporary Physics, 60(2):203–204, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRS

[3981] Manuel Vogel. Book review: The story of antimatter: matter’s vanished twin. Contemporary Physics, 60(2):204, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2019:BRS

[3982] Peter V. E. McClintock. Book review: Size really does matter: the nanotechnology revolution. Contemporary Physics, 60(2):204–205, 2019. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2019:BRP

[3983] M. P. Gururajan. Book review: Principles of thermodynamics. Contemporary Physics, 60(2):205–206, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2019:BRL

[3984] Stephen J. Blundell. Book review: The life, science and times of Lev Vasilevich Shubnikov. Contemporary Physics, 60(2):206–207, 2019. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tennyson:2019:BRB

[3985] Jonathan Tennyson. Book review: A brief history of astronomy and astrophysics. Contemporary Physics, 60(2):207, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2019:BRA

[3986] Robert Appleby. Book review: Accelerator physics, 4th edition. Contemporary Physics, 60(2):208, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2019:BRF

[3987] Gerard Gilmore. Book review: The formative years of relativity: the history and meaning of Einstein’s Princeton lectures. Contemporary Physics, 60(2):208–209, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Appleby:2019:BRB

[3988] Rob Appleby. Book review: Beams and accelerators with MATLAB, (with companion media pack). Contemporary Physics, 60(2):209, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2019:BRG

[3989] Gerry Gilmore. Book review: Grand challenges for science in the 21st century. Contemporary Physics, 60(2):210, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Willetts:2019:SCC

[3990] Katherine Willetts, Lance Farr, and Liberty Foreman. From stellar composition to cancer diagnostics. Contemporary Physics, 60(3):211–225, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Roffe:2019:QEC

[3991] Joschka Roffe. Quantum error correction: an introductory guide. Contemporary Physics, 60(3):226–245, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2019:SMM

[3992] Tom Lancaster. Skyrmions in magnetic materials. Contemporary Physics, 60(3):246–261, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRIa

[3993] B. Ishak. Book review: Introduction to astrochemistry: chemical evolution from interstellar clouds to star and planet formation, by Yamamoto S., Japan, Springer, 2018, 286 pp., £99.99 (hardback), ISBN: 978-4-431-54170-7. Contemporary Physics, 60(3):262, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRMa

[3994] B. Ishak. Book review: Mathematics for physicists: introductory concepts and methods, by A. Altland and J. von Delft, Cambridge, Cambridge University Press, 2019, 720 pp., £39.99 (hardback), ISBN: 978-1-108-47122-0. Contemporary Physics, 60(3):262, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2019:BRM

[3995] Miguel A. F. Sanjuán. Book review: Mathematical methods in science and engineering, 2nd edition, by Selçuk S. Bayin, Hoboken, NJ, Wiley, 2018, 864 pp., £125.00 (hardback), ISBN 978-111-9425-39-7. Contemporary Physics, 60(3):263, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2019:BRA

[3996] Miguel A. F. Sanjuán. Book review: Applied symbolic dynamics and chaos, 2nd edition, B. Hao and W. Zheng, Singapore, World Scientific, 2018, 500 pp., $168.00 (hardback), ISBN 978-981-323-642-4. Contemporary Physics, 60(3):263–264, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRMb

[3997] Manuel Vogel. Book review: Michelson–Morley experiments — an enigma for physics and the history of science, by Maurizio Consoli and Alessandro Pluchino, Singapore, World Scientific, 2018, 196 pp., £70.00 (hardback), ISBN: 978-981-3278-18-9. Contemporary Physics, 60(3):264– 265, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lakshminarayanan:2019:BRB

[3998] Vasudevan Lakshminarayanan. Book review: Biophotonics science and technology, by Yin Yeh and V. V. Krishnan, Singapore, World Scientific, 2018, 292 pp., £95.00 (hardback), ISBN: 978-981-323-568-7. Contemporary Physics, 60(3):265–266, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRR

[3999] Eric Howard. Book review: Relativistic quantum mechanics and quantum fields, 2nd edition, by W. Y. P. Hwang and T.-Y. Wu, Singapore, World Scientific, 2018, 402 pp., £130.00 (hardback), ISBN 978-981-327-002-2. Contemporary Physics, 60(3):266–267, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blakely:2019:BRI

[4000] Jonathan Blakely. Book review: An introduction to quantum optics and quantum fluctuations, by P. W. Milonni, New York, Oxford University Press, 2019, 544 pp., £65.00 (hardback), ISBN: 978-0-19-921561-4. Contemporary Physics, 60(3):267–268, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRMa

[4001] Eric Howard. Book review: Modern condensed matter physics, S. M. Girvin and K. Yang, Cambridge, Cambridge University Press, 2019, 714 pp., £59.99 (hardback), ISBN: 978-1-107-13739-4. Contemporary Physics, 60(3):268, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRCb

[4002] Eric Howard. Book review: Classical field theory, by Horaţiu Năstase, Cambridge, Cambridge University Press, 2019, 436 pp., £59.99 (hardback), ISBN: 978-1-108-47701-7. Contemporary Physics, 60(3):269, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRSb

[4003] B. Ishak. Book review: Solving Fermi’s paradox, by D. H. Forgan, Cambridge, Cambridge University Press, 2019, 426 pp., £120 (hardback), ISBN: 978-1-107-16365-2. Contemporary Physics, 60(3):269–271, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRC

[4004] Manuel Vogel. Book review: Chaos in nature, 2nd edition, by Christophe Letellier, Singapore, World Scientific, 2019, 436 pp., £130.00 (hardback), ISBN: 978-981-12-0119-6. Contemporary Physics, 60(3):271–272, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRD

[4005] Peter J. Bussey. Book review: Do we really understand quantum mechanics? 2nd edition, by Franck Laloë, Cambridge, Cambridge University Press, 2019, pp. xviii + 530, £44.99 (hardback), ISBN 978-1-108-47700-0. Contemporary Physics, 60(3):272–273, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2019:BRO

[4006] A. H. Harker. Book review: Optical effects in solids, by David B. Tanner, Cambridge, Cambridge University Press, 2019, 12 + 400 pp., £64.99 (hardback), ISBN: 978-1-107-16014-9. Contemporary Physics, 60(3):273– 274, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2019:BRA

[4007] Peter V. E. McClintock. Book review: Asteroseismic data analysis: foundations and techniques, by Sarbani Basu and William J. Chaplin, Princeton, Princeton University Press, 2017, 337 pp., £58.00 (hardback), ISBN: 978-0-691-16292-8. Contemporary Physics, 60(3):274, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRTa

[4008] Eric Howard. Book review: Theory of groups and symmetries: Finite groups, Lie groups and Lie algebras, by A. P. Isaev and V. A. Rubakov, World Scientific, 2019,458 pp., £121.00 (hardback), ISBN 978-981-323-685-1. Contemporary Physics, 60(3):275, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2019:BRS

[4009] Stephen J. Blundell. Book review: Solid state physics, volume 68, edited by R. E. Camley and R. L. Stamps, Cambridge, MA, Academic Press, 2017, 182 pp., £161.50 (hardback), ISBN 978-0-12-811991-4. . Contemporary Physics, 60(3):275–276, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2019:BRC

[4010] Juan A. Añel and Alan Robock. Book review: Crossing the red line: the nuclear option, by G. E. Marsh, Singapore, World Scientific, 2018, 252 pp., $68.00, (hardback), ISBN 978-981-3276-82-6. Contemporary Physics, 60(3):276, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2019:Ca

[4011] Anonymous. Correction. Contemporary Physics, 60(3):277, 2019. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anonymous:2019:Cb

[4012] Anonymous. Correction. Contemporary Physics, 60(3):278, 2019. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tsang:2019:RSQ

[4013] Mankei Tsang. Resolving starlight: a quantum perspective. Contemporary Physics, 60(4):279–298, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Brown:2019:EQM

[4014] Harvey R. Brown. Everettian quantum mechanics. Contemporary Physics, 60(4):299–314, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:ERA

[4015] B. Ishak. Essay review: The atlas of Mars: mapping its geography and geology, edited by K. S. Coles, K. L. Tanaka & P. R. Christensen, Cambridge, Cambridge University Press, 2019, 300 pp., £39.99, (hardback) ISBN: 978-1-107-03629-1. Contemporary Physics, 60(4):315–317, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2019:BRW

[4016] Stephen H. Ashworth. Book review: Waves: a very short introduction, by Mike Goldsmith, Oxford, Oxford University Press, 2018, £7.99 (paperback), 135 pp., ISBN 978-0-19-890378-9. Contemporary Physics, 60 (4):318, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McClintock:2019:BRI

[4017] Peter V. E. McClintock. Book review: Introduction to the theory of complex systems, by Stefan Thurner, Rudolf Hanel and Peter Klimek, Oxford, Oxford University Press, 2018, 431 pp., £49.99 (hardback), ISBN 978-0-19-882193-9. Contemporary Physics, 60(4):318–319, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRPb

[4018] Peter J. Bussey. Book review: Particles, fields and forces, by Wouter Schmitz, Switzerland, Springer, 2019, xi + 322 pp., £59.99 (hardback), ISBN 978-3-030-12877-7. Contemporary Physics, 60(4):319–320, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2019:BRD

[4019] Thomas Peters. Book review: Data-driven science and engineering: machine learning, dynamical systems, and control, by S. L. Brunton and J. N. Kutz, 2019, Cambridge, Cambridge University Press, 472 pp., £49.99 (hardback), ISBN 978-1-108-42209-3. Contemporary Physics, 60 (4):320, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRIb

[4020] Andrew Resnick. Book review: Imaging optics, by J. Braat and P. Török, Cambridge, Cambridge University Press, 2019, 984 pp., £79.99 (hardback), ISBN 978-1-108-42808-8. Contemporary Physics, 60(4):321, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2019:BRQ

[4021] A. H. Harker. Book review: Quantum theory of materials, by E. Kaxiras and J. D. Joannopoulos, Cambridge, UK, Cambridge University Press, 2019, 653 pp., £64.99 (hardback), ISBN 978-052-1117-11-1. Contemporary Physics, 60(4):321–322, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRIc

[4022] Andrew Resnick. Book review: Introductory quantum mechanics with MATLAB for atoms, molecules, clusters and nanocrystals, by J. R. Chelikowsky, Weinheim, Wiley-VCH Verlag GmbH & Co. KGaA, 2018, 224 pp., £70.00 (paperback), ISBN: 978-3-527-40926-6. Contemporary Physics, 60(4):322, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRN

[4023] Manuel Vogel. Book review: No shadow of a doubt, by D. Kennefick, Princeton, Princeton University Press, 2019,416 pp., £24.00 (hardcover), ISBN 978-0-691-18386-2. Contemporary Physics, 60(4):322–323, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRTb

[4024] Eric Howard. Book review: Topology and physics, by C. N. Yang, M. L. Ge and Y. H. He, Singapore, World Scientific, 2019,232 pp., $28.00 (paperback), ISBN 978-981-327-850-9. Contemporary Physics, 60(4):323– 324, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRA

[4025] Eric Howard. Book review: Antimatter, Frank Close, Oxford, Oxford University Press, 2018,166 pp., £8.99 (paperback), ISBN 978-0-19-883191-4. Contemporary Physics, 60(4):324–325, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2019:BRT

[4026] M. P. Gururajan. Book review: Thermodynamics: fundamentals and engineering applications, by W. C. Reynolds and P. Colonna, Cambridge, Cambridge University Press, 2018, 401 pp., £84.99 (hardback), ISBN: 978-0-521-86273-8. Contemporary Physics, 60(4):325, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2019:BRN

[4027] Francisco Caruso. Book review: Newton the alchemist: science, enigma, and the quest for nature’s ‘secret fire’, by William R. Newman, Princeton and Oxford, Princeton University Press, 2019, 560 pp., US$159 or £30 (hardcover), ISBN 978-0-691-17487-7. Contemporary Physics, 60(4): 326–327, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BRT

[4028] Peter J. Bussey. Book review: TASI: 2016, anticipating the next discoveries in particle physics, edited by Rouven Essig and Ian Low, Singapore, World Scientific, 2018, 623 pp., £175.00 (hardback), ISBN 978-981-3233-331. Contemporary Physics, 60(4):327–328, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRSa

[4029] Eric Howard. Book review: String fields, higher spins and number theory, by Polyakov Dimitri, Singapore, World Scientific, 2018, 170 pp., $88.00 (hardcover), ISBN 978-981-3233-39-3. Contemporary Physics, 60 (4):328–329, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2019:BRGb

[4030] Salvatore Esposito. Book review: George Placzek — a nuclear physicist’s Odyssey, by A. Gottvald and M. Shifman, Singapore, World Scientific, 2018, 369 pp., £85.00 (hardback), ISBN 978-981-323-691-2. Contemporary Physics, 60(4):329–330, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRE

[4031] B. Ishak. Book review: Encyclopedic atlas of terrestrial impact craters, edited by E. Flamini, M. Di Martino and A. Coletta, Cham, Springer, 2019, 691 pp., £149.99 (hardback), ISBN: 978-3-030-05449-6. Contemporary Physics, 60(4):330, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2019:BRR

[4032] Stephen J. Blundell. Book review: A review of modern ophthalmic optics, edited by José Alonso, José A. Gómez-Pedrero and Juan A. Quiroga, Cambridge, Cambridge University Press, 2019, 546 pp., £65 (hardback), ISBN 978-1-107-11074-8. Contemporary Physics, 60(4):330–331, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRFb

[4033] Manuel Vogel. Book review: Formation of the first black holes, edited by M. Latif and D. Schleicher, Singapore, World Scientific, 2019, 365 pp., £105.00 (hardback), ISBN: 978-981-322-794-1. Contemporary Physics, 60(4):331–332, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tabish:2019:BRI

[4034] Tanveer A. Tabish. Book review: Introductory physics for biological scientists, by Christof M. Aegerter, Cambridge, Cambridge University Press, 2018, 451 pp., £54.99 (paperback), ISBN: 978-1-108-46650-9. Contemporary Physics, 60(4):332, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2019:BRC

[4035] Thomas J. Bridges. Book review: Classical field theory, by Joel Franklin, Cambridge, UK, Cambridge University Press, 2017, 216 pp., £49.85 (hardback), ISBN 978-1-107-18961-4. Contemporary Physics, 60(4):332– 333, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2019:BRId

[4036] Andrew Resnick. Book review: An introduction to general relativity: spacetime and geometry, by S. Carroll, Cambridge, Cambridge University Press, 2019, 513 pp., £49.99 (hardback), ISBN 978-110-8488-39-6. Contemporary Physics, 60(4):333–334, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRFa

[4037] B. Ishak. Book review: Finding our place in the solar system: the scientific story of the Copernican revolution, by T. Timberlake and P. Wallace, Cambridge, Cambridge University Press, 2019, 378 pp., £29.99 (hardback), ISBN 978-1-107-18229-5. Contemporary Physics, 60(4):334, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blakely:2019:BRE

[4038] Jonathan Blakely. Book review: Electromagnetic radiation, by R. R. Freeman, J. A. King, and G. P. Lafyatis, 2019, Oxford, UK, Oxford University Press, 640 pp., £50.00 (hardback), ISBN 978-019-8726-50-0. Contemporary Physics, 60(4):334–335, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2019:BRE

[4039] Manuel Vogel. Book review: Exploring planetary climate, by R. D. Lorenz, Cambridge, Cambridge University Press, 2019, 320 pp., £42.99 (hardback), ISBN 978-1-108-47154-1. Contemporary Physics, 60(4):335, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRIb

[4040] B. Ishak. Book review: An introduction to radio astronomy, 4th edition, by B. F. Burke, F. Graham-Smith and P. N. Wilkinson, Cambridge, Cambridge University Press, 2019, 540 pp., £59.99 (hardback), ISBN 978-1-107-18941-6. Contemporary Physics, 60(4):335–336, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRQ

[4041] Eric Howard. Book review: Quantum fields: from the Hubble to the Planck Scale, by M. Kachelriess, Oxford, Oxford University Press, 2017, 544 pp., £55 (hardback), ISBN 978-019-8802-87-7. Contemporary Physics, 60(4):336, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRS

[4042] B. Ishak. Book review: Saturn in the 21st century, by K. H. Baines, F. M. Flaser, N. Krupp and T. Stallard, Cambridge, Cambridge University Press, 2018, 448 pp., £120.00 (hardback), ISBN 978-110-7106-77-2. Contemporary Physics, 60(4):337, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRN

[4043] Eric Howard. Book review: Nonlocal gravity, by B. Mashoon, Oxford, Oxford University Press, 2017, 256 pp., £65 (hardback), ISBN 978-019-8803-80-5. Contemporary Physics, 60(4):337–338, 2019. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRSb

[4044] Eric Howard. Book review: Scalar field cosmology, by Chervon, S., Fomin, I., Yurov, V., Yurov, A., Singapore, World Scientific, 2019, 263 pp., £85.00 (hardback), ISBN 978-981-12-0507-1. Contemporary Physics, 60(4):338–339, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2019:BREb

[4045] Peter J. Bussey. Book review: Elementary particle physics, an intuitive introduction, by Andrew J. Larkoski, Cambridge, Cambridge University Press, 2019, xx + 488 pp., £45 (hardback), ISBN 978-1-108-49698-8. Contemporary Physics, 60(4):339, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2019:BRMb

[4046] Eric Howard. Book review: Magnetism and magnetic materials, by J. M. D. Coey, Cambridge, Cambridge University Press, 2018, 617 pp., £59.99 (paperback), ISBN 978-1-108-71751-9. Contemporary Physics, 60 (4):339–340, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRFb

[4047] B. Ishak. Book review: Feynman and his physics: the life and science of an extraordinary man, by Jörg Resag, Cham, Springer, 2018, 319 pp., £24.99 (hardback), ISBN: 978-3-319-96835-3. Contemporary Physics, 60 (4):340–341, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRMb

[4048] B. Ishak. Book review: Mercury: the view after MESSENGER, edited by S. C. Solomon, L. R. Nittler and B. J. Anderson, Cambridge, Cambridge University Press, 2018, 583 pp., £42.99 (hardback), ISBN: 978-1-107-15445-2. Contemporary Physics, 60(4):341, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2020:SES

[4049] R. S. MacKay. Sustainable energy storage — with hot air, or cold air or liquid air. Contemporary Physics, 61(1):1–11, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hu:2020:SVP

[4050] Huayu Hu. Seed and vacuum pair production in strong laser field. Contemporary Physics, 61(1):12–25, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lewalle:2020:MFT

[4051] Philippe Lewalle, Sreenath K. Manikandan, Cyril Elouard, and Andrew N. Jordan. Measuring fluorescence to track a quantum emitter’s state: a theory review. Contemporary Physics, 61(1):26–50, 2020. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2020:BRS

[4052] M. P. Gururajan. Book review: Statistical thermodynamics: an engineering approach, by J. W. Daily, Cambridge, Cambridge University Press, 2019, 263 pages, £72.99 (Hardback), ISBN: 978-1-108-41531-6. Scope: textbook. Level: postgraduate, early career researcher. Contemporary Physics, 61(1):51, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2020:BRP

[4053] Andrew Resnick. Book review: Parameter estimation and inverse problems, 3rd edition. R. C. Aster, B. Borchers and C. H. Thurber, Cambridge, Elsevier, 2019, 404 pp., £76.95 (paperback), ISBN: 978-0-12-804651-7. Scope: monograph. Level: advanced undergraduate, postgraduate. Contemporary Physics, 61(1):51–52, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2020:BRA

[4054] Andrew Resnick. Book review: Airless bodies of the inner solar system, by J. A. Grier, A. S. Rivkin, Cambridge, Elsevier, 2018, 283 pp., £76.95 (Paperback), ISBN 978-0-12-809279-8. Scope: monograph. Level: postgraduate, advanced undergraduate, early career researcher, teacher, scientist. Contemporary Physics, 61(1):52–53, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2020:BRP

[4055] Stephen J. Blundell. Book review: Phase transitions for beginners, by S. M. Stishov, Singapore, World Scientific, 2018, 152 pp., £40 (hardback), ISBN: 978-981-327-417-4. Scope: textbook. Level: postgraduate. Contemporary Physics, 61(1):53, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2020:BRG

[4056] Stephen J. Blundell. Book review: Group theory for physicists, 2nd edition, by Z. Ma, Singapore, World Scientific, 2019, 655 pp., £115.00 (hardback), ISBN: 978-981-327-738-0. Scope: monograph. Level: researcher. Contemporary Physics, 61(1):53–54, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Anel:2020:BRS

[4057] Juan A. Añel. Book review: Selling science: how to use business skills to win support for scientific research, by Steven Judge and Richard Lucas, London, World Scientific, 2019, 98pp, £40.00 (hardback), ISBN 978-1-78634-572-1. Scope: general interest. Level: general readership. Contemporary Physics, 61(1):54, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2020:BRMa

[4058] Manuel Vogel. Book review: Meteoroids: sources of meteors on Earth and beyond, dited by Galina O. Ryabova, David J. Asher, and Margaret D. Campbell-Brown, Cambridge, Cambridge University Press, 2019, 318 pp., £110 (hardback), ISBN: 978-110-8426-71-8. Scope: edited book. Level: advanced undergraduates, postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 61(1): 54–55, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Taylor:2020:BRM

[4059] Fred Taylor. Book review: Mysteries of Mars, by F. V. De Blasio, Berlin, Springer-Praxis, 2019, 189 pp., £22.99 (paperback), ISBN: 978-3-319-74783-5. Scope: monograph. Level: general. Contemporary Physics, 61 (1):55, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ali:2020:BRL

[4060] Muhammad Mahmood Ali. Book review: Lightwave communications, by G. C. Papen and R. Blahut, Cambridge, Cambridge University Press, 2019, 970 pp., £79.99 (hardback), ISBN 978-1-108-42756-2. Scope: review, monograph, handbook, guide, textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 61(1):55–56, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2020:BRT

[4061] John D. Clayton. Book review: Tensor-valued random fields for continuum physics, 1st edition, by Anatoliy Malyarenko and Martin Ostoja-Starzewski, Cambridge Monographs on Mathematical Physics, Cambridge, Cambridge University Press, 2019, 310 pp., $145 USD (hardback), ISBN 978-1-108-42985-6. Scope: monograph. Level: postgraduate, specialist. Contemporary Physics, 61(1):56–58, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2020:BRIa

[4062] Stephen H. Ashworth. Book review: Introduction to electricity and magnetism, by J. D. Walecka, Singapore, World Scientific, 2018, 272 pp., £75 (hardback), ISBN: 978-981-327-206-4, Scope: review. Level: undergraduate. Specialist. Contemporary Physics, 61(1):58–59, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2020:BRL

[4063] K. Alan Shore. Book review: Lectures on astrophysics, by S. Weinberg, Cambridge, Cambridge University Press, 2020, 226 pp., £34.99 (hardback), ISBN 978-1-108-41507-1. Scope: textbook. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 61(1): 59–60, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2020:BRG

[4064] K. Alan Shore. Book review: Gravity! The quest for gravitational waves, by Pierre Binêtruy, Oxford, Oxford University Press, 2018, 245 pp., 19.99 (hardback), ISBN: 978-0-19-879651-0. Scope: monograph. Level: general readership. Contemporary Physics, 61(1):60–61, 2020. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2020:BRC

[4065] K. Alan Shore. Book review: Cosmic discovery: the search, scope and heritage of astronomy, by M. Harwit, Cambridge, Cambridge University Press, 2019, 348 pp., £34.99 (paperback), ISBN: 978-1-108-72204-9. Scope: monograph. Level: general readership. Contemporary Physics, 61 (1):61–62, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scali:2020:BRI

[4066] Stefano Scali. Book review: Introduction to quantum mechanics, 3rd edition, by D.J. Griffiths and D.F. Schroeter, Cambridge, Cambridge University Press, 2018, 495 pp., £54.99 (hardback), ISBN: 978-1-107-18963-8. Scope: textbook. Level: undergraduate, advanced undergraduate. Contemporary Physics, 61(1):62, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2020:BRN

[4067] Manuel Vogel. Book review: New perspectives on Einstein’s E = mc^{2}, World scientific, by Y. S. Kim and M. E. Noz, Singapore, World Scientific, 2018, 200 pp., $78.00 (hardback), ISBN 978-981-3237-70-4. Scope: monograph. Level: researcher, specialist, scientist. Contemporary Physics, 61 (1):62–63, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2020:BRP

[4068] Eric Howard. Book review: Physics on Ultracold quantum gases, by Y. Han, W. Yi and W. Zhang, Singapore, World Scientific, 2018, 288, $98.00 (Hardback), ISBN: 978-981-327-075-6. Scope: review. Level: postgraduate, early career researcher, researcher, specialist, scientist, engineers. Contemporary Physics, 61(1):63–64, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2020:BRS

[4069] Peter J. Bussey. Book review: Schrödinger’s mechanics: interpretation, by D. B. Cook, Singapore, World Scientific, 2019, 196 pp., £69 (hardback), ISBN: 978-1-78634-490-8. Scope: monograph; Level: specialist. Contemporary Physics, 61(1):64, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scali:2020:BRG

[4070] Stefano Scali. Book review: A guide to mathematical methods for physicists — advanced topics and applications, by M. Petrini, G. Pradisi, A. Zaffaroni, London, World Scientific, 2018, 308 pp., $48.00 (paperback), ISBN 978-1-78634-704-6. Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 61(1):64–65, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2020:BRV

[4071] Miguel A. F. Sanjuán. Book review: Vibrations, 3rd edition, by Balakumar Balachandran and Edward B. Magrab, Cambridge, UK, Cambridge University Press, 2018, 723 pp., £87.99 (hardback), ISBN: 978-1-108-42731-9. Scope: textbook, reference. Level: undergraduate, advanced undergraduate, postgraduate, teacher, engineers. Contemporary Physics, 61(1):65–66, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2020:BRM

[4072] M. P. Gururajan. Book review: A modern course in transport phenomena, by D. C. Venerus and H. C. Öttinger, Cambridge, UK, Cambridge University Press, 2018, 508 pp., £74.99 (hardback), ISBN 978-1-107-12920-7. Scope: reference, textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 61(1):66–67, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2020:BRN

[4073] Peter J. Bussey. Book review: Niels Bohr: a very short introduction, by J. L. Heilbron, Oxford, Oxford University Press, 2020, 160 pp., $11.95 (paperback), ISBN 978-0-19-881926-4. Scope: general interest, biography. Level: scientist. Contemporary Physics, 61(1):67–68, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2020:TC

[4074] G. J. Milburn. The thermodynamics of clocks. Contemporary Physics, 61 (2):69–95, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Powell:2020:EPG

[4075] Ben J. Powell. Emergent particles and gauge fields in quantum matter. Contemporary Physics, 61(2):96–131, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lahav:2020:DEI

[4076] Ofer Lahav. Dark energy: is it ‘just’ Einstein’s cosmological constant Λ? Contemporary Physics, 61(2):132–145, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scali:2020:BRP

[4077] Stefano Scali. Book review: Physical mathematics, 2nd edition, by Kevin Cahill, Cambridge, UK, Cambridge University Press, 2019, 757 pp., £64.99 (hardback), ISBN 978-1-108-47003-2. Scope: textbook. Level: undergraduate, advanced undergraduate, postgraduate, researcher. Contemporary Physics, 61(2):146, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ali:2020:BRE

[4078] Muhammad Mahmood Ali. Book review: Experimental methods for science and engineering students: an introduction to the analysis and presentation of data, 2nd edition. L. Kirkup, Cambridge, Cambridge University Press, 2019, 239 pp., £37.99 (hardback), ISBN: 978-1-108-41846-1. Scope: monograph, handbook, guide, textbook. Level: undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 61(2):146–147, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2020:BRX

[4079] Andrew Resnick. Book review: “X-ray Microscopy”, by Chris Jacobsen, Cambridge, UK, Cambridge University Press, University Printing House, 2020, $135, 579 pp., ISBN: 978-1-107-07657-0. Scope: Monograph. Level: advanced undergraduate student, graduate student. Contemporary Physics, 61(2):147–148, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2020:BRW

[4080] Andrew Resnick. Book review: Waves and rays in seismology: answers to unasked questions, 2nd edition, by Michael A. Slawinski, Singapore, World Scientific, 2018, 576 pp., £140 (hardback), ISBN: 978-981-3239-87-6. Scope: monograph. Level: advanced undergraduate, postgraduate. Contemporary Physics, 61(2):148, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2020:BRI

[4081] A. H. Harker. Book review: Introduction to graphene-based nanomaterials: from electronic structure to quantum transport, 2nd edition, by Luis E. F. Foa Torres, Stephan Roche and Jean-Christophe Charlier, Cambridge, UK, Cambridge University Press, 2020, 15pp. + 461pp., £74.99 (hardback), ISBN: 978-1-108-47699-7. Scope: textbook. Level: early stage researcher, postgraduate. Contemporary Physics, 61(2):149, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2020:BRC

[4082] David C. Barber. Book review: The cosmic mystery tour, by N. Mee, Oxford, Oxford University Press, 2019, 207 pp., £16.99 (hardback), ISBN: 978-019-8831-86-0. Scope: review. Level: general readership. Contemporary Physics, 61(2):149–150, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2020:BRE

[4083] Andrew Resnick. Book review: The essence of crystallography, by Mark Ladd, Singapore, World Scientific, 2020, 532 pp., £105 (hardback), ISBN: 978-178-6346315. Level: advanced undergraduate, postgraduate, early career researcher. Scope: textbook. Contemporary Physics, 61(2):150, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2020:BRP

[4084] Salvatore Esposito. Book review: Physics problems for aspiring physical scientists and engineers, by K. Riley, Cambridge, UK, Cambridge University Press, 2019, 335 pp., £24.99 (paperback), ISBN: 978-1-1087-0130-3. Scope: reference. Level: pre-university, undergraduate, teacher. Contemporary Physics, 61(2):150–151, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2020:BRL

[4085] Eric Howard. Book review: Lectures on the infrared structure of gravity and gauge theory, by A. Strominger, Princeton, New Jersey, Princeton University Press, 2018, 200 pp., £41.95 (Paperback), ISBN 978-0-691-17973-5. Scope: review. Level: postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 61(2):151–152, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2020:BRC

[4086] R. S. Shorter. Book review: Competitive physics: mechanics and waves, by J. Wang and B. Ricardo, Singapore, World Scientific, 2018, 832 pp., $68.00 (paperback), ISBN: 978-981-323-518-2. Scope: general interest, textbook. Level: general readership, pre-university, undergraduate. Contemporary Physics, 61(2):152–153, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Millen:2020:QEM

[4087] James Millen and Benjamin A. Stickler. Quantum experiments with microscale particles. Contemporary Physics, 61(3):155–168, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vincent:2020:IDD

[4088] U. E. Vincent and O. Kolebaje. Introduction to the dynamics of driven nonlinear systems. Contemporary Physics, 61(3):169–192, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Nurgalieva:2020:TQT

[4089] Nuriya Nurgalieva and Renato Renner. Testing quantum theory with thought experiments. Contemporary Physics, 61(3):193–216, 2020. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2020:BRR

[4090] B. Ishak. Book review: Red Dwarfs: their geological, chemical, and biological potential for life (2nd ed.). Contemporary Physics, 61(3):217, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2020:BRSa

[4091] B. Ishak. Book review: Star maps: history, artistry, and cartography, 3rd edition, by N. Kanas, Switzerland, Springer, 2019, 563 pp, £39.99 (hardback), ISBN 978-3-030-13612-3. Contemporary Physics, 61(3):217, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2020:BRP

[4092] R. S. Shorter. Book review: Principles of optics, 7th edition — 60th anniversary edition, by M. Born and E. Wolf, Cambridge, Cambridge University Press, 2019, 992 pp., £54.99 (hardback), ISBN: 978-1-108-47743-7. Contemporary Physics, 61(3):217–218, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2020:BRQ

[4093] Nikolaos E. Myridis. Book review: Quantum concepts in the social, ecological and biological sciences, by F. Bagarello, Cambridge, Cambridge University Press, 2019, 300 pp., £59.99 (hardback), ISBN: 978-1-108-49212-6. Contemporary Physics, 61(3):218–219, 2020. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2020:BRT

[4094] Toni Shiroka. Book review: Thermodynamics of magnetizing materials and superconductors, by Vladimir Kozhevnikov, Boca Raton, CRC Press, 2019, 133 pp., £50.00 (hardback), ISBN: 978-1-138-49993-5. Contemporary Physics, 61(3):219–220, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2020:BRA

[4095] Toni Shiroka. Book review: Application-driven quantum and statistical physics: a short course for future scientists and engineers, volume 2: equilibrium, by J. M. Gillet, London, World Scientific, 2019, 314 pp., $45.00 (paperback), ISBN: 978-178-6347-03-9. Contemporary Physics, 61(3):220–221, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2020:BRI

[4096] Toni Shiroka. Book review: Introduction to solid state physics, by A. Aharony and O. Entin-Wohlman, Singapore, World Scientific, 2018, 640 pp., $148.00 (hardback), ISBN: 978-981-3272-24-8. Contemporary Physics, 61(3):221–222, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2020:BRM

[4097] Robert S. MacKay. Book review: Modern thermodynamics for chemists and biochemists, by D. Sherwood and P. Dalby, Oxford, Oxford University Press, 2018, 912 pp., £45 (paperback), ISBN: 978-0-19-878470-8. Contemporary Physics, 61(3):222–223, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2020:BRT

[4098] Eric Howard. Book review: Topological phase transitions and new developments, by Lars Brink, Mike Gunn, Jorge V Jose, John Michael Kosterlitz and Kok Khoo Phua, Singapore, World Scientific, 2018, 264 pp., $78.00 (hardback), ISBN: 978-981-327-133-3. Contemporary Physics, 61(3):224, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2020:BRC

[4099] Nikolaos E. Myridis. Book review: Complexity science: an introduction, edited by M. A. Peletier, R. A. van Santen, and E. Steur, Singapore, World Scientific, 2019, 428 pp., £120.00 (hardback), ISBN 978-981-323-959-3. Contemporary Physics, 61(3):224–226, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2020:BRE

[4100] Francisco Caruso. Book review: Ether and modernity, edited by J. Navarro, Oxford, UK, Oxford University Press, 2018, 272 pp., £65.00 (hardback), ISBN 978-0-19-879725-8. Contemporary Physics, 61(3):226– 227, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

McIver:2020:DGW

[4101] Jess McIver and D. H. Shoemaker. Discovering gravitational waves with Advanced LIGO. Contemporary Physics, 61(4):229–255, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

White:2020:HCD

[4102] D. R. White. Hot and cold: defining and measuring temperature. Contemporary Physics, 61(4):256–276, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Campugan:2020:OMS

[4103] Carl A. Campugan, Kylie R. Dunning, and Kishan Dholakia. Optical manipulation: a step change for biomedical science. Contemporary Physics, 61(4):277–294, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

delaTorre:2020:CBE

[4104] Laura de la Torre and Juan A. Añel. Connections between the earth’s atmosphere and its magnetosphere. Contemporary Physics, 61(4):295– 297, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2020:BRSb

[4105] B. Ishak. Book review: A student’s guide to the Schrödinger equation, by Daniel A. Fleisch, Cambridge University Press, 2020, $24.99 (Paperback), ISBN 978-1-108-81978-7. Scope: guide, textbook. Level: advanced undergraduate. Contemporary Physics, 61(4):298, 2020. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ali:2020:BRP

[4106] Muhammad Mahmood Ali and Gabriel Leen. Book review: From photons to atoms: the electromagnetic nature of matter, by D. Funaro, Singapore, World Scientific Publishing Co. Pte. Ltd., 2019, 296 pp., £95.00 (hardback), ISBN: 978-981-12-0423-4. Scope: reference, monograph, general interest. Level: postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 61(4): 298–299, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2020:BRP

[4107] Robert S. MacKay. Book review: The physics of gamma-ray bursts, by Bing Zhang, Cambridge, Cambridge University Press, 2019, 579 pp., $79.99 (hardback), ISBN 978-1-107-02761-9. Scope: monograph, Level: researchers. Contemporary Physics, 61(4):299–300, 2020. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2020:BRM

[4108] Eric Howard. Book review: Modified gravity: progresses and outlook of theories, numerical techniques and observational tests, edited by Baojiu Li and Kazuya Koyama, Singapore, World Scientific, 2019, 313 pp., $98 (hardback), ISBN: 978-981-327-399-3. Scope: review. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 61(4): 300–301, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tennyson:2020:BRC

[4109] Jonathan Tennyson. Book review: The crowd & the cosmos: adventures in the Zooniverse, by C. Lintott, Oxford, Oxford University Press, 2019, 265 pp., £20 (hardback), ISBN: 978-0-19-884222-4. Scope: general interest. Level: general readership. Contemporary Physics, 61(4): 301–302, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2020:BRIb

[4110] Stephen H. Ashworth. Book review: The iodine molecule: insights into intra- and intermolecular perturbation in diatomic molecules, by S. Lukashov, A. Petrov, and A. Pravilov, Cham, Springer, 2018, 199 pp., £74.99 (hardback), ISBN: 978-3-319-70071-7. Scope: review. Level: researcher. Specialist. Scientist. Contemporary Physics, 61(4):302–303, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2020:BRM

[4111] Chih-Yueh Wang. Book review: Measurements, mechanisms, and models of heat transport, by A. M. Hofmeister, Elsevier, 2018, 440 pp., £65.41 (paperback), ISBN 978-0-12-809981-0. Scope: monograph. Level: postgraduate, researcher, scientist. Contemporary Physics, 61(4):303–304, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2020:BRAa

[4112] David J. Miller. Book review: Astronomical spectroscopy: an introduction to the atomic and molecular physics of astronomical spectroscopy, by J. Tennyson, Singapore, World Scientific, 2019, 264 pp., £33.00 (paperback), ISBN 978-1-78634-707-7. Scope: textbook. Level: undergraduate. Contemporary Physics, 61(4):304, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2020:BRM

[4113] Francisco Caruso. Book review: The Majorana Case: letters, documents, testimonies, by Erasmo Recami, Singapore, World Scientific, 2020, 400 pp., £29.90 (paperback), ISBN 978-981-12-0769-3. Scope: book, general interest. Level: undergraduate, advanced undergraduate, postgraduate, early career research, historian of science, teacher, and specialist. Contemporary Physics, 61(4):304–305, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2020:BRAb

[4114] David J. Miller. Book review: About stars: their formation, evolution, compositions, locations and companions, by Michael M. Woolfson, Singapore, World Scientific, 2019, $78/£75 (hardback), $38/£35 (paperback), ISBN 978-1-78634-712-1 (hardback), 978-1-78634-725-1 (paperback). Scope: textbook, with problems and specimen answers. Level: Physics or Astronomy students or professionals. Contemporary Physics, 61(4):305–307, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2020:BRSa

[4115] David J. Miller. Book review: The sun today, by Claudio Vita-Finzi, Cham, Switzerland, Springer, 2018, 111 pp., £9.99 (paperback), ISBN 978-3-030-04078-9. Scope: general interest. Level: general readership. Contemporary Physics, 61(4):307, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2020:BRMb

[4116] Manuel Vogel. Book review: Molecules in electromagnetic fields: from ultracold physics to controlled chemistry, by R. V. Krems, New York, Wiley, 2018, 384 pp., £110.00 (hardback), ISBN: 978-1-118-17361-9. Scope: monograph. Level: postgraduate, researcher, scientist. Contemporary Physics, 61(4):307–308, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2020:BRSb

[4117] David J. Miller. Book review: Stephen Hawking; a memoir of friendship and physics, by Leonard Mlodinow, New York, Allen Lane, 2020, 240 pp., £20.00 (Hardback), ISBN: 978-0-241-41113-1. Scope: biography, with added science. Level: general readers, undergraduates, high school students. Contemporary Physics, 61(4):308–309, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Cassidy:2020:BRI

[4118] David Cassidy. Book review: An introduction to cold and ultracold chemistry: atoms, molecules, ions and Rydbergs, by Jesús Pérez Ríos, Cham, Springer, 2020, 267 pp., £59.99 (E-Book), ISBN: 978-3-030-55936-6. Scope: text book. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 61(4):309–310, 2020. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2020:BRS

[4119] Eric Howard. Book review: Synthetic spin-orbit coupling in cold atoms, edited by Wei Zhang, Wei Li and Carlos A. R. Sá Melo, Singapore, World Scientific, 2018, 308 pp., £84.10 (hardback), ISBN: 978-981-327-252-1. Scope: review. Level: early career researcher, researcher. Contemporary Physics, 61(4):310, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ali:2020:BRC

[4120] Muhammad Mahmood Ali. Book review: Computational acoustics: theory and implementation, by D. R. Bergman, Hoboken, NJ, Wiley, 2018, 296 pp., £99.95 (hardcover), ISBN: 978-1-119-27728-6. Scope: reference, guide, general interest, textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 61(4):310–311, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pospisil:2020:BRP

[4121] Christina Pospisil. Book review: Principles of space-time-matter: cosmology, particles and waves in five dimensions, by P. S. Wesson and J. M. Overduin, Singapore, World Scientific, 2019, 262 pp., £85.00 (Hardback), ISBN: 978-981-323-577-9. Scope: review. Level: early career researcher. Contemporary Physics, 61(4):311–312, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2020:BRE

[4122] Matt Probert. Book review: Electronic structure: basic theory and practical methods, 2nd edition, by Richard M. Martin, Cambridge, Cambridge University Press, 2020, £64.99, ISBN 978-1-108-42990-0. Level: postgraduate, early career researcher, researcher. Scope: reference text. Contemporary Physics, 61(4):312, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2020:BRO

[4123] Peter J. Bussey. Book review: Origin and evolution of the universe. From Big Bang to exobiology, 2nd edition, edited by M. A. Malkan and B. Zuckerman, Singapore, World Scientific, 2020, xiv + 232 pp., £75 (hardback), £30 (softback), ISBN 978-981-12-0645-0. Level: undergraduate, teacher. Scope: guide, edited book. Contemporary Physics, 61(4): 312–313, 2020. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lukishova:2021:IRP

[4124] Svetlana G. Lukishova, Anatoly V. Masalov, and Victor N. Zadkov. Icons of Russian physics: From the Lebedev Scientific School in Physics to the Lebedev Physical Institute. Contemporary Physics, 62(1):1–13, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rose:2021:ADA

[4125] S. J. Rose and P. W. Hatfield. Astronomy domine: advancing science with a burning plasma. Contemporary Physics, 62(1):14–23, 2021. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

English:2021:DTF

[4126] Elizabeth Laier English, Ali Ashkhasi, and Peter Whibberley. Distributing time by fibre. Contemporary Physics, 62(1):24–38, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:MPP

[4127] Peter J. Bussey. Multiverse — physics or philosophy? Contemporary Physics, 62(1):39–41, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2021:PMP

[4128] Chih-Yueh Wang. Principia and magnificent Principia. Contemporary Physics, 62(1):42–45, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2021:BRM

[4129] Miguel A. F. Sanjuán. Book review: A modern approach to teaching classical mechanics: A review of Modern classical mechanics by T. M. Helliwell and V. V. Sahakian, Cambridge, Cambridge University Press, 2020, 704 pp., £56.00, (hardback), ISBN 978-1-108-83497-1. Contemporary Physics, 62(1):46–49, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2021:BRPa

[4130] A. H. Harker. Book review: The physics of graphene, 2nd edition, by Mikhail I. Katsnelson, Cambridge, UK, Cambridge University Press, 2020, 436 pp., $84.99 (hardback), ISBN: 978-1-108-47164-0. Contemporary Physics, 62(1):50, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pospisil:2021:BRS

[4131] Christina Pospisil. Book review: String theory in a nutshell, 2nd edition, edited by Elias Kiritsis, Princeton University Press, 2019, 888 pp., $95 (hardback), ISBN 978-0-691-15579-1. Contemporary Physics, 62(1): 50–51, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2021:BRP

[4132] Miriam R. Aczel. Book review: Practical alchemy: a memoir, Walter A. Harrison, Hackensack, NJ, World Scientific, 2018, 176 pp., $28.00 (paperback), ISBN: 978-981-327-831-8. Contemporary Physics, 62(1): 51–52, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aczel:2021:BRC

[4133] Miriam R. Aczel. Book review: The cosmos: astronomy in the new millennium, 5th edition. Contemporary Physics, 62(1):52–53, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:BRM

[4134] Peter J. Bussey. Book review: Modern quantum mechanics. Contemporary Physics, 62(1):53–54, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bengtsson:2021:BRD

[4135] Ingemar Bengtsson. Book review: The discovery of cosmic voids by Laird A. Thompson, Cambridge, Cambridge University Press, 2020, 290 pp., £39.99 (hardback), ISBN 978-1-108-49113-6. Contemporary Physics, 62 (1):54, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2021:BRNa

[4136] Manuel Vogel. Book review: The NASA Kepler mission, edited by Steve B. Howell, Bristol, IOP Publishing, 2020, 272 pp., $190 (hardback), ISBN 978-0-7503-2294-2. Contemporary Physics, 62(1):55, 2021. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2021:BRQ

[4137] Stephen H. Ashworth. Book review: Quantum space, y Jim Baggott, Oxford, Oxford University Press, 2018, £20.00 (hardback), ISBN 978-0-19-880911-1. Contemporary Physics, 62(1):55–56, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gaunt:2021:BRC

[4138] Jonathan Richard Gaunt. Book review: Coherent states in gauge theories and applications in collider physics, by M. Greco, World Scientific, 2019, 172 pp., £75 (Hardback), ISBN: 978-981-12-1389-2. Contemporary Physics, 62(1):56–57, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2021:BRI

[4139] Mario Bertolotti. Book review: An introduction to quantum optics: an open system approach, by Perry Rice, Bristol, UK, IoP Publishing, 2020, xii+225 pp. with figures, $95.00 (e-book), ISBN 978-0-7503-1713-9. Contemporary Physics, 62(1):57–58, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2021:BRE

[4140] Mario Bertolotti. Book review: Electromagnetic waves and lasers, 2nd edition, by Wayne D. Kimura, IOP Science, Bristol, UK, 2020, xv + 142 pp. with figures, $159.00 (paperback/hardback), ISBN: 978-0-7503-3521-8. Contemporary Physics, 62(1):58, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2021:BRS

[4141] Eric Howard. Book review: Semiclassical and stochastic gravity: quantum field effects on curved spacetime, Bei-Lok B. Hu and Enric Verdaguer, Cambridge, Cambridge University Press, 2020, 614 pp., £155.00 (hardback), ISBN: 978-0-521-19357-3. Contemporary Physics, 62(1):58– 59, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2021:BRB

[4142] R. S. Shorter. Book review: The Bohr atom: a guide, by Bruce Cameron Reed, Bristol, IOP Science, 2020, 104 pp., $159.00 (E-book), Online ISBN 978-0-7503-3611-6; Print ISBN: 978-0-7503-3609-3. Contemporary Physics, 62(1):59, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

DeAngelis:2021:BRC

[4143] Alessandro De Angelis. Book review: Cosmic messengers: the limits of astronomy in an unruly universe, by Martin Harwit, Cambridge, Cambridge University Press, 2021, 396 pp., $32.00 (E-Book), ISBN: 978-1-108-90616-6. Contemporary Physics, 62(1):59–61, 2021. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2021:BRP

[4144] Jesus Rogel-Salazar. Book review: Physics of digital photography, 2nd edition, by Andy Rowlands, Bristol, UK, IOP Science, 2020, 165 pp., $159.00 (E-Book), ISBN: 978-0-7503-2558-5. Contemporary Physics, 62 (1):61, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2021:BRD

[4145] Giovanni Volpe. Book review: Don’t be afraid of physics: quantum mechanics, relativity and cosmology for everyone, by Ross Barrett and Pier Paolo Delsanto, Cham, Switzerland, Springer Nature, 2021, 293 pp., £17.99 (e-book), ISBN 978-3-030-63409-4. Contemporary Physics, 62 (1):61–62, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2021:BRI

[4146] B. Ishak. Book review: Introduction to galaxy formation and evolution: from primordial gas to present-day galaxies, by Andrea Cimatti, Filippo Fraternali, and Carlo Nipoti, Cambridge, Cambridge University Press, 2019, 582 pages, $69.99 (hardback), ISBN: 978-1-107-13476-8. Contemporary Physics, 62(1):62, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2021:BRM

[4147] R. S. Shorter. Book review: Modern physics: a critical approach, edited by Canio Noce, Bristol, Institute of Physics Publishing, 2020, 450 pp., £64 (hardback), ISBN: 978-0-7503-2676-6. Contemporary Physics, 62 (1):62–63, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2021:BRPa

[4148] Mogadalai P. Gururajan. Book review: Probability and statistics in the physical sciences, 3rd edition, by Byron P. Roe, New York, Springer Nature Switzerland, 2020, 285 pp., EUR 46 (e-book), ISBN: 978-3-030-53693-0. Contemporary Physics, 62(1):63–64, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Andrews:2021:BRM

[4149] David Andrews. Book review: Mind and the cosmic order: how the mind creates the features & structure of all things, and why this insight transforms physics, by Charles Pinter, Springer, ISBN: 978-3-030-50083-2 (e-book). Contemporary Physics, 62(1):64–65, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2021:BRP

[4150] Nigel John Mason. Book review: Planetary diversity: rocky planet processes and their observational signatures, y Elizabeth J. Tasker, Yuka Fujii, Matthieu Laneuville, Cayman Unterborn, Steven J. Desch and Hilairy E. Hartnett, IOP Publishing, 2020, 266 pp., $159.00 (e-book), ISBN: 978-0-7503-2140-2. Contemporary Physics, 62(1):65–66, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2021:BRL

[4151] N. E. Myridis. Book review: The large-scale structure of the Universe, by P. J. E. Peebles, winner of the Nobel prize in Physics, Princeton University Press, 2020, 448 pp., £50.00 (paperback), ISBN: 978-0-691-20983-8. Contemporary Physics, 62(1):66–67, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Morozov:2021:SPD

[4152] Dmitry V. Morozov, Alessandro Casaburi, and Robert H. Hadfield. Superconducting photon detectors. Contemporary Physics, 62(2):69–91, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zuk:2021:BIC

[4153] Paweł Jan Zuk. A brief introduction to contemporary electrokinetics. Contemporary Physics, 62(2):92–104, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:BRD

[4154] Peter J. Bussey. Book review: Do wave functions jump? Perspectives of the work of GianCarlo Ghirardi edited by Valia Allori, Angelo Bassi, Detlef Dürr, and Nino Zanghi, Cham, CH, Springer Nature Switzerland, 2021, xv + 436 pp., £87.50 (hardcover), ISBN 978-3-0304-6776-0/ ISBN 978-3-0304-6777-7 (e-book). Scope: article collection. Level: early career researcher, specialist, scientist. Contemporary Physics, 62(2):105– 106, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2021:BRN

[4155] Salvatore Esposito. Book review: The neutrino story: one tiny particle’s grand role in the cosmos by Rabindra N. Mohapatra, Cham, Switzerland, Springer, 2021, xvi+222 pp., £17.99 (E-Book), ISBN 978-3-0305-1846-2. Scope: monograph, general interest. Level: non-specialists, undergraduate, specialist, scientist. Contemporary Physics, 62(2):106–107, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2021:BRE

[4156] A. H. Harker. Book review: The equations of materials by Brian Cantor, Oxford, Oxford University Press, 2020, 19 + 307pp., £59.99 (hardback), £29.99 (paperback), ISBN 978-0-19-885187-5 (hbk)/978-0-19-885188-2 (pbk). Scope: general interest. Level: pre-university, undergraduate. Contemporary Physics, 62(2):107, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2021:BRC

[4157] Robert S. MacKay. Book review: The cosmic microwave background, 2nd edition, by Ruth Durrer, Cambridge, UK, Cambridge University Press, 2021, 510 pp., £47.61 (hardback), ISBN 978-1-107-13522-2. Scope: textbook. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 62(2):107–108, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2021:BRPa

[4158] David J. Miller. Book review: Particle detectors: fundamentals and applications by Hermann Kolanoski and Norbert Wermes, New York, Oxford University Press, 2020, 944 pp., $100 (Hardback), ISBN 978-0-19-885836-2. Scope: monograph. Level: graduate. Contemporary Physics, 62(2):108–109, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2021:BRA

[4159] David J. Miller. Book review: Advances in nuclear physics: structure and reactions edited by R.K. Puri, J. Aichelin, S. Gautam and R. Kumar, Springer Nature, 2020, 273 pp., £129.99 (hardback), ISBN 978-981-1590-61-0. Scope: conference proceedings. Level: researcher. Contemporary Physics, 62(2):109–110, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2021:BRPb

[4160] David J. Miller. Book review: The physics of the dark photon: a primer by Marco Fabbrichesi, Emidio Gabrielli and Gaia Lanfranchi, Cham, Springer Nature, 2021, 78 pp., £44.99 (paperback), ISBN: 978-3-030-62518-4. Scope: review. Level: researcher. Contemporary Physics, 62 (2):110, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Miller:2021:BRC

[4161] David J. Miller. Book review: Classical mechanics: a professor–student collaboration edited by Mario Campanelli, Bristol, IOP Science, 2020, 376 pp., £99.00 (hardback), ISBN: 978-0-7503-2688-9. Scope: textbook. Level: undergraduate. Contemporary Physics, 62(2):110–111, 2021. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2021:BRS

[4162] A. H. Harker. Book review: Semiconductor physics: principles, theory and nanoscale by S. Tiwari, Oxford, Oxford University Press, 2020, 19 + 803 pp., £65.00 (hardback), ISBN 978-0-19-875986-7. Scope: textbook. Level: undergraduate, early postgraduate. Contemporary Physics, 62(2): 111–112, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Halttunen:2021:BRA

[4163] Krista Halttunen. Book review: The advent of relativity by Voldemar Smilga, Singapore, World Scientific, 2021, 352 pp., £60.00 (hardback), ISBN: 978-981-12-3114-8. Scope: monograph. Level: pre-university. Contemporary Physics, 62(2):112, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Osborn:2021:BRQ

[4164] H. Osborn. Book review: From quarks to pions: chiral symmetry and confinement by M. Creutz, World Scientific, 2018, 216 pp., £75 (hardback), ISBN: 978-981-322-923-5. Scope: monograph. Level: postgraduate. Contemporary Physics, 62(2):113, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lakshminarayanan:2021:BRH

[4165] Vasudevan Lakshminarayanan. Book review: The handbook of medical image perception and techniques, 2nd edition, by E. Samei and E.

A. Krupinski, Cambridge, UK, Cambridge University Press, 2018, 522 pp., £120.00 (hardback), ISBN: 978-1-107-19488-5. Scope: handbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist. Contemporary Physics, 62(2):113–114, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scali:2021:BRA

[4166] Stefano Scali. Book review: Advanced quantum condensed matter physics: one-body, many-body, and topological perspectives by Michael El-Batanouny, Cambridge, Cambridge University Press, 2020, £69.99 (hardback), ISBN: 978-1-108-48084-0. Scope: reference, textbook. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 62(2):114–115, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scali:2021:BRG

[4167] Stefano Scali. Book review: Group theory for physicists, 2nd edition, by Zhong-Qi Ma, Hackensack, NJ, World Scientific, 2019, 656 pp., £115 (paperback), ISBN: 978-981-327-796-0. Scope: textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 62(2):115–116, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2021:BRR

[4168] Chih-Yueh Wang. Book review: Relativistic fluid dynamics in and out of equilibrium: and applications to relativistic nuclear collisions by Paul Romatschke and Ulrike Romatschke, Cambridge, Cambridge University Press, 2019, 204 pp., US$140.00 (hardcover), ISBN-13: 978-1-108-48368-1. Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 62(2):116, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2021:BRI

[4169] Andrew Resnick. Book review: An introduction to fluorescence correlation spectroscopy by Thorsten Wohland, Radek Machán, and Sudipta Maiti, Bristol, IOP Publishing Ltd, 2020, 300 pages, $190.00, ISBN: 978-0-7503-2078-8. Scope: textbook, handbook. Level: advanced undergraduate student, graduate student, researcher. Contemporary Physics, 62(2):117, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2021:BRS

[4170] Andrew Resnick. Book review: Stigmatic optics by Rafael G. González-Acuña and Héctor A. Chaparro-Romo, Bristol, IOP Publishing Ltd, 2020, 308 pages, $190.00, ISBN-13: 978-0-7503-3461-7, ISBN-10: 0-7503-3461-4. Scope: textbook, handbook. Level: advanced undergraduate student, graduate student, researcher. Contemporary Physics, 62 (2):117–118, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shorter:2021:BRP

[4171] R. S. Shorter. Book review: The physics of the Manhattan Project by Bruce Cameron Reed, New York, Springer, Nature Switzerland AG, 2021, 256 pp., £43.99 (e-book), ISBN: 978-3-030-61372-3. Scope: general interest, textbook. Level: general readership, undergraduate, teacher. Contemporary Physics, 62(2):118, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2021:BRE

[4172] Manuel Vogel. Book review: Electrons in motion: attosecond physics explores fastest dynamics by F. Krausz, Singapore, World Scientific, 2019, 536 pp., $175 (hardback), ISBN: 978-981-12-0186-8. Scope: article collection. Level: researcher, specialist. Contemporary Physics, 62(2):119, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gershon:2021:BRP

[4173] Tim Gershon. Book review: Particle–antiparticle asymmetry in the meson system by Hiroaki Aihara, IOP Science, 2020, $190.00 (hardback), ISBN: 978-0-7503-3649-9, $159.00 (ebook), ISBN: 978-0-7503-3651-2. Scope: reference material, appropriate for undergraduate level. Contemporary Physics, 62(2):119–120, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shotter:2021:BRE

[4174] A. C. Shotter. Book review: Essentials of nucleosynthesis and theoretical nuclear astrophysics by Thomas Rauscher, Bristol, IOP Science, 2020, $190.00 (Hardback), 397 pp., ISBN 978-0-7503-1150-2. Level: advanced undergraduate, postgraduate, early career researcher, researcher. Scope: reference. Contemporary Physics, 62(2):120–121, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Volpe:2021:BRS

[4175] Giovanni Volpe. Book review: Statistical mechanics: entropy, order parameters, and complexity 2nd edition, by James P. Sethna, Oxford, Oxford University Press, 2021, 496pp., £58.99 (paperback), ISBN 978-0-19-886524-7. Scope: textbook. Level: undergraduate, postgraduate, advanced undergraduate. Contemporary Physics, 62(2):121–122, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2021:BRR

[4176] Dr. Manuel Vogel. Book review: Relativity and cosmology: volume 5 of modern classical physics by Kip S. Thorne and Roger D. Blandford, Princeton, Princeton University Press, 2021, 416 pp., £48 (softcover), ISBN: 978-0-691-20739-1. Scope: textbook. Level: advanced undergraduates. Contemporary Physics, 62(2):122, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:BRI

[4177] Peter J. Bussey. Book review: Introduction to the standard model and beyond: quantum field theory, symmetries and phenomenology by Stuart Raby, Cambridge, Cambridge University Press, 2021, xvi+619 pp., £60 (hardback), ISBN 978-1-108-49,419-9. Scope: textbook. Level: postgraduate. Contemporary Physics, 62(2):123, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2021:BRC

[4178] Mogadalai P. Gururajan. Book review: Continuum mechanics of solids by Lallit Anand and Sanjay Govindjee, Oxford, UK, Oxford University Press, 2020, 736 pp., £65.00 (hardback), ISBN: 978-0-19-886472-1. Scope: textbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, engineers. Contemporary Physics, 62(2): 123–124, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barnes:2021:PCE

[4179] Samuel J. K. Barnes and Aneta Stefanovska. Physics of cellular energy metabolism. Contemporary Physics, 62(3):125–143, 2021. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kumar:2021:SCI

[4180] Sunil Kumar, Manoj Kumar, and Amarjeet Singh. Synthesis and characterization of iron oxide nanoparticles (Fe_{2}O_{3}, Fe_{3}O_{4}): a brief review.

Contemporary Physics, 62(3):144–164, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Palti:2021:SST

[4181] Eran Palti. The swampland and string theory. Contemporary Physics, 62(3):165–179, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:BRG

[4182] Peter J. Bussey. Book review: General relativity: the essentials by Carlo Rovelli, Cambridge, Cambridge University Press, 2021, x+170 pp., £14.99 (paperback), ISBN 978-1-009-01369-7. Scope: textbook. Level: postgraduate. Contemporary Physics, 62(3):180, 2021. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2021:BRO

[4183] Peter J. Dobson. Book review: Optical materials, 2nd edition by Kelly Potter and Joseph Simmons, USA, Elsevier, 2020, 530 pp., £125-£155 (paperback) (E-Book), ISBN: 978-0-12-818642-8. Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 62(3): 180–181, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:BRSa

[4184] Peter J. Bussey. Book review: Statistical Physics, vol. 1 of Modern Classical Physics by Kip S. Thorne and Roger D. Blandford, Princeton, Princeton University Press, 2021, xxii+384 pp., £40 (paperback), ISBN 978-0-691-20612-7. Level: advanced undergraduate, postgraduate. scope: textbook. Contemporary Physics, 62(3):181–182, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wang:2021:BRC

[4185] Chih-Yueh Wang. Book review: Computational statistical physics by Lucas Böttcher and Hans J. Herrmann, Cambridge, UK, Cambridge University Press, 2021, 257 pp., US$59.99 (E-book), ISBN 978-1-108-84142-9. Scope: textbook. Level: advanced undergraduate, postgraduate, earlycareer researcher, researcher, scientist, engineers. Contemporary Physics, 62(3):182–183, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mason:2021:BRS

[4186] Nigel John Mason. Book review: Space science and public engagement: 21st century perspectives and opportunities by Amy Kaminski, Elsevier, 2021, 298 pp., $69.95 (E-Book), ISBN: 978-0-12-817390-9, https://www.elsevier.com/books/space-science-and-public-engagement/kaminski/978-0-12-817390-9. Scope: review. Level: general readership, non-specialists, pre-university, undergraduate, postgraduate, advanced undergraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 62(3):183–184, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2021:BRE

[4187] M. Jamil. Book review: Exploring classical mechanics: a collection of 350+ solved problems for students, lecturers, and researchers, 2nd revised and enlarged English edition by G. L. Kotkin and V. G. Serbo, Oxford, UK, Oxford University Press, 2020, 400 pp., £29.99 (E-Book), ISBN: 978-0-19-885379-4. Scope: monograph, handbook, guide. Level: undergraduate, advanced undergraduate, early career researcher, researcher. Contemporary Physics, 62(3):184–185, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2021:BRT

[4188] Miguel A. F. Sanjuán. Book review: Thermodynamics and statistical mechanics. An introduction for physicists and engineers by Samya Zain, Bristol, IOP Publishing, 2021, 372 pp., £99 (hardback), ISBN: 978-0-7503-3083-1, 978-0-7503-3081-7. Scope: textbook. Level: undergraduate, advanced undergraduate, teacher, engineers. Contemporary Physics, 62 (3):185, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2021:BRT

[4189] Andrew Resnick. Book review: Thermodynamics: principles characterising physical and chemical processes, 5th edition by Jurgen Honig, London, Elsevier, 2020, $190 (E-Book), ISBN: 978-0-12-821940-9. Scope: textbook. Level: advanced undergraduate, postgraduate, early career researcher, scientist. Contemporary Physics, 62(3):185–186, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2021:BRE

[4190] Mogadalai P. Gururajan. Book review: Elasticity & fluid dynamics, 1st edition by Kip S. Thorne and Roger D. Blandford, Series: Volume 3 of Modern Classical Physics, Princeton, USA, Princeton University Press, 2021, 450 pp. (565–1014), 50 USD (paperback), ISBN 978-0-691-20734-6. Scope: reference, textbook. Level: postgraduate, advanced undergraduate, early career researcher, researcher, teacher, scientist, engineers.

Contemporary Physics, 62(3):186–187, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:BRSb

[4191] Peter J. Bussey. Book review: Seven fundamental concepts in spacetime physics by Vesselin Petkov, Cham, CH, Springer International Publishing, 2021, xix + 81 pp., £50 (paperback), ISBN: 978-3-030-75638-3. Scope: monograph. Level: scientist. Contemporary Physics, 62(3): 187–188, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Xie:2021:MTP

[4192] Songbo Xie and J. H. Eberly. Managing the three-party entanglement challenge. Contemporary Physics, 62(4):189–198, 2021. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jiang:2021:PMB

[4193] Xiangqian Jiang and Andrew Henning. Precision metrology: from bulk optics towards metasurface optics. Contemporary Physics, 62(4):199– 216, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rider:2021:SNL

[4194] M. S. Rider and W. L. Barnes. Something from nothing: linking molecules with virtual light. Contemporary Physics, 62(4):217–232, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2021:RP

[4195] Peter J. Bussey. Resolving a paradox. Contemporary Physics, 62(4): 233–235, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Howard:2021:BRQ

[4196] Eric Howard. Book review: The quantum universe: essays on quantum mechanics, quantum cosmology, and physics in general by James B. Hartle, Singapore, World Scientific, 2021, 624 pp., £110.00 (E-Book), ISBN: 978-981-12-1634-3. Scope: edited book. Level: early career researcher. Contemporary Physics, 62(4):236, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2021:BRM

[4197] Matt Probert. Book review: Machine learning with neural networks: an introduction for scientists and engineers by Bernhard Mehlig, Cambridge, UK, Cambridge University Press, 2021, 249 pp., £39.99 (hardback), ISBN: 978-1-108-49493-9. Scope: monograph, textbook. Level: postgraduate, researcher. Contemporary Physics, 62(4):236–237, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gururajan:2021:BRPb

[4198] Mogadalai P. Gururajan. Book review: Physics of elasticity and crystal defects Oxford series on materials modelling, by Adrian P. Sutton, Oxford, Oxford University Press, 2020, 288 pp., £45.00 (hardback), ISBN: 978-0-19-886078-5. Scope: monograph, textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, scientist, engineers. Contemporary Physics, 62(4):237–238, 2021. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sharma:2021:BRI

[4199] Mohit Sharma. Book review: Introduction to quantum nanotechnology: a problem focused approach by Duncan G. Steel, Oxford, UK, Oxford University Press, 2021, 370 pp., £65.00 (hardback), ISBN 978-0-19-289507-3. Scope: reference, review. Level: postgraduate, early career researcher, researcher, teacher, scientist, engineers. Contemporary Physics, 62(4): 238–239, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2021:BRC

[4200] Nikolaos E. Myridis. Book review: Cosmology’s century: an inside history of our modern understanding of the universe Physics & Astronomy, by P. J. E. Peebles, winner of the Nobel prize in Physics, Princeton, NJ, Princeton University Press, 2020, 440 pp., $28.00 (hardback), ISBN (ebook) 978-0-691-20166-5. Scope: monograph. Level: non-specialists, undergraduate, advanced undergraduate, postgraduate, researcher, teacher, specialist, scientist. Contemporary Physics, 62(4):239–240, 2021. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:2021:BRH

[4201] Jack Reid. Book review: A handbook of mathematical methods and problem-solving tools for introductory physics (second edition) by Joshua F. Whitney, Heather M. Whitney, IOP Science, Bristol, Institute of Physics Publishing, 2021, $120.00 (hardback), $90.00 (e-book), E-Book (https://store.ioppublishing.org/page/detail/A-Handbook-of-Mathematical-Methods-and-ProblemSolving-Tools-for-Introductory-Physics-Second-Edition/?K=9780750333993), ISBN 978-0-7503-3399-3. Scope: handbook. Level: undergraduate. Contemporary Physics, 62(4):240–241, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2021:BRU

[4202] Michael Belsley. Book review: Ultrafast lasers and optics for experimentalists by James David Pickering, Bristol, UK, IOP Science, 2021, 157 pp., ISBN 978-0-7503-3659-8. Scope: handbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 62(4):241–242, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2021:BRPb

[4203] A. H. Harker. Book review: Principles of materials characterization and metrology by Kannan M. Krishnan, Oxford, UK, Oxford University Press, 2021, 20 + 849 pp., ISBN 978-0-19-883025-2 (hbk), ISBN 978-0-19-883026-9 (pbk). Scope: textbook. Level: undergraduate, early postgraduate. Contemporary Physics, 62(4):242, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2021:BRN

[4204] A. H. Harker. Book review: Nuclear materials science, 2nd edition by Karl Whittle, IOP Publishing Ltd, 2020, 218 pp., ISBN: 978-0-7503-2376-5 (ebook), ISBN: 978-0-7503-2374-1 (print). Scope: textbook. Level: undergraduate, early postgraduate. Contemporary Physics, 62(4):242– 243, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2021:BRP

[4205] Nikolaos E. Myridis. Book review: Principles of physical Cosmology by P. J. E. Peebles, winner of the Nobel prize in Physics, Princeton, NJ, Princeton University Press, 2020, 774 pp., £62.00 (e-book), ISBN 978-0-691-20672-1. Princeton series in physics & astronomy. Scope: reference. Level: non-specialists, advanced undergraduate, postgraduate, researcher, teacher, specialist, scientist. Contemporary Physics, 62(4): 243–244, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Boyack:2021:BRE

[4206] Rufus Boyack. Book review: Electrostatics of conducting cylinders and spheres by John Lekner, Melville, NY, AIP Publishing, 2021, 240 pp., $135 (paperback), https://doi.org/10.1063/9780735423350, ISBN: 978-0-7354-2332-9. Scope: monograph. Level: scientist. Contemporary Physics, 62(4):244–245, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2021:BRNb

[4207] Manuel Vogel. Book review: Nuclear data — a primer by David G. Jenkins and John L. Wood, Bristol, UK, IOP Publishing, 2021, 198 pp., USD 190 (hardcover), ISBN 978-0-7503-2672-8. Scope: textbook. Level: advanced undergraduates and above. Contemporary Physics, 62(4):245– 246, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2021:BRR

[4208] Miguel A. F. Sanjuán. Book review: A random walk in physics. Beyond black holes and time-travels by Massimo Cencini, Andrea Puglisi, Davide Vergni, and Angelo Vulpiani, Springer Nature Switzerland AG, 2021, 36.39 EUR (hardback), 224 pp. ISBN 978-3-030-72530-3. Scope: general interest. Level: general readership, undergraduate, teacher. Contemporary Physics, 62(4):246–247, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Deppisch:2021:BRF

[4209] Frank F. Deppisch. Book review: Foundations of quantum cosmology by Martin Bojowald, IOP Science, 2020, ISBN: 978-0-7503-2460-1 (E-Book). Scope: monograph. Level: physics graduate. Contemporary Physics, 62 (4):247, 2021. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2021:BRT

[4210] Stephen J. Blundell. Book review: Theory of simple glasses: exact solutions in infinite dimensions by Giorgio Parisi, Pierfrancesco Urbani and Francesco Zamponi, Cambridge University Press, 2020, 349 pages, £62.99 (hardcover), ISBN: 978-1-108-12049-4. Level: researcher. Scope: textbook. Contemporary Physics, 62(4):247–248, 2021. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2021:BRI

[4211] Stephen H. Ashworth. Book review: The ideas of particle physics by James Dodd and Ben Gripaios, Cambridge University Press, 2020, £34.99 (paperback, https://www.cambridge.org/gb/academic/subjects/physics/particlephysics-and-nuclear-physics/ideas-particle-physics-4th-edition?format=PB), ISBN 978-1-108-72740-2. Scope: review. Level: general readership, nonspecialists, scientists. Contemporary Physics, 62(4):248–249, 2021. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Keller:2022:CQS

[4212] Matthias Keller. Cavity-QED with single trapped ions. Contemporary Physics, 63(1):1–14, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chapman:2022:FLS

[4213] E. Chapman. First light: switching on stars at the dawn of time. Contemporary Physics, 63(1):15–33, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Milburn:2022:PLM

[4214] G. J. Milburn and Sahar Basiri-Esfahani. The physics of learning machines. Contemporary Physics, 63(1):34–60, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2022:BRSa

[4215] Manuel Vogel. Book review: Synchrotron radiation: sources and applications to the structural and electronic properties of materials by D. Phil Woodruff, Cambridge, Cambridge University Press, 2021, 450 pp., £74.99 (hardback), ISBN: 978-1-316-99574-7. Scope: monograph. Level: advanced undergraduate, postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 63(1):61, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sivakumar:2022:BRQ

[4216] Srinivasan Sivakumar. Book review: Quantum information processing, quantum computing, and quantum error correction: an engineering approach, 2nd edition by Ivan Djordjevic, Elsevier, 2021, 838 pp., $125.00 (paperback), ISBN 978-0-12-821987-4. Scope: textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 63(1):61–62, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beauty:2022:BRE

[4217] Aras Beauty, Laili Fitria Zulfa, and Iwan. Book review: Entropy and free energy in structural biology: From thermodynamics to statistical mechanics to computer simulation by Hagai Meirovitch, Boca Raton, CRC Press, 2020, 396 pp., £80.00 (hardback), ISBN 978-0-367-40692-9. Scope: textbook. Level: advanced undergraduate, postgraduate, researcher, scientist, engineer. Contemporary Physics, 63(1):62–63, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:2022:BRQ

[4218] Jack Reid. Book review: Quantum mechanics: introduction to mathematical formulation by Martin Pieper, Wiesbaden, Springer International Publishing, 2021, xiii + 33 pp., £19.99 (paperback), ISBN 978-3-658-32645-6. Scope: guide. Level: general readership. Contemporary Physics, 63(1):63, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2022:BRT

[4219] Miguel A. F. Sanjuán. Book review: Theory of the spread of epidemics and movement ecology of animals. An interdisciplinary approach using methodologies of physics and mathematics by V. M. (Nitant) Kenkre and Luca Giuggioli, Cambridge, UK, Cambridge University Press, 2021, 330 pp., £120 (hardback), ISBN: 978-1-108-84140-5. Scope: monograph. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 63(1):64, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shore:2022:BRF

[4220] K. Alan Shore. Book review: Fundamental properties of semiconductor nanowires edited by Fukata Naoki and Rurali Riccardo, Singapore, Springer International Publishing, 2021, 454 pp., £87.50 (E-Book), ISBN 978-981-15-9049-8. Scope: monograph. Level: postgraduate, early career researcher. Contemporary Physics, 63(1):64–65, 2022. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bussey:2022:BRE

[4221] Peter J. Bussey. Book review: Exofrontiers: big questions in exoplanetary science edited by N. Madhusudhan, Bristol, IOP Publishing, 2021, xxi+ 287 pp., £108 (hardback), ISBN 978-0-7503-1470-1. Scope: article collection. Level: researcher. Contemporary Physics, 63(1):65–66, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2022:BRD

[4222] A. H. Harker. Book review: Dimensional analysis: the great principle of similitude by Jeffrey H. Williams, Bristol, UK, IOP Publishing Ltd, 2021, 20 + 223 pp., £120 (hardback), ISBN 978-0-7503-3653-6; £99 (ebook), ISBN 978-0-7503-3655-0. Scope: textbook. Level: undergraduate. Contemporary Physics, 63(1):66–67, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:2022:BRA

[4223] Jack Reid. Book review: The ABCs of High-Pressure Science by Sergei M. Stishov, London, World Scientific Publishing, 2021, £40.00 (hardback), ISBN: 978-1-78634-955-2. Scope: reference. Level: general readership. Contemporary Physics, 63(1):67, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Myridis:2022:BRQ

[4224] Nikolaos E. Myridis. Book review: Quantitative atomic-resolution electron microscopy by Annick de Backer, Jarmo Fatermans, Arnold A. den Dekker and Sandra van Aert, vol. 217 of Advances in Imaging & Electron Physics, edited by Martin Hÿtch and Peter W. Hawkes, Elsevier, 2021, 294 pp., EUR139.13 (e-book), ISBN 978-0-323-85093-3. Scope: article collection. Level: researcher, scientist. Contemporary Physics, 63 (1):67–69, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2022:BRS

[4225] Stephen H. Ashworth. Book review: The science of learning physics by José P. Mestre and Jennifer L. Docktor, Singapore, World Scientific, 2021, 212 pp., £50.00 (hardback), ISBN 978-981-12-2654-0. also available as softback and ebook. Level: teacher, specialist. Scope: review. Contemporary Physics, 63(1):69, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ramires:2022:SAC

[4226] Aline Ramires. Symmetry aspects of chiral superconductors. Contemporary Physics, 63(2):71–86, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lian:2022:AMB

[4227] Xiaojuan Lian, Yuelin Shi, Shiyu Li, Bingxin Ding, Chenfei Hua, and Lei Wang. Applications of MXene-based memristors in neuromorphic intelligence applications. Contemporary Physics, 63(2):87–105, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ueda:2022:SBE

[4228] Masahito Ueda. Spinor Bose–Einstein condensates. Contemporary Physics, 63(2):106–115, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Withington:2022:QEF

[4229] S. Withington. Quantum electronics for fundamental physics. Contemporary Physics, 63(2):116–137, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lellouch:2022:UAI

[4230] Samuel Lellouch, Kai Bongs, and Michael Holynski. Using atom interferometry to measure gravity. Contemporary Physics, 63(2):138–155, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2022:BRM

[4231] Thomas J. Bridges. Book review: Mechanics of fluid deformations in rigid body rotations and plane channel F by Oleg V. Troshkin, Singapore, World Scientific, 2021, 284 pp., £85 (hardback), ISBN 978-981-1230-51-6. Scope: monograph. Level: postgraduate, early career researcher, researcher. Contemporary Physics, 63(2):156, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2022:BRV

[4232] Michael Belsley. Book review: Virtual and real labs for introductory physics II: optics, modern physics, and electromagnetism by Daniel Erenso, Bristol, UK, IOP Science, 2021, 380 pp., 90EUR (hardback), ISBN 978-0-7503-3713-7. Scope: textbook. Level: pre-university, undergraduate. Contemporary Physics, 63(2):156–157, 2022. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2022:BRSb

[4233] Manuel Vogel. Book review: String theory and the real world: the visible sector, 2nd edition, by Gordon Kane, Bristol, UK, IOP Publishing, 2021, 175 pp., $50 (hardback); $40 (ebook), ISBN: 978-0-7503-3581-2. Scope: monograph. Level: general readership. Contemporary Physics, 63(2): 157–158, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Resnick:2022:BRG

[4234] Andrew Resnick. Book review: Gravity, magnetic and electromagnetic gradiometry: strategic technologies in the 21st century, 2nd edition, by Alexey V. Veryaskin, Bristol, IOP Publishing Ltd, 2021, 175 pp., $125.00 (hardcover), ISBN-13: 978-0-7503-3801-1, ISBN-10: 0-7503-3801-6. Scope: monograph. Level: advanced undergraduate student, graduate student, researcher. Contemporary Physics, 63(2):158– 159, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2022:BRA

[4235] A. H. Harker. Book review: Applications of polymer nanofibers edited by Anthony L. Andrady and Saad A. Khan, Hoboken, NJ, Wiley, 2022, xvi + 408 pp., £133.99 (e-book), ISBN 978-1-119-26770-6. Scope: edited book. Level: early career researcher, researcher. Contemporary Physics, 63(2):159, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schofield:2022:QEF

[4236] Holly Schofield, Elena Boto, Vishal Shah, Ryan M. Hill, James Osborne, Molly Rea, Cody Doyle, Niall Holmes, Richard Bowtell, David Woolger, and Matthew J. Brookes. Quantum enabled functional neuroimaging: the why and how of magnetoencephalography using optically pumped magnetometers. Contemporary Physics, 63(3):161–179, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clements:2022:VPP

[4237] David L. Clements. Venus, phosphine and the possibility of life. Contemporary Physics, 63(3):180–199, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Seligman:2022:IO

[4238] Darryl Z. Seligman and Amaya Moro-Martín. Interstellar objects. Contemporary Physics, 63(3):200–232, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Quigg:2022:AL

[4239] Chris Quigg. Advice to a lecturer. Contemporary Physics, 63(3):233– 241, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2022:BRB

[4240] Peter J. Dobson. Book review: Bionanomaterials for environmental and agricultural applications edited by Ravindra Pratap Singh and Kshitij RB Singh, Institute of Physics Publishing Ltd, 2021, £120 (hardback), ISBN: 978-0-7503-3863-9. Scope: review, reference. Level: general readership, advanced undergraduate, postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 63(3):242, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beauty:2022:BRB

[4241] Aras Beauty and Afnan Taqiyarrahman. Book review: The brain-bladder axis in tissue growth and remodelling by Roustem N. Miftahof and Christian J. Cyron, Bristol, IOP Publishing, 2021, 175 pp., $190 (hardback), $159 (ebook), ISBN 978-0-7503-3567-6. Scope: monograph. Level: general readership, undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher, specialist, scientist, engineer. Contemporary Physics, 63(3):242–243, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Huda:2022:BRT

[4242] Erin Ficrah Huda and Suci Indah Putri. Book review: Tailored functional oxide nanomaterials: from design to multi-purpose applications edited by Chiara Maccato and Davide Barica, Weinheim, Wiley, 2022, 512 pp., $156.35 (hardback), ISBN 978-3-527-34759-9. Scope: textbook. Level: undergraduate, advanced undergraduate, postgraduate, researcher, scientist, engineer. Contemporary Physics, 63(3):243–244, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zutt:2022:BRC

[4243] Nicholas Zutt. Book review: Cosmic analogies: How natural systems emulate the universe by Valerio Faraoni, London, World Scientific, 2022, 262 pp., USD 98.00 (E-Book), ISBN 978-1-80061-342-3. Scope: monograph. Level: undergraduate. Contemporary Physics, 63(3):244–245, 2022. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:2022:BRP

[4244] Jack Reid. Book review: The physics of sound waves: music, instruments, and sound equipment, second edition, by Panos Photinos, Bristol, IOP Science, 2021, 269 pp., $50.00 (hardback), $40.00 (e-book), ISBN: 978-0-7503-3539-3. Scope: textbook. Level: undergraduate. Contemporary Physics, 63(3):245, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Schiller:2022:PSM

[4245] S. Schiller. Precision spectroscopy of molecular hydrogen ions: an introduction. Contemporary Physics, 63(4):247–279, 2022. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Stranks:2022:HPP

[4246] Samuel D. Stranks. Halide perovskite photovoltaics. Contemporary Physics, 63(4):280–304, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Boothroyd:2022:TEB

[4247] A. T. Boothroyd. Topological electronic bands in crystalline solids. Contemporary Physics, 63(4):305–327, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Esposito:2022:BRQ

[4248] Salvatore Esposito. Book review: The quantum internet — the second quantum revolution by Peter P. Rohde, Cambridge, Cambridge University Press, 2020, 344+xviii pp., £49.99 (hardback), ISBN: 978-1-108-49145-7. Scope: monograph. Level: postgraduate, researcher, specialist, engineers. Contemporary Physics, 63(4):328, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:2022:BRT

[4249] Jack Reid. Book review: Thermodynamics of complex systems: principles and applications by Vladimir N. Pokrovskii, London, IOP Science, 2020, 321 pp., $190.00 (hardback), ISBN 978-0-7503-3449-5. Scope: textbook. Level: undergraduate. Contemporary Physics, 63(4):329, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Humaira:2022:BRF

[4250] Humaira and Erin Ficrah Huda. Book review: Flexible supercapacitors: materials and applications edited by Guozhen Shen, Zheng Lou and Di Chen, Hoboken, Wiley, 2022, 331 pp., $195.00 (hardback), ISBN 978-1-119-50616-4. Scope: textbook. Level: undergraduate, advanced undergraduate, postgraduate, researcher, scientist, engineer. Contemporary Physics, 63(4):329–330, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Falcone:2022:BRP

[4251] Francisco Falcone. Book review: Printed and flexible sensor technology by Subhas Mukhopadhyay and Anindya Nag, Bristol, IOP Publishing, 2021, 300 pp., $190.00 (hardback), ISBN: 978-0-7503-3439-6. Scope: reference. Level: postgraduate. Contemporary Physics, 63(4):330–331, 2022. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hardi:2022:BRO

[4252] Anla Fet Hardi and Erin Ficrah Huda. Book review: Organic solar cells: materials design, technology and commercialization edited by Liming Ding, Weinheim, Wiley, 2022, 720 pp., $156.00 (hardback), ISBN 978-3-527-83366-5. Scope: textbook. Level: advanced undergraduate, postgraduate, researcher, scientist, engineer. Contemporary Physics, 63(4): 331–332, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2022:BRJ

[4253] Toni Shiroka. Book review: A journey into reciprocal space: a crystallographer’s perspective, 2nd edition, by Anthony Michael Glazer, IOP Science, Bristol, 2021, 274 pp., 145 Euro, ISBN: 978-0-7503-3875-2 (Hardback). Scope: Textbook. Level: Undergraduate, advanced Undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 63(4):332–333, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bertolotti:2022:BRN

[4254] Mario Bertolotti. Book review: Nanotechnology: trends and future applications by Muhammad Bilal Tahir, Muhammad Rafique and Muhammad Sagir, Singapore, Springer International Publishing, 2021, 143 pp., $79.99 (e-Book), ISBN 978-981-15-9437-3. Scope: monograph. Level: non-specialists. Contemporary Physics, 63(4):333, 2022. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuan:2022:BRS

[4255] Miguel A. F. Sanjuán. Book review: Set theory and its applications in physics and computing by Yair Shapira, Singapore, World Scientific, 2022, 480 pp., £100 (hardback), ISBN: 978-981-12-6177-0. Scope: monograph. Level: advanced undergraduate, postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 63(4): 333–334, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zutt:2022:BRD

[4256] Nicholas Zutt. Book review: Discovering the nature of light: the science and the story by Norval Fortson, Seattle, USA, University of Washington, 2022, 220 pp., 68 USD (hardback), ISBN: 978-981-12-4959-4. Scope: general interest. Level: general readership. Contemporary Physics, 63(4): 334–335, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2022:BRE

[4257] A. H. Harker. Book review: Energy materials by Aliaksandr S. Bandarenka, Abingdon, Routledge Taylor & Francis Group, 2022, xv + 236 pp., £74.99 (e-book), ISBN: 978-0-367-45736-5. Scope: monograph. Level: postgraduate, advanced undergraduate, early career researcher. Contemporary Physics, 63(4):335–336, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Wandi:2022:BRC

[4258] Wandi Wandi and Fani Mardianti. Book review: Cartoon physics: a graphic novel guide to solving physics problems by Scott Calvin and Kirin Emlet Furst, Boca Raton, CRC Press, 2022, 273 pp., £31.99 (paperback), ISBN 978-1-138-59878-2. Scope: general interest. Level: general readership, undergraduate. Contemporary Physics, 63(4):336–337, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2022:BRP

[4259] A. H. Harker. Book review: The physics of phonons, 2nd edition, by Gyaneshwar P. Srivastava, Boca Raton, FL/Abingdon, UK, CRC Press, 2023, xxi+434 pp., £180 (hardback), ISBN 978-0-367-68526-3. Scope: monograph. Level: advanced undergraduate, graduate, early stage researcher. Contemporary Physics, 63(4):337, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sharma:2022:BRQ

[4260] Mohit Sharma. Book review: Quantum transport in nanostructures and molecules: an introduction to molecular electronics by Colin John Lambert, Bristol, IOP Publishing, 2021, 445 pp., $190 (hardback), $159 (ebook), ISBN: 978-0-7503-3639-0. Scope: review. Level: researcher. Contemporary Physics, 63(4):337–339, 2022. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Marsh:2023:AA

[4261] David J. E. Marsh. Axions for amateurs. Contemporary Physics, 64 (1):1–18, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Mercer:2023:SPL

[4262] William J. Mercer and Yuri A. Pashkin. Superconductivity: the path of least resistance to the future. Contemporary Physics, 64(1):19–46, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Padovani:2023:ELT

[4263] Paolo Padovani and Michele Cirasuolo. The Extremely Large Telescope. Contemporary Physics, 64(1):47–64, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Perryman:2023:SA

[4264] Michael Perryman. Space astrometry. Contemporary Physics, 64(1): 65–81, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2023:BRA

[4265] Toni Shiroka. Book review: Application-driven quantum and statistical physics: a short course for future scientists and engineers, Volume 3: Transitions by Jean-Michel Gillet, London, World Scientific, 2020, £80 (E-Book), ISBN: 978-1-78634-788-6. Scope: textbook. level: undergraduate, advanced undergraduate, early career researcher, researcher, Bachelor and Master’s level engineering and physics students. Contemporary Physics, 64(1):82–83, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2023:BRF

[4266] Toni Shiroka. Book review: Fundamentals and applications of magnetic materials by Kannan M. Krishnan, Oxford, Oxford University Press, 2020, 795 pp., £40 (Paperback), ISBN: 978-0-19-886204-8. Scope: reference, handbook. level: undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher, scientist, engineers. Contemporary Physics, 64(1):83–84, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Zutt:2023:BRO

[4267] Nicholas Zutt. Book review: An object-oriented Python cookbook in quantum information theory and quantum computing by M. S. Ramkarthik and Pranay Barkataki, Boca Raton, FL, CRC Press, Taylor & Francis Group, 2023, 220 pp., 104 USD (hardback), ISBN: 978-1-032-25607-8. Scope: handbook. level: postgraduate. Contemporary Physics, 64(1): 84–85, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belsley:2023:BRI

[4268] Michael Belsley. Book review: Invitation to quantum mechanics by Daniel F. Styer, Singapore, World Scientific Publishing Co. Pte. Ltd, 2022, 289 pp., £30 (e-book), ISBN: 978-981-12-4792-7; £40 (paperback), ISBN: 978-981-12-4928-0; £80 (hardback), ISBN: 978-981-12-4790-3. Scope: textbook. level: undergraduate. Contemporary Physics, 64(1):85, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2023:BRC

[4269] A. H. Harker. Book review: Commonly Asked Questions in thermodynamics, 2nd edition, by Marc J. Assael, Geoffrey C. Maitland, Thomas Maskow, Urs von Stockar, William A. Wakeham, and Stefan Will, Boca Raton, FL, USA and Abingdon, UK, CRC Press, 2022, xxiv + 461 pp., £170 (hardback), ISBN: 978-0-367-33891-6; ISBN: 978-1-032-27589-5 (paperback); ISBN: 978-0-429-32952-4 (e-book). Scope: textbook. level: advanced undergraduate, graduate, early stage researcher. Contemporary Physics, 64(1):86, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2023:BRO

[4270] A. H. Harker. Book review: Offbeat physics: machines, meditations and misconceptions, edited by P. I. C. Teixeira, Boca Raton, FL, and Abingdon, UK, CRC Press, 2022, xvi + 258 pp., £99.99 (hardback), ISBN: 978-1-032-02116-4; ISBN: 978-1-032-03393-8 (paperback); ISBN: 978-1-003-18710-3 (e-book). Scope: article collection. level: undergraduate. Contemporary Physics, 64(1):86–87, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2023:BREa

[4271] Stephen J. Blundell. Book review: Elementary Introduction to Quantum Geometry by Jan Ambjorn, Abingdon, CRC Press, 2023, 292 pp., £74.99 (hardcover), ISBN: 978-1-032-33555-1. Scope: textbook. level: researcher. Contemporary Physics, 64(1):87, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dobson:2023:BRT

[4272] Peter J. Dobson. Book review: Transition metal oxides for electrochemical energy storage, edited by Jagit Nanda and Veronica Augustyn, Chichester, UK, Wiley, 2022, 432 pp., £138.99 (e-Book), ISBN: 978-3-527-34493-2. Scope: review, reference, monograph. level: advanced undergraduate, postgraduate, early career researcher, researcher, teacher, specialist, scientist, engineers. Contemporary Physics, 64(1):87–88, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2023:BRD

[4273] Stephen J. Blundell. Book review: Density functional theory: a practical introduction, 2nd edition, by David S. Sholl and Janice A. Steckel, Hoboken, NJ, Wiley, 2023, 204 pp., £91.25 (hardcover), ISBN 978-1-119-84086-2. Scope: textbook. level: graduate student, researcher. Contemporary Physics, 64(1):88, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2023:BRS

[4274] A. H. Harker. Book review: So simple a beginning: how four physical principles shape our living world, by Raghuveer Parthasarathy, Princeton, NJ, USA and Woodstock, Oxfordshire, UK, Princeton University Press, 2022, x + 326 pp., £30 (hardback), ISBN: 978-0-691-20040-8; £30 (e-book), ISBN: 978-0-691-23161-7. Scope: general interest. level: general reader. Contemporary Physics, 64(1):88–89, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beauty:2023:BRL

[4275] Aras Beauty and Bimo Bramantio. Book review: Life in the universe, 5th edition, by Jeffrey Bennett, Seth Shostak, Nicholas Schneider and Meredith MacGregor, Princeton, Princeton University Press, 2022, 544 pp., £84.00 (e-book), ISBN: 978-0-691-24178-4. Scope: review. level: general readership, undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher, teacher, scientist, engineers. Contemporary Physics, 64(1):89–90, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2023:BRW

[4276] Miguel A. F. Sanjuán. Book review: When things grow many. Complexity, universality and emergence in nature, by L. S. Schulman, Oxford, UK, Oxford University Press, 2022, 304 pp., £45 (hardback), ISBN: 978-0-19-886188-1. Scope: monograph. level: advanced undergraduate, postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 64(1):90–91, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Buchmueller:2023:LSA

[4277] Oliver Buchmueller, John Ellis, and Ulrich Schneider. Large-scale atom interferometry for fundamental physics. Contemporary Physics, 64(2): 93–110, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clark:2023:NCS

[4278] S. J. Clark and Z. Hawkhead. Non-collinear spin in electronic structure calculations. Contemporary Physics, 64(2):111–126, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lancaster:2023:QSL

[4279] Tom Lancaster. Quantum spin liquids. Contemporary Physics, 64(2): 127–146, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pedernales:2023:OFS

[4280] Julen S. Pedernales and Martin B. Plenio. On the origin of force sensitivity in tests of quantum gravity with delocalised mechanical systems. Contemporary Physics, 64(2):147–163, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2023:BRI

[4281] A. H. Harker. Book review: Introduction to the mathematical physics of nonlinear waves, 2nd edition, by Minoru Fujimoto, Bristol, UK, IOP Publishing, 2021, xv + 165 pp., £75.00 (hardback), ISBN: 978-0-7503-3757-1 (hbk); 978-0-7503-3760-1 (myPrint); 978-0-7503-3759-5 (ebook); 978-0-7503-3758-8 (mobi). Scope: textbook. level: advanced undergraduate, graduate. Contemporary Physics, 64(2):164, 2023. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Peters:2023:BRS

[4282] Thomas Peters. Book review: Statistical and thermal physics: with computer applications, 2nd edition by Harvey Gould and Jan Tobochnik, Princeton and Oxford, Princeton University Press, 2021, 528 pp., £75.00 (hardback), ISBN 978-0-691-20189-4. Scope: text book. level: undergraduate. Contemporary Physics, 64(2):164–165, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2023:BRH

[4283] M. Jamil. Book review: High school and undergraduate physics practicals with 3D simulations, 1st edition by Robert Lucas, UK, Taylor & Francis Group, CRC Press, 2022, 246 pp., 148 illus., $64 (£39) (Paperback), ISBN: 978-1-032-19739-5, ISBN: 978–1-032-2029-0. Scope: monograph. level: high school & undergraduate students, lecturers, researchers. Contemporary Physics, 64(2):165, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Beauty:2023:BRB

[4284] Aras Beauty and Fitri Natalia. Book review: Bioelectromagnetism history, foundations and applications, by Shoogo Ueno and Tsukasa Shigemitsu, Boca Raton, CRC Press, 2022, 346 pp., EUR150.00 (hardback), ISBN 978-1-032-00121-0. Scope: textbook. level: advanced undergraduate, postgraduate, researcher, scientist. Contemporary Physics, 64 (2):166, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gangopadhyay:2023:BRS

[4285] Gautam Gangopadhyay. Book review: The sky is for everyone: women astronomers in their own words, edited by Virginia Trimble and David A. Weintraub, Princeton and Oxford, Princeton University Press, 2022, E-Book, ISBN: 978-0-691-20710-0. Scope: general interest. level: general readership. Contemporary Physics, 64(2):166–167, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ashworth:2023:BRP

[4286] Stephen H. Ashworth. Book review: Physics behind music: an introduction, by Bryan H. Suits, Cambridge, UK, Cambridge University Press, 2023, E-Book, ISBN 978-1-108-95315-3. Scope: review. level: general readership, non-specialists. Contemporary Physics, 64(2):167–168, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bowen:2023:QLM

[4287] W. P. Bowen, Helen M. Chrzanowski, Dan Oron, Sven Ramelow, Dmitry Tabakaev, Alex Terrasson, and Rob Thew. Quantum light microscopy. Contemporary Physics, 64(3):169–193, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pietropaolo:2023:PMN

[4288] Antonino Pietropaolo. The physical mechanisms of neutron detection. Contemporary Physics, 64(3):194–223, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rosanov:2023:HCE

[4289] N. N. Rosanov, M. V. Arkhipov, R. M. Arkhipov, and A. V. Pakhomov. Half-cycle electromagnetic pulses and pulse electric area. Contemporary Physics, 64(3):224–243, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sukirno:2023:BRB

[4290] Sukirno Sukirno, Elvian Indah Nilamsari, and Aras Beauty. Book review: Biosensors for virus detection by Adil Denizli and Yeseren Saylan, Bristol, IOP Publishing, 2021, 184 pp., $190.000 (hardcover), ISBN 978-0-7503-3867-7. Scope: textbook. Level: advanced undergraduate, postgraduate, researcher, scientist. Contemporary Physics, 64(3):244–246, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gangopadhyay:2023:BRJ

[4291] Gautam Gangopadhyay. Book review: The joy of science by Jim Al-Khalili, Princeton and Oxford, Princeton University Press, 2022, 224 pp., $16.95/£13.99 (E-Book), ISBN: 978-0-691-23566-0. Scope: general interest. Level: general readership. Contemporary Physics, 64(3):247, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Maina:2023:BRS

[4292] Ezio Maina. Book review: A student’s guide to python for physical modeling, 2nd edition by Jesse M. Kinder and Philip Nelson, Princeton, NJ, Princeton University Press, 2021, 240 pp., $26.95 (paperback), ISBN 978-0-691-22365-0. Scope: textbook. Level: undergraduate. Contemporary Physics, 64(3):247–248, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bridges:2023:BRW

[4293] Thomas J. Bridges. Book review: Wavefronts and rays as characteristics and asymptotics, 3rd edition by Andrej Bona and Michael A. Slawinski, Singapore, World Scientific, 2020, 356 pp., £120 (hbk), £95 (ebook), ISBN 978-981-12-2646-5. Scope: reference. Level: postgraduate, advanced undergraduate. Contemporary Physics, 64(3):248, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jamil:2023:BRR

[4294] M. Jamil. Book review: Renewable energy: a first course, 3rd edition by Robert Ehrlich, Harold A. Geller, and John R. Cressman, Boca Raton, FL, Routledge Taylor & Francis Group, 2022, 562 pp., US$88.12 (paperback), ISBN: 978-0-367-76837-9. Scope: review. Level: undergraduate, advanced undergraduate, postgraduate, early career researcher. Contemporary Physics, 64(3):249–250, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2023:BRP

[4295] A. H. Harker. Book review: Physics and chemistry of interfaces (fourth edition) by Hans-Jürgen Butt, Karlheinz Graf and Michael Kappl, Weinheim, Germany, Wiley-VCH, 2023, xii + 458 pp., £40.99 (e-book), £65 (paperback), ISBN: 978-3-527-41405-5 (paperback), ISBN: 978-3-527-83617-8 (ePDF), ISBN: 978-3-527-83616-1 (ePub). Scope: monograph. Level: advanced undergraduate, graduate, early stage researcher. Contemporary Physics, 64(3):250–251, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2023:BRH

[4296] Stephen J. Blundell. Book review: How to be a quantum mechanic by Charles G. Wohl, Boca Raton, Routledge Taylor & Francis Group, 2022, 380 pp., £42 (paperback), ISBN: 978-1-032-25602-3. Scope: textbook. Level: undergraduate student. Contemporary Physics, 64(3):251, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sharma:2023:BRQa

[4297] Mohit Sharma. Book review: Quantum mechanics II advanced topics by S. Rajasekar and R. Velusamy, USA, Routledge Taylor & Francis Group, 2022, £69.59 (E-Book), ISBN: 978-0-367-77000-6. Scope: review. Level: early career researcher. Contemporary Physics, 64(3):251–252, 2023. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Chiesa:2023:QIP

[4298] Alessandro Chiesa, Emilio Macaluso, and Stefano Carretta. Quantum information processing with molecular nanomagnets: an introduction. Contemporary Physics, 64(4):253–281, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rudelius:2023:IWG

[4299] Tom Rudelius. An introduction to the Weak Gravity Conjecture. Contemporary Physics, 64(4):282–295, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Pajic-Lijakovic:2023:PCS

[4300] Ivana Pajic-Lijakovic, Milan Milivojevic, and Peter V. E. McClintock. Physics of cancer spreading through epithelium. Contemporary Physics, 64(4):296–314, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2023:BRD

[4301] Miguel A. F. Sanjuán. Book review: Dynamics of planetary systems. A review of by Scott Tremaine, Princeton, NJ, Princeton University Press, 2023, 604 pp., £125 (paperback), ISBN 978-0-691-20712-4. Scope: monograph, textbook. Level: postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 64(4):315–318, 2023. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Liliya:2023:BRN

[4302] Frolova Liliya. Book review: Nanotechnology for light pollution reduction by Tuan Anh Nguyen and Ram K. Gupta, Boca Raton, FL, CRC Press, 2022, 330 pp., £135.00 (E-Book), ISBN: 978-1-032-02120-1. Scope: review. Level: scientist. Contemporary Physics, 64(4):319, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gangopadhyay:2023:BRP

[4303] Gautam Gangopadhyay. Book review: Philosophy of physics: a new introduction by Robert P. Crease, Bristol, UK, IOP Publishing, 2023, 62 pp., $40 (e-Book), ISBN: 978-0-7503-2636-0. Scope: Review. Level: General readership, non-specialists. Contemporary Physics, 64(4):319– 320, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2023:BREb

[4304] Stephen J. Blundell. Book review: Essentials of signals and systems by E. Martins, Hoboken, Wiley, 2023, 313 pp., £60 (paperback), ISBN: 978-1-119-90923-1. Scope: textbook. Level: undergraduate. Contemporary Physics, 64(4):320, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hazra:2023:BRP

[4305] Susmita Hazra. Book review: Physics of the Sun by Dermott J. Mullan, Routledge Taylor & Francis Group, 2022, £84.99 (hardback), ISBN: 978-0-367-71039-2. Scope: review. Level: postgraduate. Contemporary Physics, 64(4):321, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Salasnich:2023:BRM

[4306] Luca Salasnich. Book review: A multidisciplinary approach to quantum field theory, volume 1 by Michael Ogilvie, Bristol, UK, IOP Science, 2022, 300 pp., $89.91 (hardcover), ISBN: 978-0-7503-3227-9. Scope: textbook. Level: advanced undergraduate. Contemporary Physics, 64(4):

321–322, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Casazza:2023:BRP

[4307] Marco Casazza. Book review: Physics behind music by Brian H. Suits, Cambridge, UK, Cambridge University Press, 2023, 422 pp., EUR87.52 (hardback), EUR35.00 (paperback), EUR35.00 (E-book), ISBN: 978-1-108-94870-8. Scope: monograph. Level: non-specialists, undergraduate, researcher, teacher. Contemporary Physics, 64(4):322, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sharma:2023:BRQb

[4308] Mohit Sharma. Book review: Quantum mechanics non-relativistic and relativistic theory by Lukong Cornelius Fai, London, Routledge, Taylor & Francis Group, 2022, 552 pp., £115.00 (E-Book), ISBN: 978-1-032-22146-5. Scope: review. Level: advanced undergraduate. Contemporary Physics, 64(4):323–324, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Harker:2023:BRE

[4309] A. H. Harker. Book review: Experimental techniques in magnetism and magnetic materials by Sindhunil Barman Roy, Cambridge, UK, Cambridge University Press, 2022, xv + 311 pp., £110, ISBN: 978-1-108-48998-0 (hbk), ISBN: 978-1-009-27655-9 (ebook). Scope: monograph. Level: advanced undergraduate, graduate, early career researcher. Contemporary Physics, 64(4):324, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2023:BRS

[4310] Miguel A. F. Sanjuán. Book review: Stochastic thermodynamics: an introduction by Luca Peliti and Simone Pigolotti, Princeton, NJ, USA, Princeton University Press, 2021, 272 pp., £65 (hardback), ISBN: 978-0-691-20177-1. Scope: textbook. Level: postgraduate, early career researcher, researcher, specialist. Contemporary Physics, 64(4):324–325, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2023:BRM

[4311] Miguel A. F. Sanjuán. Book review: More than the sum of the parts. Complexity in physics and beyond by Helmut Satz, Oxford, Oxford University Press, 2022, 160 pp., £21.99 (Hardback), ISBN: 978-0-19-286417-8. Scope: general interest. Level: general readership, non-specialists, undergraduate, researcher, teacher, scientist. Contemporary Physics, 64(4):

325–326, 2023. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Blundell:2024:E

[4312] Stephen Blundell. Editorial. Contemporary Physics, 65(1):1, 2024. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Heinosaari:2024:CQC

[4313] Teiko Heinosaari and Mark Hillery. Can a qudit carry more information than a dit? Contemporary Physics, 65(1):2–11, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ellis:2024:ARF

[4314] Simon Charles Ellis and Joss Bland-Hawthorn. Astrophotonics: recent and future developments. Contemporary Physics, 65(1):12–39, 2024. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kxfchne:2024:NMR

[4315] H. Kühne and Y. Ihara. Nuclear magnetic resonance spectroscopy in pulsed magnetic fields. Contemporary Physics, 65(1):40–61, 2024. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reid:2024:BRB

[4316] Jack Reid. Book review: Back to the moon: the next giant leap for humankind by Joseph Silk, Princeton, Princeton University Press, 2022, 287 pp., $29.95 (hardback), ISBN: 978-0-691-21523-5. Scope: general interest. Level: general readership. Contemporary Physics, 65(1):62, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ardiansyah:2024:BRN

[4317] Abd Aziz Ardiansyah and Devi Yulianty Surya Atmaja. Book review: No-frills physics: a concise study guide for algebra-based physics by Matthew D. McCluskey, Boca Raton, CRC Press, 2019, 260 pp., $32.22 (paperback), ISBN 978-1-138-58387-0. Scope: handbook; guide; textbook. Level: general readership; pre-university. Contemporary Physics, 65(1):63, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kamran:2024:BRM

[4318] Muhammad Kamran. Book review: Mathematical physics for nuclear experiments by Andrew E. Ekpenyong, Boca Raton, FL, CRC Press, 2022, 280 pp., £58.99 (Paperback), £150.00 (Hardback), £53.09 (E-book), ISBN: 978-1-032-10499-7. Scope: reference, textbook. Level: advanced undergraduate, postgraduate. Contemporary Physics, 65(1):63–64, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kamran:2024:BRQ

[4319] Muhammad Kamran. Book review: Quantum mechanics I the fundamentals by S. Rajasekar and R. Velusamy, Boca Raton, FL, CRC Press, 2022, 607 pp., £86.99 (hardback), £78.29 (eBook), ISBN: 978-0-367-76998-7. Scope: reference, textbook. Level: advanced undergraduate. Contemporary Physics, 65(1):64–65, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2024:BRM

[4320] Miguel A. F. Sanjuán. Book review: A mind over matter by Andrew Zangwill, Oxford, UK, Oxford University Press, 2021, 416 pp., £30.99 (hardback), ISBN 978-0-19-886910-8. Scope: biography. Level: general readership, non-specialists. Contemporary Physics, 65(1):65–66, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Muki:2024:BRP

[4321] Beata Graceshela Muki and Sumiati. Book review: Primary physical science education (PPSE): an imaginative approach to encounters with nature Hans U. Fuchs and Federico Corni, Switzerland, Springer Cham, 2023, xiv + 347 pp., EUR39.99 (paperback), Series: 1, ISBN: 978-3-031-43955-1. Scope: general interest. Level: general readership, undergraduate, researcher, teacher. Contemporary Physics, 65(1):66–67, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kamilah:2024:BRF

[4322] Dendy Siti Kamilah and Sumiati. Book review: Fascinating problems for young physicists: discovering everyday physics phenomena and solving them by Nenad Vukmirović and Vladimir Veljić, Cambridge, Cambridge University Press, 2022, 243 pp., $34.99 (Hardback), ISBN: 978-1-009-16028-5. Scope: general interest. Level: general readership, preuniversity, undergraduate, teacher. Contemporary Physics, 65(1):67–68, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hidayat:2024:BRS

[4323] Taufik Hidayat and Eduwin Saputra. Book review: Surfacefunctionalized ceramics: for biotechnological and environmental applications, 1st edition, by Laura Treccani and Fabian Meder, Weinheim, WI-LEY VCH, 2023, 464 pp., US$156.00 (Ebook), ISBN: 978-3-527-69805-9.

Scope: review, reference, monograph. Level: undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher, scientist, engineers. Contemporary Physics, 65(1):68–69, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Aryani:2024:BRU

[4324] Siska Dewi Aryani and Sri Zakiyah. Book review: Understanding the physics of toys: principles, theory and exercises by S. Rajasekar, R. Velusamy and Miguel A. F. Sanjuan, Singapore, World Scientific, 2023, 470 pp., $58.00 (paperback), ISBN: 978-981-12-6800-7. Scope: general interest. Level: general readership. Contemporary Physics, 65(1):69– 70, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fritz:2024:BRG

[4325] Alexander Fritz. Book review: ”Great mysteries in astrophysics“: a guide to what we don’t know by Nicole M. Lloyd-Ronning, Bristol, IOP Publishing, 2022, 164 pp., EUR46.30 (e-book), ISBN: 978-0-7503-4051-9 (e-book); ISBN: 978-0-7503-4049-6 (print). Scope: textbook. Level: undergraduate, advanced undergraduate, postgraduate. Contemporary Physics, 65(1):70–71, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ping:2024:BRF

[4326] Yow Ai Ping. Book review: Fascinating problems for young physicists by Nenad Vukmirović and Vladimir Veljić, Cambridge, Cambridge University Press, 2022, 300 pp., $84.99, ISBN: 978-1-009-16027-8. Scope: reference. Level: high school and undergraduate students, lecturers. Contemporary Physics, 65(1):71, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fritz:2024:BRT

[4327] Alexander Fritz. Book review: A traveler’s guide to the stars by Les Johnson, Princeton, Princeton University Press, 2022, 219 pp., $27.95/£22.00 (hardcover), $18.95/£15.99 (e-book), ISBN: 978-0-691-24007-7. Scope: monograph. Level: general readership, undergraduate. Contemporary Physics, 65(1):71–73, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2024:BRQ

[4328] Francisco Caruso. Book review: Quantum drama: from the Bohr– Einstein debate to the riddle of entanglement by Jim Baggott and John L. Heilbron, Oxford, UK, Oxford University Press, 2024, 352 pp., $32.99 (hardcover), ISBN 978-0-19-193849-8. Scope: review. Level: postgraduate, early career researcher, researcher, teacher. Contemporary Physics, 65(1):73–74, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Croon:2024:GWP

[4329] Djuna Lize Croon and David James Weir. Gravitational waves from phase transitions. Contemporary Physics, 65(2):75–96, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Belhouari:2024:NTP

[4330] Aissa Belhouari. Numerical tools in particle physics. Contemporary Physics, 65(2):97–111, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barbieri:2024:QME

[4331] Marco Barbieri, Ilaria Gianani, Aaron Z. Goldberg, and Luis L. Sánchez-Soto. Quantum multiphase estimation. Contemporary Physics, 65(2): 112–124, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Shiroka:2024:BRJ

[4332] Toni Shiroka. Book review: John David Jackson: a course in quantum mechanics A review of by Robert N. Cahn and John David Jackson, Hoboken, NJ, USA, Wiley, 2024, 400 pp., £76.50 (hardback), ISBN: 978-1-119-88038-7. Scope: textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher, scientist. Contemporary Physics, 65(2):125–127, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Clayton:2024:BRC

[4333] John D. Clayton. Book review: Crystalline solid state physics: an interactive guide by Meng Lee Leek, Bristol, UK, IOP, 2023, 247 pp., 120 USD (hardback), ISBN: 978-0-7503-5217-8. Scope: textbook. Level: undergraduate. Contemporary Physics, 65(2):128–129, 2024. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sharma:2024:BRA

[4334] Prince Sharma. Book review: Atomistic simulations of glasses: fundamentals and applications edited by Jincheng Du and Alastair N. Cormack, Hoboken, John Wiley & Sons, Inc., 2022, 560 pp., $169 (e-book), ISBN: 978-1-118-93907-9, (e-book), ISBN: 978-1-118-93906-2 (hardback).

Scope: edited book. Level: postgraduate, early career researcher, researcher, scientist, engineers. Contemporary Physics, 65(2):129–130, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Vogel:2024:BRS

[4335] Manuel Vogel. Book review: Supermassive black holes by Andrew King, Cambridge, UK, Cambridge University Press, 2023, 308 pp., GBP 68.55 (hardcover), ISBN: 978-1-108-48805-1. Scope: monograph. Level: undergraduate and above. Contemporary Physics, 65(2):130–131, 2024. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

MacKay:2024:BRI

[4336] Robert S. MacKay. Book review: Introduction to Arnold’s proof of the Kolmogorov–Arnold–Moser theorem by Achim Feldmeier, Boca Raton, Routledge Taylor & Francis Group, 2022, £44.99 (e-book), ISBN: 978-1-032-26065-5. Scope: monograph. Level: specialist. Contemporary Physics, 65(2):131–132, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reddy:2024:BRC

[4337] Y. S. Reddy. Book review: Crystalline solid state physics by Meng Lee Leek, Bristol, UK, IOP, 2023, xii + 244 pp., $120.00 (e-book), ISBN: 978-0-7503-5217-8. Scope: textbook. Level: undergraduate. Contemporary Physics, 65(2):132–133, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2024:BRLa

[4338] B. Ishak. Book review: The little book of exoplanets by Joshua N. Winn, Princeton, Princeton University Press, 2023, 304 pp., £18.99 (hardback), ISBN: 978-0-691-21547-1. Scope: general interest. Level: general readership, non-specialists. Contemporary Physics, 65(2):133, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2024:BRP

[4339] B. Ishak. Book review: Particle cosmology and astrophysics by Dan Hopper, Princeton, Princeton University Press, 2024, 448 pp., £62 (hardback), ISBN: 978-0-691-23504-2. Scope: reference, textbook. Level: advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 65(2):133–134, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2024:BRS

[4340] B. Ishak. Book review: The sixth element: how carbon shapes our world by Theodore P. Snow and Don Brownlee, Princeton, Princeton University Press, 2024, 264 pp., £22.00 (hardback), ISBN: 978-0-691-12588-6. Scope: general interest. Level: general readership, non-specialists. Contemporary Physics, 65(2):134, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2024:BRM

[4341] J. Rogel-Salazar. Book review: Machine learning for physicists by Sadegh Raeisi and Sedighe Raeisi, Bristol, UK, IOP, 2023, 94 pp., $95 (EBook), ISBN: 978-0-7503-4957-4. Scope: review. Level: undergraduate. Contemporary Physics, 65(2):134–135, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2024:BRLb

[4342] B. Ishak. Book review: Life and language beyond earth by Raymond Hickey, Cambridge, Cambridge University Press, 2023, 695 pp., £25.00 (hardback), ISBN 978-1-009-22641-7. Scope: general interest. Level: general readership, non-specialists. Contemporary Physics, 65(2):135, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barber:2024:BRI

[4343] D. C. Barber. Book review: An introduction to the physics of nuclear medicine (second edition) by Laura Harkness-Brennan, IOP, 2023, 125 pp., £73.40 (hardback), ISBN 978-0-7503-5197-3. Scope: review. Level: general readership, pre-university, undergraduate, advanced undergraduate, postgraduate. Contemporary Physics, 65(2):136, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2024:BRE

[4344] Gerard Gilmore. Book review: The Einsteinian revolution: the historical roots of his breakthroughs by Hanoch Gutfreund and Jürgen Renn, Princeton University Press, 2023, 272 pp., (e-book), ISBN: 978-0-691-16876-0. Scope: review. Level: general readership. Contemporary Physics, 65(2):136–137, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2024:BRA

[4345] B. Ishak. Book review: Atlas of the Messier objects: highlights of the deep sky, 2nd edition, by Ronald Stoyan, Cambridge, Cambridge University Press, 2023, 372 pp., £59.99 (hardback), ISBN: 978-1-009-36406-5. Scope: general interest. Level: general readership. Contemporary Physics, 65 (2):137, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2024:BRLc

[4346] B. Ishak. Book review: The large scale structure of space-time, 50th anniversary edition, by Stephen W. Hawking and George F. R. Ellis, Cambridge, Cambridge University Press, 2023, 410 pp., £27.99 (hardback), ISBN: 978-1-009-25315-4. Scope: monograph. Level: specialist. Contemporary Physics, 65(2):138, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2024:BRB

[4347] Gerard Gilmore. Book review: Big Science in the 21st century by Panagiotis Charitos, Theodore Arabatzis, Harry Cliff, Gunther Dissertori, Juliette Forneris and Jason Li-Ying, IOP, 2023 (E-Book), ISBN 978-0-7503-3631-4. Scope: reference. Level: specialist. Contemporary Physics, 65(2):138, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Gilmore:2024:BRQ

[4348] Gerry Gilmore. Book review: Quantum drama: from the Bohr–Einstein debate to the riddle of entanglement by Jim Baggott and John L. Heilbron, Oxford, OUP, 2024, 336 pp., GBP26.99 (hardback), ISBN: 978-0-19-193849-8. Scope: monograph. Level: specialist. Contemporary Physics, 65(2):139, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Probert:2024:BRA

[4349] Matt Probert. Book review: AI for physics by Volker Knecht, London, Routledge Taylor & Francis Group, 2022, £22 (e-book), ISBN: 978-1-032-15169-4. Scope: reference. Level: undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 65(2):139–140, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Sanjuxen:2024:BRD

[4350] Miguel A. F. Sanjuán. Book review: Data science for complex systems by Anindya S. Chakrabarti, K. Shuvo Bakar and Anirban Chakraborti, Cambridge, Cambridge University Press, 2023, 304 pp., £54.99 (hardback), ISBN: 978-1-108-84479-6. Scope: monograph. Level: advanced undergraduate, postgraduate, early career researcher, researcher, specialist, scientist. Contemporary Physics, 65(2):140–141, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2024:BRC

[4351] J. Rogel-Salazar. Book review: Classical mechanics and relativity, 2nd edition by Harald J. W. Muller-Kirsten, WSPC, 5 Top Tuck Link, Singapore, 2024, 588 pp., £70 (Softcover), ISBN: 978-981-12-8741-1. Scope: review. Level: advanced undergraduate. Contemporary Physics, 65(2): 141–142, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rogel-Salazar:2024:BRP

[4352] J. Rogel-Salazar. Book review: Physics behind music: an introduction by Bryan H. Suits, Cambridge, CUP, 2023, 406 pp., £29.99 (E-Book), ISBN: 978-1-108-94870-8. Scope: review. Level: undergraduate. Contemporary Physics, 65(2):142, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Kamran:2024:BRF

[4353] Muhammad Kamran. Book review: Fourier Transform and its applications using Microsoft EXCEL (second edition) by Shinil Cho, Bristol, IOP, 2023, 189 pp., US$92.70/GBP70 (hard cover), ISBN: 978-0-7503-6044-9 (ebook); 978-0-7503-6042-5 (print); 978-0-7503-6043-2 (mobi). Scope: reference. Level: undergraduate/beginners. Contemporary Physics, 65(2):142–143, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Reddy:2024:BRP

[4354] Y. S. Reddy. Book review: Physics of matter by George C. King, Wiley, UK, 2023, xviii + 446 pp., $56.00 (E-Book), ISBN: 978-1-119-46852-3. Scope: textbook. Level: undergraduate. Contemporary Physics, 65(2): 143–144, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Fritz:2024:BRW

[4355] Alexander Fritz. Book review: When galaxies were born: the quest for cosmic dawn by Richard S. Ellis, Princeton, NJ, Princeton University Press, 2022, 223 pp., $33.00/£28.00 (hardback), ISBN: 978-0-691-21130-5. Scope: biography. Level: general readership, non-specialists. Contemporary Physics, 65(2):144–145, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Caruso:2024:BRE

[4356] Francisco Caruso. Book review: The Einsteinian revolution: the historical roots of his breakthroughs by Jürgen Renn and Hanoch Gutfreund, Princeton and Oxford, Princeton University Press, 2023, 272 pp., $32.00/£28.00 (e-book), ISBN: 978-0-691-16876-0. Scope: review. Level: general readership, undergraduate, advanced undergraduate, postgraduate, early career researcher, researcher. Contemporary Physics, 65(2): 145–148, 2024. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Baggott:2009:AFW

[4357] J. E. Baggott. Atomic: the First War of Physics and the Secret History of the Atomic Bomb, 1939–49. Icon, London, UK, 2009. ISBN 1-84831-044-7 (hardcover). xiv + 576 + 16 pp. LCCN QC773 .B24 2009.

Baggott:2010:FWP

[4358] Jim Baggott. The First War of Physics: The Secret History of the Atomic Bomb, 1939–1949. Pegasus Books, New York, NY, USA, 2010. ISBN 1-60598-084-6. xiv + 576 pp. LCCN QC773 .B24 2010.

Bernstein:2014:CBO

[4359] Jeremy Bernstein. A Chorus of Bells and Other Scientific Inquiries. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2014. ISBN 981-4578-94-0 (hardcover). xi + 275 pp. LCCN Q173 .B53 2014.

Bishop:2017:FPA

[4360] Christopher J. Bishop and Y. (Yuval) Peres. Fractals in Probability and Analysis, volume 162 of Cambridge studies in advanced mathematics. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-13411-0 (hardcover), 1-316-86692-0, 1-316-46023-1 (e-book). ix + 402 pp. LCCN QA614.86 .B57 2017.

Briggs:2011:SDC

[4361] Asa Briggs. Secret Days: Code-Breaking in Bletchley Park. Frontline Books, London, UK, 2011. ISBN 1-84832-615-7. xix + 202 + 26 pp. LCCN D810.C88 B75 2011.

Chelikowsky:2018:IQM

[4362] James R. Chelikowsky. Introductory Quantum Mechanics with MATLAB For Atoms, Molecules, Clusters, and Nanocrystals. Wiley-VCH, Weinheim, Germany, 2018. ISBN 3-527-40926-2 (paperback), 3-527-65501-8 (e-book), 3-527-65499-2 (Mobipocket ebook), 3-527-65500-X (ePub ebook), 3-527-65501-8 (PDF ebook). 240 pp. LCCN QC174.17.D37 .C445

2019. URL http://swbplus.bsz-bw.de/bsz506870707cov.htm;
http://www.wiley-vch.de/publish/dt/books/ISBN978-3-527-40926-
6/.

Cooper:2017:PDH

[4363] Malcolm (Malcolm J.) Cooper and Jim Grozier. Precise Dimensions: a History of Units from 1791–2018. IOP expanding physics, [IOP release 4]. Institute of Physics Publishers, Bristol, UK, 2017. ISBN 0-7503-1485-0 (hardcover), 0-7503-1486-9 (mobi), 0-7503-1487-7 (e-book). ISSN 2053-2563. 181 (est.) pp. LCCN QC86 .C667 2017eb. URL http://

iopscience.iop.org/book/978-0-7503-1487-9.

delaPena:1996:QDI

[4364] Luis de la Peña and Ana María Cetto. The Quantum Dice: an Introduction to Stochastic Electrodynamics, volume 75 of Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-3818-9 (hardcover). xvi + 509 pp. LCCN QC631 .P383 1996. URL http://www.

loc.gov/catdir/enhancements/fy0813/95040168-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0813/95040168-t.html.

DiBartolo:2014:CSL

[4365] Baldassare Di Bartolo and Richard C. (Richard Conger) Powell. Crystal Symmetry, Lattice Vibrations and Optical Spectroscopy of Solids: a Group Theoretical Approach. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2014. ISBN 981-4579-20-3 (hardcover). xviii + 515 pp. LCCN QC176.8.L3 D53 2014.

Donoghue:1992:DSM

[4366] John F. Donoghue, Eugene Golowich, and Barry R. Holstein. Dynamics of the Standard Model, volume 2 of Cambridge monographs on particle physics, nuclear physics, and cosmology. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-36288-1 (hardcover). xviii + 540 pp. LCCN QC794.6.S75 D66 1991. URL

http://www.loc.gov/catdir/description/cam032/91019007.html;
 http://www.loc.gov/catdir/toc/cam032/91019007.html.

Ellis:2016:HCP

[4367] George Ellis. How Can Physics Underlie the Mind?: Top-Down Causation in the Human Context. Frontiers Collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-49807-3. xxvi + 482 pp. LCCN QC6 .E45 2016. URL http://link.

springer.com/10.1007/978-3-662-49809-5.

Forgan:2019:SFP

[4368] Duncan Forgan. Solving Fermi’s Paradox, volume 10 of Cambridge astrobiology. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-107-16365-X (hardcover), 1-316-68151-3 (e-book). xvii + 413 pp. LCCN QB54 .F77945 2019.

Fox:2014:WWN

[4369] Michael H. Fox. Why We Need Nuclear Power: the Environmental Case. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-934457-4 (hardcover). xii + 306 pp. LCCN TK9153 .F69 2014.

Frankel:2012:CDC

[4370] Henry Frankel. The Continental Drift Controversy. Cambridge University Press, Cambridge, UK, 2012. ISBN 0-521-87504-8 (vol. 1: hardcover), 0-521-87505-6 (vol. 2: hardcover), 0-521-87506-4 (vol. 3: hardcover), 1-107-01994-X (vol. 4: hardcover), 0-521-87507-2. LCCN ???? 4 volume set.

Fultz:2014:PTM

[4371] B. (Brent) Fultz. Phase Transitions in Materials. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06724-3. xxiv + 563 pp. LCCN QC175.16.P5 F86 2014. URL http://www.loc.

gov/catdir/enhancements/fy1408/2013046223-b.html;   http://
www.loc.gov/catdir/enhancements/fy1408/2013046223-d.html;
http://www.loc.gov/catdir/enhancements/fy1408/2013046223-t.
html.

Girifalco:2008:UFG

[4372] Louis Girifalco. The Universal Force: Gravity — Creator of Worlds. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-922896-5, 0-19-152797-1 (e-book), 1-281-15020-7. xx + 297 pp. LCCN QC178 .G555 2008.

Giustino:2014:MMU

[4373] Feliciano Giustino. Materials modelling using density functional theory: properties and predictions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-966243-6 (hardcover), 0-19-966244-4 (paperback), 0-19-163942-7 (e-book). xi + 286 pp.

LCCN TA403.6 .G58 2014. URL http://catdir.loc.gov/catdir/
enhancements/fy1503/2013953616-b.html;   http://catdir.loc.
gov/catdir/enhancements/fy1503/2013953616-d.html;   http://
catdir.loc.gov/catdir/enhancements/fy1503/2013953616-t.html.

Graham-Smith:2013:UCUa

[4374] Francis Graham-Smith. Unseen Cosmos: the Universe in Radio. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966058-1 (hardcover). vii + 238 + 8 pp. LCCN QB477 .G733 2013; QB477 .G74 2013.

Green:2016:CMC

[4375] Dan Green. Cosmology with MATLAB: with companion media pack. World Scientific Publishing Co. Pte. Ltd., Singapore, 2016. ISBN 981-310-839-8 (hardcover), 981-310-840-1 (paperback). xi + 250 pp. LCCN QB981 .G74 2016.

Hagar:2014:DCQ

[4376] Amit Hagar. Discrete or continuous? The quest for fundamental length in modern physics. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06280-2, 1-139-95816-X (e-book), 1-139-96027-X (e-book), 1-107-47734-4 (e-book), 1-139-95921-2 (e-book). xi + 267 pp. LCCN QC173.59.S65 H34 2014.

Hirose:2011:QTC

[4377] Kenji Hirose and Nobuhiko Kobayashi. Quantum transport calculations for nanosystems. Pan Stanford, Singapore, 2011. ISBN 981-4267-32-5 (hardcover). xvii + 523 pp. LCCN QC175.23 .H576 2014.

Iben:2012:SEP

[4378] Icko Iben. Stellar Evolution Physics. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-60253-X (hardback: set), 1-107-01657-6 (hardback: v. 2). xvi + 1495 (vol. 2) pp. LCCN QB806 .I24 2012.

Jackson:2004:KPC

[4379] Kenneth A. Jackson. Kinetic Processes: Crystal Growth, Diffusion, and Phase Transitions in Materials. Wiley-VCH, Weinheim, Germany, 2004. ISBN 3-527-30694-3. xv + 409 pp. LCCN QC174.9 .J33 2004. URL http:

//www.loc.gov/catdir/enhancements/fy0620/2004303465-b.html;
 http://www.loc.gov/catdir/enhancements/fy0620/2004303465-d.
html;   http://www.loc.gov/catdir/toc/wiley051/2004303465.html.

Jacobs:2014:QMT

[4380] Kurt (Kurt Aaron) Jacobs. Quantum Measurement Theory and its Applications. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02548-6 (hardcover). xii + 544 pp. LCCN QC174.17.M4 J33 2014.

Jaroszkiewicz:2014:PDT

[4381] George Jaroszkiewicz. Principles of Discrete Time Mechanics. Cambridge monographs on mathematical physics. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-03429-9. xiv + 365 pp. LCCN QC174.13 .J37 2014. URL http://www.loc.

gov/catdir/enhancements/fy1408/2013039555-b.html;   http://
www.loc.gov/catdir/enhancements/fy1408/2013039555-d.html;
http://www.loc.gov/catdir/enhancements/fy1408/2013039555-t.
html.

Koberlein:2013:ATC

[4382] Brian Koberlein and David D. Meisel. Astrophysics through computation: with Mathematica support. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-01074-8 (hardcover), 0-511-86317-9 (e-book). ix + 372 pp. LCCN QB462.2 .K63 2013.

Kraftmakher:2007:EDP

[4383] Yaakov Kraftmakher. Experiments and demonstrations in physics. Barllan Physics Laboratory. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2007. ISBN 981-256-602-3, 981-270-538-4 (rúst.). x + 533 pp. LCCN ????

Layzer:1990:CGO

[4384] David Layzer. Cosmogenesis: The Growth of Order in the Universe. Oxford Paperbacks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-506908-0. ix + 322 pp. LCCN ????

Lemons:2017:SGD

[4385] Don S. (Don Stephen) Lemons. A Student’s Guide to Dimensional Analysis. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-16115-0 (hardcover), 1-316-61381-X (paperback), 1-316-67616-1 (ebook). xii + 102 pp. LCCN TA347.D5 L46 2017.

Lenton:2011:RME

[4386] Tim (Timothy) Lenton and A. J. (Andrew J.) Watson. Revolutions That Made the Earth. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-958704-3. xii + 423 pp. LCCN QH343.4 .L46 2011. URL http://www.loc.

gov/catdir/enhancements/fy1110/2010038457-b.html;   http://
www.loc.gov/catdir/enhancements/fy1110/2010038457-d.html;
http://www.loc.gov/catdir/enhancements/fy1110/2010038457-t.
html.

Lewis:2017:FUL

[4387] Geraint F. Lewis, Luke A. Barnes, and Brian Schmidt. A fortunate universe: life in a finely-tuned cosmos. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-15661-0 (hardcover), 1-316-71648-1 (e-book), 1-316-71634-1, 1-316-71606-6, 1-316-66141-5. xvii + 373 pp. LCCN Q172.5.C45 L4845 2017.

Monk:2012:ICL

[4388] Ray Monk. Inside the Centre: the Life of J. Robert Oppenheimer. Jonathan Cape, London, UK, 2012. ISBN 0-224-06262-X (hardcover). xiv + 818 + 32 pp. LCCN QC16.O62 M66 2012.

Ohanian:1976:GS

[4389] Hans C. Ohanian. Gravitation and Spacetime. Norton, New York, NY, USA, 1976. ISBN 0-393-09198-8. xiv + 461 pp. LCCN QC178 .O35.

Ohanian:1994:GS

[4390] Hans C. Ohanian and Remo Ruffini. Gravitation and Spacetime. Norton, New York, NY, USA, second edition, 1994. ISBN 0-393-96501-5. xv + 679 pp. LCCN QC178 .O35 1994.

Ohanian:2013:GS

[4391] Hans C. Ohanian and Remo Ruffini. Gravitation and Spacetime. Cambridge University Press, Cambridge, UK, third edition, 2013. ISBN 1-107-01294-5 (hardcover). xvi + 528 pp. LCCN QC178 .O35 2013.

OMeara:1997:DSC

[4392] Stephen James O’Meara. Deep-Sky Companions — The Messier Objects. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-55332-6, 0-933346-85-9. xiv + 304 pp. LCCN QB65.O44 1997. Foreword by David H. Levy.

OMeara:2014:DSC

[4393] Stephen James O’Meara. Deep-Sky Companions. Cambridge University Press, Cambridge, UK, second edition, 2014. ISBN 1-107-78562-6, 1-107-01837-4 (hardcover). xiv + 409 pp. LCCN QB65 .O44 2014.

Purrington:2018:HAC

[4394] Robert D. Purrington. The Heroic Age: the Creation of Quantum Mechanics, 1925–1940. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-065517-8 (hardcover). xv + 403 pp. LCCN QC173.98 .P87 2018.

Rau:2014:BPPa

[4395] A. R. P. (A. Ravi P.) Rau. The beauty of physics: patterns, principles, and perspectives. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-870991-9 (hardcover), 0-19-178018-9 (e-book). xii + 219 pp. LCCN QC21.3 .R38 2014.

Raymer:2017:QPW

[4396] Michael G. Raymer. Quantum Physics: What Everyone Needs to Know? What Everyone Needs to Know. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-025071-2 (paperback), 0-19-025072-0 (hardcover), 0-19-025074-7 (e-book). xxii + 306 pp. LCCN QC174.12 .R387 2017.

Resag:2018:FHP

[4397] Jörg Resag. Feynman and His Physics: The Life and Science of an Extraordinary Man. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-96835-1 (print), 3-319-96836-X (e-book), 3-319-96837-8. ISSN 2365-0613 (print), 2365-0621 (electronic). xii + 319 pp. LCCN QC16.F49 R4713 2018.

Schrijver:2015:LSH

[4398] Karel Schrijver and Iris Schrijver. Living with the stars: how the human body is connected to the life cycles of the Earth, the planets, and the stars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2015. ISBN 0-19-872743-7 (hardcover). ix + 202 pp. LCCN ????

Schrijver:2018:OTB

[4399] Karel Schrijver. One of ten billion earths: how we learn about our planet’s past and future from distant exoplanets. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186486-2, 0-19-879989-6 (hardcover), 0-19-256021-2 (e-book). xi + 460 + 16 pp. LCCN QB820 .S357 2018.

Schwartz:2017:LMW

[4400] David N. Schwartz. The Last Man Who Knew Everything: the Life and Times of Enrico Fermi, Father of the Nuclear Age. Basic Books, New York, NY, USA, 2017. ISBN 0-465-07292-5 (hardcover), 0-465-09312-4 (e-book). xxiii + 453 pp. LCCN QC16.F46 S39 2017.

Stuewer:2018:AIN

[4401] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Svistunov:2015:SSM

[4402] Boris Svistunov, Egor Babaev, and Nikolay Prokofiev. Superfluid States of Matter. Taylor and Francis Group, Boca Raton, FL, 2015. ISBN 1-4398-0275-0 (hardcover). xxi + 561 pp. LCCN QC175.4 .S85 2015.

Thorne:2014:SI

[4403] Kip S. Thorne. The Science of Interstellar. W. W. Norton and Company, New York, NY, USA, 2014. ISBN 0-393-35137-8 (paperback). x + 324 pp. LCCN PN1997.2.I635 T49 2014; PN1995.9.S26 T49 2014. Foreword by Christopher Nolan.

Thorpe:2006:OTI

[4404] Charles Thorpe. Oppenheimer: the Tragic Intellect. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-79845-3 (hardcover), 0-226-79846-1 (paperback). xx + 413 + 12 pp. LCCN QC16.O62 T56 2006. URL http://www.loc.

gov/catdir/enhancements/fy0666/2006015223-b.html;   http://
www.loc.gov/catdir/enhancements/fy0666/2006015223-d.html;
http://www.loc.gov/catdir/enhancements/fy0666/2006015223-t.
html.

Tripon-Canseliet:2014:NME

[4405] Charlotte Tripon-Canseliet and Jean Chazelas. Nanoscale microwave engineering: optical control of nanodevices. FOCUS: Nanoscience and Nanotechnology Series, FOCUS Series. ISTE, London, UK, 2014. ISBN 1-84821-587-8, 1-118-92539-4 (e-book). ISSN 2051-249X. xiii + 120 pp. LCCN TK7876 .T757 2014.

Unser:2014:ISS

[4406] Michael A. Unser and Pouya Tafti. An Introduction to Sparse Stochastic Processes. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-05854-6 (hardcover). xviii + 367 pp. LCCN QA274.23 .U57 2014.

Vigdor:2018:SAV

[4407] Steven E. Vigdor. Signatures of the Artist: the Vital Imperfections That Make Our Universe Habitable. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-185295-3, 0-19-881482-8 (hardcover), 0-19-254676-7 (e-book). xv + 345 pp. LCCN QB461.3 .V54 2018.

Wasserman:2018:PTT

[4408] Ryan Wasserman. Paradoxes of Time Travel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-879333-2 (hardcover), 0-19-184198-6 (e-book), 0-19-251181-5 (e-book). x + 250 pp. LCCN QC173.59.S65 W37 2018; BD632 .W38 2018.

Weaire:1999:PF

[4409] Denis Weaire and Stefan Hutzler. The Physics of Foams. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-850551-5, 0-19-851097-7. xiii + 246 pp. LCCN QD549 .W254 1999.

Wickramasinghe:2015:SOC

[4410] Chandra Wickramasinghe. The Search for Our Cosmic Ancestry. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2015. ISBN 981-4616-96-6 (hardcover), 981-4616-97-4 (softcover). xx + 196 pp. LCCN QH326 .W534 2015.

Woodward:2013:MSS

[4411] James F. (James Francis) Woodward. Making Starships and Stargates: the Science of Interstellar Transport and Absurdly Benign Wormholes. Springer-Praxis books in space exploration. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-5622-3 (paperback). xxvi + 279 pp. LCCN TL790 .W66 2013.

Woolfson:2007:FSS

[4412] M. M. Woolfson. The formation of the solar system: theories old and new. Imperial College Press, London, UK, 2007. ISBN 1-86094-824-3 (hardcover). xxii + 318 pp. LCCN QB503 .W659 2007.

Zewail:2014:VMR

[4413] Ahmed H. Zewail. 4D visualization of matter: recent collected works. Imperial College Press, London, UK, 2014. ISBN 1-78326-505-1 (paperback), 1-78326-504-3, 1-78326-506-X (e-book). xvi + 409 pp. LCCN QC173.59.S65 Z45 2014.

Brink:2014:NLP

[4414] Lars Brink, editor. Nobel Lectures, Physics, 2006–2010. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2014. ISBN 981-4612-67-7 (hardcover), 981-4612-68-5 (paperback). xii + 371 pp. LCCN QC71 .N63 2014.

Cherry:2015:EVB

[4415] Simon R. Cherry, Ramsey D. Badawi, and Jinyi Qi, editors. Essentials of In Vivo Biomedical Imaging. CRC Press/Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4398-9874-X (hardcover). xvi + 272 pp. LCCN RC78.7.D53 E88 2015.

Davies:2014:INR

[4416] P. C. W. Davies and Niels Henrik Gregersen, editors. Information and the Nature of Reality: From Physics to Metaphysics. Cambridge University Press, Cambridge, UK, canto classics edition, 2014. ISBN 1-107-68453-6 (paperback). xvii + 487 pp. LCCN BD701 .I49 2014.

Dine:2012:STA

[4417] Michael Dine, Tom Banks, and Subir Sachdev, editors. String theory and its applications: TASI 2010, from meV to the Planck scale: Proceedings of the 2010 Theoretical Advanced Study Institute in Elementary Particle Physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2012. ISBN 981-4350-51-6 (hardcover). LCCN QC793 .T44 2010.

Enz:2002:NTB

[4418] Charles P. (Charles Paul) Enz. No Time to be Brief: a Scientific Biography of Wolfgang Pauli. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-856479-1. viii + 573 pp. LCCN QC16.P37 E59 2002. URL http://www.loc.gov/

catdir/enhancements/fy0614/2002726902-d.html;   http://www.
loc.gov/catdir/enhancements/fy0614/2002726902-t.html.

Gan:2014:HMO

[4419] Fuxi Gan and Shouyun Tian, editors. History of Modern Optics and Optoelectronics Development in China, volume 1 of Series on archaeology and history of science in China. World Scientific/World Century, Hackensack, NJ, USA, 2014. ISBN 981-4518-75-1 (hardcover). xx + 372 pp. LCCN QC352 .H57 2014.

Jones:2004:IGC

[4420] Mark H. (Mark Henry) Jones, Robert J. Lambourne, D. J. (David John) Adams, et al., editors. An Introduction to Galaxies and Cosmology. Open University, Milton Keynes, UK, 2004. ISBN 0-521-83738-3, 0-521-54623-0 (paperback). vi + 442 pp. LCCN QB857 .I67 2004. URL http://www.

loc.gov/catdir/description/cam051/2004301655.html;   http:/
/www.loc.gov/catdir/toc/cam051/2004301655.html.

Jones:2015:IGC

[4421] Mark H. (Mark Henry) Jones, Robert J. Lambourne, and Stephen Serjeant, editors. An Introduction to Galaxies and Cosmology. Cambridge University Press, Cambridge, UK, second edition, 2015. ISBN 1-107-49261-0 (paperback). xviii + 432 pp. LCCN QB857 .I67 2015.